A Complete Bibliography of Publications in Historia Mathematica

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

10 December 2024 Version 1.52

Title word cross-reference

#8 [Bar74].

$10.50 [Tan74, Eis75]. $11.95 [McK82]. $116 [Bee06]. $12.50 [Zas83b, Ber74, Bla76, Cib76, Dav75, Hir74, Wil75a]. $12.80 [Bie74a]. $124.70 [Swe79]. $13 [Enr85b, Enr85a]. $13.20 [Hod75]. $13.25 [Bar74]. $14 [Ang88]. $14.00 [Bay11, Hod75]. $14.50 [Ano74n, Ano82f]. $14.95 [Ano85h]. $140.00 [Ano95f]. $149.50 [Gra94b]. $15 [Tro74, Wis76]. $15.00 [Ano84i, Pye75, Yok75]. $16.50 [Jon77c]. $17.50 [Bir77, Tat75]. $18.00 [Ano86e, Zas83b]. $18.50 [Vek75]. $19.50 [vM84]. $19.95 [Ano86d, Lev97]. $199 [McL12]. $2.50 [Ano82f]. 2/n [Abd08, Gil78]. $20 [Ano74o, Ano74-28, Gin83, Haw77b]. $20.80 [Das81]. 20 × 26 [Maz21]. $215 [Sch13b]. $22 [Enr85b, Enr85a]. $22.50 [Ano81l]. $229.50 [Gra06]. $23.00 [Ano89e]. $23.50 [Enr85a]. $23.90 [Bun76a]. $24.95 [Ano87e, Kla84]. $25 [Ano74y, Yon75]. $25.00 [Ano87k, Ano89c, Hym78]. $26.50 [Enr80]. $27.50 [Ano88e, Haw81]. $27.85 [Har77a]. $27.95 [Dav90, Fer04b]. $28 [Haw77a].

$29.50 [Ano77s, Tro83]. $29.50/Dfl [Ano84g]. $29.95 [BT84]. 2n [BS77b]. $3.00 [Sai74]. $3.50 [Ped75b]. $3.95 [LY79]. 3/2 [vBS20]. $30 [Ano88m]. $30.00 [Ano84i]. $30.95 [Bru77]. $300 [Moo03]. $32.50 [Hel83]. $34.00 [Ang88, Syl83]. $34.95 [Lew90]. $35.00 [GG86, Lei79, Lew95]. $36.50 [Ano84h]. $36.95 [Ale04]. $37.50 [Ano91e]. $38.50 [Ano74z]. $39 [Haw03]. $39.50 [Ewi92, vM84]. $39.95 [Dav90, Kin86b, Wal17]. $4.95 [Ano74n]. $40.00 [Row90a]. $41.50 [Ano87f]. $42.00 [Ano84j]. $44.00 [How18, Jon85]. $45.00 [Ano89l, Ung85]. $47.50 [Bow84]. $49 [Haw03, vdW76a]. $49.50 [Ano88g, Ric82]. $49.80 [Wol91]. $49.95 [Ano92n, Shu80]. $5.00 [Gra77]. $5.25 [Ano81l]. $5.95 [Wil75b]. $50.00 [Ano88i, Ano89d]. $51.95 [Ayo06]. \frac{51}{2} [Mia08]. $54.50 [Sag91]. $55.00 [Ano88f, Per10]. $56 [Hog03]. $58.00 [Ano91f]. $59 [San03]. $6.25 [Ano74u]. $60 [Ano77m]. $60.00 [Ano92o, Ano93f]. $62.50 [Coh82]. $65.00 [Ano88f]. $7.95 [Rus75]. $70 [Ano88k]. $72.00 [Hog85b]. $76.00 [Gra18]. $79.00 [Gra94a]. $8.00 [Bid74]. $8.50 [Ano74r]. $8.95 [Ano87k]. $80.00 [Rid88]. $83.95 [Buc77]. $9 [Ano74x]. $9.95 [Ano82d, Par89]. $95.00 [Neu86]. $98 [Nor03]. $98.00 [Ano87g, Ano93g]. $99.00 [Fra10]. ^{c} [Van98]. ^{∘} [Gar98]. ^{e} [Ano79l, Ano90-34, Ano92-50, Ano92-27, BH14, Dje97, Gol82, GG76a, Gra01, Hog01a, Kit81, Pei17, Sim96, Tat86]. ^{ème} [DP88, Kno88a]. ^{o} [Ano77j]. ^{⊃} [Hog86]. ^{th} [Rab24, Zhu20]. Aρíμητικη [Chr91]. · [Dje98]. χ^{2} [Pie83]. † [Dje98]. e [Fol14, RP19]. ≫ [Høy07, Bra09b]. L [Gra18]. ≪ [Bra09b, Høy07]. n [DR03]. ω [GA20]. OR → CJ [Ano74l]. p [Ano79w, Joh84]. π [CS81, TG98, Vol97, War15]. ρ [Ano92a]. Στoιχϵιωσις [Chr91]. \sqrt{2} [Buc23]. V [Kle17]. y [Aus16].

-/U.S [Bow84]. -adic [Ano79w]. -approximation [vBS20]. -Functions [Gra18]. -ten [Joh84]. -th [Joh84].

. [Gra10b].

/Atlantic [Ano77s]. /U.S [Bow84].

0 [Abe06, Abe09, AH16, And07, Bel07a, Bel08, Bor14, Bra07, Bra09b, Cal08, Con06, Dru07, Eva13, Gis14, Gra13, GG05c, Gra18, Hun07, Kat07, Lan07, Lor16, Maa07, Maz21, Mel06, Nor07, Par07, Plo06a, Rec07, Rid88, Row23, San05, Sch07, Sch05, Sen06, SS07, Ste05, Thi06, Tur08, Wil05, Wil07, vM07]. 0-19-508030-0 [GG05c]. 0-19-852309-2 [Ste05]. 0-19-852495-1 [Bee06]. 0-19-853930-4 [Gra06]. 0-19-856484-8 [vM07]. 0-19-856601-8 [Maa07]. 0-19-856737-5 [Nor07]. 0-262-16216-4 [Con06]. 0-387-25284-3 [Bra07]. 0-387-95303-5 [Mel06]. 0-444-50871-6 [SS07]. 0-521-48124-4 [Sch05]. 0-674-00979-7 [Sen06]. 0-674-32449-8 [Fer04b]. 0-691-05069-4 [Wil05]. 0-691-09259-1 [Plo06a]. 0-691-11445-5 [Hun07]. 0-691-11822-1 [Lan07]. 0-8018-7109-3 [San05]. 0-8018-7434-3 [Abe06, Lor16]. 0-8018-8011-4 [Wil07]. 0-8018-8235-4 [Bel07a]. 0-8176-4040-1 [Wol03]. 0-8218-3776-1 [Rec07]. 0-8218-3788-5 [Sch07]. 0-8218-4376-1 [AH10]. 0-8223-3814-9 [Bel08]. 0-8240-9284-8 [Rid88]. 0-88385-178-4 [Kat07]. 0-88385-546-1 [Ayo06]. 0-88385-547-X [Thi06]. 0-88385-549-6 [And07]. 0-88385-550-X [Par07]. 0-88385-559-3 [Cal08]. 0-913510-24-6 [McK82]. 0-930326-14-8 [Dru07]. 0-9887449-2-9 [AH16]. 0.50 [Gri76]. 000 [Ano85f, Isr85, Ken82, Ysh90]. 0221-3664 [Mal97]. 08 [Tob06].

1 [All09, Atz06, Bru07, Cog15, Fra06, Gol13, Gon08, Gra08, How18, Høy09a, Høy23, Jam13, Jon77b, Lan07, Nau03, Plo06a, Rec07, Sau06a, Ser18, Sid06a, Sid06b, Wat09, Wol03]. 1-4020-0081-2 [Sid06a]. 1-4020-2697-8 [Atz06]. 1-4020-2777-X [Sau06a]. 1-86232-211-2 [San06]. 1-905122-09-8 [For10]. 1.00 [BT81, Jon76a]. 10.00 [Ano74s]. 10.4324 [Gur24]. 10.4324/9780429506680 [Gur24]. 100 [Row85a]. 100th [Ano92-31]. 102 [Don05]. 10th [Mat74]. 11th [Hog86, Hog95, Mat74, Ai04]. 11th-century [Hog86]. 12 [Pur86a, Lév99]. 12.80 [Lem74]. 121 [Sal78]. 1283 [Moz13]. 12th [Fol15]. 12th-century [Fol15]. 1316 [Ano74q]. 1350 [Maz77]. 139.99$ [DR19]. 144p [Tho76]. 147c7 [Høy90c]. 148d7 [Høy90c]. 14th [Ano93a, Fol83a, Fra88, Rab74]. 1500th [Gup76a, Gup77a]. 150th [Ano77a]. 1551 [Sch16]. 1551-1840 [Sch16]. 1593 [Bee09]. 1593-X [Bee09]. 15th [Hug87, Van24]. 15th-century [Hug87, Van24]. 16 [Ano87-27]. 160 [Gol76]. 1600 [Swe00]. 1632 [Wis76]. 1665 [Ano74-28]. 1666 [Bru07]. 1673-1676 [Cri17]. 1676 [Bar78b]. 1685 [Bee06]. 1687 [Nag88]. 1687/1987 [Nag88]. 1693 [Gar05]. 16th [Bir78, Gol00, Mes94, iÔ75, Rey06]. 16th- [Gol00, Mes94]. 16th-Century [Rey06]. 17 [SY90]. 1700 [Per10]. 1706-1749 [Pei15]. 17264 [Cav85]. 1749 [Fra04a]. 1750 [Kil83, Sen91]. 1752 [Kle17]. 1766 [Müü75]. 176pp [GG80]. 17e [Ano92-28]. 17S [Aup76]. 17th [Cal83, Cas80, Che10, Gol00, Lüt20, Mah84, Mes94, SD08]. 17th- [Che10, Gol00]. 17th-Century [Mes94]. 18.50 [vM84]. 1800 [Sag91]. 1803 [Müü77]. 1807 [Her75a]. 1807/08 [Bie74b]. 1814 [Hal78]. 1820s [Fer08, San09]. 1826 [Phi79a]. 1830 [Boi89]. 1830/1885 [Sch06b]. 184 [Ano81i]. 1840 [Par92a, Sch16]. 1844 [Law05]. 1847 [Kno88a]. 1856 [GG13]. 1860s [Ehr18]. 1864 [Jon85]. 1868-1951 [Sau14]. 1871 [GG86]. 1875 [Dum18]. 1884 [Ano89k]. 1885 [Bel08]. 1891 [Ano82l]. 18e [Ano92-28]. 18th [Ano91-31, Beb96, BF23, Bla14, Cas80, Dom04, Fin11, Ken76d, Per80, Scr88a, Böt12, BS13, Bul09b, Che10, Gol00]. 18th-century [Bla14, Fin11, Bul09b, Che10, Gol00]. 1900 [BKM^{+}82, Sch91b, Sch91c]. 1902 [Row07]. 1904 [Ebb07]. 1907 [Tob06, Row23]. 1907/08 [SS08, Tob06, Tob06]. 1908.By [Gar98]. 1910 [Ano88e]. 1918 [Sel90]. 192 [Ter89]. 1920 [Cal75]. 1920s [SS93]. 1925 [Dum15, Dur21]. 1927 [Has10a]. 1930s [Sen04]. 1931 [Fer04b, Moo77]. 1932 [Ano84f]. 1933 [Cas80, Daw16b, Seg90]. 1935 [Krö06, Row83a]. 1936 [Har77a]. 1938 [Shu80]. 1940 [SS07]. 1944 [Tob06, SS08]. 1944/45 [Tob06, SS08]. 195-379 [Str82a]. 1950 [Gra96]. 1951 [Jur07]. 1960 [Gra91a]. 1964 [Ano92m]. 1970 [Ano79k]. 1972 [Bos77]. 1972/4 [Hod75]. 1973 [Ano74-51, Ano77f, Dem74, Pye74]. 1973/1974 [Ano74-37]. 1974 [Ano74-48, Ano74-76, Fre76, Rog74, Tot79, GG90b, Hei81]. 1975 [Ano75k, Ano75p, Ano76e, Ano76v, Ano76x, Ano76-28, Ano77p, Ano81m]. 1976 [Ano76h, Ano76i, Ano76u, Ano76w, Ano76-29, Ano76-34, Ano77d, Ano79l, Kit78, Rog77, Ano81s]. 1977 [Ano77-29, Ano77-30, Bie78, Kot79, Obe78, dBR76]. 1978 [Ano78t, Ano78y, Buc80, Edw80, Str82a, vdW80, Ano81e]. 1978-3 [Str81]. 1978-3/4 [Str81]. 1979 [Ano79-31, Vol81]. 198 [Hes90]. 1980 [Ano80p]. 1981 [Ano82k, Ano84k, Bin83, Bir83, Cri82, Dho85, Scr82a]. 1982 [Con85]. 1983 [Ano83n, Pyc84a, Sen84]. 1984 [Ano83f, Ano85-28, Ano85-30, Arb86, dMC86a, Che87, Lo 85, Pur86b]. 1985 [Ano85n, Ano85p, Ano87-36, Dru85, Hou89, Pur86a, Scr84, Fre92]. 1986 [Ano86m, Ano87m, Bin90, Bow87, Gra87, Høy90b, Mai87, Shi88]. 1987 [Ano86s, Ano87c, Ano87n, Ano87-34, Ano87-27, Ano88-31, Ano88r, BDD^{+}88, GG87b, Mic88, Nag88, Ott88, Pur88, Rod88]. 1988 [Ano87l, Ano87-31, Ano88o, Ano88v, Ano89-34, Ano89t, Ano90f, BJ89a, Dho88b, Pur90, Roe89, Row89, Seg90, Sel91]. 1989 [Ano88u, Ano89m, Ano89o, Ano89-35, Ano89-30, Ano90c, Ano90g, Ano90k, Ano90m, Ano90q, Ano90-32, Ano90-36, Bin91, Boi89, Dho90, Gia90, GG90d, Pep90, Ric91, Scr87, Str94]. 1990 [Abe91, Ano89r, Ano90i, Ano90j, Ano90l, Ano90s, Ano90a, Ano90t, Ano90-35, Ano91c, Ano91k, Ano91n, Ano91o, Ano91p, Ano91q, Ano91u, Ano92-27, Ano92-49, Art91, Bre91, Mue96, Nag91b, PP91, Pur91, Toe96]. 1991 [Ano90e, Ano90u, Ano90-33, Ano91h, Ano91s, Ano91t, Ano92h, Ano92f, Ano92s, Ano92v, Ano92-44, BLF92, GG92b, Mil93, Pep92, Swe92]. 1992 [Ano91j, Ano91-33, Ano92g, Ano92e, Ano92q, Ano92-50, Ano92u, Ano92a, Ano92x, Ano92-28, Ano92-31, Ano92-30, Ano92-34, Ano92-35, Ano92-47, Ano92-48, Ano93k, Ano93m, Ano93x, Bir97, Vol93a]. 1993 [Ano92w, Ano93j, Ano93s, Ano93w, Ano93a, Ano94g, Ano94b, Coo96, Koe94, Ric98, Sim96, Vog94, Lew99]. 1994 [Ano74-46, Hoc96, SS97a, Vil96]. 1995 [Gra98a, Lew98]. 1999 [Atz04]. 19th [Ano80-31, Ano86h, Ano87-29, Ano88a, Ano91-31, Bir78, Cal83, Des09, Fri21, Gra94b, Gra94c, Gra97a, GN92, Hog90, Ken76d, Mar20, Pyc84b, Taz01, Tur81, Wan22, Cra99, Cra00, Cra07, Cri99, Fol76, Ine94, Kë96, Sar08]. 19th-century [Hog90, Pyc84b, Cra99, Cra00, Cri99]. 1pp [Ano11f]. 1r [Sel74].

2 [Ano83k, Bre07, Bru07, DR19, GG08, Gra08, Har06, Kle17, Kno07, Lon21, Lüt15, Mar14, Maw06, Moo75a, Plo06b, Sid06a, Ste05]. 2-02-153070-1 [Maz24]. 2-271-06128-8 [Har06]. 2-7116-1787-4 [Kno07]. 2-7117-5309-3 [Maw06]. 2-7302-1138-1 [Gra08]. 2-84174-275-X [Krö06]. 2-85367-221-2 [Bre07]. 2-85367-230-1 [Bru07]. 2-volume [Enr85a]. 2.00 [GG90a]. 2.5.1987 [Pur88]. 2.75 [Kil83]. 20 [Pur90, Tai14]. 200 [Ano77g]. 2002 [BGG06a, Rom04]. 2003 [MP06]. 2005 [Syl07]. 2009 [Dau12, Jah12, Kat11].

200th [Ano90e, Fel83]. 2013 [Aus13, Fra13]. 2017 [Sør17]. 2018 [Pei22]. 2021 [Ano21a, Ano21j, Ano21m, Ano21k, Ano21l, BG21a]. 2022 [Ano22j, Ano22i, Ano22k, Bul24]. 2023 [Ano23i, Ano23j, Ano23k, Ano23h, Lê24, Sid24]. 2024 [Ano24h, Ano24f, Ano24g, Ano24i]. 20th [BC04, Bro75, Dau93a, GN92, Nik18, RW03, Hun96]. 20th-Century [Hun96]. 21 [Ano91-30]. 211 [Syl03]. 22 [Sch76a]. 24 [GG98, Gra98b]. 24.50 [Pye76]. 247 [Eva94]. 25 [Gab76, San06]. 250th [Ano78a]. 26 [Hug05]. 27 [Dho90, Pur88]. 28. [Pur88]. 28th [Ano85d, Pur88]. 29 [Del03a, GG03, Krö06]. 29. [Pur90]. 29th [Pur90]. 2nd [Ano88a, Ano94a, Gra08, Seg90]. 2pp [Ano09h, Ano10g, Ano10h, Ano10i, Ano10j, Ano11g, Ano11h, Ano11i, Ano12f, Ano12g, Ano12h].

3 [Abe06, Ber07, Bos03, Bra07, Cal08, Fen08, Fra06, Ges18, GG07, Hei09, Høy09a, Lor16, Lüt09a, Maw06, Mos16, Müü75, Pan08, Pan18, Pec20, Plo06b, Row06a, San05, Scr03, Scr07, Scr08, SS08, Wol03, Sch90b]. 3-05-003471-82003 [Gar05]. 3-05-004003-3 [Bla05]. 3-406-47427-6 [Don05]. 3-447-04644-9 [BB04]. 3-515-08223-9 [Van06]. 3-515-08620-X [GG07]. 3-515-08713-3 [Ber07]. 3-534-15894-6 [Lüt09a]. 3-540-20836-4 [Scr07]. 3-540-42224-2 [Plo06b]. 3-7 [Vol91]. 3-7001-2984-X [Scr03]. 3-7643-1476-1 [Fra06]. 3-7643-4040-1 [Wol03]. 3-7643-5048-2 [Gra97a]. 3-7643-7538-8 [Scr08]. 3-7705-4126-X [Hei09]. 3-8265-4804-3 [Bos03]. 3-928186-41-8 [Has04]. 3-930430-68-1 [Høy09a]. 3-936905-07-X [Dho11]. 3-936905-21-5 [SS08]. 3-938016-04-3 [Row06a]. 3-Mannigfaltigkeiten [Krö06]. 3.75 [Tan74]. 3/4 [Str81]. 30 [Moz13, Ano77g]. 30. [Pur91]. 30/U.S [Ano84i]. 308 [Gra94c]. 30th [Pur91, Ano92b]. 319 [Mac76]. 322 [MW17, Rob01]. 3300 [Sid12]. 34/35 [Str81]. 348 [Sai94]. 349 [Rid94]. 35 [Goe09]. 36 [Gui09a]. 37 [Has10b]. 38 [Ano82e]. 395r [Ray23]. 398 [Par92a]. 3rd [Ano86a, vM84, Ano90a].

4 [Arc09, Bel07a, Con06, Kat07, KS13a, Kno07, Lor18, Mar09, Maz24, Müü77, Pei15, Rei15, Sau09, Sch16, Sch05, Scr07, SG09, Wil05, Wil07, Rid94].

4. [Ano94b, BLF92]. 40 [Has04]. 40th [Ano76x]. 43.30 [Ano77o]. 45 [Tob06, SS08]. 450th [Ano90j]. 46 [Atz04]. 460 [GW94b]. 47 [CM19a]. 472 [BKM^{+}82]. 479 [Cri24a]. 48 [Gra90, Grö86]. 4th [Ano81i, CM19a, CM19b, Fre89]. 4th-7th [CM19a].

5 [Ale14, Høy09b, Hun07, Jur07, Mel06, Nor07, Pan09, San09, Sch07, SS08, Thi08, Tim09, Zit07a]. 50 [Row85b]. 54a [AS93]. 55 [Row93]. 56 [Ano89f]. 568 [Gon08]. 5è [Ano91q]. 5th [Jay24a, Jay24b]. 5th-century [Jay24a, Jay24b].

6 [And07, Cap13, Dum15, Hug07, Krö15, Lei09, Lew09, Lüt09a, Mal06a, Mro16, Oli08, Pur91, SS07, SS17, Wil08a]. 6.5-12.5 [Pur91]. 60 [Fra04a, Taz09, Yan83]. 60/U.S [Ano84i]. 6000 [Ken78b]. 606 [Eli24]. 60th [Ano79-42, Ano81-35, GY77]. 62 [Gau13]. 64.00 [Tat11]. 65.00 [Ano84g]. 6500 [Ano81f]. 65f. [Ano77n]. 65th [Ano95f, Bur75]. 68 [Van06]. 686r [Ray23]. 686r-686v [Ray23]. 686v [Ray23].

7 [Bul16a, Dav19, Fol14, Gil15, Krö17, Lüt08, Net24, Neu08, Nic13, Sau14, Sch17, Sch13c, Sen06, Ste09]. 70th [ADG^{+}88, Ano89s, DPY91, Mar10b]. 721/1321 [Dje98]. 7289 [Buc23, FR98]. 75 [Ano74p]. 75.90 [Dra84]. 75pp [LY79]. 75th [Fol84]. 78 [Ano88n]. 7th [Ano92l, CM19a, CM19b, lLW13].

8 [Atz06, Aus16, Bor18, Bul22, Dru07, Duc16, Fol15, Gil81, GG13, Har06, Maa07, Pei17, Rid88, Sch09, Scr08, Taz15, vM07]. 8000 [Van80]. 80k [Sel74]. 80th [You83]. 817th [Dru85]. 84 [Pan08]. 84-9769-173-3 [Pan08]. 842 [Nak03]. 85 [Lei09]. 85-7496-061-6 [Lei09]. 85.00 [Gra10a]. 85th [Ano91a]. 87 [Sid06b]. 87-7289-815-1 [Sid06b]. 88 [Hug07, Mal06a, Rab08]. 88-222-5529-1 [Taz09]. 88-467-1043-6 [Mal06a]. 88-467-1351-6 [Hug07]. 88-548-0626-9 [Rab08]. 88-9007876-1-7 [GG05a]. 88-9007876-2-5 [GG05a]. 88-9007876-7-6 [GG05a]. 8th [BLF92].

9 [AH16, Bel08, De 20, Des09, Her75a, Krö14, MF10, Rab08, Row07, Row15, Sai05, Sid13]. 90 [Mue80, Sai05]. 90-04-11363 [Hog03]. 90-04-11977-9 [Sai05]. 90th [Fre93]. 949 [War15]. 95Ffr [Ano92k]. 978 [Abe09, Ale14, All09, Arc09, Aus16, Bor14, Bor18, Bra09b, Bul16a, Bul22, Cap13, Cog15, Dav19, De 20, Des09, Duc16, Dum15, DR19, Eva13, Fen08, Fol14, Fol15, Ges18, Gil15, Gis14, Gol13, Gra13, GG08, GG13, Gra18, How18, Høy09b, Jam13, Jur07, KS13a, Kle17, Krö14, Krö15, Krö17, Lew09, Lon21, Lor18, Lüt08, Lüt15, Mar14, Mar09, Maz21, Maz24, Mos16, Mro16, Net24, Neu08, Nic13, Oli08, Pan09, Pan18, Pec20, Pei15, Pei17, Rei15, Row07, Row15, Row23, San09, Sau09, Sau14, Sch16, Sch17, Sch09, Sch13c, Ser18, SG09, Sid13, SS17, Ste09, Taz15, Thi08, Tim09, Tur08, Wat09, Wil08a, Zit07a]. 978-0-19-856793-6 [Oli08]. 978-0-19-856947-3 [Mos16]. 978-0-19-870305-1 [How18]. 978-0-19-921312-2 [Gra10a]. 978-0-19-968961-3 [Pec20]. 978-0-19-968962-0 [Pec20]. 978-0-19-995104-8 [Duc16]. 978-0-387-72930-5 [Høy09b]. 978-0-387-73467-5 [San09]. 978-0-387-74640-1 [Wat09]. 978-0-387-84825-9 [Sid13]. 978-0-387-87856-0 [Bel12]. 978-0-387-87868-3 [Cog12]. 978-0-393-06027-0 [Abe09]. 978-0-521-75994-6 [Bay11]. 978-0-674-04661-0 [Gra11]. 978-0-691-11880-2 [Fra10]. 978-0-691-12973-0 [Eva13]. 978-0-691-14014-8 [Hun12]. 978-0-691-14041-4 [Sen10]. 978-0-691-15478-7 [Bul16a]. 978-0-691-17958-2 [Lon21]. 978-0-7083-2526-1 [Jam13]. 978-0-8018-8814-4 [SG09]. 978-0-8018-8823-6 [Per10]. 978-0-8018-9755-9 [Hud12]. 978-0-8176-4684-4 [Mar09]. 978-0-8218-3457-2 [D’A10].

978-0-8218-3860-0 [Tur08]. 978-0-8218-4172-3 [Fen08]. 978-0-8218-4343-7 [Neu08]. 978-0-8218-6914-7 [Nic13]. 978-0-85729-537-8 [Ser18]. 978-0-85729-928-4 [KS13a]. 978-0-88385-182-1 [All09]. 978-0-88385-564-5 [Thi08]. 978-0-88385-565-2 [GG08]. 978-0-930326-25-8 [Dru12]. 978-0-934223-93-5 [Tim09]. 978-1-107-01684-2 [Sch13b]. 978-1-107-59464-7 [Dav19]. 978-1-4471-2691-1 [Ser18]. 978-1-4704-1076-6 [Mro16]. 978-1-4704-3584-4 [Lor18]. 978-1-4825-5168-6 [Dum15]. 978-1-56881-259-5 [Zit07a]. 978-1-59691-522-0 [Gra13]. 978-1-84564-092-7 [Ste09]. 978-1-84628-790-9 [Des09]. 978-2-02-153070-4 [Maz24]. 978-2-200-92598-7 [Gau13]. 978-2-251-22001-7 [Fol14]. 978-2-271-08749-2 [Kle17]. 978-2-296-56553-1 [Gol13]. 978-2-503-54129-7 [Gil15]. 978-2-503-58638-0 [Maz21]. 978-2-7056-8106-7 [Krö17]. 978-2-7056-8697-0 [Gis14]. 978-2-7302-1584-8 [Taz15]. 978-2-7351-1414-6 [Krö15]. 978-2-7453-1911-1 [Bré12]. 978-2-8143-0032-3 [Gra12]. 978-2-8143-0125-2 [Lüt15]. 978-2-8143-0180-1 [Cog15]. 978-2-84516-354-6 [Mar10a]. 978-2-84559-115-8 [Pei17]. 978-2-84788-233-9 [Row15]. 978-2-85367-247-4 [Fer11]. 978-2-88049-241-0 [Bra09b]. 978-2-88074-498-4 [Rei15]. 978-2-906053-92-2 [Yus10]. 978-2-916097-26-8 [Fol12]. 978-2-9700636-5-0 [Bor14]. 978-3-0348-0251-2 [DR19]. 978-3-03719-059-3 [Tat11]. 978-3-03719-146-0 [Gra18]. 978-3-319-03939-8 [Fol15]. 978-3-319-30845-6 [SS17]. 978-3-347-05156-0 [Row23]. 978-3-487-13585-4 [Sch16]. 978-3-487-14534-1 [Her14]. 978-3-487-15685-9 [De 20]. 978-3-487-15822-8 [Bul22]. 978-3-515-09176-3 [Sch11]. 978-3-515-09296-8 [Lef10]. 978-3-515-09638-6 [KS12]. 978-3-515-13573-3 [Høy24b]. 978-3-515-13581-8 [Høy24b]. 978-3-540-64373-9 [Row07]. 978-3-642-14868-2 [Che14b]. 978-3-642-17853-5 [Ale14]. 978-3-642-21824-8 [Bor18]. 978-3-642-21825-5 [Bor18]. 978-3-642-25803-9 [Krö14]. 978-3-642-32486-4 [Pei15]. 978-3-662-44149-7 [Sch17]. 978-3-7643-6555-4 [Arc09]. 978-3-7643-7136-4 [Sau09]. 978-3-7643-7257-6 [Lew09]. 978-3-7643-7299-6 [Wil08a]. 978-3-7643-8637-5 [Bra09a]. 978-3-7643-8859-1 [GG10]. 978-3-8353-1206-7 [Sau14]. 978-3-926269-34-8 [Sch09]. 978-3-928186-44-5 [Jur07]. 978-3-937219-57-8 [GG13]. 978-4-320-01910-2 [Sid12]. 978-7643-7701-4 [Gra10b]. 978-81-7304-809-8 [Knu11]. 978-82-7099-442-7 [Lüt08]. 978-84-15271-88-8 [Aus16]. 978-88-08-52129-3 [Pan18]. 978-88-222-5990-5 [Coz12]. 978-88-222-6003-1 [Rem12]. 978-88-222-6008-6 [Cap13]. 978-88-3315-476-3 [Net24].

978-88-3315-478-7 [Net24]. 978-88-498-1904-5 [Pan09]. 978-88-498-2156-7 [Sch13c]. 978-88-498-2461-2 [Mar14]. 978-88-7642-404-5 [Gal12]. 978-88-7916-388-0 [Oak12]. 978-90-04-43656-5 [Hog24]. 978-90-481-2739-9 [Sch12]. 978-90-481-3541-7 [Zep12]. 978-90-481-8117-9 [McL12]. 978-972-9376-31-3 [Ges18]. 978-9979-793-99-1 [Kat11]. 9780429506680 [Gur24]. 98 [Gra97a]. 9th [Ai04, Ano76u, BLF92].

= [òP09]. āAryabhaòtāiya [BK06]. Āpastamba [Ano88m]. Āryabhata [Gup76a, Gup77a].

âge [Rei15]. âges [Mar90b].

A-G [Pec20]. A. [AR76, Ano81a, Ano91a, Bie80, Bur86, GG02, GG03, Gre88, Isr90, Moo76, Tat88, VHK16, Wei04, You83, Gra94b]. A.-C. [Gre88]. A.-T. [Tat88]. A.D [Wag85, Wag79]. Aḣmad [Ano92i, Ros92]. a.M [Høy90b]. A8 [Kit78]. Aaboe [Ste07b]. Aachen [Bos03, BLF92, Str81]. Aant [Tho76]. Abāu [Ai04]. Abacus [iÔ75, Fol83a]. Abaris [Gin83]. abbaci [Sch96, Han07, Spi00a, Spi00b]. Abbas [Rom04]. Abbott [Bay11]. Abdullah [Sal78]. Abe [Gra91b, Wol03]. Abel [Del03a, Del02, Del06, Nik09, Pes03, Sør05, Con06]. Abeles [Dru07, Dru12]. abendländischen [Høy90b]. Aberystwyth [Cho92]. Abgrall [Bre07]. Abhandlung [Fol83a]. Abhandlungen [Ano89j, GG07, Haf16, Mat75, Sau14, Ull03]. Abolghassem [Bag02]. Abraham [SS17, Ber84b]. Abramovich [Vog10]. Abrégé [BKM^{+}82]. abridgement [Hug10]. Abriss [Lem74]. abroad [Hog91]. absolu [Fre77]. absolute [Ano84f]. absoluten [Ano95h]. Abstract [Ano85a, Ano88b, Ano89a, LZ88, Mar09, Hol15, Kle07, PV22, Pro22]. Abstracted [Ano95q, Ano96c, Ano97c, Ano98f, Ano99a, Ano00a, Ano01a, Ano02d, Ano74-54, Ano75o, Ano76-27, Ano77-32, Ano79-28, Ano80o, Ano81-34, Ano82n, Ano83m, Ano84p, Ano85m, Ano86q, Ano87q, Ano87-33, Ano88t, Ano89q, Ano90p, Ano91m, Ano92z, Ano93r, Ano94l, Ano03m, Ano04k, Ano05m, Ano06i]. Abstracts [Ano74a, Ano74b, Ano74c, Ano74d, Ano75a, Ano75b, Ano75c, Ano76a, Ano76b, Ano76c, Ano76d, Ano77b, Ano77c, Ano78b, Ano79b, Ano79a, Ano81b, Ano81c, Ano82a, Ano84a, Ano85b, Ano86b, Ano87b, Ano87a, Ano88c, Ano89b, Ano91b, Ano92c, Ano92d, Ano93b, Ano93c, Ano93d, Ano93e, Ano94c, Ano94d, Ano94e, Ano94f, Ano95a, Ano95b, Ano95c, Ano95d, Ano96a, Ano97a, Ano98a, Ano98b, Ano98c, Ano98d, Ano02a, Ano02b, Ano02c, Ano03a, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c, Ano04d, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano06b, Ano06c, Ano06d, Ano10a, Ano11a, Ano24a, Dav91, DMP07a, DMP07b, DMP07c, DMP07d, DMP08a, DMP08b, DMP08c, DMP08d, Lew78a, Lew78b, Lew78c, Lew79a, Lew79b, Lew80a, Lew80b, Lew80c, Lew81a, Lew81b, Lew82a, Lew82b, Lew82c, Lew83a].

Abstracts [Lew83b, Lew83c, Lew83d, Lew84a, Lew84b, Lew85a, Lew85b, Lew86a, Lew86b, Lew87, Lew88, May77a, May77b, MMP09a, MMP09b, MMP09c, MMP09d, MMP10a, MMP10b, MMP11a, MMP11b, MMP11c, MMP12a, MMP12b, MMP12c, MMP12d, MMP13d, MMP13a, MMP13b, MMP13c, MMP14c, MMP14a, MMP14b, MMP15a, MMP15b, MMP15c, MMP15d, MMP16a, MMP16b, MMP16c, MMP16d, MMP17a, MMP17b, MMP17c, MMP17d, MMP18c, MMP18a, MMP18b, MMP19d, MMP19a, MMP19b, MMP19c, MMP20a, MMP20b, MMP20c, MMP20d, MMP21a, MMP21b, MMP21c, MMP21d, MP22a, MP22b, MP22c, MP22d, MP23a, MP23b, MP23c, MP23d, MP24a, MP24b, Van00a, Van00b, Van00c, Van00d, Van01a, Van01b, Van01c, Van01d, Van02, Zit90a, Zit89a, Zit89b, Zit90b, Zit90c, Zit90d, Zit91, Zit92a, Zit92b, Zit96a, Zit96b, Zit96c, Zit97a, Zit97b, Zit97c, Zit99a, Zit99b]. Abstracts [Zit99c, Zit99d]. abstraite [Ano91s]. Abu [Bre07, Mat74, Tee78, Ray12, Rom04]. Abu’l [Ken84]. Abu’l-wafa’ [Ken84]. abuse [Gra84a]. Academia [Ano94g, Bir74, Nos12]. Academic [Ano88g, Atz06, Bar74, Bla76, Ewi92, Høy95, Lew90, Pla76, Sau06a, Sen84, Sid06a, Spi76, Tan74, Tro83, Cas80, Tob06]. Académie [Ano77j, Mah78]. Academy [Ano74f, Ano74m, Ano88k, Ano88m, Ano89-33, Ehr11, GG76a, Ken75, Lüt08, Man89, MY99, Sar08, Ste80, Tee84, ABC^{+}88, Arn88a, Kas03, ST16, San03]. Accademia [Ano82e, Ano82-28, Ano90f]. accept [Ano79-34]. acceptance [Hol15, Ung20]. Access [Böt12]. accessible [Vol06]. according [Fre75, dMC92]. Account [Ste01, Wil94]. Accounting [Car18]. accroissements [Ano81j]. accuracy [Car18, Tan79]. accurate [Mae89, Pie83]. Achievements [Nik18]. Ackerberg [SG09]. Ackerberg-Hastings [SG09]. Ackerman [Haw77a, Haw77b]. Acknowledgement [Ano78c, Glu78]. Acknowledgment [Ano79c, Ano80a, Ano81d, Ano82b, Ano83a, Ano84b, Ano85c, Ano88d]. acque [Cap13, Maf10]. Acta [Ano77g]. Actes [Gra01]. Activities [Ano80p, BS77a]. Activity [Rob96, Sac75]. actuarial [Mar20]. actuel [Enr76]. acuta [Hug07]. ad [Bar77, Rid83a, Rid83b, Sch00b, lLW13]. Ada [Hym78, Par89]. Adam [Ano92e, Høy09a, Høy23, Ano92e, Høy23]. Adam-Ries-Bund [Høy09a, Høy23]. Adam-Ries-Bundes [Høy09a]. Adams [Cha76]. ADAPT [Høy07]. ADAPT-SNES [Høy07]. adapté [Jaë19]. Addendum [Ano78d, Wil94]. Additional [Gar75]. additions [Bar78a]. Additive [Pet85, Sam24]. Address [Bur75, Ing11]. Adelard [Eli18]. adic [Ano79w]. adiuvante [Net24]. Admiralty [Roy17]. admits [Bur12]. Adnotationes [Net24]. Adolf [Ano79l, Osw17, Scr04]. Adolph [Ter89, Ano95e, BGM^{+}76, BDG^{+}86, Dem86]. Adolphe [Aub14]. Adrian [Hog77]. Adult [GG90c]. Advanced [Ano74e, Bat06, Zit07a]. advancement [Ano93s, Lew76]. advances [Ano74-27]. adventure [Ano74p]. Adventures [Bir77]. Advisory [BEJ^{+}75]. affiliated [Des07]. affine [ABBS09]. affinity [Win82]. AFIPS [Ano78e, Ano79d]. Africa [Ano74-70, Ger94, Shi86, Wil75a]. African [Shi86, Shi88, Ano86c, Cam76, Cro75a, FG90]. After [Daw16a, Enr85a]. again [Ber90]. against [Cas80, Fek14, Rad03, Sac75, SS88]. Agathe [Wil08a]. Age [Ano93w, Cri04, iÔ75, Syl07, Ano89n, Cre16, Cri06, Por04, Ano93p, Ses13, Hun07, Wil07, iÔ75]. agencies [Wea77]. agendas [Sør14]. Ages [Kin86a, Syl81, Fol83c, Kna88, Ses13]. Agnesi [Fin11, Roe15]. ago [Bau74]. agony [Wil08c]. Ägyptische [Imh03a, BB04]. Ahmad [Sai74, Tee78]. Ahmed [Høy07]. ai [Par97]. Aida [Ung20]. aide [Ano87i]. Aids [Ano74w, Ano74-55]. Air [Rid94, Roy17]. aircraft [Roy17]. Aiton [Gin83, Vek75, Sch00b, Wil91]. Ajima [FC22]. Akademie [Ano81x, Ano84e, Ano92l, Ano93x, Bar78b, Bla05, Cal75, Coo95, Cri17, Gar05, Haf16, Scr03, Seg90, Str78, Ull03, Vol81]. Akademie-Verlag [Ano84e, Ano92l, Bar78b, Cal75, Coo95, Cri17, Seg90, Str78, Vol81]. Akademik [Lin74]. Akademische [Ano74v]. Akihito [iÔ75]. Akivis [Haw96]. Al [Ano76-28, Kin86b, Sal78, Ano81p, Ano87g, Ano92s, Dje98, For10, Hil83, Hog86, Hog89, Ken76d, MP06, Mat74, Ray12, SI19, Sim96, Tee78, Toe96, AF94, Bag97b, BV05, Hog95, Roz76, Ai04, AH20, CH90, Dje98, Ham80, Hog89, Hog02b, Kc74, Oak19, Ras89, Sam24, SI19, Bre07, DS99c, Pin80, Rom99, Sim96, AR76, Ber07, DS99b, Sai74]. al-Asad [DS99c]. Al-B [Sai74]. al-Bīrūnī [Tee78, Kc74]. Al-Bahir [AR76]. al-Balishtar [AH20]. al-Bannā [Dje98]. al-Buzajani [Mat74]. al-Dīn [Hog89]. Al-Daffa [Kin86b, Sal78]. al-Din [SI19]. Al-Farabi [Ano76-28]. Al-Farghani [Ber07]. Al-Farisi [AF94, Oak19]. Al-Furqan [For10]. al-h [Dje98]. al-H*āsib [Kc74]. al-Hassan [Ano81p]. al-Haytham [Ano87g, For10, Roz76, Hog02b, Ras89, Sim96]. al-Hwarizmi [Rom99]. Al-Kāshī [Bag97b]. Al-Karaji [Mat74]. al-Kashi [Ham80]. al-Khāwarizmi [CH90]. al-Khwarizmi [MP06]. Al-Kuhi [BV05]. al-M [Hog86]. al-Murrākusī [Dje98]. Al-Mu’taman [Hog95]. al-Qāuhāi [Ai04]. al-Quhi [Bre07]. al-Rayhān [Tee78]. Al-Sijzi [DS99b]. al-Tūsī [Hog89]. al-Tusi [SI19]. al-Uqlidisi [Pin80]. al-Wafa [Mat74, Ray12]. al-Zanjani [Sam24]. Alabama [Ano87-30]. Alan [Daw16a]. Albany [Ano95k, Ano95l, Lev97]. Albatros [Ano74p]. Albers [Dav19]. Albert [Ano77n, Ano84f, Bre07, Bru07, Fer11, GW94a, Law05, Ped77, Row06a, Sch76a, Wue05, Ano75w, Ano79b, Ano81b, Ano81c, Ano84a, Ano85b, Ano85a, Ano86b, Ano87b, Ano87a, Ano88c, Ano88b, Rob96, Row06b, Wue05]. Alberti [Gra09b]. Alberto [Sch99a]. Albertus [Lo 85, Lo 83, Lo 85]. Albin [Ano95i]. Albis [Ano91-30]. Albree [San03]. Album [SS97a]. Alcuin [BLF92]. Aldershot [Mor96]. Aldo [Isr85, Par97, Ysh90]. Aldrich [SG09]. Aleae [Bel05]. Alejandro [Aus16]. Aleotti [Vér01]. Aleppo [Ano78x, Ano81p, Ano88-31, Tee78]. Alessandra [Rem12]. Alessandrino [Coz12, Tra10]. Alex [Des09]. Alexander [Gam96, Gra11, H^{+}08, Plo06a, Sch11, Str78, Vol81, Wat09, H^{+}08]. Alexander-von-Humboldt-Forschung [Vol81]. Alexander-von-Humboldt-Forschungsstelle [Vol81]. Alexanderson [Dav19]. Alexandre [Gab76, Lor16]. Alexandria [Moy17, CO13, Kno93b, Tra10, Yus95]. Alexandrov [Ano74v]. Alexis [Gre84]. Alfonsi [Gol13]. Alfonsine [Gol80]. Alfred [Swe00, Lê20, Row21, Ano88e]. Algebra [AB02, Ano81o, Bec80, Bee06, Est06, Kje17, Kle01, Kno88b, KK90, MP06, Man06, Mas08, Mel06, Sch17, Sch05, Toe96, Win82, Abh75, Ano74n, Ano76e, Ano79f, Ano84l, Bal04, CO13, CM19a, CM19b, Die75, Dub77, Fla03, FR83, Glu77, Gra94b, Gra94c, GP07, Har11, Hee14, Hol15, Hor91, Høy02a, Høy06, Hug87, Kle07, Mar95, Moo95, Moy19, OA05, OA07, Oak19, Øhr85, Par04, Pyc81, Pyc82, Pyc84b, Pyc97, Rey06, Ric80, Sam24, Ste02, Swe86, Van83a, Ano81i, Høy90b, Hud12, Mar09, Neu08, Bir74, Sai74, Par97, vdW75]. Algebraic [Bir96, Epp95, Fuk87, HK98, Jah93, Mar04, Rey09, Wag15, Abh75, Bas81, BC04, Fra87, Gra91a, Kno06, Sil08]. algébraique [Dra84]. algebraische [Høy09a, Bot80a]. algebraischen [Haf17]. algebraist [Fra88]. Algebras [GG02, Haw80, GG03]. Algebrass [Høy24a]. Algèbre [BKM^{+}82, AR76, Cor96, Fla03, Har06, Pan88, Pet74]. algebrice [Ano81r]. algébrique [Fre76]. algébriques [Lê21]. Algeria [Ano90a]. Algérie [Mar90b]. Algiers [Ano90a, Ano92-49]. algorismi [Fol09a]. Algorismus [Fol09a]. Algorismus-Vorlesung [Fol09a]. algorithm [Kan88, Sha94, vBS20, Sch95]. Algorithmen [BB04, Imh03a]. algorithmic [Coo11]. algorithms [Bul16b, Imh03a]. Ali [Kin86b, Sal78]. Aliabraa [Hug05]. Alice [Ano89l, Eis75]. Alicia [PB08]. Alireza [Ano85e]. alla [Bee09, DN12, Fol14, Hug07, Taz09, Ven02]. alle [Ano82e]. alleged [Kno93c]. allemande [Ano77j]. Allen [Clu76]. Aller [Kni01]. Allgemeinen [Row06a, Wue05]. Allison [SS17]. Alma [Ano76-28]. Almagest [Gol76, Ano77g, Ano88j, Sid13, PJ11, Sid14]. Almanack [Per79]. Alphonse [Bur01]. Also [Ano85h, GG05a, Lew95]. Alta [Pep90, Pep92]. alternate [Ger85]. alternative [Gon08]. älteste [Rom99]. Altmann [Tur08]. altri [Taz09]. Alvi [Ano88k]. always [Lau92]. Am [GG13, Ano82d, Ano93x, Ber04]. AM289 [Fle75]. a‘māl [Dje98]. amateur [Hog90, Inö06]. Amaury [Cog15]. ambition [Gis18]. Ambroise [Ano89g]. Ambrose [BTB^{+}03, Fra06]. Ambrosetti [Oak12]. America [All09, And07, Ano74w, Ano78t, Ano81s, Ayo06, Bid74, Cal08, Cas77, Dru07, Dru12, Enr85b, Enr85a, GG08, Kat07, Lew95, Par07, Sch89, Thi06, Thi08, AA88, Glu05, Hun96, Ken08, Kul76, Lok80, Rei86, Ano85j, Sel91]. American [AH10, Ano74-27, Ano81s, Ano87-31, Ano88h, Arb86, Bid74, Cas77, Coo96, D’A10, Dru85, Fen08, Gra77, Haw96, Haw03, Hod77, Law05, Lo 84, Lor18, Mro16, Neu08, Nic13, Rec07, Row85c, San03, Sch07, Sel91, SG09, Syl81, Tur08, Vek75, Vol77, AG77, Ano74f, Ano74g, Ano74h, Ano76e, Ano77d, Ano78t, Ano80-33, Ano92f, AH99, BC04, Bur87, Bur93, Coc98, D’A79, Fef97, Fis82, Glu11, Hog77, Hog81a, Hog90, Hog91, Jor81, KAHR08, Lew76, Lew78d, Rob96, SS94, Tim04]. Americas [Per10, Bur09]. Amherst [Ano91t]. amicable [BH89, Hog85c]. Amir [Gra11]. Amirouche [Kat20]. Amma [Yan83]. Amman [Mat74]. Amodeo [Taz09]. Among [Lüt20].

Ampère [Fle86a]. Amphora. [Ano95f]. ample [Mat74]. Amsterdam [Ano79-31, Ano81e, Ano84g, Ano84h, Ano95j, Buc77, Enr76, SS07, Str82a, Too80, Ano84-27]. Amsterdam-New [Buc77]. Amsterdam/New [Ano84h]. amusante [Sch76a]. Amy [All09, GW94b, Kat07, GW94a]. an-Nayrīzīs [Rom04]. analemma [Kc74]. Análise [Lei09]. analitica [For75]. analitiche [Roe15]. Analogia [Ano90-32]. analogie [vdW80]. Analyse [Bru07, Ano86g, Ano91s, BKM^{+}82, eFD88, Dug76, FG12, Kit81, Lüt15, Pan05, Tat86]. Analysis [Ano88i, Bir78, Bos77, Cra98, Cra99, Cra00, IM98, Kje00, Rid94, Sch95, SS88, Spa89, Ano77r, Arn88b, AR24, Aup76, Bel15, BK84, Bro75, BS77b, Cal92, Die75, Eli24, FK90, FP03, Gil78, Goo77, GG75c, Jah93, Lau92, Mac75, Mai21, Mik75, Moo08, Pan05, RP19, RW03, SS10a, Sha94, SS12a, Sma16, Sti74, TC05, Tim04, Van84, Zyg75, Arn88b, Bul24, HNE21, Ano74n, Ano77f, Ano81k, Ano84g, GG77]. Analysiskonzeptionen [Ano88i]. analyst [Cra12]. Analytic [EF20, Gra97a, BR16, For77, IN84, Kë96, Ros17, BR16]. Analytica [Dho87, BDD^{+}88]. Analytical [Atz04, Enr83, For75, Lum97, Rei76, Scr83]. analytiques [Pet74]. Anaritu [Vit96]. Anastácio [Rod88]. Anatoliennes [Yus10]. ancienne [Auj76]. Ancient [Est79, Imh21, Kas90a, Ken76d, Mel02, Moy17, Yan83, Ano94h, Che09, Chi22, Dav75, De 09, Eng77, Ger85, Hir74, HP20, Kan88, Kno82, LY79, LYTS86, lLW13, Mar95, Mol76, RS85, SS10a, Sid04, Swe79, Swe86, Wan22, Ano92g, Gup91, Mel09, RT02, Sin85a]. Anderen [SS89]. Anders [Sen91]. Andersen [Gra09b, Toe96]. Anderson [Ayo06]. Andoni [Ano93i]. André [Ano85h, Edw77b, Syl03]. Andrea [Rem12]. Andreas [Pei22]. Andreevich [Zai95]. Andrei [Ano95e]. Andrewes [Lew99]. Andrews [Ano81v]. Andrezj [Moo82b]. Anfänge [Bow84, SD94, Ano89k]. anfängen [Str85]. Ang [Mar95]. Ange [Auv13, Gis14]. Angel [Pan08]. Angela [Bo24]. Angeles [Knu83a]. Angelo [Arc09, Gia90]. Angewandte [Ano90l]. angle [Hog81b, Mah78, Von93]. angled [TS78]. anglice [Net24]. Anglo [Lok80]. Anhang [Lem74]. Anmerkungen [Ano92j, Fri91]. Ann [Ale04, Ano86d]. Anna [GC82]. Annaberg [Ano92e, Høy09a, Høy23]. Annaberg-Buchholz [Høy09a, Høy23]. Annaberg-Buchhoz [Ano92e]. Annals [Ano79e, Ano78e]. Anne [Gab76, Neu15]. Année [Kno88a, Tat86]. années [Dho88b]. Annette [BB04]. annexe [Vol02]. anni [Ben95]. Annick [Ebe96]. Anniversaries [Ano92-31, Bau74]. Anniversary [Ano77a, Ano91c, Dau93a, DPS^{+}02, Gup77a, ADG^{+}88, Ano78a, Ano90e, BGM^{+}76, DPY91, Fel83, GY77, Gup76a, Sau10]. Annotated [Jah12, Plo06a, Ano88j, BJ00, Buc80, DS99b, Lei79, McK82]. annotations [Lo 83]. annoté [Lê21]. Announcement [Ano90b, Kno90]. Announcements [Ano75d]. Annual [Ano76f, Ano85d, Ano87-32, Ano90l, Ano91-33, Ano92h, Ano92f, Ano92w, Ano92-47, Ano93k, Ano94i, Ber85, Gup81, Jon77a, Jon81, Abe91, Ano76-34, Ano85i, Ano90-33, Dun89, Gra76c, Lew76]. annuel [Cha81, Cha87]. annuel-1986 [Cha87]. annular [Moz13]. anomaly [GG14]. anonymous [DR16, Sti76a]. Ansätze [Sch82b]. Anschauung [Sch90b]. Answers [Mon11]. ante [Hog81a]. ante-bellum [Hog81a]. antecedents [Pow76]. Anthology [And07, Ayo04, Gri76]. Anthon [Fol78]. Anthony [Ano89d]. anticipated [Das81]. anticipation [Cri78, Smi82]. Antike [Høy90b]. Antiken [Mue80, Høy90b]. antiquité [Ano89-30]. Antiquity [Kje17, Ano89-33, Bie74a, CO13, Dau85b, Rid88]. Anton [Knu83a, Knu83b]. Antonia [Tan74]. Antonio [Ano76h, Bur01, Fra06, Ser90, Fra88, Ros20]. Antrittsrede [Row85e]. Antwerp [Mes94]. Anuj [Gur24]. Anwendungen [Row19, Ora82]. any [LY80]. AO [Cav85]. août [Bor89]. Apeiron [Høy95]. Apollonius [FU01, Sai05]. app [Ber87, Sal87]. appeal [BEJ^{+}75]. Appear [Ano95z, Ano96h, Ano96i, Ano96j, Ano96k, Ano97h, Ano97i, Ano97j, Ano98k, Ano98l, Ano98m, Ano99g, Ano00g, Ano01g, Ano02k, Ano02l, Ano02m, Ano02n, Ano79-38, Ano79-39, Ano79-40, Ano80-27, Ano80-28, Ano80-29, Ano80-30, Ano81-41, Ano81-42, Ano81-43, Ano81-44, Ano82u, Ano82v, Ano82w, Ano82x, Ano83r, Ano83s, Ano83t, Ano83u, Ano84t, Ano84u, Ano84v, Ano84w, Ano85w, Ano85x, Ano85y, Ano85z, Ano86x, Ano86y, Ano86z, Ano86-27, Ano87x, Ano87y, Ano87z, Ano88y, Ano88z, Ano88-27, Ano88-28, Ano89z, Ano89-27, Ano89-28, Ano89-29, Ano90-28, Ano90-29, Ano90-30, Ano90-31, Ano91z, Ano91-27, Ano91-28, Ano91-29, Ano92-39, Ano92-40, Ano92-41, Ano92-42, Ano93-29, Ano93-30, Ano93-31, Ano93-32, Ano94s, Ano94t, Ano94u, Ano94v, Ano95w, Ano95x, Ano95y, Ano03s]. appear [Ano03t, Ano03u, Ano03v, Ano04q, Ano04r, Ano04s, Ano04t, Ano05s, Ano05t, Ano05u, Ano05v, Ano06o, Ano06p, Ano06q, Ano06r, Ano07j]. appearing [CM22]. Appell [RL16]. Appendices [Yok75, Sch91b]. Appendix [ARR12, Asp83, Wil83, Ano78j, Ken82]. Applicate [Ene18]. Application [Aub14, Car22, Ger74]. Applications [Rad02, Ano79f, Ano87-29, Ano91-31, Boy74, Kan88, LYK85, Sen91, Wal74, Goo77]. Applied [Ano77d, Ano93g, Ken84, Glu05, SS93]. apport [Che94, YDD83]. appraisal [Asp83, HMR17]. Appreciation [Fow98, BDS07, RFS89, GG08]. Approach [WDB96, Abe93, Abe01, Eis80, Gra86, PV22, Pul05, Pyc82, Sau06b, Smi79a, Spi00b, Van22, Ano95j, Dru07]. approche [Spi00a, Spi00b]. appropriation [Maf10]. Appropriazione [Cap13, Maf10]. approximate [Pie89]. Approximation [Plo96, Plo01, SS88, Sto76, TG98, Glu76, Gup74, Mae89, RP19, vBS20]. Approximations [FR98, Eds84]. Approximationssatzes [SS88]. April [Ano74-46, Ano77-29, Ano77-30, Ano81p, Ano85n, Ano86s, Ano87l, Ano88-31, Ano90l, Ano91t, Ano92-35, Ano93j, Ano93w, Che87, Kat14, Müü77, Pur88, Vog94, Gra84b]. Apropos [Bie80]. apse [Cha77]. ARA [Eli24]. Arab [Amb08, Sai74]. Arabe [Ano92-27, Høy07]. arabes [Ano79l, Ano87-28, Ano92-27, Ano93v, Mar90a]. Arabic [Ano77i, Ano78x, Ano78w, Ano81p, Ano88k, Ano88-31, Ano90a, Ano92-49, Ber04, Hog85b, Mat74, Oak12, Tee78, Too80, Van05, AH20, Ano78q, Ano80h, Ano92i, Bre19, De 84, De 05, De 15, Eli18, Eli24, Kno83b, OA05, OA07, Yus95, dY92, Mat74]. Arabic-Islamic [Oak12]. arabisch [Gol76]. arabisch-lateinscher [Gol76]. arabische [Ano91u, Yus95]. arabischen [Rom04]. arabo [Amb08, Oak12]. arabo-islamica [Amb08, Oak12]. Aracne [Rab08]. Aragon [Ano82-27, Ano82-27]. Arbeiten [Ano89k, Böt12, Pei15]. Arbeitstagung [Spa89]. arbitary [Bot78]. Arbor [Jay83]. Arbuthnot [Sho87]. Archer [BM74]. Archief [Ano81e]. Archimede [Sid12, Ano90c, DN12, Fol14]. Archimedea [Net24]. Archimedean [Rab74]. Archimedeo [DN12, Fol14]. Archimedes [Atz06, Sau06a, Sch13b, Zep12, Ber77, Hog02a, Kno78b, Kno93c, MF10, N^{+}11, Sch93a, Wat95]. Archimedis [Net24]. Architectonica [Van10]. architects [BT81]. Architecture [DS99c, Law98, Oli08, FG06, RS85, Law98]. Archiv [Bar78b, Neu93, Ano89k]. archival [BEJ^{+}75]. Archive [Ano74-67, Ano79f, Ano79g, Ano82-28, Bor14, BD11, Luc21, Ano74-75, Haw76, Neu93, Ano74i, Ano78f, Gia01]. Archives [Ano74j, Ano74-44, Ano78l, Ano79r, Ano79q, Ano86k, Bur87, Bur93, Gra75b, Bil85, Fri74a, Lew78d]. Archivio [GG05a]. arctangent [Abe93]. Ardmore [Zas83b]. Area [Hog02a, Ger85, Kic10, Nau03, dV22]. Argenteuil [GW94a]. Argentina [Ano74-61]. Argibra [Hug05]. argument [Sho87]. Argumente [Osw17]. arguments [And11]. arismetica [Sil08]. Aristarchus [Sid24]. Aristoteles [Mue80]. aristotelischen [Mue80]. Aristotle [Cri24a]. Aristoxenos [Høy90b]. Arithmêtikê [Kno93b]. Arithmetic [AÖ01, Ano77s, Bas81, Daw16b, Zla23, AH20, BS13, Bul09b, CV22, Cor16, Dru86, EF20, ESH^{+}13, Gol92, Haf17, Hen75, Kna88, Mar95, Mat74, Pin80, Rad03, Rey06, Sho85, Wan22, dW10, AF94]. Arithmetica [Hog85b, Tho11, dW10]. Arithmeticae [VJRZ87]. Arithmetical [Sch76a, FT83, Hug89, Sai19]. Arithmetick [Sch05, Pyc97]. Arithmetik [Ano81i, Kna88]. Arithmétique [Sen99]. arithmétiques [Ber87, Sal87]. Arithmetische [Cri17, Høy09a]. Arithmetization [Gol00]. aritmetica [Ano85f]. Arlettte [Mah78]. Arlington [Ano77d]. Armand [Gau13]. Armano [GG05a]. Armin [Bul22]. Armonie [Sch99a]. Army [Rid94]. Arney [San03]. Arnold [Eck13, Jur07, Sau14]. Arnzen [Rom04]. Árpád [Bow84]. Arrighi [Ano91k, Mal06a]. Arrow [Lüt19]. Ars [Dho87, BDD^{+}88, Mat14]. Art [Bel07a, Fra17b, Gra09b, Wil77b, BS06, Cro75a, Høy07, Lab01, Ric80, RS85, Shu92, Tra10]. arte [Sil08]. Artemis [Dil81]. Artemis-Verlag [Dil81]. Arthur [Don98, Gil75, Hug76, Cri99, Cri06, CWW17, Wil07]. Arti [Ser90, Amb08, Cap13, Maf10, Oak12]. article [Cog15, ET14, Gra08]. Articles [Ano95p, Ano96b, Ano97b, Ano98e, Ano99b, Ano99a, Ano00b, Ano00a, Ano01b, Ano01a, Ano02e, Ano02d, Ano74-53, Ano75n, Ano76z, Ano77-31, Ano78v, Ano79-27, Ano79-34, Ano80n, Ano81-33, Ano82m, Ano83l, Ano84o, Ano85l, Ano86p, Ano87p, Ano88s, Ano89p, Ano90o, Ano91l, Ano92y, Ano93q, Ano94k, Ano03n, Ano03m, Ano04l, Ano04k, Ano05n, Ano05m, Ano06j, Ano06i, Krö08]. artifacts [Che09]. Artificial [Gra84a]. Artillerymen [Glu11]. Artin [Dum15, Fen08, Gra18, AR07, Art13, FS10]. artis [KO09, Lef10, Syl81]. Artistarium [Vit96]. Arts [Ano74f, Özd00, Amb08, Dho88b, Maf10, Oak12, Maz77]. Arukimedesu [Sid12]. Aryabha [Hay97]. Aryabhatīya [Kel05]. Aryabhata [Ano77e].

Aryabhatiya [Wil08a]. as-Samaw’al [Sai74]. Asad [DS99c]. ascesa [DF10, Rem12]. Aschendorff [Grö86]. Ascher [Kat14, Par92b]. ascribed [Eli24]. Asger [Ste07b]. Asot [Ter89]. Aspects [Fer98b, Rad00, Sch19, Van10, Dad84, Kim75, Mik75]. Aspekte [Sch19]. Aspray [Ano91d]. Assembled [Gab76]. assessment [De 05, De 09, Gui24]. assigner [Maz16]. Assistance [Ano88k, Ano77f, Gab76, GG76a, Gra91b, Wol03]. assisted [Eis75]. ASSN [Ano74-46]. Associados [Lei09]. Associate [Fra10]. Association [All09, And07, Ano74w, Ano78t, Ano79n, Ano79k, Ano79p, Ano81s, Ano89n, Ano91q, Ayo06, Bid74, Cal08, Enr85b, Enr85a, GG08, Gri76, Kat07, Lew95, Par07, Sch89, Thi06, Thi08, Wil90, Ano93o, Lew76, Sen87, Ano76g, Ano90d]. Astrolabe [Van05, Van10, FL05, Ber07]. Astrologer [lLW13]. Astronomers [Kin86a]. astronomi [Kin86a]. Astronomica [Don05, Rid83a, Rid83b]. astronomicae [Bar77, Rid83b, Sch00b]. Astronomical [MY99, Swe01, For78]. Astronomico [Ano87c]. Astronomie [For10, Ano91n]. Astronomiegeschichte [H^{+}08, Sch11]. Astronomiehistoriker [Bie83]. Astronomy [Gup91, Hog01b, Mor96, Mor19, Ano79u, Ano92i, GG87b, GG88, H^{+}08, Hog24, Kin74b, Mar00, Swe79, Van05, Wil05, Ano89h, Ano91f, Pin76]. asymptotic [Sac20]. Ata [Ano76-28]. atelier [eFD88]. Athenæum [DR16]. Athens [Ano92-30, Kas90a, Zer75]. Atlanta [Ano78t]. Atlantic [Ano75p, Ano77s, Pye74]. Atlanticus [Ray23]. Atmospheres [Sho02]. Atomists [Buc77]. Atomphysiker [Eck13, Sau14]. attack [Sti77]. Attempt [Atz04, Gal10]. Attempts [Bec99, Jor81, Ken08]. Atti [Ser90]. attivitá [Ano93j]. attributed [Hog85b, Row21]. attribution [May75a]. aube [Maz21]. Aubin [Dur21]. Audin [Ale14, KS13a, Krö14]. Auditorium [Ano90l, Ano92-50, Ano92-28]. Aufarbeitung [GG07, TC05]. Aufgaben [Høy24a]. Aufgabentexten [Imh03a, BB04]. Aufklärung [Ano88i]. Auflage [Lem74]. Auftrag [SS89]. Auftrage [Str78]. augmenti [Høy24b]. Augsburg [Dho11, Don08, Høy24a, SS08]. August [Ano74-48, Ano74-76, Ano76u, Ano78y, Ano79-31, Ano80p, Ano82k, Ano83k, Ano85-30, Ano87-36, Ano90u, Ano91o, Ano92w, Ano92-31, Ano92-30, Art91, BJ89a, Gra87, Mic88, Neu15, Pur86a, Pyc84a, Rog77, Scr84, Scr87, SY90, Swe92, Ano87-35, Ano90-37, Ano21j, Ano22j, Ano23i, May77c, Scr82b, Str94]. Auguste [Row83a]. Augustin [GW94a, GW94b, Gra94a]. Augustin-Louis [GW94a, GW94b, Gra94a]. Augusto [Ken76b]. Augustus [GG90a, DR16, Pan03]. Aujac [Ano95n]. Aula [Lei09]. auquel [Mal13]. Ausbau [Nag88]. Ausbildungsforschung [Tob90]. Ausdehnungslehre [Law05]. Ausejo [Ric98]. Ausgaben [Yus95]. ausgehenden [Jay83]. ausgewählter [HB^{+}13]. Ausnahmeregeln [Ber90, Lic97]. ausrechnen [KS12, Tob10]. Austin [Ano88h, Ano93f]. Australia [Ano82f, Ano74-73, Tee82]. Australian [Ano81-31, Ano83b, Ano85d]. Austria [Ano90-36, Bin90, Bin91]. Austrian [Ano92-48, Ano93-34, Bin90, Bin91, Ano86l]. Ausz [Ped75a]. Auszüge [Jen86]. auteur [Mar93b]. Author [Row85a, Ano03n, Ano03m, Wol03]. Authority [Mor19]. Authors [Ano95p, Ano95q, Ano96b, Ano96c, Ano97b, Ano97c, Ano98e, Ano98f, Ano99b, Ano99a, Ano00b, Ano00a, Ano01b, Ano01a, Ano02e, Ano02d, Ano74-53, Ano74-54, Ano75n, Ano75o, Ano76z, Ano76-27, Ano77-31, Ano77-32, Ano78v, Ano79-27, Ano79-28, Ano80n, Ano80o, Ano81-33, Ano81-34, Ano82m, Ano82n, Ano83l, Ano83m, Ano84o, Ano84p, Ano85l, Ano85m, Ano86p, Ano86q, Ano87p, Ano87q, Ano88s, Ano88t, Ano89p, Ano89q, Ano90o, Ano90p, Ano91l, Ano91m, Ano92y, Ano92z, Ano93q, Ano93r, Ano94k, Ano94l, Ano04l, Ano04k, Ano05n, Ano05m, Ano06j, Ano06i]. authorship [Dal92]. Autographen [Fol09a]. automathography [Ano87f]. Autónoma [Gar98]. Autores [Lei09]. autres [Vil96]. Auvinet [Gis14]. aux [Ai04, Ano87i, Bre07, Jaë19]. auxilio [Pei22]. av [Bré12]. available [Krö08]. avait [Ano89i]. avalanche [SS10b]. Avant [Mac76, Fol12, Rom10]. Avant-propos [Mac76]. aventura [Ano74p]. Aventures [FG12, Lüt15]. Avril [Kno88a]. awakening [Pin76, vdW76b]. Award [Ano74-66, Ano10k, Par05b]. awarded [Fol94]. Awarding [Ano21a, BG21a, Fra13, Fra17a]. Axel [Fol78, Fol78]. axiom [Dru91, Mos79]. Axiomatics [Lüt09a, Pul05]. Axiomatik [Pul05, Lüt09a]. Axiomatization [Cor04, Sau06a, Moo76, Moo95, Rad03]. Axiomen [Rom04]. axioms [GG75c]. Axworthy [Bo24]. Aylmer [Edw92]. Ayoub [And07].

B [Ano74m, Ano74-45, Ano75v, Ano76k, Ano82d, Ano89k, Ano89m, Ano90g, Boi89, Bor18, Cas77, Cro76, D’A10, Dil81, Dyc75a, Enr76, Gra91b, Haw96, Jay83, Kat92, Kin86a, Kli76, Mue80, Net24, Row85b, Mar93b, Tee78, But81, Sai74]. B. [Ano94h, Isr90]. B.C [Abe91, Høy82]. B.I [Toe96]. B.I.-Wissenschafts [Toe96]. Bīrūnī [Tee78, Kc74]. Babbage [Ano90e, For75, GG92b, Ano82p, Bec80, Dub77, Tee82, Wil77a, Enr80, For75]. Babbage/Faraday [GG92b]. Babin [Her14]. Babylon [Ayo06, AKW04, Rob01]. Babylonian [Mel06, Ano86a, FR98, Fri81a, Høy02a, Høy02b, MW17, Nis89, Pow76]. Babylonians [Høy90c]. babylonienne [Cav85]. Babylonischen [Nis89]. Bachir [Ric91]. back [BGSS16]. background [Cas80, Haw80, Kno78b, Len79, SS88]. Bad [Ano88i, Kit78, Ine91, Tuo23]. Baden [Fol87]. Baden-Baden [Fol87]. Baenst [Hes90]. Baenst-Vandenbroucke [Hes90]. Baer [Ebb04]. Bag [Ano88m]. Baghdad [Moy17, Ken84]. Bahir [AR76]. Bailhache [Ano81m]. Baker [Cro77]. Bakker [Cri24a]. Baldi [Ano82e]. Baley [Ano79k]. Balishtar [AH20]. Ball [Van14]. ballistics [Glu11]. Baltimore [Abe06, Ano88e, Bel07a, Bra12, Hel83, Hud12, Mon17, Per10, San05, SG09, Wil07]. Baltimore/London [Hel83]. Banach [Kar22, Ale04]. Banarsidass [Yan83]. Banchoff [Bay11]. Bancroft [Eis75]. Band [Ano81g, Ano81i, Ano88n, Ano89j, Bar78b, Ber07, Bla05, Bur08, Cri17, Fol87, Gar05, Grö86, Hes90, Høy90b, Jay83, Jur07, Plo06b, Row85b, Sch16, Scr07, SS89, Van06, Vol81]. Bane [Ano74r]. Bank [Mat75]. Bannā [Dje98]. Baotou [Swe92]. Baozong [Ano92-31]. Baptism [Cra98]. Baptiste [Wan22]. Barberini [Mal06a]. Barbiche [Mah78]. Barbilian [BS07]. Barcelona [Gra94a, MP06]. Bari [Isr85, Ken76b, Ysh90]. Barnabas [Gra83, Høy09b]. barocca [Rom05]. Baron [Hog76, War15]. Barrier [Aub02]. Barrow [Fra10, Dia99]. Barrow-Green [Fra10, Dia99]. Bartel [Pin76, DS97, Fre93]. Bartels [Bie74b, Bie74b, Lum97]. Barthélemy [Spi00a, Spi00b]. Bartholomäus [Ano90j]. Based [Her75a, Ang79a, Car18, HNE21, Jay24a, Jay24b, Spi00b]. basée [Spi00a, Spi00b]. Basel [Ano77m, Ano79j, Ano86d, Ano92m, Ano92j, Ano93h, Ano95f, Ano95h, Ano95m, Arc09, Bar77, Bra09a, Bul24, Cas77, Dho88a, Edw80, Fra06, Gis96, GG10, Gra91a, Gra94b, Gra94c, Gra97a, GS91, Hes90, Lew09, McC90, Mon75, Nab04, Nag88, Ora82, Par91, Par92a, Pei22, Row83a, Row83b, Row90a, Sau09, Sca94, Scr08, SS17, Syl83, Syl03, Ter89, Tob90, Ume76, Wil08a]. Basel-Stadt [Dho88a]. Basel/Boston [Ano86d]. Basel/Boston/Berlin [Ano95f, Ano95h, Ano95m, Gra97a, Hes90, Nab04, Scr08, Syl03]. Basel/Boston/Stuttgart [Dho88a, Hes90]. Basel/Stuttgart [Ano79j]. bases [Abe76, RS85]. Bashmakova [Git74, Ano81-35, DPY91, DPS^{+}02, DV07]. Basic [Jon77c]. basis [Wan22]. Basque [Ano91p, dMC86a]. Bath [Eli18]. Batterson [Zit07a]. Battista [Ser90]. Baudhāyana [Ano88m]. Baudhayana [Kic06]. Baum [Bun74]. bayerischen [Jay83]. Bayes [Bel02, Dal88, Dal91, Gil87, Smi80]. Bayesian [Gil87]. Bd [Mue80, Tob90, Toe96]. be [Ano74-74, Ano87f, Das94, Imh21]. bear [Lew75]. Bearbeitet [Ano92j, SS89]. Beatrice [Net24]. Beaune [Mah78]. Beauregard [Zas83b]. beautiful [Tob10]. became [Grc11]. Becchi [Fra06]. Beck [Don05]. Becker [Hei09]. bedeutender [Kno78a]. Bedford [Asp83]. Bedini [Hod77]. Beeley [Maa07, Mos16]. been [Blå17a, Buc23, Pie83]. Beery [Tat11]. before [Ang79b, Bul09b, Enr85b, Maz77, OS19, Rom10, San05, San06, Sch91b, Sen91, Ano93f]. Beginn [Ano85e, Ano93x]. Beginnenden [Fol83a]. Beginning [Ken76d, Ano78-31, Fri21, Høy06, Sti76b]. Beginnings [BM96, Epp95, Bow84, Hog74, Lew04a, SD94, Str85]. Begriffsbildungen [Nis89]. Begriffsschrift [Vil98]. begründung [Sch82b]. behalf [Ano87n]. behandeltes [Ine91]. behind [Imh21]. Beiheft [Mon75, Ume76]. Beijing [Ano85-30, Ano92-31, Ano94g]. Being [Sid06b, ET03, TE03]. beiten [Ano89k]. Beitrag [Bos77, Jay83, Row85d]. Beiträge [Dho88a, H^{+}08, Jay83, Sch11, Tob90, Vol81]. Beiträgen [Hes90]. Belgrade [Ano78a, Ano78r, Sti78]. Belhoste [Gra94a]. Beliefs [How18]. believed [Sch93b]. Belin [Ano89g, Ano89i, Ano92k, Mal97, Ric91]. Belknap [Dau12]. Bell [Lew95]. Bellé [Net24]. Belles [Ber87, Eva94, Fol14, Sim96, Sal87]. bellum [Hog81a]. Belmont [Dav90]. Belna [Fer98a]. Beltrami [Gra99a, BR16, Ros17, Taz97, Thi97, Vac20]. Beltz [Tob90]. Bemerkung [Fel78]. Bemerkungen [Ine94, Joh82]. Bendegem [Ano95k]. Benedetti [Ser90]. Benedetto [FR83]. Benin [Cro75a]. Benjamin [Hun12, Tim09, Hog08]. Bennett [BT84]. Benno [Van06]. Benoit [Gis96]. Bento [Sil08]. Benutzung [Hes90]. Benvenuto [BTB^{+}03, Law98, Fra06]. Beppo [Mos79]. ber [Ano89k]. Berg [Kit78]. Bergbeamter [Ano92e]. Berggren [Ano89e, Plo06a, Moy17]. Berghe [Ste09]. Berichte [Sau14]. Berkeley [Gra83, Lor18, Scr84, Scr85a, And11, Ano95g]. Berlin [Ano77i, Ano78j, Ano81i, Ano83h, Ano84e, Ano84j, Ano86a, Ano87j, Ano87g, Ano88j, Ano89f, Ano92l, Ano92m, Ano92j, Ano92v, Ano93g, Ano93h, Ano93i, Ano95f, Ano95h, Ano95m, Bar78b, Bor18, Cal75, Che14b, Coo95, Cri17, Das81, Edw77a, Edw77b, Fre92, Gra94a, Gra13, Gra97a, Has04, Has10b, Has10a, Hes90, Hog85b, Krö14, Lem74, Lew09, Nab04, Par91, Par92a, Pei15, Plo06b, Row85d, Row07, Sau09, Sca94, Sch17, Scr07, Scr08, Seg76, Seg90, Str78, Str81, Swe79, Syl03, Vol81, Wil08a, Ano93k, Has10a, SS93]. Berlin-Heidelberg-New [Edw77b]. Berlin/Diepholz/Stuttgart [Has04]. Berlin/Heidelberg [Ano89f]. Berlin/Heidelberg/New [Ano87g, Fre92, Plo06b, Scr07, Swe79]. Berlin/New [Ano81i]. Berliner [Ano81x, Cal75, Seg90]. Bern [Rid83a, Rid83b]. Bernard [Ano92-27, Gab76, Gra83, Gra06, Høy07, Kit78, Mac76, Mah78, Maw06, Mue96, Sen99, BR04, Sch93b]. Bernardino [Ano82e]. Bernardo [Ges18]. Bernd [Høy23]. Bernhard [Lau99, Lo 83, Sch82a, Wol03]. Bernhardt [Ano74v]. Bernoulli [Ano92j, Hes90, Nag88, Syl03, BFG94, Dal88, Mat14]. Bertin [Str82a]. Bertram [Cog12, Has04, Sch09]. Bertrand [Ano87-32, Ano93h, Ano75e, Ano81z, Ano84c, GG78, Lew85c]. Berücksichtigung [Ano85e, Row85d]. besoin [Ano89i]. besonderer [Ano85e, Row85d]. Bessel [Neu93, Neu93]. Bessel-Hagen [Neu93, Neu93]. Besserat [Ano93f]. bessmertie [Dyc75a]. Bestrebungen [Row85d]. Beta [Ano74-66]. Bethlehem [Tim09]. Betti [Bot77]. betweeen [Ano07e]. Between [Arc09, How18, Law98, Ano79-30, Ano80p, Ano80-31, Ano86i, Ano87-29, Ano91-31, Arn88b, Axw18, Boo86, BC04, Ehr11, FS22, Fer93, Fla03, Fre89, Gue88, GM04, GN05, Gui01, Gup91, Hog10, Isr90, IM98, LY80, Men86, Özd00, PGN18, PV22, Sch13a, Scr83, SS94, SS10b, Sti76b, TC05, Taz18, Zho21]. Bewegung [Sch17]. Beweis [Joh84, SS88]. Beyond [Gra18, Kno06, Aud11, Pyc84b, Sar15, VF22, Wat83, Lor18, Ale14]. Bézout [Gol13, Alf11, Sar15]. Bhāaskara [BK06]. Bhārata [Gup77b]. Bhāskara [Kel05]. Bhaskara [Wil08a, Sel75]. BI [Sch96]. Bialas [Don05, Don08]. Bianchini [Ros81]. Biard [Syl07]. Bibhutibhusan [Gup80, Jon76c]. bibl [Ber87, Sal87]. Bibliografia [Fio17]. bibliographic [Gis18, SS84, You86]. Bibliographie [Ull03, Fol83b]. Bibliographies [BG75a, Cas77]. bibliographique [Gis18]. bibliographiques [Mar93b]. Bibliographisches [Ric84]. Bibliography [Ano79h, Ano92i, Ano94h, CJ84, Cro76, Enr84, Kat92, Ric84, Row85a, vM84, Ano74z, Ano76-36, Ano81l, Ano87i, Ano88k, Ano95j, Ano95l, Buc80, Dau85b, GG75a, Lei79, Sch91a, Sch91b, Scr75a, Thi07, Wag78, Ano88h, Asp83, Wil83, Neu86, Rid88, Ano74o]. Biblioteca [Ano82e, Ges18, Hug05, Rom97, Rom05]. Bibliothek [Fol09a]. Bibliothéque [Fre76, Ano92-50, Ano92-28]. Bicentenary [FG90, GG92b]. Bicentennial [Ano87c, Ano94j, Cas77, dBR76, Ano76i, Ano77z, Ano77-27, Ano78r, Ano81s, Ano81e, BS77a, Str82a, Ano76h, Ano90e]. bicomplex [Cer17]. Bicylinder [Hog02a]. Bielefeld [Ano84k, Ano92h]. Bielefelder [Tob90]. Bielfeld [Yca16]. Bienayme [Das81]. Biermann [Ano89s, Cal75, Seg90, Str78, Vol81, Ano79-42, Ano89s, FS03]. Biggs [Har77a]. Bildung [Ano84k]. Bildungsgeschichte [Ano88n, Sch90b]. Bilinski [Ano82e]. Bina [Tee84]. Binary [Knu83b, Knu83a]. binomial [Yad80]. biobibliographic [Ano89l, Scr75a]. biobibliography [Wal74]. Bioesmat [Gra12]. Bioesmat-Martagon [Gra12]. Biografie [Eck13, Sau14]. biographical [SS10b, SS12b, SS22b, Tob06]. Biographie [Bie74b, Bos03, Fol83b, GG13, Gra12]. Biographien [Kno78a, Row85b, BMW08, Sch09]. biographies [BMW08, Kno78a, Kno78a]. biographische [Ano89f]. Biographisches [Tob06, SS08]. Biography [Ano74k, Bie74b, Gar01, Tim09, Von88, Ait86, Eck13, Hog08, Ken74b, Ken79b, Moo75b, NT05, Pee95, Pet09b, Gra94a, GG10, Wal17]. Biometric [Ano76u, Nor75]. biometry [Ano76u]. Birkhäauser [Ano95f]. Birkhauser [Che84, Hil83, Kla84, Sca94, Ano79j, Syl83, Wil08a, Ale04, Ano77m, Ano86d, Ano91g, Ano92m, Ano92j, Ano93h, Ano95h, Arc09, Bra09a, Bul24, Dho88a, DR19, Edw80, Fra06, Gis96, GG10, Gra91a, Gra94b, Gra94c, Gra97a, Gra10b, GS91, Hes90, Lew09, Lüt20, Mar09, McC90, Mon75, Nab04, Ora82, Par91, Par92a, Pei22, Row83a, Row90a, Sau09, Scr08, SS17, Syl03, Ter89, Ume76, Wol03]. Birkhduser [Ano95m]. Birkhoff [Ano77f, Ano75v, Ven02, Zun02]. Birth [Ano77a, DPS^{+}02, Gup77a, ADG^{+}88, Ano90e, DPY91, Edw92, For77, GY77, Mat14, Pin76]. Birthday [Ano79-42, Ano89s, Ano90j, Fre93, Scr96, Ano81-35, Ano91a, BDG^{+}86, Bur75, Fol84, Kno93a, Mar10b, Thi07, You83, Ano95f]. Births [Ano92-31]. bis [Bos77, BMW08, Che84, Fel85, Høy90b, Lüt09a, Pul05, Pur86a, Sch09, SS08, Tob06]. bisected [Car22]. bisher [Fol83a]. Bishop [KS75]. Bjarnadóttir [Kat11]. Björnbos [Fol78]. Bkouche [Ano92k]. Black [Fol76, Zas83b]. Blaise [Mar10a, Syl07, Des10]. Blanchard [Ano77n, Bre07, Bru07, Fer11, GW94a, Ped77, Sch76a, Vol16]. Blanton [Cal92]. Blick [Böl93]. blind [Tat92]. blizhnem [Ken75]. Blocher [Row83a]. blooming [SS12b]. Bloomington [GG86]. Bloomsbury [Gra13]. Blue [Sho02]. BM [Gon08]. Board [Ano09a, Ano10b, Ano10c, Ano10d, Ano10e, Ano11b, Ano11c, Ano11d, Ano11e, Ano12a, Ano12b, Ano12c, Ano12d, Ano18a, Ano18b, Ano78m, Ano03i, Ano03j, Ano03k, Ano03l, Ano04g, Ano04h, Ano04i, Ano04j, Ano05i, Ano05j, Ano05k, Ano05l, Ano06e, Ano06f, Ano06g, Ano06h, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09b, Ano09c, Ano09d, Ano13a, Ano13b, Ano13c, Ano14a, Ano14b, Ano15a, Ano15b, Ano15c, Ano15d, Ano16a, Ano16b, Ano16c, Ano16d, Ano17a, Ano17b, Ano17c, Ano18c, Ano18d, Ano19a, Ano19b, Ano19c, Ano19d, Ano20a, Ano20b, Ano20c, Ano20d, Ano20i, Ano21b, Ano21c, Ano21d, Ano21e, Ano22a, Ano22b, Ano22c, Ano22d, Ano23a, Ano23b, Ano23c, Ano23d, Ano24b, Ano24c, Ano24d, Ano24e]. Boccard [Yus10]. Bochner [Gil89]. Bochum [Ano86h]. Bodil [Atz04]. Bodleian [War15]. Body [Dia99]. ‘Boethiu [Mat75]. Boethius [Ano89j, Ber07, Bur08, GG07, Gra90, Høy24b, KS12, Mat75, Sch11, Van06]. Boethius-Texte [Ano89j]. Bogolyubov [Por75, Thi08]. Bohemia [Bos10]. Bohigan [Ano74l]. Boi [Gra99a]. Bois [RC85]. Bois-Reymond [RC85]. Bokhove [Bul22]. Boktryckeriet [Tho76]. Boletín [Ano91-30]. Bollati [Sai94]. Bölling [SS97a]. Bölling. [Coo95]. Bologna [Arc17, Ken82, Pan18, Mar99]. Boltzmann [Sau10]. Bolyai [Ano74p, Ano89k, Kis99, Ano74p]. Bolzano [Kit78, BR04, GA20, RKL05, Rus80, Sch93b, Gra06]. Boncompagni [Ano78f, Ano79i, Isr90]. Bonn [Mon83, Neu93]. Boockmann [Don05, Don08]. Book [Abe06, Abe09, AH10, AH16, Aga75, Ale04, Ale14, All04, All09, And07, Ang79a, Ang88, Ano74o, Ano74p, Ano74r, Ano74m, Ano74q, Ano74s, Ano74u, Ano74t, Ano74-27, Ano74l, Ano74w, Ano74v, Ano74n, Ano74x, Ano74y, Ano74-28, Ano74-29, Ano74z, Ano74-30, Ano77i, Ano77j, Ano77k, Ano77h, Ano77q, Ano77l, Ano77m, Ano77n, Ano77r, Ano77s, Ano77p, Ano77o, Ano77g, Ano77f, Ano78g, Ano78h, Ano78i, Ano78j, Ano79j, Ano79k, Ano79l, Ano79m, Ano80b, Ano80i, Ano81g, Ano81s, Ano81t, Ano81k, Ano81p, Ano81e, Ano81f, Ano81h, Ano81v, Ano81i, Ano81j, Ano81l, Ano81n, Ano81o, Ano81r, Ano81q, Ano81m, Ano81u, Ano82c, Ano82f, Ano82d, Ano82e, Ano84e, Ano84f, Ano84g, Ano84h, Ano84j, Ano84d, Ano84i, Ano84k, Ano85e]. Book [Ano85f, Ano85h, Ano85g, Ano86d, Ano86e, Ano87d, Ano87e, Ano87j, Ano87g, Ano87k, Ano87h, Ano87i, Ano87f, Ano88e, Ano88n, Ano88f, Ano88g, Ano88h, Ano88j, Ano88i, Ano88k, Ano88m, Ano88l, Ano89d, Ano89c, Ano89e, Ano89f, Ano89g, Ano89h, Ano89j, Ano89i, Ano89k, Ano89l, Ano90f, Ano91d, Ano91e, Ano91f, Ano91g, Ano92k, Ano92n, Ano92l, Ano92m, Ano92j, Ano92o, Ano93f, Ano93g, Ano93h, Ano93i, Ano95f, Ano95g, Ano95h, Ano95j, Ano95k, Ano95i, Ano95m, Ano95l, Arc96, Arc99, Arc09, Arc17, Asp83, Atz04, Atz06, Aub14, Aup76, Aus16, Ayo06, Bar74, Bar77, Bar78b, Bar82a, Bar82b, BKM^{+}82, Bay11, Bee09, Bee06, Bel07a, Bel08, Bel12, Ber74, Ber87, Ber07, Bid74, Bie74a, Bir74, Bir77, Bir78, Bir80, Bir83]. Book [Bir97, Bla76, Bla05, Bo24, Bor14, Bor18, Bos77, Bos03, Bow84, Boy74, BB04, Bra07, Bra09b, Bra09a, Bra12, Bra17, Bré12, Bre07, Bru07, Bru77, Buc80, Buc77, Bul16a, Bul24, BT81, BT84, Bun74, Bun76a, Bun76b, Bur08, Cal75, Cal83, Cal92, Cal08, Cap13, Cas77, Cas80, Cav95, Cha76, Che94, Che84, Cib76, Cif00, Clu76, Cog12, Cog15, Coh82, Con06, Coo95, Coo96, Coo99, Cor96, Coz12, Cri17, Cri24b, Cro77, D’A10, Das81, Dau93b, Dau12, Dav75, Dav76, Dav19, Dav90, Daw16b, De 20, Des09, Dho85, Dho11, Dia99, Dil81, DS99b, DS99c, Don05, Don08, Dra84, Dru91, Dru07, Dru12, Duc16, Dum15, Dum18, DR03, DR19, Dur21, Dyc75a, Dyc75b, Ebe96, Edw77a, Edw77b, Edw80, Eis75, Enr76, Enr80, Enr85b]. Book [Enr85a, Eva94, Eva13, Eve76, Ewi92, Fen08, Fer98a, Fer11, Fer04b, Fol87, Fol09a, Fol12, Fol14, Fol15, For10, For75, Fra04a, Fra06, Fra10, Fra17b, Fre92, Fre76, Fre77, Gab76, Gal12, Gam96, Gar05, Gar98, Gau13, Ges18, Gil15, Gil75, Gin83, Gis96, Gis99a, Gis14, Git74, Giu99, Glu12, Gol76, Gol82, GW94a, GW94b, Gol13, Gra77, Gra94a, Gra76a, Gra83, Gra10a, Gra11, Gra13, GG05a, GG76a, GG77, GG80, GG86, GG90c, GG90b, GG90a, GG05b, GG05c, GG07, GG08, GG10, GG13, Gra90, Gra91b, Gra91a, Gra94b, Gra94c, Gra96, Gra97a, Gra98a, Gra99a, Gra99b, Gra06, Gra08, Gra09b, Gra10b, Gra12, Gra18, Gre85, Gri76, Gri82, Grö86, GS91, Gui96, Gur24, Haf16, Hal78, Hal84]. Book [Har77a, Har06, Has04, Haw77a, Haw81, Haw96, Haw03, Hei81, Hei09, Hel83, Her14, Her75a, Hes90, Hil83, Hir74, Hoc96, Hod77, Hod75, Hog85b, Hog03, Hog24, How18, Høy90a, Høy90b, Høy95, Høy07, Høy09a, Høy09b, Høy23, Høy24b, Hud12, Hug76, Hug05, Hug07, Hun07, Hun12, Hym78, Isr85, Jah12, Jam13, Jay83, Jon76a, Jon77c, Jon85, Jur07, Kat92, Kat07, Kat11, Kat20, KS12, KS13a, Kel05, Ken76a, Ken75, Ken76d, Ken76c, Ken76b, Ken78b, Ken82, Kil83, Kin86a, Kin86b, Kit78, Kit81, Kje17, Kla84, Kle17, Kno80, Kno88a, Kno07, Knu11, Knu74, Knu83a, Kot79, Krö06, Krö14, Krö15, Krö17, Lan07, Law98, Law05, LY79, Lê21, Lê24, Lef10, Lei79, Lei09, Lem74, Lev97, Lév99, Lew90, Lew95, Lew98, Lew99, Lew09, Lin74, Lo 84, Lo 85]. Book [Lon21, Lor16, Lor18, Lüt08, Lüt09a, Lüt15, Lüt20, Maa07, Mac76, Mah78, Mal97, Mal06a, Mar10a, Mar14, Mar09, Mar95, MF10, Mat74, Mat75, Maw06, Maz77, Maz16, Maz21, Maz24, McC90, McG74, McK82, McL12, Mel06, Mon75, Mon83, Mon17, Moo75a, Moo77, Moo82a, Moo82b, Moo03, Mor96, Mor19, Mos16, Moy17, Mro16, Mue80, Mue96, Nab04, Nak03, Nee79, Net02, Net24, Neu86, Neu08, Nic13, Nor07, Nor03, Oak12, Oli08, Ols76, Ora82, iÔ75, Pan08, Pan09, Pan18, Par89, Par91, Par92a, Par92b, Par97, Par07, Pec20, Ped77, Ped75b, Pei15, Pei17, Per10, Pet82, PP78, Phi79a, Pin76, Pin80, Pla76, Ple78, Plo06a, Plo06b, Por05, Por75, Pri80, Pye75, Pye76, Pys81, Rab08, Rec07, Rei15, Rei76, Rem12, Ric98, Ric82, Ric84]. Book [Ric91, Rid83a, Rid83b, Rid88, Rid94, Rom97, Rom05, Rom99, Rom04, Ros74, Row83a, Row83b, Row85b, Row85c, Row85d, Row85a, Row90a, Row93, Row06a, Row07, Row08, Row15, Row19, Roz90, Rus75, Sag91, Sai74, Sai94, Sai05, Sal78, Sal87, San03, San05, San06, San09, Sau06a, Sau09, Sau14, Sca94, Sch76a, Sch99a, Sch02, Sch91a, Sch91b, Sch91c, Sch11, Sch13b, Sch16, Sch17, Sch89, Sch07, Sch95, Sch96, Sch90b, Sch05, Sch09, Sch12, Sch13c, Sch19, Sch00b, Scr03, Scr07, Scr08, Seg76, Seg90, Sel74, Sel90, Sel91, Sen84, Sen91, Sen99, Sen06, Sen10, Ser18, Ser90, SG04, SG09, Shu80, Sia23, Sid06a, Sid06b, Sid12, Sid13, Sid24, SS89, SS97a, SS07, SS08, SS17, Sim96, Smi79b, Soi09, Spi00a, Spi76]. Book [Ste05, Ste09, Ste80, Str78, Str82a, Swe79, Swe00, Syl81, Syl83, Syl03, Syl07, Tan74, Tat75, Tat11, Taz09, Taz15, Tee78, Tee84, Ter89, Thi06, Thi08, Tho76, Tim09, Tob90, Toe96, Too80, Tot79, Tro74, Tro83, Tru75, Tur08, Ull03, Ume76, Ung85, Van05, Van80, Van06, Vek75, Vil96, Vil97, Vit96, Vol81, Vol02, Vol16, Vol77, Wal17, Was20, Wat09, Wil75a, Wil83, Wil75b, Wil77b, Wil08a, Wil05, Wil07, Wil08b, Wis76, Wol91, Wol03, Yan83, Yin11, Yok75, Yon75, Ysh90, Yus10, Zas83b, Zas83a, Zep12, Zer75, Zho21, Zit07a, vM84, vM07, vdW76a, vdW80, AH20, Wal75, vH67, vdW75, vdW76b, Ber77, BV05, Erl03, Fow92a]. Books [Ano85h, Ber87, Gin83, Jon77c, iÔ75, Sal87, Tan74, Too80, GG90a, Her75b, Jon75a, Por19, Sai19, Xu05, Gol00, Hog85b]. Bookshelf [McK82]. Boole [Ano87e, Jon85, PB08, Par02, Sen92, Wol08, Gri76, Was20, Ric91]. Boolean [Gri76]. Booth [GG90c]. Bordeaux [Neu89]. Bordoni [Ros20]. Borel [Lüt15, Maz21, FG12, Gra76b, LBB^{+}04, NB77, Maw06]. Borgato [Rom05]. Boringhieri [Sai94]. Boris [ADG^{+}88, GY77, Vog10]. Borri [Car07]. Bos [Str82a, Gui24, Mar10b]. Bosanquet [Ano81u]. Boscovich [Ano90f, Ano87c, Ano87l, DF18]. Bose [Ano74m]. Boston [AH16, Ale04, Ano76u, Ano77r, Ano82c, Ano84i, Ano86d, Ano91g, Ano92m, Ano92j, Ano93h, Ano95f, Ano95h, Ano95m, Bow84, Che84, Cri24a, Dho88a, Fra06, Gra91a, Gra94b, Gra94c, Gra97a, Gra10b, Hes90, Hil83, Kla84, Lew90, Lew09, Mar09, McC90, McG74, Nab04, Par91, Par92a, Pin80, Row83a, Row90a, Sau06a, Sau09, Sca94, Scr08, Ste09, Syl83, Syl03, Wil75a, Wil08a, Wol03, vM84, vdW76a, Lew76]. Boston/Basel [Fra06, Row90a]. Boston/Basel/Stuttgart [Row83a]. Boston/San [Lew90]. both [Hog24]. Bottazini [Sch99a]. Bottazzini [Arc17, Ken82, Arc99]. Böttcher [Pei15, Sch09]. Boulder [Pri80]. Bound [vM84]. Boundaries [Hun96]. Bourglinster [Ano92q, Ano92u]. Bowley [Wil77b]. Box [Ano74w, BT81, Gri82]. Boyer [Ano76w, Ano76k, Cro76, Kli76]. Bradley [GG08, Jah12]. Braduardini [Syl07]. Bradwardine [Rom13]. Brahmagupta [Kic10, Kic12]. Brakel [Ott88]. Branch [Epp95, Law05]. Brandt [Jen86]. Branner [Atz04]. Brauer [Haw03]. Braunschweig [Ano85e, Bie74b, SS89, SS97a, Bie74b]. Braunschweig/Wiesbaden [SS89]. Brazil [Ano87-31, Ano89o, Ano90m, Ano74-63, dS90, dS99a]. BRD [Ano84k, Ken76c]. BRDO [Ano93a, Vog94]. BRDO/Manetin [Vog94]. bread [Imh03b]. Breaking [Gre88]. Bremen [Ano91c]. Brendan [Nak03]. Brennan [Ros74]. Brepols [Gil15, Maz21]. Brera [Ano87c]. Brescia [Ano85g, Ano88r]. Brest [Ano92a]. Brewer [Row83b]. Brian [Ano95i, Knu74]. Brief [Böl93, Bra17, Neu88, SS84]. Briefe [Bar78b, Bla05, Gar05, Ano80b]. Briefen [Bul22]. Briefwechsel [Ano77m, Ano81x, Bar78b, Bie80, Gar05, Haf16, Hes90, Jur07, Sch14, SS89, SEM04, Coo95, Str78, Vol81]. Brieskorn [Mon83, Plo06b]. Brigaglia [Isr85, Par97, Ysh90]. Brill [Hog03, Hog24, Sai05, Cri24a]. brilliant [Hog95]. Bringing [Lew75]. Britain [Ano88l, Sag91, Cra99, Des07, Fle75, Fri74a, Mai21, Pan87]. British [Ano87n, Des09, Sch05, Ano79n, Ano80c, Ano83w, Ano84x, Cra07, Cra12, Cri82, Del03b, Fle86b, GG87a, Gra76c, GS77, Gra78b, Pan03, Pyc81, Pyc84b, Pyc97, RW03, Ric80, Ver15, Wal74, Wil90]. Britton [Ste11]. Brno [Sis02]. Brodén [Pej07]. Broke [Aub02]. Bronislaw [Ano82e]. Brookline [Goo77, Haw77a, Haw77b]. Brooks [Par92b]. Brooks/Cole [Par92b]. Brouwer [Ser18, Wil83, Fra95, Fre75, Van79, Wil83]. Brown [Ano74s, Eis75, SS17]. Brownian [Kah75]. Bru [Maw06, Sen99]. Bruce [Ano87j, Ano91f, Per10]. Bruckheimer [Eve76]. Brüder [Nag88]. Bruderschaft [Dil81]. Bruins [Kno91a]. Brummelen [Bra07, Eva13, Fra17b, Moy17, Ano03a, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c, Ano04d, Ano05a, Ano05b, Ano05c]. Brun [Ano80-36]. Bruna [Ano92n]. Bruno [Gra94a, GG05b, Str81]. Brush [Buc77]. Brussels [GG76a]. Bryan [Moo75a]. BSB [Row85b]. Bt [Ano90-34]. Bucarest [Dug82]. Bucciarelli [Ano84i]. Bucharest [Ano82l, Ano82k, Scr82a]. Buchgesellschaft [Ano81g, Lüt09a]. Buchholz [Høy09a, Høy23]. Buchhoz [Ano92e]. Buchreihe [Grö86]. Buchwald [Bul16a]. Budan [Bor91]. Buddhist [Jay24b, Jay24a]. Buffalo [Ano74-40]. Bühler [Ano89f]. Building [De 19]. BULAC [Eli24]. bulletin [Gis18]. Bullettino [Fio17]. Bulmer [Ano80q]. Bulmer-Thomas [Ano80q]. Bund [Høy09a, Høy23]. Bundes [Høy09a]. Bundesstr [Ano90t]. Bunsenstraße [Ano90-33]. Burali [MG81]. Burali-Forti [MG81]. Burdick [Per10]. Burgers [Alb94]. Bürgi’s [FLT16]. Burgundy [Ano81q]. Burning [Ano77i]. Burns [D’A10]. Burnside [Haw03, Wag78]. Burova [GG80]. Burt [Ols76]. Burton [Pri80]. Busard [Bur08, Too80, Eva98, Fol09b]. Bushmen [Tuo23]. Business [Ros74, Scr82a, Wat09]. Bustamante [Maz21]. Butterworth [Ano74u]. Buzajani [Mat74]. Buzón [Sid24]. Bychkov [Dav76]. byed [Høy24a]. Byron [Hym78]. Byways [Bra12]. byzantine [Chr98, Chr98].

C [Ano74-27, Ano74-30, Ano79b, Ano81b, Ano81c, Ano84a, Ano85b, Ano85a, Ano86b, Ano87b, Ano87a, Ano87d, Ano88c, Ano88b, Ano89d, Ano89k, Ano89m, Ano90g, Ano92k, Ano95f, BKM^{+}82, Bie80, Boi89, Bré12, Cal08, Cib76, Das81, Dav90, Don05, Edw80, GG75a, GG86, GG94, GG08, Hal78, Jah12, Joh84, Jon85, Law05, Mon75, Neu86, Nic13, RH11, San03, Scr96, Sid24, Str78, Ume76, Zer75, Høy82, Tat88]. C. [Abe79, Ano76i, Ano76j, Ano86w, Bie80, Bie83, BLW77, Bul09b, Eis78, Gre88, Joh84, Jon77b]. C2 [Ano09a]. CA [Knu83a, McK82, Par92b, AH20, Rom97]. caché [Bré12]. Caen [Ano95n]. cafe [Kla84]. Cahier [Krö06]. Cahiers [Ano80d, Ano92k, Mal97]. Cairo [Mat74]. Caius [Ano93m, GG87b, GG90d]. calcolo [Amb08, Ano92s, BFG94]. calcolodell’Europa [Oak12]. Calcul [Jaë19, Maz16, Cré87, Pan88]. calculate [Tob10]. calculated [Buc23]. calculates [Ken84]. Calculating [FG90, Imh03b, Amb08, Car22, FLT16, Rei86]. Calculation [Dal92, Jaë19, CS81, Oak12, Vol97, War15]. Calculations [Fin11, Moz13, RP19]. calculs [Bré12]. Calculus [AB02, Bec99, Cra99, Kat07, Nag88, Thi97, And11, Blå12, Cal83, Dom08, Fer04a, Fra87, Fra92, Fra09, Gre84, Gui87, Kat87, Ken77, KK90, Man89, Mar20, Pan87, Ros77, Sag91, SGJ05, Ste20, Tat75, Bra09a, Dho85, Glu77]. Calcutta [Ano77e]. Calendario [Don05]. calendrier [Bré12]. Cali [Arb86]. California [Bun74, Dav90, Gra83, Hug74]. Calinger [Ano86e]. call [Buc77, Par03]. called [Blå17a, Sin85a]. Cambridge [Aga75, Ano74n, Ano74y, Ano77f, Ano81l, Ano81u, Ano84d, Ano92n, Ano92o, Ano93m, Bay11, Bir83, BT84, Coh82, Con06, Dau12, Dav19, Des09, Dho85, Dum18, Enr80, Fer04b, Gab76, Gra11, GG87b, GG90d, GG92b, Haw81, Her75a, Hir74, Hoc96, Ken76a, Ken08, Kno80, Lew98, Lew99, Moo77, Nak03, Par89, Rus75, Sag91, Sch91b, Sch13b, Sch05, Sen06, Sia23, Tro74, Ung85, Wil83, Wil75b, Yok75, Yon75, BGG06b, Cra07, Cri99, Cri04, CH90, Enr83, Swe00]. Cambridge/New [Ano92o, Hoc96, Lew98, Sag91]. came [Das94]. Camille [Bil85, Joh77]. Campan [Ano74p]. Campanus [Cor16, Bur08]. Campbell [Ano89l, Dav90, Por05, Wee91]. Campbell-Kelly [Por05]. Campinas [Lei09]. Campus [Ano91p]. Can [Ano74o, Pye75]. Canada [Ano83k, Pye74, Ano79-43]. Canadian [Abe91, Ano87-32, Ano74-31, Ano74-32, Ano74-33, Ano74-47, Ano75f, Ano76f, Ano77t, Ano78k, Ano85i, Ber85, Cha81, Cha87, Jon77a]. Canadienne [Ano74-32, Ano74-33, Cha87, Jon77a, Ano75f, Ano78k, Cha81]. Canguilhem [Bun76b]. Cannstatt [Ano88i, Kit78]. canons [Høy04, Mah84]. Cantelli [Ben88]. Cantor [Ano89k, Arn88b, Fer98a, GG07, Haw81, SS89, Arn88b, Déc10, Fer93, Fer95, Jen76, TC05, Che14b]. Cantorian [Ano89c]. Cantors [Jen76]. Caramalho [Bra09a]. Carathéodory [Thi97]. Carbondale [Ano82d]. Cardano [Ano79j, Ano81f, Bel05]. Cardinality [TC05]. Cardinals [TC05]. career [Ano74-28, SD94]. Carignano [GG05b]. Carinthia [Too80]. Carl [Ano76w, GG13, Ine94, Joh82, Str78, Ano76k, Cro76, DGD04, For78, Gar79, Joh82, Kli76, Sen94, Ano74q, Neu86]. Carleman [Alm87]. Carlini [Sac20]. Carlo [Ano85g, Tro83]. Carlomagno [Hug07]. Carlton [Ano87n]. Carlyle [Cra00]. Carman [Sid24]. Carnegie [Fen03]. Carnot [Hei81, Tat75]. Carolyn [Ano77s, Pys81, Ket82]. Carroll [Abe09, Dru07, Dru12, Ped75b, Abe76, Wil08c]. Carruccio [Ano81-30]. Cartan [Ano84f, Ano89m, Ano90g, Boi89, Fre77, Haw96, CM18]. Cartesian [Kno06, Mar10b]. cartography [Jon76b]. cas [BH14, DP88, Gis99b, Dug76]. Case [Asc95, Asc97, Coc98, Fen97, Gra98a, IM98, Rob96, Ven02, Ano87j, CV22, Ene18, Glu05, Hee14, Ols76, Par04, RL16, Taz18, Tuo23]. case-study [Ene18]. cases [Mar04, Zhu20]. Caso [Ven02]. Casorati [PR22]. Caspar [Atz04]. Cassano [Ano92s]. Cassiano [Fol14, DN12]. Castelnuovo [IN83, Taz09, Gar01, Sch99a]. Castle [Wil77b]. Catalan [Gis99a, BF23, Rey06, Rey09]. catalog [Cro75a, Nag91a]. Catalogue [Sch13b]. Catalonia [Rey06]. Catania [Ano90c, Ing11]. categories [Høy04]. Category [McL12, Ott88]. catenary [RP19]. Catherine [Che94, Coo99]. Catholic [TC05]. Catholique [Kit81]. Cattle [Wat95, Sch93a]. Cauchy [Ano88u, Dho85, GW94a, GW94b, Gra94a, GG76a, Bar11, Bur12, Dho87, Dho90, Fis78, Gra78a, Haw75a, Lau87, LN94, Meo04, Sør05, Jah12]. Causality [Hod75]. Cavalieri [LYK85]. Caveing [Mue96]. Cavendish [vM07, MS05]. Cayley [Wil07, Cri78, Cri86, Cri88, Cri99, Cri06, CWW17]. Cazenave [Ano79l]. CBI [Ano79d]. CBMS [BEJ^{+}75]. CCCLXI [Ano81g]. ccvi [Ber87, Sal87]. ccxlii [Rid83b]. CD [GG05a, Her14, Jur07, Rem12]. CD-ROM [Jur07, Rem12]. CD-ROMs [GG05a]. CE [CM19a, LP10, CM19b]. Cecilia [Maz21]. Cedex [Ano89m]. Celebrate [Cas77]. celebrates [Ano91a]. Celebrating [Ano89-35]. Celebration [Ano94j, Edw92, Ano77z, Ano81e]. celebrations [Ano77e, Ano78r, Ano87i]. céleste [Bré12, Fra04a]. celestial [Sho03]. Census [Vol77]. Cent. [Rab74]. centenaire [Ano92r, dSG10, Row15]. Centenary [Mil93, Edw92]. Center [Ano78y, Pet82]. Centers [Ano78h]. Centina [Rem12]. Central [Bel12, Cha76, Blå12, Kah75, Fis11]. Centre [Ano78l, Ano79r, Ano79q, Ano83o, Ano86k, Ano89-33, Ano90-34, Che87, GG76a, Hei81, Hou89, Koe94, Pei17, dBR76, Str82a, Too80, Fre89]. Centro [Ano81w, Ano82e, Ano83c, DG90, Ges18, Hug05]. Centrum [Ano81e, Too80]. Centuries [Ano81h, Ano89-35, BLF92, Dem74, Kin86a, Cas80, Fri81b, Gra09a, GN92, Zit07b, Ai04, Ano74z]. Century [Ano74-27, Ano86h, Ano88a, Beb96, Bo24, Gra97a, Hog95, Jam88, Ken76d, Lew98, lLW13, Lüt20, Pep92, Pye76, Taz01, Wag85, Yon75, Ano80-31, Ano82q, Ano87-29, Ano91-31, BF23, Ben95, Bla14, Bow87, BC04, Bro75, Car07, Cha90, Che17, CM19a, CM19b, DF21, Dom04, Fin11, Fol15, Fri21, Gar86b, Gol92, Gra94b, Gra94c, Hog90, Hog86, Høy83, Hug87, Jan11, Jay24a, Jay24b, Lüt09b, Mah84, Mar20, Mat74, Moo82a, Nee79, Nik18, iÔ75, Per80, Pyc84b, Rab24, Rey09, RW03, SS86, Sel91, Sep05, Sin85b, Sor13, Sti76b, Tra21, Tur81, Ung78, Van10, Van24, Ver15, Wag79, Wan22, War08, Yca16, Zhu16, Zhu20, Böt12, BS13, Bul09b, Che10, Cra99, Cra00, Cra07, Cri99, Del03b, Dij04, Ehr11, FP03, Fol83a, Fra88, Glu05, Gol00, Hun96]. Century [Ine94, Ken84, Kë96, Lok80, Mai09, Mar04, Mes94, Plo96, Rey06, Sar08, SD08, Tra10, WDB96, AH16, Ano82c, Atz06, Bir78, Buc77, Cog12, Des09, Kje17, Lüt08, Row90a, Cal83, Tro83]. cercle [Che94]. Cerisy [Ano90i, Bre91]. Cerisy-la-Salle [Ano90i, Bre91]. certain [GG90a]. Cesare [Cap13]. Cetraro [Roe89, Roe89]. Ceulen [Hog10, dW10]. Ch [Vol81]. Château [Ano92q, Ano92u]. Châtelet [Böt12, Böt12, Pei15]. chair [GC21]. Chairman [Dau94, Scr79b, Scr85b]. Chakravala [Sel75]. Challenges [BF00, Pre24]. Challis [Wil77b]. Cham [Fol15]. champ [Fre77]. Champion [Bré12]. chance [dMC92, Ano92o]. Chances [Bel07b]. Chancrin [Mah78]. Chandler [Ano87j]. Chandrasekharan [Fre92]. Chang [Lê24, Yon74, Yon75]. change [Cro75c, Cro75b]. changed [Lüt19]. Changes [Del03b, PH97, Glu11]. Changing [Mah84, Pul98, Abe76, Cas80, Lüt09b]. Chantilly [Pet82]. Chao [Gil77]. Chaos [AD02, Ano91q]. Chapront [Fra04a, Wil08b]. Chapter [Wil08a, BK06, Row83c, War15, Yon74]. chapters [BJ00, Shu92, Plo06a, Ano81e]. Characteristica [Deb13]. Characters [Ano86r, Mic88, Shi93, Zho21]. Charles [Ano74-30, Ano77s, Dru07, Dru12, Enr80, GG86, Haw03, Hod77, Hug76, Jon76b, Kat20, MS05, Ped75b, Pys81, Rid94, Tat75, Tro74, vM07, Abe93, Abe01, Abe10, Ano74-34, Ano82p, Ano90e, Auv13, Eis82, Nag82, Tee82, Gis14]. Charles-Ange [Auv13, Gis14]. Charlier [SS93]. Charpit [Ano80i, GGE82]. charts [Ano93f]. chasing [May75b]. Chasles [Sma16]. Chatterjee [Tee84]. Chatterji [Plo06b]. Chebyshev [BJ89b, Eds84]. chemists [Ano07e]. Chemla [Gis96, Lê24, Vol10]. Chen [XX10]. Chern [Fel85]. Cheshire [Hal84]. Chetverikov [She93]. Cheyney [Zas83a]. chez [Fer98a]. Chicago [Ano74x, Ano88f, Ano91e, Ano95g, Cri24b, Cro77, Dru85, Jam13, Shu80, Ano74-58, Par04]. chief [GG86, GG90b]. Chiefly [Atz04]. Chiffrieren [KO09, Lef10]. chimistes [Ano07e, Sau09]. Ch’in [Yon75]. China [Ano85-30, Ano91g, DS^{+}02, Swe92, Van06, Vog83, Yok75, Ano82j, Che09, Che12, Che10, Che17, Kan88, LYTS86, LYK89, Swe74, Vog83, Zhu16, Zhu20, LY79, Mar95, Mor19]. Chinese [Ano85-30, Ano86m, Hud12, LY79, Mic88, Nee79, iÔ75, Ano83i, Ano86r, Cul07, DZ97, Dun89, Har11, Jam88, LY80, LYK85, Mar97, SS86, Shi93, SS84, TS78, Wag79, Xu05, You86, ZD88, Zho21, Ano89d, Hir74, Yon75]. Ch’ing [Gil77]. chinois [Bré12]. chinoise [Bré12, Che94]. Chinoises [Ano87h, Che94, Mar93b, Ple78, Dau93b]. chisel [Ano81o, Sel74]. Chisholm [Ano84h]. chiu [Yon75, Yon74]. chiu-chang [Yon75]. Chiu-Shao [Yon75]. choice [Dru91, GG75c, Mos79]. choisies [Moo82b]. Choix [GW94a, Gra76a]. chords [DF21]. Chorlay [Krö17]. Christa [Ano88f]. Christiaan [Ano79-31, Atz06, Dij04, vM85]. Christian [Ano77-30, GG07, Lor16, Lüt15, Mar06, Vol81, Sid24, FG12]. Christiania [Bir96]. Christianidis [Sid06a]. Christianity [CJ84]. Christine [Yus10]. Christoffel [Str81, Ano79-29, But81]. Christoph [Maa07, Mos16, Ano93j, Bee14, FK89]. Christophe [Cog15]. Christopher [BT84, Mor19, San06]. Christos [McK82]. chronological [GG86, SS84]. Chronology [Ano76k, Scr75a, Van84]. Chün [Gil77]. Chuprov [Sen87, She93]. Chuquet [Dra84]. Church [Hal84]. ci [Arc17]. ciencia [Ano82-27]. Ciencias [Ano90-32]. Cifoletti [Mal97]. Ciliberto [Par97]. cinquecento [Ser90, Tra10, Coz12]. ciphering [KO09]. ciphers [Shu80]. circa [Syl07, Fol09a]. Circle [Gui01, LYTS86, Lüt20, Coo11, Eng77, Ger85, Gre88, Hog10, Kic06, Kno91b, Scr83, Str23, Von93]. Circles [FT99, Ped75a, Zho21, Gil77]. Circolo [CB21, Isr85, Par97, Ysh90]. circular [Joh84]. Circulation [Ene18, Ste20]. Circuli [DN12, Fol14]. circulo [Kno91b]. circumference [Coo11]. circumradius [Gup77c]. Cirino [Bee09]. CIRM [eFD88, Dho83]. Ciro [Par97]. citations [Tuo23]. cited [Høy95]. Citrabhanu [HK98, Wag15]. City [Mro16]. Ciudad [Ano92-44]. Civilization [BLF92, Ano92i, Bul16a, Ber74]. civilizations [Swe86]. Civita [Goo77, Cog16, NT05]. Clagett [Mur07]. Clairaut [Lew98, Gre88, Kat81]. Clara [GG05a]. Clarendon [Ano89d, Ano89c, Bru77, Har77a, Jon85, Moo03, Nor03, Pec20, Ric82, Row93]. clarifications [Chr07]. Clas [Scr92]. Clas-Olof [Scr92]. Class [Gra18, Ang79b, Ang79b]. Classic [Ano86f, SS86]. Classical [Bel12, Hod75, Pul98, Ano77f, Das80, Fis11, Bar74, Sch91a]. Clássicos [Ges18]. Classics [Ano86e, Sid06a]. classique [Ano89n, BKM^{+}82]. classroom [Ano82-29, SGJ05, Kat07]. Claude [Ano87h, Bor14, Bré12, Dau93b, Dur21, Fre76]. Claudia [Wil75a]. Claudius [Gol76]. Clavius [Ano93j, Axw18, Kno07]. clé [Kno07]. Clebsch [Lê20, Row21]. Cleonice [Mal97]. Clermont [Mar10a]. Clermont-Ferrand [Mar10a]. Clifford [Fra06, Bos03]. CLM [Jay83]. closed [Moo08]. Closely [Ken08]. Closs [Ano88h]. closure [Lê18]. Cloth [Ano74s, Ano81l, Ano84i, Bul22, Ang88, Ano74n, Ano89e, Ber87, Dav90, Sal87, Zas83b]. Club [RWG95]. Cluj [Ano74p]. clvii [Bul22]. cm [Maz21]. CNRS [Ano86g, Fra04a, Har06, Kle17, Lor16, Wil08b, Ano91r]. CNRS-UPR [Ano91r]. Co [Ano74u, Ano77r, Aup76, Bun74, Kat92, Ols76, Par92b, vdW76a]. Coach [BT81]. Codazzi [Phi79b]. codes [Shu80]. codex [Hug87, Ray23]. coefficients [Fer04a]. cognition [Jay24a, Jay24b]. Cogwheels [Edw04, Abe06]. Cohen [Gab76, SS17, Was20, Ano74-45, Ano75v]. Cohen-Mansfield [SS17]. Cole [Par92b]. Colección [Gar98]. Colin [Gau13, Hal84]. coll [Dur21]. collaboration [BKM^{+}82, Bie74a, Cog15, Cri24a, Her75a, SS10b, Taz09]. collana [Rom97]. colleagues [Sch12, VMT09]. Collected [Ano89h, Bar77, Edw77b, GG90b, HB^{+}13, Mor96, Par91, Pec20, SS22a, Ano78-32, Ano83d, Ano93u, Gra76b, Hal84]. Collection [Ano77j, Ano95n, Ber87, Bre07, Fre77, Gra12, MY99, Sal87, Vil97, Høy07, Kno78a]. Collections [Bur93, Mah78, Luc18]. College [All04, Ano74-46, Ano77-30, Ano79i, Ano83f, Ano85-28, Ano86s, Ano91t, Ano93m, Cas80, Cri82, GG87b, GG90d, iÔ75, Row89, Zas83a, Che94, Ano87h, Maz21, Ple78, Del03b]. colleges [Ano74x, Ano83z]. Collegio [Rom97]. Collette [Ano77k]. Collingwood [GG75b]. Collins [Arc96]. Colloque [Ano92q, Boi89, Bun76b, DG90, Obe78, Ano87-28, Ano92-50, Ano92a, Ano93v, Gra01, Mar90a, Ano86g, Ano88o, Ano89m, Ano89n, Ano89t, Ano90i, Ano90g, Ano90h, Ano92p, Ano92r, Bie78, eFD88, Hou89]. Colloque-atelier [eFD88]. Colloquio [Ano92s]. Colloquium [Ano79o, Ano83h, Gar86a, Gar92, Ano82o, Ano93n, Fol76, Ano90j, Vol93a]. Colloquy [Ano93x, Ano91n, Ano92e, Ano92-28, BDD^{+}88]. Colombia [Arb86]. Colombian [Ano93n]. color [Ano74w, Sid12, Wil76]. Colorado [Pri80, Pri80]. Colorful [Soi09, Wat09]. Coloring [Soi09, Wat09]. Colour [BLW77]. Coloured [Ano87d]. Columbia [Ros74]. Columbus [Hug76]. combinations [Ste07a]. Combinatorial [Bul24, Ano87j, SK24]. Combinatorics [Ano78u, Big79, KB74]. come [Ken78b]. comedy [GG77]. Comisión [Ano85j]. comité [Bab77]. commemorate [Ano90e]. Commemoration [Ano83e, Ano87c, Dun89]. comment [Eis78, Sør05]. commentaar [Lo 85]. commentaire [Cog15, ET14, Ped77, Gra76a]. commentaires [Mue96]. Commentaries [Ine88, Jay24b, Sch05, Vol10, Bre19, Che17, Gra99a, Mat74, Pyc97, Vil96, Jay24a]. commentarires [Ano81m]. Commentary [Gin83, Knu11, KS75, Lo 85, Mon17, Row85e, Sch13b, Ste80, Tee78, Vit96, ALB10, Ang79a, Ano77i, Asp19, Coo95, De 19, Kel05, Ken76c, Lo 83, Oak19, Syl81, Tee84, VHK16, Yon74, Bay11, Ano88m]. Commentatio [FK90]. Commentationes [Bar77, Sch00b, Rid83a, Rid83b]. commentator [De 19]. commented [Sen99]. Comments [Gra76b, Ine94, SI19, BS77b, Gil78, She93, Wil79a, You86, Zun83, Joh82].

Commercium [Ano77m, Pei22, Ter89]. Commission [Ano08e, Bor89, Dau86, Dau93a, Dau94, Kno94, May75c, May77c, Par03, Pre24, Scr79b, Scr85b, Scr85a, Dau82a, Scr82a, Scr82b, Scr84]. Committee [Ano85q, Ano87n, BEJ^{+}75, Wil90]. common [GG90c]. commonly [Sch93b]. Commonwealth [Big19]. Communication [Ano74-56, Ano74-40, Tro75]. communities [Ano74x]. Community [Coc98, Cri04, Par03, Tim04]. Como [Ano89-35]. compacité [Pie80]. compact [Ano81n]. Companion [Fra10, GBGL08, GG98, Gra98b, Kat97]. Company [Ano74-27, Ano82f, Ano86e, Zas83b]. Comparative [SS12a]. comparée [Ano78i]. Comparison [Bec01, PZ11]. compasso [Fri21]. compendium [Zas83a, Spi00b]. Compendy [Spi00a, Spi00b]. compensating [And11]. compilation [Hog86]. Compiled [Ano74n, Neu86]. completed [Ano79t, Ano79v, Ano79u, Ano83g]. completeness [Bur92]. Complètes [Wil08b, Fra04a, GG76a, Kle17]. Completing [Ace11, Daw85]. Completion [Hog02b, Ano87g]. Complex [EU99, Gam96, Ale95, Fla03, Glu77, Wan22, FS22]. complexen [Osw17]. complexes [DR03, Fla03, Har06]. Complutense [Ano90-32, Ano92-44]. Component [Ano79-45]. computability [Por19]. Computation [Ful99, Zhu16]. Computational [Høy02b]. computer [Ano79d, Tro74, Wil77a]. Computers [Kat07, Knu74, SGJ05, Asp83]. Computing [Kid09, Par89, Ano79e, Smi79a, Tro83, Ano90e, Ano78e]. Comunale [Hug05]. Comune [GG05a]. Con [Fol14, DN12]. Concavity [GM04]. Concept [Ano86a, Zor96, Ano91n, Gra80a, Gra91b, Hof92, Ken74a, LYK85, Moo07, Nic93, Phi79c, Sch79, Yad80]. Conception [Ano95l, Wol03, Lam19, Lau99, Mar95]. Concepts [Ano86h, Sen06, Ano78g, Nis89, Wil79a, Yca16]. Conceptual [Høy04, Cro75b]. Conceptualism [Gui03]. Concerning [Bee06, Cro75c, Cro75b, EG81, Gro87, Ste02, Tat75]. concise [Ano74m, Ano89d]. concordance [Bur86]. Concorde [Ano91n]. concrete [Bro75, PV22, Van22]. concursu [Vil96]. condition [Nik09, Sch92]. conditions [Fra09]. Condorcet [Ano88o, Cré87, Gra76a]. Condorcet. [Cro77, Sen99]. Conference [Ano74-46, Ano75g, Ano76u, Ano76-28, Ano77-30, Ano80l, Ano81-31, Ano85k, Ano85-30, Ano86m, Ano88-29, Ano89o, Ano90e, Ano90m, Ano90s, Ano91t, Ano92t, Ano92x, Ano92-29, Ano93j, BGG06a, Dem74, DV80, DK86, Eng81b, Eng81a, GG92b, Gup81, Kas90a, Kat11, Mac78a, Nag91b, Nag91c, Pye74, Swe92, ZD88, Ano75v, Ano76-34, Ano85-28, Bre91, Ros89, Ano75e, Ano79p, Ano83f, Ano86h, Che87, Lew85c]. Conferences [Rog74, Ano76v]. Conferenze [Ano82e]. Conflict [Gol00]. Confluences [Sør14]. confraternal [Gis18]. confraternelle [Gis18]. confronté [Mal13]. congeneric [Cer17]. Congratulations [Ano74-35]. Congrès [Cha81, dMC86a, Dug82, Cha87]. Congress [Ano80k, Ano82l, Ano82k, Ano87-31, Ano89-34, Ano91o, Ano94x, Ebb07, Gup85, Rog77, Scr84, Scr87, Shi86, Shi88, Ano74-76, Ano76i, Ano85-31, Ano87-35, Ano87-36, Ano91j, Ano93s, Ano93t, Ano77-42, Ano80-39, Ano83x, Ano86i, Ano86j, Ano90n, Ano90-37, Ber75, Scr85a, Scr88a]. Congress-1985 [Gup85]. conic [Ano80e, Ano80-34, Vac20, Mai09]. Conica [FU01, Sai05]. coniche [Mai09, Mar14]. Conics [Hog02b, Ano87g]. Conjectandi [Mat14]. Conjecture [San05, Fri81b, Llo90, Pet85, Tai14, Wil76, BLW77]. conjectures [Sma16]. Conjecturing [Bel07a, BS06]. conjunction [Ano90n]. conjuntos [Gar98]. connaissance [Fol12, Rom10]. connaissances [Cav95]. connected [Ken08]. Connecticut [Ano75p]. connecting [Alb94]. connection [Ano74-31, CM18, LY80, Sch77, All04]. Connections [Özd00]. Conoids [Kno93c]. Conservation [Cog12, Gra08]. Conservatoire [Dho88b]. Considerazioni [BFG94]. Constance [Edw77a, Lew95]. constitution [Lê18]. construct [Gal10]. Constructing [GG02, GG03, Luc18]. Construction [Erl97, Ano88l, Bré12, Fer11, Ped75a]. Constructions [Bur92, Blå23, Sid04]. constructivism [Sep05]. Consuegra [Ano93h]. consulting [Glu05]. consummatio [Syl81]. contacts [BJ89b]. Conte [Sch99a]. contemporaine [Fre76]. contemporaries [Sch93b]. Contemporary [Ano74i, Ano79q, Ano86k, Ano91h, Ano92-29, Hod75, Ano85-27, Bis75, KKS18, Mug87, VHK16, Ano78l, Ano79r]. Content [Dau99, Sil08]. Contents [Ano03e, Ano03f, Ano03g, Ano03h, Ano04e, Ano04f, Ano04u, Ano04v, Ano05e, Ano05f, Ano05g, Ano05h, Ano22l]. Context [Dau99, FR98, Sai05, Ano76-29, Ano79-32, FU01, Ott89]. Contexts [Mar10b]. contextualism [Gui03]. Contextualized [Kje00]. continental [Pan87, Sho87]. Continu [Ano90i, Ano86g, Bre91, Bré12]. ‘Continuation [lLW13]. Continued [Kic15, Cre14]. Continuing [GG90c]. continuités [Lor16]. Continuity [Fer00, Arn88b, DF18, RKL05]. continuous [Cog14]. continus [Cog15, ET14]. continuum [Cox75]. Contributing [GG94]. Contribution [AB02, Høy24b, Lum97, Roz90, Chr84, Lüt82, Por82, Sal78, SM18, Vac20, YDD83]. Contributions [Edw77b, Rad00, Taz01, Thi97, Ano80f, Ano89u, Eng80, Fis94, Gar79, H^{+}08, Mar93a, Pie93, Pin76, Row94, Spr78, Ste20, Thi06, Zun83]. contributo [Par97]. Contributors [Ano95u, Ano96d, Ano96e, Ano96f, Ano96g, Ano97d, Ano97e, Ano97f, Ano97g, Ano98g, Ano98h, Ano98i, Ano98j, Ano99c, Ano99d, Ano99e, Ano99f, Ano00c, Ano00d, Ano00e, Ano00f, Ano01c, Ano01d, Ano01e, Ano01f, Ano02g, Ano02h, Ano02i, Ano02j, Ano79-35, Ano79-36, Ano80w, Ano80x, Ano80y, Ano80z, Ano81-39, Ano81-40, Ano82s, Ano82t, Ano83p, Ano83q, Ano84r, Ano84s, Ano85s, Ano85t, Ano85u, Ano85v, Ano86t, Ano86v, Ano86u, Ano87u, Ano87v, Ano87w, Ano88x, Ano89x, Ano89y, Ano90x, Ano90y, Ano90z, Ano90-27, Ano91v, Ano91w, Ano91x, Ano91y, Ano92-36, Ano92-37, Ano93y, Ano93z, Ano93-27, Ano93-28, Ano94o, Ano94p, Ano94q, Ano94r, Ano95v, Ano95s, Ano95t, Ano03o, Ano03p, Ano03q, Ano03r, Ano04m, Ano04n, Ano04o, Ano04p, Ano05o, Ano05p, Ano05q]. contributors [Ano05r, Ano06k, Ano06l, Ano06m, Ano06n, Ano07f, Ano07g, Ano07h, Ano07i, Ano08f, Ano08g, Ano08h, Ano08i, Ano09e, Ano09f, Ano09g, Ano09h, Ano10g, Ano10h, Ano10i, Ano10j, Ano11g, Ano11f, Ano11h, Ano11i, Ano12e, Ano12f, Ano12g, Ano12h, Ano13d, Ano13e, Ano13f, Ano13g, Ano14c, Ano14d, Ano14e, Ano14f, Ano15e, Ano15f, Ano15g, Ano15h, Ano16h, Ano16e, Ano16f, Ano16g, Ano17d, Ano17e, Ano17f, Ano17g, Ano18e, Ano18f, Ano18g, Ano18h, Ano19e, Ano19f, Ano19g, Ano19h, Ano20e, Ano20f, Ano20g, Ano20h, Ano21f, Ano21g, Ano21h, Ano21i, Ano22e, Ano22f, Ano22g, Ano22h, Ano23e, Ano23f, Ano23g]. Contro [Ano92l]. control [Ano89-35, SS94, Rus75]. controversie [Ben95]. controversies [Ben95]. controversy [IN89, Scr83, Sho87]. Convegno [Ano90c, Ser90]. Convention [Ano90l]. Convergence [Fer07, Ano86d, LN94, Nik09]. converging [Scr83]. Convolutions [Par92a]. Cooke [Gra97a]. Cooper [Bun74]. Coordinated [Gis96]. Copenhagen [Ano77g, Ano87d, Atz04, Sid06b, BPL17]. Copernic [Mal13]. Copernicana [Ste80]. Copernicus [Ken76a]. Copper [Sch99b]. copy [Ano74-56, Moy19]. Copyright [Pri80, Glu12]. Cora [Ano81u]. Cordula [Row23]. Coriolis [GG15]. Cork [Was20]. Cornelius [Bla76]. Cornell [Ano81u, Tru75, Ano78f, Coc98]. corollary [Erl94]. Corp [Ros74]. corporibus [Van80]. corporum [Vil96]. Corpus [DN12, Fol14]. Corradi [Fra06]. Corrado [Gia01]. corrected [Moo77]. Correction [Ano75h, Ano75j, Ken78a]. Corrections [GG76b]. Correspondance [Hah11, Kle17, Ter89, Ano07e, Bil85, Sau09, Vol02, Gil15]. correspondances [Ano92-50, Ano92-28]. Correspondence [Alm87, Ano76l, Ano77u, IN83, Maa07, Mos16, Pec20, Ser18, SEM04, Ang79a, Ano74-27, Ano07e, Bie80, CB21, Dau80, Daw85, Ebb04, Fle91, GG79, GN92, HMR17, Isr90, Jen86, Joh77, Jon85, MS05, Men86, MB20, Sch86a, Sch13a, SB12, TC05, vM07, Sch14]. Correspondent [Hog76]. Corrigendum [CM19a, Del03a, Gui09a]. Corry [Bra17, Sau06a]. corte [Hug07]. Cortona [Pep90, Pep92]. Cosenza [Roe89]. cosmicam [Sch00b]. cosmographicum [Gin83]. Cosmology [Ano83n, Ano85n, Ano91e]. Cosmos [H^{+}08]. Coß [Høy23]. Costabel [Ano87i, Hes90, Mac76, Mah78, Ter89, Tru75, Tat91]. Cotes [Gow95]. Couchoud [Cav95]. Coulston [Tat75]. counterexamples [Sør05]. Counterparts [Bec01]. Countess [Hym78]. Counting [Ano93f, BB21]. countries [Ano74-60, Bus76]. Country [Ano91p]. counts [Wil75a]. coupe [d’E01]. couple [GG14]. Courant [Wil80, Edw77a]. Cours [Jah12, Maz21, Bar11]. course [Ano89-32, Cam76, Eis80, Fle75, Hug75, Mal79, Pet78c, Wil79b]. courses [Ano74-71, BR16, Ros17]. Court [Bel07a, Law05, BS06]. Courtine [Ano77j]. covariación [Pan08]. Cover [Ano09a, Bor18, BT81, Haw03]. cradle [Fre75]. Craft [Bra07, De 19, VK05]. Craik [Des09]. Cramer [Hed99]. Crane [Ano74u, Ano74x]. Creating [Hoc96]. Creative [Ano74r]. Creativity [Dau12]. Creators [Soi09, Wat09, All04]. Credibility [Dal92]. Cremona [BD11, Men86]. Crépel [Sen99]. crest [Kat92]. Crilly [Wil07]. criminal [Ano74l]. Crippa [Lüt20]. crisis [Bis75, CB21, Fle86b, KKS18]. Cristoforo [Car07]. crítica [Ges18]. Critical [Fel85, Giu86, Her75a, Sen99, Sim00a, Sim00b, Tee84, Gra83, Gra99a, HMR17, LD22, Nee79, PV22]. Critiche [Ven02]. Criticism [Pul98, Ven02, Pyc82]. critique [Ano77j, Ano81m, Das79, Høy04, Mac76]. Croarken [Por05]. Croom [Sal78]. crops [Fie22]. Crosby [Swe00]. cross [Hee14]. cross-cultural [Hee14]. Crosse [Ano90s]. Crossley [Ano82f, Ano89d]. Crossroads [DR19]. Crowe [Wil79a]. cryptanalysis [Lei77]. cryptography [Lei79]. CS [Ano89h, Mor96]. CSHPM [Ano75i, Ano75j, Ano75k, Ano76m, Ano76n]. CSHPM/SCHPM [Ano75i, Ano75j, Ano75k, Ano76m, Ano76n]. Cube [Ano79s, Joh11, Tai14, Sai95]. cubic [Hof87, Hog89, Kic15]. Cullen [Mor19]. Culmann [Has04]. cult [Ken79a]. Cultivating [Tur20]. cultura [Coz12, Tra10, Ser90]. cultural [Hee14, Rai92, Roe15, Wil74]. Culture [Ano76-28, Eck13, Kni01, Tra10, Lev97]. Culturel [Ano89t]. Cultures [AD02, Ano92g, Asc84, Asc88, Zas79, Zhu16]. culuslibet [Syl81]. cuneiform [Gon08, Pro22, Ano93f]. Cuneo [GG05a]. Cunha [Rod88]. cura [Rom05]. curaverunt [Net24]. Cures [Lan07, Nah06]. Curie [Ano81j, Ano81n, Ano89m, Ano91q, Ano91s, Kno88a, Tat86]. current [Ano76o]. Curriculum [Ano79-45, Scr75b, Sti78]. Curtis [Haw03]. curvature [Dar15]. curve [McC94]. curves [Bas81, Fuk87, Row21]. Cusanus [Grö86]. Cusanus-Gesellschaft [Grö86]. cxxxiv [Ber87, Sal87]. Cybernetics [Kot79]. Cyclic [Ung23, Gup77c, Kic10, Kic12]. Cylinder [MF10, Ber77]. cylindrical [MA10]. Czech [Ano93a]. Czechia [Ano93a, Vog94].

D [Ano74r, Ano74m, Ano77r, Ano77o, Ano87n, Ano89h, Ano95f, Cal92, Cas77, Cri24a, Des09, Don08, Fer11, Fol09a, Gra77, Gri76, Hes90, Pet82, Plo06b, Wil83, vM84, vdW76a]. D. [AB02, vdW76b]. D.C [Ano86s, Enr85b, Enr85a, Lew95]. Dīn [Hog89]. d’abaco [Van80]. d’accumulation [Ano81n]. d’Aepinus [DP88]. Daffa [Kin86b, Sal78]. Dahan [GW94a, GW94b]. daily [Imh03b]. d’al [AD87, Hug05, Bee09, Dje97]. d’al-Haṡṡār [AD87]. d’al-Mu’taman [Dje97]. D’Alembert [Kle17, Fra04a, Wil08b, HS15, Bal04, Das79, Müü75, Van83b, Lam19]. Dalen [Ano77o, Ser18, Van06, Wil83]. d’Alessandro [Fol14]. d’Alexandrie [Chr91, Mue96]. Dalla [Ano92s, Taz09]. dall’Ateneo [Rem12, DF10]. Dalmedico [GW94a, GW94b]. Dalton [Ano81s, Gra77]. Daly [Van80]. Damas [Sai74]. D’Ambrosio [Sco22]. Dan [Ano77g]. d’analyse [Jah12, Bar11]. Daniel [Cri24a, Gam96, Hes90, Kla84]. Daniela [Row06a]. Daniele [Fol14, Net24]. Daniels [Ols76]. Danish [Ano78y, Ram00a, Ram00b]. dans [Ano92p, Ano92-50, Ano92-28, Ano92-27, Ano92-32, Bil85, Bre07, Cog15, ET14, Lam19, Mal13, Pan88, Ple78, Spi00a, Spi00b]. Dante [Ine88]. Dante-Kommentare [Ine88]. D’Antonio [GG08]. d’après [Tat86]. Darboux [Ale95]. d’Archimède [Ano91n]. d’architecture [d’E01]. Dardi [Hug05, Van83a]. dargestellt [Høy90b, Bos77]. d’Aristote [Ano86g]. d’Aritmètica [Lab01]. Darkness [Beb96]. Darmstadt [Ano81g, Lüt09a, Mai87, Spa89]. Darstellung [Lau92]. d’as-samaw’al [AR76]. dash [Gab79]. Daston [Sch91a]. Data [Sid06b, SI19, Syl83, ET03, HF84, TE03]. Date [Hed99]. datis [Gra83]. datta [Gup80, Jon76c]. Dauben [Ano88g, Haw81, Rid88, Van06]. daughter [Hym78]. daughters [Cra03]. d’Augustin [GG76a]. Daumas [Ano74z]. D’Autolycos [Ano95n]. Davenport [Ano84n]. David [Ano88n, Dur21, Ewi92, Hog03, Lei79, San03, SG09, Vol77, Wil77b, Wue05, Zun02, Ano81y, Ano88c, Ano88b, Ano89a, Ano89b, Ano92c, Ano92d, Cer04, Cor04, Don98, ESH^{+}13, HEH^{+}04, Plo07, Row06b, Wue05, Daw16b, Row07, Sau06a]. Davide [Lüt20]. Davis [Hil83, Van80]. Dawn [Ale10, Gra11, VF22]. Day [Ano91k, Cho86]. days [Gia90, Pie86, Roe89]. DC [All09, And07, Ayo06, Cal08, GG08, Kat07, Thi06, Thi08]. DDR [Bar78b, Vol81, Tob81, Str78]. Dear [Ang79a, Ang79a]. Death [Dyc75a, FG90, Ano78a, Ano83e, Fel83, Sau10]. débat [Pan88]. Debate [Lüt20, Axw18, Ben95, Man89, Rad03]. Debates [Gra84a]. Debever [Ano84f]. déblais [RL16]. Deborah [McK82]. Debus [Clu76]. débuts [De 20, Gis99b]. Dec [Ano74-51, Tat86]. decade [Bla76]. decades [BC04, Str82a]. December [Ano76w, Ano87-34, Ano90a, Ano92-49, DGL^{+}15, Jam97, PP91, Ano86a, Ano24h, Gar92]. décembre [Ano89t, Ano90k, Mar90a]. Decimal [Bul09a, Van24, Ano76y]. decline [Cri88]. Decoding [Bel05]. decomposing [Abd08]. Découverte [AD87, Ano95n, Gra99a]. Dedalo [Isr85, Ken76b, Ysh90]. Dedekind [Ano79m, Fer98a, Haf16, Lê21, Sch14, SS89, Ano79x, Dru86, Dug76, Fer93, Gan85, Haf17]. d’Edition [Sal87, Ber87]. d’Edo [Ebe96]. Dedomena [ET03, TE03]. d’Édouard [Déc14]. Dedron [Jon76a]. deduction [And83]. deductive [Ung85]. deduttivo [Ken78b]. Dee [Clu76]. Defending [Gou10, Zep12]. Defining [Par04]. definirte [Osw17]. Definition [Zor96, Gan85, Auj86]. Definitionen [Rom04]. Definitions [Bo24]. definizione [Gan85]. deflection [Pou07]. degli [Coz12, Tra10]. degree [LY80]. degrees [Ano74e]. Dehn [Ano78j]. deistvitel’nogo [Smi79b]. Dekker [Cas77, Row83b]. Del [Rem12, Taz09, Amb08, Ano85f, Ano91k, Arc17, Bab77, Ben95, Coz12, DN12, Fol14, Fri21, Hug05, Oak12, Pan18, Par97, Ser90, Tra10, Vér01]. Delambre [Enr76, Ano76-35, Wan22]. Delaware [Neu86]. Délégation [Ano87i]. Delhi [Ano74m, Ano88k, Ano88m, Knu11, Tee84, Yan83, Gup77a]. Delhi/Varanasi/Patna [Yan83]. dell [Van80]. Della [Gra18, Ano81r, Ano83c, Ano90f, Arc99, DN12, DF10, Fol14, Gal12, GG05b, Hug05, Pan18, Rab08, Rem12, Roe89, Ser06]. dell’algebra [FR83, Ano81f]. dell’arti [Coz12, Tra10]. dell’Età [Mal06a]. dell’Europa [Amb08]. dell’ottocento [Arc99]. Delone [D’A10]. demi [Cha90]. demi-siècle [Cha90]. Demidov [Ano95f]. d’Émile [Maz21]. Democratic [Ano90j, Seg76, SS12b]. Democritus [Hof87]. Demography [Bel08, Sch06b]. Demokratischen [Bos77]. Demokrit [Hof87]. démonstration [DP88]. démonstrations [Pet74]. Demotic [Ano74s]. Dénes [WM14]. Denis [KS13b, Ano87i, Taz15]. Denise [Ano93f]. Denmark [Ano78y, Buc80, Gra87, Sør14]. Denominate [Kar22]. density [Von93].

Department [Coc98, GG05a, Pri80, Ano79k, Ano84l, ABC^{+}88, Roy17]. Départments [Ano90-34, Pri80]. d’Epistémologie [Ano90q, Bor89]. depuis [Enr76]. d’équations [Ple78]. Derek [Gui09a, Gui09b]. derivation [Kic10, Phi79b, Wag79]. derivative [Gra78a]. derive [Fu91]. Desargues [Ano92r, Ano90h, Deb13, Fio01]. Desarguian [Cer04]. Descartes [MP06, Mar10a, Bos10, For77, IN89, Len79, MA10, Man06, Mar10c, Mol76, Rab10, SD08, Ste12, War08, Sen84]. descendants [Cri04]. Descent [PS97]. Descotes [Mar10a]. described [Ano80-31, HZ85]. description [Hug87]. descriptive [d’E01]. design [Sti74]. Desmond [Ano87e, Was20]. Despeaux [Ano05a, Ano05b, Ano05c, Ano05d, Ano06b, Ano06c, Ano06d]. destin [Bor14, PR12]. details [Nau03]. d’été [Bor89, Dho91, Ano93-33]. determination [BF23]. determined [Che09]. determining [Coo11, Tat87, Zho21]. déterminisme [Ano91q]. Detlef [Wol03]. detta [Ano90f]. d’étude [Pei17]. d’Etudes [Yus10]. d’Eugenio [Gra99a]. DEUGSSM [Kno88a, Tat86]. d’Euler [HS15]. Deutsch [Ano85h, Scr03]. Deutsche [Ano87m, Ano91c, Ano92h, Ano93k]. deutschem [Jay83]. Deutschen [Ano92-46, Bos77, SS89, Ano88i, Bie74b, SS08, Tob06]. Deutscher [Lem74, Seg76]. Deutsches [Sau14]. Deutschland [Ano94b, Has10b, Has10a]. deux [Ano89t, Gau13, Mar93b]. Deuxième [Mar90a]. Developing [FP03]. Development [Ano76-28, Ano86a, Ano92u, AB84, Glu12, GM04, San09, Ai04, Ano80-31, Ano87-29, Dau75, Daw16a, Dru91, Fer08, GG75c, GG77, Ine94, Kno83a, Mai21, Mik75, MB20, Nic93, Phi79c, Rei76, Ros77, Sag91, Sch79, Sch77, Sti76b, Tur81, Zyg75, dS90, Ano93h, Ano77o, Ano88g, Yok75, Ano74u, Ano74t, Dav76, Gra96, Seg76]. developments [Lan75]. Développement [Bre07, Ai04, Ano92q]. Développements [May76]. devices [Kid09]. devoted [Ano80-34, Ano90v]. Devreese [Ste09]. Dewulf [Aïs96]. DF [Aus16]. Dfl [Ano84i, Bow84, Hog03]. Dhahran [Kin86b]. D’histoire [Ano83o, Ano90q, Ano90-34, Ano92k, Bor89, Dho91, Hou89, Kno88a, Mah78, Mal97, Ano74-33, Ano75f, Ano77j, Ano78k, Ano80d, Ano85p, Ano88v, Ano91r, BKM^{+}82, Bun76b, Cha81, Cha87, dMC86a, Dho88b, Gau13, Tat86, Ano74-32, Ano80-32, Jon77a, Gis96]. Dhombres [Ano89g]. Diagne [Ric91]. Diagrams [De 05, De 09, Car18, Edw04, Gil77, Kel05, Rey09, Abe06]. dialectic [GG80]. Dialectics [Vuc99, Vuc00, Vuc02]. Dialogue [Gup91, dW10]. Diamond [Ano81u]. Diana [Ano74x]. diary [CM22, Joh82, Joh86, AB09, Per79]. diastema [Sid04]. dibattito [Ben95]. Dichters [Bos03]. Dick [Kat07, Row83a]. Dickson [Fen97]. dictionary [Ano92-38, Tob06, Zer75]. Did [Coo11, FT99, GG96, GG97, Lai75, ARR12, Fu91, Gil81, Hof87, Spa11]. Didaktik [Ano84k, Toe96, Ano81g]. didattico [Mai08, Sch13c]. Diderot [Gil75]. Diego [Ewi92, Lew90, McK82]. Diepholz [Has04, Jur07]. Dieter [Fol87]. Dieudonné [Ano79m, Ano84g, BKM^{+}82, Gra91a, Gra01, Gre85]. difference [Mia08]. Differences [Hay97]. Different [Zhu16, Sai19]. Differential [Taz97, ABBS09, Eng80, FK90, Fer04a, Gra91a, Gro87, Haw77b, Kat81, Ken77, Nik18, Por82, Rei76, Sac20, Sch77, Ste20, Zun83, Jaë19, Boy74]. Differentialgeometrie [Syl03]. Differentialkalkül [Ano95h]. Differentials [Fer04a, Bot78]. differentiation [Ano80g]. différentiel [Jaë19, Pan88]. différentielle [BKM^{+}82]. différentielles [Krö17, Tou03, Fer11]. Differenzen [Bla05]. differenziali [Bot78]. difficult [Mat14]. Diffusion [Ano74x, Mal97, Yca16, Mal97]. Dig [Kat11]. Digging [Ung85]. Digital [Asp83, Cre16, Knu74]. Dijksterhuis [Atz06]. dimensio [DN12, Fol14]. dimension [Dau75, PB08, Van83b]. Dimensional [Aub02, Gra80a]. dimenticata [Net02]. diminutionis [Høy24b]. Din [SI19]. d’Infinis [Ano92a]. dings [Ped75a]. dintorni [Taz09]. Dio [Ano92s, Gan86]. Diocles [Ano77i]. Diofant [Git74]. diofantovy [Git74]. Diophante [Chr91, Chr98, Ber87, Sal87]. Diophantos [Tho11]. Diophantu [Hog85b]. Diophantus [CM19a, Ace11, Bas81, Chr98, Chr07, CO13, CM19b, Kno93b, Wat93]. dioptrique [Sim96]. diorism [SS10a]. diplomacy [Dau80]. diplomatic [Shu80]. Direction [Atz04, Hog03, BKM^{+}82, Bun76b, GG76a]. Directory [Ano85r, Ano95-27, Ano79-48, Ano80-38, Ano81-45, Ano83i, Ano84m, Ano84z, Ano92-51]. Dirichlet [Fis94, Sch86b, GG77]. dirigée [Høy07]. Diritto [Oak12]. Dirk [Ken82, Lem74, Ser18, vM84, Alb94, Ano76-44, Bos02, Row94]. Disciplinary [Ano85-29]. Disciplines [AD02]. Disciplining [Sch06b, Bel08]. Discours [Fre77]. Discourse [Bee06, Ste02]. Discovered [Bel02, EU99, CWW17, GG78, Moy19, Neu88, Roz75]. discoveries [Kil83, Moo75a]. Discovery [Eli24, Hog86, Bal23, Dra74, Fer95, RT81, Row06b, Wue05, Sch88, Ano81l]. discret [Bré12]. discussion [Gar75, Gil77, Meh76, Ano87n]. Discussions [Lok80, Bul09b]. disease [Fri91]. disem [Ped75a]. d’Islam [Roz90]. Disquisitiones [VJRZ87]. dissemination [Hee14, Sco16]. Dissertation [Ano74-36, Ano79t, Ano79v, Ano79u, Ano79w, Ano79x, Ano80f, Ano80g, Ano80e, Ano81y, Ano82g, Ano83g, Vol91, Ano81z, Jen76, Van79, Vol06]. Dissertations [Ano74-37, Ano80h]. Distance [Hog03, Ken84]. Distances [Sid24]. distinguishing [GG04]. Distributed [Ple78, Jam13]. distribution [Ano74-40, Joh84]. Distributions [Syn83, Ano84j, BPL17]. distributivity [Cor16]. distributivity-like [Cor16]. distributor [BT81]. Disturbed’ [Tra21]. Disziplinen [Bor18]. divergence [Høy04, Sto78]. divergence-canons [Høy04]. divergent [BL76]. diversi [Scr03]. divine [Sho87]. Division [May75c, May77c, Scr82b, Høy82, Høy90a, Scr84]. divulgación [Ano87r]. divulgazione [GG05b]. Djafari [Ano85e]. Djebbar [Høy07]. Dkr [Ano87d]. DM [Ano74q, Ano81i, Ano85e, Ano88n, Ano89f, Ano89k, Gol76, Gra90, Grö86, Hes90, Lem74, Par92a, Ano89j, Ano95m, Bir80, Jay83, Pys81, Sch90b]. DM.80 [Mat75]. DMV [Nab04]. DMV-Seminar [Nab04]. do [Ano77l, Gil81, Ken76d, May75d, Sti76b, Zhu20]. Dobrovolskii [Rei76]. Docent [AH16, Glu12]. doctor [Ano80q]. Doctoral [Ano74-38, Ano81z, Gup76b, Gup93]. Doctrine [McK82, Tat11, BS09, Bel07b, Mah78]. doctrines [Bun76b]. Documents [Gar01]. Dodgson [Dru07, Dru12, Ped75b, Abe79, Abe93, Abe01, Abe10, Kat20]. does [Scr81]. Dohse [Thi06]. DOI [Gur24]. Dokumente [SS89]. Dolciani [Enr85b, Enr85a]. Dold [Hog15, Van06]. Dold-Samplonius [Hog15, Van06]. dollars [Lor18]. Domaine [Ano90-35]. Dominance [SS97b]. Domingues [Bra09a]. Domínguez [Gar98]. Dominique [DR03, Har06, Krö15, Mar10a]. Donald [Ano88l, Dav19, Row93]. Done [Ber97, Ped75a]. Donostia [Ano91p]. Donostia-San [Ano91p]. d’Or [Cav95]. d’Oran [Mar90b]. Dordrecht [Ang88, Ano77r, Ano84i, Atz06, Bel12, Bow84, Cog12, Fol15, Jah12, Pin80, Sau06a, Sch12, Sid06a, Zep12, vM84, vdW76a]. Dordrecht-Holland [Ano77r, vdW76a]. Dordrecht-Holland/ [vdW76a]. Dordrecht-Holland/Boston [Ano77r]. Dordrecht/Boston [Bow84, Pin80]. Dordrecht/Boston/London [Ano84i, Sau06a, vM84]. Doris [Hym78]. Dorothy [Par89]. Dorrance [Zas83b]. d’Orsay [Ano90-34]. double [Scr83]. Doubling [Sai95]. Doubts [Ben95]. Douglas [Ano95g, Dav90]. douloureux [Bor14, PR12]. Douwes [Tat87]. Dover [Ped75b]. D’Ovidio [Taz09]. Down [Ano82f]. Dozenten [Cal75, Seg90]. Dr [Ano79o, Ano79-42, Ano86h, Dho11, Don08]. Dr. [Nah06, Lan07]. Draft [Asp83]. Drake [Ano78h]. dramatic [BG10]. Drawing [Hun96]. drawings [Ano78i, Ano88h, Ano93f]. drevneishikh [Ano77l, Ken76d]. drevnem [Vol91]. Dritte [Bar78b]. Droz [Ano77q, Gol82]. Drucker [Sca94]. Dry [Duc16]. Dualität [GS91]. Dubbey [Ano74u, Enr80]. Dubbiezze [Ben95]. Dublin [All04, Ano79i]. Duchesneau [Vil97]. Duckworth [Ang79a]. Duda [Mro16]. Dudley [Bie74a]. due [Lo 84]. Duel [Ale10, Gra11]. Dugac [Ano79m, Ano81j, Ano81n, Ano87i, BKM^{+}82, Kit81, Kno88a, Maw06, Tat86]. Duke [Bel08]. Dulaurens [Ste12]. d’Ulm [Dur21]. Dumbaugh [Gra18]. Dumézil [Yus10]. Dummett [Ano89c]. Dún [Ano87e, AD87, Ano77n, Auv13, Bré12, Gis14, Høy07]. Duncan [Gin83]. Duncker [Row85d]. d’une [Ano86g, DP88, Gra12, Lê21]. Dunham [Dav19]. Dunnington [Thi06]. duo [Rom13]. duplication [LGG74]. Dupuis [McK82]. Durand [Maz16, Pie89]. durch [Nag88, Osw17, Row06a, Str78, Wue05]. Durée [AD02]. Duren [Coo96, Sel91]. Dürer [Ped75a, Sch99b, Wei04, Ano84e]. Durham [Ano87-34, Bel08, GG75b]. during [Ano86h, Ano87-29, Ano91-31, Cas80, Fuk87, Kas03, Pep90, Pep92, Roy17, SS93]. Dutch [Bec99, Bec01]. Duvillard [Maz16]. Dworsky [Ano84i]. dyadik [Ano77h]. Dyck [Row08]. Dynamical [AD02]. Dynamics [Gam96, Gui96, Ale95, Bra89, Tru75]. dynamique [Vil96, Vil97]. Dýnamis [Fow92b, Høy90c].

E-Book [Høy24b, Bor18]. e-ISBN [Bor18]. E. [But81, Gre77, KK90]. each [Ewi92]. Eames [Tro74]. Earlier [Hed99, Bec19]. earliest [Her75b, Hug74, Phi79b]. Early [Ano92-30, Ano95l, Cal08, Coc98, Cog14, Cra99, Cra00, Cri24a, Eva98, Eva13, Hod77, Hod75, Kje93, Kje17, Lüt08, Mor19, Pyc82, San07, San09, Wea77, Zit07a, vdW76a, Ano82c, Ano84c, Bat06, Ben95, Blå12, CM19a, CM19b, Cre14, Cre16, Cri04, Cul07, Dal85, Dau75, Eng80, Fer08, Fol83a, FT83, Glu05, GG93, GA20, HF19, Høy82, Hug87, Hug10, Jay24a, Jay24b, JET84, KK90, Mai21, MS05, Mas08, Mes94, Pan03, Pow76, Rei86, Rey06, Rey09, RW03, Ros76, Sai95, Sho83, Ste20, Sza74, Tho76, Tim04, Tur81, Van09, Wag79, vM07]. Earth [Eva13, Lew98, Gol84, Van09, Mac77]. East [Ano76-28, Hog24, Ken75, San06, Hai86, Bus76]. Easter [Ano76v, Rog74, Ber90, Lic97]. Eastern [BLF92]. Eberhard [Ano84e, Ano92l, Bar82a, Bar82b, Don05]. eccentricity [Car22]. échangées [Vol02]. Echeverria [Ano93i]. Eckert [Asp83, Jur07, Sau14]. Eckes [Cog15]. éclairages [Déc14]. eclipse [Moz13]. École [Ano87i, Ano89g, Dho91, Dho11, Gra08, Gre85, GW94a, GW94b, Hei81, Taz15]. Economia [Oak12]. Economic [Ano92n]. Economics [GM04, Lüt19]. écrit [AD87]. écrites [Ano78i]. Ed [All04, All09, And07, Ber07, Bor14, Cog15, Dru07, Dum15, Eva94, Fen08, Fol09a, Fol15, Gil15, GW94b, GG98, GG05b, GG13, Gra98b, Hei09, Høy09a, Høy09b, Kle17, Lüt15, Maz21, Pec20, Rec07, Sch07, Scr07, Sel91, Ser18, Sid06a, SS07, Taz15, Yus10, Ano81i, Ene18, Seg90, Ser90, vM84]. Edgerton [Jon77c]. Edicão [Ges18]. Ediciones [Gar98]. Ediderunt [Pei22]. ediert [Rom04]. edificio [Ano92-44]. Edinburgh [Ano88l, Mai21]. Edited [Ang88, Ano74m, Ano74q, Ano74v, Ano74-29, Ano77b, Ano77c, Ano77i, Ano77l, Ano77s, Ano77p, Ano77f, Ano78-32, Ano79b, Ano81b, Ano81c, Ano81s, Ano81k, Ano81p, Ano81i, Ano81l, Ano81o, Ano81q, Ano81u, Ano82c, Ano84a, Ano84d, Ano85b, Ano85a, Ano86b, Ano86e, Ano87b, Ano87a, Ano87i, Ano88c, Ano88b, Ano88g, Ano88h, Ano89a, Ano89b, Ano89l, Ano91d, Ano92c, Ano92d, Ano93i, Ano95f, Ano95k, Ano03a, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c, Ano04d, Ano05a, Ano05b, Ano05c, Ano05d, Ano06b, Ano06c, Ano06d, Ayo06, Bar77, Bar78b, BKM^{+}82, Ber87, Bid74, Cas77, Cas80, Coo95, Coo96, Coo99, Dav90, Don05, Dru12, DR03, Edw77b, Eis75, Ewi92, Fol12, Fra04a, Fra10, Fre92, Gal12, Giu99, Gra77, Gra10a]. Edited [GG05a, GG76a, Gra96, Gra97a, Gra99b, Gre85, Hal84, Her14, Hir74, Høy95, Hug05, Kat11, Ken76a, Ken76c, Kit78, Kla84, Knu11, Knu74, Law98, Lef10, Lew99, Lo 84, Mah78, Mar10a, Mat74, Moo77, Moo82b, Nab04, Nak03, Par97, Pet82, Pys81, Ric98, Rid88, Rom99, Row83b, Row93, Sai74, Sal87, Sca94, Sch99a, Sch11, Sch13b, Sch89, Sch00b, Seg76, Ser90, SS97a, SS17, Str82a, Syl81, Syl03, Syl07, Tat11, Ter89, Toe96, Too80, Tro83, Van06, Vit96, Wil83, Zas83b, Ale04, Ano74z, Ano82d, Ano84f, Ano92l, Cib76, Gab76, Goo77, Ken76b, McK82, Moo03, Pla76]. Edith [Cri24a]. Editing [Cre16]. Edition [Ano80i, Ano81m, Bay11, Gab76, Gab79, GG13, Her75a, Mac76, Sim00a, Sim00b, Ste80, Tot79, Ull03, Vil96, ALB10, Ano74u, Ano74v, Ano92-51, Enr76, Gra83, GG86, Gra08, McK82, Pep75, Sen99, Sen06, Shu92, Tee84, Thi06, VJRZ87, Ano77j, Bie74a]. Editions [Bra09b, DR03, Fra04a, GW94a, Gol00, Har06, Hog02b, Lüt15, Maz16, Giu86, Neu12, Yus95, Ano77k, Cav95, Dur21, Gra08, Gre85, Krö15, Lor16, Maz24, Wil08b, Hei81, Kle17, Row15]. Editor [GW94a, GG86, GG90b, Wil83, Dau79, Hal90, Par95, Ano02f, Dau82a, Fra04b, Gui99]. Editor-in-chief [GG90b]. Editore [Mar14, Net24, Pan09, Sch13c, Van80]. Editores [Aus16, Ric98]. Editorial [Ano76p, Ano76q, Ano77v, Ano09a, Ano10b, Ano10c, Ano10d, Ano10e, Ano11b, Ano11c, Ano11d, Ano11e, Ano12a, Ano12b, Ano12c, Ano12d, Ano18a, Ano18b, AG13, AG14, AG15, ASS16, BGG10, Dau83, Dau84, Dau85a, KP86, MS22, MO23, RP94, SCM19, SSS17, SSS20, vDBG07, Ene18, Gis18, Gra91b, Wol03, Ano78m, Ano03i, Ano03j, Ano03k, Ano03l, Ano04g, Ano04h, Ano04i, Ano04j, Ano05i, Ano05j, Ano05k, Ano05l, Ano06e, Ano06f, Ano06g, Ano06h, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09b, Ano09c, Ano09d, Ano13a, Ano13b, Ano13c, Ano14a, Ano14b, Ano15a, Ano15b, Ano15c, Ano15d, Ano16a, Ano16b, Ano16c, Ano16d, Ano17a, Ano17b, Ano17c, Ano18c, Ano18d, Ano19a, Ano19b, Ano19c, Ano19d, Ano20a, Ano20b]. Editorial [Ano20c, Ano20d, Ano20i, Ano21b, Ano21c, Ano21d, Ano21e, Ano22a, Ano22b, Ano22c, Ano22d, Ano23a, Ano23b, Ano23c, Ano23d, Ano24b, Ano24c, Ano24d, Ano24e]. éditoriale [Gis18]. Editors [Fra10, Gra94b, KP85, Kno17, Neu12, vD09, PH99]. Editrice [Ano85g]. Editura [Ano74p]. edizione [DN12, Fol14]. Edizioni [Gal12, Hug07, Isr85, Mal06a, Oak12, Ysh90]. Edmonton [Ano75k, Høy95]. Edmund [CB21, Mai21, Wef89]. edn [Fer04b]. Edo [Fuk87]. Edoardo [Fra06, Law98]. Edouard [Sch76a]. Edouardo [Bul24]. Eds [Atz04, Bla05, Bra07, Bul22, Daw16b, Don08, Dum18, Fol14, Fra06, Gar05, GG08, Gra94c, How18, Høy23, Jam13, Jur07, Kat07, Kat20, Krö15, Lor16, Lüt08, Maa07, Mal06a, Maw06, Mos16, Moy17, Neu08, Plo06b, Row07, Sch19, Taz09, Thi08, Tur08, Van05, Wil08b, Cri24a, Lê24]. Eduard [Edw77b, Ter89]. Eduardo [Tur08]. Educacion [Ano90-32]. Education [Ano74-39, Ano75l, Ano76r, Ano76s, Ano76t, Ano77w, Ano77x, Ano77y, Ano78n, Ano78o, Ano78p, Ano81-27, Ano82h, Ano86r, GG90c, Kat11, Mic88, Rog77, Sai74, Shi93, Swe92, Ano74-60, Ano80-31, Del03b, Jan11, Kid09, Pre24, Sco16, Swe74, Yok75, Zas79, Ano80k, Ano95k, Ano81q]. Educational [GG92a, Tan74, GG94, Mai08, Sch13c, Del03b]. EDVAC [Asp83]. Edward [Ano74l, Ano84d, Cal08, GG75b, GG86, Jah12, Moo82b, Pye76, Tim09, Cra03, vD10]. Edwards [Abe06, Spi76]. Edwin [Bay11]. Egbert [Plo06b, Thi06]. Egham [Ano74-46]. Egypt [Che12, Eng77, Kno82, Mel04, RT02]. Égypte [Cav95]. Egyptian [Abd08, Ano74-40, Coo11, De 09, Ger85, Gil79, Gil81, HP20, Imh03a, RS85]. Egyptians [Lum80]. Égyptiennes [Cav95]. Egyptological [Ano74s]. eight [Wil08c, Cam77]. Eighteenth [AH16, Ano74-27, Ano74z, Lew98, Pye76, Che17, DF21, Sor13, Yca16, FP03, Jam88, Lok80]. Eighteenth-Century [Ano74-27, Lew98, Che17, FP03, Lok80]. Eighth [Rid94]. Eightieth [DPS^{+}02, BDG^{+}86]. Eighty [Sch92, Ano76-44]. eighty-two [Ano76-44]. Eiji [Sid12]. einem [Fel78, Lem74]. Einfluss [Ano88i]. Einführung [Ull03, Toe96]. Eingeleitet [Rom04, Bul22]. einige [Osw17, Fri91]. einmal [Ber90]. einschliesslich [HB^{+}13]. Einstein [Ano84f, Ano88f, Bla76, Dyc75a, Row06a, Wue05, Zit07a, Ano75w, Ano90k, Bat06, Row06b, Sau06b, Wue05]. Einstein. [Dyc75a]. Eisele [Ano77s, Ket82, Pys81]. Eisenstein [Pat90]. Elaine [Nic13]. elasticity [Ano84i]. Eleanor [Gra10a, Por05]. electromagnetism [Fle86a]. elem [Rab74]. element [Wil80]. Elementa [Rom04]. Elementarmathematik [Ano81i]. Elementary [Zor96, Ano78g, CV22, Wat91, GC21]. Elemente [Bus76, Mon75, Ume76]. Elementen [Lo 85]. Elements [Jon77b, Ano77s, Ano90u, Lo 85, Too80, Ung85, Mar93b, Clu76, Ano80h, Ano87d, BV05, Bur08, Bus76, DZ97, De 84, Eis78, Eli18, Eli24, Fow92a, Gol00, GG96, GG97, Hog87, Lo 83, Mar97, Sai19, Ung85, Xu05, dY92, vdW76a, Gra12, Mue96, AG77]. Elena [Ric98]. Eleventh [Van10, Sin85b]. Eleventh-century [Van10, Sin85b]. Elfving [Sel90]. Elie [Ano81l, Fre77, Ano84f, Haw96]. elimination [Grc11]. Elizabeth [Bos10]. Ellen [Gra13]. Ellipse [Erl97]. elliptiques [BKM^{+}82]. Ellis [Ver13]. Ellison [BKM^{+}82]. Else [Buc80]. Elsevier [Ano95j, SS07, Vek75]. Elsvier [Ano74-27]. elusive [Cra13]. Elwin [Str81]. Elzinga [Tho76]. Emancipation [SS97b, SS93]. Emanzipation [Row85d, Ano95m]. ème [Ano92a]. Emergence [Coc98, Ano82f, Gar85, Kno80, Mar10b, Moo08]. Emerging [Kat09]. Emigrant [Sør14]. Emigration [Nos12]. Emil [Fen08, AR07, Art13, FS10, Dum15, Gra18]. Emile [Maw06, Gra76b, LBB^{+}04, Maz21]. Emilio [Rab08]. Emmanuel [Gol82, Maz16]. Emmanuel-Etienne [Maz16]. Emmel [Bul22]. Emmy [Jen86, Ano74k, Cho86, Jen86, ST16, Row83a, Row83b, Sch17]. Emmylou [Lê21]. emphasis [Buc80]. Empire [BS13]. empiricism [Lüt09a, Pul05]. Empirie [Lüt09a, Pul05]. Enciclopedia [Ano90f]. Encyclopedia [GG98, Gra98b, Kat97]. encyclopédique [Gis99b]. end [BS13]. Endemann [Ken76c]. Endowment [Ano79k, Ano90u]. Enfants [Ano92p]. Enfield [Ano74-56, Ano74-59]. engagements [Auv13, Gis14]. Engel [Haw76]. Engfer [Ano88i]. Engine [For75]. engineer [BG21b]. engineering [Buc80]. engineers [Ano07e, Vér01]. England [Ano74-46, Bel08, Eve76, GG87b, GG92b, Hal84, Kil83, Rog74, Ano74-56, Ano74-59, Ano76v, Ano78l, Ano79r, Sch06b, Ste01, Lew90, Tan74]. English [Ano74u, Ano77i, Ano84f, Ano88m, Bee06, Mon17, Ste80, Tee84, Ano74-41, Bur92, Gol00, Row85e, Ste02, Zer75]. Engraving [Sch99b, Wei04]. Eniac [Asp83, Ano79t]. Enigmatic [Sch99b]. enlightenment [Sch91a]. Enrico [Sai94, Bot77]. Enriques [Sch99a, IM98, LD22]. ENS [Row15]. Entanglements [Sch05, Pyc97]. Entdeckung [Row06a, Sch88, Wue05]. enter [Zit07b]. Entirely [Mat74]. Entirety [Del02, Del03a]. entities [Sam24]. Entre [DR03, Har06, Ano89t, Ano92q, Ano92-50, Ano92-28, Ano07e, Fla03, Gau13, Sau09, Vol02, Law98]. entreprise [Gis99b]. entry [Joh86, Wat91]. Entstechung [Grö86]. Entstehung [Sch02, Tob90, Ano88n, Bir80]. Entwickelung [Bos77]. Entwicklung [Ano88i, Ano95h, Bor18, Ine94, Kno83a, SS88]. enunciations [OA05]. Episode [Ful99, Gui96, BG10, Gui87]. Episodes [GP07, Ano89e, Neu08]. Epistemological [Ano93i, Tur81, Car07, Ano82c]. Epistémologie [Ano89m, Ano90g, Ano92a, Boi89]. Epistemology [McL12]. epistolicum [Ano77m, Ter89, Pei22]. epokhi [Sti76b]. époque [Ano92p, Ebe96]. Epple [Plo06b, Sch02]. épreuve [PP91]. Épures [d’E01]. equality [Tho11]. equation [Ano89-35, CWW17, Pat90, Sac20]. Equations [Bir96, Fri81a, LMP01, LYK89, Boy74, Cer17, Cra13, Eng80, Gro87, Hog89, Kic15, LY80, Lê15, Lüt82, Nik18, OA07, Phi79b, Row06b, Sac20, Sau06b, Sch77, Wue05, Fer11, Tou03, Rei76]. Équatoires [Gol82]. equazioni [Ano81r]. equiangular [Erl92, Wil94]. équilibre [Ano81m]. equilibrium [Ano92n]. équipe [Ano89n]. equitable [Ano74-40]. Era [Vuc00, Vuc02]. Eratosthenes [Gol84]. ergodic [Zun02]. Erhabene [Gra10b]. Erhard [Che84, GS91, Lê24, Nab04, Plo06b, Toe96, Fel85]. Eric [Ano74-27, Ano95i, HP20, Sch00b, Fow84, Wil91]. Erich [Neu93, Neu93]. Erlanger [Haw84, Lê15, Row83c, Row85e]. Erlichson [Wil94]. Ernst [Moo75b, Moo76, Pee95, Edw77b]. eroi [Arc17]. Erone [Coz12, Tra10]. Errata [Ano74-42, Ano74-43, Bar78a]. Erratum [Ano75m, Ano80j, Ano81-28, Ano81-29, Ano82i, Ano87o, Ano88p, Ano88q, Ano91i, Ano93l, Goe09, GG03, Has10b]. Error [Plo01, Cam77, Kno93c]. errors [And11, Bel15, GG77]. erschienenen [HB^{+}13]. ersten [Böl93]. Erster [Bar78b, Hes90]. erweiterte [Lem74]. Erwin [Dho11, Don08, Høy24a, SS08, Ull03, Stu80]. Es-Senia [Mar90b]. Escritos [Aus16, dL01]. Escuela [Ano91p]. espace [Gra12]. Espagnole [dMC86a]. España [Ano90-32, Ric98]. Esprit [DP88]. Essai [Cor96]. Essay [Ano75e, Con06, DS99c, GG13, Høy24a, Ano74-30, Ano84i, Pes03, SS12b, SS22b, Tat75, Wis76, Imh21]. Essays [Ano88g, Oli08, Pys81, Smi79b, Ano92k, Cra13, FG06, BTB^{+}03, Fra06, Rei76]. Essen [Rom04, Kno88b]. Essence [Knu11, òP09]. Essential [IM98]. essor [Hei81]. est [Net24]. Establishment [Lew78d, BK84]. Estate [Sch98, HB^{+}13]. estestvoispytatel [Dyc75b]. Esther [Sch89]. estimate [Buc23]. estimates [MB20]. Estonian [Dem74]. Estudios [Gar98, Ano85j]. Estudos [Ges18]. etá [Ano93j]. étage [Ano90-34]. Etas [For75]. État [Bor14, Enr76]. etc [Høy90b]. etc. [Enr76]. Ethel [Zas83a]. Ethnomathematics [Ano89-32, Asc95, Asc97, Asc88, Par92b]. Etienne [Maz16, Alf11, Sar15, Gol13]. ETS [Hug07, Mal06a]. Ettore [Ano81-30]. étude [Ebe96]. Etudes [Ano87h, Che94, Ple78]. Euclid [Ano87d, Ano90u, Bur08, Clu76, Lo 85, Too80, Ung85, Vit96, Ano80h, BV05, Bus76, DZ97, De 84, De 15, Dod13, Eli18, Eli24, ET03, Fis79, Fow92a, FT99, Gol00, GG96, GG97, Hen75, HF84, Hog87, Lo 83, Mar97, Sai19, SI19, TE03, The84, Tra21, Wag83, Xu05, dY92, Syl83, Ped75b, Sid06b]. Euclide [Gal12, Kno07, Lo 85, Auj86, Mue96]. Euclidean [Ano74p, Bo24, Gra91b, Sch95, vdW76a, De 95, De 05, Gra79, Haw80, Pej07, Sch79, Sha94, War24, Zor96]. Euclides [Sai94]. euclidienne [Gra99a, Mar93b]. euclidiennes [Roz90]. Euclids [Rom04]. eudoxischen [Mue80]. Eudoxos [Mue80]. Eugene [Don98, Fen97, Aïs96, Gis99a]. Eugenio [BR16, Ros17, Vac20]. Eukleides [Høy90b]. euklidische [Gra78d]. euklidischen [Ano89k]. Euklids [Bus76]. Euler [Ano74-27, Ano77m, Ano90q, Bar77, Bra09b, Cal92, Dho88a, Fri81b, Lan07, Spa11, Ter89, Ano80g, Ano83h, Ano84-28, Ant95, BL76, BMY07a, BMY07b, BDS07, Bur86, Cal96, Cre14, Fel83, Fel07, Fer98b, Fer00, Fer20, Fra92, GG93, Krö08, Müü75, Nah06, Pet12, Pie93, San07, Spa11, Cal08, Scr08, GG08, Thi08]. Euleri [Pei22, Rid83a, Rid83b, Sch00b, Bar77]. Eulerian [Fer04a]. Eulers [Ano81x]. EUR [Bla05, Bor18, BB04, Don05, Fra04a, Ges18, GG05a, Has04, Krö06, Maz21, Pan18, Row23, Syl03, Tat11, Taz09, Bré12, Cog12, Coz12, Dur21, Fol12, Gal12, Gar05, Gau13, Gra12, KS12, Lef10, Mar10a, Oak12, Rem12, Row19, Sch11, Sch12]. EUR. [Che14b, Her14]. Euro [Hog03, Van06, Hog24]. Europa [Bus76]. Europe [Ano92-35, Oak12, Amb08, Ano92-32, Bec01, Bus76, Hai86, Str94, Van10, Yca16, Ano95l, Coo99]. European [Ano81p, Ano91j, Gra18, Kat92, Ric98, Tat11, Ano93m, BJ89b, Ehr18, Fol83c, Pep90, Pep92, War24]. Européenne [Ano93-33]. Eva [Ano74-30]. evaluation [AB84, Pie83]. Evangelos [CK92]. Evans [Ano05a, Ano05b, Ano05c, Ano05d, Ano06b, Ano06c, Ano06d]. Évariste [Rad02, Neu12]. even [Bec19]. événements [Maz16]. event [Ano90e, IN84]. events [Pre24]. Evert [Kno91a]. Eves [Enr85b, Enr85a]. Evesham [Glu12]. Evidence [San05, Ung20, SS88]. Evolution [Gra91b, Moo02, Row90a, Sho02, Ano74f, Che10, Moo07, Syn83, Tur13, Wil79a, vdW76a, Ano78g]. Evora [Dho91]. Ewald [Daw16b, Moo03, RFS89, RFS89]. ex [Net24]. Exact [Ken75, Ano88w, Kin86b, MW17, Vol77]. exacten [Ano89j]. Exactes [Mar90b]. exactness [Pan11]. Exakten [Mat75, Grö86]. examination [GC21, LY79, Sai19]. examinations [GG90c]. examiner [GG90c]. Example [Plo96, Ano80-31, BS13, CMT79, Fio17, FT83, GG93, Ung22]. examples [Lei77]. exams [Ros20]. excavated [Cul07]. exception [Ber90, Bur12]. Exceptions [Sør05]. Excerpta [Mar10c]. Excerpts [Jen86]. Exchange [Ano95l]. exchangeability [Dal85, Sen87]. exchanges [Ehr18, Hog10]. exclusion [Lic97]. Exclusive [BT81]. Executive [Ano85q]. exhaustive [Tee78]. exhibit [Ano77-28]. Exile [AH99, EI12]. existing [Her75b]. Expectation [Sho83, Das80, Sen87]. experience [BG21b, Seg86, Hil83]. experiences [Ano91g]. experimental [Cam76]. experiments [Pie83, Sti74]. explained [Ano78-32]. explanation [Car18, Hod75]. explanatory [Høy04]. Exploration [Hir74]. Explorations [Che14a, Ano93i]. exploratory [BM77]. Exponential [Wüs02]. Exposition [AR07, Fen08]. Expositions [Enr85b, Enr85a]. Expounding [BK06, Wil08a]. expresion [Ano90-32]. expression [Bec19, HF19]. extension [Sto76, Law05]. extensional [GP91]. extensions [Ano74l]. exterior [Glu11]. externalism [Pyc84b]. extraction [Joh11, Zhu16]. extrapolation [Dut84, Eng79, Eng79]. extremal [Fen89]. extreme [Høy90a]. Exzerpte [Rom04]. Eyes [Ale04].

F [Abe06, AB02, Ano77j, Ano77o, Ano81h, Ano89k, Ano89m, Ano90g, Aup76, Bay11, Bie80, Bie83, Boi89, Boy74, Bul09b, Don08, Dru07, Dru12, Edw80, GG77, Hal78, Joh84, Mal06a, MB20, Smi79b, Str78, Taz09, Tot79, Wil77b, Roz90]. F. [PP91, YDD83]. F.R.S [Wee91]. F1 [Ano77o]. Fa [Nee79]. Faà [GG05b]. Fabricius [Car24]. Fabrizio [Net24]. Fabulous [Lan07, Nah06]. Fachmedien [Sch19]. Fachsektion [Tob81, Ano92-46]. Facsimile [Sch76a, Tat75]. Facultad [Ano90-32, Ano92-44]. Faculté [Eva94]. Faculty [Zit07a, Bat06]. fading [Hog90]. fail [Fu91]. fairest [AS93]. faith [Fin11]. Faksimile [Fol09a]. Fall [Bul24, Lew98]. False [Han07]. Falsirol [RFS89]. Familie [Hes90]. family [Roe15]. Famous [Bag97a, BGSS16, Fri81b]. Fang [Bie74a]. fantasmas [Pan08]. Fantastical [Abe09, Wil08c]. Farabi [Ano76-28]. Faraday [GG92b, Fle86a]. farewell [Scr85b]. Farghani [Ber07]. Farisi [AF94, Oak19]. Fates [Sen10, SS09]. Fatou [Gam96, Aud11, Ale14]. Faulhaber [Cif00, Man06]. Fauvel [Gra02, Ste05]. Favorable [San03]. February [Ano76-34, Nag91c, Ano21m, Ano22i, Ano23j, Lew76]. Federal [Ano89-34, Ano90-33, Ano91-33, Ott88, Scr87, Per93]. Federico [Sen84, Taz09]. Federigo [Sch99a, LD22]. Fedor [Zai95]. feedback [Rus75]. Feferman [GG90b, GG05c, Pec20]. Feingold [Bul16a]. Feke [Lon21]. Feldbau [Aud12, Krö14]. Felgner [Plo06b]. Felice [Ano85g, Ken78b, PR22]. Felix [Row85d, Ano77n, Gis99b, Haw84, Hel23, Lê15, Row83c, Row85e, Sch13a, SS22a, Kot79, Mon83, Plo06b, Row85b, Row90a, Row19, Scr07]. Fellmann [Ano77j]. Fellowship [Ano80r, Ano84q, Roe15]. Feltrinelli [Ano81f]. Female [WW80]. fencing [Hog10]. Fennica [Sel90]. Fenny [Jam13]. Ferdinand [Ano74-30]. Fermat [Ano74-28, Lüt15, Mal97, Mar10a, And83, FG12, Fle89, Fle91, IN89, LP10, SD08, Ste12, De 20]. fermented [Fer95]. Fern [BKM^{+}82]. Fernandes [Sil08]. Fernel [Rom13]. Ferney [Pei17]. Ferney-Voltaire [Pei17]. Fernparallelismus [Sau06b]. Ferrand [Mar10a]. Ferraro [San09]. Ferreirós [Gar98, Lê24, Oli08]. Féry [Lüt15]. Festschrift [Ano95f, Don08]. Fettweis [RFS89, RFS89]. feud [Chi22]. Février [Kno88a]. FF [Cav95, DR03, Gol82, Vil97]. FF60 [Mal97]. Fibonacci [FH08, Glu76, Hug10, RT02, Sin85a, Høy09b]. Fichant [Vil96]. fictions [Rab24]. Field [Ano91e, Gra18, Jon76a, Sau06b, Thi97, Gui24, Gup93]. Fields [Ine94, Nic13, Fra09, Tro76, Zhu20, RH11]. Fierz [Ano79j]. Fifteenth [Plo96]. Fifteenth-Century [Plo96]. Fifth [Ano89-34, Bow87, Kno93a, Par05b]. fifty [Bid74]. Figs [Ano74q, Ano95m, Swe79]. Figure [Ber04]. Figures [Dur21, Lor15, Ste05, BT84, FFW00, Gol82, Mar90b, Moo75a, Van14]. filling [Nau03]. film [Ano74w]. Filosofia [Ano90-32, Ano87r]. filosofii [Sti76b]. fin [Mar93b, Tat86]. Fincke [Tra21]. find [Hof87, Tob10]. Finding [Hog03, CG92, Kre03]. Fine [Ros76]. Fingers [Wil97]. finis [Ano81j]. finite [Gan85, Gra80a, Sen87, Wil80]. finite-dimensional [Gra80a]. finito [Gan85]. Fink [Hei09]. Finland [Sel90]. Finnish [Gra10b]. Fiocca [Rem12]. Firenze [Cap13, Coz12, Ken78b, Rom97, Taz09, FR83, Høy06]. First [Ano81p, Ano81-31, Ano83i, Ano86l, Ano93n, Asp83, Cal96, Gar86a, Kas90a, Mar99, Ano82j, BGSS16, Bid74, Bin90, Böl93, BC04, CMT79, Eng80, Gui24, Jan11, Kas03, Sar93, Xu05, Ano88r, Ano91j, CB21]. first-order [Eng80]. Fischer [Ano89j, Ano95k, Bel12]. Fischler [Høy90a]. Fisher [Gri82, Edw92]. Fisiche [Fio17]. fits [Wil08c]. Five [Ano78q, Wil97, CM22]. Flächentheorie [Ano89k]. Flament [DR03, Har06, Krö15]. Flammarion [Ano78i]. Flatland [ALB10, Bay11]. flaw [Wei92]. Fleckenstein [Hes90]. Fleeing [Sen10, SS09]. Fleeting [Mar95]. Flegg [Ano85h]. Flexner [Zit07a, Bat06]. Flood [Por05, Ste05]. Florence [Rem12, Rom05, Del05, FR83, Moy19]. Florensky [For98]. Florica [Ano74p]. Florimond [Mah78]. fluxions [Bla14, Smi80]. Foce [Fra06]. focus [Por19]. Foden [GG90c]. Fokko [Atz06]. Folge [Sau14]. Folgen [Bla05]. Folkerts [Ano95f, Høy23, Høy24a, Mat75, Van06, Rom99]. following [Eli24]. fols [Ray23]. fonctionnelle [BKM^{+}82]. Fonctions [BKM^{+}82, Lê21, YDD83]. fondamental [Pet74]. fondements [Ano79m, Kit81, Pan88, Tat86]. Fontaine [Gre84]. Fontana [KO09]. Fontanas [KO09, Lef10]. fonti [GG05b]. footnote [KKS18]. footsteps [Fra16, Mar95]. Forbes [Ano74-27]. Force [Ano78h, Pou07, Rid94, Wil74]. Ford [Ros74]. foreign [Ano88k]. Forest [BT81, Fol76, Glu11]. foreward [Ols76]. Foreword [Ano89c, Vos75, Ano89l, Che94, Law05]. forewords [Mon83]. Forgotten [Fra17b, Lê15, AD78, Cra12, Llo79, PZ11, Row83c]. Forlag [Lüt08]. Form [Nor07, Gup74, Lüt05]. formal [Des07, Fer07]. formalist [Ton82]. formation [Daw16a]. forme [Rab08, Ser06]. formed [Roe15]. former [Wal94]. Forms [HK98, Ant95, Bre19, Kat81, Ser06]. Formula [Fer98b, Lan07, Nah06, Ber90, Ger85, Krö08, LN94, Lic97]. Formules [Maz16]. Forschung [Ano81g, Vol81]. Forschungen [Ano88i]. Forschungsinstitu [Pur90]. Forschungsinstitut [Ano92b, Art91, Eng81b, Eng81a, Pur86a, Pur88, Pur91]. Forschungsstelle [Vol81]. Forthcoming [Ano95z, Ano96h, Ano96i, Ano96j, Ano96k, Ano97h, Ano97i, Ano97j, Ano98k, Ano98l, Ano98m, Ano99g, Ano00g, Ano01g, Ano02k, Ano02l, Ano02m, Ano02n, Ano79-38, Ano79-39, Ano79-40, Ano80-27, Ano80-28, Ano80-29, Ano80-30, Ano81-41, Ano81-42, Ano81-43, Ano81-44, Ano82u, Ano82v, Ano82w, Ano82x, Ano83r, Ano83s, Ano83t, Ano83u, Ano84t, Ano84u, Ano84v, Ano84w, Ano85w, Ano85x, Ano85y, Ano85z, Ano86x, Ano86y, Ano86z, Ano86-27, Ano87x, Ano87y, Ano87z, Ano88y, Ano88z, Ano88-27, Ano88-28, Ano89z, Ano89-27, Ano89-28, Ano89-29, Ano90e, Ano90-28, Ano90-29, Ano90-30, Ano90-31, Ano91z, Ano91-27, Ano91-28, Ano91-29, Ano92-39, Ano92-40, Ano92-41, Ano92-42, Ano93-29, Ano93-30, Ano93-31, Ano93-32, Ano94s, Ano94t, Ano94u, Ano94v, Ano95w, Ano95x, Ano95y]. forthcoming [Ano03s, Ano03t, Ano03u, Ano03v, Ano04q, Ano04r, Ano04s, Ano04t, Ano05s, Ano05t, Ano05u, Ano05v, Ano06o, Ano06p, Ano06q, Ano06r, Ano07j]. Forti [MG81]. Fortwirken [Mue80]. forza [Bee09]. Fossès [Ano92-27]. Fotoalbum [SS97a]. Found [Bag97b, Del02, Del03a, Del06]. Foundation [AB02, For10, Blå17a, Dru86, Kas03]. foundational [Blå12, Man89, Sac75]. Foundations [Ano77-30, Ano92v, Daw16b, Moo03, Pul98, Rad00, Row07, Tob81, AGSL76, Ano74l, Ano81u, Ano89u, Cer04, Dru86, ESH^{+}13, FP03, Fer04a, Fre80, Gar86b, Gre83, HEH^{+}04, Ken77, Mar93a, Mol76, Pej07, Rai92, Sel06, SS10c, Van79, Ver13]. founded [Ano82o]. founders [Fle86a]. Founding [RWG95, Shi94]. Four [BGG06a, Ine94, lLW13, Hay87, Pie93, Wil76, BLW77]. four-color [Wil76]. Four-Colour [BLW77]. Four-Fields [Ine94]. four-squares [Pie93]. Fourcy [Ano89g]. Fourier [Bru77, Her75a, GG90d, SS88, Zyg75]. Fourieranalysis [SS88]. Fourteenth [Ano92v]. Fourth [Ano88-31, Bow87, Van83b, Ano80k, Ano86m, Ano92w, Ano92x, Ano93n, Ano94i, Gup81]. Fr [Ber87, Sai74, Sal87, Dra84, Sch76a]. fractales [Cha90]. Fractals [GG90d, Cha90]. fractional [Ros77, Van24]. fractions [Abd08, Ano74-40, Cre14, Gis96, Kno82, Rit88]. Fraenkel [SS17]. Francais [Yus10, Dho90, Mar10a, SD08]. Francaise [Mal97, Ano77j, Dho88b, Dra84, Pei17, Ano80-32, GW94a, GW94b]. francaises [Ano89t]. France [Ano87h, Ano88u, Ano89m, Ano90k, Ano90-35, Ano91q, Ano91s, Ano91r, Ano92k, Ano92p, Ano92-28, Ano92-27, Bel08, Ber87, Bre91, Che94, Dho88b, Fre76, Gau13, Her75a, Hou89, Maz21, Mue96, Pet82, Ple78, Sal87, Tur08, AO05, Ano83o, Ano90-34, DF10, Pan88, Ple78, Sch06b, Rem12]. Frances [Nic13]. Francesc [Lab01]. Francesca [Van80]. Francesco [Ano88r, Ken76b, Ben88, Luc21, Sac20, GG05b]. Franchis [CB21]. Franci [Ano81r, Ano85f, Hug05, Hug07, Mal06a]. Francine [Dru07, Dru12]. Francis [Ano77p]. Francisci [Net24]. Francisco [Ano74-51, Ano92f, Ano93h, Bun74, Ges18]. Franco [BDD^{+}88]. Franco-Italian [BDD^{+}88]. Francois [Eva94, Gis99a, Vil97, Bor91, Car22]. francs [Ano89g, Ano89i]. Frank [Ano81q, Gra94a, GG90c, LY79, Yok75]. Frankfurt [Ano77h, Ber04, Høy90b]. Frankfurt/M [Ano77h]. Franklin [Ols76, San05]. Frantsuzskaya [Ano81h]. Franz [Ano89j, Bar82a, Bar82b, Ber07, Bur08, GG07, Gra90, Høy24b, Lef10, Mat75, Pet82, Row08, Van05, Van06]. Franzosen [Che14b, Déc10]. Frau [Böt12, Pei15, Row23]. Frauen [Kni01, Row23]. Frauke [Pei15]. Frédéric [Ter89]. Frederick [San03]. Frederico [Fra06]. Fredholm [Vol02]. Free [Dum15]. Freeman [Bun74]. Frege [Fer04b, Gab79, Moo77, Dru86, Gab79, SD94, Vil98, vH67, Fer98a]. Freges [SD94]. Freiherr [HB11]. Freiman [Ano82d]. French [Ano81h, Ano88u, McK82, Par92a, Sch12, Vil96, Ai04, Ano07e, Cha90, CV22, Chr98, DD87, Déc10, Dho83, Dho90, Ehr11, Fio01, Fla03, Gis18, Jaë19, Kas90b, LBB^{+}04, Pan03, Pan05, Pie86, Ras89, Rom10, Ros76, SD08, Ses13, SS88, Spi00b, Ste12, VMT09, YDD83, dMC92]. Frenicle [Fle91]. frères [Cav85]. Freudenthal [Bos92]. FRG [Ano77-28]. Fried [Sai05]. Friederike [Don05]. Friedrich [Ano74q, DGD04, For78, Gar79, GG13, Has10b, Has10a, Joh82, Kit78, Neu86, Pye75, Scr03, SS97a, Str78, Has10a, Haw76].

Friedrich-Wilhelms-Universität [Has10b, Has10a]. Friend [Bel07a, BS06]. Fritz [Grö86, Row85b, Thi06]. Fritz-Egbert [Thi06]. Frobenius [Haw03]. Fromm [Kit78]. Fromman [Ano88i]. Fromman-Holzboog [Ano88i]. Frontiers [Goo77]. FRS [BM74, SS10b, Ped77]. frühen [Ano88i, Høy09a]. Frühgeschichte [Sza74]. Ftr [GW94a]. Fuji [iÔ75]. Fuller [Ano77p, FG90]. Fulvia [Kat11]. Function [EU99, Gra97a, Lau92, Taz01, Wüs02, CMT79, Kë96, SS10a, Scr75b, Spa11, Spi76]. Functional [Fer00, Ano84g, BK84, Bro75, Cra13, Aup76]. Functionen [Osw17]. Functions [Ano81h, Epp95, Fer98b, Fer00, IN84, Kat87, Nik18, Phi78, Smi79b, YDD83, Gra18]. fundament [GG14]. Fundamenta [dW10]. Fundamental [AÖ01, Bal04, CR22, Gol92, Hen75, AF94, Phi79b]. Fundamentals [Lau92]. fundamentos [Aus16, dL01]. Funds [Fen03]. Fünfte [Lem74]. Funktion [Lau92]. Funktionen [Haf17]. Funktionentheorie [Ull03, Ull03]. Funktionsbegriffe [Spa11]. funktsii [Smi79b, Ano81h]. Furinghetti [Kat11]. Furqan [For10]. Further [Gai75, Wei04, Wil80]. fusion [Haw75b]. Fuss [Phi79a, Müü77]. Füssli [Bar77]. future [Bul16b]. fysiske [Atz04].

G [And07, Ano74v, Ano74n, Ano77i, Ano77h, Ano79k, Ano81p, Ano81u, Ano87n, Ano88j, Ano89k, Ano90f, Ano90-35, Ano92l, Ano92t, Bar82a, Bar82b, BKM^{+}82, Buc77, Clu76, Dav19, Dyc75a, Fel78, For75, Gar05, Git74, GG75a, Gri76, IN83, Kin86a, Luc21, Maz77, Pec20, Pet82, Pie83, Pye76, Roe15, Row85b, San06, Ste09, Tan79, Tee79, Thi06, Thi08, Vil96, Vol16]. G. [Ano90l, Dau80, Fel78, HNE21, KB74, RW03, Ano77j, Ano80b]. G.D.R [Bar78b]. Gaṅita [Gup77b]. Gaṅitasarakaumudī [Knu11]. gab [Spa11]. Gädeke [Bla05]. Gaetana [Fin11]. gained [BG21b]. Galdeano [Hor91]. Galilée [Ano91n, Rom10, Fol12, Mac76]. galiléenne [Eva94]. Galileian [EG81]. galileiana [Sai94]. Galileo [Ano93j, Rab08, Wis76, Dra74, Mac77, Rom10, Ser06, Ano78h, Wis76]. Gallaccini [Vag82]. Gallery [Wil77b]. Gallia [Ano89-35]. Galliffet [Ano89t]. Galois [Ano81f, Hel23, Mar99, Neu12, Rad02]. Galuzzi [PG24]. Gambler [Sho86]. games [dMC92]. Gandt [Eva94]. Gaònitasāarakaumudāi [òP09]. García [Hor91]. Garciadiego [Aus16]. Gardabaer [Kat11]. Gardner [Fow84]. Gareth [Jam13]. Gario [Sch99a]. Garland [Lei79, Rid88]. Garrett [Ano77f, Ano75v]. Garry [Ano87b]. gases [Buc77]. Gaston [Ale95]. Gatto [Rom97]. gauging [Mes94]. Gauss [Ano74q, Ano89m, Ano90g, Bie74b, Boi89, Joh84, Neu86, Str78, Ano89k, GG13, Joh86, ARR12, Ano77t, Ano77z, Ano77-27, Ano77-28, Ano77-29, Ano78r, Ano89o, Ano90m, Ano94a, Bie74b, BS77a, Bie80, Bie83, Bul09b, Cam77, Cox77, eFD88, DGD04, For78, Gar79, Gra78d, Hal78, Joh82, Joh84, Lic97, Müü77, Muü78, dBR76, Rob80, Spr78, Sti77, Wat79, Wat84, Wat86, Wat91, Zor96, Edw80, Thi06, Ano89f]. Gauss-Komitees [Str78]. Gaussian [Ber90, Grc11, PR22]. Gaußsche [Ano89k]. Gauss’schen [Ber90, Lic97]. Gaußunf [Ano89k]. Gaußzum [GG13]. Gauthier [Fre77, GG76a].

Gauthier-Villars [Fre77, GG76a]. Gazette [Ano74l]. GDR [BS77a]. Gebhardt [Høy09a, Høy23]. Geburtstag [Ano95f, Str78, Vol81, Ano89s, Fre93, Don08]. Geburtstag/Festschrift [Ano95f]. Geburtstages [Str81]. Gedächtniss [GG13]. Gedankenstrich [Gab79]. Gedenkband [Dho88a]. Gedenkschrift [Str81]. Geest [Ano74v]. Gegenwart [Høy90b]. gegründet [Høy90b]. GEHIMAB [Ano93o]. Geiringer [SS93, EI12]. Geiringer-von [SS93]. gekommen [Hof87]. Gellasch [All09, Kat07]. gemacht [Ped75a]. Genealogy [Cri24b]. genera [Ant95]. General [Nab04, Ano77r, Dor95, Gra76c, Row06b, Wue05]. générale [Ano81m, Kle17, Mue96]. generalised [Cri78]. generalists [Ano74-74]. generalized [Rob80]. Genèse [Høy07]. geneseologie [Spa89]. genesis [Dor95]. Genetic [Bo24, Pie83]. Geneva [Ano77q, Bra09b, Gol82, PP91]. Geneviève [Ano78i]. génie [PR12, Bor14]. Genius [Ano87e, Zit07a, AB09, Bat06]. Genoa [Bot80b, BD11]. Genocchi [Gia90, Isr90]. Genootschap [Ano81e, Str82a]. genre [Lê20]. genuine [Row06b]. Geography [Plo06a, BJ00]. geomagnetism [Gar79]. Geomatikum [Ano90t]. geometer [BG21b, Cox77, PB08]. Géomètre [Gis99a, Bra09b]. Geometria [Sch13c, Ano92s, Fri21, LD22, Mai08, Par97, Taz09]. Geometric [Nor07, Sch05, Zho21, Abe93, BS07, DB23, Gra91b, HNE21, Lüt05, Pyc97, Sch79]. Geometrica [Deb13, dW10]. Geometrical [DS99b, Lê15, McL12, Ano77r, Ano91e, DF18, Hog86, Mae89, Pan11]. Geometrie [Ano89k, Mat75, Sch13c, Bre07, Hug10, Lo 84, Mai08, Ano90g, Mar90b, Ai04, Ano89m, BKM^{+}82, Boi89, Dra84, Fla03, For10, Gra99a, Hog01a, Len79, Mar93b, PP91, d’E01, Ano88n, Gra78d, Sau06a, Sza74, Lo 85, Ano84e, Clu76, Høy09b, Ano78j, Ric84, De 20, Eva94, Har06, Cor04, Vol93b, AG77, Krö17, Sim96, FH08, Syl81]. Geometries [Sch13c, Ano74p, Cer04, Gra12, Lüt15, Roz90]. geometriilor [Ano74p]. geometrischen [Gra90]. geometrization [Len79]. Geometry [BGG06b, BGG06a, Cra98, Cra00, DS99c, GG02, GG03, Gra97a, Gra09b, IM98, Lum97, Row07, Sch13c, Swe86, Zor96, vdW76a, Ai04, Abh75, ABBS09, Ano89u, Bar11, Blå17a, Blå23, BC04, Cer04, Cog16, Cro75a, De 95, De 09, Est06, FK90, Fla03, For77, Gra79, Gra91b, Gre83, Haw80, Hog81b, Kno83b, Kë96, Lew90, lLW13, Lüt21, Mai08, Mar93a, Mar06, Mar10b, Mol76, Ols76, PZ11, Pej07, Por82, Sai95, Sch79, SS12a, Syl81, Sza74, Wie23, Zun83, dW10, Lo 85, Yan83]. Georg [Ano74q, Ano81g, Ano92w, GG07, Her14, Jen76, Sch16, Fer93, Jen76, Mal13, Mar06, Sis02, TC05, Vol06, Haw81]. Georg-August-Universität [Ano92w]. George [Boy74, Bun76b, Eis75, Kat92, Lév99, Row85a, Was20, Ano81v, Ano94j, Don98, Hog76, Pan87, Par02, Pyc81, Wee91, Wol08, Zun02, Ano87e, Wil77b]. Georges [Yus10]. Georgii [Por75]. Gerald [Dav19, Ano74-40]. Gerdes [D’A16]. Gergonne [BGSS16, Lor15, Sti74, Sti76a]. Gerhard [Plo06b]. Gericke [Ano81i, Fol84, FR08]. Germain [Ano84i, DD87, Del05, LP10]. Germaine [Ano95n]. German [Ano74v, Ano90j, Bul24, Eve76, Seg76, Ano89s, Ano91-32, Ano92-52, Arn88b, Ber90, Bie74b, Bie80, Böl93, BMW08, Böt12, Bul09a, Bus76, Che14b, Déc10, Eck13, EI12, Eng79, Eng81b, Eng81a, Fel78, Fle86b, Fol78, Fol83a, Fre93, Fri91, Gab79, Gre83, Has10a, HB^{+}13, H^{+}08, Hof87, Hof92, Imh03a, Ine88, Ine91, Ine94, Jen76, Jen86, Joh82, Joh84, Kna88, Kni01, Kno78a, Kno83a, KO09, Krö08, Lau92, Lic97, Neu93, Nis89, Nos12, PS83a, Ped75a, Pet78b, Pul05, Pur86a, Pur88, Pur90, Pur91, RFS89, Sch88, Sch14, SD94, Scr81, Seg86, SS94, SS97b, SS12b, SEM04, Spa11, Str85, Sza74, TC05, Tob81, Tob06, Tob10, Vog83, Wue05, Yus95]. German-speaking [Nos12]. Germanische [Lév99]. Germany [Ano89-34, Ano90-33, Ano91-33, Ano92h, Ano93x, BLF92, Ott88, Scr87, Sen10, Ano92b, Bul09b, Has10a, Jah93, SS09, Tra21, Row85c, SS17]. Gernet [Che94]. Gerolamo [Gal12]. Gersonides [Sim00a, Sim00b]. Gerstenberg [Bir80, Jay83]. Gert [Ano86h, Dho11, Kat11, Lei09, Tob90]. Gervais [Row15]. Gesammelte [Don05, Plo06b, Scr07, HB^{+}13]. gesammelten [Hes90]. Gesamtausgabe [Kit78]. Geschichte [Ano91-32, Ano92b, Ano92-46, Ano92-52, Ano94b, Bin89, Bos77, BMW08, Eng79, Eng81b, Eng81a, Fol87, GG07, Gra90, Hof92, Ine91, Jay83, Jur07, Lem74, Mat75, Pur88, Pur90, Pur91, Sch90a, Sch17, Sch09, Scr03, SS89, Tob81, vdW80, Pur86a, Ano81i, Ano85e, Che84, Fel85, Ora82, Ric84, Toe96]. Geschichtsschreibung [Spa89]. Geschicted [Ano89j]. Geschlecht [Lê20, Lê20]. Gesellschaft [Ano74q, Ano86h, Ano90l, Ano90t, Grö86, Tob81]. gesuiti [Rom05]. gesuitica [Ano88r]. Gesuitico [Rom97]. get [Scr81]. Getaldić [Dad84]. Geyer [Lo 83]. Gheorghe [ABBS09]. Gheverghese [Kat92]. Ghorbani [Bag02]. Ghori [Ano88k]. Ghost [Fen09]. Giacardi [GG05b, Gra99a]. Giambattista [Vér01, Rom05]. Gian [Tro83]. Gian-Carlo [Tro83]. Gibbs [Gai75, Ust74]. Gibraltar [Hog24]. Giessen [Haw76]. Gil [Yin11]. Gilain [Ano92k, Lor16, FG12, Lüt15]. Gillies [Row93]. Gilling [Gil78]. Gillings [Ano74y, Ano74-73, BS77b]. Gillispie [Tat75]. Gimlekollen [BJ89a]. Gino [Ano91k, Mal06a]. Giochi [Hug07]. Gion [FC22]. Giorgio [Ano92n, Pan18, Ray23]. Giornata [Ano91k]. Giornate [Roe89]. Giovan [Ser90]. giovani [Hug07]. Giovanna [Mal97]. Giovanni [San09, BFG94, BF23, Ros81]. Gipson [Zas83b]. Gipuzkoa [Ano91p]. Girard [Ano90h, Ano92r]. Girolamo [Ano79j]. Gisela [Pei22]. Gispert [Ano92k]. Gittleman [Hug76]. Giuseppe [Ang88, Cib76, GG05a, Ano83e, FR82, Ken74b, Mon75, Ume76]. Giusti [Sai94]. give [Spa11]. given [Cas80, ET03, TE03, Dho11, Fer11, For10, Fre92, Gra10b, Kat11, Knu11, Yus10, Sid06b]. Glaser [Knu83a, Knu83b]. Glasgow [Gra02]. Gld [Pye76]. Gleichungen [Scr03]. Glen [Bra07, Eva13, Fra17b, Moy17, Ano03a, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c, Ano04d, Ano05a, Ano05b, Ano05c]. Glenie [Cra09]. Glimpse [Cri99]. Global [Sen10, SS09]. Glossaren [Rom04]. Glossary [vM84]. GmbH [Ano89j, Gra90]. GNT [Has04]. GNT-Verlag [Has04]. goal [Lau92]. God [Bar11, Gan86]. Gödel [Fer04b, GG90b, Moo77, Pec20, Ano83d, Ano93u, Chr84, Daw85, Daw86, GG79, vH67]. Gods [How18].

Goethe [Ber04, Dyc75b, Dyc75b]. going [Kat11]. Goladhyaya [Gur24]. Goldberg [Haw96]. Golden [Coo96, Bec19, HF19, Ano93p, Ano93w, RT02]. Goldstein [Coo99]. Goldstine [Ano92j, Cal83, Bir78]. Gonseth [Ano74-30, PP91]. Gonville [Ano93m, GG87b, GG90d]. good [May75d]. Goodstein [Ano74t]. Gordon [Jon85, Pye75]. Görlitz [Ano90j]. Gosen [Ano92v]. Gösta [Vol02, Dau80]. Gottfried [Cri17, HB11, Pet78a, Bar78b, Bla05, Gar05, Her14]. Göttingen [Ano74q, Ano88n, Ano90-33, Ano91-33, Ano92w, Ano92-47, Ano94i, Edw77a, Row06a, Sau14, Sch90b, Ull03, Ano77z, Ano77-27, Row90b, ST16, Ano74w]. Gottlob [SD94, SD94]. Goulding [Zep12]. Gowers [Fra10]. Grabiner [Dho85]. Graduate [Gra77]. Graeco [Ano87d]. Graeco-Latina [Ano87d]. Graham [Ano85h, Dau12]. grand [LP10, lLW13]. Grande [Dur21]. Grant [Ano84d, Gra91b, Wol03]. graph [WM14, Har77a]. graphic [RS85]. graphical [Chi22]. graphies [May76]. graphique [Tou03]. graphische [Has04]. Graphs [Asc88, KQS85]. Graßmann [GG10, Ano92t, Ott89, Pet09b, Rad00, Rad03, Lew09]. Gratis [Hug05]. Grattan [Ang79a, Ano74-56, Ano74-59, Bra12, DGL^{+}15, GG98, Gra98b, Her75a, Par92a, SS07]. Grattan-Guinness [Ang79a, Ano74-56, Ano74-59, Bra12, DGL^{+}15, Gra98b, Her75a, Par92a, SS07]. Grave [Ano92j, Hes90, Law98]. gravitation [Ano90q, Row06b]. gravitational [Wue05]. Gravitationsgleichungen [Row06a, Wue05]. Gravity [Ano78h]. Gray [Coo99, Neu08, Oli08, Ric82, Thi06, Wal17, Lê24]. Great [Ale14, Fle75, Fri74a, Gol00, Aud11, Enr85b, Enr85a]. Grèce [Auj76]. Grecheskaya [Ano78s]. greco [Net02]. Greece [Ano92-30, Bow87, Kno82]. Greek [Ano77i, Ano92-30, Bow84, CM19a, McK82, Moy17, Sia23, Sid06a, Zer75, vdW76a, Art91, Ber84a, CM19b, FT83, Hog81b, Kas90a, Kas03, Kno83b, Sai95, SS10a, Sid04, SS12a, Wan22, BT81]. Greek-speaking [CM19b]. Greek/French [McK82]. Green [AH10, Fra10, Wil77b, Ano94j, Dia99, Taz01]. Greenberg [Lew98, GG05d]. Greenwood [Ano89l]. greetings [Dau90]. Gregoriano [Don05]. Gregorius [Van84]. Gregory [Dru91, Lüt20, AB02, Scr83]. Griechischen [Bow84]. Griffiths [Kil83]. Grigori [Ano82d]. Grigorijan [Ter89]. Grigoryevna [Ano81-35, DPY91, DPS^{+}02, DV07]. GRIMM [Bek88]. Grinstein [Ano89l]. Groningen [Ano77o]. Gronwall [Glu05]. Grootendorst [Str82a]. Grössenlehre [Kit78]. Grössing [Fol87]. Group [Ano79g, Ano79-30, Ano80p, Ano86n, Ano92v, Ano87j, Ano89-31, Bek88, Boo86, Haw77a, Meo04, Nic93, Nur92, Per93]. Groupes [Cog15, ET14]. Groups [Goo77, Wüs02, Cog14, Sch06a]. Grove [BT81, Par92b]. Growth [Ano80-37, Yok75, Yos75, Yos76]. Grundlagen [Lau92, Tob81, Vol93b, Ano88n, Sau06a, Cor04]. Grundlehren [Ano78j]. Grundzüge [Plo06b]. Grüner [Mue80]. Gruppe [GS91]. Gruyter [Ano81i, Ano93i, Haf16]. Guarujá [Ano89o, Ano90m]. Guérindon [BKM^{+}82]. Guerraggio [Arc09]. Guerre [Dur21]. guerres [Ano89t, Gau13]. Guggenheimer [Ano74-30]. Guglielmo [DF10, Rem12]. Guicciardini [Sag91]. guide [Ano87d, SS84]. Guides [Ano89c]. Guido [Isr85, Sch99a, Ysh90, Gar01]. Guilbaud [Lor16]. Guillaume [BKM^{+}82, Ehr11]. Guimarães [Sar97]. Guinness [Ang79a, Ano74-56, Ano74-59, Bra12, DGL^{+}15, GG98, Gra98b, Her75a, Par92a, SS07]. Guitel [Ano78i]. Gülru [DS99c]. Guo [Shu92, Vol10]. Gupta [Scr96]. Gustav [Sel90]. Gutenbergplatz [GG13]. Guy [Ano95o]. Guyot [BH14]. guys [Row06b, Wue05]. Gyldén [Vol02]. Gyula [Ano80f].

H [Ano74v, Ano74-30, Ano77o, Ano78-32, Ano82c, Ano89k, Ano89m, Ano90g, Ano92k, Ano92j, Ber74, Bie78, Bir78, Boi89, Bos77, Bur08, Cal83, Cho92, Dau80, Dav19, DS99c, Don05, Don08, Dru91, Edw80, Eva98, Glu05, Glu12, GG75a, Gre77, Hes90, Lef10, Lew99, RW03, Row83a, Str82a, Too80, Vol81, Ysh90, Zas83b, Dje98, Pec20]. H*āsib [Kc74]. H*abash [Kc74]. H-Z [Pec20]. H. [Ano84n, Tan79, Tee79]. Hôtel [Ano89t]. Haṡṡār [AD87]. Haase [Ano80m]. Habicht [Ano77m, Bar77]. Habil [Ano79-42]. Habilitation [Has10b, Has10a, Row23]. Habsburg [BS13]. Hacking [Ano92o, Kno80]. Hadamard [Gra76b, LN94]. Hadden [Ano95l]. Haffner [Lê21]. Hagen [Neu93, Neu93]. Hague [Ano77s, Pye76, Pys81]. Hague/Paris/New [Pys81]. Hahn [Gil15, Lo 84]. Haig [Ano74l]. Halbgruppenbegriffs [Hof92]. Hald [Sen91]. Half [Ano89-35, Cha90, Moo82a, iÔ75]. half-century [Cha90]. halftone [Fer04b]. Hall [Ano87n, SG04]. Halle [Bos77, Jen76]. hallenser [Jen76]. Hallett [Ano89c, Row07]. Halley [Gow95]. Halmos [Ano87f]. Hamburg [Ano90t, Ano90-37, Haf16, Row23, Scr87, SY90]. Hamburg-Munich [Ano90-37]. Hamel [Sis02]. Hamilton [Cri78, Øhr85, Win82, Hel83, Nak03]. Hammurapi [Gra86]. Hand [EU99, Ano92n]. Handbook [Gra10a, Hog24, RS09, Ano74n]. Handle [GG96, GG97]. Hands [SG07, All09]. Handschrift [Fol09a, Jay83]. Hankins [Hel83]. Hannover [Bar78b, Pet82]. Hanover [Ano89-34, Ano90l]. Hans [Ano77h, Ano81g, Ano84k, Ano88i, Ano95f, Bel12, GG10, Lew09, Mue80, vdW80, Ano80-35, Bos92, SS12b]. Hans-Georg [Ano81g]. Hans-Joachim [GG10, Lew09, Mue80]. Hans-Jürgen [Ano88i]. Hao [Bun76a]. Harald [Sch11]. hardbound [Ano87k]. Hardcover [Ser18, Enr85b, Enr85a, Haw03, Hud12]. Hardy [Dav19, Gra91b, Wol03, Dau80, RW03]. Harley [CWW17]. Harmattan [Gol13]. Harmonic [Wat93, Mac75]. Harmonies [Gra13, KK11]. harmonique [Ano91s]. Harold [Spi76]. Harper [Arc96, Hym78]. Harrap [Pye75]. Harrassowitz [BB04, Gol76]. Harriot [Ano87-34, Tat11, BS09, Fen89, Pep75, Ste07a]. Harrison [Coh82]. Harry [Rob96]. Hart [Hud12]. Hartman [Moo82b]. Hartmut [Sch11]. Harvard [Ano77f, Ano84d, Dau12, Dum18, Fer04b, Gab76, Gra11, Haw81, Lew99, Moo77, Sch91b, Sen06, Tro74, Ano74-44, Ano74-45]. Harward [Moo76]. Hashagen [Row08]. Hassan [Ano81p]. Hasse [FS10, Seg80]. Hastinapur [Ano85n]. Hastings [SG09]. Hauppauge [Bie74a]. hauptschriften [Ano77h]. Hausdorff [Mon83, Plo06b, Scr07, HP05, SS10c, SS22a, Sch07].

Hautes [Ano87h, Che94, Ple78]. Haven [Ano74-29, Ano75p]. Having [Mon11]. Hawig [Ano95m]. Hayashi [Sid12]. Haytham [Ano87g, For10, Hog02b, Ras89, Roz76, Sim96]. hazard [dMC92]. Heart [Tim09, Hog08]. Heat [Her75a, BR16, Buc77, Ros17]. Heavenly [Fra17b, Mor19]. Heavens [Eva13, Van09]. Hebrew [Eli18, Eli24, Lan96]. Hecht [Sch11]. Hegel [Pet91]. Heidelberg [Ale14, Ano77i, Ano84j, Ano87j, Ano87g, Ano88j, Ano89f, Ano93g, Bel12, Bor18, Cog12, Edw77a, Edw77b, Fol15, Fre92, Gra94a, Hog85b, Krö14, Pei15, Plo06b, Row07, Row19, Sch17, Sch12, Scr07, Swe79, Vol93a, Zep12, Ebb07]. height [Fri74b]. Heijenoort [Fer04b, Moo77, Ane88]. Heike [Bos03]. Heimatblätter [Str81]. Heims [Bir83]. Heinekamp [Pet82]. Heineman [Tan74]. Heinrich [Bul22, Gra78c, Haf16, Jen86, Lê21, Nor07, Sch14, Vol81, Vol93a, Vol16, Haf17, Lüt05, Pon78, Scr79a]. Heinz [Gar05, Her14, Row85d, Sch96, Sch09]. Heinz-Jürgen [Her14]. held [Cas77, MP06]. Hele [Ano87n]. Helena [Sch05]. helix [MA10]. Hellenistic [LH84]. Hellmann [Fol09a, Høy24a]. Helm [Sal78]. Helmholtz [SS89, Vol93b]. Helmut [Lüt09a, Ano80m, FS10, Seg80]. Helmuth [Ano81i, Fol87, Fol84, FR08]. Helsingfors [Sel90]. Helveticae [Rid83a, Rid83b]. Henk [Gui24, Mar10b]. Henri [Ano77n, Ano89m, Ano90d, Ano90k, Ano91q, Ano91s, Ano93t, Bie78, Obe78, Row15, Sau09, Vol02, Ano07e, Maw06, Wal17]. Henrici [BG21b]. Henrique [Ges18]. Henry [Bra09b, Jen86, Ust74]. Hensel [Ano79w]. Her [Ber97, DPS^{+}02, GC82, Ano81-35, DPY91, Row83b]. Herausgeber [Ano77m]. Herausgegeben [Ano81g, Bar78b, Bul22, Fol87, SS89, Str78]. Herbart [Sch82a]. Herbert [Bir80, Mon83, SS22b]. Herdade [Dho91]. Hereditary [Wil74]. heretofore [Ros76]. Herglotz [Ano89k]. Hérigone [Mas08]. Heritage [Ano74r, Ano77d, Cas77, For10, Gup88, Amb08, Ano74g, Ano74h, Ano76e, GG04, Ken76a, Luc21, Oak12]. Herivel [Bru77]. Herman [Ano92j, Bir78, Cal83]. Hermann [BKM^{+}82, Bos77, Bun76b, Cog15, Gis14, Goo77, Gra76a, Haw77a, Haw77b, Jen86, Krö17, Nag91b, Nag91c, Str85, Too80, Tru75, Ano92t, ET14, Gui96, Jen86, Nag91a, Ott89, Pet09b, ST16, Str85, Fre92, GG10, Nab04]. Hermelink [Scr79a]. Hero [Fek14, Kno93b]. Heroes [Ale10, Gra11]. Heron [Tai14, Tra10]. Herrlich [Plo06b]. herrliche [KS12, Tob10]. Hersh [Hil83]. Hertz [Nor07, Lüt05]. Hervé [Dra84]. hervorragender [Row85b]. Herz [Høy90a]. Herz-Fischler [Høy90a]. Heß [Gar05, Her14]. heuristic [Kno78b]. Heybech [CG92]. Heyde [Das81]. Hfl [Mue80, vM84]. Hibeh [FT83]. Hibner [Ano86d]. Hidden [KK11, Gra13]. Hiebert [Stu80]. Higgins [Dav90]. High [Abh75, Syl81]. High-school [Abh75]. Higher [Sai74, Sco16, PB08, Swe74]. Highlands [Ano77s]. Highways [Bra12]. Hilary [Ols76]. Hilbert [Daw16b, Moo03, Row07, Sau06a, Ung85, Ano81y, BGSS16, Cer04, Cor04, ESH^{+}13, HEH^{+}04, Moo02, Pat14, PK02, Pio85, Row06b, Wue05, Row06a]. hilbertische [Ano74v]. Hilberts [Ano88n]. Hilda [EI12, SS93]. Hildesheim [Ano74q, Bir80, Bul22, De 20, Her14, Jay83, Sch16]. Hill [Boy74, Lor18]. him [Ano92i]. himself [Roz76]. Hintergrund [SS88]. Hintikka [Ano77r]. hir [Sai74]. Hirayama [Ano78-32]. Hirose [Ano78-32]. Hirsch [BKM^{+}82, Ano95o]. Hirst [BM74]. Hist [Jay24b]. Histoire [Ano89n, Ano90g, Ano91q, Ano92a, Ano92-44, Ano93-33, Bre07, Fla03, Ano87-28, Ano89m, Ano90d, Ano93v, Boi89, Cor96, Dho91, Dug76, Gis99b, Høy24b, Mar90a, Pet74, Pic89, YDD83, Ano77k, Ano78i, Ano81j, Ano89g, Dau93b, Gis96, Har06]. Histoires [Gra12, Lüt15]. Historia [Ano85j, CM19a, Dau82a, Ano82-27, Ano87r, Del03a, GG98, Gra98b, SY90, Has10b, Ano92-33, BF00, Dau92, Dau94, Dau99, Goe09, GG03, Gui09a, KP85, Kno17, PH96b, PH96a, PH97, Scr79b]. Historian [Bos02, Bra07, You83, Ano81a, DF10, Gil89, Gra75a, Gup80, Jon76b, VK05]. Historians [Ano74-67, Ano85r, Ano95-27, Ano74-74, Ano74-75, Ano81-45, Ano83i, Ano92-51, Ano93a, Par03, SS10b, Wal75]. Histórica [Lei09, Ano77g]. Historical [AÖ01, Ano93i, Ano95j, Bow84, Hoc96, Hun12, Kje00, Par92a, Vac20, Wüs02, Ano74f, Ano76-36, Ano79-34, Ano82-29, Ano88g, Aup76, Bak75, Boy74, Cam77, GG75c, Lau92, Lei77, Lew75, Mai08, Mik75, Pla76, Pul05, Sch13c, Scr75b, Smi79a, Sti76b, TC05, Van84, War10, vBS20]. historical-educational [Mai08, Sch13c]. historically [Che09]. históricos [Aus16]. Histories [Bul16b, GG76c]. Historiographic [May75b, May75a]. historiographical [Van22]. Historiography [Ano92-29, Sar97, Ano87-27, Ano91h, Gui03, Meh76, Sma16]. Historique [Pie80, Bab77, Enr76, Ano76-35]. historiques [Ano89n, Ano92-35]. historisch [Bos77, Høy90b]. historisch-kritischer [Bos77]. historisch-systematische [Høy90b]. historische [Lau92]. historischen [Ano92j]. History [Abe91, Ale14, All09, Ang88, Ano74i, Ano74-32, Ano74-47, Ano74-50, Ano74-49, Ano74-52, Ano75g, Ano75p, Ano76v, Ano77d, Ano78e, Ano78k, Ano78q, Ano78u, Ano78t, Ano78x, Ano78w, Ano79n, Ano79p, Ano79z, Ano79y, Ano79-45, Ano79-46, Ano79-47, Ano80c, Ano80l, Ano80p, Ano80r, Ano80-39, Ano81p, Ano81-32, Ano82l, Ano82k, Ano83w, Ano83x, Ano83y, Ano84h, Ano84q, Ano84x, Ano85k, Ano85-30, Ano86c, Ano86l, Ano86m, Ano86n, Ano86o, Ano86r, Ano87n, Ano87-31, Ano87-32, Ano88-31, Ano88a, Ano89-33, Ano90n, Ano90a, Ano90-37, Ano91f, Ano91o, Ano91-33, Ano92h, Ano92w, Ano92x, Ano92v, Ano92-47, Ano92-48, Ano92-49, Ano93-34, Ano94i, Ano94x, Ano95j, Ano08e, Arb86, Arc96, ABC^{+}88, Atz06, AD02, Bec01, BJ89a, Bel12, Ber84a, Ber04, Bin83, Bin91, Bir78, BF00]. History [BMW08, Bra12, Bra17, BLF92, Cha81, Clu76, Coc98, Con85, Cri82, D’A79, D’A10, Das81, Dau86, Dau93a, Dau94, Dem74, DV80, DK86, Dru85, Dum18, Eva13, Fen08, Fle75, Fol15, Fra06, Ful99, Gam96, Gar86a, Gar92, Glu12, Goo77, Gra10a, GG76c, GG98, GS77, Gra78b, Gra84b, Gra87, Gra98b, Gra09b, Gui96, Gup75a, Gup78, Gup79, Gup81, Gup85, Gup86, Haw03, Hog85b, Høy09b, Imh21, Jah12, Jon77a, Kas90a, Kat97, Kat07, Kat11, Ken75, Ken76d, Kje93, Kje17, Kno94, Kno88b, Knu83b, Lê24, Lew76, Mac78a, Mar09, Mat93, May75c, May77c, McL12, Meh80, Mel09, Mic23, Mic88, Mol81, Mon11, Mro16, Neu08, Nic13, Nor75, Oli08, Ott88, Par89, PH96b, Par03, Por05, Pur86b, Pur88, Pur90, Pur91, Pye74, Rec07, Rid94, Rit88, RS09]. History [Row90b, San03, Sar88, Sau06a, Sca94, Sch79, Sch07, Scr79b, Scr85b, Scr85a, Scr85c, Scr87, Scr88a, Sen06, Ser18, SG04, Shi86, Shi93, Sia23, Sid06a, Sti78, Swe92, Tee78, Tob81, Ven02, Wil79b, Wil97, Wis76, Zep12, ZD88, Ale95, AKW04, Ano74e, Ano74g, Ano74h, Ano74o, Ano74m, Ano74-33, Ano74-55, Ano74-64, Ano74-71, Ano74-76, Ano75f, Ano76f, Ano76e, Ano76o, Ano76-34, Ano78-28, Ano78-34, Ano79d, Ano79e, Ano79v, Ano79-30, Ano79-33, Ano79-37, Ano79-43, Ano79-48, Ano80t, Ano80u, Ano80-38, Ano81e, Ano81-31, Ano81-36, Ano81-46, Ano82j, Ano82-27, Ano83j, Ano83z, Ano84j, Ano84i, Ano84m, Ano85h, Ano85i, Ano85-27, Ano85-31, Ano86j, Ano87j, Ano87-30, Ano87-36, Ano88-30, Ano89-32, Ano90v, Ano90w, Ano90-33, Ano91k, Ano91-32, Ano92i, Ano92n, Ano92-52, Ano93n, Ano93o]. history [Ano94g, AB84, Arn88a, BTB^{+}03, Bek88, Ber85, Bin90, Boo86, BM77, BS07, Buc77, Cal83, Cas77, Cas80, CV22, Cha87, Che87, Che14a, CM19a, CM19b, Cog14, Cro75c, Cul07, Dau82a, Dau85b, Dho83, Dun89, Eis80, Eng79, Ewi92, FG06, Fis11, Fol76, Fol83c, FR83, Fri74b, Gar86b, Ger94, Giu86, Goe08, Goe09, Gou10, Gra75a, GG87a, GG04, Gra76c, Gra80a, Gra86, Gra91b, Gra91a, GP07, Gui24, Gui87, Gup76b, Gup93, Haw84, H^{+}08, Hof92, Hog24, Høy90a, Hug75, Ine91, IN84, Jon75b, Kat81, Kle07, Lüt09a, Lüt19, Mac78b, Mal79, Mar20, Mar90c, May75d, McG74, Nur92, PS83a, Per93, Pet78c, Pow76, Pri80, Pye75, Rid88, RH11, Roe89, Row83c, Row89, Row94, Sar93, Sch91b, Sch89, Scr82a, Scr82b, Scr84, Sel90, Sen91, Sen92]. history [SGJ05, SG07, Shi94, SS12b, SS22b, Smi79b, Sto78, Swe79, SS84, Sza74, Tro76, Tro83, Van09, Vog94, YDD83, Zas83a, dS90, vD09, vM84, Ano74-46, Ano74-48, Ano74-51, Ano76u, Ano76w, Ano76x, Ano77-30, Ano82l, Ano82k, Ano83k, Ano90s, BEJ^{+}75, Eng81b, Eng81a, Fla03, Jon81, Pur86a, Pyc84a, Rog74, Rog77, Ayo06, San06, Zep12, Ano84h, Ano89d, Ros74, Wol91, Aga75, Ano79k, Ano81k, Ano84g, Ano91d, Eve76, Hug76, Knu83a]. hitherto [De 15, LH84]. Hjelmslev [Lüt21]. Hlavatý [DS19]. Hobbes [Ped77, Gui01]. Hochschule [Mai87, Row85d, Spa89]. Hochschulen [SS08, Tob06]. Hoe [Ple78]. Hoffman [Nic13]. Hofmann [Bar78b, Bur75, Scr75a, Scr85c]. Hogan [Tim09]. Hogendijk [Ano87g, DS99b]. Högskola [Tur20]. höheren [Bos77]. Holbrook [Ano74z]. Hölder [Rad03]. Holdred [Ful99]. Holkema [GG77]. Holland [Ano84g, Ano84h, Ano95j, Buc77]. Holland/ [vdW76a]. Holland/Boston [Ano77r]. Holloway [Ano74-46]. Holmes [Ayo06, AKW04, Rob01]. Holt [Ber74]. Holzboog [Ano88i]. homeomorphism [Moo07]. homine [Ped77]. Hommage [Dug82]. homme [DR03, Krö15]. homology [Bal23]. Homöomorphieproblem [Krö06]. Hong [Ano93g, Gra94a]. honneur [FG12, Lüt15]. Honor [Moy17, Ano76w, Ano79o, Ano91k, Eis75]. Honorary [Ano80q]. Honoré [Bré12]. Honour [Eva98, Gia90, Mar10b]. Hooke [Chi22, Erl97, Nau98]. Hopkin [Des09]. Hopkins [Abe06, Ano88e, Bel07a, Bra12, Hel83, Hud12, Mon17, Per10, San05, SG09, Wil07, Cra07]. Horiuchi [Ebe96]. Horizons [Beb96, Vuc02]. Hormigón [Ric98]. Horner [Ful99]. horologerie [Gol82]. Horovitz [vdW80]. Horst [Plo06b, Seg76]. Hot [Sho02]. Hotel [Ano91n]. Hoüel [Gra99a]. Hourya [Cor96]. House [Ano87n, Bir97, Syt91]. Houston [All04, Ano74r, Ano81k]. Houzel [Ano92k, BKM^{+}82]. Howard [Enr85b, Enr85a]. Howlett [Tro83]. howling [Hog74]. Howson [Ano74n]. Høyrop [Buc80]. Høyrup [Lev97, Mel06, Høy93]. HPM [Boo86]. Hrsg [Hes90]. Hua [Ano91g]. Huber [Pin76]. Hubert [Ang88, Cib76, Mon75, Ume76]. Hubertus [Fol09b]. Hud [Hog86, Hog95]. Hughes [Gra83, Høy09b]. Hui [Fu91, Nee79, SS86, Wag79, Wag85, Yon74]. Huillier [Dra84]. Humaines [Eva94]. Human [Dal92, Tim09, Hog08, SS93]. Humanism [Tan74]. Humanist [Ano92e]. Humanistische [Fol87]. Humanities [Ano77s, Ano90u, Bun76a, Lei79, Sal78]. Humblot [Row85d]. Humboldt [H^{+}08, Sch11, Str78, Vol81, Bie80, H^{+}08]. Humboldts [Bie80]. Humphries [Ano74s]. Hunayn [dY92]. Hund [Pye75]. hundertsten [Sau10]. hundred [Jan11, Kul76]. hundredth [Sau10]. Hungarian [Has10a]. Hungary [Mik75]. Hunger [Kje17, Neu08, Nor03]. Hurwitz [Osw17]. Husek [Plo06b]. Huygens [Ano79-31, Atz06, Lüt20, Dij04, Fre80, Scr83, Sho86, vM84, vM85, vM84]. Hwarizmi [Rom99]. Hybridizing [AH20]. Hyde [Ano74w]. hydre [DR03]. Hygiène [Bra09b]. Hypatia [Zep12, Gou10]. hypergeometric [Sen94]. Hypothesis [Sch99b, Wei04, Car22].

I. [Ano74-45, Ano74-56, Ano75v, Das81]. I.39 [Wat93]. I.C.M.E [Ano80p]. I.R.E.M [Ano92a]. I.VII.17 [Hug05]. Iacopo [Fol14, DN12]. Ian [Ano92n, Ano92o, Høy95, Ung85, Kno80]. Ibarra [Ano93i]. Iberia [Ano93p, Ano93w]. Iberian [Hog24]. Ibero [D’A79]. Ibl [Roz76]. IBM [Ros74]. Ibn [For10, Sim96, BH89, Hog85b, Hog85c, Hog86, Hog95, Hog01a, Van98, dY92, Dje97, Dje98, Hog02b, Ras89, Ano87g]. Ibrahim [AH20, Hog01a]. Iceland [Dum15, Kat11, Art13]. ich [KS12, Tob10]. ICHM [Ano85q, Dau90, Par03]. ICMI [Jan11]. icosahedron [Hog87]. ICTP [Cho86]. idea [BS07, Row90a]. ideals [Wat84]. Ideas [Gra99b, GO99, Asc84, Daw16a, Ott89, VHK16, DS^{+}02, Van06, Ano87j, Par92b, Ric82]. identities [EF20]. Ideology [Fef97, SS93]. IFC [Ano10b, Ano10c, Ano10d, Ano10e, Ano11b, Ano11c, Ano11d, Ano11e, Ano12a, Ano12b, Ano12c, Ano12d]. Ignatieff [Bos03]. ignored [Ros76]. ihre [Cal75, Seg90]. ihrer [Bos77, Lic97, Ora82]. II [Ano81p, BKM^{+}82, Buc77, GG76a, GG90b, Hod75, Jon76a, Jur07, Kit78, Lo 85, Maa07, Mac76, Pin76, Plo06b, Rid94, Sch16, Str82a, Ano75n, Ano81-46, Bin89, Cro75a, Fis79, Gar92, Kje00, May75b, PS83b, Ros17, Sel75, Sim00b, Mat75, Ter89, Isr85, Ysh90, vdW76b]. II.vi.5 [CH90]. III [Ber87, Don05, Gar05, Mos16, Sal87, Ano81k, Ano76z, Ano76-27, Ano81x, dMC86a, Tai14]. ijāb [Dje98]. Ikh [Kin86a]. Ilfracombe [Kil83]. Ilkhani [HZ85]. ill [Pec20]. Illinois [Ano74-30, Ano82d, Ano86e, Dru85]. Ills [Nah06, Lan07]. Illus [Shu80, Ber87, Cog12, Dav19, Lei79, Sal87, Ano81e]. illustrated [Syl81, Tro74, Ano84e, Gri82, Mon83]. Illustrations [Cas77, Pye75, Ano85f, Ano89i, BT81, Ano85g]. illustrés [Mar93b]. im [Ano81x, Ano85e, Ano86h, Ano88i, Ano91u, Bie74b, Böt12, Che14b, Eng81b, Eng81a, Gra90, Has10b, Has10a, Ine94, Jay83, Kna88, Krö14, Neu93, Pei15, Row85d, Sch83, Sch17, SS89, Str78, Tob90]. imagery [Ste07a]. Images [Nor07, Sch13b, Lüt05]. imagined [Lor15]. Imhausen [BB04]. immaginazione [Rom97]. Immagini [Arc99]. Immanuel [Rab74]. immer [Lau92]. Immortality [Dyc75a]. Impact [Kje00, EI12, KK90, Roy17, SS09, Sen10]. imperative [Tur20]. Imperial [Mor19]. Importance [Gup86, Sid06b, ET03, Lüt09b, TE03]. important [Kno78a, Roe15]. Imposition [Ano76y]. Impossibility [Lüt20, Lüt09b, Lüt19, Mic23]. Impossible [Sch05, Pyc97]. improvement [Pet78b]. Impulse [Erl97, Spa89]. Imre [Fra10, Ano81l]. incl [Swe79]. Including [Ano78h, Ano95j, Del06, HB^{+}13, Kat92, Par92b, Sch91a, Sch91b, Sch91c]. Incógnitas [Pan08]. Incommensurable [vdW76a]. incomparable [Bra09b]. incompletability [GG79]. Indalecio [AGSL76]. Independence [AH99]. indeterminate [Sma16]. Indétermination [Bré12]. Index [Ano74-53, Ano74-54, Ano75n, Ano75o, Ano76z, Ano76-27, Ano77-31, Ano77-32, Ano78v, Ano79-27, Ano79-28, Ano80n, Ano80o, Ano81-33, Ano81-34, Ano82m, Ano82n, Ano83l, Ano83m, Ano84o, Ano84p, Ano85l, Ano85m, Ano86p, Ano86q, Ano87p, Ano87q, Ano88s, Ano88t, Ano89p, Ano89q, Ano90o, Ano90p, Ano91l, Ano91m, Ano92y, Ano92z, Ano93q, Ano93r, Ano94k, Ano94l, Ano95p, Ano95q, Ano96b, Ano96c, Ano97b, Ano97c, Ano98e, Ano98f, Ano99b, Ano99a, Ano00b, Ano00a, Ano01b, Ano01a, Ano02e, Ano02d, Ano03n, Ano03m, Ano04l, Ano04k, Ano05n, Ano05m, Ano06j, Ano06i, Høy95, Mac76, Yok75, vM84, Ano76-35, Ano87-33, Ano88w, Ano93i, Ber87, Bot78, Bow84, Joh84, Moo03, Pye75, RFS89, Sal87, Sca94, Cas77, Kat92, Par92b, Ano74-27, Ano87i, Ano95j]. index [Ano95l, Sch91a, Sch91b, Sch91c]. indexes [Ano81l, Ano95k, Ric84, Row85a]. India [Ano85n, Ano88k, Gup76a, Ano74-50, Che12, De 95, Gup75a, Gup86, Gup88, Gup91, Mon11, Sin85a, Ano74m, Yan83]. Indian [Ano74m, Ano88k, Ano88m, Gup81, Tee84, Gup74, Gup76b, Gup78, Gup80, Gup85, Gup93, Hug74, Plo01, Sin85b]. Indiana [GG86, Ano74-67, Ano74-75]. indice [Bot78]. indices [Syl81, Tee84]. indifference [Car18]. indische [Rom99]. Individual [Sen10, SS09]. Industrie [Row85d]. Industry [DR19]. INED [Maz16]. inedita [Fre75]. inédites [Maz21]. inédits [Sen99, dMC86b, dMC92, Vil96]. inequality [Sen92]. Inexplicable [Fer98b]. inference [Dav75]. Infinite [Ano86h, Moo02, Cal92, Cog14, Dal91, Déc10, Haw81]. Infinitely [Lau87, Fis78]. infinitesimal [Blå17a]. infinitésimales [For10, Giu99]. Infinitesimals [AR24, Sel06, Chr84, Dom04, Lai75, Lok80]. infinitesimo [Bee09]. Infinitos [Aus16]. Infinity [Dau12, Wil97, DF18, Wat79]. Inflexions [Høy07]. Influence [For98, dS99a, Bie78, Deb13, GG77, Jam88, Sch82a, Dru91]. influences [Nor75]. información [Ano87r]. informal [Pet78b]. information [Ano83b, Lew04b, Row06b]. informes [Bun76b]. ingegneri [Vér01]. ingegnose [Coz12, Tra10]. ingenieurs [Ano07e, Sau09]. ingenious [Tra10]. Ingenium [Lo 85]. Ingo [Sch11]. Ingrao [Ano92n]. inicium [Fol09a]. initial [Sch77]. Inner [Shi94, Swe92]. Innovation [Hog03, Ano92-27]. inquiries [Pan03]. inquiry [Lê20, Ols76]. Insanity [Har77b]. insbesondere [Krö06, Lév99]. insegnamento [GG05b]. inserted [GG90a]. insieme [Gan85]. inspired [Che09, Zhu20]. Instant [Har77b]. Instantaneous [Erl97]. Institut [Ano87h, Ano89m, Ano89t, Ano90k, Ano90-33, Ano91q, Ano91s, Ano91-33, Ano92w, Ano92-27, BKM^{+}82, Bun76b, Che94, Her75a, Mar90b, Mon83, Ple78, Rem12, Ric84, Yus10, Ano84k, DF10]. Institute [Ano76x, Ano81p, Ano89-33, Ano94g, BD11, Eng81b, Eng81a, Fol76, Maz77, Nic13, Pur86a, Pur88, Pur90, Pur91, Tee78, Zit07a, Ano78x, Ano78w, Bat06, Ber04, DF10, MY99, Shi94]. Institution [Fen03, Fri74a, Kin74b]. Institutions [Fef97, Gra77]. Instituzioni [Roe15]. Instruction [Pre24]. Instructional [Ano74-55]. Instrument [AH16, Ano84-27, Ano95n, Sor13]. instrumentis [KO09, Lef10]. instruments [Ano74z, Gol82]. Insurance [San06]. integers [Tai14]. Integral [Jaë19, Gui87, LBB^{+}04, Lüt82, Phi79c, Jaë19]. integrala [Phi79c]. intégrale [LBB^{+}04, Maw06]. integration [Ano77o, Bar74, Dut84, Kno83a, Pie89]. Integrationstheorie [Kno83a]. integrity [Roy17]. intellectual [Cas80, Wis76, Ano88f]. intelligence [Gra84a]. intelligibility [Mah84]. intended [Wag83]. Intensional [GP91]. INTER [Bor89, Ano92a]. inter-I.R.E.M [Ano92a]. INTER-IREM [Bor89]. interaction [Bow87]. Interactions [How18, BM77]. interest [Mug87, TS78]. Intermediate [Rus80]. Internalism [Pyc84b]. International [Ano74-76, Ano76u, Ano76x, Ano77-42, Ano79-29, Ano79-30, Ano80k, Ano80-39, Ano81p, Ano82l, Ano82k, Ano83o, Ano85-30, Ano85-31, Ano86m, Ano86r, Ano88-31, Ano89-34, Ano89o, Ano90g, Ano90m, Ano90n, Ano90-35, Ano90-37, Ano92-29, Ano92-28, Ano93j, Ano93x, Ano94x, Ano08e, Ber75, Bie78, Boi89, Boo86, Bre91, Che87, Dau82a, Dau86, Dau93a, Dau94, Dav90, Dun89, Gar86a, Gar92, Hou89, Kno94, MP06, May75c, May77c, Mic88, Mug87, Obe78, Par03, Pin76, RW98, Rog77, Ros74, Scr79b, Scr82b, Scr84, Scr85b, Scr87, Scr88a, Shi93, Swe92, Ano76i, Ano86g, Ano87-28, Ano87-35, Ano87-36, Ano89m, Ano90h, Ano92p, Ano92-50, Ano93n, Ano93v, Bab77, Dau80, Luc18, Pei17, Taz18, Ano76-28, Ano79-31, Ano80p, Ano83n, Ano85n, Ano88u, Ano89r, Ano91n, Ano91o, Ano91p, Ano92-31]. International [Ano92-30, Ano93s, Ano93t, Mil93, Pre24, Scr82a, Scr85a]. Internationale [Mah78, Ano77j, Ano92-44]. internationales [Gis99b]. internationalisation [Fio17]. internationalism [CB21]. internationalization [BPL17, SS18]. Internazionale [Ano90c, DG90, Lam87, Ano92s, Ser90]. Interplay [PGN18, Taz18, SS18]. interpolation [Ger74, HZ85]. Interpretation [Lic97, Chr98, Coo10, FC22, Glu77, Gra79, Sai95, Wag83, Wat91, Chr98]. interpretatus [Net24]. Interpreting [AR24]. interrelations [Ano87-29].

Intersections [GG02, GG03]. introduce [Ken08]. Introducing [Sch06a]. Introduction [Ano89g, Bir96, Bir75a, Bir75b, Bir75c, Bir75d, Bir75e, Clu76, Dra84, Gil77, Gin83, Gra99a, Knu11, Nab04, PV22, Ste80, Ano78h, Ano82d, Bar11, Che14a, Coo95, Ken76c, Mac76, Mat74, Mue96, NLL08, Pan87, Par03, Sai74, Sch76b, Scr75b, Swe74, Tee84, VHK16, Vil96, Arc96, Cal92]. Introductory [Die75]. introduit [Lê21]. Introduzione [Ano85f]. intuition [DS19]. intuitionism [VHK16, Wil83]. intuitionist [Ton82]. Intuitionistic [Fra95]. intuitive [Høy83]. Invariance [Cog12, Gra08, Dau75]. invariant [Cri86, Cri88, Haw77b, Wol08, Pie86]. Invariante [Gra08]. invariantes [Pie86]. invention [Deb13, VV14]. inventioni [Scr03]. Inverse [Wei98, Dal88, Wei92]. Inverse-Square [Wei98, Wei92]. inverses [Rob80]. inversion [Abe79]. investigation [Lüt09a]. Investigations [Høy82, Dal85]. invisible [Ano92n, Ano74x]. invitation [Fow92a]. involume [Ano80n]. Ioan [SS10b]. Iowa [Eis75]. ipotetico [Ken78b]. ipotetico-deduttivo [Ken78b]. Iran [Ano93s]. Iranian [Bag97a]. Iraq [Mel09]. IREM [Bor89]. Irène [Kle17]. Iris [KS12, Tob10, Tob10, DR19]. Irish [All04]. Irrational [Bur92, Von93]. irreducibility [Joh84]. irreduzibilität [Joh84]. irresistibile [DF10]. irresistible [DF10]. irresitibile [Rem12]. Irreversible [Buc77]. Irrisolto [Ven02]. Isḣāq [dY92]. isāb [Dje98]. Isaac [Coh82, Duc16, Ait86, Dry14, Gab76]. Isabella [Ano81-35, DPY91, DPS^{+}02, DV07]. ISBN [Abe06, Abe09, AH10, AH16, Ale14, All09, And07, Arc09, Atz06, Aus16, Ayo06, Bay11, Bee09, Bee06, Bel07a, Bel08, Bel12, Ber07, Bla05, Bor14, Bor18, Bos03, BB04, Bra07, Bra09b, Bra09a, Bré12, Bre07, Bru07, Bul16a, Bul22, Cal08, Cap13, Che14b, Cog12, Cog15, Con06, Coz12, D’A10, Dav19, De 20, Des09, Dho11, Don05, Dru07, Dru12, Duc16, Dum15, DR19, Eva13, Fen08, Fer11, Fer04b, Fol12, Fol14, Fol15, For10, Fra06, Fra10, Gal12, Gar05, Gau13, Ges18, Gil15, Gis14, Gol13, Gra76a, Gra10a, Gra11, Gra13, GG05a, GG05c, GG07, GG08, GG10, GG13, Gra97a, Gra06, Gra08, Gra10b, Gra12, Gra18, Har06, Has04, Hei09, Her14, Hog03, Hog24, How18, Høy07, Høy09a, Høy09b]. ISBN [Høy24b, Hud12, Hug07, Hun07, Hun12, Jam13, Jur07, Kat07, Kat11, KS12, KS13a, Kle17, Kno07, Knu11, Krö06, Krö14, Krö15, Krö17, Lan07, Lef10, Lei09, Lew09, Lon21, Lor16, Lor18, Lüt08, Lüt09a, Lüt15, Maa07, Mal06a, Mar10a, Mar14, Mar09, Maw06, Maz21, Maz24, McK82, McL12, Mel06, Mos16, Mro16, Net24, Neu08, Nic13, Nor07, Oak12, Oli08, Pan08, Pan09, Pan18, Par07, Pec20, Pei15, Pei17, Per10, Plo06a, Plo06b, Rab08, Rec07, Rei15, Rem12, Rid88, Row06a, Row07, Row15, Row23, Sai05, San05, San06, San09, Sau06a, Sau09, Sau14, Sch11, Sch13b, Sch16, Sch17, Sch07, Sch05, Sch09, Sch12, Sch13c, Scr03, Scr07, Scr08, Sen06, Sen10, Ser18, SG09, Sid06a, Sid06b, Sid12, Sid13, SS07, SS08, SS17]. ISBN [Ste05, Ste09, Tat11, Taz09, Taz15, Thi06, Thi08, Tim09, Tur08, Van06, Wat09, Wil08a, Wil05, Wil07, Wol03, Yus10, Zep12, Zit07a, vM07]. ISBN-10 [Maz24]. ISBN-13 [Maz24]. ISBS [BT81]. ISHEM [Mic88]. Islam [Ber97, Che12, Ano89e, Kin86b]. Islamic [Ano92g, Ano93s, Ber04, DS99c, For10, Hog03, Kin86a, Moy17, Oak12, Amb08, Ano78q, Ano92i, De 95, Hog81b, Hog01b, Kin74a, Kin74b, Kin79, Mat93, Özd00, Yad80, Ano89h, Mor96]. islamica [Amb08, Oak12]. Island [AH10, Ano74s, Nic13, Sel91, SS86]. Islands [Asc95]. Isoperimetric [Fra92]. isou} [Auj86]. Israel [Ano92n, Mar09, Pan18, Enr76, Ano82o]. ISSN [Mal97]. issue [Ano81e]. Issues [Ano95z, Ano96h, Ano96i, Ano96j, Ano96k, Ano97h, Ano97i, Ano97j, Ano98k, Ano98l, Ano98m, Ano99g, Ano00g, Ano01g, Ano02k, Ano02l, Ano02m, Ano02n, Jah93, Ano79-38, Ano79-39, Ano79-40, Ano80-27, Ano80-28, Ano80-29, Ano80-30, Ano81-41, Ano81-42, Ano81-43, Ano81-44, Ano82u, Ano82v, Ano82w, Ano82x, Ano83r, Ano83s, Ano83t, Ano83u, Ano84t, Ano84u, Ano84v, Ano84w, Ano85w, Ano85x, Ano85y, Ano85z, Ano86x, Ano86y, Ano86z, Ano86-27, Ano87x, Ano87y, Ano87z, Ano88y, Ano88z, Ano88-27, Ano88-28, Ano89z, Ano89-27, Ano89-28, Ano89-29, Ano90-28, Ano90-29, Ano90-30, Ano90-31, Ano91z, Ano91-27, Ano91-28, Ano91-29, Ano92-39, Ano92-40, Ano92-41, Ano92-42, Ano93-29, Ano93-30, Ano93-31, Ano93-32, Ano94s, Ano94t, Ano94u, Ano94v, Ano95w, Ano95x, Ano95y]. issues [Ano03s, Ano03t, Ano03u, Ano03v, Ano04q, Ano04r, Ano04s, Ano04t, Ano05s, Ano05t, Ano05u, Ano05v, Ano06o, Ano06p, Ano06q, Ano06r, Ano07j, Mar10b]. Istanbul [Mat74, Yus10]. Istikmal [Hog86]. Istituto [Ano90f, Pep90, Pep92, Ser90, IN83]. istoricheskogo [Sti76b]. istorii [Ken75, Smi79b]. Istoriya [Ano77l, Ken76d]. István [Ora82]. Italia [Ano89-30, Ben95, Ano87-27]. Italian [Ano77q, For75, Amb08, Ano86i, Ano87-27, Ano91k, Ben95, Bot78, BDD^{+}88, BC04, DF10, FR83, Gan85, Gan86, Gue88, GN92, GN05, Høy06, Maf10, Mai07, Mai08, Mai09, Roe89, Ser06, Tra10, Ven02, Vér01, Arc09]. Italiana [Ano90f, Coz12, Taz09, Tra10]. italiano [Par97]. Italien [Ano89t, Bot80a]. italiennes [Ano89t]. Italy [Ano87l, Ano89-35, Ano90c, Ano92s, Gia90, Ken76b, Pep90, Pep92, Roe89, Ano87-27, BF23, Ben95, Fin11, Mar99, Mar04, PR22, Van24]. Itard [Ano85o, Jon76a]. Iterative [Plo96, Gre88]. Ithaca [Ano81u, Tru75]. Itinéraires [Gis14, Auv13]. itinerary [Sch13c]. Ito [Syl83]. IV [Ano80p, Ber87, Hog85b, Mos16, Mue96, Pec20, Pei22, Sal87, Ano87i, Lo 85, Scr03, Ano77-31, Ano77-32, Vit96]. IV. [Tob81]. IVA [Ter89]. Ivan [Bar74]. Ivanovich [Phi79a]. Ivo [Cif00]. Ivor [Bra12, Gra98b, Par92a, Ano74-59, Ano80q, DGL^{+}15]. Ivory [Bos02, She94]. IVth [Ano84-27, Tob81]. IX [Grö86, Mat75, Ano77i, Bay11, Bir80, Das81, Gra10b, Haw81, Sch12, Syl83, Van06, Vit96]. IXe [Bre07, For10, Giu99, Ai04]. IXe-XIe [Bre07]. Iz [Ken75]. Izd [Dav76, Sel74]. Izdatel’stvo [Ano81t, Smi79b, Dyc75a, Dyc75b].

J [Ano74u, Ano74y, Ano74-73, Ano76-44, Ano77i, Ano77j, Ano77h, Ano77r, Ano79m, Ano81k, Ano81q, Ano81m, Ano87b, Ano89e, Ano89k, Ano89l, Ano91e, Ano92k, Ano95f, Arc96, Bar78b, BKM^{+}82, Bee14, Ber84b, Bie74a, Bie78, Bré12, BS77b, BT84, Bun74, Cha76, Cor96, Das81, Dav19, Dra84, Ebe96, Enr76, FK89, Fra95, Gil78, Gin83, Gra77, GG94, GG98, Gra98b, Gre85, Gui24, Har77a, Her75a, Hes90, Hil83, Hog85b, Jon76a, Ken82, Kje17, Kno07, Lan07, LY79, Lem74, Lew99, Lo 85, Maa07, Mac76, Mah78, McK82, Neu08, Nic13, Pat75, Plo06a, Sch07, Sch00b, Sen84, Ser18, Ste09, Str82a, SY90, Tro83, Vek75, Vil96, Vit96, Wil79a, Wil77b, vM84]. J. [Bur75, Pie83, RH11, Sch95, Scr75a, Scr85c]. J.-C.-1644 [Bré12]. J.-L [BKM^{+}82]. J.-P [Ano92k]. J.J.M [Cri24a]. J.L [Moy17]. Jacob [Ano92j, Nag91b, Nag91c, Eli18, Eli24, Mat14, Nag91a]. Jacobi [McC94, Pul98]. Jacobson [Par91]. Jacopo [Høy06]. Jacqueline [Bee06, Gra10a, Lon21, Tat11, vM07, Neu15]. Jacques [Che94, Fol15, Rei15, Aud12, Gra76b, Krö14]. Jahnke [Ano82c, Ano84k]. Jahr [Has10b, Has10a]. Jahre [Bie74b, FK89]. Jahren [Ano89k]. Jahrestagung [Ano87m]. Jahrgang [Str81]. Jahrhundert [Ano86h, Ine94, Row85d, Sch83, Tob90, Böt12, Fol83a, Ano88i, Gra90, Pei15]. Jahrhunderts [Lem74, Ano84k, Fol87]. j’ai [LP10]. Jain [Ano83n, Ano85n]. Jaina [Gup75b, Ano79u]. Jakob [Gui96, Syl03, BFG94, Yca16]. Jamaica [Cra98]. James [Gar05, GG90c, Row83b, San05, SS10b, Ano79-44, Cra09, Ken79a, Ken79b, Nor03]. Jami [Che94, Van98]. Jan [Ano87g, Atz06, DS99b, Hoc96, Kit78, Row94, Alb94]. Janet [Tat11]. Janis [Ano89i]. János [Kis99]. Januar [Spa89]. January [Ano76h, Ano78t, Ano92f, Gau13, Moz13]. January-June [Gau13]. Janvier [Ano89-30, DG90]. Jaouiche [Ano79l, Roz90]. Japan [Ano91o, Mic88, Fuk87, Yos75, Yos76]. Japanese [iÔ75, Ano78-29, Mar90c, Ung23]. japonaises [Ebe96]. Jardine [Kat07]. JAS [Pye74]. Javier [Ano93i]. Jayadeva [Sel75]. Jean [Ano77k, Ano79m, Ano84g, Ano87h, Ano89g, Ano95k, Ano95n, Bie78, Bor14, Bré12, Dau93b, Dau12, Fer98a, Fer04b, Fre76, Gra91a, Gra96, Høy07, McL12, Moo77, Obe78, Ros74, Sid06a, Wil08b, Ane88, Ano85o, Gra01, Rom13, Wan22, Fra04a, Wil08b, Kle17]. Jean-Baptiste [Wan22]. Jean-Claude [Ano87h, Bor14, Bré12, Dau93b, Fre76]. Jean-Henri [Bie78]. Jean-Michel [Dau12]. Jean-Paul [Ano77k, Gra96]. Jean-Pierre [Fer98a, McL12]. Jeanne [AH10]. Jed [Bul16a]. Jeff [SG04]. Jefferson [Tat87]. Jens [Lev97, Mel06, Høy93]. Jérôme [Gis14]. Jeremy [Coo99, Lê24, Neu08, Oli08, Ric82, Thi06, Wal17]. Jersey [Bul16a, Sal78, SG04]. Jerzy [Ken76a]. Jesper [Ano84j, Ano93g, Atz04, Nor07, Mic23]. Jesseph [Ano95g]. jésuites [Ano92-32]. jesuiti [Ano93j]. jeux [dMC92]. jew [Ano82d]. Jewish [Dav75, SS17, Has10a]. Ji [Lê24]. Jihe [Mar93b]. Jim [Ano84f, Coo99, Gis96]. Jiska [SS17]. Jiu [Shu92, Vol10]. Jiuzhang [Yin11]. Jñanaraja [Mon17]. Joachim [Ano89j, GG10, Gra18, Lew09, Mue80]. Joan [Gri82, Lew90]. João [Bra09a]. Jock [Ple78]. Joe [San03]. Joel [Syl07]. Joella [Zas83b]. Johann [Ano92j, Ber04, Bul22, Hes90, Sch82b, Vol16, Gra78c, Pon78].

Johann-Heinrich [Pon78]. Johannes [Ano81i, KO09, Lef10, Høy24a, Cif00, Don05, Gin83]. John [Ano77p, Ano81l, Ano82f, Ano89d, Asp83, Bru77, BT81, Cal92, Clu76, Coh82, Dav90, Enr80, Gra77, Gri82, Has10b, Has10a, Kin86b, Lew95, Lew98, Maa07, Moo75a, Mos16, Par07, Ste01, Ste05, Wil91, Wol91, Ano75x, Big19, CS81, Cra98, Cra00, GG05d, Gra02, Hal87, Has10a, KO09, Kre03, Mai07, MS05, Réd05, SB12, Ste12, Ste11, VV14, Von88, Zun02, Bir83, Bir97, Pan09, Rec07, vM07]. Johns [Abe06, Ano88e, Bel07a, Bra12, Hel83, Hud12, Mon17, Per10, San05, SG09, Wil07]. Johnson [GC82, McG74]. joint [vD09, Ano75p, Pye74]. Jones [Ano74r, Plo06a, Pri80, Wil77b, Eis78]. Jongmans [Gis99a]. Jordan [Bil85, Joh77]. Jordanian [Mat74]. Jordanus [Cor16, Hug89, Gra83]. Jornadas [Ano90-32]. Jos [Shi88]. José [Gar98, Lê24, Oli08]. Josef [Zla23]. Joseph [Ano88g, Haw81, Kat92, Nor03, Rid88, Van06, Ehr11, Fra87, Hog10, Jam97, Lüt82, Neu89, Ano93g, Bru77, Her75a]. Jost [FLT16]. jour [Ric91]. Jourdain [Ang79a, Ano74-56, Ano79h]. Jourdian [Ang79a]. Journal [Ano79-33, Art13, Dum15, Ano78-28, Ano78-31, Ano79-34, Ano80t, Ano80u, Ano81-36, Kno17, Cri04]. Journals [Ric98, vD09, Ano74j, BM74, Des07, Ehr18, Gra75b, Luc18, Mil93, PGN18, SS18, Taz18]. Journée [Dho88b, Koe94]. Journées [Ano83o, Ano85p, Ano88v, Ano90q, Ano91q]. Journey [Wil97]. jours [Pie86]. Jr [Ano81k, Ber74, Jon77c]. Jubilee [Gup77b]. Juden [Sch16]. Judgments [Has10a]. jüdischen [Aud12, Has10b, Has10a, Krö14]. Judith [Dho85]. Judy [AH10]. Juillet [eFD88, Dho91, Dho90, Mar90b]. Juin [Ano83o, Ano88o, Ano90q, Ano92-50, Hei81, Hou89]. Jules [Ano84f, Gis99b, Gra99a]. Julia [Ale14, Aud11, Gam96]. Julio [Mor96, Hog24]. Julius [Fra16, SM18, Wie23]. July [Ano76-29, Ano83k, Ano87-31, Ano87-36, Ano88u, Ano89o, Ano90e, Ano90m, Ano90-33, Ano91j, Ano91-33, Ano92e, Ano92w, Ano92-47, Bee14, Dem74, Dho90, GG92b, Gra02, Pyc84a, Row90b]. June [Abe91, Ano75k, Ano76i, Ano87n, Ano87-31, Ano89-35, Ano90u, Ano90-35, Ano91h, Ano91n, Ano91u, Ano92q, Ano92u, Ano92-28, Ano93a, Bow87, DGL^{+}15, Dho83, Dia99, Fra10, Gau13, Hog15, Kat11, Kat14, Nag88, Ott88, Pep92, dBR76, Row89, Ano74-47, Ano85i, Ano24f, Kno88b]. Jungnickel [Ano88f]. Juni [Mai87]. Junior [iÔ75]. Jürgen [Ano88i, Gar05, Her14, Sch19]. Jushkevich [Ano77m]. Juskevic [Ter89]. justice [Ano74l]. justification [Chr84]. Justifier [FN11, Krö15]. Justinian [BT81]. Justus [Hog10, Rad00].

K.-G [Ano74v]. Kāmil [AD87]. Kāshī [Bag97b]. Kātyāyana [Ano88m]. Kaedmon [Cha76]. Kagan [Tot79]. Kahane [Ano92k]. kak [Dyc75b]. Kaluza [Ale04]. Kan [iÔ75]. Kanaev [Dyc75b]. Kann [Row23]. Kannenberg [Law05, Law05]. Kanovei [Plo06b]. Kansas [Ano79k]. Kant [Moo03, Win82]. Kantian [Gre83]. Kantons [Dho88a]. Kantor [Dau12]. Kanzan [Won22]. Kao [LY79]. Kaplan [Gra13]. Kappa [Ano74-66]. Karaji [Mat74]. Karamata [Nik09, Nik18]. Kardinäle [GG07, TC05]. Kardinalität [TC05, GG07]. Karen [Kje17, Neu08, Nor03]. Karin [Ano81i, Ano95h, GG13, Thi07]. Karine [Gis96, Lê24, Vol10]. Karl [Cas80, Coo95, Jur07, Mar06, Row85d, Sch09, Dug76, Gra80b, Ken77, MB20, Phi79b, Por04, Has04, Hun07, Ken76c, Ken76b, Sch19]. Karl-Heinz [Row85d, Sch09]. Karlsruhe [Rog77]. Karmi [Ano81p]. Karpinski [Jon76b]. Kartoniert [Ano89f]. Kashi [Ham80]. Katahiro [Ebe96]. katholischen [GG07, TC05]. Katrin [Haf16]. Katscher [Scr03]. Katz [Arc96, Ayo06, GG98, Gra98b, Kje17]. Kaufmann [Ano89f]. Kaunzner [Jay83]. Kaza’n [Bie74b, Bie74b]. Kazimierz [Moo82a, Moo82b]. Keble [Ano83f, Ano85-28]. Kegan [Bun76a]. Keith [Cro77]. Keller [Wil08a]. Kelly [Por05]. Kelvin [Pet09a]. Kempisty [JMT19]. Ken [All04, Sid12]. Keng [Ano91g]. Kennedy [Ang88, Cib76, Mon75, Tee78, Ume76, Van05, vD10]. Kenneth [Ano74o, Bid74, Bra07, Ano77b, Ano77c, Ano80r, Ano84q, Ano08e, Ano10k, DR78, Dra78, Enr84, Fol94, JET84, Lew04b, Par05b, VK05]. Kepler [Don05, Gin83, Bra89, Car22, Car24, Mar00, Nau03, Sac20, Ano91e, Ano91f, Wil05]. Keplerian [Mae89]. Kepleriana [Don08]. Kerle [Row06a, Wue05]. Keynes [BGG06a]. Khāwarizmi [CH90]. Khalil [Roz90]. Khan [Ano88k]. Khaqani [Ham80]. Khinchin [VHK16]. Khosrow [Bag97a]. Khwarizmi [MP06]. Kidwell [SG09]. Kiev [Rei76]. Killing [Haw80]. Kim [Ano05d, Ano06b, Ano06c, Ano06d]. Kimé [Krö06]. kin [Høy02a, Mel06]. kind [Buc77]. Kinds [FT99]. kinetic [Buc77]. King [Ano87n, Ano89h, Hog86, Hog95, Hog03, Bar11, Big19]. King-Hele [Ano87n]. kinship [Gre77]. Kinyon [Bra07]. Kirchenbibliothek [Fol09a]. Kirk [SY90]. Kirsti [Gra09b, Toe96]. Kiryu [Mic88]. Kishinev [Dav76]. Kitaie [Vol91]. Kitcher [Ano87k, Ano91d]. Klasse [Scr03]. Klaus [Krö06, Ric84, Sch90b, Che14b]. Klein [Kot79, Row85b, Row90a, Row19, Ano89k, Gis99b, Haw84, Hel23, Lê15, Row83c, Row85e, Sch13a]. Kleiner [Mar09]. Kleinert [Pei22]. Kleins [Row85d]. Klose [Ano85e]. Klostermann [Ano77h]. Klügel [Vol06]. Kluwer [Atz06, Sau06a, Sid06a]. knew [Rab10]. Knobloch [Ano92l, Bar82a, Bar82b, Bla05, Don05, Blå17b]. Knorr [vdW76a, Fow98]. Knot [Epp95]. Knotentheorie [Sch02]. Knowledge [BG21b, Mac78b, Ano74x, Böt12, Rom10, Gra98a, Ano87k, Ano88l]. Known [RT02, Hug74, Lew95]. Knudsen [Mon17]. Koblitz [Ano86d]. Kochina [PP78]. Kodama [iÔ75]. Koenigsberger [Hou87]. Koepke [Plo06b]. Koerner [Fol87]. Koetsier [Ano95j]. Kolmogorov [Ano81o, Gra94b, Gra94c, Gra97a, You83]. Köln [Rom04]. kombinatorik [Bar82a, Bar82b]. Komitees [Str78]. Kommentar [Rom04]. Kommentare [Ine88]. kommentiert [Ano92j, Bul22, Scr03]. Kommission [Scr03]. kommt [Scr81]. Kompas [For75]. Kong [Ano93g, Gra94a]. König [Row85b, Sch19, Ano86h, WM14]. Königliche [Ano74q]. König’s [Fra16]. Konstruktionen [Sch02]. Kontext [Krö14, Pei15, Che14b, Sch17]. Kontexte [Sch02]. Kontinua [Ano89k]. Kontinuitätsprogramm [Arn88b].

Kontinuum [Mue80]. kop [Lin74]. Kopecks [GG80, Ano81o]. Koreuber [Sch17]. Korfu [Ano92-29, Ano91h]. Korrespondenz [FS10, GG07, Jen86, TC05]. Kosmann [Cog12, Gra08, Taz15]. Kosmann-Schwarzbach [Cog12, Gra08, Taz15]. Kosmos [H^{+}08, Sch11]. Kostant [Ale04]. Kothmann [Don08]. Kotz [Pla76]. Kovalevskaia [Ano86d]. Kovalevskaya [BN22, KS13a]. Kovalevsky [Coo86]. Kowalewskaja [Ano95m, Böl93, Coo95]. Kowalewskajas [Böl93]. Koyré [Gab76, Koe94]. kr [Ano77g]. Krankheit [Fri91]. Kranz [Lef10]. Kreises [Str81]. Kreisquadratur [Cri17]. kreisteilungspolynoms [Joh84]. Kretkowski [Pio85]. Krise [Sch90b]. Kristiansand [BJ89a]. Kristín [Kat11]. kritischer [Bos77]. Kronberg [Ken76c]. Kronecker [SS89, Edw78, Gan86]. kubischen [Scr03]. Kuhi [BV05]. Kuhn [Kje00, Meh76]. Kuibyshev [Roz75]. Kujang [Yin11]. Kulturbote [Eck13, Sau14]. kulturellen [Sch17]. kulturgeschichtliche [Lév99]. Kultusministeriums [Row23]. Kummer [Edw77b]. Kungl [Tho76]. Kunitzsch [Gol76, Van05]. Kunst [Ano84e]. Kuratowski [Moo82a, Moo82b]. Kurt [Ano79o, Ano79-42, Ano81i, Cal75, Seg90, Str78, Vol81, Ano83d, Ano89s, Ano93u, Daw86, Fol83b, Fol86, FS03, GG79, GG90b, Pec20]. Kurt-R [Cal75, Str78]. Kurt-Reinhard [Ano79-42, Ano89s, FS03]. Kusch [Gra98a]. Kushyar [Van98]. Kuznetsov [Dyc75a]. Kyoiku [Ano78-32]. Kyoritsu [Sid12]. Kyoto [Ano74-76, Ano91o].

L [Abe79, Abe93, Ano74r, Ano74t, Ano74-29, Ano78g, Ano84i, Ano86w, Ano89e, Bar77, BKM^{+}82, Ber74, Bor18, Bur08, BJ89b, Dav75, Dav19, Dil81, Eva98, Fol09b, GG08, Gra08, Gre77, Har77a, Hel83, Hes90, Kat20, Ken78b, Ken82, Kil83, Lew90, Lew98, Lew04a, Lo 84, Mal06a, Nee79, Par05a, Par07, Ped75b, Pin76, Pla76, Ser18, Ter89, Too80, Tur08, Wil77b, Zas83b, Rom04, Ano94h]. l-‘Abbas [Rom04]. L. [Fra95]. Lūqā [Hog85b]. Labban [Van98]. Labor [Bec01, Bab77]. Laboratoire [Ano91r]. Laboratory [Ros74, Mai21]. Labours [Pet09a]. Lacroix [Dom08, Tat86, Bra09a]. lacuna [Ber77, MF10]. Ladie [AB09]. Ladies [Per79]. LaDuke [AH10]. Lagrange [Bos77, GG87b, Bec99, Eng80, Fra87, Fra92, GG88, LMP01, Pul98, Ros89]. Lagutinskii [DLS98]. Laird [Net24]. Laisant [Gis14, Auv13]. Lakato [Ano95j]. Lakatos [Ano81l]. L’Algèbre [Høy07]. Lalla [Tee84]. Lam [Mar95, Nee79]. Lambert [Bie78, Obe78, Gra78c, Vol16]. Lamberts [Bul22]. Lambrecht [Ano91-31]. Lana [Ano88r]. Lanczos [Bla76]. land [vM84]. Landau [CB21, Wef89]. Länder [Bus76]. Landesbibliothek [Bar78b]. Landmark [GGC^{+}05, SS07]. Lane [Hal84]. Lang [Høy90b]. Langevin [Maz21]. Langins [Ano89i]. Langley [Hym78]. language [Bur92, Her75b, Mas08, Pet78b]. languages [War24, Ano81p]. langue [Ano77j, Dra84]. Lanka [Jay24b, Jay24a]. l’Antiquité [DG90]. Laoghaire [Ano87e]. Laplace [Gil15, Fri74b, GP91, Hah11]. L’Archivio [GG05a].

Large [Sen92, Jay24a, Jay24b]. L’arithmétique [Sch76a]. Last [Kat07, Mar90c, iÔ75, RL16, SGJ05, Xu05, LP10]. late [Che17, CO13, Dom04, Mar20, Mes94, Per80]. lateinische [Ano91u, Rom99]. lateinischen [Rom04]. lateinscher [Gol76]. later [Daw16a, Mah84]. Latin [Ano87-31, Ste80, Syl83, Too80, Vit96, AG77, AA88, Arb86, Eli18, Fri81b]. Latina [Ano85j, Ano87d]. Latino [Fol14, DN12]. Latinus [Ray12]. latitude [Tat87]. latter [Tat75]. lattices [SS22b]. Laufbahn [SD94]. Laugwitz [Wol03]. Launching [Des07]. Laura [Ano81r, Ano85f]. Laureate [Wil07, Cri06]. Laurence [Ano84h]. Laurent [Sch12, BPL17]. Laurier [Høy90a]. Lausanne [Rei15]. Law [NB77, Sen92, And83, Nau03, NB77]. Lawlor [McK82]. Lawrence [Ano77q, Ano79k, How18]. Laws [Cog12, Fer00, Cro75c, Cro75b]. Lay [Mar95, Nee79]. Lazare [Tat75]. Leader [Fra10]. Leamington [Gra02]. Learning [Ano74w, Bra12, Sel90, Ste01, Böt12, Gur24, Nos12, Che10]. least [Ger74]. Leather [Gil81, Gil79]. Leben [Ano95m, Gra10b, Dho88a]. Lebens [Sch19]. Lebesgue [Kno83a, Maw06, Kno83a, Pie86, Ano77n]. l’École [Bil85, Dug82]. Lecture [Ale14, Ano74-58, BF89]. Lectures [Bir96, Cas80, Daw16b, Mar99, Row07, Ano81u, ESH^{+}13, HEH^{+}04, Neu88, VK05, Wil83, Bra07]. Leeds [GG90c]. Leendert [DS97, Fre93]. Leffler [Vol02, Dau80, Tur13]. left [Imh21]. Legacy [Fol78, Gra99b, Pet09a, Ano87n, Ano87s, Edw78, Gab79, Gui24, Mar06, Neu88, Neu93, Por19, Sch86b, Par89]. Legendre [AG77, Gra86, Sti77]. legitimate [Hym78]. legs [Gil77]. Lehigh [Tim09]. Lehrbücher [Toe96]. Lehre [Mue80]. Lehrers [Bie74b]. Lehto [Gra10b]. lei [Ano74p]. Leibnitiana [Pet82, Bar82a, Bar82b]. Leibnitz [Aga75]. Leibniz [Ano77j, Ano77-42, Ano80b, Ano89-34, Ano90l, Ano90-32, Ano90-35, Ano92l, Bar78b, Bar82a, Bar82b, Bee09, Bla05, Bos77, Cri17, Fel78, Gar05, Her14, HB11, Lam87, Nag88, Pet82, Vil96, Ano80g, Ano84-29, AR24, Blå17a, Chr84, Deb13, DB23, Fel78, H^{+}08, HNE21, KB74, Kno06, Mug87, Ott89, Pet78a, RP19, Str23, dMC86b, dMC92, Ano77h, Lüt20, Sch11, Vil97, Pet82, Tru75]. Leibniz-Archiv [Bar78b]. Leibniz-Society [Pet82]. Leibnizian [Blå12, Gro87]. Leiden [Cri24a, Hog03, Hog24, Mat74, Sai05, Pye76]. Leiden-Boston [Cri24a]. Leipzig [Ano74v, Ano89k, Ano93x, Bos77, GG13, Row85b, Sch96, Sch90a]. Leitão [Ges18]. L’élite [Dur21]. L’émergence [Ano91s]. Lemma [Cam78, Pes01]. lendemains [LBB^{+}04, Maw06]. Lengths [Høy02a, Mel06]. lenguaje [Pan08]. Lenhart [Hog90]. Leningrad [Ano76-34, Ano79p, Ano81t, Ano89-33, Dyc75b, Sel74]. Leningradskoe [Ano81t]. Lennart [Plo06a]. L’Enseignement [Gis18]. Lenses [Dij04, Atz06]. Leo [Bra17, Sau06a, Taz09]. Leonard [Fen97, Pat75, Spi00a, Spi00b]. Leonardo [Glu76, Han07, Ray23, Spi00b, Sch96]. Leonhard [Ano81x, Cal92, Cal08, Spa11, Ter89, Ano84-28, Cal96, Fel83, Fel07, Fri81b, Pet12, San07, Spa11, Ano77m, Bra09b, Dho88a, Scr08, Ter89]. Leonhardi [Bar77, Pei22, Rid83a, Rid83b, Sch00b]. Léopard [Cav95]. L’Eredit [Oak12].

L’eredità [Amb08]. Lernen [Böt12, Pei15]. Lerner [Mac76]. Leroy [Ano84f]. Lesevergnügen [Sch96]. Leslie [Ver13, Cra00]. l’état [Ano95i]. Letter [Ano02f, Bag97b, Bel07a, Gui99, Hal90, Kno17, Ano74-59, Ano83b, BS06, BN22, Böl93, Kc74, Müü77, Müü75, Pan87]. Lettere [Oak12, Sch99a, Ser90, Taz09]. Letters [Del05, Dru07, GG94, Lüt08, Abe01, Ano84f, Ano84n, Bos10, Bot80b, CWW17, GG90a, Hal84, HP20, LBB^{+}04, Réd05, Nor03, Rec07]. lettre [HS15]. Lettres [Ber87, Eva94, Fol14, Maw06, Sal87, Sim96, Gra99a, LBB^{+}04, Mar93b, Vol02]. letzte [Lau92]. leur [Enr76]. L’Europe [Coo99]. level [Ken08]. Levi [Goo77, Cog16, Mos79, NT05]. Levi-Civita [Goo77, Cog16, NT05]. Lewin [San06]. Lewis [Dru07, Dru12, Law05, Ped75b, Row85c, Abe76, Ano75q, Ano79b, Ano81b, Ano81c, Ano84a, Ano85b, Ano85a, Ano86b, Ano87b, Ano87a, Ano88c, Ano88b, Wil08c, Abe09]. Lexikon [SS08, Tob06]. Leyden [Pin76]. Lezioni [LD22]. l’Histoire [Gra01]. l’homme [Ped77]. L’hypothèse [Pan88]. Li [Ano89d, Ano92-31, Dje98, Xu07]. liar [Ano74-29]. Libbrecht [Yon75]. Libby [Bel08]. libellus [Van80]. Liber [Høy24b, Sch96, Ses14, Ste80, Bel05, Fol15, Han07, Moy19, Spi00a, Spi00b]. Liberal [AH99]. Libermann [BKM^{+}82]. Librairie [Ano77j, Ano77q, Bre07, Cor96, Ebe96, Gol82, Kno07, Mac76, Mah78, Vil96]. Library [Ano77r, Bow84, Buc80, CH90, DS99c, Eve76, GG90a, Lei79, Bot80b, Coh82, GG75b, Luc18, Vag82, Ano79i, Luc18, Roz75, GG90a]. Libri [Ken76b, Rem12, Rom13, DF10, Ehr11]. Lie [Cog15, Goo77, Haw77a, Haw77b, Row90a, ET14, Fri91, Haw80, Wüs02, Kot79]. Liebermeister [Ine94, Ine94, Sen94]. Lies [Fri91]. Liévano [AGSL76]. Life [Ale04, Ano79-31, Ano90f, Bag97b, DR19, Dyc75a, Her75a, Hun07, Jam13, Nag91c, Nic13, Nor03, Pet09a, Sch19, Soi09, Wat09, Ano82p, Ano87e, Ano87l, BM74, But81, Cha76, DF21, Fis82, Gri82, GC82, JET84, Mai07, Meo04, Nag91b, Par89, Pep75, Por04, RH11, Row83b, Wee91, Wil08c, Abe09, Lon21, Ang88]. Light [Beb96, GG05c, Was20, Chi22, Cul07, Fef98, GG75d, Rai92, Wil76]. lignes [Vol16]. like [Cor16, Tob10]. Liliane [Gol13]. Limit [Bel12, Cha76, Fis11, Pou01, Erl03, Moo08]. limitation [Ano89c]. Limitations [Ano85-29]. Limite [Ano81n]. Limited [Ano87e]. limites [Maz16]. Limits [Lor18, Dom04, Kno06]. lincei [Ano82-28]. linchpin [Cri04]. Lindegaard [Mon17]. Lindgren [Bay11]. Lindsay [SG09]. line [Ano88h, Ano93f, Gil81]. linéaire [Bré12]. linéaires [Cog15, Spi00a, Spi00b, ET14]. Linear [Hud12, LYK89, Ano79f, Ano79v, Ano83j, Har11, Jon77c, Moo95, Sch77, SM18, Spi00b]. lineare [Ano89k]. L’infinito [Lam87]. link [FK90]. L’intégration [Tou03]. Lionville [Ano93g]. Liouville [Ehr11, EF20, Lüt82, Neu89]. Lipschitz [SS89, Sch86a]. Lire [Ano85f, Isr85, Ysh90, Ken76b]. lisant [Kno88a]. Lisboa [Ges18]. Lisbon [Rod88, Sar08]. Liscia [Cri24a, Don08]. Lise [Gra12]. list [Ano92i, GG90a, Roz76]. L’istikmal [Dje97]. lists [Wal75]. Lit [Van80]. literarischen [HB^{+}13]. literary [HB^{+}13]. Literature [WDB96, Dav75]. Liu [Fu91, SS86, Wag79, Wag85]. lived [Ene18]. Lives [Gra13, Ano86d, KK11]. Livia [GG05b, Gra99a]. Living [OS19]. livre [AD87]. livres [Mue96]. Livros [Lei09]. Lizhen [Lê24]. LLC [Wat09]. Lloyd [Har77a, Law05]. Lobacevskij [Bie74b]. Lobatchevski [Tot79]. Lobatschewski [Ano89k]. local [Taz18]. locus [Ehr18]. L’Oeuvre [Eva94]. Loève [BKM^{+}82]. logarithmic [Cra03]. Logarithms [Bur01, NLL08, War08]. Logic [Ano75e, Ano80l, Ano88-29, Ano89c, BM96, Daw16b, Dru85, Dru12, Fer04b, Fra95, GG02, GG03, Gra97b, McL12, Moo97, OS19, Par02, Ven02, Wei98, Abe10, Aga75, Ane87, Ang79a, Ano79y, Ano81l, Ano86f, Ano87-35, Ben95, ESH^{+}13, Fef98, Gra94b, Gra94c, Moo77, Pan03, vH67, GG05c, Sca94, Ano74t]. Logica [Ven02, Bee09, Ben95]. Logical [Abe09, GG75d, Pan03, Wil08c, May75a]. logician [Put82]. logicians [OS19]. logika [Ano81o]. logique [BKM^{+}82, Hog01a, Pan03]. logistika [Ano78s]. Logos [Ano74t, vdW80]. Loi [Gre85]. lois [Gra08]. London [Abe06, Abe09, AH16, Ang79a, Ano74s, Ano74u, Ano74t, Ano74-29, Ano77p, Ano78g, Ano81l, Ano84i, Ano85h, Ano87n, Ano88e, Ano89h, Ano90e, Ano93g, Ano95g, Bel08, Bel12, Bir83, Bla76, BT81, BT84, Bun76a, Cas80, Cog12, Coh82, Con06, Coo96, D’A10, Des09, Enr80, Eve76, Fen08, Fol15, For10, Gra94a, Gra11, Gra13, GG90a, Gra98a, Gri76, Gur24, Haw81, Hel83, Her75a, Jon76a, Jon85, Kat92, KS13a, Ken76a, Law05, Lei79, Moo75a, Neu08, Par89, Pla76, Pye75, Rec07, Ric82, Rus75, Sal78, San03, Sau06a, Sch07, Sch12, Ser18, Sid24, Tan74, Tru75, Ung85, Vek75, Zep12, vM84, Bla14, Del06, RWG95, Ric96, RW98, RW03, Sor13]. long [Ehr15, Gil77]. long-term [Ehr15]. Longitude [Lew99]. Longo [Van80]. Longue [AD02]. Loo [Ano91g]. Loo-Keng [Ano91g]. look [Scr83, Wil80]. L’opera [GG05a]. L’opere [Tra10]. Lorch [Ber07, Van05, ZvDF22]. Lord [Bel07b]. lore [Bus76]. Loren [Dau12]. Lorenzenhof [Ano92b]. Lorey [Ano87s]. Lorraine [Lüt15, Sch91a]. Losano [For75]. Lost [Mro16, Ano77i, War15]. Lothian [San06]. Louis [Ano81m, Ano84i, Fra87, GW94a, GW94b, GG14, Jon76b, Gra94a]. Louise [Ano89l]. Louvain [Kit81]. Louvre [Cav85]. Lovelace [HMR17, Hym78]. Lowe [Ano88e]. Lower [Bin90, Bin91]. Ltd [Hal84, Jon76a, Kat92]. Lubos [Ano90r, Bir74]. Luca [Rey06]. Lucas [Sch76a, Déc14]. Lucasian [Tat92]. Luciano [Gra99a]. Lucienne [Ano77n]. Lucio [Net02]. Ludo [Bel05]. Ludolf [Her14]. Ludolph [Hog10]. Ludvig [Bir96]. Ludwig [Sau10]. Luigi [Ano85g, Mar14, Pan09, Sch13c, BD11, Men86]. Lumières [Alf11, Gol13]. Luminy [Che87, eFD88]. Lun [Ano89d]. lunar [Ano74-27, Cha77, Gol80]. Luncheon [Ano92-33]. Lund [Ano74s]. Lundi [Ano83o]. Lune [Lan96]. Lüneburg [Sch96]. L’univers [Mal13]. lunyue [Mar93b]. Lutwidge [Abe01, Abe10, Dru07, Dru12]. Lützen [Ano84j, Ano93g, Atz04, Mic23, Nor07]. Luxembourg [Ano92q]. Luzin [For98, Phi78]. Lvov [Mro16]. lx [Bar77, Taz09]. LXXIV [Net24]. LXXIV-474 [Net24]. LXXV [Ano77r, Bar78b]. lxxxv [Lo 84]. Lynn [Wil75b]. Lyon [Row15]. Lysenko [Phi79a].

M [Ano74m, Ano74u, Ano77h, Ano78j, Ano79l, Ano81t, Ano82c, Ano87d, Ano88n, Ano88k, Ano89c, Ano92k, Ano95f, Ano95g, BKM^{+}82, Bie74b, Bir77, Bow84, Dav90, Enr76, Fen08, Gil75, Gin83, Gra80b, Gra10b, Gre85, Gui24, Haw96, Hog86, Hun07, Ken76a, Kot79, Lo 85, Mat74, Phi79b, Pys81, Row90a, Sch91b, Sch91c, Sch07, Sen06, Spi76, Str82a, Ter89, Vit96, Wil75b, Wil77b, Zas83a]. M. [Bie74b, Roe15, Wil79a]. M.-M [Ano88n]. M.I.T [Ano74y, Bir83, Wal94]. Mānava [Ano88m]. MA [Ano84d, Ano92n, Asp83, Con06, Fer04b, Par89, Sch91b, Sen06, Zit07a]. MA/London [Con06, Par89]. MAA [Ano76h, Lor18]. macchina [For75]. Macdonald [Vek75]. MacHale [Ano87e, Was20]. machinarum [Rid83a]. Machines [Ros74]. MacKenzie [Ano88l]. Maclaurin [Fer98b, Hal84]. MacMahon [Llo90]. Macmillan [Cas80]. Macquarie [Ano79z]. Macrae [Bir97]. Macrocosmos [Sho03]. Macrocosmos/microcosmos [Sho03]. Maddison [Tru75]. Madeline [Ano74r]. Madison [Ano78h]. Madrid [Ano90-32, Ano92-44, Aus16, Gar98, Ric98]. Maestro [FR83, Hug05]. Maffioli [Cap13]. Maghreb [Aïs96]. Maghrebian [Ano90a]. maghrébin [Ano93v, Mar90a]. Maghrib [Hog24]. Maghribian [Ano92-49]. magic [DV08, Hay87, Ste09, VV14, DV08, Ste09]. Magisteria [Tat11, BS09]. Magisterio [Ano91p]. Magistri [Syl07]. Magna’ [Tat11, BS09]. Magnitude [Gol00, Mal06b]. Magnitudes [GG96, GG97, vdW76a]. Magnus [Ano87j, Lo 85, Lo 83]. Mahameleth [Ses14, Fol15]. Mahoney [Ano74-28]. Mai [Dho88b]. Maierú [Sch13c, Mar14, Pan09]. Main [Ber04]. Mainardi [Phi79b]. Mainz [Ano94b]. Mainzer [Ric84]. Maison [DR03, Krö15]. Maistrov [Pla76]. Maitte [Høy07]. Majer [Row07]. majorizability [Nik09]. Majutsushi [Sid12]. maker [BG21b]. Makers [AH16, Sor13]. Making [Kel05]. Mal [Böl93, OA05]. Malagasy [Asc97]. Malayalam [Wag15]. Malcolm [Ano81u, vM07]. male [Kni01]. Malevich [BN22]. Mallikārjuna [Tee84]. Malte [Her14]. Malte-Ludolf [Her14]. man [Ano81-30, Ano88e, BGSS16, Bru77, Gan86]. managing [Par95]. Manara [Ano85g]. mandate [Des07]. Manegold [Row85d]. Manetín [Ano93a, Vog94]. manipulation [Fer07]. Mannelli [Bee09, Mar14, Pan09, Sch13c]. Männerkultur [Kni01]. Mannheim [Ric84, Sch96, Toe96]. Mannheim/Leipzig/Wien/Zürich [Sch96]. Mannigfaltigkeiten [Krö06]. Mannigfaltigkeitsbegriffs [Fel85, Che84]. Manno [Ser90]. Manohar [Knu11]. manoscrito [Hug05]. Manoscritti [Ken76b]. Mansfield [SS17]. manual [Ano74o, SS86]. Manuscript [BF89, Del02, Nau98, Ano80-35, Ano82q, Car18, Cul07, Del03a, Des10, GG75d, Haw96, Jen76, Mah78, Rey06, Rey09, Roz75, Vag82, vM85]. Manuscripta [Don05]. Manuscripts [Bel02, Dry14, MY99, Par02, Aro83, De 15, Del06, GGE82, Gro87, HNE21, IN83, Van84, dMC92, Duc16]. manuscrit [Cré87]. Manuscrits [dMC92, Aïs96]. manuskript [Jen76]. manuskripte [Ken76c]. Manuskripten [Bul22]. Many [Lan07, Nah06, Ang79b]. Maps [Asc95, Hog03]. Maracchia [Ano81f, Ken78b]. Marburg [Ano87m]. Marc [Høy24b]. Marcel [Cas77, Row83b]. Marceli [Moo82b]. March [Ano75p, Ano87-27, Ano89-33, Ano90t, Ano91s, Ano92g, Ano93w, Dru85, Jam97, MP06, Müü75, Nag91b, Pye74, Shi88, Ano24g, Kas90a]. Marcia [Par92b, Kat14]. Marco [Bru07]. Marczewski [Moo82b]. Margaret [Van80]. marginal [EG81, ST16]. marginalia [Ano77j]. Maria [Rom05, Fin11]. Mariano [Ric98]. Marie [Ano81j, Ano81n, Ano89m, Ano91q, Ano91s, Kno88a, Kno91a, Ped77, Tat86]. Marin [Mac76, Pan08]. Mario [For75, Ano89u, Ing11, Mar93a, Mar06]. Mark [GG10, How18, Str78]. Märker [Jur07]. Markus [Ano79j]. Markush [Lin74]. Markushevich [Ano81a]. Marlow [Ayo06]. Marquis [McL12]. Marquise [Böt12, Pei15, Böt12]. Mars [Ano88v, Car24]. Marseille [Ano83o, Ano85p, Che87, Hou89]. Marseilles [Dho83]. Marshall [Asc95, Mur07]. Martagon [Gra12]. Martens [Wil05]. Martha [Maz21, Row83b]. Martin [Ano74-29, Fol09a, Gra98a, Høy24a, Kat92, Pei22, Por05, Pys81, Syl03, Lum97]. Martina [Bor18]. Martinus [Pye76]. Martyrs [Ale10, Gra11]. Martzloff [Ano87h, Bré12, Dau93b]. Marx [Ken76c, Ken76b, Aro83, Ken77]. Marxism [Sen04]. Mary [Ano74z, Gri76, Por05, Sch05, Ste20]. Maryland [Ano86s]. März [Str81]. marzo [Ano87-27]. Masani [Ano92m]. Mascheroni [Fri21]. Masotto [Isr85, Ysh90]. Mass [Aga75, Ano74y, Ano77f, Bir83, Dho85, Gab76, Gra11, Haw77a, Haw77b, Haw81, Her75a, Hir74, Ken76a, Rus75, Tro74, Ung85, Yok75]. Mass./London [Haw81]. Massachusetts [AH16, Goo77, Lew99, Wil75b, Yon75]. Massimo [Cri24b, Fra06, PG24]. Masson [Dau93b]. Master [Ano93g, Hog10, Van83a]. Masters [And07, Ayo04]. mastery [Ano88f]. Matarrese [Ken76b]. Matematica [And75, Ano83c, Ano84y, DG90, Hug05, Pep90, Ano84g, Lei09, Ano81w, Arc99, DF10, Gan86, Ken78b, Mal06a, Pep92, Rem12, Roe89, Ano74-57, Ano85j, dL01, Ken82]. Matemáticas [Ges18, Ano92-44, Ano74-57, Ano87r, Aus16, Ano91-30]. Matematiche [Ene18, Fio17, Ano87-27, Ano91k, Maf10, Rab08, Rom97, Ser06, Cap13]. Matematicheskaya [Ano81o]. matematici [Ano82e, Arc17, Hug07, Ken76b]. Matematico [CB21, IN83, Mil93, DF10, Isr85, Par97, Rem12, Ysh90, Rod88]. matemáticos [Pan08]. Matematik [Atz04]. Matematik-fysiske [Atz04]. Matematika [Ken76d, Vol91, Kë96, Ano81o]. matematiki [Ano77l, Ken76d, Sti76b, Kin86a]. Materia [Ano80s]. material [Ano82-29, vM85]. Materialien [Ano88i, Tob90]. Materials [Fol83c, San03, Ano88k, AB84, BEJ^{+}75, Tee82]. Matériaux [Gra01]. Materie [Nab04]. Math [Ano92x, Goo77, Haw77a, Haw77b, Jay24b, Ano78j, Ano90s, Ano95k, Goe09]. Math-History [Ano90s]. Math. [Gui09a]. Matheematical [Kim75]. Mathematica [CM19a, Eva98, Ano80s, Gab76, Dau82a, Del03a, GG98, Gra98b, SY90, Gol76, Don05, Has10b, Ano77j, Pei17, Ano74-56, Ano92-33, BF00, Dau92, Dau94, Dau99, GG03, KP85, Kno17, Mar10c, PH96b, PH96a, PH97, Scr79b, Fel78]. Mathematical [Abe09, AH10, Ale14, All09, And07, Ane87, Ano74w, Ano74-46, Ano77d, Ano78t, Ano78y, Ano79-46, Ano80k, Ano81s, Ano81v, Ano85d, Ano86h, Ano86r, Ano87-36, Ano88-29, Ano90t, Ano92f, Ano93h, Ano93p, Ano95l, Asc84, Atz06, Ayo06, Bag97a, Bar74, BM96, Beb96, Bec01, Bid74, Bre19, Bur09, Cal08, Cas77, Coc98, Con06, Coo96, Coo99, Coo10, Cri04, D’A10, Dau12, Dho83, Don98, Dru85, Dru07, Eis82, Enr85b, Enr85a, Fef97, Fen08, Fer04b, Fol76, Gil79, Gil81, GG75a, GG92a, GG98, GG02, GG03, GG08, Gra94b, Gra94c, Gra97b, Gra98b, Gra06, Gra09b, Gra18, Gup79, Haw75b, Haw96, Haw03, Hog76, Hun96, Jah93, Jay24a, Kat97, Kat07, Kat20, Ken08, Kle01, Knu11, Lan07, Law05, Lew95, Lew98, Lor18, Mac78b, Mai21]. Mathematical [MY99, Mic23, Moo97, Moy17, Mro16, Nah06, Nak03, Neu08, Nic13, Par07, Por05, Pre24, Rad00, Rai92, Ram00a, Rec07, RWG95, Ric98, RW98, RW03, Rob96, RS85, Rog77, San03, Sca94, Sch89, Sch07, Sel91, Sen91, She96, Shi86, Shi93, Sid14, SS97b, Soi09, Ste01, Ste05, Ste80, Sti78, Str82a, Tat11, Thi06, Thi08, Tob81, Too80, Tur08, Van98, Van06, WM14, Wat09, Wil90, Wil08a, Zep12, Abe01, Aga75, AH20, AB09, And83, AKW04, Ano74g, Ano74h, Ano74s, Ano74t, Ano74l, Ano74-28, Ano76e, Ano78g, Ano81l, Ano82q, Ano85-31, Ano87k, Ano88u, Ano89h, Ano89r, Ano91g, Ano91p, Aro83, Aud11, Ayo04, BG21b, Bek88, Bel15, BK06, BR04, Bot77, Böt12, BT84, Che09, Che14a, Che17, Clu76, Cre16, Dad84, De 95]. mathematical [De 19, Des10, Des07, Dho90, Dij04, DS^{+}02, Ehr15, Ehr18, Enr80, FFW00, Fek14, Fis94, Fol15, Fri74a, Gou10, GG75b, GG75c, GG77, Gra75b, Hil83, Hog85a, Hol15, HMR17, Høy90a, Imh03a, Jam88, KB74, Lan75, Lüt09a, Lüt19, Mah84, Meo04, Mia08, Mik75, Mil93, Moo77, Nee79, Ney78, Ott89, Par91, Par92b, Par04, PGN18, Pes03, Pet12, Pro22, Pul05, Put75, Pys81, Ric80, Ros76, Roz75, Roz76, Sch86a, Sco16, Scr84, SS86, Sep05, Shu92, SS94, SS18, Swe79, Syn83, Tee84, Tur13, VV14, Van80, Van84, Wil79a, Wil08c, Yin11, Zer75, Zhu20, vH67, Eng81b, Eng81a, Jay24b, Pur86a, Pur88, Pur90, Pur91, Ser06, Ses13, Ano88g, Kil83, Lew90, Per10]. Mathematician [Aub02, Has10a, Par07, SS17, Wil07, Ano87f, Bie74b, Big19, BMW08, Cri06, DF10, DR16, Glu05, Gra75a, Gra78c, Hog77, Hog95, Inö06, Kno78a, MS05, Par05a, Sin85b, Zhm89, Zit07b, Jam13, Bir77, vM07]. Mathematicians [Ano90n, Ano91o, Ber75, Dau80, Fri74b, Kin86a, Shi86, Shi88, SS09, Taz01, VMT09, Ano74-74, CB21, EI12, Est79, Fle86b, Glu11, Nos12, SD08, SS10b, Sør14, Ton82, Zas83b, Jon76a, Ano84h, Bie74a, How18, Sch12, Sen10]. Mathématicien [Ano77n, DLS98, Gol13, Aïs96, Alf11, Auv13, Dje97, Gis14]. Mathématiciens [SD08, Mar10a]. Mathematics [Abe91, AH10, Ale10, Ale14, All04, AO05, Ano74-47, Ano74-50, Ano74-49, Ano74-60, Ano75g, Ano76-29, Ano77d, Ano78y, Ano79g, Ano79n, Ano79z, Ano79-32, Ano79-45, Ano79-47, Ano80p, Ano80r, Ano83f, Ano83k, Ano83n, Ano83y, Ano84g, Ano85k, Ano85r, Ano86c, Ano86i, Ano86l, Ano86n, Ano86o, Ano86r, Ano86a, Ano87j, Ano88f, Ano88a, Ano89o, Ano90m, Ano90a, Ano91f, Ano91o, Ano91t, Ano91-33, Ano92h, Ano92w, Ano92v, Ano92-30, Ano92-47, Ano92-48, Ano92-49, Ano93i, Ano93m, Ano93w, Ano95j, Ano95k, Ano95-27, Arc96, Arc09, AH99, Ber97, Bin83, Bin91, Bos02, BF00, BMW08, Bur93, Cal83, Cal08, Che12, Coc98, Cri82, Cri99, Cri24b, D’A10, Das81, Dau86, Dau93a, Déc10, Dem74, DV80, DK86, Des09, Dra74, Dum18, DR19, Eng81b]. Mathematics [Eng81a, Eva98, Eva13, Fef97, Fen97, Fen08, Fen09, Fol15, FR98, Fra17b, Gar86a, Gar92, GBGL08, GG05a, GG88, GG02, GG03, Gra84b, Gra87, Gra94b, Gra96, Gra99b, Gri76, Gue88, GM04, GC21, Gup78, Gup85, Gup86, Gup91, Haw03, Hoc96, Hog85b, Hog01b, How18, Høy09b, Jah12, Kas90a, Kat07, Kat09, Kat11, Ken76d, Kla84, Kni01, Kno94, Knu11, Kot79, Lê24, Lev97, Lon21, Luc18, Mac78b, Mac78a, May75c, May77c, McK82, Meh80, Mel09, Mic88, Mon11, Moo97, Mro16, Neu08, Nic13, Oli08, Özd00, PH96b, Par03, Pul98, Pur86a, Pur86b, Pur88, Pur90, Pur91, Pyc84a, Rad00, Ram00b, Rec07, Ric96, Ric98, Rid94, RS09, Rog74, Rog77, Rom10, Row85c, Row90b, San03, San07, Sar88, Sar97, Sar08, Sch89, Sch07, Scr84, Scr85a]. Mathematics [Seg86, Sen04, Ser18, SG09, Shi93, Sid12, SS07, Soi09, Swe79, Swe92, oSTT79, Tob81, Ung22, VK05, Vuc99, Vuc00, Vuc02, WDB96, Wat09, Wil97, Zas79, ZD88, Zho21, dS99a, Alb94, Ano74e, Ano74j, Ano74-48, Ano74-55, Ano74-64, Ano74-71, Ano74-70, Ano75f, Ano76f, Ano76v, Ano77s, Ano77-30, Ano78-27, Ano79k, Ano79-30, Ano79-33, Ano79-34, Ano79-37, Ano79-43, Ano79-48, Ano80t, Ano80u, Ano80-31, Ano80-33, Ano80-38, Ano81s, Ano81e, Ano81q, Ano81u, Ano81-31, Ano81-32, Ano81a, Ano81-36, Ano81-45, Ano82j, Ano82l, Ano82k, Ano83i, Ano83z, Ano84h, Ano84j, Ano84l, Ano85i, Ano85n, Ano85-27, Ano85-28, Ano86f, Ano87-29, Ano87-27, Ano87-30, Ano88g, Ano88-30, Ano89d, Ano89e, Ano89l, Ano89-32, Ano90n, Ano90-33, Ano91j, Ano91k, Ano91-31, Ano91-32, Ano92g, Ano92i]. mathematics [Ano92u, Ano92-43, Ano92-51, Ano92-52, Ano93n, Ano93o, Ano93a, Ano94g, AB84, Arn88a, Art91, BJ89a, BS05, Ber84a, Ber85, Bin90, Bis75, Boo86, BM77, Bro75, Buc80, BT81, Bun74, Bun76a, Bur87, Cam76, Car07, CJ84, Che87, Con85, Coo88, Cra07, Cra09, Cro75c, Cro75b, Cul07, D’A79, Dau82a, Dau82b, Dau85b, Dav76, Daw16a, DF10, Del03b, Dho83, Dia75, Dun89, EI12, Eis80, Enr83, Enr85b, Enr85a, Eve76, Fen03, Fen89, Fer20, FG06, Fin11, Fle75, Fol76, Fol83c, Fri81a, Gal10, Gan86, Gar86b, Ger94, Gil89, Gis18, Giu86, Glu05, Goe08, Goe09, Gra75a, GG76c, GG87a, GG04, GGC^{+}05, Gra76c, GN92, GN05, Gui24, Gup75a, Gup76b, Gup80, Gup93, Haw80, Haw81, Haw84, H^{+}08, Hee14, HEH^{+}04, Hog74, Hog81a, Hog90, Hog91]. mathematics [HP20, Hug75, Jan11, Joh11, Jon75b, Jon76b, Kah75, Kar22, Kas90b, Kas03, Ken84, Ken08, KAHR08, Kid09, Kno06, Kop75, Kul76, Lau99, Law05, Lew75, Lew76, Lew78d, Lew04b, Lo 83, Maf10, Mal79, Mar93a, Mar90c, Mas08, Mat93, Meh76, Moo82a, Nis89, Nur92, PS83a, Par92a, Pep90, Pep92, Per93, Per80, Pet78c, Pet85, òP09, Pow76, Pre24, Pri80, Rab10, Rab24, Rai92, Rid88, RH11, Roe89, Row85a, Row89, Row94, Sar93, Sch76b, Scr75b, Scr82a, Scr82b, Scr85c, Seg76, Sel90, Sel91, SGJ05, SG07, Shi86, SS93, SS10c, SS12b, SS22b, Smi82, Ste07a, Sti76b, Sti78, Str82a, Swe74, SS84, Thi75, Tim04, Tob06, Tra21, Tur20, Ung85, Van09, Van79, Van14, Vog94, Wal74, Wal75, War10, Wea77, Wee91, Wil74]. mathematics [Wil79b, Yad80, Yon75, Yos75, Yos76, You83, You86, Zit07b, dS90, Ano74-32, Ano75q, Ano76g, Ano78k, Ano80c, Ano83w, Ano84q, Ano84x, Ano87-32, Ano91t, Ano92-34, Ano93-34, Ano94i, Ano08e, ABC^{+}88, Cha81, Dau94, GG87b, GS77, Gra78b, Gra97b, Gup81, Jon77a, Kë96, Scr79b, Scr85b, lLW13, VHK16, Ano74f, Ano74-33, Ano93n, Cha87, All09, Ano93g, Ano95g, Bra12, Coo96, Fra10, Gra10a, Gra11, Hun12, Imh21, Jon77b, Moo03, Mro16, SG04, Sia23, Sid06a, Wol03, Ano74o, Ano74u, Ano77q, Ano86e, Ano88h, Ano91d, Bow84, Cro77, Ewi92, Gra77, Hug76, Kat92, Moo75a, Row93, Sal78, Sch89, Wil75b, Ano74y, Asp19, Ber74, Bra07, Dav90, Gra94c, Gra97a, Jon76a, Ken08, Tur08, Wol91, Yok75, KKS18]. Mathematics-à [Bin91]. Mathematics-Stimulating [Ano86l, Bin90]. Mathematics-Theme [Kas90a]. Mathematics/Société [Ano74-32, Ano78k, Jon77a, Cha81, Ano74-33, Cha87]. Mathematik [Ano84k, Ano88n, Ano89k, Ano90l, Ano92b, Ano92-46, Ano94b, Ano95m, Bie83, Bos77, BMW08, Bow84, Che14b, Déc10, GG07, H^{+}08, Jay83, Kni01, Krö14, Lem74, Mon75, Pur86a, Pur88, Row08, Row19, Sch90a, Sch11, Sch17, Sch90b, Sch09, Scr03, Seg90, SS89, SS08, Str81, Tob81, Tob06, Toe96, Ume76, Cal75, PS83a, PS83b, PS83c, Pur90, Pur91, Sch82b, Ano91-32, Ano92-52, Ano93x, Bin89, Nis89, Pei15, Hei09, Ano81g, Sch16]. Mathematik- [Bie83, H^{+}08, Sch11]. mathematik-Die [Sch82b]. Mathematiker [Ano90l, Ano91c, Ano92h, BMW08, Has10b, Has10a, Hes90, Kno78a, Sch09, SS89, Ano87m, Mon83, Ano92-46, Ano93k, Ano85e]. Mathematiker-Vereinigung [Ano91c, Ano92h, SS89, Ano92-46, Ano93k]. Mathematikerkongreß [Ano90-36]. Mathematikers [Aud12, Bie74b, Bos03, Krö14, Sch96]. Mathematiklehrerberufs [Tob90]. Mathématique [Ano90i, AD87, Ano87h, Ano92k, Ano92-35, BKM^{+}82, Bil85, Bré12, Bre91, Coo99, DP88, Dho90, Dug76, Gis18, Gra76a, Hou89]. mathématique/Mathematical [Coo99]. Mathematiques [Ano90-34, Ano83o, Ano90-34, Ano91q, Ano92a, Ano92-44, Bor89, Che87, Déc14, Dho91, Lor16, Rom10, Yus10, Ano79l, Cha81, Ano74-33, Ano75f, Ano78k, Ano80d, Ano87-28, Ano89n, Ano89t, Ano90d, Ano90g, Ano91n, Ano92q, Ano92-27, Ano93-33, Ano93v, BKM^{+}82, BH14, Boi89, Cav95, Cha87, Dau93b, Dur21, Ebe96, Enr76, FN11, For10, Fra04a, Gau13, Gis99b, Giu99, Gre85, Hou89, Høy24b, KS13b, Lam19, Mar90a, Pei17, Rei15, Ses13, Taz15, Wil08b, Ano74-32, Gra01, Jon77a, Lai10, Ano79m, Ano77k, Che94, Krö15, Maz24, Cav95, Fol12]. Mathématisation [Ano88v, Bun76b]. Mathématisations. [GW94a, GW94b]. Mathematisch [Ano81e, Scr03]. Mathematisch-Naturwissenschaftliche [Scr03]. Mathematische [Bla05, Ken76c, Too80, Böt12, Pei15, Ano90t, Ano91u, Ano92l, Cri17]. Mathematischen [Lüt09a, Pul05, Tob81, Bar82a, Bar82b, BB04, Eng81b, Eng81a, Fol87, Her14, HB11, Imh03a, Sch02, vdW80]. Mathematischer [Bar78b, Gar05, Ano88i]. Mathematisches [Ano90-33, Ano91-33, Ano92w, Ano92b, Mon83, Pur86a, Pur88, Pur90, Pur91, Art91, Hal78]. Mathematon [Høy95]. Mathemetik [Eng81b, Eng81a]. Mathesis [Ano87r, Vil97, Lê21]. Mathieu [Spi00a, Spi00b]. matrices [Haw75a]. Matter [Gol00, VHK16]. Matthew [Smi82]. Matthias [Kno93a, Thi05]. Mattmüller [Pei22, Syl03]. mature [Bra89]. Matvievskaya [Kin86a]. Matz [Fol09a, Fol09a]. Mauchly [Asp83]. Mauldin [Kla84]. Maupertuis [Ter89]. Maurer [Has04, Sch09]. Maurice [Ano74z, Mue96]. Maurin [Ped77, Gra99b]. Maurolyci [Net24, Net24]. Max [Men86]. Maxim [Eve76]. Maximum [Ano90l]. Maxine [GG90a]. Maxwell [Fle86a]. May [Abe91, Ano74g, Ano74o, Ano80r, Ano83h, Ano84q, Ano86m, Ano88a, Ano90f, Ano90j, Ano91h, Ano91-31, Ano92a, Ano92-29, Ano92-34, Ano92b, Ano93s, Ano93a, Ano08e, Ano10k, Bid74, Bow87, Bra07, Cas77, Fol94, Gia90, Gra02, Hog15, Koe94, Par05b, Pur88, Pur91, SY90, Vol93a, Buc23, Ano77b, Ano77c, Ano21k, Ano23k, BG21a, DR78, Dra78, Enr84, Fra13, Fra17a, JET84, Lew04b, VK05]. Mayer [Ano74-27]. Mayr [Rus75]. Mazliak [Sch12]. Mazzinghi [Fra88]. Mazzini [BD11]. Mazzotti [Cri24b]. McArthur [Rid94]. McCartney [How18]. McCleary [Ewi92]. McCormmach [Ano88f]. McGilvray [Ano92n]. McGraw [Boy74]. McGraw-Hill [Boy74]. McKinnon [Nic13]. McLean [Tan74]. me [Fer95]. mean [Høy90a, Sho85]. Meaning [Con06, Pes03, Tho11, Ano85h]. means [Blå17a, Coo11, Wat93]. Measure [Lev97, Swe00]. measurement [Gol84, Sch91b, Vog83]. measurements [LYTS86]. measures [Pie86, Sen87]. Measuring [Fie22]. Mécanique [Law98, Ano90q, Fra04a, DD87]. Mecca [Hog03, Ken84]. meccanica [Pan18]. Meccanicismo. [Pan18]. mechanica [Rid83a, Rid83b]. mechanicae [Bar77, Rid83a, Rid83b, Sch00b]. mechanical [BR16, Ros17, KO09, Ros17]. Mechanics [AH16, Ano80-37, Aub14, Dem74, Fra06, Law98, Nau98, Pul98, BTB^{+}03, GG87b, GG88, GG90e, GG14, Sel06, Sho03, Sor13, Tat75, DD87, Nor07, Lüt05]. Mechanik [Ano89k, Ano90l, Gra90, Hes90]. mechanischen [Ora82]. Mechanisches [KO09, Lef10]. Mechanistic [Lüt05, Nor07]. mechanization [BS05, Kna88]. Mechthild [Sch17]. Méconnu [DLS98]. Medal [Ano10k, Par05b, Tro76, RH11, Nic13]. Meddelser [Atz04]. Media [Wat09, Che10]. Mediaeval [Maz77, Moy19]. Mediasenter [BJ89a]. medical [Ine94]. medicinalium [Ano77g]. Medicine [Ano87n, Bra09b, Ano84m, vD09]. Medieval [Ano84d, Ano89e, Cri24a, Eva98, Gup91, Hod75, Ken75, Mor96, Moy17, Oak12, Özd00, Ste01, Yan83, Ano88k, Ano92i, Ano94h, Cor16, Dav75, Hai86, Hog24, Joh11, Kin74a, Kin74b, Kin79, Kin86b, Kno91b, Mat93, Roz75, Syl83, Yad80, Yus95, Amb08, Ber97, Lan96, Sin85a, Lo 84]. medievale [Amb08, Oak12]. médiévales [Høy24b]. médiévaux [Aïs96]. Medioevale [Ano83c, Hug05, Ano81w]. medioevo [Ano85f]. Mediterranean [Che87]. Méditerranée [Ano89n]. Medium [Wei98]. Medizin [Scr03]. Mediziner [Row85b]. medizinischen [Ine94]. Medvedev [Ano81h, Smi79b, YDD83, Zai95]. Meersseman [Gra08]. Meerut [Ano85n]. Meeting [Ano74-47, Ano74-51, Ano75k, Ano76-29, Ano77t, Ano85d, Ano86o, Ano87-32, Ano92h, Ano92f, Ano92v, Ano92b, Ano93k, Ber85, Dru85, Pur88, Pur90, Pur91, Scr82a, Tob81, Abe91, Ano76f, Ano76h, Ano76-45, Ano85i, Ano87n, Ano92-46, Ano92-52, Art91, Dun89, Gra76c, Jon77a, Jon81, Lew76, Scr85a]. Meeting-1984 [Ber85]. Meetings [Ano74-50, Ano74-61, Ano75r, Ano75s, Ano76-30, Ano77-33, Ano78z, Ano83w, Ano94m, Ano94n, Ano95r, Ano78t, Ano84x]. Megara [Clu76]. Mehmet [Inö06]. Mehrtens [Bir80, Mon83, SS22b]. Mei [Ano87h, Dun89]. Meidong [XX10]. meilleur [Ano90-35]. Meissel [Pee95]. mélanges [FG12, Lüt15]. Melbourne [Coh82, Hoc96]. Melencolia [Sch99b, Wei04]. Melo [Ges18]. Melon [Van05]. Melon-Shaped [Van05]. Melvyn [Ano82d]. members [Ano20i]. Memoir [Del02, Del03a, RL16]. mémoire [Gra01, Kno88a, Ano80i]. Memoires [Che94, Ano87h, Ple78, Fra04a, Wil08b]. Memoirs [Syl81, PP78, Sar08]. memorabilia [Hal87, Wef89]. Memorial [Ano75e, Fel83, Dra78, Syt91, Gia90]. Memoriam [BB97, DS97, Fow84, Jam97, Ano80m, Ano80-36, Ber84b, GG79, Kli76, Scr79a, ZvDF22]. memorie [KO09, KO09, Lef10]. Memorieren [KO09, Lef10]. memorization [KO09]. Memory [Fra06, BTB^{+}03]. Men [Des09, Cra07, Hod77, Asp19, Gra77, Moo75a]. Mendel [Pie83]. Mendeleev [She96]. Menelaos [Yus95]. Menelaus [Yus95]. Mengenlehre [Ano89k, Plo06b]. Menger [Maz24]. Mengoli [Est06, Mas97, NS94]. Menso [Høy23, Høy24a, Mat75, Rom99, Van06]. mensuration [Lo 84]. mental [MS05, vM07]. mente [Hug07]. Mentoring [Tur20]. mercantile [Ano85f]. Mercator [War08]. Merchants [Ano95l]. Mercure [Mal13]. meridian [Gow95]. merito [Rom05]. Merrill [Hug76]. Merrington [GG90a]. Mersenne [Mac76, EG81, Mac77]. meru [SK24]. Merzbach [Ano77f, Neu86]. Mesopotamia [Che12, Mel02, Mel04, VF22]. Message [Dau93a, Kno94, Dau86, Dau92, Dau94, Scr79b, Ano77n]. messenger [Eck13]. Messengers [Ric98]. mesure [Ano95i]. Mesures [Pie86]. Meta [Fek14, SS22a, DN12, Fol14]. Meta-mathematical [Fek14]. meta-review [SS22a]. metafisica [Bee09]. metaphysics [Man89]. meteorology [Fri74b]. Method [Hog02a, PS97, Plo96, Abd08, Ano74-30, Ano77r, Car22, Chi22, Chr07, CS81, FLT16, Ger74, Gre88, Kre03, Mar10c, Ray12, Tai14, Tat87, Wil80]. Methode [Dho11, Fre77, Dho91, Mal97]. Methoden [Ano88i]. Methoden-Modelle [Ano88i]. Méthodes [Che94, Lam19]. Methodist [Cas77]. methodological [Dad84, Len79]. methodology [Ano74-30, Ano87-35, Eis82, Dru86]. Methods [Ano79-46, Fri81a, LYK89, Lum97, Van98, Ano91g, EF20, Ger85, Glu76, Yin11, Zho21, Sen06]. Methodus [Ano84-29]. Methuen [Tru75]. métiers [Dho88b]. Métivier [Ano87i]. metric [vBS20]. Metrika [Tai14]. métrologie [Bab77]. Metropolis [Tro83, Ric96]. Metter [Pet82]. Meudon [Ano90q]. México [Aus16, Ano85-27]. Mezzo [Mal06a]. Miami [Ano92x]. Michael [Ano74-28, Ano88h, Ano89c, Bra07, Cro77, Enr80, Haw77a, Haw77b, Jur07, McC90, Row07, Sai05, Sau14]. Michel [Ano87i, Ano95i, Dau12, Mac76, Mar10a, Maz24, Vil96]. Michel-Paul [Mac76]. Michèle [KS13a, Krö14]. Michèlexunbr [Ale14]. Michelle [Fra04a, Wil08b]. Michelstadt [Fol09a]. Michigan [Hod75]. microcosmos [Sho03]. microform [Ano88w]. Micronesia [BB21]. mid [Cri04, Del03b, Ver15]. mid-Nineteenth [Del03b]. mid-nineteenth-century [Ver15]. mid-Victorian [Cri04]. Middle [Hog24, Ken75, Kin86a, Wis76, Ehr11, Syl81, Fol83c, Imh03a, Kna88, Ses13]. Middlesex [Ano74-56, Ano74-59]. Midwest [Ano92x]. Miedzyzdroje [Ano88a]. Mighty [Ano88f]. Mikhail [DLS98]. Mikhailov [Thi08]. Miklós [Rec07]. Mikosch [Sch11]. Milan [Ano81r, For75]. Milano [Ano81f, Ano87c, Oak12]. Military [ABC^{+}88, Arn88a, Kas03, San03]. Millennium [BF00]. Miller [Gra77, Wil77b, GG94]. Mills [Ano79-44, Hal84, Ken79a, Ken79b]. Milton [BGG06a]. Mind [Abe06, Ano91t, Edw04, Ken76a]. Minerals [Kin86b]. Ming [iÔ75, iÔ75]. mini [Lan75]. mini-study [Lan75]. miniature [Cam77]. Miniconference [Ano86c]. Ministry [Sai74]. Minkowski [Ano89k, Ano74-62, Str85]. Minkowski-Welt [Ano89k]. Minkowskis [Str85]. Minneapolis [Ano91d]. Minnes [GG10]. Minnesota [Ano91d]. Minoan [Str82b]. Mint [Big19]. minutes [Ano74w]. Mir [Tot79]. miracle [Høy83]. Mirek [Plo06b]. Mirror [WDB96, DR16, Ebb04]. Mirrors [Ano77i]. Miscellanea [Don08]. Miscellany [Ken08, Hog85a, Luc21, Ken08]. miserie [Pan18]. Mises [EI12, SS93, SSZ07, SS10c, Sis02]. Misra [Gur24]. missed [SS10b]. Missing [Sim00a, Sim00b, FK90, Jon77b, Lor15]. Mitarbeit [Toe96]. mithartum [Fow92b]. Mito [Ano90c]. Mitra [Dho91]. mitraille [Dur21]. Mittag [Vol02, Dau80, Tur13]. Mittag-Leffler [Vol02, Dau80, Tur13]. mittelägyptischen [BB04, Imh03a]. Mittelalter [Ano85e, Ano91u, Kna88, Jay83]. Mittelalter-Der [Ano91u]. mittelalterliche [Yus95]. Mitteln [Høy90b]. mnozhestv [Ano81h]. möchte [KS12, Tob10]. mode [Bin91]. model [Car22, Car24, Ene18]. modèles [Cor96]. modèles-Essai [Cor96]. Modeling [Ano91g, Fen97]. Modelle [Ano88i]. Models [Asc95, Ano74l, BG21b, Vol19]. Modena [Ano91k]. Modern [Ale10, Ang88, Ano74w, Ano85k, Ano91-33, Ano92w, Ano92-47, Ano94i, Ano95l, Bec80, Bel12, BMY07b, Chi22, Cri24a, Dod13, Epp95, Eva98, Gra11, Hoc96, Kle01, Neu08, Oli08, Row90b, Thi08, Ang79b, Ano74u, Ano84i, Ano90-33, Ano91d, Ano92-52, Bar74, Bir74, BMY07a, Cre16, Dub77, Ewi92, FG06, Fis11, Fre75, FC22, GP07, Lew75, Mas08, Ped75b, Row89, Tho76, Yos75, Yos76, vM84, vM07]. moderna [Ano88r, Taz09, Net02]. moderne [Sch17, vdW75]. modernen [Ano92-52, Ano94b, Høy90b, Sch02]. modernes [Ano91q, De 20, Ano90d]. Modernist [Fen09]. modernity [SS22b]. Modes [Goe08, Goe09]. Modular [Bul09b]. mögliche [Jay83]. Mohammad [DS99c]. moitié [Pan88]. Moktefi [Kat20]. Moldavian [Dav76]. Molk [Gis99b]. Moment [Kje93, MB20]. Momenti [Ano81f]. momentis [Net24]. moments [Enr85b, Enr85a]. Monastyrsky [McC90]. Monatsbuch [Bul22]. mondaines [BH14]. Monde [Ano92-27, Maz24, Ano95n, Dho90, Fol12, Rom10]. mondes [Ano90-35]. mondo [Pan18]. Monge [RL16, Gra09b]. Mongolia [Shi94, Swe92]. Mongolian [DZ97, Mar97]. Mongré [HB^{+}13, HB^{+}13]. Monna [Ano77o, Aup76, GG77]. Monnier [Ken78b]. Monograph [Her75a, Tat75]. Monographien [Toe96]. Monographs [Ano78-27]. Monschau [Str81]. Monte [Ano92s]. Montel [Ale14, Aud11]. Montréal [Ano77k, GG76a, Pye74, Ano75q]. Montreuil [Fol12]. Montucla [Ano10f, Ano21a, Aus13, Sør17]. monument [AB09, Ano76j]. Moon [Sid24]. Moonlight [Knu11, òP09]. Moore [Ano86e, Dru91, GG86, Hym78, Par07, Gre77, Lew04a, Par05a, Ano74r]. Mordechai [Bul16a]. Morgan [GG90a, Jon85, Moo75a, BLW77, DR16, HMR17, Pan03]. Morgen [KS12, Tob10]. Moritz [Plo06b, Sch02, Sch13a]. Mormann [Ano93i]. Moscow [Ano77l, Ano81h, Ano81o, Ano87-35, Coo96, Git74, GG80, Ken76d, Kin86a, Kot79, Lin74, PP78, Phi79a, Por75, Smi79b, Tot79, Coo10, Phi78]. Moskva [Dyc75a]. most [Tra10]. Mostowski [Moo82b]. Mota [Ges18]. Motilal [Yan83]. Motion [Van98, Buc77, Cha77, Kah75, Mae89, Ano77p, Bo24]. motions [Vek75]. Moufang [Cer04]. Moulton [Glu11]. Mouton [Ano77s, Pys81]. mouvement [Ano81m, Ano86g, Lai10]. movimento [Bee09]. Moyen [Syl07, Rei15, Ses13]. Moyon [Høy24b]. Mr [Cra07, Des09]. Mr. [AD78, Llo79]. MR2609069 [Has10b]. MS [DS99c, CH90, Mat74, Eli24, War15]. Mss [Mat74]. much [Ped75a]. Müchen [Jay83]. Mueller [Høy95, Ung85]. Muhammad [Tee78]. Mulino [Ken82]. multicultural [Par92b]. Multiplication [Abe76, FS22, Kin74a, Kin79]. München [Hei09, Row08]. Munich [Ano90-37, Bla05, Dil81, Don05, Gar05, Scr87]. municipal [Vag82]. Münster [Grö86]. Muqarnas [DS99c]. Murrākusī [Dje98]. Murray [Moo75a]. Murthy [Ano88k]. Muse [Sza74]. musée [Cav85]. Musei [Ano87d]. Museum [Ano87d, Ano90j, Sau14, Sid06b, Wil77b]. music [Ano92-34, Kim75]. Musical [War08]. Musikgeschichte [Høy90b]. Musings [Ayo04, And07]. Muslim [Sal78]. Mussia [Ano81r]. musulmanskogo [Kin86a]. Mu’taman [Hog95, Dje97]. Mutual [Kan88, Zho21]. Mutual-subtraction [Kan88]. my [Str94, vdW75, vdW76b]. Mycenaean [Str82b]. mysteries [DF18]. Mysterium [Gin83]. mystery [Dar15, Erl94]. Mysticism [Dau12]. Mythes [Ano92-35].

N [Aga75, Ano74m, Ano81p, Ano81o, Ano82c, Ano82f, Ano88m, Ano89d, Ano89k, Bar74, D’A10, Dav75, For98, GG80, GG02, GG03, Gra94b, Gra94c, Gra97a, Har77a, Por75, Sai05, Stu80, Taz09, Thi08, Tro83, Tru75, You83, Ano77j, Gar98]. N. [For98]. N.J [Ano77s]. N.Y [Ano74w, Ano81u]. Nāser [Bag97a]. Nāser-e [Bag97a]. Nabonnand [Krö15, Vol02]. nach [Bie74b, Bus76, Rom99, Vog83]. nachala [Ano77l, Ken76d, Sti76b]. nachgelassenen [Bul22]. Nachlass [Fol78, Neu93, Sch98, Ano87s, Edw78, HB^{+}13, Neu88, Sch86b]. Nachlasses [Gab79]. Nachlese [Gab79]. nacimiento [Gar98]. Nacional [Ges18]. Nadia [Oak12]. Nadir [Inö06]. Nadis [Dum18]. Nagel [Grö86]. Nahen [Bus76]. Nahin [Lan07]. Naini [Ano85e]. naissance [Pan88]. Nakayama [Hir74]. Nam [Thi75, Yin11]. Name [Ano95k, Ric84, Ano81l, Høy95, Moo03, Scr81]. named [KP85]. Namen [Scr81]. names [Ano93i, Sca94]. Naming [Dau12]. Namnum [Ano81p]. Nan [Lo 84]. Nancy [Ano84i, Asp83, Cog15, Gra12, Lüt15, Maz77, BPL17]. Nantwich [Hal84]. Naples [Ano92s]. Napoleon [Ano89j]. Napoleonic [Pep90]. Napoleons [Scr81, Scr81]. Napoletano [Rom97]. Napoli [Taz09]. Napolitani [Fol14, Net24]. Nasir [SI19]. Nastasi [Arc09, Arc17]. Nat [Ano79-42]. Natalia [GG05a, Sch13b]. Nathan [Hir74, Moy17, Par91, Par91]. Nathanson [Ano82d]. Nation [Ste01]. National [Ano74m, Ano76j, Ano87n, Ano88k, Ano88m, Ano92-50, GG76a, Gup88, Hei81, Pet82, RWG95, RW98, Tee84, ZD88, Ano76-34, Seg86, Ano90u]. nationale [Ano92-28]. nationales [Ano87i]. Native [Ano88h]. natura [Rab08, Ser06]. Natural [Dyc75b, Kni01, She96, Cro77, Lüt09a, Pul05]. naturalis [Gab76, Pei17]. Naturalium [Rid83a, Rid83b, Ano77g]. Nature [Ano95l, Ano87k, Ano88f, Ser06, Ano86g]. naturelle [Pei17]. naturels [Maz16]. Naturphilosophen [Bos03]. Naturphilosophie [Lüt09a, Pul05]. naturphilosophische [Böt12, Pei15]. Naturwissenschaft [Fol87]. Naturwissenschaften [GG07, Jur07, Scr03, Kni01, Ano89j]. Naturwissenschaftler [Ano90l, Row85b]. Naturwissenschaftliche [Scr03]. naturwissenschaftlicher [Bar78b, Gar05]. nauk [Ken75]. Nauka [Ano77l, Ano81t, Ano81h, Ano81o, Dyc75a, Dyc75b, Git74, GG80, Ken76d, Kin86a, PP78, Phi79a, Por75, Sel74, Smi79b]. nautical [Gow95]. navalem [Bar77, Rid83b]. Navarra [Pan08]. Nayrīzīs [Rom04]. Nazi [Sen10, SS09]. Nazionale [Ano90f, Pep90, Pep92, Ano82-28]. Neale [Tan74]. Near [Hog24, Ken75, Ano92v, Hai86, Spi00b]. nearby [Bus76]. Necipoglu [DS99c]. Needham [Jam97]. neeuclidiene [Ano74p]. Negatives [Rab24]. neglected [Wat86]. Negro [Zas83a]. nei [Ben95]. Neißstraße [Ano90j]. Neither [Rob01]. nel [Mai09, Mar14]. nell [Ano93j]. nella [Coz12, Gan86, Sai94, Ser90, Tra10]. nelle [Amb08, Oak12]. nell’età [Rom05]. nell’Italia [Arc99, Arc17]. Nemorarius [Cor16]. nemore [Gra83, Hug89]. Neo [Gre83, Gal12]. Neo-Kantian [Gre83]. neoclassico [Ben88]. Neohumanism [Row85c]. Nervo [GG05a]. Netherlands [Str82a, Bec99]. Networks [Gis96, Par92a]. Netz [Sch13b, Sia23]. Neu [Bul22, Str78]. Neue [Sau14]. Neuenschwander [Ull03]. neuere [Ano78j]. neueren [PS83c]. Neues [Row06a, Wue05]. Neugebauer [Swe79, Ano88w, HP20]. Neuhofen [Ano92-48, Bin90, Bin91]. Neumaier [Høy90b]. Neumann [Asp83, Bir83, Has10b, Has10a, Lüt09a, Rec07, Ano75x, Hal87, Has10a, Pul05, Réd05, Ven02, Von88, Zun02, Bir97]. Neustadt [Ano91-31]. Neustadt/Weinstraße [Ano91-31]. Neuzeit [Ano85e, Ano93x, Høy09a]. Nevanlinnas [Gra10b]. Never [Ait86]. Newell [Zas83b]. Newly [Bag97b, CWW17, Moy19, Roz75]. News [Ano74-63, Ano75t, Ano76-31, Ano76-32, Ano76-33, Ano77-34, Ano77-35, Ano77-36, Ano77-37, Ano81-38, Ano81-37, Ano82r, Ano87t, Ano89v, Ano83b]. Newsletter [Ano74-64, Ano74-34]. Newton [Ano87n, Ano90q, eFD88, Duc16, Gab76, GG87b, Lew98, Lüt09a, Pul05, Sch05, Ait86, AA88, Ano78a, Ano86s, Blå23, Bra89, Chi22, Dry14, Erl92, Erl94, Erl97, Erl03, Fel78, Gal10, GG88, Gui96, Gui03, Jor81, Lai75, Lok80, Pan05, Pes01, Pou01, Pou07, Pul05, Pyc97, Sel06, Ser06, War08, Wei92, Wei98, Wil94, Coh82, Rab08, Bru07, Bul16a]. Newtoni [Ano77j]. Newtonian [Sag91]. Newtonianism [Pet91]. newtonienne [Ano90q]. Newtons [Fel78, Nau03]. Neyman [Ken76a]. Niccolò [Sag91]. Nice [Ano91n, Eva94, Gra01]. Nicholas [Tat92]. Nicholaus [CG92]. nicht [Ano89k, Gra78d]. nicht-euklidische [Gra78d]. nicht-euklidischen [Ano89k]. Nichtstandard [Spa89]. Nichtstandard-Analysis [Spa89]. Nicolaevich [Phi78]. Nicolai [Phi78]. Nicolas [Cra12, Dra84]. Nicolaus [Fol09a, Grö86]. Nicolaus-Matz-Bibliothek [Fol09a]. Nicolo [Scr03]. Niedersächsichen [Bar78b]. Niels [Ano84k, Bul22]. Nieuw [Ano81e]. Nigel [Sch13b]. Nigeria [Shi88]. Nihonjin [Ano78-29]. Nijhoff [Pye76]. Nijmegen [Lo 85, vM84]. Nikoladze [Por75]. Nikolaevich [DLS98, Por75]. Nikolai [KS08, Phi79a]. Nikolajevich [KS08]. Nine [Shu92, Vac20, Xu05]. nine-point [Vac20]. Nineteenth [Ano82c, Buc77, Row90a, Ver15, Del03b, Ehr11, WDB96]. Nineteenth-Century [WDB96]. Ninth [Ano76-34]. Nippur [dV22, Yus10]. Nityananda [Gur24]. NJ [Ano84f, Hun07, Plo06a, Sch91a, Sch91c, Wil05]. Nl [BT81]. No [Ano81k, Ano84g, Ano84h, Ano87-36, Ano92k, Dho11, Enr85b, Enr85a, Fer11, For10, Fre92, Gra77, GG90c, GG90b, GG10, Gra10b, Høy95, Hug05, Kat11, Knu11, Kot79, Mah78, Mal97, Par97, Pys81, Scr03, Yus10, Ano78-29, DV08, iÔ75, Sid12, Ste09]. Noble [Bul24]. Noch [Ber90]. Noel [Sch13b, vM07]. Noether [Cho86, Cog12, Gra08, Jen86, Row83a, Row83b, Sch17, Ano74k, Jen86, Men86, ST16]. Noether-Schule [Sch17]. nombre [DR03, Fer98a]. nombres [BKM^{+}82, DR03, Fla03, Har06, Pic89, Ras89, Spi00a, Spi00b]. Nominalism [Sep05]. Nomography [Glu12]. Non [AR24, Cer04, Gra79, Kat92, Ano74p, Asc84, Gra91b, Gra99a, Sch79, Zhu20, Haw80, Roz90]. Non-Desarguian [Cer04]. Non-Euclidean [Gra79, Ano74p, Gra91b, Sch79]. non-euclidiennes [Roz90]. Non-European [Kat92]. non-mathematical [Zhu20]. Non-standard [AR24]. non-Western [Asc84]. Nondecimal [Knu83b, Knu83a]. Nonlinear [Kje00]. nonlogical [Gar85]. nonstandard [Arn88b, Arn88b]. Nonstandard-Analysis [Arn88b]. Noordhoff [Pin76]. nor [Rob01]. Nora [Bla05]. Norbert [Bir83, Ano92m]. Nordhoff [Ano77o]. Normal [HK98, Lew98, NB77, Shi94]. Normale [Bot77, Gal12, Gre85]. normaliennes>> [Dur21]. normaliens [Dur21]. normals [Mar10c]. Norman [Ano81u, Bir97, Ols76]. Norstedt [Tho76]. North [Ano84g, Ano84h, Ano88e, Ano95j, Buc77, Dru07, Dru12]. North-Holland [Ano84g, Ano84h, Ano95j]. Norton [Abe09]. Norway [BJ89a]. Norwegian [Lüt08]. nos [Pie86]. Notas [Ano84g, Lei09]. notation [FT83, Pow76, Pro22]. Note [Ano80v, Ano03w, Bur10, Dau79, Fra04b, Høy02b, Mar97, Par95, SSS17, Vol19, Abe76, Abe79, Ano75q, Edw91, EG81, FP03, Fri74a, GG92a, Gra80b, Hou87, ST16, Sch93a, Scr75a, SCM19, SSS20, Wat84, Wat86, vBS20]. notebook [Aub14]. Notes [Ale14, Ano79-35, Ano79-36, Ano80w, Ano80x, Ano80y, Ano80z, Ano81-39, Ano81-40, Ano82s, Ano82t, Ano83p, Ano83q, Ano84r, Ano84s, Ano85s, Ano85t, Ano85u, Ano85v, Ano86t, Ano86v, Ano86u, Ano87u, Ano87v, Ano87w, Ano88x, Ano89w, Ano89x, Ano89y, Ano90x, Ano90y, Ano90z, Ano90-27, Ano91v, Ano91w, Ano91x, Ano91y, Ano92-36, Ano92-37, Ano93y, Ano93z, Ano93-27, Ano93-28, Ano94o, Ano94p, Ano94q, Ano94r, Ano95v, Ano95s, Ano95t, Ano95u, Ano96d, Ano96e, Ano96f, Ano96g, Ano97d, Ano97e, Ano97f, Ano97g, Ano98g, Ano98h, Ano98i, Ano98j, Ano99c, Ano99d, Ano99e, Ano99f, Ano00c, Ano00d, Ano00e, Ano00f, Ano01c, Ano01d, Ano01e, Ano01f, Ano02g, Ano02h, Ano02i, Ano02j, Ano03o, Ano03p, Ano03q, Ano03r, Ano04m, Ano04n, Ano04o, Ano04p, Ano05o]. Notes [Ano05p, Ano05q, Ano05r, Ano06k, Ano06l, Ano06m, Ano06n, Ano07f, Ano07g, Ano07h, Ano07i, Ano08f, Ano08g, Ano08h, Ano08i, Ano09e, Ano09f, Ano09g, Ano09h, Ano10g, Ano10h, Ano10i, Ano10j, Ano11g, Ano11f, Ano11h, Ano11i, Ano12e, Ano12f, Ano12g, Ano12h, Ano13d, Ano13e, Ano13f, Ano13g, Ano14c, Ano14d, Ano14e, Ano14f, Ano15e, Ano15f, Ano15g, Ano15h, Ano16h, Ano16e, Ano16f, Ano16g, Ano17d, Ano17e, Ano17f, Ano17g, Ano18e, Ano18f, Ano18g, Ano18h, Ano19e, Ano19f, Ano19g, Ano19h, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano21f, Ano21g, Ano21h, Ano21i, Ano22e, Ano22f, Ano22g, Ano22h, Ano23e, Ano23f, Ano23g, Bay11, BH89, Kis99, Men86, ALB10, Ano78h, Ano81u, Ano89g, Ano95l, BF89, Dra84, Fri91, GG75c]. notes [Gra99a, Kin79, Law05, Mac76, Maz21, Neu88, Sai74, Tee84, Vol02, Wil83, Boy74, Tee84, Gil77]. Notice [Cho75, Jay24b, Fre77]. notices [Ano74-65, Ano75t, Ano76-31, Ano76-32, Ano76-33, Ano77-34, Ano77-35, Ano77-36, Ano77-37, Ano81-38, Ano81-37, Ano82r, Ano87t, Ano89v, Mar93b]. Notion [Fer98a, Pou01, Chr84, Cog16, GA20, Pie80]. notions [Mal06b]. Nottingham [Wil77b]. notwithstanding [Kni01]. nouvelle [Eva94, Spi00a, Spi00b]. nouvelles [Mac76]. Nova [Ano84-29]. novae [Mal13]. Novara [Bur08, Cor16]. Novecento [Ben95]. November [Ano76e, Ano83n, Ano89-34, Ano92-44, Ano92-48, Ano94b, Arb86, Bin90, Lam87, Lew99, Pet82, Pur90, Str81, Ano92-52, Ano21l, Ano22k, Ano23h, Kas90b]. Novembertagung [Ano92-52, Ano94b]. novogo [Ano77l]. Novus [Lüt08]. Novy [Bir74, Ano90r]. Nu [Yon74]. Nuclear [Sho02, Eck13]. nuevo [Ano92-44]. nuit [Ric91]. Number [D’A10, Edw77b, GG97, Gra86, Kis99, Lev97, Bak75, Bur86, Fie22, Gra94b, Gra94c, H^{+}08, Hog89, Ken74a, Lê20, Mal06b, Neu88, Spi00b, Kna88, Ano82f]. numbered [Ber87]. Numberland [Abe09, Wil08c]. Numbers [Ano81k, GG96, GG97, Mor19, Sch05, Tat11, AGSL76, Ano79w, BS09, BF23, BH89, Bur92, Cer17, FS22, Fer95, Fla03, Glu77, Hog85c, Hug74, Jay24a, Jay24b, Kar22, PV22, Pro22, Pyc97, Ras89, Sin85a, Tuo23, Van22, Wan22, NB77, Sen92, Bra17, Ano85h]. Numeracy [VF22, Rei86]. numerals [Str82b]. numeration [Ano76y, Van24, Knu83b, Knu83a]. numérations [Ano78i]. Numerical [Bir78, LMP01, LY80, Mai21, dV22]. numéricas [Pan08]. numeris [Gra83, Ace11]. números [dL01]. Nuncius [Rom97, Rom05]. nuova [Taz09]. nutation [Wil08b]. NY [Ano91f, Ano95l].

O [Ano74o, Ano77b, Ano77c, Ano80r, Ano81k, Ano84q, Ano08e, Ano10k, Ber74, Bid74, Bra07, BT81, DR78, Dra78, Enr84, Fol94, Hes90, JET84, Lew04b, Par05b, SY90, Swe79, VK05, Ano74-57, Rod88]. O.F.M [Gra83]. Oak [Ano86e]. Oberschelp [Lef10]. Oberwolfach [Ano92b, Art91, Fol76, Pur88, Pur90, Pur91, Bin83, Bur75, Con85, Eng81b, Eng81a, Pur86a, Pur86b, Pur88, Pur90, Pur91, Scr85c]. Oberwolfach-Walke [Ano92b]. Obituaries [Ano75u]. Obituary [Ano74-65, PG24]. object [Kah75]. objective [Das94, Ver15]. objectivity [Pat14]. objets [Lam19]. Obras [Ges18]. O’Brien [Smi82]. Observations [Hog87, She96, She94]. Observatoire [Ano90q]. Obtain [Bec99]. obtaining [Ger85]. Occasion [Ano79-42, Ano90j, Ano90l, Ano95f, Ano81e, Ano81-35, BGM^{+}76, BDG^{+}86, Fol84, Shi94, Str82a, You83]. Occidental [Ano92g, Che94]. Ocherki [Smi79b]. oCorps [Cor96]. oct [Tat86]. October [Ano77d, Ano84-27, Ano88r, Ano90c, Ano90s, Ano91k, Ano91q, Ano92h, Ano92s, Ano92x, Ano92v, Ano93x, Ano94g, Bee14, Bin91, BDD^{+}88, Pep90, Rod88]. Octobre [dMC86a, Ano85p, Ano92-27]. Odense [Ano77g]. Odessa [DK86]. Odyssey [Dry14, Duc16]. oeuvre [Ano87h, Ano90h, Bie78, Bor91, Cog15, ET14]. OEuvres [Fra04a, Kle17, Moo82b, Wil08b, GG76a]. office [Tro74]. officielle [Bré12]. ogni [Gal12]. O’Hara [Gar05]. Ohio [Ano92x, Hug76]. Ohm [Ano88f]. oiseau [Ric91]. Oklahoma [Ano92g]. Oktober [Ano94b]. Oktober-2 [Ano94b]. Olaf [Ano77g, BB97]. Olaus [BG21b]. Old [FR98, Høy02b, Mel06, Fri81b, Høy02a, Jon75a, Pow76, Sho03]. Oldenbourg [Haf16]. Olinde [Tur08, AO05]. Olindo [RFS89, RFS89]. Olli [Gra10b]. Olms [Ano74q, Bul22, De 20, Her14, Sch16]. Olof [Scr92]. Olschki [Cap13, Coz12, Rem12, Rom97, Rom05, Taz09]. Omnia [Ter89, Ano77m, Bar77, GG05a, Pei22, Sch00b]. omnimetre [Kid09]. Omnisciene [Fol12]. On-going [Kat11]. Once [Ber90]. One [And11, HK98, NB77, PP78, Wag15, Ang79b, Del06, Lê20, Roy17]. only [Erl92]. onore [Ano91k]. Ontario [Høy90a, Nic13, dBR76]. Oosthoek [Aup76, GG77]. op [Lo 85]. Open [Law05, Fle75, Moo08, BGG06a]. Opening [Ano75v, Kli75]. Opera [Ano77m, Net24, Pei22, Rid83a, Rid83b, Ter89, Bar77, Sch00b]. operations [Ano74l, Rid94]. opératoire [Pan88]. opere [Coz12]. opportunity [SS10b]. Opposite [IM98]. opposition [PV22]. Optics [Atz06, Dij04, PR22, Ros76, Wie23]. optimalization [Ano79v]. Option [Kno88a, Tat86]. Opuscula [Ano87d]. Oral [Ano79-37, Tro75]. Orbital [Nau98]. Orbits [Wei98, Wei92]. order [Eng80, Gup74, HZ85, Hay87, Sch77]. Ordered [Sch07, HP05]. ordinaires [Tou03]. ordinary [Grc11, Sch77]. Orell [Bar77]. Oresme [Von93]. Oreste [Coz12]. Organisé [Ano89n]. Organisée [Bor89]. Organized [Ano87n]. Orient [Ano85e]. Oriental [Hog24, MY99]. Orietta [Fra06]. Origin [Bul16a, PS97, Ano77r, Cam78, HNE21, Tur13, Wil76]. Original [Ano84f, Ano89k, Dav90, Ano78g, Enr76, Glu12, Gra08, Mah78, Ano77i]. Original- [Ano89k]. originator [PZ11]. origines [Bru07, Pan05]. Origins [Ano93h, GM04, MP06, Sha94, ABBS09, CR22, Cre14, Dho85, Dru91, Fre76, GP91, Gra78a, GG90c, Knu74, Mar10c, MG81, Pan03, Pan05, Pyc81, Rus75, Sai19, Tro76, Vac20, Wol08, Sch95, Bir74]. Ornament [DS99c]. Oronce [Ros76]. orthogonal [SM18]. Ortiz [Tur08]. Osaka [Ano78-32]. Osen [Wil75b]. Oskar [Hei09]. Oslo [Bir96, Lüt08]. Osram [KS12, Tob10, Tob10]. Ossa [Maz21]. Osservatorio [Ano87c]. Ostens [Bus76]. Osterformel [Lic97, Ber90]. Österreichische [Scr03]. Österreichischer [Ano90-36]. Österreichisches [Bin89]. ot [Sti76b]. Otdelenie [Ano81t]. Other [Ayo06, AKW04, CB21, GG76c, Knu83a, Knu83b, Law05, Zas79]. otros [Pan08, dL01]. Ottawa [Ano77-29]. Otte [Ano82c]. Otto [Gol76, Rus75, HP20, Rad03]. Ottoman [Inö06]. Our [Ste01, GG04, Pie86]. Outgoing [PH99]. Outline [Pee95, Ano77o, But81, Dor95, Wil79b]. ouverte [PP91]. ovals [Mar10c]. Overcomes [Bec01]. overlooked [Lüt09b]. overspecification [Car18]. overview [Mar20]. Owen [Cas77]. Own [EU99, Ste01]. Oxford [Ano78l, Ano79r, Ano83f, Ano85-28, Ano87k, Ano89d, Ano89c, Ano92x, Ano92-34, Bee06, Bra17, Bru77, Buc77, Cri82, Duc16, Eva13, Gil75, Gra10a, GG90b, GG05c, Gra06, Har77a, How18, Hun12, Jon85, Kat20, Kje17, Maa07, Mic23, Moo82a, Moo03, Mor19, Mos16, Nor07, Nor03, Oli08, Pec20, Pin76, Por05, Ric82, Row93, Sen10, Ste05, vM07, Ano92-38, FFW00, RS09, Ste05]. Oxford/New [Ano89c]. Oxyrhynchus [Swe01]. Ozhigova [Sel74].

P [Ano74v, Ano77l, Ano77m, Ano79l, Ano81o, Ano88h, Ano91a, Ano92k, Ano92m, BKM^{+}82, BT81, Dav76, DS99b, Gra94b, Gra94c, Gra97a, Hes90, Jon76a, Ken76d, Kil83, Kin86a, Nak03, PP78, Sel74, Sen84, Sid24, Ste11, SY90, Tat75, Ter89, Thi08, Tho76, Van05, Vit96, ZvDF22, Zer75, Ano74q, Ano74n, Ano74x, Ano74y, Ano74-28, Ano77i, Ano77n, Ano77s, Ano77o, Ano77g, Ano82f, Bid74, Bie74a, Bun76a, Cal75, Cas77, Cib76, Clu76, Cog12, Dav75, Eis75, Gol76, Goo77, Gri76, Her75a, Hir74, Hug76, Ken76b, Lin74, Mon75, Moo75a, iÔ75, Ped75b, Pin76, Pye75, Pye76, Rus75, Sai74, Sch76a, Sch12, Ser18, Str78, Ume76, Vek75, Vol77, Wil75a, Yok75, Yon75, vdW76a, Ano82f, Aga75, Ano74t, Ano74v, Ano74-29, Aup76, Bir74, Cha76]. p [Cro77, Dav76, Edw77b, For75, Fre77, Gil75, Ken76a, Ken76d, Ken76c, Knu74, Maz77, Por75, Rei76, Ros74, Seg76, Tru75, Vol77, Wil77b, Zer75]. P. [Bot78, BJ89b, For98, Ano77k, Ano77h, Ano77l, Ano77r, Ken75]. P.-L [Ter89]. P.O [Ano74w]. PA [Zas83b, Zas83a, Ano81q]. Pacific [Par92b]. Pacioli [Rey06]. Padua [Maz77]. Paganica [Ano90f]. Pages [Ano21j, Ano21m, Ano21k, Ano21l, Ano22j, Ano22i, Ano22k, Ano23i, Ano23j, Ano23k, Ano23h, Ano24h, Ano24f, Ano24g, Ano24i, Ano95j, Bul22, Fer04b, Ges18, GG97, Høy95, Lor18, Maz24, Pri80, Tat11, Ano91d, Bra17, Kat20, Mor19]. Paideia [Bie74a]. pair [Hog85c]. Palace [DS99c]. Palaiseau [Gra08, Taz15]. Palaisseu [Ano87i]. Palazzo [Ano92s, GG05b]. Palermo [CB21, Par97, Mar04, Isr85, Ysh90]. palimpsest [N^{+}11, Sch13b]. Palladino [Taz09]. Pamphlets [Dru07, Dru12, Abe01, Abe10]. Pamplona [Pan08]. Pan [Shi86, Shi88]. Pan-African [Shi86, Shi88]. pandiagonal [Hay87]. Panza [Bru07]. Paola [Sch99a]. Paolo [Ben88, Fol14, PZ11]. Papadoupoulos [Vol16]. Paper [Ano82f, Ano84i, BT81, Sch95, Vol93b, Ano74n, Ano81l, Ano86e, BL76, Dav90, Goo77, Meh76, Osw17, Rus80, Sti74, VHK16, Zas83b, Ang88, Goo77, Haw77a, Haw77b]. Paperback [Ano78g, Pec20, Sen06, Ano85h, Enr85a, Fer04b, vM84, Ano87k, Enr85b, Enr85a, Glu12]. Paperbound [Pri80]. Papers [Ano75w, Ano75x, Ano79-38, Ano79-39, Ano79-40, Ano80-27, Ano80-28, Ano80-29, Ano80-30, Ano81p, Ano81-41, Ano81-42, Ano81-43, Ano81-44, Ano82u, Ano82v, Ano82w, Ano82x, Ano83r, Ano83s, Ano83t, Ano83u, Ano84t, Ano84u, Ano84v, Ano84w, Ano85w, Ano85x, Ano85y, Ano85z, Ano86x, Ano86y, Ano86z, Ano86-27, Ano86w, Ano87x, Ano87y, Ano87z, Ano88y, Ano88z, Ano88-27, Ano88-28, Ano89z, Ano89-27, Ano89-28, Ano89-29, Ano90-28, Ano90-29, Ano90-30, Ano90-31, Ano91z, Ano91-27, Ano91-28, Ano91-29, Ano92-39, Ano92-40, Ano92-41, Ano92-42, Ano93-29, Ano93-30, Ano93-31, Ano93-32, Ano94s, Ano94t, Ano94u, Ano94v, Ano95w, Ano95x, Ano95y, Ano95z, Ano96h, Ano96i, Ano96j, Ano96k, Ano97h, Ano97i, Ano97j, Ano98k, Ano98l, Ano98m, Ano99g, Ano00g, Ano01g, Ano02k]. Papers [Ano02l, Ano02m, Ano02n, Ano03s, Ano03t, Ano03u, Ano03v, Ano04q, Ano04r, Ano04s, Ano04t, Ano05s, Ano05t, Ano05u, Ano05v, Ano06o, Ano06p, Ano06q, Ano06r, Ano07j, Dry14, Duc16, Nak03, Ane88, Ano74-62, Ano79i, Ano82l, Ano84n, BGG06a, Bel07b, Bot77, Daw86, Edw77b, Eis75, Fri74a, GG75b, Neu89, Pat75, Knu74, Par91]. papiers [dMC86b]. Papyri [Swe01, Ano74s]. Papyrus [Coo10, Gil78, FT83, Mia08, BS77b, Gil79]. para [dL01]. parabolae [DN12, Fol14]. Paradigme [Ano95n]. paradojas [Aus16]. paradox [Ano74-29, Fra16, GG78, MG81, PK02, RT81]. paradoxes [Ano81z, Gar85, GG80]. parallel [Ung78]. Parallelen [Lau92]. parallèles [Roz90, Vol16]. parallelism [Ano84f, Cog16]. parallelisme [Fre77]. parallels [Lau92]. Parame’svara [Gup77c]. Parameter [Fri81a]. Parameters [Taz97]. Paraná [Per93]. parfaits [Pic89, Ras89]. Pariṡad [Gup77b]. Paris [Ano77j, Ano77n, Ano78i, Ano79l, Ano79m, Ano81j, Ano81n, Ano87h, Ano87i, Ano88o, Ano88u, Ano89g, Ano89i, Ano89m, Ano90g, Ano90k, Ano90q, Ano90-34, Ano91j, Ano91q, Ano91s, Ano91r, Ano92k, Ano92-50, Ano92-28, Ano92-27, Ano92-35, Ano93g, Ano95i, BKM^{+}82, Ber87, Bré12, Bre07, Bru07, Bun76b, Cav95, Che94, Cor96, Dau93b, Dho88b, DS^{+}02, Dra84, DR03, Dur21, Ebe96, Eli24, Enr76, Eva94, Fer11, Fol14, Fra04a, Fre77, Gau13, Gis14, GW94a, Gol13, Gra94a, Gra76a, GG76a, Gre85, Har06, Hei81, Høy07, Kle17, Kno88a, Kno07, Koe94, Krö06, Krö15, Krö17, Lew98, Lor16, Mac76, Mah78, Mal97, Maw06, Maz16, Maz24, Mue96, Ped77, Pet82, Ple78, Pys81, Ric91, Roz90, Sal87, Sch76a, Sim96, Syl07, Tat86, Taz15, Tru75]. Paris [Van06, Vil96, Vil97, Vol16, Wil08b, Yus10, Ano80d, Ano89-31, Dho90, Dug82, Man89, Vog83]. Paris-Sud [Ano90-34]. Parisian [Del03a, Del02]. Park [Ano74w, Ano86e, Ano86s]. Parker [Ano74s, Par07]. Parme [Syl07]. Parshall [Kje17, Neu08, Nor03]. Part [Gra97b, Lo 84, Moo97, iÔ75, Sch98, Str82a, Tee84, Nag91c]. partes [Lo 84]. partial [Ano80g, Eng80]. partially [Buc80]. Participants [Ano75y]. particular [Gra76b]. parting [Scr85b]. partir [Bré12, Mar93b]. partis [dMC86b]. partition [Ant95]. Partitions [Ano81v]. parts [Sch86b, Van79, Swe79]. Pascal [Mar10a, Taz09, Des10, Fri74b, LY80, SD08, Ste12, Mar10a, San05]. Pasch [Ano78j, Sch13a]. Paso [Ano76e]. Passage [Erl03, Kno06]. passages [Høy95]. Passeron [Kle17]. Past [Ano90t, Bul16b, GG04, Dau94]. Pastos [Ano91-30]. pasturage [Jor81]. path [Mai08, Wie23]. Pathways [Bra12]. Patna [Yan83, Gup76a]. patria [Arc17]. Patrice [Ano81m]. Patricia [Ano92j, Law98]. patrie [Gis99a]. patronage [Hou89]. pattern [Daw16a]. patterns [Cro75a, Cro75c]. Paul [Ano77k, Ano77q, Ano87f, Ano89l, Ano95k, BT81, Bun76a, Cri24a, Gis96, Gol76, Gra96, HB^{+}13, Lan07, Lo 85, Mac76, Maz21, Ped77, Row15, GGE82, HB^{+}13, RC85, RL16]. Paul-Marie [Ped77]. Paulette [BKM^{+}82]. Paulo [Ano87-31]. Paulus [D’A16]. Pavlovich [Ano95e, BGM^{+}76, BDG^{+}86, Dem86]. Pays [dMC86a, Roz90]. pbk [Pec20]. peacock [Kat92, Bec80, Dub77, Pan87, Pyc81]. Peano [Taz09, Ano83e, Dru86, GG05a, Ken74b, Ken74a, Aga75, Ang88, Cib76, GG05a, Mon75, Ume76]. Pearls [Mro16]. Pearson [Cas80, Hun07, MB20, Por04]. pebbles [Kim79]. Peckhaus [Hei09]. Pédagogique [Ano77k]. Pedagogy [Ano80p, Ano86n, Ano79-30, Ano93n, Boo86]. Pedemonte [Fra06]. Pedersen [Ano77g]. Peet [HP20]. Peggy [SG09]. Peiffer [Hes90]. Peirce [Ano76i, GG86, Pys81, Ano74-34, Ano79-44, Ano79-41, BLW77, Dau82b, Eis82, Fis82, Gan85, Hog91, Hog08, Jon77b, Ken79a, Ken79b, Ket82, Nag82, Put82, Ano77s, Tim09]. Peletier [Axw18]. Pelikan [Zla23]. Pell [GC82, vM07, MS05]. Pembroke [Cri82]. pendulum [She94]. Peninsula [Hog24]. Pennsylvania [LY79]. pensées [Mac76]. Penser [Gre85]. Pensiero [Arc99, Net02]. Pensions [San06]. People [Ano91g, Dav90, Pre24, Swe92, Rei86]. perception [Ric80]. percorso [Mai08, Sch13c]. Percussion [Ano78h]. perdu [Chr91]. peremennogo [Smi79b]. Perfect [Sor13, Zla23, Ras89, AH16]. Perga [Sai05, FU01].

Pergamon [Moo82a]. Peri [Høy95]. Period [Pep90, Fuk87, Wis76, Gre83, Vuc99]. periodic [Hol15]. periods [Bul09a]. Permanence [Toa21]. permanente [Bré12]. perpendiculars [Kic12]. perpetuity [VV14]. Persian [Bre19, Her75b, Ano88k]. persischer [Ano85e]. Persistence [Row85c]. Person [Ano76-35]. Personae [Asp19]. personal [Ano91g, Dau80, Moo03, NT05, Row85a]. Personen [SS08, Tob06]. Perspectiva [Ste80, DB23, Ste80]. perspectivae [Ste80]. Perspective [Ano77-30, Deb13, Gra09b, Ano74-27, Ano83j, DB23, Hoc96, Aup76, Jon77c, Tro74]. Perspectives [Mol81, Ano88g, Lüt08, Sca94]. perspectivist [HNE21]. pertinentes [Bar77, Rid83a, Rid83b, Sch00b]. Peruana [Ano84y]. Perugia [Fre89]. Pervyi [Ano77l]. Pesic [Con06]. Pesin [Bar74]. Peter [Ano95m, Con06, Coo96, Høy90b, Pin76, Plo06b, Sel91]. Peters [Zit07a]. Petersburg [D’A10, Cal96]. Petersburger [Ano81x]. Peterson [Gra80b, Phi79b]. Petroleum [Kin86b]. Petsche [GG10, Lew09]. Peurbach [Mal13]. Pezzo [Taz09]. Pfalzakademie [Ano91-31]. Ph.D. [Tob06]. pharaohs [Ano74y]. pharaonique [Cav95]. PhD’s [AH10]. phenomenon [Gai75, Ust74]. Pheru [Knu11]. Phi [Ano74-66]. Philadelphia [Ano76w, Lo 84, Row85c, Syl81, Vol77]. Philip [Ang79a, Ano87k, Ano91d, Hil83, Maa07, McG74, Ano79h]. Philippe [Bra09b, Bre07, Krö15, Vol02]. Phillip [Mos16]. Phillips [Sch89, Wil77b]. philology [Ung22]. Philomaths [WW80]. philosopher [You83]. philosophers [Fek14]. Philosophia [Krö06]. Philosophiae [Pei17, Gab76]. Philosophical [Ano93i, Ano95k, Gra98a, Jah93, Lo 84, Rad00, Row85c, Syl81, Vol77, Böt12, HB^{+}13, Ott89]. Philosophie [Ano74-32, Ano91q, Ano92k, Che14b, Déc10, Fre76, Hei09, Jon77a, Mal97, Mue80, Tob81, Ano74-33, Ano75f, Ano78k, Ano89n, Ano90d, Cha81, Cha87, Gre85, Pei17, PP91, Ano88i]. Philosophique [Ano77j, Cor96, Ebe96, Kno07, Mac76, Mah78, Pon78, Vil96]. Philosophiques [Syl07, Lam19]. philosophischen [HB^{+}13]. philosophischer [Ano88i]. Philosophisches [Scr07]. Philosophy [Abe91, Ano74-32, Ano74-47, Ano77r, Ano78k, Ano79p, Ano85k, Ano86o, Ano87-32, Ano91t, Ano92v, Ano93h, Ano95g, Ano95j, Atz06, Cha81, Cri24a, Gar86a, Gar92, GG98, GG02, GG03, Gra98b, Gup79, Hoc96, Hog03, Jon77a, Kas90a, Kat97, Lon21, May75c, May77c, McL12, Mol81, Ott88, Par02, Sau06a, Scr82b, Tob81, Ung85, Wil05, Zep12, Ano74-33, Ano75f, Ano76f, Ano76-34, Ano79y, Ano84c, Ano85i, Ano87-35, Ano90v, Ano91d, Ano93n, Ber85, Bow87, Cha87, Cro77, Déc10, FG06, Fin11, Haw81, Lüt09a, Mar00, Nag82, Pul05, Pys81, Scr84, Sep05, Sti76b, Sti78, Ver13, Oli08, Bun76a, Bun74]. Photo [SS97a]. photocopy [iÔ75]. Photographic [Ano74-67, Ano74-75]. photographs [Ano87f, Ano88h, Moo82a, Moo82b]. photomechanic [Bos77]. Photos [Moo75a, Ano93f]. Phragmén [Vol02]. Physical [Ano75p, Ano87j, Ano91f, Cal83, Das81, Fol15, Hog85b, Høy09b, Jah12, Ser18, Swe79, Ano84j, Mah84]. physicam [Sch00b]. physicians [Ano07e]. physicien [Bor14, PR12]. physiciens [Ano07e, Sau09]. physicist [Eck13, Bru77]. physico [Roz75, Roz76]. physico-mathematical [Roz75, Roz76]. Physics [Ano89o, Ano90m, Buc77, Cor04, Fef97, Hoc96, Pye74, Sau06a, Ano88f, Aub14, BGG06a, Dra74, HEH^{+}04, Gra99b, McC90, Tho76, Pye76]. Physik [Cor04, Sau06a]. Physiker [Hes90]. Physique [Ano91q, eFD88, Ano86g, Ano90d, Ano90q, Ano91n]. Pi [Abe93, CS81]. Piatkiewicz [BM96]. Piazza [Ano90f]. pickings [Gab79]. Pieces [Dru07, Abe01, Abe10, Dru12]. Pieper [Edw80]. Pier [Ano85g, Fol14, Net24, Gra96]. Pierce [Ano76j]. Pieri [Ano89u, Ing11, Mar93a, Mar06]. Piero [Van80]. Pierre [Ano74-28, Ano79m, Ano81j, Ano81n, Ano87i, Ano89m, Ano91q, Ano91s, Fer98a, Gil15, Hes90, Kit81, Kno88a, Mac76, Mah78, Maw06, Maz24, McL12, Sen99, Tat86, Ter89, Tru75, Hah11, Mas08, Tat91]. Pierre-Michel [Maz24]. Pierre-Simon [Hah11]. pierres [d’E01]. Pietenpol [Hes90]. Pieter [Ano87g]. Pietro [Arc09, Arc17, Ano87-27, Est06, NS94]. Piles [Sho02]. Pincherle [BF89]. Pingree [Plo07, Vol77, vdW76b]. pioneer [Wil77a]. pioneering [SM18, AH10]. Pioneers [Haw03]. Pisa [Gal12, Hug07, Mal06a, Nag91b, Nag91c, Net24, Bot77, DF10, Han07, Mar04, Spi00b, Van83a]. Pisani [Sch96]. pisano [DF10, Rem12, Ray23]. Pise [Spi00a, Spi00b]. Pitane [Ano95n]. più [Coz12, Tra10]. Pizzamiglio [Ano85g]. Pizzetti [PZ11]. place [Dau82b, Fis82, Haw84, Ket82, Nag82, Pow76, Pro22, Scr75b]. place-value [Pro22]. plan [LP10]. planar [KQS85]. Plane [Atz04, Zor96, Ano82y, Row21]. planétaire [Gol82]. planetarum [Mal13]. Planetary [Van98, Vek75]. planetes [Gol82]. planned [Ano81-46]. Planning [Fef97]. plans [BT81]. plate [Ano77m, PP78]. Plates [DR19, Ano74s, Ano74z, Ano78i, Gol76, Gol82, Sid12, Swe79, Van80, vM84, Ano77i, Bar82a, Bar82b, Hod77, Pin76, Ste80]. platform [lLW13]. Plato [Ano90u, Bun74, Hoc96, AS93, Fen09, oSTT79, McK82]. play [Asp19]. Plaza [Aus16]. pleasing [Ken76a]. plein [Ric91]. Plimpton [Fri81a, Don98, MW17, Rob01]. Plofker [Ano05d, Ano06b, Ano06c, Ano06d]. Plotkin [Sch07]. Plücker [Lor15, Wie23]. plural [Ano75q]. Pneumatiká [Coz12, Tra10, Tra10]. Poet [Bag97a]. poeta [Rod88]. Pogrebysski [Lem74]. Poincaré [Ano89m, Ano90d, Ano90g, Ano90k, Ano91q, Ano91s, Boi89, GG87b, Sau09, Vol02, Wal17, Ano93t, Ano07e, Bas81, GG88, IN84, IM98, Kat81, Che84, Fre76, Dia99, Fel85]. Poincaré-Volterra [IN84]. Poinsot [Ano81m, GG14]. Point [Lor18, McL12, Ano81n, Lüt09a, Vac20, Van24]. Points [Epp95, IM98, Wol03, Lau99, Moo08]. Poisson [Ano87i, Taz15, Dug82, KS13b]. Poitiers [Bor89]. Polacca [Ano82e]. Poland [Ano88a, BM96]. Poles [Mel04]. policy [Ano76p]. Polish [Moo82a, Ste80]. Polishchuk [Ano81t]. politica [Arc17]. Political [Dru07, Abe01]. politics [Seg86]. politique [Sen99]. polygonis [Ace11]. Polygons [Atz04]. polyhedra [LH84, Sen84]. Polyhedron [Sch99b, Krö08, Wei04]. Polylogarithms [Cra13]. polynômes [Ple78]. polynomial [Sti74]. polynomials [Joh84]. Polytechnic [Ano74-56, Ano74-59]. Polytechnique [Ano87i, Dug82, Gra08, Hei81, Taz15, Bil85, Ano89g, Dho11]. polytechnique-Palaiseau [Taz15]. Polytechnique-Palaisseu [Ano87i]. Polytechniques [Rei15]. Poncelet [Bal23, Lor15]. Pont [Bor14, Fre76]. Pontifical [Maz77]. ponyattiya [Phi79c]. Ponzio [Ken76b]. Popp [Eve76]. Popper [Ven02]. Popularizing [Ano91g]. Porter [Hun07, Sch91c]. Portig [Ano74v]. portions [Lo 83]. Portrait [Mel06, Coo86, Høy02a]. portraits [Est79, Cal75]. Portugal [Dho91, Ges18, Rod88, Beb96, Car07, Dom04, GO99, Mar20, Sar15]. Portuguese [Sar88, Sar93, Sar15]. position [Han07, Kim79]. positive [Hog89]. Positivism [dS99a]. possibilité [Ano86g]. possible [Car22, Rey09]. Post [Vuc02]. Post-Stalin [Vuc02]. posterius [Rid83a, Rid83b]. postulate [Ung78]. Postulaten [Rom04]. postulates [Bur92]. Potenzsummenformeln [Sch83]. Poughkeepsie [Row89]. Poulle [Gol82]. £150 [Sch13b]. Pour [Pic89, Ano90d, Ano91q, Ano92r, Bab77, Che94, Gra01, Kno07, Maz16]. power [LN94]. Pp [Mal97, AH10, Ale04, Ano74o, Ano74p, Ano74r, Ano74s, Ano74u, Ano74-27, Ano74z, Ano77m, Ano78i, Ano79k, Ano81k, Ano81e, Ano81f, Ano81i, Ano81j, Ano81l, Ano81n, Ano81o, Ano82f, Ano82d, Ano82e, Ano84e, Ano84g, Ano84h, Ano84j, Ano84i, Ano85e, Ano85f, Ano85h, Ano85g, Ano86d, Ano86e, Ano87d, Ano87e, Ano87g, Ano87k, Ano87i, Ano87f, Ano88e, Ano88n, Ano88g, Ano88h, Ano88i, Ano88k, Ano88m, Ano89d, Ano89c, Ano89e, Ano89g, Ano89j, Ano89i, Ano89k, Ano89l, Ano91e, Ano91f, Ano92k, Ano92n, Ano92o, Ano93f, Ano93g, Ano93i, Ano95f, Ano95k, Ano95m, Ano95l, Asp83, Ayo06, Bar74, Bar78b, Bar82a, Bar82b, Bee06, Ber74, Ber87, Bir77, Bir80, Bla05, Bow84, BB04, Bra09a, Bré12, Bru77, Buc77, BT81, BT84]. pp [Cav95, Che14b, Coh82, Coz12, Cri24a, D’A10, Das81, Dav19, Dav90, Dho11, Don05, Dra84, DR03, DR19, Dur21, Enr80, Enr85b, Enr85a, Ewi92, Fer11, Fol12, For10, Fra04a, Fra10, Fre92, Gab76, Gal12, Gar05, Gau13, Gin83, Git74, Glu12, Gol82, Gra10a, Gra11, GG86, GG90c, GG90b, GG90a, GG10, Gra90, Gra97a, Gra06, Gra10b, Gra12, Gra18, Gri82, Grö86, Gur24, Har77a, Has04, Haw77a, Haw77b, Haw81, Haw03, Hel83, Her14, Hod77, Hog85b, Hog03, Hog24, How18, Høy24b, Hud12, Hug05, Hun12, Isr85, Jay83, Jon77c, Jon85, Kat92, Kat11, KS12, Ken78b, Ken82, Kil83, Kla84, Knu11, Krö06, Lef10, Lei79, Lem74, Lev97, Lew90, Lew95, Lo 84, Lo 85, Lon21, Mac76, Mar10a, Mar95, Mat74, Mat75, McK82]. pp [McL12, Mon83, Moo77, Moo82a, Moo03, Mue80, Nak03, Net24, Neu86, Nor03, Oak12, Pan18, Par89, Par92a, Par92b, Pec20, Ped77, Per10, PP78, Ple78, Pys81, Rem12, Ric82, Ric84, Rid88, Row85a, Row90a, Row93, Row19, Roz90, Sag91, Sal87, San03, San06, Sca94, Sch91a, Sch91b, Sch91c, Sch11, Sch90b, Sel74, Sen10, Shu80, Sid12, Ste80, Str82a, Swe79, Syl81, Syl83, Syl03, Tan74, Taz09, Tee78, Ter89, Tro74, Tro83, Ung85, Van80, Van06, Vil97, Wal17, Wil83, Wil75b, Wis76, Wol91, Wol03, Yan83, Ysh90, Yus10, Zas83b, vM84, Ang79a, Ano74m, Ano74l, Ano74-30, Ano78g, Ano78h, Ano78j, Ano78-32, Ano79m, Ano81s, Ano81t, Ano81p, Ano81h, Ano81v, Ano81r, Ano81u, Ano82c, Ano87j, Ano88f, Ano88j, Ano88l, Ano89h, Ano91g]. pp [Ano92l, Ano92m, Ano95g, Ano95n, Bar77, Boy74, Bun76b, Cal83, Cal92, Cas80, Che84, Dho88a, Dil81, Dru91, Dyc75a, Edw77a, Enr76, Eve76, Gra76a, Gra83, GG77, Gra91b, Gra91a, Gre85, GS91, Hei81, Hil83, Høy90a, Kin86a, Kit81, Kno88a, Knu83a, Mah78, McC90, McG74, Nee79, Ora82, Par91, Phi79a, Pin80, Pla76, Row83a, Row83b, Sch89, Sel90, Sen91, Ser90, Smi79b, Str81, Tat86, Tee84, Zas83a]. pp. [Ano77j, Ano77q, Ano79j, Ano81g, Ano84f, Ano84d, Ano87h, Ano92j, Ano93h, Ano95h, Ano95i, Bel12, Bra12, Bun74, Che94, Coo95, Dau93b, Dru12, Dum18, Dyc75b, GG76a, Hal84, Høy07, Høy23, Lüt20, Nab04, Ols76, Par97, Pet82, Rid83a, Rid83b, Rom97, Rom05, Row85b, Row85c, Row85d, Sch19, SG04, Spi76, Too80, Ull03, Was20, Zep12, Kno80]. Practica [Høy09b, FH08, Hug10]. Practical [Syl81]. Practice [Özd00, Imh03b, Toa21, Spi00b]. practices [Zhu20]. Practicing [CO13]. praeface [Clu76]. Praeger [Ano74z]. Prag [Scr04]. Prager [EI12]. Prague [Bir74, GC21]. praticque [Spi00a, Spi00b]. pratike [Syl81]. Pre [AH10, Vuc99]. Pre-1940 [AH10]. Pre-Stalin [Vuc99]. Preceptors [Del03b]. Précession [Wil08b]. précis [Che94]. Precursor [Enr83, Glu77]. predecessors [Sam24]. prediction [Moz13]. Preface [Ano85g, BG75b, Fre82, Moo82a, Ped77, Ano79m, Høy07, Ano79l, Dur21, Wil83, Mar93b]. préhistoire [Roz90]. prehistoric [Pet85]. Prehistory [Ine88, Ano84j, OA05]. Préhoude [Spi00a, Spi00b]. Preliminary [GG75c, Bie74a, De 05]. Premier [Ano87-28, Cré87]. Premiere [Dra84, Ano93-33, Dho91]. Premiers [Fra04a, Pic89]. premises [Lê15]. Prentice [SG04]. Prentice-Hall [SG04]. preparation [Ano82-29, Ano92-51]. Present [Ano90t, Gup86, Bul16b, Dau85b, Rid88, Bun74]. Presentation [Vol02, DP88]. Presented [Her75a, Ano82l]. preservation [BEJ^{+}75]. preserved [Del05, Eli18]. Press [Abe06, AH16, Aga75, All04, Ano74o, Ano74s, Ano74t, Ano74l, Ano74n, Ano74x, Ano74y, Ano74-28, Ano74-29, Ano74-45, Ano77s, Ano77g, Ano77f, Ano78h, Ano81l, Ano81q, Ano81u, Ano82d, Ano84f, Ano84d, Ano87d, Ano87e, Ano87k, Ano88e, Ano88f, Ano88g, Ano88h, Ano88l, Ano89c, Ano91d, Ano91e, Ano92n, Ano92o, Ano93f, Ano95g, Ano95k, Ano95l, Asp83, Bar74, Bay11, Bee06, Bel07a, Bel08, Bie74a, Bir83, Bla76, Bra12, Bra17, Bru77, Bul16a, BT84, Bun76a, Cib76, Coh82, Con06, Cri24b, Cro77, Dau12, Dav75, Dav19, Dho85, Duc16, Dum15, Dum18, Eis75, Enr80, Eva13, Ewi92, Fer04b, Fra10, Fra17b, Gab76, Gil75, Glu12, Goo77, Gra83, Gra10a, Gra11, Gra13, GG86, GG90b, GG05c, Gra06, Har77a, Haw77a, Haw77b]. Press [Haw81, Hel83, Her75a, Hir74, Hoc96, Hod75, How18, Høy90a, Hud12, Hun07, Hun12, Jam13, Jon85, Kat92, Kat20, Ken76a, Kje17, Lan07, LY79, Lev97, Lew90, Lew98, Lon21, Lor18, Maa07, Mic23, Mon17, Moo77, Moo03, Mor19, Mos16, Nak03, Nee79, Nor07, Oli08, iÔ75, Par89, Pec20, Per10, Pin76, Pla76, Plo06a, Por05, Ric82, Row93, Rus75, Sag91, Sal78, San05, San06, Sch91a, Sch91b, Sch91c, Sch13b, Sch05, Sen84, Sen06, Sen10, SG09, Sia23, Sid06b, Spi76, Ste05, Ste09, Syl83, Tat75, Tim09, Tro74, Tro83, Tru75, Ung85, Wal17, Was20, Wil83, Wil75b, Wil05, Wil07, Yok75, Yon75, vM07]. Press-Oxford [Ano89c]. Press/Birkhaüser [Syl83]. Presse [Fre76]. Presses [Cog15, Gra12, Lüt15, Mar10a, Mue96, Rei15, Ple78]. presso [Rom97]. pressure [Fri74b]. pressure-height [Fri74b]. Preuß [Plo06b]. Preußen [Tob90]. Preußischen [Row23]. prevails [SS10c]. previously [Fol83a]. Price [Ano79k, Hog24, Sch13b, Ano81k, Ano88k, Dav90, Dho11, Fer11, For10, Fre92, GG90c, GG90b, GG10, Gra10b, Kat11, Knu11, Yus10]. Primary [San03]. Primeras [Ano90-32]. primi [Ben95]. primitive [Ano76y, CMT79, Van22]. primus [Ste80]. Princess [Bos10]. Princeton [Ano74-28, Ano84f, Bul16a, Eva13, Fra10, Fra17b, Hun07, Hun12, Kje17, Lan07, Lon21, Plo06a, Sch91a, Sch91c, Sen10, Tat75, Wal17, Wil05, GBGL08]. Princeton/Oxford [Kje17]. principalement [Mar93b]. Principes [Maz16, Pei17]. principia [Ano77j, Gab76, Pei17, Ano86s, Ano87n, Cha77, Erl94, Gui96, Gui03, Pou07, Wei92, Wei98, Fel78, AA88, Ano74-56, Bra89, Erl03, Fel78, Nau03]. Principien [Bos77]. principios [Aus16]. principle [DF18, GG77, LYK85, Toa21]. Principles [Nor07, Rad02, Aub14, Lüt05]. Prindle [McG74, Wil75a]. print [Høy24b]. Printed [Per10, Sai74, Bur09, Ros76]. Printing [Høy95, Moo77]. Prinzipien [Ora82]. prior [Yos75, Yos76]. priority [Zun02, May75b]. prius [Bar77]. Privatdozentin [Row23]. private [Wil83]. Prize [Ale14, Ano75e, Ano08e, Ano21a, Aus13, Sør17, SY90, Aud11, Ano74-45, Ano10f, Ano10k, Fol94, Par05b]. Prizes [BG21a, Fra13, Fra17a]. Probabilista [Ben88]. Probabilistic [Das80]. probabilists [Sho83]. probabilitá [Ben95, BFG94]. Probabilités [Maz16, BKM^{+}82, Cré87]. probabilities [Das94, GP91]. Probability [Bar74, Bel12, Hoc96, Ros20, San05, Ano74g, Ano74h, Ben95, Das79, Das80, Fis94, Fis11, Fre80, Gra94b, Gra94c, Høy83, Kno80, Ney78, Sch76b, Sch91a, Sen91, SS10c, Ver13, Ver15, Cas77, Dav75, Pla76]. Probablement [Ano89m]. Problem [Coo10, DS99b, Ine91, Kje93, Lew98, Wat95, Blå12, Bra89, CO13, Deb13, Erl92, Fen89, FS22, Fri74b, Fri81b, FC22, Høy82, Imh03a, Jor81, Mac77, Mia08, NS94, RL16, Sch77, Sch93a, SSZ07, Vog83, vBS20, Cam77, Gon08, PS83a, Pio85, Sho86, Dia99]. problem-solving [CO13]. Probleme [Ano91u, Lam87, Cav85, dMC86b, Ano74v]. Problèmes [Dug76, Spi00a, Spi00b]. Problemgeschichite [PS83b, PS83a]. Problemgeschichte [PS83c]. Problemi [Hug07, Lam87]. Problems [Bag97a, Dau75, Sim00a, Sim00b, Wag15, Ano82c, Bul09b, Che09, CM22, Dav90, Fra92, Hee14, lLW13, Spi00b, Tur81, Ung23, Wat83, Zho21, dV22]. Probst [Bla05]. Proceedings [Cas77, Kat11, Lew99, GG76b, Ano81p]. process [Ano80-31, Sel75, Sti76b]. Processes [Buc77]. proches [Spi00a, Spi00b]. Prodigy [FG90]. produced [Ano74w]. Prof [Ano86h, Ano91k]. Prof. [Ano91k]. professionalisierung [Tob90].

Professor [Mon83, Str81, Tat92, Ano79o, Ano79-42, Cri99, SS10b, Sti76a]. Profiles [Ano87e]. profilo [Ken78b, Ken82]. progetto [Mai07, Pan09]. program [Ano76h, Ano79d, Ano79-37, Ano89-32, Arn88b, Tho76]. Programm [Haw84, Lê15, Row83c]. Programming [Kje00]. progrès [Enr76]. Progreso [Mil93]. Progress [Kop75, Ano74-36, Ano79w, Ano79x, Ano81y, Ano82g, Ano92-45, GG75d, Haw75b, Hol15]. progresses [Dia75]. progression [Pet85]. proiettiva [LD22]. Project [Ano76-36, Ano79-43, Ano85-27, Ano87-29, Ano76o, Ano80-31, Ehr15, Kin74b, Mai07, Rai92, Wal75]. projectif [Gra12]. Projective [GG02, GG03, Mar06]. Projects [Ano74-68, Ano76-37, Ano76-38, Ano76-39, Ano77-38, Ano77-39, Ano77-40, Ano78-30, Ano82z]. Projet [Ano83v]. Prolegomena [Ano82e]. Promotion [Ano74-69]. promovierter [SS08, Tob06]. prompted [Imh21]. Proof [Con06, Gui96, Bal04, Joh84, Nau03, Pes03, Ung78, Von93, Wag85]. Proofs [Oak19, Ano81l, Llo90, Ung23]. Prop [Kno93c, Ace11]. Propagation [Her75a, Cam77]. proper [Hog91]. property [Erl92]. proportion [Rom13]. proportional [Cra09]. proportionality [CV22]. Proportionentheorie [Mue80]. proportionibus [Rom13]. Proportions [GG96, GG97, Ano89-30, Ano91-30, DG90, Kic15, Koe94, Mas97]. Proportionum [Syl07]. proporzioni [Sai94]. propos [Mac76]. proposal [Ton82]. proposed [Lan75, Pep75]. Proposition [Gui96, Pou07, Erl94, Erl03, MF10, Nau03]. propositions [HF84, Kic12]. prosopographical [Tim04]. prospective [Ken74b, Moo75b]. Prosthaphaeresis [Tho88]. Protection [Nos12]. protsesse [Sti76b]. Proust [Yus10]. prove [LP10]. Providence [AH10, Ano74s, Coo96, D’A10, Fen08, Haw96, Haw03, Lor18, Neu08, Nic13, Rec07, San03, Sch07, Sel91, Tur08, Sho87]. Providence/London [D’A10, Fen08]. Prozeßder [Tob90]. Ps [Rom04]. Ps.- [Rom04]. Pseudonym [HB^{+}13]. pseudosphère [Gra99a]. Psychologism [Gra98a]. Ptolemaios [Høy90b]. Ptolemäus [Gol76]. Ptolemee [Gol82]. Ptolemy [BJ00, Fek14, Sid14, Ano88j, Lon21, Plo06a]. Pub [Hal84]. Pública [Pan08]. Publication [Ano78-31, Clu76, GN92, Ano83c, Ano83v, Imh21, Tim04]. Publications [Ano79-44, Ano95q, Ano96c, Ano97c, Ano98f, Ano99a, Ano00a, Ano01a, Ano02d, Eva94, GG90b, Ped75b, Tan74, Wis76, Ano74-54, Ano75o, Ano76-27, Ano77-32, Ano79-28, Ano80o, Ano81-34, Ano82n, Ano83m, Ano84p, Ano85m, Ano86q, Ano87q, Ano87-33, Ano88t, Ano89q, Ano90p, Ano91m, Ano92i, Ano92z, Ano93r, Ano94h, Ano94l, Ano03m, Ano04k, Ano05m, Ano06i, Dem86, Mar90c, Scr75a]. publish [Ano78e]. Published [Ano85f, iÔ75, Ano92i, HB^{+}13, Lo 83, Sti77]. Publisher [Ano03w, Bur10]. Publishers [Ano77s, Jon76a, Knu83a, Lo 85, Maz21, Moo82a, Pys81]. Publishing [Ano74-27, Ano77r, Ano86e, Aup76, BT81, Høy95, Lei79, Par92b, Pin76, SS97b, vdW76a]. Pulte [Lüt09a]. Punktmannigfalttigkeiten [Ano89k]. Purchase [Ano74w]. Pure [Ano93g, Ene18, Glu05, Row85c, SS10c, Øhr85, Win82]. Purity [Cri24b]. Purkert [Ano74v, Plo06b]. purpose [Roe15]. purposes [Bre19]. Pursuit [Bat06, Lew90, Zit07a]. Pursuits [San03]. Pushing [Lor18]. put [Tat88]. Putnam [Ols76]. puzzle [Ano79s]. PWN [Moo82a, Moo82b]. Pycior [Sch05]. Pyenson [Row85c, Ano75q]. Pygmies [Tuo23]. Pyong [Yin11]. Pyong-Gil [Yin11]. pyramid [Hof87, Wag79]. Pyramide [Hof87]. Pythagoras [LY79, Zhm89]. Pythagorean [Gra13, Fri81a, Gon08, KK11, Lum80, Wag85]. Pythagoreans [Sza74]. Pythagoreer [Dil81, Sza74].

Qāuhāi [Ai04]. Qûhî [Sim96]. Qian [Ano92-31]. qibla [Kc74]. Quaderno [Hug05]. quadrando [Kno91b]. quadrangles [Ano87d]. quadrans [Lo 84]. Quadratura [DN12, Fol14]. Quadrature [Eng77, Lan96, Hog10, Kic06, Kno91b, Scr83]. quadratures [Blå12, Erl94]. quadrilateral [Gup77c, Kic10]. quadrilaterals [Kic12, Ung23]. Qualitative [IM98]. qualunque [Bot78]. quanta [Ano90k, Maz21]. Quantenmechanik [Bor18]. Quantification [Swe00, GG75c]. Quantifier [Wat83]. Quantifying [Cri24a]. Quantistica [Ven02]. Quantitative [WDB96, Yos75, Yos76]. quantities [FS22, GA20, Lau87]. Quantum [Ano80-37, Ven02, Pye75, Sch06a, Sho03]. quarrel [Ehr11]. quarrels [Ste12]. quarta [Ano77m]. Quasi [GM04, Mas97]. Quasi-Concavity [GM04]. quaternions [Ken79a]. quatrième [Ano92r, Ano93v]. Quattri [Ano85f]. Quattrocento [DN12, Fol14]. Qué [dL01, LP10]. Queens [Cam77]. Quelle [Jay83]. Quellen [Høy90b]. quellenkritische [Ull03]. Quelques [dMC86b, dMC92, Déc14]. query [Per80]. Quesiti [Scr03]. Quest [Lew99]. question [Tan79, Zun02]. Questiones [Syl07]. Questions [Ano74-70, Ano76p, GG92a, Gar75]. Quetelet [Aub14]. Quhi [Bre07]. quinque [Van80]. quintic [CWW17, Pat90]. Quo [Imh21]. quotation [Ano92-38]. quotations [De 19]. quote [Ano74-71, BGSS16, Tan79]. quotidien [Bré12]. quotient [Nic93]. Qurra [BH89, Hog85c]. Qustā [Hog85b].

R [Ano74r, Ano74t, Ano74-29, Ano74-40, Ano79l, Ano81k, Ano81u, Ano87f, Ano87n, Ano92k, Ano92m, Aus16, Ber74, Bor18, Cal75, Gil78, GG08, Gre85, Har77a, Her75a, Hes90, Kla84, Mal06a, Mue80, Øhr85, Pys81, Sch89, SS08, Str78, Tim09, Tru75, Van05, Vol81, Wil77b, Wis76]. R. [Ano74-73, Ano86w, Ano87c, BS77b, Gre77, Lew04a, Par05a, Roe15, Scr96, Ano90f, Gri82, Par07]. Rabbi [Eli18, Eli24]. Rabbinical [TG98]. Rabinovitch [Dav75]. raccolta [Ano85g]. races [SSZ07]. Radelet [Ano92j, Hes90, Law98]. Radelet-de [Ano92j, Hes90]. Radelet-de-Grave [Law98]. Rademacher [Ano80-35]. radiation [Sch92]. Radiative [Sho02]. Radius [FT99]. Raf‘ [Dje98]. Raffaela [Ano81r]. Raffaele [Bee09]. Raffaella [Ano85f, Hug07]. Raffaelle [Hug05]. Ragland [Gra94a]. Rahman [Ano88k]. Rainer [Høy09a, Høy23]. Ralph [Shu80]. Ramchandra [Rai92]. Ramist [Tra21]. ramp [lLW13]. Rampinelli [Roe15]. Ramus [Zep12, Gou10]. Randell [Knu74]. Random [Bir97]. Ranking [Abe79]. ranks [Zit07b]. rapides [Che94]. rapport [Auj86, Che94, Ano76-35, Enr76]. rarely [Ine91]. rares [Sen99]. Rashed [Ber87, De 20, For10, Gra76a, Sai74, Sal87, Sim96, Giu99]. Ratio [Ano89-30, DG90, Van10, Høy90a]. Ratio-La [DG90]. Rationale [Sel75]. rationality [Das80]. Rations [Imh03b, Mia08]. Ratios [GG96, GG97, Sen87]. ratnakara [SK24]. Raum [Nab04]. Räume [Ano89k]. Rauminhalt [Hof87]. Rauner [Dho11, Don08, Høy24a, SS08]. Ravello [Nag91c]. Ravenna [Van80]. Ravetz [Her75a]. Ray [Glu11, Tro74]. Rayhān [Tee78]. Raymond [And07, Ano78g, Por05, Ste05]. Razvitie [Phi79c, Sel74]. razvitiya [Sti76b]. Re [Sai19, Gra79, HMR17]. re-appraisal [HMR17]. Re-examination [Sai19]. re-interpretation [Gra79]. Reactionary [Cri24b]. réactions [Mar93b]. Reactors [Sho02]. Read [Hun12, ARR12, Fow92a, Wan22, War10]. Reader [Dav19]. readers [Dau92, Dau94, Scr79b]. Reading [Rey06, Cre16, Tra21]. readings [Sch91b, Gab76]. Real [EU99, AGSL76, Ano78y, Bur92, Lor15, Rom10, Smi79b, Wue05]. realism [Ols76]. Réalités [Ano92-35]. Reality [Swe00, Lüt21]. reappraisal [MG81]. reasons [Mug87]. Reassessment [Rob01]. recalled [DF21]. receiving [SY90]. Recently [Bel02, EU99, Del06, Neu88, Ped75a]. récents [May76]. Reception [Vil98, PR22, Tra10, Tur13]. recezione [Coz12, Tra10]. Rechenbrett [Fol83a]. Rechenbuch [Jay83]. Rechenmeister [Ano92e, Cif00]. Recherche [GG76a, Hei81]. Recherches [Ano89n, Cav95, Ano87h]. Rechnen [Rom99]. reckoning [Cul07, Her75b]. recognition [Lê18]. Recollections [SS17]. reconstruction [Fle91]. record [Hug74]. Recorde [Jam13]. Recording [LGG74]. Recordings [Ano74-72]. recovered [IN83]. recreation [Che17]. recreational [Hee14]. Récréations [BH14, Déc14, Ses13, Che14a, Sch76a, Ses13, WM14, Rei15]. rectification [Blå12]. recto [Gil79]. recursive [CMT79]. rédaction [Dje97]. Rédei [Rec07]. Redfield [Llo90]. rédigées [Mar93b]. Rediscovery [Zep12, Gou10, Jon77c]. réductionniste [Pan88]. Reece [Pye75]. réel [Fol12, Rom10]. réelle [Cor96]. Reevaluating [Gon08]. reexamined [Wei92]. Reference [Lei79, McK82, Mat74]. references [Buc80, Gai75]. referring [Wat84]. refers [Ray23]. reflected [Dau80]. Reflections [Che09, Haw84, Ayo04, Høy04, Zhu20, And07, Moo82a]. réflexion [Mar93b]. Reform [Des09, Cra07, Sar15]. reformatus [Sai94]. réforme [Bré12, Vil96]. réfoulé [Dho11]. refraction [And83]. Refugee [Fle86b]. refutations [Ano81l]. regarding [Pet78b]. Regards [Gau13]. Regensburg [Ano76-29]. Regina [Sch11]. Reginald [Ano74r]. Regiomontanus [Høy24a]. Regional [Roz75]. Regions [Shi93]. regression [Sti74, SM18]. regular [LH84, Sti78]. regularibus [Van80]. Rehbein [Pon78]. REHSEIS [Ano89n, Ano89-31]. Reich [Ano81i, Ano95h, GG13, Thi07]. Reid [Edw77a, Lew95, Ols76]. Reidel [Ang88, Ano77r, Ano82c, Ano84i, Bow84, Pin80, vM84, vdW76a]. Reihe [Bar78b, Bla05, Cri17, Gar05, Jay83]. Reihen [Osw17, Bla05]. Reine [Kit78]. Reinhard [Ano79-42, Coo95, Lüt08, Sen10, Ano89s, FS03]. reinterpretation [Yin11]. reist [Vog83]. rejected [Van79]. Related [Ber04, Dru07, Dru12, Abe01, Abe10, Buc80, Jay24a, Jay24b]. relation [Ano79-30, Ano91-31, Bro75, Che10]. relational [Sam24]. Relations [Ano80p, Fer00, Abe93, BC04, Fer93, SS94, Zho21]. relationship [Ano80-31, Boo86, Sti76b]. Relationships [IM98]. Relativitätstheorie [Row06a, Wue05, Ano95h]. Relativity [Row06b, The84, Wue05]. Relevance [Gup86, Wal75]. relevanten [Scr03]. religion [Sen04]. Religiöse [Dil81]. Religious [Dau12, How18, Cas80]. remainder [Bul09b]. remark [Fel78, Fis79]. remarkable [McC94]. Remarks [BC04, Joh86, Sac75, SY90, Wüs02, Ano75v, Bak75, BM77, Die75, Fel85, Kli75, Mac75, Tho11]. remblais [RL16]. Remembering [KS13a]. remembrance [Scr85c]. Remembrances [Moo82a]. Remes [Ano77r]. reminiscence [Sto76]. Reminiscences [Wal94]. Renaissance [Ano77q, Ano83f, Jon77c, Zep12, Ano85-28, Ano92-32, Ano92-43, Fol83c, Gou10, Mal06b, Mug87, Ros76, Sti76b, Vér01]. Renate [DR19, KS12, Row85b, Row19]. Renatus [Gra90]. Renaud [Krö17]. Rencontre [Ano92-44]. Rencontres [Ano83o, Che87, Hou89]. rendere [Hug07]. Rendiconti [Par97]. renewal [Enr83]. renewed [BEJ^{+}75]. renounce [Lai75]. Renouveau [Ano77k]. Rental [Ano74w]. Rentals [Ano74w]. Reorganization [Ano76q]. Repertorium [Hou87]. Reply [Blå17b, Gra97b]. Report [Asp83, Bin91, Dau82a, Fol76, Mac78a, May75c, May77c, Scr82b, Thi75, Gra87, Neu88, Bin90]. Reporter [Ale04]. Representation [Atz04, Haw03, Goe08, Goe09, Lau92, ET14, Cog15]. Reprint [Ano74q, Ano78j, Enr76, Ano74u, Bos77, Sch76a]. reprinting [Ano78g]. Reprints [Ano89h]. reproduced [Ano79k, Ano81j, Ano81n]. reproduction [Tat75]. Republic [Ano89-34, Ano90j, Ano90-33, Ano91g, Ano91-33, Ott88, Sai74, Scr87, Swe92, Ano90u, SS12b, Ton82, Seg76]. républicain [Gis99a]. Republics [AH99]. Republik [Bos77]. république [Ano89i, Auv13, Gis14, Gis99a]. request [Ano82y]. Rer [Ano79-42]. Research [Ano80-31, Ano88k, Ano92-45, Bek88, Coc98, Eng81b, Eng81a, Fol76, Kat11, Mat93, Pet82, Pur86a, Pur88, Pur90, Pur91, SS97b, Ano74o, Ano74l, Ano89-31, Ano89-32, Ber84a, Bul09a, Cri04, Des07, Fio17, Ken08, Lew75, Mar04, Nur92, Par04, Per93, Tho76, Ton82, Tur20, Zit07b, Bot78, Ano74-30]. research-level [Ken08]. Researcher [Lum97]. Researches [Kis99]. Residence [Ano78-29]. Resisting [Wei98]. Resnikoff [Ber74]. resolution [IN89]. resolving [Row21]. Resorgimento [Arc17]. Resource [All09, Ano81-32, SG07]. Resources [Ano87-30, Arn88a, Neu86]. Response [Ano74-73, Gil78, GG98, Gra98b, Fle86b, vD09]. rest [Ait86]. Restauracionis [Moy19]. Restivo [Ano95k]. result [Lüt09b, Tur13]. Resultatismus [Spa89]. results [Cor16, Dav90]. Rethinking [Pan11]. Retour [Row15, Dho11, dSG10]. retrieval [Lew04b]. retrospection [Str82a]. retrouvée [HS15]. returning [FP03]. Reuben [Hil83]. rev [vM84]. Reviel [Sch13b, Sia23]. Review [Abe06, Abe09, AH10, AH16, Aga75, Ale04, Ale14, All04, All09, And07, Ang79a, Ang88, Ano74o, Ano74p, Ano74r, Ano74m, Ano74q, Ano74s, Ano74u, Ano74t, Ano74-27, Ano74l, Ano74w, Ano74v, Ano74n, Ano74x, Ano74y, Ano74-28, Ano74-29, Ano74z, Ano74-30, Ano77i, Ano77j, Ano77k, Ano77h, Ano77q, Ano77l, Ano77m, Ano77n, Ano77r, Ano77s, Ano77p, Ano77o, Ano77g, Ano77f, Ano78g, Ano78h, Ano78i, Ano78j, Ano79j, Ano79k, Ano79l, Ano79m, Ano80b, Ano80i, Ano81g, Ano81s, Ano81t, Ano81k, Ano81p, Ano81e, Ano81f, Ano81h, Ano81v, Ano81i, Ano81j, Ano81l, Ano81n, Ano81o, Ano81r, Ano81q, Ano81m, Ano81u, Ano82c, Ano82f, Ano82d, Ano82e, Ano84e, Ano84f, Ano84g, Ano84h, Ano84j, Ano84d, Ano84i, Ano84k, Ano85e]. Review [Ano85f, Ano85h, Ano85g, Ano86d, Ano86e, Ano87d, Ano87e, Ano87j, Ano87g, Ano87k, Ano87h, Ano87i, Ano87f, Ano88e, Ano88n, Ano88f, Ano88g, Ano88h, Ano88j, Ano88i, Ano88k, Ano88m, Ano88l, Ano89d, Ano89c, Ano89e, Ano89f, Ano89g, Ano89h, Ano89j, Ano89i, Ano89k, Ano89l, Ano90f, Ano91d, Ano91e, Ano91f, Ano91g, Ano92k, Ano92n, Ano92l, Ano92m, Ano92j, Ano92o, Ano93f, Ano93g, Ano93h, Ano93i, Ano95f, Ano95g, Ano95h, Ano95j, Ano95k, Ano95i, Ano95m, Ano95l, Arc96, Arc99, Arc09, Arc17, Asp83, Atz04, Atz06, Aub14, Aup76, Aus16, Ayo06, Bar74, Bar77, Bar78b, Bar82a, Bar82b, BKM^{+}82, Bay11, Bee09, Bee06, Bel07a, Bel08, Bel12, Ber74, Ber87, Ber97, Ber07, Bid74, Bie74a, Bir74, Bir77, Bir78, Bir80]. Review [Bir83, Bir97, Bla76, Bla05, Bo24, Bor14, Bor18, Bos77, Bos03, Bow84, Boy74, BB04, Bra07, Bra09b, Bra09a, Bra12, Bra17, Bré12, Bre07, Bru07, Bru77, Buc80, Buc77, Bul16a, Bul24, BT81, BT84, Bun74, Bun76a, Bun76b, Bur08, Cal75, Cal83, Cal92, Cal08, Cap13, Cas77, Cas80, Cav95, Cha76, Che94, Che84, Cib76, Cif00, Clu76, Cog12, Cog15, Coh82, Con06, Coo95, Coo96, Coo99, Cor96, Coz12, Cri24a, Cri17, Cri24b, Cro77, D’A10, Das81, Dau93b, Dau12, Dav75, Dav76, Dav19, Dav90, Daw16b, De 20, Des09, Dho85, Dho11, Dia99, Dil81, DS99b, DS99c, Don05, Don08, Dra84, Dru91, Dru07, Dru12, Duc16, Dum15, Dum18, DR03, DR19, Dur21, Dyc75a, Dyc75b, Ebe96, Edw77a, Edw77b, Edw80, Eis75, Enr76]. Review [Enr80, Enr85b, Enr85a, Eva94, Eva13, Eve76, Ewi92, Fel85, Fen08, Fer98a, Fer11, Fer04b, Fol87, Fol09a, Fol12, Fol14, Fol15, For10, For75, Fra04a, Fra06, Fra10, Fra17b, Fre92, Fre76, Fre77, Gab76, Gal12, Gam96, Gar05, Gar98, Gau13, Ges18, Gil15, Gil75, Gin83, Gis96, Gis99a, Gis14, Git74, Giu99, Glu12, Gol76, Gol82, GW94a, GW94b, GA01, Gol13, Gra77, Gra94a, Gra76a, Gra83, Gra10a, Gra11, Gra13, GG05a, GG76a, GG77, GG80, GG86, GG90c, GG90b, GG90a, GG98, GG05b, GG05c, GG07, GG08, GG10, GG13, Gra90, Gra91b, Gra91a, Gra94b, Gra94c, Gra96, Gra97a, Gra98a, Gra98b, Gra99a, Gra99b, Gra06, Gra08, Gra09b, Gra10b, Gra12, Gra18, Gre85, Gri76, Gri82, Grö86]. Review [GS91, Gur24, Haf16, Hal78, Hal84, Har77a, Har06, Has04, Haw77a, Haw81, Haw96, Haw03, Hei81, Hei09, Hel83, Her14, Her75a, Hes90, Hil83, Hir74, Hoc96, Hod77, Hod75, Hog85b, Hog03, Hog24, How18, Høy90a, Høy90b, Høy95, Høy07, Høy09a, Høy09b, Høy23, Høy24b, Hud12, Hug76, Hug05, Hug07, Hun07, Hun12, Hym78, Imh21, Isr85, Jah12, Jam13, Jay83, Jon76a, Jon77c, Jon85, Jur07, Kat92, Kat07, Kat11, Kat20, KS12, KS13a, Kel05, Ken76a, Ken75, Ken76d, Ken76c, Ken76b, Ken78b, Ken82, Kil83, Kin86a, Kin86b, Kit78, Kit81, Kje17, Kla84, Kle17, Kno80, Kno88a, Kno07, Knu11, Knu74, Knu83a, Knu83b, Kot79, Krö06, Krö14, Krö15, Krö17, Lan07, Law98, Law05, LY79, Lê21, Lê24, Lef10, Lei79, Lei09, Lem74, Lev97, Lév99]. Review [Lew90, Lew95, Lew98, Lew99, Lew09, Lin74, Lo 84, Lo 85, Lon21, Lor16, Lor18, Lüt08, Lüt09a, Lüt15, Lüt20, Maa07, Mac76, Mah78, Mal97, Mal06a, Mar10a, Mar14, Mar09, Mar95, Mat74, Mat75, Maw06, Maz77, Maz16, Maz21, Maz24, McC90, McG74, McK82, McL12, Mel06, Mic23, Mon75, Mon83, Mon17, Moo75a, Moo77, Moo82a, Moo82b, Moo03, Mor96, Mor19, Mos16, Moy17, Mro16, Mue80, Mue96, Nab04, Nak03, Nee79, Net02, Net24, Neu86, Neu08, Nic13, Nor07, Nor03, Oak12, Oli08, Ols76, Ora82, iÔ75, Pan08, Pan09, Pan18, Par89, Par91, Par92a, Par92b, Par97, Par07, Pec20, Ped77, Ped75b, Pei15, Pei17, Per10, Pet82, PP78, Phi79a, Pin76, Pin80, Pla76, Ple78, Plo06a, Plo06b, Por05, Por75, Pri80, Pye75, Pye76, Pys81, Rab08]. Review [Rec07, Rei15, Rei76, Rem12, Ric98, Ric82, Ric84, Ric91, Rid83a, Rid83b, Rid88, Rid94, Rom97, Rom05, Rom99, Rom04, Ros74, Row83a, Row83b, Row85b, Row85c, Row85d, Row85a, Row90a, Row93, Row06a, Row07, Row08, Row15, Row19, Roz90, Rus75, Sag91, Sai74, Sai94, Sai05, Sal78, Sal87, San03, San05, San06, San09, Sau06a, Sau09, Sau14, Sca94, Sch76a, Sch99a, Sch02, Sch91a, Sch91b, Sch91c, Sch11, Sch13b, Sch16, Sch17, Sch89, Sch07, Sch95, Sch96, Sch90b, Sch05, Sch09, Sch12, Sch13c, Sch19, Sch00b, Scr03, Scr07, Scr08, Seg76, Seg90, Sel74, Sel90, Sel91, Sen84, Sen91, Sen99, Sen06, Sen10, Ser18, Ser90, SG04, SG09, Shu80, Sia23, Sid06a, Sid06b, Sid12, Sid13, Sid24, SS89, SS97a, SS07, SS08]. Review [SS17, Sim96, Smi79b, Spi00a, Spi76, Ste05, Ste09, Ste80, Str78, Str82a, Swe79, Swe00, Syl81, Syl83, Syl03, Syl07, Tan74, Tat75, Tat11, Taz09, Taz15, Tee78, Tee84, Ter89, Thi06, Thi08, Tho76, Tim09, Tob90, Toe96, Too80, Tot79, Tro74, Tro83, Tru75, Tur08, Ull03, Ume76, Ung85, Van05, Van80, Van06, Vek75, Vil96, Vil97, Vit96, Vol81, Vol02, Vol16, Vol77, Wal17, Was20, Wat09, Wil75a, Wil83, Wil75b, Wil77b, Wil08a, Wil05, Wil07, Wil08b, Wis76, Wol91, Wol03, Yan83, Yin11, Yok75, Yon75, Ysh90, Yus10, Zas83b, Zas83a, Zep12, Zer75, Zit07a, vM84, vM07, vdW76a, vdW80, AR76, Ano74-41, Eis78, Høy24a, SS22a, Wil79a, vdW76b]. Reviewed [Ano95q, Ano96c, Ano97c, Ano98f, Ano99a, Ano00a, Ano01a, Ano02d, Høy93, Ano74-54, Ano75o, Ano76-27, Ano77-32, Ano79-28, Ano80o, Ano81-34, Ano82n, Ano83m, Ano84p, Ano85m, Ano86q, Ano87q, Ano88t, Ano89q, Ano90p, Ano91m, Ano92z, Ano93r, Ano94l, Ano03m, Ano04k, Ano05m, Ano06i]. reviewing [SS94]. Reviews [Ano76-40, Ano76-41, Ano95p, Ano96b, Ano97b, Ano98e, Ano99b, Ano00b, Ano01b, Ano02e, BG00, Jon75a, Row00, Sch00a, Thi00, Ano74-53, Ano75n, Ano76z, Ano77-31, Ano78v, Ano79-27, Ano80n, Ano81-33, Ano82m, Ano83l, Ano84o, Ano85l, Ano86p, Ano87p, Ano88s, Ano89p, Ano90o, Ano91l, Ano92y, Ano93q, Ano94k, Ano03n, Ano04l, Ano05n, Ano06j, DR16]. Revised [Thi06, Gra08]. revision [BV05, Øhr85]. Revisited [Wis76, Tho88]. revisiting [Bar11]. Revista [Ano87r]. Revolution [AD02, Kas90b, Wis76]. revolutionary [Ano86d]. Revolutions [Row93]. Revue [Gau13]. Rewriting [Cul07, Gar86b]. Reykjavik [Kat11]. Reymond [RC85]. Reynolds [Big19]. Rhees [Ano81u]. rhetoric [Fek14]. Rhind [Gil78, BS77b, Gil79, Mia08]. Rhode [AH10, Ano74s, Nic13, Sel91]. Rhonda [Wil05]. RI [Haw03, Rec07, San03, Sch07]. Riccardi [Ano87-27]. Riccardo [Net24]. Riccati [Ano89-35, Roe15]. Ricci [Goo77]. Riccioli [Rom05]. Rice [Ano81k]. Ricerca [DG90, GG05b]. Ricerche [Bot78]. Rich [Zas83b]. Richard [AH16, Ano74s, Ano74y, Ano95l, Ber07, Fow98, Haf16, Lê21, Sch14, vdW76a, Ano79x, Dug76, EI12, Fer93, Haf17, SSZ07, Sis02, Wil80, ZvDF22, Ano79m]. Richards [Lew90]. Richardson [Dut84, Eng79]. Richardson-Extrapolation [Eng79]. Richness [Lê24]. Rickey [San03]. Riechardt [Ano89k]. Riehm [Nic13]. Riemann [Ano89k, Ano89m, Ano90g, Boi89, Che84, Fel85, Gra99b, Kno83a, Row15, Wol03, Ano77a, Dar15, EU99, FK90, GO99, Kno83a, LN94, Lau99, Neu88, Por82, Sch82a, Zun83, dSG10, McC90, Spi76]. Riemannian [Gra99b, Blå17a, Cog16, DS19, EU99]. Riemanns [Ull03]. Riemannsche [Ano89k]. Riemannschen [Ull03]. Ries [Ano92e, Høy09a, Høy23]. Rigatelli [Ano85f, Mal06a]. Rigetelli [Ano81r]. right [LY79, TS78]. right-angled [TS78]. Rigor [Bec99]. rigorous [Blå17a, Dho85]. rigueur [DP88]. Rinascimento [Vér01, Ano85f]. Rinehart [Ber74]. Riposte [Sch99a]. Rise [Ale10, Bul24, Gra11, Kle01, Lew98, San09, Cri86, DF10, Fer08, RW03, Sch91c, Tan74]. Risi [Gal12]. Rising [Ano74-40]. Ritter [Ano84f, Gis96, Coo99]. Ritz [Bor14, PR12]. Rivals [Dod13, Ped75b]. riviste [GG05a]. rivoluzione [Net02]. Rizzetti [BFG94, BF23]. Road [Ano74w, BT81, SS18]. Robert [Ano84f, Bun74, D’A10, Goo77, Gra13, Haw77a, Haw77b, Jah12, McK82, Zep12, Chi22, CWW17, Hog77, Rad03, Ver13, Jam13]. Roberta [Net24]. Roberts [Jam13, SG09]. Robin [Abe09, Ayo06, Kat20, Ste05]. Robson [Gra10a, Por05]. Rochester [Ano74w]. Rochot [Mac76]. Rockefeller [Wea77]. Rodolfo [Sid24, Sar97]. Rodrigues [Tur08, AO05]. Rodriguez [Ano93h]. Rodriguez-Consuegra [Ano93h]. Roero [GG05a]. Roger [Gil15, Gra97a, Høy90a, Hud12]. Rohr [vdW80]. Roland [Ano95k]. Role [Fen97, Gil81, Moo02, Pul98, Taz97, Des07, Jon75b, KK90, Mai21, RW03, Tro75, Zyg75]. Rolf [Gra10b]. roll [Gil79]. ROM [Jur07, Rem12]. Roma [Ano82e, Ano90f]. Roman [Ale04, Moo82b, Mro16]. Romandes [Rei15]. Romania [Ano82l, Ano82k, Scr82a]. Romanian [Ano74p]. Romans [Spi00a, Spi00b]. Romantic [Gre83]. romantique [Ano92p]. Romberg [Dut84]. Rome [Rab08, Ano87l, IN83, Lam87, Mar04, Nur92]. Rommevaux [Kno07, Syl07]. Rommeveaux [Fol12]. ROMs [GG05a]. Romy [Sch11]. Ronald [Ano86e, Edw92]. Ronchi [Ped77]. Rond [Fra04a, Kle17, Wil08b]. Ronde [Hei81]. Root [FR98, Ful99, Aub14, Joh11, Zhu16]. Roots [Hud12, Big79, Har11, Hog89, Kat92, Lau92, Tai14]. Rose [Ano77q]. Rosen [Fen08]. Rosenfeld [Gra91b, Haw96, Kin86a, GY77, Vog10]. Roshdi [Ber87, De 20, For10, Giu99, Gra76a, Sai74, Sal87, Sim96]. Roskilde [Ano78y, Buc80]. Rosmorduc [Høy07]. Rossana [Gra99a, Sch12]. rossii [Sel74]. Rota [Tro83]. rotating [Mac77]. rotations [Von93]. Roth [Man06]. roubles [Ano81o]. Rouse [Van14]. route [Gre84]. Routes [Bra12]. Routledge [Bun76a, Gra98a, Gur24, Sid24]. Row [Hym78]. Rowan [Hel83, Nak03]. Rowe [Ano95f, Ewi92]. Roy [Net24]. Royal [AH16, Ano74-46, Fri74a, Sor13]. Rozenfeld [ADG^{+}88]. Rs [Ano88k, Ano88m, Yan83]. Rubbettino [Bee09, Mar14, Sch13c]. Rubettino [Pan09]. rubezhe [Ano81h]. Rüdiger [Høy23, Rom04]. Rudolf [Sch86a, SS89]. Rue [Dur21, Ano89m, Ano89t, Ano91q, Ano91s, Ano92-28, Ano92-27]. Ruestow [Pye76]. Ruhr [Ano86h]. Ruin [Sho86]. Rule [Hay97, Hed99, BH89, Gon08, Gup77c, Kic06, Edw91]. rules [Ber90, Lic97, Pie89]. Runge [DR19, KS12, Tob10, Tob10]. Ruprecht [Ano88n, Sch90b, Ull03]. ruptures [Lor16]. Rush [Ano81u]. Russ [Gra06, Lin74]. Russak [Ano74u]. Russell [Ang79a, Ano75e, Ano87-32, Ano88f, Ano93h, Lew85c, Ano74-56, Ano81z, Ano84c, GG75d, GG78, RT81]. Russia [OS19]. Russian [Aga75, D’A10, Gra97a, Haw96, Kot79, Phi79a, Phi79c, Sti76b]. Russo [Net02]. Ruth [Cer04].

S [AD87, AR76, Ano74m, Ano74v, Ano74-34, Ano76i, Ano76j, Ano77s, Ano87k, Ano88k, Ano88m, Ano89f, Ano89l, Ano92l, Ano95f, BLW77, Bir77, Cap13, Cas80, CK92, Dav76, Eis82, Fel85, FR82, Gam96, GG86, Hes90, Jon77b, Ken75, Kin86b, Lin74, Mat74, Meh76, Pin80, Pys81, Ric98, Row90a, Row23, Scr03, Taz09, Tee78, Van05, Vol81, vD10, Ano77l, Ano92-32, Dje97, Haf17, Ken76d, Tob90]. S. [Fel85]. S.S.R [Dem74, Ken75]. Sūri [Tee84]. sa [Mal97]. Saṅgita [SK24]. Saṅgita-ratnakara [SK24]. Sabetai [Sai05, Ste80, Fri24]. Sabine [Fol12, Kno07, Syl07]. Saccheri [Gal12]. Sachen [KS12, Tob10]. Sachs [Seg76, Ber84b]. Sächsische [Ano93x]. Sacrobosco [Ano95n]. Sadi [Hei81]. Sadleirian [Cri99]. Saecula [Fol87]. Saharan [Ger94]. Sahl [Ai04, Bre07, Sim96]. Saidan [Mat74, Pin80, Ano92i, Ros92, AR76]. sailing [The84]. Saint [Ano92-27, Row15]. Saint-Bernard [Ano92-27]. Saint-Gervais [Row15]. Saito [Sid12]. SaKHYa [Knu11]. Sal [Ano95k]. Salah [Sai74]. salesman [vBS20]. Salim [Ano92i, Ros92]. Salle [Ano90i, Ano90-34, Bre91]. Salomon [Gil89]. Salvatore [BF89]. Samarkand [Bag97b]. Samaw’al [Sai74, AR76]. same [Lê20]. Sammlung [Kno78a]. Samos [Sid24]. Samplonius [Hog15, Van06]. Samsó [Hog24, Mor96]. samt [RFS89]. Sämtliche [Ano80b, Bar78b, Bla05, Gar05]. Samuel [Jon77c, Pla76, Ano91-30]. samyoga [SK24]. samyoga-meru [SK24]. San [Ano74-51, Ano76h, Ano89r, Ano91p, Ano92f, Bun74, dMC86a, DN12, Ewi92, Fol14, Lew90, McK82]. sanctity [Fin11]. Sancto [Van84]. Sandbjerg [Gra87]. Sandifer [Cal08, GG08, Jah12]. Sang [Cra03]. sangaku [Won22, CM22]. Sannia [Taz09]. sans [Gis99a]. Sanskrit [Ano88k, Plo96, Sma16, Wag15, Vol77]. Santcliment [Lab01]. São [Ano87-31]. Sapienza [Nur92]. Saragossa [Hog86, Hog95]. Sarah [Duc16, Wol03]. Sarasa [Bur01]. Sarasvati [Yan83]. Sardinia [Sco16]. Sarma [Ano88k]. Sartaq [Dje97]. Sartorius [GG13]. Sarvasiddhantaraja [Gur24]. satire [SSZ07]. Satz [Scr81]. Sätze [Sch88]. sau [Ano74p]. Saunderson [Tat92]. savage [Tuo23, Pat75]. savant [Tat75]. savants [Ano92p, Ano92-50, Ano92-28, Ano89i]. Savile [Zep12, Gou10]. savoirs [Fol14]. Saxony [Kre03]. Scaife [Nak03]. Scales [Sid12, PGN18]. Scaliger [Hog10]. Scandinavian [Lüt08, SS18]. scelti [Mal06a]. scènes [Gis99b]. Schafer [Ano89l]. Scharlau [SS89]. Scheel [Haf16]. Scheltema [GG77]. scheme [HZ85]. Schicksal [Aud12, Krö14]. Schinzel [Moo82b]. schlechte [Ine91]. Schmandt [Ano93f]. Schmandt-Besserat [Ano93f]. Schmidt [Bos77, McG74, Wil75a, BB97]. Schneider [Bor18, Cif00]. Schocken [Ano85h]. Scholar [Che17, Dum15, Hog10]. Scholarly [Neu86]. Scholars [Gup91]. Scholz [Che84, GS91, Lê24, Nab04, Plo06b, Toe96, Fel85]. School [Ano78-29, Ano79g, Ano81h, Ano88-29, D’A10, GG90c, Mro16, Abh75, Ano88a, Ano91-31, Ano93a, CV22, Gup75a, Kar22, Lew04a, Nor75, Par04, Phi78]. schools [Ano83z, BC04, Mar04, Eve76]. Schouten [Cog16, CM18]. SCHPM [Ano75i, Ano75j, Ano75k, Ano76m, Ano76n]. Schramm [Kno93a, Thi05]. Schrift [Rom99]. Schriften [Ano80b, Bar78b, Bla05, Bul22, Cri17, Gar05, HB^{+}13, Høy09a, Ano92l]. Schriftenreihe [Vol81]. schriftenverzeichnis [RFS89]. Schröder [Ano84e, Gam96]. Schubring [Dho11, Kat11, Lei09, Tob90]. Schule [Dil81, Sch17]. Schulen [Fol87]. Schultz [Sch82b]. Schultze [Lüt08, Sen10]. Schumacher [Vol81, Vol81]. Schuppener [Lév99]. Schur [Haw03]. Schwartz [BPL17]. Schwarz [Sch11]. Schwarzbach [Cog12, Gra08, Taz15]. Schweber [Bel08]. Schwermer [Gra18]. Sci [Goo77, Haw77a, Haw77b]. Science [Ang88, Ano74-47, Ano74-51, Ano74-67, Ano76w, Ano76-28, Ano77r, Ano78x, Ano79p, Ano80-39, Ano81p, Ano82l, Ano82k, Ano83x, Ano85-29, Ano85-30, Ano86m, Ano87n, Ano87-31, Ano88k, Ano88m, Ano88-31, Ano89-33, Ano90-37, Ano91d, Ano92-29, Arb86, Atz06, Ber04, BMY07b, BLF92, Clu76, Cri24a, DR19, Eva94, Gis96, Gup85, Hod75, Hog03, Jon81, Lew98, Lüt08, May75c, May77c, McL12, Mol81, Nos12, Ott88, Par92a, PH96b, PP78, dBR76, San03, San05, Sau06a, Sch13c, Scr82b, Scr87, Scr88a, Sel90, She96, Tee78, Tee84, Wat09, Wis76, Zep12, Amb08, Ano74m, Ano74-74, Ano74-75, Ano74-76, Ano76-34, Ano78q, Ano78-28, Ano80q, Ano80-31, Ano82-27, Ano84i, Ano84m, Ano85-31, Ano86g, Ano86j, Ano86f, Ano87i, Ano87-35, Ano87-36, Ano90v, Ano90w, Ano91h, Ano93s, Ano94h]. science [BMY07a, Bow87, Buc80, BT84, Dij04, Dru86, DGD04, Gil89, Hir74, Hol15, HNE21, KS13b, Lew76, Lüt09a, Øhr85, Pul05, Ric80, Scr84, Shi94, Smi79a, Tan74, Ver13, vD09, vM84, Ano78w, Ano89-33, Ano94x, Mai08, Hog03, Lew98, Ser18, Thi06, Thi08, Ano84d, Ano92n, Hod77, Taz15, Ano88k, Pin76, vdW76b]. Science-Nauka [PP78]. Sciences [Ale14, Ano74f, Ano75p, Ano80-32, Ano87j, Ano87-36, Ano89n, Ano89-33, Ano90t, Ano91f, Ano92k, Ano92p, Bre07, Cal83, Das81, Eva94, Fol14, Fol15, GG76a, Gup79, Hog85b, Høy09b, Jah12, Kat97, Ken75, Kni01, Kno88a, Lüt08, Mah78, Mal97, Mar90b, MY99, Moy17, Ste80, Swe79, Ano77j, Ano78-34, Ano82c, Ano85p, Ano85-31, Ano88v, Ano88w, Ano91r, Ano92-32, Ano93p, Aud11, Bun76b, dMC86a, Dho88b, DR03, Enr76, FFW00, Gau13, GG76c, Kin86b, Krö15, Maz77, Nor75, Oak12, Sco16, SS12b, Tat86, Tur81, Vol77, Ehr11, GG98, Gra98b, Man89, Ste05, Ano78h, Ano84j, Pet82, Lor16, Sar08]. Scientia [DB23]. scientiæ [Krö06]. scientiam [Bar77, Rid83b]. Scientiarum [Rid83a, Rid83b, Sel90, Ano77g, Jay83]. Scientific [Ano76-28, Ano78l, Ano79r, Ano79q, Ano86k, Ano88r, Ano89-33, Ano90f, Ano94w, Bag97b, Gup88, Hun07, Mar95, Moo82a, Nab04, Pet82, Wal17, AH20, Ano74x, Ano84-27, Ano87l, Ano88l, Ano92-38, Cas80, CB21, Fre89, Gab79, Hod75, NT05, Pee95, Por04, Pys81, Roe15, SD94, Str85, TC05, Tat75, SS94, SEM04, Sti78, Ano74z]. scientifica [Ano93j, GG05b]. scientifico [Net02, Rom05]. Scientifique [Bre07, GG76a, Hei81, Ano92-50, Ano92-28, Dho88b]. scientifiques [Cho75, Fre77]. Scientist [Ano86d, Dyc75b, Ano81-30, Gra78c, Gri82]. Scientists [Gup91, BJ89b]. Scienza [Ano90c, Mai08, Ano88r, Net02, Rom97, Sch13c]. Scienze [Ano82e, Ano90f, Fio17, Amb08, Oak12, Ser90]. Scotland [San06, Cra98, Cra00]. Scottish [Kla84]. Scriba [Ano95f, Bee14, FK89, Maa07, Mos16]. scribe [Gil81]. scribes [Coo11]. Scribner [Hod77]. Scribners [Bir77]. Scritti [Mal06a]. Scroll [DS99c]. Scultetus [Ano90j]. Scuola [Ano85g, Ano88r, Sai94, Bot77]. SCV2 [Cal75]. Se [Mar95]. sea [SS86]. Sean [Ano74-28]. Search [Don98, CM18, Ehr15, Lew95]. Sebastian [Ano91p, Ano89r, dMC86a]. sec [Hug05]. Secant [Plo96]. sechster [Cri17]. Second [Ano83x, Ano92-46, Fol94, Moo77, Shi93, ZD88, HZ85, Sch77, Won22, Ano87-31, Ano90-33, Ano92-52, Bin91, DV80, Shi86, Shi88]. second-order [HZ85]. Secondary [Mac78a]. seconde [Pan88]. Secret [McK82, Gin83]. Section [Ano79p, Ano89-33, Ano92h, Hog24, Roy17, Tob81, Ano76-34, Bec19, HF19, Shi86, RT02]. Sections [Ano94w, Ano80e, Ano80-34, Mai09]. Sector [Ber04]. Sector-Figure [Ber04]. secunda [Rid83a, Rid83b, Sch00b]. see [SY90]. Seed [Wil08a, BK06]. seems [Ano82d]. Segre [Gia01, Taz09, Cer17]. segregation [Pie83]. Seicento [Mai09, Mar14]. Seidel [Kno88a]. Seillac [Ano90-35]. seinem [Ano95f, Scr81]. seinen [Scr03]. seiner [GG07, Mon83, TC05, Ull03]. seines [Gab79, Sch19]. Seiten [Fol87, SS89]. Seki [Ano78-32, Ebe96, Ano78-32]. Selected [Dem86, Gri76, Knu74, Moo03, Par02, Rec07, Ser18, HB^{+}13, Réd05, Thi07, Cib76]. selection [AR07, Fen08]. Selections [Wil83]. Selective [Ber97, Dau85b, Rid88]. Selenius [Scr92]. Sella [Ken82]. selon [Gol82, dMC92]. selten [Ine91]. Semevsky [BN22]. semi [Cog15, ET14]. semi-simples [Cog15, ET14]. semigroups [Hof92]. Séminaire [Gre85, Ano80d, Ano90-34, Mar90b]. Seminar [Ano75p, Ano79-45, Ano83n, Ano83y, Ano87-34, Ano88r, Ano90u, Arb86, Fre89, Gup88, Gup91, Nab04, Sar88, Sch90a, Ano81-32, Ano85n, Ano94g, Coo88, Ano82-27, Gup75b]. Seminario [Ano82-27]. semiotics [Ano74-52]. Sen [Ano74m, Ano88m]. Seneta [Das81]. Senia [Mar90b]. sept [Pic89]. September [Ano76x, Ano76-28, Ano82k, Ano83f, Ano85-28, Ano87c, Ano87m, Ano89m, Ano89r, Ano90g, Ano90-32, Ano90-36, Ano91c, Ano91o, Ano91p, Ano93k, Ano93m, BGG06a, BDD^{+}88, Bre91, BLF92, Cri82, GG87b, GG90d, Maa07, Mil93, Nag91c, Neu15, Roe89, Scr82a, Ano24i, Erl97, GG88]. Septembre [Ano90i, Bie78, Obe78, Boi89, Bor89]. sequence [Scr83, Tuo23]. sequences [Ger74, SS10c]. Ser [Ano92l, Par97]. Serfati [Mar10a]. Sergio [Rab08]. Série [Fra04a, Wil08b, GG76a, Kle17]. Series [Ano77m, Ano88i, Ano89h, Ano91g, Bar74, Bar77, Fer98b, Kit78, McK82, Mor96, Par89, Pei22, Plo01, Rid83a, Rid83b, RT02, San09, Sen91, Ter89, Vol77, Ano74-58, Ano78q, Ano83c, Ano90v, Ano92k, BL76, Dal91, FP03, Fer07, Fer08, GG93, Gup74, LN94, Nik09, SS93, Zyg75]. Sernesi [Par97]. Serra [Ano92s, Net24]. service [Ano95n, Big19, KS13b, Taz15]. Sesiano [Fol15, Hog85b, Rei15]. session [Ano76x, Ano87-32, Dru85]. sessions [Cas80]. s’est [Mal13]. Set [Moo02, Ang79b, Ano81z, Ano89c, Dau75, Enr85a, Gan85, McG74, Moo76, Scr81, dV22]. Sets [Ano81h, Bel07a, SS10c, BS06, Gar85, GG80, HP05, Moo08, Neu88, Sch07, Ano77o]. Settlement [Bie74b]. Seuil [Gre85, Maz24]. Seventeenth [Ano74z, Atz06, Pye76, Car07, Gol92, Sep05, War08, Dij04, Mai09]. seventeenth-century [Car07, Sep05]. seventh [Zhu16]. seventieth [BGM^{+}76]. several [Gre84]. Sexagenarian [Ano90r]. sexagesimal [MW17, Pro22]. Sexton [Kid09]. sezione [Mar14]. sezioni [Mai09]. sFr [Gra97a]. shadows [Tee78]. Shadowy [BS05]. Shaker [Bos03]. Shaker-Verlag [Bos03]. Shallit [Sch95]. Shao [Yon75]. Shape [Lew98]. Shaped [Van05, Sma16]. Sharaf [Hog89]. Shea [Wis76]. Sheard [Wil77b]. Sheet [EU99]. Shell [All09, Kat07]. Shell-Gellasch [All09, Kat07]. Shenitzer [Gra91b, Wol03]. Sheppard [MB20]. Sherlock [AKW04, Rob01, Ayo06]. Shifting [Mol76]. Shigeru [Hir74]. Shimodaira [Ano78-32]. Shing [Dum18]. Shing-Tung [Dum18]. Shiran [Ano89d]. Shiva [Hal84]. Shkola [Rei76, Ano81h]. shocked [Car24]. Short [Dav76, Ene18, Gil77]. short-lived [Ene18]. Should [Ano74-74, May75d]. Shoulders [Ano95l]. showing [Wat86]. Shrine [Won22, FC22]. Shtiinstsa [Dav76]. shu [Zho21, Cul07, Yon75, Shu92, Yon74, Cul07, Vol10]. Shu-shu [Yon75]. Shuchun [Shu92, Vol10]. Shulman [Lei79]. Shuntaro [Syl83]. Shuppan [Sid12]. si [And75]. Śiṡyadhīvṙddhida [Tee84]. Sibirskii [Dav76]. Sicily [Taz18]. Siddhantasundara [Mon17]. side [SS93]. sides [Hog24]. Sidoli [Moy17]. Siebert [Sch11]. Siècle [Ano92p, Ano89m, Ano89-30, Ano90g, BH14, Boi89, Cha90, DP88, DG90, Dug76, For10, Gra08, Hog01a, Mar10a, Mar93b, Pan88, Pei17, SD08, Sim96, Giu99, Gol82, Kit81, Kno07]. Siècle. [Gra01]. Siècles [Bre07, Ai04, Ano79l, Ano92-50, Ano92-28, Ano92-27, Dje97, Lai10, d’E01].

Sieg [Daw16b]. Siegmund [Bla05, Lüt08, Sen10]. Siegmund-Schultze [Lüt08, Sen10]. Siena [Ano85f, Hug05, Vag82]. Sierpiński [Moo82b]. Siglo [Ric98]. sign [Kna88]. signals [Ano89-35]. Significance [Bos02, vdW76a, GG75c, Høy83, Pie83, Sai95, Ano77r]. significant [IN84]. Sijzi [DS99b]. Sikidy [Asc97]. Sikorski [Moo82b]. Silver [Gup77b]. Silvia [GG05a]. Silvio [Ano81f, Hod77, Ken78b]. Siméon [KS13b, Ano87i, Taz15]. Siméon-Denis [KS13b, Ano87i, Taz15]. similar [dV22]. Simmons [Boy74]. Simon [Gil15, Ste09, Tur08, DV08, Hah11]. simple [Mae89]. simples [Cog15, ET14]. Simpliciu [Rom04]. simplified [Bra89]. Simplifying [OA07]. Simposio [Lam87]. Simpson [Bla14, Sho85]. Sinaceur [Cor96]. Sinan [Hog01a]. since [Ano94h, Ott89]. Sine [Hay97, Plo01, Gup74]. Sine-Differences [Hay97]. sines [FLT16]. Singapore [Mar95, Nee79]. Singular [Ano75q, Sac20]. singularities [Row21]. Sinica [Ano94g]. Sion [Bor14]. Sir [Nak03, vM07, Edw92, GG75b, MS05, Hel83]. Sirasi [Maz77]. Sirazhdinov [Ken75]. sirven [dL01]. Sisana [Net24]. sistema [Ken78b]. Sisters [Ber97]. Sivin [Hir74]. six [Cav85, NS94]. six-square [NS94]. Sixteenth [Bo24, Rey09, Tra21, Abe91, Tra10]. Sixteenth-Century [Bo24, Rey09, Tra21]. Sixth [Høy83, Ano10k]. Sixth-century [Høy83]. sixty [Kno93a]. sixty-fifth [Kno93a]. Siyuan [Ple78]. size [Ano89c]. Sizes [Sid24]. Sketch [EU99, Dav76, Pla76]. Sky [Sho02]. Slave [FG90]. Slaves [Cra98]. Sloan [Ano05a, Ano05b, Ano05c, Ano05d, Ano06b, Ano06c, Ano06d]. Sluse [Hal86]. Slutsky [She93]. Smale [Aub02]. Small [dV22, Fis78, Lau87]. smert [Dyc75a]. Smilka [Coo96]. Smirnov [Ano77m]. Smith [Jam13, Jon85, Row83b, Don98]. Smithsonian [Kin74b]. Smyrna [McK82]. Smythies [Ano81u]. Snedecor [Eis75]. Snellius [dW10]. SNES [Høy07]. Snezana [How18]. Snienton [Wil77b]. so-called [Sin85a]. sobre [dL01]. Social [Ano79-47, Ano95k, Mac78b, Meh80, Mel09, San06, Tur08, AO05, Ano76-29, Ano82c, Ano88l, Aub14, Cro77, Nor75, SS22b, Tur81, dS90]. socialism [Alb94]. socialist [Ano74-60, Seg86, Ano81q]. Sociedad [Ano90-32, Ano84y]. Societas [Sel90]. Societatis [Rid83a, Rid83b]. Société [Ber87, Mal97, Sal87, dMC86a, Ano75f, Ano92k, Dho88b, Gra76a, Hou89, Ano74-32, Ano74-33, Ano78k, Ano80-32, Cha81, Cha87, Jon77a]. societies [Jay24a, Jay24b, Van22]. Society [Abe91, AH10, AH16, Ano74-32, Ano74-47, Ano74-51, Ano76f, Ano76w, Ano77t, Ano78k, Ano78t, Ano78y, Ano80c, Ano81v, Ano83w, Ano83x, Ano84x, Ano85d, Ano87-32, Ano92f, Bec01, Cha81, Coo96, Cri82, D’A10, Dho83, Dru85, Dru07, Dru12, Enr83, Fen08, GS77, Gra78b, Gra18, Gup78, Gup81, Haw96, Haw03, Jon77a, Law05, Lo 84, Lor18, Mro16, Neu08, Nic13, Pet82, Ram00a, Rec07, RWG95, RW98, Row85c, San03, Sch07, Sel91, Sor13, Swe00, Syl81, Tat11, Tob81, Tur08, Vol77, Law05, Ano74-33, Ano75f, Ano76-34, Ano78-34, Ano85i, Ber85, BM77, Cha87, GG87a, Gra76c, Aub14, Jon81, Nos12, RWG95, RW98, RW03]. Society/London [D’A10, Fen08]. Sociology [Gra98a, Mac78b]. Sofia [Ano86d, Ano95m]. Sofja [Coo95]. soft [Bor18, Haw03]. soft-cover [Bor18, Haw03]. Softcover [Dav19, Ser18]. Sofya [KS13a]. Sohn [SS97a].

Soifer [Wat09]. Solar [Gol80, Moz13]. solid [lLW13]. solide [Mah78]. Solomon [GG90b, GG05c, Pec20]. Solution [LMP01, Bra89, Chr07, Erl92, Fri81b, FC22, Kic15, LY80, Mac77, Sac20, Sho86]. Solutions [Ung23, Sch77]. solve [Jor81]. Solving [Atz04, LYK89, CO13, DS99b]. Some [Ane88, Ano91g, Bak75, Bel02, BS77b, Chr07, Cor16, Dad84, Fer98b, Gil78, GG75d, Jah93, Lei77, Mac75, Mik75, She93, Taz01, Thi97, Tho11, Wil79a, You86, Zun83, BM77, Dal85, Gar85, Nau03, Fri91, dMC92]. Somerville [Ste20]. Sommerfeld [Eck13, Jur07, Sau14, Eck13, Sch92]. Son [Mal97, Ano87i, Fio01, Hal86, YDD83, dL01]. Sonderforschungsbereich [Ano86h]. Soner [Tho76]. Sons [Gri82, Hod77, Kin86b]. Sophia [Coo86]. Sophie [DD87, Del05, LP10, Ano84i]. Sophus [Fri91, Kot79, Row90a, Fri91, Haw77a, Haw77b]. Sorcerer [Sid12]. Sørensen [Lüt08]. Sorrenson [AH16]. Sossinsky [Row90a]. Souchay [Wil08b]. Souleymane [Ric91]. Source [Fer04b, Moo03, Ano77f, Ano84d, Ano88k, Moo77, Xu05, vH67]. Sourcebook [Kat09, Che12, Ano89l]. Sources [Ano75z, Ano76-42, Ano76-43, Ano77-41, Ano78-33, Con06, Fol15, Hog85b, Høy09b, Jah12, Ram00b, Ser18, Che09, GG75d, Pes03, Rey09, Sma16, Spi00a, Spi00b, vdW75]. Sous [BKM^{+}82, Bun76b, Dur21, Hou89]. Southampton [Ano81q, Ste09]. Southern [Ano82d, Cas77]. Soveria [Bee09, Mar14, Pan09, Sch13c]. Soviet [Ano76-34, Ano79p, DV80, Ane87, Hol15, Sen04, Vuc99, Vuc00, Vuc02]. Sozial [Ano88n, Sch90b]. Sozial- [Ano88n, Sch90b]. SP [Lei09]. Spa [Gra02]. Space [Ano93i, Cox75, Dor95, DS19, Gra80a, Sau06b, Sch79, Gra91b, Ric82]. space-time [Cox75]. spaces [BS07]. Spain [Ano89r, Ano91p, Ano92-44, Mil93, NLL08, Mor96]. spake [CH90]. Spanish [AH20, Ano78-28, Ano78-34, Ano82-27, Ano83x, AH99, VJRZ87]. spanners [BS05]. spead [Rei86]. speak [Kel05]. speaking [CM19a, CM19b, Nos12]. Special [Ano81e, Por19, Krö06, Ano87-32, Luc18]. spectral [Haw75a]. speculo [Ros76]. Speiser [Hes90]. Spektrum [Pei15, Sch17]. Spence [Cra13]. Sphaerica [Yus95, Yus95]. Sphere [MF10, Ano95n, Ber77, Fu91, Mal13]. Spheres [Gur24]. Spherical [Atz04, Fra17b, Che10, Hai86]. sphérique [For10, Auj76]. sphéropée [Auj76]. Spiral [Wei98, Erl92, Wil94, EG81]. spirals [Kno78b]. spirit [Hog91]. Spiritalia [Fol87]. spoiled [Hee14]. Sprach [Lév99]. Sprach- [Lév99]. Sprachbereich [Ano91u]. spread [Sco16]. Spreadsheets [Por05]. Sprekels [Sch19]. Springer [Ale14, Ano77i, Ano78j, Ano84j, Ano87j, Ano87g, Ano87f, Ano88j, Ano89e, Ano89f, Ano91f, Ano93g, Bel12, Bor18, Bra07, Cal83, Cal92, Che14b, Cog12, Das81, Daw16b, Des09, Dru91, DR19, Edw77a, Edw77b, Fol15, Fre92, Gra94a, Gra91b, Gra09b, Hog85b, Høy09b, Jah12, KS13a, Knu74, Krö14, Lê24, McL12, Mel06, Moy17, Pei15, Plo06b, Row85a, Row07, Row19, San09, Sch17, Sch12, Sch19, Scr07, Ser18, Sid13, SS17, Swe79, Wat09, Wol91, Zep12]. Springer-Verlag [Ale14, Ano77i, Ano78j, Ano84j, Ano87j, Ano87g, Ano87f, Ano88j, Ano89e, Ano89f, Ano91f, Ano93g, Bor18, Bra07, Cal83, Cal92, Das81, Dru91, Edw77a, Edw77b, Fre92, Gra94a, Gra91b, Hog85b, Knu74, Mel06, Plo06b, Row07, Scr07, Ser18, Swe79, Wol91]. Springer/Birkhäuser [DR19, SS17]. Springfield [Ano74-30]. Spriptor [Ken76c]. Square [FR98, Wei98, Fow92b, Fri81b, Hay87, NS94, Wei92, Zhu16, dV22]. squared [Str23]. squares [Ger74, Gil77, Pie93, Zho21]. Squaring [Gui01, Lüt20]. ’Srîpati [Sin85b]. srednem [Ken75]. srednevekovom [Ken75]. srednevekovya [Kin86a]. Sri [Jay24b, Jay24a]. Srishti [Plo06b]. SSR [Dav76]. St [BT81, D’A10, Kat92]. St. [Cal96]. Staatsbibliothek [Jay83]. Stability [Ano77p]. Stadt [Dho88a, Fol09a]. Stalin [Vuc99, Vuc00, Vuc02]. Stamatis [CK92]. Stand [Kat11]. standard [AR24]. Stanhope [Bel07b]. Stanislaw [Moo82a, Moo82b, BM96]. Stanley [Per10]. Stanton [Ano81k]. Stark [Moo82b]. State [Ano74-40, Ano95k, Ano95l, Eis75, LY79, Lev97, Rei76, Zas83a]. stated [Ano81k, GG90c, GG90b, GG10]. States [ABC^{+}88, Arn88a, Cas77, San03, Shu80, SS97b]. Statik [Has04]. Station [San03]. Statistical [Ano76x, Buc77, Das81, Hun07, Sen06, Dav75, Høy83, Por04, Sch91c, SSZ07, Yca16, Eis75]. Statistics [Bar74, Bel08, Hun96, Ano74g, Ano74h, Ano76x, Cas77, Cas80, Ine94, Ney78, Nor75, Sch91b, Sch06b, Sen91, Spr78, Tan79, Tee79, Ano88l, Sen06]. Statistik [Ine94]. status [Car07]. statut [Mal97]. Staudt [Mar06]. Stedall [Bee06, Gra10a, Neu15, Tat11, vM07]. Steepest [PS97]. Stefan [Ale04, JMT19, Kar22]. Stegmaier [Scr07]. Stein [Par89]. Steiner [Ano81g, Ano89j, Bar82a, Bar82b, Ber07, Bur08, GG07, Gra90, Høy24b, KS12, Lef10, Mat75, Pei22, Pet82, Row08, Sch11, Van05, Van06]. Steiner-Verlag [Bur08]. Steklov [Lin74]. Stella [Hal84]. Stephen [Buc77, Sch91b, Sen06, Syl81, Aub02]. Stephenson [Ano91f]. Stern [Asp83, Gra10b]. Steve [Bir83, Dum18, Gra06, Zit07a]. Stevin [vM84, DV08, Ste09]. Stigler [Sch91b, Sen06]. Stillman [Ano78h]. Stillwell [Wol91]. stimulated [Ano86l, Bin90]. Stimulating [Ano86l, Bin90]. Stochastics [Ine88, Ine91, Mat14]. Stochastik [Ine88, Ine91]. Stockholm [Tho76]. Stockholms [Tur20]. Stockwell [Kil83]. stoicheiôsis [Kno93b]. stoletiya [Ano77l, Ken76d, Ken76d]. Stones [Mel02]. Storia [Fio17, Ano81f, Ano91k, FR83, GG05b, Roe89, Ano81r]. Storico [Ven02, DF10, Mai08, Rem12, Sch13c, Ken78b, Ken82]. storico-didattico [Mai08, Sch13c]. storiografia [Ano87-27]. Story [Abe06, Dau12, Edw04, Nik09]. Stott [PB08]. Strachey [Sma16]. Straddling [Zit07b]. strait [Hog24]. Strange [Dry14, Duc16, RL16]. Strassburg [Str81]. Strategical [Haf17]. Stratification [Rob96]. Straub [Hes90]. street [BGSS16]. Streitschriften [Ano92j]. stress [Wil74]. string [Van14]. Strong [Hog81a, Sen92]. Stroyls [Kin86b]. structural [Roy17]. structure [Bré12, Fer20, Ung85]. Structures [Ano89r, Kle01, Ano91p]. struggle [VHK16]. struggles [Zit07b]. Struik [Ken82, Lem74, vM84, Alb94, Ano76-44, Bos02, Row94]. Stubblefield [Zas83b]. Student [Eve76, RWG95, Wal94]. students [Ehr18, Scr75b]. Studi [Ano82e, Ano83c, Ano90c, GG05b, Hug05, Ano81w]. Studia [Bar82a, Bar82b, Pet82, Ste80]. studie [Ano89f]. Studien [Ano88n, Ano88i, Høy90b, Mue80, Sch90b, Tob90, Bar82a, Bar82b]. Studienreform [Tob90]. Studies [Ang88, Ano74s, Ano77r, Ano81k, Ano84g, Ano84h, Ano84j, Ano84i, Ano87j, Ano89h, Ano91d, Ano91f, Ano95j, Ano95k, Atz06, Cal83, Das81, Eva98, Fol15, Gra77, GG90c, Hog24, Høy09b, Jah12, Lev97, MY99, Maz77, Mor96, Moy17, Par92a, Pys81, Sau06a, Sch89, Ser18, Swe79, Zep12, vM84, Ket82, KB74, Gup75b, Kin86b, Pys81, Sch89]. studio [Ano91k, Ser90]. studium [Maz77]. Study [Ano79g, Ano79-30, Ano80p, Ano80-33, Ano86n, Ano92v, Coc98, Fre89, Gra98a, IM98, McL12, Rob96, vdW76a, Ano87j, Ano91k, Asc88, Boo86, Cam77, Dal85, Ene18, Hee14, Imh03a, Jay24a, Jay24b, Lan75, LD22, Nee79, Ray12, Ric80, Ton82, Tuo23, Bat06, Zit07a, Ano78g]. Stuloff [KS08]. stupende [Coz12, Tra10]. Stuttgart [Ano79j, Ano88i, Ber07, BMW08, Bur08, Dho88a, Edw80, GG07, Gra90, Has04, Hes90, Høy24b, KS12, Kit78, Lef10, McC90, Ora82, Row83a, Row08, Sch11, Sch09, Syl83, Van05, Van06, BMW08]. Stuttgart-Bad [Ano88i, Kit78]. Stuttgarter [BMW08, Sch09]. Stuutgart [Ano89j]. Styazhkin [Aga75]. Style [Ott88]. styles [Bre19]. Suan [Nee79, Shu92, Yon74, Vol10, Cul07]. Suanshu [Zho21, Yin11]. Sub [Ger94]. Sub-Saharan [Ger94]. Subbarayappa [Ano74m]. Subject [Ano87-33, Ano81l, Ano95k, Høy95, Ric84, Row85a, Sti78, VHK16]. subjective [Das94, Ver15]. subscription [Wal75]. Subseries [Ale14]. Substitutions [Bir96, Jaë19]. subtext [FU01, Sai05]. subtraction [Kan88]. subtractive [Sam24]. Sud [Ano90-34]. Sufficient [Fra09]. Sugaku [Sid12, Ano78-29]. suggested [Sch91b]. suggestion [Ano76o]. Sugino’o [Won22]. sui [Bot78]. suivi [Fre77]. sukunnûm} [Fie22]. sukunnûm}-number [Fie22]. sul [BFG94]. śulbasūtras [Ano88m]. sulhae [Yin11]. Sum [Dum18]. Sumer [Por05]. Sumerian [Høy82]. Summa [Lab01, Rey06]. summable [Nik09]. Summary [Bir96, Ano74-41]. Summation [Fer98b]. Summer [Ano83k, Ano90u, DV80, DK86, Ano93a, Gup75a, Pyc84a, Ano88-29, Ano91-31, Dem74]. sums [Blå17a]. Sun [Sid24, vM84, Ano91-30]. Sundman [BG10]. Sunschrift [vM84]. sunt [Lo 84]. suo [Ser90]. supérieure [Gre85]. Superiore [Bot77]. superposition [Axw18, Wag83]. Supplement [Ano84z, Sch95, Ano85r]. Supplementa [Ano87d, Pet82, Vit96, Bar82a, Bar82b]. Supplementary [Kin79, Scr75a]. Supplementband [Bul22]. Supplemento [Par97]. support [Fen03, Wea77]. supported [Nos12]. surd [Cer17]. Surface [Hog02a, Phi79b, CR22]. Surfaces [Mel06, DD87, Høy02a, Row15, dSG10]. Surjectivity [Wüs02]. Surrey [Ano74-46]. Survey [AÖ01, De 09, Her75a, Ric96, Sid13, Ano77g, Ber84a, Mar90c, PJ11]. Surveys [Moy17, Hol15]. surviving [Del06]. Susan [Eis75]. Susanne [Lüt15]. Susceptible [Aub14]. Suzuki [SG04]. Swansea [Jam13]. Swetz [Ano81q, LY79, Yok75]. SwF [Ter89]. Switzerland [PP91]. Sydney [Ano74-73, Gra13]. Sylla [Cri24a]. Sylow [Sch88, Bir96, Sch88]. Sylvester [Nor03, Hog91]. Sylvia [Cav95]. Symbolic [Mas08]. Symbolical [AB02, Pyc81, Pyc82]. Symbolism [Ste07a]. Symbols [Pyc97, Sch05]. Symmetrie [Mai87, GS91]. Symmetry [Gra09a, Row90a]. Symposion [Ano92-48, Nag88, Mai87]. Symposium [Ano77d, Ano77t, Ano77-29, Ano79-31, Ano79-46, Ano81p, Ano85-29, Ano85-31, Ano86l, Ano86r, Ano87-36, Ano88-31, Ano90m, Ano90l, Ano90a, Ano91o, Ano92-31, Ano92-30, Ano92-49, Ano93-34, Ano94a, Bin89, Bin91, BLF92, Cas77, Gup79, Gup86, Lew99, Mic88, Pet82, dBR76, Shi93, Vog94, Ano74g, Ano74h, Ano76e, Ano79-29, Ano82j, Ano84-27, Ano88u, Ano89o, Ano89r, Ano90n, Ano91p, Ano92-33, Bin90, Dun89, Mil93, Shi94, Ano90-35, Gra87, Ott88, Row89, Scr84]. Syntaxis [Gol76]. Synthese [Ano77r, Bow84]. synthetischen [Dho11]. Syo [iÔ75]. Syracusa [Ano90c]. Syracusa-Catania [Ano90c]. Syria [Sai74, Ano78x]. Syrian [Sai74]. system [Ano74l]. systematische [Høy90b]. Systematizations [Bul09b]. systèmes [Ple78]. Systems [AD02, Ano89-35, Ano81m]. Syuzan [iÔ75]. syzygy [CG92, Kre03]. Szabó [Bow84, Ora82].

T [Ano74p, Ano77p, Ano88j, Eis75, Gra77, Ken76a, LY79, Lew95, Ste09, Ter89, Yan83, Pet82]. T. [Glu05, Meh76, Tat88]. Tūsī [Hog89]. Table [Ano04u, Ano04v, Ano22l, Hay97, Hei81, Sen06, BS77b, CG92, Gil78, Jon77b]. Tables [Por05, Ros81, Van98, Wil90, Ano74q, Ano88h, Ano93f, Bel15, Bul09a, Cra03, Ham80, Kin74a, Kin79, MB20, Sid14, Gol80, Ano78i]. tablet [Gon08]. tablette [Cav85]. Tablettes [Yus10]. taboos [Høy04]. taboos-and [Høy04]. Tactics [Ehr15]. Tafeln [Vol81]. Tagebuch [Hal78, Wat91]. Tagebuchnotiz [Joh82]. Tagung [Pur88, Pur90, Pur91, Tob81, Ano91-32, Eng81b, Eng81a]. Tahta [Gri76]. tails [Sen94]. Taine [Lew95]. Taisbak [Ano87d]. Takakazu [Ebe96, Ano78-32, Ano78-32]. take [Imh21]. Takebe [Ebe96]. Tales [Ayo06, AKW04]. talk [KS75]. taman [Hog86]. taming [Ano92o, Kje17]. tangent [Ped75a]. Tantalizer [Har77b]. Tantra [Tee84]. Tapp [GG07]. Tardy [Bot78, Bot80b]. Tartaglia [Scr03, Kic15]. Tartu [Dem74, Muü78]. Tarwater [Ano81s, Gra77]. Tashkent [Ken75, MY99]. Taton [Ano79l]. Taunus [Ken76c]. Tauris [Kat92]. Taylor [Ano77p, Mai21, Gup74, Plo01]. Tazzioli [Sch12, Gra99a]. Tchernetska [Sch13b]. Teacher [Zas79, Ano76v, Bie74b, Par05a, Par07]. Teachers [Gri76, Mac78a, Rog74, AB84, Ehr18, Hug75, Mal79]. Teaching [All09, Ano88-30, BS13, Cri99, dS99a, Ano74r, Ano74-64, Ano81-32, Ano89u, Arb86, Gis18, Hel23, Kas03, KAHR08, SG07, Vog94, SG09]. Tech [Gra77]. technical [Ano84c, Syl81, Tob06, Roy17]. Technik [Jur07, Row85d, Row08]. Techniker [Ano90l, Row85b]. Technique [Bre07, And83, Dho88b]. Techniques [Ano80-32, Høy02b, Kno82, Mal97, Ano91r, Dho88b]. Technische [Ano90l, Mai87, Row85d, Spa89]. Technischen [SS08, Tob06]. technischer [Bar78b, Gar05]. Technisierung [Kna88]. Technology [Ano87n, Ano87-31, Ano89-33, Atz06, Sau06a, Zep12, Ano76-34, Ano82-27, Ano84m, Ano88k, Ano93s, vD09, Ano79p]. técnica [Ano82-27]. tecniche [Ser90]. Ted [Lor18]. Tee [Ano87b]. Teheran [Ano93s]. Teichmüller [Ano86-28]. Teil [Ano81x, Hes90]. Telefunken [Tob10, KS12]. teleparallel [Sau06b]. tell [Str23]. Temple [Row85a]. tempo [Ser90]. temps [Bré12, Fio01, Hal86, Ano87i]. Ten [Cro75c, Cro75b, Joh84]. Tenbin [Sid12]. Tenn [Ano95j]. tennis [BS06, Bel07a]. Tension [IM98]. tensor [FK90, Goo77]. Tensorkalküls [Ano95h]. tentative [Car18]. Tenth [Van10, Ano87d, Ken84]. Tenth- [Van10]. tenu [Bun76b]. Teofilo [Vag82]. teorema [BFG94]. teoria [Ano81r, Sai94, Gar98]. teorii [Ano81h, Sel74, Smi79b]. Teoriya [Ano81o]. Teresa [Rom05]. term [Ehr15, Sid04]. terme [Sch82b]. Termessos [Row06a]. terminologia [Lam87]. Terminologie [Lam87]. Terminus [vdW80]. terms [Ano74n, Spa11, War24, Syl81]. Terrace [Ano87n]. Terzi [Ano88r]. Test [Ine94, Pie83]. Testimonies [CM19a, CM19b]. Testimony [Dal92]. Teubner [Ano89k, Row85b, Ano89k]. Teubner-Archiv [Ano89k]. Texas [Ano74r, Ano76e, Ano76h, Ano77d, Ano81k, Ano88h, Ano93f, Gra77]. Text [Plo96, Sai05, Ano84f, Ano88m, BT84, CH90, Dra84, FU01, Ung20]. Textband [Bar82b]. textbook [BS13]. textbooks [Bek88, Bur92, Kar22, Lau87]. Texte [Ano89j, GG07, HB^{+}13, Jay83, Lê21, Mat75]. Textes [Sen99, Syl07, Fra04a, Gra76a, Lam19, Vil96, Dho91]. Texts [Ber04, Lan96, Ano92i, Cre16, HB^{+}13, Imh03a, Pro22]. Textstellen [Scr03]. textual [De 84]. TH [Row08, Sch90b, Joh84]. Thābit [BH89, Hog85c]. Thakkura [Knu11]. Thanks [Ano75-27, PH99]. thatsächlichen [Vol93b]. Theaetetus [Høy90c]. Their [Ano84h, Ano85h, How18, Kin86a, Ano93p, Bul09a, Dau80, EI12, GP91, Rey09, Sch77, Sen91, Ung23, Zas83b, Zho21]. Thekal [vdW80]. Them [GG96, GG97]. Thema [Edw80]. Theme [Kas90a]. Theodore [Hun07, Sch91c, Hog81a]. Theologen [GG07, TC05]. theologians [TC05]. Theology [Hog03]. Theon [McK82]. Theorem [AF94, AÖ01, Bel12, Fel78, Kje00, Phi79b, Sør05, AD78, Bal04, Bur12, CR22, Cri78, Dal88, DF21, Fle89, Gol92, GG79, Hen75, IN84, KK11, KQS85, Lüt19, McC94, Meo04, Pie93, Row21, Sto78, Yad80, Fis11, Rus80, SS88, Sto76, Tur13, Wag85, Cha76, Gra13, LP10]. theoreme [Ano81j, Pet74, Row15, dSG10]. Théorèmes [Gra08]. Theorems [Cog12, Lê18, Sch88, Zun02]. Theoretical [Ano88f, Ano89o, Ano90m, Plo06a, Ano87-29, BJ00]. Theoretiker [Høy90b]. theoretischer [Sch82b]. theoriam [Rid83a]. Theoricae [Mal13]. théoricien [Bor14, PR12]. Theorie [Høy90b, Sch82b, Ano81m, Ano89-30, BKM^{+}82, DD87, DG90, Fre77, Gol82, Koe94, Lai10, May76, Maz21, Roz90, Vol16, YDD83, Cog15, Sch02, Haf17, ET14, Lê21]. Theories [Ken76a, Ano89r, Ano91p, GG14, Meh76, SS22b, Bar74]. Theory [Ano81h, EU99, Epp95, Fer98b, Gra09b, Gra18, Haw03, Kis99, Law05, Mar99, Moo02, Özd00, San09, Sho02, Thi97, vdW76a, Ang79b, Ano80e, Ano80-34, Ano81z, Ano84j, Ano84i, Ano87j, Ano89c, Bak75, BPL17, BR16, Buc77, Bur86, Car24, Cer17, Cha77, CR22, Cri86, Cri88, DD87, Das79, Das80, Das81, Dau75, Dor95, Eng80, Fer07, Fer08, Fis94, Fis11, Gar85, Gra78a, GG80, Gra86, Gra94b, Gra94c, Har77a, Haw75a, Hel23, Imh03b, IN84, Kno83a, Kë96, LY79, Lüt82, Lüt09a, Moo76, Ney78, Phi78, Pla76, Pul05, Rei76, Ros17, Sau06b, Sch06a, Sho03, Smi79b, Syn83, Vek75, Ver13, Ver15, WM14, Wol08, YDD83, Phi79b, Row06b, Bel12, D’A10, Edw77b, Gra94b, Gra97a, McL12, Ano74-27, Gra94c, McG74, Pye75]. Théret. [Vol16]. thermodynamique [Hei81]. theses [Ano74-38, Gup76b, Gup93]. Thesis [Ano80-34]. Thessaloniki [Kas90b]. things [Tob10]. Thinkers [Hod77]. thinking [Sch91c]. Third [Ano92-51, DK86, Gab76, Pio85, Rog77, Wag85, Dun89, Gup74, Joh86, SS86, Wag79, dY92, Ano77-42, Ano85-30, Ano91-33, Ano92-47, Ano92-48, Ano92-49, Ano93-34]. Thirteenth [Yon75, Nee79, Ung78]. thirteenth-century [Nee79]. Thirty [Scr85c]. Thom [Gre85]. Thomas [Ano74-30, Ano87-34, Ano93i, Bay11, Gui09a, Gui09b, Hel83, Pei22, Sca94, Ano80q, BS09, Bel02, Bla14, BM74, Cra00, Dal91, FG90, HP20, Pep75, Rom13, Sho85, Smi80, Ste07a, Tat87, Tra21, Ols76, Ped77, Tat11]. Thome [Syl07]. those [Gra76b, Lo 83]. Thought [Moo02, Cas80, Eis82, Fre89, Len79, Mug87]. Thoughts [Kre03, PH99, Pet78b]. thousand [Ano74-35]. threat [vD09]. Three [Dia99, Fle86a, Ger85, Por19, GG05a, Sch86b]. Three-Body [Dia99]. Thron [Pri80]. Thuillier [Cog15]. Thus [CH90]. Tian [Mar95]. Timaeus [AS93]. time [Ano74y, Böl93, Cox75, Fio01, Kid09, Øhr85, Pou07, Sau06b, TC05, Win82, Ano74-30]. Times [Ano84h, Ano87-34, Gra13, GG92a, Gup86, Jam13, Ken76d, Nic13, AO05, Cha76, GG94, KK11, Kre03, RH11, Tur08, Del03b]. Timothy [Fra10]. Tinbergen [Alb94]. tinkers [Hod77]. Titan [Thi06, DGD04]. Titeica [ABBS09]. Tiziana [GG05a]. Tobies [DR19, KS12, Row85b, Row19, SS08]. tochnykh [Ken75]. today [Bie74a, Ano81v]. Todestag [Sau10]. Toepell [Ano88n]. together [Bel07a, BS06, RFS89]. Toke [Mon17]. Tokyo [Ano74-48, Ano74-76, Ano87g, Ano88j, Ano91o, Ano93g, Gra94a, iÔ75, Sid12, Syl83]. Tokyo/Boston [Syl83]. Tollmien [Row23]. Tom [Ano77l, Dum15, Ken76d]. Tomash [Knu83a]. tome [GG76a]. Tomorrow [Tob10]. tomorrows [LBB^{+}04]. Tomus [Net24, Net24]. Ton [Høy95]. Tonsystem [Høy90b]. Tonsysteme [Høy90b]. Tony [Wil07]. Too [Pie83]. tool [SK24]. Tools [KAHR08, Goe08, Goe09, SG09]. Toomer [Ano77i, Ano88j]. topic [Bul09a]. Topics [Eve76, Buc80]. Topkapi [DS99c]. Topologe [Aud12, Krö14]. Topologie [Krö06, Krö17, BKM^{+}82, Fre76, Pon78]. topology [Dau75, Die75, Fre75, Gra91a, Lew04a, McC90, Moo08]. Torch [Ano88f]. Torino [GG05b, Taz09]. torno [Aus16]. Toronto [Ano74o, Ano77p, Ano83k, Cib76, Dav75, Maz77, Nic13, Pyc84a, dBR76, Ano74-47, Ano80r]. Torricelli [Eva94]. Torsten [Pej07]. Tortoriello [Taz09]. Tosho [Ano78-32]. total [Ber87]. Toti [Ano81r, Ano85f, Mal06a]. Toulis [McK82]. tournant [Ano92-50, Ano92-28]. Tournès [Fer11]. Touzé [Fra04a, Wil08b]. Tov [Rab74]. Tov-elem [Rab74]. Tower [Bos02]. Tr [Wol03]. tra [Bee09, Rom97]. Trabucco [Coz12]. Tracing [CM19b, Mar95, BGSS16, CM19a]. tract [Rab74]. Tractatum [Syl07]. Tractatus [KO09, Lef10, KO09]. tractionnelle [Fer11]. Tracts [Too80]. Tradition [Ano88r, Ano92g, Che94, Hog03, AH20, De 95, De 05, Eli18, Jam88, Ott89, Rom13, Bo24, Hir74]. Traditional [Gup91, LYK89, Mar90c]. traditione [Net24]. traditions [Che17, De 84, Fri81a]. Tradizione [Ano90c]. traducão [Ges18]. Traduction [Ano77j, Cog15, Mue96, Ped77, ET14, Gra08, Pei17]. traduit [Lê21]. Trained [Gup91]. Traité [Ped77, Chr91, Tat86]. Traités [Fra04a, Wil08b]. Trans [Ano92n, Høy24a]. Transactions [Lo 84]. transcendence [Pet12]. transcendental [Kno06]. transcribed [Mah78]. Transcription [Row85e]. Transcriptions [Sch13b]. Transfer [Sho02]. transfinite [Fer95]. Transformation [Fen09, Ano92g, Che10, Haw77a, Maf10, Mol76]. transformations [Cog15, ET14]. transit [Sør14]. Translated [Aga75, Ale04, Ano74z, Ano74-30, Ano79l, Ano82d, Ano88j, Che14b, Cib76, Cog12, D’A10, DS99b, Eve76, Goo77, Gra94a, GG10, Gra91b, Gra10b, Haw96, Ken76b, Ken82, Law05, McK82, Pla76, Pye75, Row83a, Row90a, Tru75, Ano74v, Ano89d, Bow84, Cal92, Gra97a, Her14, Jon76a, SS17]. Translation [Gin83, Knu11, Mon17, Plo06a, Row85e, Ste80, Tee78, Vit96, vM84, Ano77i, Ano78h, Ano84f, Ano88m, BJ00, BK06, CH90, Eli18, Eli24, Gil77, Gra83, Hog85b, Krö08, Lo 83, Rus80, Syl81, Syl83, Tee84, Too80, VHK16, Vol10, Wil08a, Xu05, dY92, Jah12]. Translations [DZ97, Mar97, AG77, Haw77a, Haw77b, Vil96]. Transmission [Ano87-36, GO99, Van06, Ano85-31, Ano92g, Ano92-27, Auj86, Car18, DS^{+}02, Kno83b, Neu88, Scr84]. transmitted [Hog81b]. transnational [Ehr18]. Transworld [Ano78g, Eve76, Jon76a]. trasformazione [Cap13, Maf10]. Tratado [Sil08]. trattato [Van80]. travail [Ano92-50, Ano92-28]. travaux [Ano77j, Cho75, DD87, Ebe96, Fre77, Mah78]. travel [CM22]. traveling [vBS20]. travels [Vog83]. travers [Mar90b]. Traylor [Ano74r]. Treasures [Don98]. treated [Ine91]. treating [Tat75]. Treatise [DS99b, Ano80e, Ano80-34, Fol83a, Fol15, LH84, Rai92, Rom13, Ros76, Tee78, Nee79]. treatises [AH20, Van80]. Treatment [She96, She94]. tredition [Row23]. Trends [Ano92-29, Ano86j, Ano91h]. Trento [DG90, Ano89-30]. tretii [Ken76d]. trial [SS10c]. Triangle [Fri81a, Hug89, LY79, LY80, PZ11]. triangles [AS93, TS78, Glu77]. Triangular [Tat11, BS09]. tribes [Ano76y]. Tribute [Lê24, Scr96, Ano81s, Ano84-28, Dra78, Row83b, Thi07]. Tricentenary [Ano84-29]. Tricomi [Luc21]. tried [Str23]. triggered [SS10b]. trigonometric [GG93, Ham80, Kat87]. trigonométrie [Che94, For10]. trigonométriques [Ros81]. trigonometry [Che10, Hai86, MW17, Ung20, Van09, Eva13, Fra17b]. Trinity [Ano79i]. Trionfi [Pan18]. trip [Str94, Won22]. Triples [Fri81a, Lum80]. trisections [Hog81b]. Troisième [Gis14, Auv13]. Tropfke [Ano81i]. Trotz [Kni01]. troublesome [GG93]. trouvé [LP10]. trudy [Kin86a]. True [Dau12, Dry14, Duc16, CG92, Kre03]. Truesdell [Bos03, Fra06, Ano86w]. truth [Put75, Ric80]. truths [Ser06]. Tschirnhaus [KK90]. t·a [Hay97]. Tucci [Net24]. Tucker [Kje00]. Tuckwell [San06]. Tudor [Jam13, Tan74]. Tullio [NT05]. Tummers [Lo 85, Vit96]. Tung [Dum18]. Tunis [Mar90a]. Turbulent [RH11, Nic13]. Turici [Bar77]. Turin [Gia90, GG05a, GG05b, Sai94, Luc18]. Turing [Daw16a, Por19]. Turkey [EI12, Inö06]. Turkish [EI12]. Turn [Ano81h, BF00, DF21, Gar86b, Mar04, Wan22]. Turner [Zas83a, Fow84]. Turnhout [Gil15, Maz21]. Turning [Wol03, Lau99]. Tuschmann [Ano95m]. Tusculani [Ano87d]. Tusculanum [Ano87d, Sid06b]. Tusi [SI19]. Twentieth [Cog12, Kje17, Lüt08, Ben95, Tro83, Glu05, Mar04]. Twentieth-Century [Glu05]. Twenty [HK98, Wag15]. Twenty-One [HK98, Wag15]. Twisted [Wei98]. Two [Ano86i, Ano86-28, Ano89-35, Arc09, De 15, FT99, Gro87, Gue88, Hog02b, HP20, Kul76, Moo03, Zho21, Ano76-44, Ano78h, Bul22, Che17, Fri81b, GN05, Sch91b, Sch13b, SS94, SSZ07, Sma16, Ung23, Row06b, Wue05, Yus95, Str82a]. two-centuries-old [Fri81b]. TX [Ano93f]. typed [Mah78]. typescript [Ano79k, Ano81j, Ano81n]. Tzelekis [Zer75].

U [Ano77r, Ano81r, Bie74a, Ano89k]. u.a [Bul22]. U.K [Moo75a, Tan74]. U.S [Ano84g, Ano84i, Ano95f, Bow84, Bun76a, Ped75b, Rid94, SS94, Tan74, Tat75, Wil75a]. U.S.$60 [Gab76]. U.S.A [BT81]. Überlieferung [Bus76, Gol76]. übersetzt [Scr03]. Übersiedlung [Bie74b]. Ubiratan [Sco22]. Übungen [Høy24a]. UK [Sch05]. Ulam [Bir77, Moo82a]. Ulf [Row08]. Ulrich [Plo06b, Row07, Yon75]. ultimate [Lau92]. Umberto [Arc99, Arc17, Ken82, Sch99a]. Umbruchs [Cif00]. Umfrage [Row23]. unbekannte [Fol83a]. uncertainty [Sch91b]. undergraduate [Ano78-27]. undergraduates [Eis80, Pet78c]. understand [Zhu20]. Understanding [Sør05, McK82, Nic93, Pet12, oSTT79]. Underway [PH97]. unedited [CB21]. Unendliche [Che14b, Ano89k, Lam87, Osw17, Déc10]. unendlichen [Sch82b]. Ungar [Bow84]. ungarisch [Has10b, Has10a]. ungarisch-jüdischen [Has10b, Has10a]. Unguru [Sai05, Ste80, Fri24]. unified [Sau06b]. uniform [RKL05]. Uniformisation [dSG10, Row15]. uniformity [Wat83]. Union [May75c, May77c, Scr82b, Scr84, Ane87, DV80, Sen04, Shi86, Vuc99]. unit [Abd08, Ano74-40]. unitaire [Fre77]. unitary [Gal10]. United [Cas77, San03, ABC^{+}88, Arn88a, SS97b, Shu80]. units [CV22]. Unity [McL12]. Univ [Ano84f, Ano84d, Ano89c, Ano95l, BT84, GG05c, Hoc96, Høy90a, Sch91a, Sch91c, Sch05, Sen06]. Univac [Asp83, Ano79t]. Universal [Sch05, Gis18, Pyc97]. universe [Gin83]. universelle [Gis18]. Universidad [Ano90-32, Ano92-44, Gar98, Pan08]. Università [GG05b]. Universitaire [Fre76]. Universitaires [Cog15, Gra12, Lüt15, Mar10a, Rei15, Mue96, Ple78]. Universitaria [Ano92-44]. Universitarie [Oak12]. Universität [Ano84k, Ano86h, Ano87m, Ano90l, Ano90t, Ano90-33, Ano91-33, Ano92w, Ano94b, BMW08, Cal75, Has10b, Has10a, Neu93, Sch09, Seg90, Row85d].

Universitäten [SS08, Tob06]. Université [Ano81j, Ano81n, Ano90-34, Bun76b, Kit81, Kno88a, Tat86, Ano75q, Ano93-33, Bor89, Mar90b]. Universités [Ber87, Sal87]. universities [Tob06]. University [Abe91, Abe06, All04, Ano74i, Ano74o, Ano74r, Ano74s, Ano74n, Ano74x, Ano74-28, Ano74-29, Ano74-40, Ano74-45, Ano76e, Ano77d, Ano77g, Ano77f, Ano78h, Ano78x, Ano78y, Ano79k, Ano80r, Ano81k, Ano81l, Ano81u, Ano82d, Ano83k, Ano85f, Ano87k, Ano88e, Ano88f, Ano88h, Ano88l, Ano88-31, Ano90s, Ano91d, Ano91e, Ano92g, Ano92o, Ano92x, Ano92-30, Ano93f, Ano93s, Ano95g, Ano95k, BGG06a, Bay11, Bee06, Bel07a, Bel08, Ber04, Bra12, Bra17, Buc80, Bul16a, Cas77, Cas80, Cib76, Coh82, Cri24b, Cro77, CH90, Dau12, Dav75, Dav19, Duc16, Dum18, Eis75, Enr80, Eva13, Fer04b, Fra10, Fra17b, Fre89, Gab76, Gil75, Gra77, Gra83, Gra10a, Gra11, GG05a, GG86, GG90c, GG90b, GG90a, GG92b, Gra06, GC21, Gup91]. University [Has10a, Haw81, Hel83, Hod75, How18, Hud12, Hun07, Hun12, IN83, Jam13, Jon85, Kat11, Kat20, Kin86b, Kje17, Lan07, LY79, Lev97, Lew98, Lew99, Lon21, Maa07, Mat74, Mic23, Mon17, Moo77, Mor19, Mos16, Nak03, Nee79, Nor07, Oli08, Per93, Per10, Pin76, Plo06a, Por05, Pri80, Pyc84a, Rei76, Ric82, Ros74, Sag91, Sai74, San05, Sch91b, Sch13b, Sen10, SG09, Shi88, Sia23, Ste05, Sti78, Syl83, Tat75, Tee78, Tim09, Tro74, Tru75, Wal17, Was20, Wil83, Wil05, Wil07, vM07, Ano74-58, Ano84l, Des07, Fle75, Nur92, Zit07b, dS90, Ano74-44, Ano74-67, Ano74-75, Ano78f, Ano79z, Ano87-30, BMW08, DF10, GG75b, Ing11, Muü78, Neu93, Par04, Sar15, Shi94, SS93]. university-affiliated [Des07]. University-Trained [Gup91]. Unknown [Kje17, Sch98, vM85, De 15, Des10, Fol83a, LH84, Sch93b]. Unpublished [FR82, Ane88, Ano80-35, Aub14, HNE21, Jen86, Neu89, Osw17, RP19, Tat75, dMC92]. Unscrambling [Gre88]. unsentimentalised [Rai92]. unsolvability [Pes03, Con06]. Unsolved [Ven02]. Unterricht [Row19]. Untersuchung [BB04, Imh03a, Lüt09a, Pul05]. Untersuchungen [Bot80a, Lév99]. until [Mar20]. Untiring [Bec01]. unveröffentlichten [Jen86]. Unviability [AR24]. uomo [Gan86]. Upcoming [Ano76-45]. updated [Ano85r]. upon [Vol10]. UPR [Ano91r]. Upside [Ano82f]. Uqlidisi [Pin80]. uravneniya [Git74]. Urbino [Ano85f]. Urteile [Has10b, Has10a]. US$14.95 [Hug76]. US$169.00 [Bra09a]. US$23.95 [Hun12]. US$28.95 [Gra11]. US$49.50 [Sen10]. US$62.00 [D’A10]. US$65 [Hud12]. US$79.00 [AH10]. US$8.95 [Clu76]. USA [Ano74-49]. usage [Bec19, HF19]. Use [Ano83z, Ano74-40, Ano82-29, GG93, Gra84a, Haf17, Pet78b, Pet12, Sch77, Sid04, dW10]. used [Ano74n]. useful [McK82, oSTT79]. uses [Ano93p, BGG06a]. Using [Ano86r, Kat07, Shi93, Erl92, Mic88, SGJ05]. USSR [Ano89-33, Ano74-37, Ano87-35, Ano89-33, DK86, Hol15]. ustorio [Ros76]. Uta [Ano77f, Neu86]. utilisation [Auj86]. Utopias [Tur08, AO05]. Utrecht [Aup76, GG77]. Uzbek [Ken75]. Uzbekistan [MY99, Mat93].

V [Ano74m, Ano77m, Ano88k, Ano91e, Ber87, Dav76, Dho85, Eis78, For10, Haw96, Jon76a, Lin74, Pec20, Phi79a, Rei76, Sal87, San03, Tot79, Ano91f, Bie80, Sel74, Sti76b, Vol91, Auj86]. Va [Arc99]. Vaclav [Zla23, DS19]. vadis [Imh21]. Valais [Bor14]. Valdés [Aus16]. Valentin [Fol87]. Validity [Pes01]. Valla [Ray23]. Vallesia [Bor14]. valuable [AB09]. Value [Rus80, Pro22]. Vanden [Ste09]. Vandenbroucke [Hes90]. Vandenhoeck [Ano88n, Sch90b, Ull03]. Vandermonde [Tat88]. Varāhamihira [Hay87]. Varanasi [Yan83]. Varenne [Ano89t]. Varia [Ano95n]. variable [GA20, Smi79b, Lê21]. Variables [Dom04, Gre84, Pan08]. variación [Pan08]. variant [Gab76]. variational [Fra92]. Variationen [Edw80]. Variations [Thi97, Cal83, Fra09, Maz16, dV22]. Variationsprobleme [Gra08]. variationsrechnung [Ano92j]. Varieties [VV14, GG90e]. Varignon [Hes90]. Variorum [Ano89h, Mor96]. Various [Bec01, PGN18, SS18]. Vasco [Ped77]. Vassar [Row89]. VEB [Lem74, Seg76]. vecteurs [DR03]. vector [Dor95, Gra80a]. vectorial [Smi82]. veka [Ano81o, Sti76b, Kë96]. velocities [Gol80]. vendicato [Gal12]. Vendredi [Ano83o]. Veneto [Ser90]. Venezia [Ser90]. Venice [Ser90]. Venn [Abe06, Edw04]. Venti [Ano85f]. Veränderlichen [Haf17]. verbandsteorie [Bir80]. Vereinigung [Ano91c, Ano92h, SS89, Ano87m, Ano92-46, Ano93k]. verfassten [GG13]. Verità [Ser06, Rab08]. Verlag [Ale14, Ano74q, Ano77i, Ano78j, Ano84e, Ano84j, Ano85e, Ano87j, Ano87g, Ano87f, Ano88j, Ano89e, Ano89f, Ano89j, Ano91f, Ano92l, Ano93g, Ano95f, Ano95h, Bar78b, Bar82a, Bar82b, Ber07, Bir80, Bla05, Bor18, Bos03, BB04, Bra07, Bur08, Cal75, Cal83, Cal92, Coo95, Cri17, Das81, Dho88a, Dho11, Dil81, Don05, Don08, Dru91, Edw77a, Edw77b, Fol87, Fra06, Fre92, Gar05, Gra94a, GG07, Gra90, Gra91b, Has04, Hei09, Her14, Hog85b, Høy24a, Høy24b, Jay83, Jur07, Ken76c, Kit78, Knu74, Lef10, Lem74, Mel06, Mon75, Mue80, Ora82, Pet82, Plo06b, Row07, Row08, Sau14, Scr07, Scr08, Seg76, Seg90, Ser18, SS08, Str78, Swe79, Syl03, Toe96, Ume76, Van05, Van06, Vol81, Wol91]. Verlagsgesellschaft [Ano74v, Ano89k]. Verley [BKM^{+}82]. Vermessungsproblem [Vog83]. vernacular [Høy06]. Veröffentlichungen [Scr03]. veroyatnostei [Ano81o]. vers [Pon78]. Verschueren [Maz24]. versehen [Bul22, Rom04]. version [Eli24, Gol92]. versions [Ano80h]. versus [Cra00, Ful99, SS10c]. Vertices [Erl97]. Verzaglia [FR82]. Verzeichnissen [Bul22]. Vestigia [Eva98]. vetus [Lo 84]. VI [Kno88a, Ple78, Ter89, Ano81e, Ano84g, Fre92, Kat11, Gol00, Pou07]. Via [Ano92s, Cap13, Maf10, Row21, SS22b]. Viaggiando [Taz09]. viata [And75]. Vicarious [Car22]. vices [May75b, May75a]. Victor [Ano88e, Arc96, Ayo06, Kje17, Syl81, GG98, Gra98b, Ano91-30]. Victoria [Abe91, Ano82f]. Victorian [Lew90, Wil07, Cri04, Cri06, DR16, Ric96]. vie [Ano90h, Bie78, Bor91]. Vienna [Ano90-36, Ano79-41]. Vienno [Scr03]. Vierfeldertest [Ine94]. Viet [Thi75]. Viète [Car22, Dho87, Glu77]. View [IM98, McL12, Böl93, CB21, Cra09, Gal10, Lüt09a, Øhr85, Par92b]. viewed [Hol15]. Vieweg [Gam96, SS89, SS97a]. viewpoint [Zhu16, Row85a].

Viganò [Ano85g]. Viggo [Ano80-36]. VII [Ano74s, Ber87, Bla05, Hog85b, Sal87, Scr07, Too80, BKM^{+}82, Gra10a, GG76a, GG77, GG90c, Gra90, Grö86, Jay83, Row90a]. VIIe [Ano89t]. VIII [Ano79l, Ano92-27, Kin86a, Ser18, Ano78j, Ano81s, Ano87j, Ano88m, Ano88l, Ano89g, BT84, De 20, Enr76, Enr80, Fer11, Høy07, Knu11, Lüt20, Moo82a, Nak03, Pri80]. VIIIth [Ano87-35]. vil [Ped75a]. Vilant [Cra12]. Villa [Ano89-35]. Villars [Fre77, GG76a]. Vincent [AD78, Llo79]. Vincentio [Van84]. Vincenzo [Gal12]. Vinci [Ray23]. Virginia [Zas83b]. Virginio [Ken82]. visages [DR03]. viscous [Gre88]. Vishcha [Rei76]. visibles [Ols76]. vision [BS05, Fra87]. visione [Pan18]. Visionen [Row19]. visions [Lew90]. visual [Car18, Ste07a]. visualization [Erl94]. vita [Arc17, Mai07, Pan09]. Vital [Bel08, Sch06b]. vite [Ano82e]. Viterbo [Dur21]. vitesse [Ano91n]. Vito [Ano81t, Ano81t]. Vitrac [Mue96]. Vittorio [Ano77h]. Vivienne [Ano92-28]. Vladimir [Plo06b]. Vladimirov [Lin74]. Vleck [NB77]. Vogel [Ano81i, Fol83b, Ano79o, Fol86]. Voici [LP10]. voie [Pon78]. Voigt [Mon83]. Vol [Ano84i, Ano87d, Ano87e, Ano87j, Ano87h, Ano88f, Ano88i, Ano89k, Ano91g, Ano92l, Ano93f, Bar82a, Bar82b, Bie74a, Buc77, Che94, Ges18, Haw03, Hod75, Hog85b, Hog03, Hog24, Høy95, Jon76a, Kle17, Lei79, Lo 84, Lo 85, Mac76, Par89, Par92a, Pec20, Pei22, Pet82, Rid83a, Rid83b, San03, Sch89, Sel90, Ste80, Swe79, Syl81, Syl03, Ter89, Vol77, vM84, Ale14, Ano77m, Ano77r, Ano77g, Atz06, Ber04, D’A10, Don05, Fen08, For10, Gol82, GG05b, GG07, Haf16, Høy09a, Jon77b, KS12, Law05, Mro16, Mue96, Neu08, Ple78, Rec07, Sau06a, Sau14, Sch11, Sch07, Sch00b, Zep12]. volgarizzato [Ray23]. Volker [Don05, Don08, Hei09]. Volkert [Krö06, Sch90b, Che14b]. Vols [Dav90, Par91, Par92a, Ano74q, Bur08, Ewi92, Fol15, Gal12, Høy23, Kin86a, Lo 85, Moo82b, Sel91]. Voltaire [Pei17]. Vol’terra [Ano81t, Ano81t, Sch12, Ano82-28, IN84, Jaë19, VMT09]. Volume [Ano77s, Ano81k, Ano81p, Ano88e, Ano91d, Ano95j, Ano95p, Ano95q, Ano96b, Ano96c, Ano97b, Ano97c, Ano98e, Ano98f, Ano99b, Ano99a, Ano00b, Ano00a, Ano01b, Ano01a, Ano02e, Ano02d, Bar77, Ber87, Edw77b, For10, Fra04a, Goo77, GG90b, GG97, Ken76d, Maa07, Sal87, Sch13b, Tee78, Wil08b, Ano74-53, Ano74-54, Ano75n, Ano75o, Ano76z, Ano76-27, Ano77-31, Ano77-32, Ano78v, Ano79-27, Ano79-28, Ano80o, Ano81-33, Ano81-34, Ano82m, Ano82n, Ano83l, Ano83m, Ano84o, Ano84p, Ano85l, Ano85m, Ano86p, Ano87p, Ano87q, Ano88s, Ano88t, Ano89p, Ano89q, Ano90o, Ano90p, Ano91l, Ano91m, Ano92y, Ano92z, Ano93q, Ano93r, Ano94k, Ano94l, Ano03n, Ano03m, Ano04l, Ano04k, Ano05n, Ano05m, Ano06j, Ano06i, Enr85a, Fel83, Fu91, GG86, Hof87]. volume [Wag79]. Volumen [Bar77, Rid83a, Rid83b]. Volumes [Mos16, Ano86q, Ano87-33, BF89, Bul22, Gab76, Moo03, Sch13b]. Vom [Ano95h, Mue80, Spa89, vdW80]. voor [Ano81e]. vor [SS88]. Vorarbeiten [Hes90]. Vorgeschichte [Ine88]. Vorlesung [Fol09a]. Vorlesungen [Ano89k, Ull03, Ano78j]. Voronoy [Syt91]. vortex [Vek75]. Vospominaniya [PP78]. vosrozhdeniya [Sti76b]. vostoke [Ken75]. vremen [Ano77l, Ken76d]. vremeni [Ano77l]. Vrin [Ano77j, Ano79l, Ano79m, Ano81m, Dra84, Ebe96, Kno07, Lê21, Mac76, Mah78, Roz90, Syl07, Vil96, Vil97, Cor96]. vs [GP91]. VT [Mor96]. vu [Mar93b]. Vuibert [Høy07, Maw06]. Vuibert/ADAPT [Høy07]. Vuibert/ADAPT-SNES [Høy07]. vuole [Arc17]. VV [Ano81h, Kin86a, Vol91]. Vzaimosvyaz [Sti76b].

W [Abe06, Ano74v, Ano77j, Ano77h, Ano77m, Ano80b, Ano88g, Ano89d, Ano90l, Ano90-35, Ano92l, Ano95l, Bar82a, Bar82b, BKM^{+}82, Bos77, Bul22, Eis75, Fel78, Haw03, HNE21, KB74, KK90, Lef10, Pet82, Pri80, Rid88, Rid94, Row83b, Str82a, Swe00, Tru75, Van06, Van14, Vil96, Wil77b]. W. [Cho92, GG75a, GG94, MB20, Øhr85, Van14]. Wabkanawi [Moz13]. Waclaw [Moo82b]. Wadsworth [Dav90]. Waerden [Bor18, Dil81, Fre93, Pin76, Ano94h, DS97, Fre93]. Wafa [Mat74, Ray12, Ken84]. Waldo [Thi06, Zas83b]. Wales [Jam13]. Walke [Ano92b]. Wallace [Hod75, Cra99, Cra00, Pan87]. Wallis [Maa07, Pan09, Gui01, Mai07, SB12, Ste01, Ste12, Mos16]. walls [Mel04]. Wallstein [Sau14]. Wallstein-Verlag [Sau14]. Walter [Ano81i, Ano89f, Ano92l, Bar77, Eve76, Net24, Plo06b]. Waltershausen [GG13, GG13]. Walther [Bor14, PR12, Row08]. Wang [Ano91g, Bun76a, Lê24, lLW13]. want [Ano87f]. Wantzel [Lüt09b]. War [Ano81-46, Dau80, Gui01, Kje00, Rid94, Roy17, Sch12, VMT09, CB21]. Ward [CS81, Hog91]. Wardhaugh [Hun12]. Waring [Edw91]. Warren [BT81, Haw81]. Wars [Ano86i, Gue88, GN05, SS94, Arc09]. Warsaw [Ano76x, Moo82a, Moo82b]. Was [Gil87, Dav90, Lüt09b, Høy90b, LY79]. wasan [Mar90c, Ung20, Ung22]. Waschkies [Mue80]. Washington [All09, And07, Ano86s, Ayo06, Cal08, Enr85b, Enr85a, GG08, Kat07, Lew95, Thi06, Thi08, Fen03]. Waterloo [Høy90a]. waters [Maf10]. Watson [Ros74, Tan74]. Waves [Atz06, Dij04]. Way [Lon21, Chr07, Maf10]. Ways [BB21]. Weaving [Bla14]. Web [Krö08, PH96a, Vol06]. Weber [Lê21, McG74, Shu80, Vol93a, Wil75a, Haf17, Sch14, Haf16]. week [Dho83]. weekends [Ano86-28]. Wege [Ano81g]. Weierstrass [SS89, Coo95, Lau92, SS97a, Böl93, Dug76, SS88, Sto76, Böl93, Lau92, Sch98, Sch19, SS97a]. Weierstrassian [Haw80]. Weierstraßschen [SS88]. Weighing [Mel02]. Weight [Lev97]. Weil [Edw77b, Syl03, Bur86]. Weimar [Ton82]. Wein [Ric84]. Weinheim [Tob90]. Weinheim/Basel [Tob90]. Weinstraße [Ano91-31]. Weisbach [SM18]. Weissenborn [Bos77]. Welche [Spa11]. welfare [Lüt19]. Wellesley [Zit07a]. Wells [Ano81k, Ber74, Tan79, Tee79]. Welt [Ano89k, Cif00]. Welten [Gra10b]. Wending [Ano87h, Dun89]. werden [Row23]. Were [RT02, Hog81b]. Werk [Scr07, Dho88a]. Werke [Don05, HB^{+}13, Hes90, Plo06b, Scr07, Syl03, Ano74q]. Werner [Scr07]. Werther [Sch11]. Wessel [Atz04]. West [Lor18, Bus76, Cra98]. West-Europa [Bus76]. Western [BLF92, SS07, Swe00, Asc84, BJ89b, Che10, Fol83c, GGC^{+}05, Jam88]. Westminster [Ano91n]. Weyl [Cog15, Fre92, Jen86, Nab04, ET14, Jen86, ST16]. Wheaton [Ano77-30]. Wheeler [GC82, Rob96]. Where [Kat11]. which [Bur12, CMT79, Spa11]. White [Gra77]. Whitehead [Ano88e, GG03, Ano74-56, GG02]. Whiteside [Gui09a, Gui09b]. Whitman [Gab76]. Whittaker [Mai21]. Who [Aub02, Scr88b, Tat88, Imh21, May75d]. wichtigsten [Ora82]. Wide [Krö08, PH96a, Vol06]. wider [OS19]. widespread [Yad80]. Widmanns [Jay83]. Widths [Mel06, Høy02a]. Wie [Scr81, Hof87]. wieder [KS12, Tob10]. Wiederaufstieg [Dho11]. Wiederkehr [Str81]. Wien [Sch96, Toe96]. Wien/Zürich [Toe96]. Wiener [Ano92m, Fol87, Bir83]. Wiesbaden [Ano89j, Bar82a, Bar82b, BB04, Gam96, Gol76, Gra90, Mat75, Pet82, Sch19, SS89, SS97a]. Wilbraham [Gai75, Ust74]. Wilbur [vdW76a, Fow98]. Wilder [Ano78g, Ano86w]. Wilderich [Ano95m]. wilderness [Hog74]. Wiley [Ano77p, Gri82, Kin86b, Sen91]. Wilfrid [Høy90a]. Wilfried [Daw16b, Høy90b]. Wilhelm [Ano87j, Bar78b, Bla05, Cri17, Gar05, Has10a, Hei09, Her14, HB11, Pet78a, Ano87s]. Wilhelmian [Row85c]. Wilhelms [Has10b, Has10a]. Wilkins [Wil77b, VV14]. will [Imh21]. William [Ano74r, Ano91d, Bay11, Cha76, Dav19, Daw16b, Hel83, Hod75, Lew99, Moo03, Nak03, Sch13b, Wis76, Cra99, Cra00, Cra13, EI12, Hog90, Pan87, Wag78]. Willig [Bos03]. Wilmington [Neu86]. Wilson [Abe09, Ayo06, Gil75, Har77a, Kat20, Sch13b, Ste05]. Wine [Mes94]. Winfried [SS89]. Winston [Ber74]. Winter [Ano92f, Ter89]. wirdt [Ped75a]. Wirkens [Sch19]. wirkliche [Row06a, Wue05]. Wirkungsgeschichte [Ull03]. Wisconsin [Ano78h, Ano90s]. Wiskunde [Ano81e]. Wiskundig [Ano81e, Str82a]. Wissen [Ano86h]. Wissenchaften [Ano89j]. Wissenschaft [Dil81, Ano84k]. Wissenschaften [Ano74q, Ano78j, Ano81x, Ano93x, Bar78b, Haf16, Lem74, Mat75, Scr03, Seg76, Str78, Ull03, Vol81, Grö86]. Wissenschaftliche [Ano81g, Lüt09a, Bos03]. wissenschaftlichen [Gab79, Str85]. Wissenschaftlicher [Jur07, SD94, SEM04]. Wissenschafts [Ano88n, Sch90b, Toe96]. Wissenschafts- [Ano88n, Sch90b]. Wissenschaftsgeschichte [Jay83]. Wissenschaftshistorische [GG07, TC05]. Wissenschaftsorganisation [Row08]. wissenschaftstheoriegeschichtliche [Lüt09a, Pul05]. Wissenschaftsverlag [Sch96]. Wissenszugänge [Böt12, Pei15]. WIT [Ste09]. Witelo [Ste80]. Witelonis [Ste80]. Withdrawal [Jay24b]. without [Des07, Haw75b]. Wittgenstein [Ano81u]. Wizards [McK82]. WKB [Sch77]. Wladysław [Pio85]. wohl [Hof87]. Wojbor [Ale04]. Wolfenbüttel [Ano91u]. Wolfgang [Ber04, Bla05, Jay83, Ken76c, Pri80, Sch19, GG13, GG13]. Wolters [Ano77o]. woman [Böt12, Per79]. Women [AH10, Ano76g, Per80, Ano80-33, Ano84m, Coo88, Roy17, Kni01, Ano89l, Buc80, Wil75b]. wonder [VV14]. wonderful [DV08, Ste09, Tra10]. Woodhouse [Bec80]. word [Gra09a]. Work [Ano79-31, Ber97, Cra99, Her75a, IM98, Nag91c, Nor03, Sch19, Taz97, AGSL76, Ano82p, Ano88e, Böt12, Bur86, But81, Cre14, Dal91, Enr80, Fin11, For78, GC82, Haw80, Jam88, KK90, Nag91b, NB77, Pej07, Ros20, Tat75, Ver13, WM14, Wat86, Ano90f, Nab04, Ano87e, Row83b]. Working [GG15, Nag91b, Nag91c, Goe08, Goe09]. Works [Bar77, Gra06, HB^{+}13, Kin86a, Pec20, Per10, Sar97, SS22a, Ang88, Ano79h, Ano83d, Ano86f, Ano87l, Ano93u, BR04, Bur09, Cib76, Dad84, DD87, GG90b, Gra76b, GA20, Hor91, Nag91a, Pep75, Roz76, Tra10, Ano78-32, Law05, Zas83b]. Workshop [Ano78y, Ano79-47, Ano80-37, Ano82-29, Ano83k, Ano86a, Ano89-35, Ano90-33, Ano91-33, Ano92w, Ano92-47, Ano94i, Gra84b, Kno88b, Mac78b, MP06, Meh80, Row90b, Ano74f, Ano75y, Ano79-32, GG76b, Mug87, Pyc84a, Ano91-32, Nis89]. World [Ano79-48, Ano80-38, Ano81-45, Ano85r, Ano86i, Ano92-51, Ano93s, Ano95-27, Arc09, CB21, GG02, GG03, Gue88, Kat20, Krö08, Mar95, Özd00, Rid94, Sch12, Ano78y, Ano83i, Ano84z, Ano88u, CM19a, CM19b, DV08, Dho90, MS05, OS19, Rom10, Ste09, Str23, vM07, Ano81-46, BPL17, Dau80, GN05, Kje00, PH96a, Roy17, SS94, Vol06, VMT09, Hog03]. World-Maps [Hog03]. World-Wide [Vol06]. Worlds [Ano95k]. Worrall [Ano81l]. would [Tob10]. Woyczyński [Ale04]. Wren [BT84]. writer [Ano86d]. Writing [Ano93f, AD02, Mon11, VF22]. Writings [SS07, Cro76, FR82, GGC^{+}05, Gri76, HB^{+}13, Neu12, RFS89, Tat75, Cul07, GG86]. written [Roz76]. Wroclaw [Ano82e, Ste80]. WS [SS08, Tob06, Tob06]. Wuensch [Row06a]. wujūh [Dje98]. Würdigung [RFS89]. Würfel [Ine91]. Wurzeln [Lau92]. Wussing [Ano74v, Ano95f, SS12b]. Wuttke [Fol87]. Wyoming [Ano74i].

X [Dho11, Krö06, Sim96, Ano84h, Ano85h, Asp83, BKM^{+}82, Gra97a, Lew95, Sid12, Bee09, Fow92a, GG07, Hei09, Hog01a, Par07, Sau06a, Scr03, Thi06]. XI [Ano91d, Ano88e, Ano95f, Gra83, Gra91b, Moo77, Ter89, Yan83, Dje97, Bar82a]. Xiaotong [lLW13]. XIe [Bre07, For10, Giu99, Ai04]. xii [Ano79k, Ano81l, Ano84i, Ano87k, Ano88f, Ano95g, Bar82b, Bra12, Bru77, Dav19, Enr85a, Gri82, Hog85b, Høy95, Par92b, Pys81, Sag91, Sch91c, Sen91, Syl81, Wil83, vdW76a, Too80]. XII. [Ano90-36]. XIIe [AD87]. XIII [Gol82, Gra90, Ano77q, Ano87e, Ano89d, Ano92n, Ano92o, Bra17, Cal92, Cog12, Hud12, Kla84, Knu83a, Lew90, McC90, Dje97]. XIV [Net24, Ano81i, Ano88n, Ano89e, Ano89j, Coh82, Dru91, Enr85b, For10, Hug05, Kat20, Mor19, Pan18, Ano82q, Dje97]. XIVe [Lai10]. XIVth [Ano74-76]. XIX [Ano77l, Ano81h, Ano81o, Dem74, Ken76d, Kit81, Kë96, Ano93g, Cas80, Hil83, Hod77, Ols76, Par89, Tro83]. XIXe [d’E01]. XIXème [Ano89-30, DG90, Dug76]. XIXth [Ano94x]. XL [Ano90f, Ano77s]. XLIV [Rom05]. xlix [Yus10]. xlviii [Ano81g, GG86]. XV [Ano79l, Ano92-27, D’A10, GG90b, Ora82, Ung85, GG76a]. XVI [Gol82, Høy24b, Ano93h, Ano93i, Bar82a, Ewi92, Gol76, Høy90a, Kat92, Lei79, Mar95, McK82, Row85a, San06, Bar82b]. XVIe [Kno07, Ano92-32, Lai10, d’E01]. XVII [Kin86a, Pet82, Ano77m, Ano84f, Cal83, Jon77c, Pin80, Rid83a, Wol03].

XVIIe [Mar10a, Ano92-32, SD08]. XVIII [Ken76d, Pei17, Tat86, AH10, Ano78-32, Ano84d, Ano93f, Ano95l, Hes90, Lev97, Moo03, Par91, Rom05, Sch91a, Sch91b, Shu80, SS89, Tee78, BH14, DP88]. XVIIIe [Pan88]. XVIIIth [Ano90-37, Scr87]. XVIIth [Scr84, Ano85-31, Ano87-36]. XVIth [Ano82l, Ano82k, Ano80-39]. XVlle [Mar93b]. XVth [Pep92]. XX [Ano81h, Dem74, Ano86d, Ano91e, Dau93b, Dru12, GG10, Hal84, Nee79, Zep12, vM84, Gra01, Sti76b]. XXe [Gra08]. XXI [Hog24, Ric98, Gra91a, Hel83, Swe79]. XXII [Syl07, Ano89c, Pec20, Rem12, Van80]. xxiii [Ano88f, Sca94]. XXIst [Ano90n]. xxiv [Dav90, Rid94, Str82a]. XXIX [Høy95, Buc77, Hog03]. XXth [Sti76b]. XXV [Ano77r, Ano86e, Ano92j, Neu86, Sen10, Syl03]. XXV/Synthese [Ano77r]. xxvi [Pec20]. xxvii [Ano78h, Hes90]. xxviii [Tee84]. xxx [Gra06]. xxxi [Ano88k]. XXXII [Che94, Sch90b]. xxxix [Ano92l].

Ya [PP78, VHK16]. Yaglom [Kot79, Row90a]. Yale [Ano74-29]. Yamagata [Won22]. Yamaguchi [CM22, Won22]. Yan [Ano89d, Ano92-31]. Yang [Dum18, Nee79, Yon74]. Yasuaki [Ung20]. Ybbs [Bin90, Bin91]. YBC [Buc23, FR98]. Year [Dau90]. yearbook [Ano90w]. Years [Ano90t, Cal96, Coo96, Dau99, Dum18, Kat07, Ram00a, Row85a, Ste05, Van06, Ano92-33, Bau74, Ben95, Bie74b, DS^{+}02, FFW00, Fer95, Jan11, Kas03, Kul76, Mar90c, Sch92, Scr85c, SGJ05, Bid74]. Years/Context [Dau99]. = Y1500 [iÔ75]. Yesterday [Ano81v]. Ying [Yon74]. Yong [Mar95, Nee79]. Yonkers [Ano74l]. Yorick [Ano81u]. York [Abe09, Ano74q, Ano74u, Ano74-27, Ano74l, Ano74z, Ano74-40, Ano77i, Ano77p, Ano81i, Ano81l, Ano84h, Ano84j, Ano85h, Ano87j, Ano87g, Ano87k, Ano87f, Ano88g, Ano88j, Ano89c, Ano89e, Ano89l, Ano91f, Ano92o, Ano93g, Ano93i, Ano95k, Ano95l, Arc96, Bar74, Bay11, Bel12, Ber74, Bie74a, Bir77, Bir97, Bla76, Boy74, Bra07, Buc77, BT84, Bun76a, Cal83, Cal92, Cas77, Cas80, Cha76, Clu76, Cog12, Coh82, Dru91, Dru07, Dru12, Duc16, Edw77a, Edw77b, Enr80, Fol15, Fre92, Gin83, Gra94a, Gra13, GG90b, GG05c, Gra91b, Gra98a, Gra09b, Gri82, Hoc96, Hod77, Hog85b, Høy09b, Hym78, Jon77c, Jon85, Kat92, Kin86b, Knu74, Lei79, Lev97, Lew98, Mel06, Ols76, Ped75b, Pin76, Pla76, Plo06b, Pys81, Ric82, Rid88, Row83b, Row85a]. York [Row89, Row07, Sag91, San09, Sch16, Sch12, Scr07, Sen84, Sen91, Sid13, Spi76, Swe79, Tan74, Tro83, Vek75, Wat09, Wis76, Wol91, Zep12, Ano74w, Edw77a]. York-Oxford [Buc77]. York/Berlin/Heidelberg/London/Paris/Tokyo/Hong [Ano93g]. York/Berlin/Heidelberg/Tokyo [Ano88j]. York/Cambridge/London [BT84]. York/Heidelberg/Berlin [Ano84j, Ano87j, Hog85b]. York/London [Abe09, Enr80, Lei79, Ric82]. York/Melbourne [Hoc96]. York/Oxford [Ano87k]. York/Tokyo [Ano87g]. Young [Ano84h, AH99, Cho92, GG75a]. Youqin [Vol97]. Youschkevitch [Ano79l, Tat75]. Yuan [Ano91g]. Yujian [Ple78]. Yurie [Bos03].

Yushkevich [Ano77l, Ano81o, Gra94b, Gra94c, Gra97a, Ken76d, Thi08, Ano91a, Ano95e, BGM^{+}76, BDG^{+}86, Dem86]. Yusuf [Hog86]. Yvette [Cog12, Taz15]. Yvonne [Van06, Was20, Hog15].

Z [Bul16a, Pec20]. Zīj [Van98]. Zach [War15]. Zacher [Ano77h]. Zahar [Ano81l]. Zahl [H^{+}08, Lév99, Sch11, Kna88]. zahlenlehre [Kit78]. zahlentheoretisches [Edw80]. Zahlentheorie [Ano85e]. Zahlwörter [Lév99]. Zanichelli [Arc17]. Zaniclelli [Pan18]. Zanjani [Sam24]. Zaragoza [Mil93, Fol94]. Zaslavsky [Wil75a]. Zdravkovska [Coo96]. Zealand [Ano81v, Tee82]. Zeichen [Kna88]. Zeit [Nab04, TC05, GG07, Mon83]. Zeitschriftenartikel [HB11, Her14]. zenshu [Ano78-32]. Zentralantiquariat [Bos77]. Zermelo [Daw85, Ebb04, Ebb07, GG79, Moo75b, Moo76, RT81, Dru91]. Zerner [Ano92k]. Zero [NB77]. zeta [Spi76]. Zeuner [Hou87]. Zhang [Shu92, Vol10]. Zhao [Vol97]. Zhigeng [Mar93b]. Zhizn [Dyc75a]. Ziegler [Gra90]. Ziel [Lau92]. Zij [HZ85, Ham80]. Zij-i [HZ85, Ham80]. Zitarelli [Ano88c, Ano88b, Ano89a, Ano89b, Ano92c, Ano92d]. Zloty [Ano82e]. Znanie [Kot79, Lin74]. Zorn [Cam78]. Zum [Fre93, Sau10, Str78, Vol81, Ano89s, Don08, Hof87, Kni01, Tob90, Böl93, Joh84]. Zur [Bie74b, Fol87, Kno83a, Sch17, Sza74, Ano77h, Ano88n, Ano88i, Ano89j, Ano89k, Ano91-32, Ano92-52, Ano94b, Ano95h, Bar82a, Bar82b, Bin89, Bos77, Gab79, GG07, H^{+}08, Høy90b, Høy24a, Ine94, Jay83, Kna88, Lév99, Lüt09a, Mat75, Mue80, Pul05, Pur88, Pur90, Pur91, Row85d, Row06a, Row23, Sch90a, Sch11, Sch82b, Sch90b, SS89, Str81, Tob90, Toe96, Ull03, Vol81, Wue05, vdW80, Eng79, Hof92, Lic97]. Zurich [Dil81, Ric84, Sch16, Sch96, Str81, Tat11, Toe96, vdW80]. Zwei [Row06a, Wue05, Bor18, Yus95]. Zweite [Ano92-52]. zwischen [Arn88b, Coo95, GG07, Str78, TC05, Vol81, Bor18].

References

Albis:1988:NPL

[AA88] Víctor S. Albis and Luis C. Arboleda. Newton’s Principia in Latin America. Historia Mathematica, 15(4):376–379, November 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

         tronic). URL http://www.sciencedirect.com/science/article/
         pii/031508608890033X.

Arcavi:1984:DEM

[AB84] A. Arcavi and M. Bruckheimer. Development and evaluation of materials on the history of mathematics for teachers. Historia Mathematica, 11(2):232, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

         com/science/article/pii/S031508608480021X.

Allaire:2002:SAF

[AB02] Patricia R. Allaire and Robert E. Bradley. Symbolical algebra as a foundation for calculus: D. F. Gregory’s contribution. Historia Mathematica, 29(4):395–426, November 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508600292358X.

Albree:2009:VMM

[AB09] Joe Albree and Scott H. Brown. “A valuable monument of mathematical genius”: the Ladies’ Diary (1704–1840). Historia Mathematica, 36(1):10–47, February 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008000967.

Agnew:2009:GTO

[ABBS09] Alfonso F. Agnew, Alexandru Bobe, Wladimir G. Boskoff, and Bogdan D. Suceava. Gheorghe Titeica and the origins of affine differential geometry. Historia Mathematica, 36(2):161–170, May 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008000694.

Arney:1988:HDM

[ABC^{+}88] David C. Arney, Robert L. Ball, Warren C. Chellman, William E. Ebel, Jr., and Martin L. Vozzo. History of the Department of Mathematics of the United States Military Academy. Historia Mathematica, 15(4):379, November 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086088900341.

Abdulaziz:2008:EMD

[Abd08] Abdulrahman A. Abdulaziz. On the Egyptian method of decomposing 2/n into unit fractions. Historia Mathematica, 35(1):1–18, February 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086007000274.

Abeles:1976:MCB

[Abe76] Francine Abeles. Multiplication in changing bases: a note on Lewis Carroll. Historia Mathematica, 3(2):183–184, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900355.

Abeles:1979:RIN

[Abe79] Francine Abeles. Ranking by inversion: a note on C. L. Dodgson. Historia Mathematica, 6(3):310–317, August 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079901290.

Abeles:1991:SAM

[Abe91] Francine Abeles. Sixteenth annual meeting: Canadian Society for History and Philosophy of Mathematics: University of Victoria, B.C., 31 May–1 June, 1990. Historia Mathematica, 18(2):178–179, May 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609190501N.

Abeles:1993:CDG

[Abe93] Francine F. Abeles. Charles L. Dodgson’s geometric approach to arctangent relations for pi. Historia Mathematica, 20(2):151–159, May 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508608371013X.

Abeles:2001:PPL

[Abe01] Francine F. Abeles, editor. The political pamphlets and letters of Charles Lutwidge Dodgson and related pieces: a mathematical approach, volume 3 of The pamphlets of Lewis Carroll. Lewis Carroll Society of North America, New York, NY, USA, 2001. ISBN 0-930326-14-8. xx + 260 pp. LCCN PR4611.A4 A24 2001.

Abeles:2006:BRC

[Abe06] Francine F. Abeles. Book review: Cogwheels of the Mind. The Story of Venn Diagrams, A. W. F. Edwards. The Johns Hopkins University Press, Baltimore and London (2004), ISBN 0-8018-7434-3. Historia Mathematica, 33(3):374–376, August 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086005000984.

Abeles:2009:BRL

[Abe09] Francine F. Abeles. Book review: Lewis Carroll in Numberland. His Fantastical Mathematical Logical Life, Robin Wilson. Norton, New York/London (2008), ISBN 978-0-393-06027-0. Historia Mathematica, 36(3):287–289, August 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008001080.

Abeles:2010:LPC

[Abe10] Francine F. Abeles, editor. The logic pamphlets of Charles Lutwidge Dodgson and related pieces, volume 4 of The pamphlets of Lewis Carroll. Lewis Carroll Society of North America, New York, NY, USA, 2010. ISBN 0-930326-25-3 (hardcover). xx + 271 pp. LCCN QA9 .D63 2010.

Abhyankar:1975:HSA

[Abh75] Shreeram S. Abhyankar. High-school algebra in algebraic geometry. Historia Mathematica, 2(4):567–572, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075901226.

Acerbi:2011:CDP

[Ace11] Fabio Acerbi. Completing Diophantus, De polygonis numeris, prop. 5. Historia Mathematica, 38(4):548–560, November 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086011000437.

Akritas:1978:FTM

[AD78] Alkiviadis G. Akritas and Stylianos D. Danielopoulos. On the forgotten theorem of Mr. Vincent. Historia Mathematica, 5(4):427–435, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086078902112.

Aballagh:1987:DEM

[AD87] M. Aballagh and A. Djebbar. Découverte d’un écrit mathématique d’al-haṡṡār (XIIe s.): Le livre I du Kāmil. Historia Mathematica, 14(2):147–158, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086087900188.

Aubin:2002:WHD

[AD02] David Aubin and Amy Dahan Dalmedico. Writing the history of dynamical systems and chaos: Longue Durée and revolution, disciplines and cultures. Historia Mathematica, 29(3):273–339, August 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086002923517.

Aleksandrov:1988:BRA

[ADG^{+}88] A. D. Aleksandrov, S. S. Demidov, A. T. Grigoryan, G. P. Matvievskaya, S. S. Sirazhdinov, and A. P. Yushkevich. Boris A. Rozenfeld: On the 70th anniversary of his birth. Historia Mathematica, 15(1):1–8, February 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086088900456.

Agargun:1994:AFF

[AF94] Ahmet G. Agargün and Colin R. Fletcher. Al-Farisi and the Fundamental Theorem of Arithmetic. Historia Mathematica, 21(2):162– 173, May 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086084710159.

Albis-Gonzalez:1977:LAT

[AG77] Victor S. Albis-González. Latin–American translations of Legendre’s “Éléments de Géométrie”. Historia Mathematica, 4(3):339–340, August 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086077900611.

Archibald:2013:E

[AG13] Tom Archibald and Niccolò Guicciardini. Editorial. Historia Mathematica, 40(1):1–2, February 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086012000808.

Archibald:2014:E

[AG14] Tom Archibald and Niccolò Guicciardini. Editorial. Historia Mathematica, 41(1):1–2, February 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508601400007X.

Archibald:2015:E

[AG15] Thomas Archibald and Niccolò Guicciardini. Editorial. Historia Mathematica, 42(2):131, May 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086015000208.

Agazzi:1975:BRH

[Aga75] Evandro Agazzi. Book review: History of mathematical logic from Leibnitz to Peano: By N. I. Styazhkin. Translated from the Russian.

Cambridge, Mass. (MIT Press), 1969. 341 p. Historia Mathematica, 2(3):361–365, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086075900919.

Albis-Gonzalez:1976:WIL

[AGSL76] Victor S. Albis-González and Luis I. Soriano-Lleras. The work of Indalecio Liévano on the foundations of real numbers. Historia Mathematica, 3(2):161–166, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508607690032X.

Ausejo:1999:MIS

[AH99] Elena Ausejo and Mariano Hormigón. Mathematics for independence: From Spanish liberal exile to the young American republics. Historia Mathematica, 26(4):314–326, November 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508609992253X.

Ackerberg-Hastings:2010:BRP

[AH10] Amy Ackerberg-Hastings. Book review: Pioneering Women in American Mathematics: the Pre-1940 PhD’s. By Judy Green and Jeanne LaDuke. Providence, Rhode Island (American Mathematical Society). 2009. ISBN 0-8218-4376-1. xviii + 350 pp. US$79.00. Historia Mathematica, 37(1):124–128, February 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508600900069X.

Ackerberg-Hastings:2016:BRP

[AH16] Amy Ackerberg-Hastings. Book review: Perfect Mechanics: Instrument Makers at the Royal Society of London in the Eighteenth Century, Richard Sorrenson. Docent Press, Boston, Massachusetts (2013), ISBN: 0-9887449-2-9. Historia Mathematica, 43(1):104–105, February 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086015000634.

Ageron:2020:IAB

[AH20] Pierre Ageron and Hmida Hedfi. Ibrahim al-Balisht.ar’s book of arithmetic (ca. 1575): Hybridizing Spanish mathematical treatises with the Arabic scientific tradition. Historia Mathematica, 52(??):26–50, August 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086020300021.

Abgrall:2004:DGI

[Ai04] Philippe Abgrall and Wayjan ibn Rustam Qåuhåi [approximately 1000AD]. Le développement de la géométrie aux IXe–XIe siècles: Abāu Sahl al-Qāuhāi. (French) [The development of geometry in the 9th to 11th Centuries]. Collection Sciences dans l’histoire, 1258-0996. Blanchard, Paris, France, 2004. ISBN 2-85367-221-2. 354 pp. LCCN QA443.5 .A24 2004.

Aissani:1996:MED

[Aïs96] Djamil Aïssani. Le mathématicien Eugène Dewulf (1831–1896) et les manuscrits médiévaux du Maghreb. Historia Mathematica, 23 (3):257–268, August 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086096900270.

Aiton:1986:NRB

[Ait86] E. J. Aiton. Never at rest. A biography of Isaac Newton. Historia Mathematica, 13(1):78–81, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086086902399.

Anderson:2004:SHB

[AKW04] Marlow Anderson, Victor J. Katz, and Robin J. Wilson, editors. Sherlock Holmes in Babylon: and other tales of mathematical history. Mathematical Association of America, Washington, DC, USA, 2004. ISBN 0-88385-546-1. x + 387 pp.

LCCN QA21 .S48 2004. URL http://www.loc.gov/catdir/
enhancements/fy0632/2003113541-d.html;   http://www.loc.
gov/catdir/toc/cam051/2003113541.html.

Alberts:1994:CSM

[Alb94] Gerard Alberts. On connecting socialism and mathematics: Dirk Struik, Jan Burgers, and Jan Tinbergen. Historia Mathematica, 21(3):280–305, August 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086084710263.

Abbott:2010:FEN

[ALB10] Edwin Abbott Abbott, William F. Lindgren, and Thomas Banchoff. Flatland: an edition with notes and commentary. Cambridge University Press, Cambridge, UK; New York, NY, USA, 2010. ISBN 0-521-76988-4 (hardcover), 0-521-75994-3 (paperback). ix + 294 pp. LCCN QA699 .A13 2010.

Alexander:1995:GDH

[Ale95] Daniel S. Alexander. Gaston Darboux and the history of complex dynamics. Historia Mathematica, 22(2):179–185, May 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086085710166.

Alexander:2004:BRT

[Ale04] Daniel S. Alexander. Book review: Through a Reporter’s Eyes, The Life of Stefan Banach: By Roman Kaluza. Translated and edited by Ann Kostant and Wojbor Woyczyński. Boston (Birkhäuser). 1996. 168 pp. $36.95. Historia Mathematica, 31(2):222–226, May 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086003000612.

Alexander:2010:DDH

[Ale10] Amir Alexander. Duel at Dawn: Heroes, Martyrs, and the Rise of Modern Mathematics. Harvard University Press, Cambridge, MA, USA, 2010. ISBN 0-674-04661-7. 307 pp. LCCN QA10.7 .A44 2010.

Alexander:2014:BRF

[Ale14] Daniel S. Alexander. Book review: Fatou, Julia, Montel: the Great Prize of Mathematical Sciences of 1918, and beyond, vol. 2014, Michèlexunbr Audin, in: Lecture Notes in Mathematics, History of Mathematics Subseries. Springer-Verlag, Heidelberg (2011), ISBN 978-3-642-17853-5. Historia Mathematica, 41(2):224–232, May 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086013000232.

Alfonsi:2011:EBM

[Alf11] Liliane Alfonsi. Étienne Bézout (1730–1783): mathématicien des Lumières. Collection Histoire des sciences. Série études. L’Harmattan, Paris, France, 2011. ISBN 2-296-56553-0 (paperback). 414 pp. LCCN QA29.B525 A44 2011.

Allaire:2004:BRC

[All04] Patricia R. Allaire. Book review: Creators of Mathematics: the Irish Connection, Ken Houston (Ed.). University College Dublin Press, Dublin (2000). Historia Mathematica, 31(4):499–501, November 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086004000576.

Allaire:2009:BRH

[All09] Patricia R. Allaire. Book review: Hands on History: a Resource for Teaching Mathematics, Amy Shell-Gellasch (Ed.). The Mathematical Association of America, Washington, DC (2007), ISBN 978-0-88385-182-1. Historia Mathematica, 36(3):291–292, August 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009000044.

Almkvist:1987:CC

[Alm87] Gert Almkvist. Correspondence of Carleman. Historia Mathematica, 14(3):284, August 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086087900486.

Ambrosetti:2008:LAI

[Amb08] Nadia Ambrosetti. L’eredità arabo-islamica nelle scienze e nelle arti del calcolo dell’Europa medievale. (Italian) [The Arab–Islamic heritage in science and arts of calculating in Medieval Europe]. Studi e ricerche. LED, Milano, Italy, 2008. ISBN 88-7916-388-4. 407 pp. LCCN QA303.2 .A52 2008.

Andonie:1975:MSV

[And75] George St Andonie. Matematica si viata. Historia Mathematica, 2(3):356–358, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086075900889.

Andersen:1983:MTF

[And83] Kirsti Andersen. The mathematical technique in Fermat’s deduction of the law of refraction. Historia Mathematica, 10(1):48–62, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900320.

Anderson:2007:BRM

[And07] Marlow Anderson. Book review: Musings of the Masters: an Anthology of Mathematical Reflections, Raymond G. Ayoub (Ed.). The Mathematical Association of America, Washington, DC (2004), ISBN 0-88385-549-6. Historia Mathematica, 34(3):356–358, August 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086006000966.

Andersen:2011:OBA

[And11] Kirsti Andersen. One of Berkeley’s arguments on compensating errors in the calculus. Historia Mathematica, 38(2):219–231, May 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086010000510.

Anellis:1987:MLS

[Ane87] Irving H. Anellis. Mathematical logic in the Soviet Union, 1917–1980. Historia Mathematica, 14(3):285–287, August 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086087900498.

Anellis:1988:SUP

[Ane88] Irving H. Anellis. Some unpublished papers of Jean van Heijenoort. Historia Mathematica, 15(3):270–274, August 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900675.

Angelelli:1979:BRD

[Ang79a] Ignacio Angelelli. Book review: Dear Russell–Dear Jourdain. A commentary on Russell’s logic, based on his correspondence with Philip Jourdian: By I. Grattan-Guinness. London (Duckworth). 1977. 234 pp. Historia Mathematica, 6(4):464–466, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079900417.

Angelelli:1979:COC

[Ang79b] Ignacio Angelelli. “Class as one” and “class as many” before modern set theory. Historia Mathematica, 6(3):305–309, August 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079901289.

Angelelli:1988:BRL

[Ang88] Ignacio Angelelli. Book review: Life and works of Giuseppe Peano: Edited by Hubert Kennedy. Studies in the History of Modern Science 4. Dordrecht (Reidel). 1980. $34.00 cloth, $14 paper. Historia Mathematica, 15(3):291–294, August 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900791.

Anonymous:1974:Aa

[Ano74a] Anonymous. Abstracts. Historia Mathematica, 1(1):113–126, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074901967.

Anonymous:1974:Ab

[Ano74b] Anonymous. Abstracts. Historia Mathematica, 1(2):219–252, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074900305.

Anonymous:1974:Ac

[Ano74c] Anonymous. Abstracts. Historia Mathematica, 1(3):351–386, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074900871.

Anonymous:1974:Ad

[Ano74d] Anonymous. Abstracts. Historia Mathematica, 1(4):481–510, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074900561.

Anonymous:1974:ADH

[Ano74e] Anonymous. Advanced degrees in history of mathematics. Historia Mathematica, 1(3):326, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074900767.

Anonymous:1974:AAA

[Ano74f] Anonymous. American Academy of Arts and Sciences workshop on the historical evolution of mathematics. Historia Mathematica, 1 (4):461–462, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074900470.

Anonymous:1974:AMHa

[Ano74g] Anonymous. American mathematical heritage symposium on history of statistics and probability, 27–29 may 1974. Historia Mathematica, 1 (2):186, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074900123.

Anonymous:1974:AMHb

[Ano74h] Anonymous. American mathematical heritage symposium on the history of statistics and probability. Historia Mathematica, 1(4): 458–460, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074900469.

Anonymous:1974:ACH

[Ano74i] Anonymous. Archive of Contemporary History, University of Wyoming. Historia Mathematica, 1(3):324, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL

http://www.sciencedirect.com/science/article/pii/0315086074900731.

Anonymous:1974:AMJ

[Ano74j] Anonymous. Archives of mathematics journals. Historia Mathematica, 1(4):448, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074900366.

Anonymous:1974:BEN

[Ano74k] Anonymous. Biography of Emmy Noether. Historia Mathematica, 1(3):324, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074900743.

Anonymous:1974:BRF

[Ano74l] Anonymous. Book review: OR → CJ. The foundations and mathematical models of operations research with extensions to the criminal justice system: By Haig Edward Bohigan. Yonkers, New York (Gazette Press), 1971. 305 pp. Historia Mathematica, 1(2):199– 200, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074900202.

Anonymous:1974:BRCb

[Ano74m] Anonymous. Book review: A concise history of science in India: Edited by D. M. Bose, S. N. Sen, and B. V. Subbarayappa. New Delhi (Indian National Science Academy), 1971. 689 pp. Historia Mathematica, 1(2):208–212, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074900263.

Anonymous:1974:BRHb

[Ano74n] Anonymous. Book review: A handbook of terms used in algebra and analysis: Compiled by A. G. Howson. Cambridge (Cambridge University Press). 1972. 247 p. $14.50 (cloth), $4.95 (paper). Historia Mathematica, 1(2):214, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074900287.

Anonymous:1974:BRBa

[Ano74o] Anonymous. Book review: Bibliography and research manual of the history of mathematics: By Kenneth O. May. Toronto (University of Toronto Press), 1973. 827 pp. Can. $20. Historia Mathematica, 1(2):192–194, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074900172.

Anonymous:1974:BRBb

[Ano74p] Anonymous. Book review: Bolyai, sau aventura geometriilor neeuclidiene [(Romanian) [Bolyai, non-Euclidean geometries and adventure]]: By Florica T. Campan. Cluj (Editura Albatros), 1971. 139 pp. lei 6,75. Historia Mathematica, 1(2):194, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074900184.

Anonymous:1974:BRCc

[Ano74q] Anonymous. Book review: Carl Friedrich Gauss: Werke: Edited by the Königliche Gesellschaft der Wissenschaften zu Göttingen, 1863– 1933. Reprint by Georg Olms Verlag, Hildesheim and New York, 1973. 12 vols. 7874 p. Figs., tables. DM 1316. Historia Mathematica, 1 (4):468–470, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074900512.

Anonymous:1974:BRCa

[Ano74r] Anonymous. Book review: Creative teaching: Heritage of R. L. Moore: By Reginald D. Traylor with William Bane and Madeline Jones. Houston, Texas (University of Houston), 1972. 478 pp. $8.50. Historia Mathematica, 1(2):188–192, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074900160.

Anonymous:1974:BRDa

[Ano74s] Anonymous. Book review: Demotic mathematical papyri: By Richard A. Parker. Providence, Rhode Island (Brown University Press, Brown Egyptological Studies VII) and London (Lund Humphries), 1972. 100 pp., 27 plates. Cloth £10.00. Historia Mathematica, 1(2):195– 199, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074900196.

Anonymous:1974:BRDc

[Ano74t] Anonymous. Book review: Development of mathematical logic: By R. L. Goodstein. London (Logos Press), 1971. 157 p. Historia Mathematica, 1(2):212–214, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074900275.

Anonymous:1974:BRDb

[Ano74u] Anonymous. Book review: Development of modern mathematics: By J. M. Dubbey. New York (Crane, Russak, & Co.), 1973 [reprint of English edition: London (Butterworth & Co.), 1970]. 145 pp. $6.25. Historia Mathematica, 1(2):202–204, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074900238.

Anonymous:1974:BRHa

[Ano74v] Anonymous. Book review: Die hilbertische probleme: Edited by P. S. Alexandrov. German edition translated by H. Bernhardt and W. Purkert and edited by H. Wussing. Leipzig (Akademische Verlagsgesellschaft Geest & Portig K.-G.), 1971. 306 p. Historia Mathematica, 1(2):207–208, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074900251.

Anonymous:1974:BRG

[Ano74w] Anonymous. Book review: Göttingen and New York: a film in color (43 minutes) produced by the Mathematical Association of America, 1966. (Rental from Modern Film Rentals, 2323 New Hyde Park Road, New Hyde Park, N.Y. 11040. Purchase from Modern Learning Aids, P.O. Box 302, Rochester, N.Y. 14603.). Historia Mathematica, 1(2): 202, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074900226.

Anonymous:1974:BRI

[Ano74x] Anonymous. Book review: Invisible colleges. Diffusion of knowledge in scientific communities: By Diana Crane. University of Chicago Press, 1972. 223 p. US $9. Historia Mathematica, 1(4):477–479, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074900548.

Anonymous:1974:BRMa

[Ano74y] Anonymous. Book review: Mathematics in the time of the pharaohs: By Richard J. Gillings. Cambridge, Mass. (M.I.T. Press), 1972. 297 p. US $25. Historia Mathematica, 1(4):463–468, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074900500.

Anonymous:1974:BRS

[Ano74z] Anonymous. Book review: Scientific instruments of the Seventeenth and Eighteenth Centuries: By Maurice Daumas. Translated and edited by Mary Holbrook. New York (Praeger), 1972. 361 pp., plates, bibliography. US $38.50. Historia Mathematica, 1(4):476–477, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074900536.

Anonymous:1974:BRE

[Ano74-27] Anonymous. Book review: The Euler–Mayer correspondence (1751– 1755): a new perspective on Eighteenth-Century advances in the lunar theory: By Eric C. Forbes. New York (American Elsvier Publishing Company, Inc.), 1971. 125 pp., index. Historia Mathematica, 1(2):200–202, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074900214.

Anonymous:1974:BRMb

[Ano74-28] Anonymous. Book review: The mathematical career of Pierre de Fermat 1601–1665: By Michael Sean Mahoney. Princeton University Press, 1973. 427 p. US $20. Historia Mathematica, 1(4):470–476, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074900524.

Anonymous:1974:BRP

[Ano74-29] Anonymous. Book review: The paradox of the liar: Edited by R. L. Martin. New Haven and London (Yale University Press), 1970. 149 p. Historia Mathematica, 1(2):215–218, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074900299.

Anonymous:1974:BRT

[Ano74-30] Anonymous. Book review: Time and method: an essay on the methodology of research: By Ferdinand Gonseth. Translated by Eva H. Guggenheimer. Springfield, Illinois (Charles C. Thomas), 1972. 453 pp. Historia Mathematica, 1(4):479–480, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508607490055X.

Anonymous:1974:CC

[Ano74-31] Anonymous. The Canadian connection. Historia Mathematica, 1(3): 253, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074900597.

Anonymous:1974:CSHa

[Ano74-32] Anonymous. Canadian society for the history and philosophy of Mathematics/Société canadienne d’Histoire et de philosophie des Mathématiques. Historia Mathematica, 1(1):78, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074901694.

Anonymous:1974:CSHb

[Ano74-33] Anonymous. Canadian society for the history and philosophy of mathematics/Société canadienne d’histoire et de philosophie des mathématiques. Historia Mathematica, 1(3):327–332, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074900779.

Anonymous:1974:CPN

[Ano74-34] Anonymous. The Charles S. Peirce newsletter. Historia Mathematica, 1(1):27, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074901645.

Anonymous:1974:CT

[Ano74-35] Anonymous. Congratulations to the thousand! Historia Mathematica, 1(1):1–2, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074901608.

Anonymous:1974:DP

[Ano74-36] Anonymous. Dissertation in progress. Historia Mathematica, 1 (4):455, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074900421.

Anonymous:1974:DU

[Ano74-37] Anonymous. Dissertations in the USSR 1973/1974. Historia Mathematica, 1(4):455, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508607490041X.

Anonymous:1974:DT

[Ano74-38] Anonymous. Doctoral theses. Historia Mathematica, 1(1):90, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074901797.

Anonymous:1974:Ec

[Ano74-39] Anonymous. Education. Historia Mathematica, 1(4):462, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074900494.

Anonymous:1974:EUU

[Ano74-40] Anonymous. The Egyptian use of unit fractions for equitable distribution: a communication from Gerald R. Rising, State University of New York at Buffalo. Historia Mathematica, 1(1):93–94, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074901840.

Anonymous:1974:ESR

[Ano74-41] Anonymous. English summary of the review. Historia Mathematica, 1(1):112, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074901955.

Anonymous:1974:Ea

[Ano74-42] Anonymous. Errata. Historia Mathematica, 1(3):254, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074900603.

Anonymous:1974:Eb

[Ano74-43] Anonymous. Errata. Historia Mathematica, 1(3):254, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074900615.

Anonymous:1974:HUA

[Ano74-44] Anonymous. Harvard University archives. Historia Mathematica, 1(1):86, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074901761.

Anonymous:1974:HUP

[Ano74-45] Anonymous. Harvard University Press Prize to I. B. Cohen. Historia Mathematica, 1(1):90, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074901803.

Anonymous:1974:HMA

[Ano74-46] Anonymous. History at the Mathematical ASSN. Conference Royal Holloway College, Egham, Surrey, England 8–11 April 1994. Historia Mathematica, 1(1):84, February 1974. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074901748.

Anonymous:1974:HMJ

[Ano74-47] Anonymous. History of mathematics at the 3–5 June 1974 Toronto meeting of the Canadian society for the history and philosophy of science. Historia Mathematica, 1(1):78, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074901700.

Anonymous:1974:HMT

[Ano74-48] Anonymous. History of mathematics at Tokyo 19–26 August 1974. Historia Mathematica, 1(3):262, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074900640.

Anonymous:1974:HMU

[Ano74-49] Anonymous. History of mathematics in the USA. Historia Mathematica, 1(1):78, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074901712.

Anonymous:1974:HMM

[Ano74-50] Anonymous. History of mathematics meetings in India. Historia Mathematica, 1(3):262, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074900639.

Anonymous:1974:HSS

[Ano74-51] Anonymous. History of Science Society Meeting San Francisco 27–29 Dec. 1973. Historia Mathematica, 1(1):12, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074901621.

Anonymous:1974:HS

[Ano74-52] Anonymous. History of semiotics. Historia Mathematica, 1(1):90, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074901827.

Anonymous:1974:IAA

[Ano74-53] Anonymous. Index of authors of articles and reviews in volume I. Historia Mathematica, 1(4):511, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074900573.

Anonymous:1974:IAP

[Ano74-54] Anonymous. Index of authors of publications reviewed and abstracted in volume I. Historia Mathematica, 1(4):512–516, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074900585.

Anonymous:1974:IAH

[Ano74-55] Anonymous. Instructional AIDS for the history of mathematics. Historia Mathematica, 1(2):184, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508607490010X.

Anonymous:1974:JCW

[Ano74-56] Anonymous. Jourdain’s copy of Whitehead and Russell’s Principia Mathematica: a communication from I. Grattan-Guinness, Middlesex Polytechnic at Enfield, England. Historia Mathematica, 1(2): 187, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074900147.

Anonymous:1974:MM

[Ano74-57] Anonymous. La matemática o las matemáticas? Historia Mathematica, 1(4):457–458, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074900457.

Anonymous:1974:LSU

[Ano74-58] Anonymous. Lecture series at the university of Chicago. Historia Mathematica, 1(3):316, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074900718.

Anonymous:1974:LIG

[Ano74-59] Anonymous. A letter from Ivor Grattan-Guinness, Middlesex Polytechnic, Enfield, England. Historia Mathematica, 1(4):456, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074900445.

Anonymous:1974:MES

[Ano74-60] Anonymous. Mathematics education in socialist countries. Historia Mathematica, 1(4):455, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074900408.

Anonymous:1974:MA

[Ano74-61] Anonymous. Meetings in Argentina. Historia Mathematica, 1(3): 262, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074900652.

Anonymous:1974:MP

[Ano74-62] Anonymous. Minkowski papers. Historia Mathematica, 1(1):90, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074901815.

Anonymous:1974:NB

[Ano74-63] Anonymous. News from Brazil. Historia Mathematica, 1(3):290, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508607490069X.

Anonymous:1974:NTH

[Ano74-64] Anonymous. Newsletter on teaching history of mathematics. Historia Mathematica, 1(1):27–28, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074901657.

Anonymous:1974:ON

[Ano74-65] Anonymous. Obituary notices. Historia Mathematica, 1(2):187, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074900159.

Anonymous:1974:PBK

[Ano74-66] Anonymous. Phi Beta kappa award. Historia Mathematica, 1(2): 180, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508607490007X.

Anonymous:1974:PAH

[Ano74-67] Anonymous. Photographic archive of historians of science at Indiana University. Historia Mathematica, 1(1):86, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074901773.

Anonymous:1974:Pb

[Ano74-68] Anonymous. Projects. Historia Mathematica, 1(2):186, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074900135.

Anonymous:1974:Pa

[Ano74-69] Anonymous. Promotion. Historia Mathematica, 1(2):180, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074900068.

Anonymous:1974:QMA

[Ano74-70] Anonymous. Questions on mathematics in Africa. Historia Mathematica, 1(4):454, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074900391.

Anonymous:1974:QHM

[Ano74-71] Anonymous. A quote on history of mathematics courses. Historia Mathematica, 1(2):150, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074900032.

Anonymous:1974:R

[Ano74-72] Anonymous. Recordings. Historia Mathematica, 1(4):448, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074900378.

Anonymous:1974:RRJ

[Ano74-73] Anonymous. Response by R. J. Gillings, Sydney, Australia. Historia Mathematica, 1(1):94, February 1974. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074901852.

Anonymous:1974:SWM

[Ano74-74] Anonymous. Should we be mathematicians, historians of science, historians, or generalists? Historia Mathematica, 1(2):127–128, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074900019.

Anonymous:1974:UIP

[Ano74-75] Anonymous. University of Indiana photographic archive of historians of science. Historia Mathematica, 1(4):448, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074900354.

Anonymous:1974:XIC

[Ano74-76] Anonymous. XIVth International congress of the history of science Tokyo and Kyoto, 19–27 August 1974. Historia Mathematica, 1(4):462, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074900482.

Anonymous:1975:Aa

[Ano75a] Anonymous. Abstracts. Historia Mathematica, 2(1):102–126, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086075900622.

Anonymous:1975:Ab

[Ano75b] Anonymous. Abstracts. Historia Mathematica, 2(2):224–252, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075901603.

Anonymous:1975:Ac

[Ano75c] Anonymous. Abstracts. Historia Mathematica, 2(3):376–424, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900956.

Anonymous:1975:Ad

[Ano75d] Anonymous. Announcements. Historia Mathematica, 2(4):624, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086075901366.

Anonymous:1975:BRM

[Ano75e] Anonymous. Bertrand Russell Memorial Logic Conference: Essay prize. Historia Mathematica, 2(3):320, August 1975. CODEN HI-MADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0315086075900749.

Anonymous:1975:CSH

[Ano75f] Anonymous. Canadian society for the history and philosophy of mathematics société canadienne d’histoire et de philosophie des mathématiques. Historia Mathematica, 2(1):75–76, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900476.

Anonymous:1975:CHM

[Ano75g] Anonymous. Conference on the history of mathematics. Historia Mathematica, 2(1):83, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075900506.

Anonymous:1975:C

[Ano75h] Anonymous. Correction. Historia Mathematica, 2(3):271, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900646.

Anonymous:1975:CS

[Ano75i] Anonymous. CSHPM/SCHPM. Historia Mathematica, 2(4):617, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086075901317.

Anonymous:1975:CSC

[Ano75j] Anonymous. CSHPM/SCHPM correction! Historia Mathematica, 2 (2):199, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086075901512.

Anonymous:1975:CSM

[Ano75k] Anonymous. CSHPM/SCHPM Meeting in Edmonton 6–7 June 1975. Historia Mathematica, 2(3):336–338, August 1975. CODEN HI-MADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0315086075900798.

Anonymous:1975:Ea

[Ano75l] Anonymous. Education. Historia Mathematica, 2(2):192–193, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075901470.

Anonymous:1975:Eb

[Ano75m] Anonymous. Erratum. Historia Mathematica, 2(4):617, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075901329.

Anonymous:1975:IAA

[Ano75n] Anonymous. Index of authors of articles and reviews in volume II. Historia Mathematica, 2(4):619, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075901342.

Anonymous:1975:IAP

[Ano75o] Anonymous. Index of authors of publications reviewed or abstracted in volume I. Historia Mathematica, 2(4):620–624, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075901354.

Anonymous:1975:JAS

[Ano75p] Anonymous. Joint Atlantic Seminar on the History of the Physical Sciences New Haven, Connecticut, March 21–22, 1975. Historia Mathematica, 2(1):82, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508607590049X.

Anonymous:1975:MSP

[Ano75q] Anonymous. ‘Mathematics’ — singular or plural?: a note from Lewis Pyenson, Université de Montréal. Historia Mathematica, 2(2):197– 199, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086075901494.

Anonymous:1975:Ma

[Ano75r] Anonymous. Meetings. Historia Mathematica, 2(2):??, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075901457.

Anonymous:1975:Mb

[Ano75s] Anonymous. Meetings. Historia Mathematica, 2(3):318–320, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900737.

Anonymous:1975:NN

[Ano75t] Anonymous. News and notices. Historia Mathematica, 2(1):30, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086075900336.

Anonymous:1975:O

[Ano75u] Anonymous. Obituaries. Historia Mathematica, 2(2):152, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075901408.

Anonymous:1975:ORC

[Ano75v] Anonymous. Opening remarks for the conference by Garrett Birkhoff and I. B. Cohen. Historia Mathematica, 2(4):433–436, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900993.

Anonymous:1975:PAE

[Ano75w] Anonymous. Papers of Albert Einstein. Historia Mathematica, 2(1): 73, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086075900452.

Anonymous:1975:PJN

[Ano75x] Anonymous. Papers of John von Neumann. Historia Mathematica, 2(1):73, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086075900440.

Anonymous:1975:PW

[Ano75y] Anonymous. Participants in the workshop. Historia Mathematica, 2(4):431–432, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086075900981.

Anonymous:1975:S

[Ano75z] Anonymous. Sources. Historia Mathematica, 2(3):272, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900658.

Anonymous:1975:T

[Ano75-27] Anonymous. Thanks. Historia Mathematica, 2(4):617, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075901330.

Anonymous:1976:Aa

[Ano76a] Anonymous. Abstracts. Historia Mathematica, 3(1):110–132, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086076900276.

Anonymous:1976:Ab

[Ano76b] Anonymous. Abstracts. Historia Mathematica, 3(2):244–258, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900616.

Anonymous:1976:Ac

[Ano76c] Anonymous. Abstracts. Historia Mathematica, 3(3):362–384, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076901221.

Anonymous:1976:Ad

[Ano76d] Anonymous. Abstracts. Historia Mathematica, 3(4):500–508, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086076900938.

Anonymous:1976:AMH

[Ano76e] Anonymous. American mathematical heritage symposium on the history of algebra at the University of Texas at El Paso, November 15–16 1975. Historia Mathematica, 3(2):220, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086076900495.

Anonymous:1976:AMC

[Ano76f] Anonymous. Annual meeting of the Canadian society for the history and philosophy of mathematics. Historia Mathematica, 3(3):334–336, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086076901105.

Anonymous:1976:AWM

[Ano76g] Anonymous. Association for Women in Mathematics. Historia Mathematica, 3(3):324, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076901026.

Anonymous:1976:BPM

[Ano76h] Anonymous. Bicentennial program at the MAA meeting in San Antonio, Texas 24–26 January 1976. Historia Mathematica, 3(1):85, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508607690015X.

Anonymous:1976:CPB

[Ano76i] Anonymous. C. S. Peirce bicentennial international congress June 16–20, 1976. Historia Mathematica, 3(4):416, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900707.

Anonymous:1976:CPN

[Ano76j] Anonymous. C. S. Pierce National Monument. Historia Mathematica, 3(3):327, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508607690104X.

Anonymous:1976:CCB

[Ano76k] Anonymous. Chronology of Carl B. Boyer. Historia Mathematica, 3(4):395–396, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076900641.

Anonymous:1976:C

[Ano76l] Anonymous. Correspondence. Historia Mathematica, 3(4):440, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086076900732.

Anonymous:1976:CSa

[Ano76m] Anonymous. CSHPM/SCHPM. Historia Mathematica, 3(1):49, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086076900057.

Anonymous:1976:CSb

[Ano76n] Anonymous. CSHPM/SCHPM. Historia Mathematica, 3(2):202, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900379.

Anonymous:1976:CHP

[Ano76o] Anonymous. A current history project suggestion. Historia Mathematica, 3(1):20, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076900033.

Anonymous:1976:EQP

[Ano76p] Anonymous. Editorial: Questions of policy. Historia Mathematica, 3(1):1–4, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508607690001X.

Anonymous:1976:ER

[Ano76q] Anonymous. Editorial: Reorganization. Historia Mathematica, 3(4):385, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076900628.

Anonymous:1976:Ea

[Ano76r] Anonymous. Education. Historia Mathematica, 3(1):86, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900161.

Anonymous:1976:Eb

[Ano76s] Anonymous. Education. Historia Mathematica, 3(3):333, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076901099.

Anonymous:1976:Ec

[Ano76t] Anonymous. Education. Historia Mathematica, 3(4):473–476, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086076900811.

Anonymous:1976:HBI

[Ano76u] Anonymous. History of biometry at the 9th International Biometric Conference Boston, 22–27 August 1976. Historia Mathematica, 3(4):402, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076900677.

Anonymous:1976:HMT

[Ano76v] Anonymous. History of mathematics at teachers’ conferences in England Easter 1975. Historia Mathematica, 3(2):160, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900318.

Anonymous:1976:HSSa

[Ano76w] Anonymous. History of Science Society to honor Carl Boyer, Philadelphia, 28 December 1976. Historia Mathematica, 3(4):396, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900653.

Anonymous:1976:HSSb

[Ano76x] Anonymous. History of statistics at the 40th session of the International Statistical Institute Warsaw 1–9 September 1975. Historia Mathematica, 3(4):416, November 1976. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086076900690.

Anonymous:1976:IDN

[Ano76y] Anonymous. Imposition of decimal numeration on primitive tribes. Historia Mathematica, 3(3):296, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086076900999.

Anonymous:1976:IAA

[Ano76z] Anonymous. Index of authors of articles and reviews in volume III. Historia Mathematica, 3(4):509, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508607690094X.

Anonymous:1976:IAP

[Ano76-27] Anonymous. Index of authors of publications reviewed and abstracted in volume III. Historia Mathematica, 3(4):510–514, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086076900951.

Anonymous:1976:ISC

[Ano76-28] Anonymous. International Scientific Conference on Al-Farabi and the Development of Science and Culture in the East Alma Ata, 11–13 September 1975. Historia Mathematica, 3(2):220, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900501.

Anonymous:1976:MMS

[Ano76-29] Anonymous. Meeting on “Mathematics in a social context” at Regensburg 2–3 July 1976. Historia Mathematica, 3(4):470–471, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086076900793.

Anonymous:1976:M

[Ano76-30] Anonymous. Meetings. Historia Mathematica, 3(3):337–338, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076901117.

Anonymous:1976:NNa

[Ano76-31] Anonymous. News and notices. Historia Mathematica, 3(2):212, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086076900422.

Anonymous:1976:NNb

[Ano76-32] Anonymous. News and notices. Historia Mathematica, 3(3):339– 340, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086076901129.

Anonymous:1976:NNc

[Ano76-33] Anonymous. News and notices. Historia Mathematica, 3(4): 479–480, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076900835.

Anonymous:1976:NAC

[Ano76-34] Anonymous. Ninth annual conference of the Leningrad section of the Soviet national society for the history and philosophy of science and technology, 4–5 February 1976. Historia Mathematica, 3(2):219, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900483.

Anonymous:1976:PID

[Ano76-35] Anonymous. Person index for Delambre’s Rapport Historique of 1810. Historia Mathematica, 3(3):321–324, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076901014.

Anonymous:1976:PHB

[Ano76-36] Anonymous. Project for historical bibliography. Historia Mathematica, 3(4):440, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076900720.

Anonymous:1976:Pa

[Ano76-37] Anonymous. Projects. Historia Mathematica, 3(1):76, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900112.

Anonymous:1976:Pb

[Ano76-38] Anonymous. Projects. Historia Mathematica, 3(2):216, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900458.

Anonymous:1976:Pc

[Ano76-39] Anonymous. Projects. Historia Mathematica, 3(3):332, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076901087.

Anonymous:1976:Ra

[Ano76-40] Anonymous. Reviews. Historia Mathematica, 3(1):87, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900173.

Anonymous:1976:Rb

[Ano76-41] Anonymous. Reviews. Historia Mathematica, 3(3):341, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076901130.

Anonymous:1976:Sa

[Ano76-42] Anonymous. Sources. Historia Mathematica, 3(3):331, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076901075.

Anonymous:1976:Sb

[Ano76-43] Anonymous. Sources. Historia Mathematica, 3(4):472, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508607690080X.

Anonymous:1976:DJS

[Ano76-44] Anonymous. To Dirk J. Struik at eighty-two. Historia Mathematica, 3(2):133, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076900288.

Anonymous:1976:UM

[Ano76-45] Anonymous. Upcoming meeting. Historia Mathematica, 3(2): 182, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086076900343.

Anonymous:1977:ABR

[Ano77a] Anonymous. 150th anniversary of the birth of Riemann. Historia Mathematica, 4(1):6, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086077900301.

Anonymous:1977:AEKa

[Ano77b] Anonymous. Abstracts: Edited by Kenneth O. May. Historia Mathematica, 4(1):115–126, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086077900544.

Anonymous:1977:AEKb

[Ano77c] Anonymous. Abstracts: Edited by Kenneth O. May. Historia Mathematica, 4(3):371–378, August 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508607790074X.

Anonymous:1977:AMH

[Ano77d] Anonymous. American Mathematical Heritage Symposium on the History of Applied Mathematics at the University of Texas at Arlington, 2–3 October 1976. Historia Mathematica, 4(1):6, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900295.

Anonymous:1977:ACC

[Ano77e] Anonymous. Aryabhata celebrations at Calcutta. Historia Mathematica, 4(3):347–350, August 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086077900659.

Anonymous:1977:BRSd

[Ano77f] Anonymous. Book review: A source book in classical analysis: Edited by Garrett Birkhoff with the assistance of Uta Merzbach. Cambridge, Mass. (Harvard University Press). 1973. Historia Mathematica, 4 (3):364–365, August 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086077900726.

Anonymous:1977:BRSc

[Ano77g] Anonymous. Book review: A survey of the Almagest. Acta historica scientiarum naturalium et medicinalium, vol. 30: By Olaf Pedersen. Copenhagen (Odense University Press). 1974. 454 p. Dan. kr. 200. Historia Mathematica, 4(1):110–113, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900520.

Anonymous:1977:BRHb

[Ano77h] Anonymous. Book review: Die hauptschriften zur dyadik von G. W. Leibniz: By Hans J. Zacher. Frankfurt/M. (Vittorio Klostermann). 1973. 392 p. Historia Mathematica, 4(1):114, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900532.

Anonymous:1977:BRD

[Ano77i] Anonymous. Book review: Diocles On Burning Mirrors. The Arabic translation of the lost Greek original: Edited, with English translation and commentary by G. J. Toomer. Berlin, Heidelberg, and New York (Springer-Verlag). 1976. ix + 249 p. 24 plates. Historia Mathematica, 4(3):360–363, August 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086077900714.

Anonymous:1977:BRG

[Ano77j] Anonymous. Book review: G. W. Leibniz marginalia in Newtoni principia mathematica: Édition critique en langue allemande par E. A. Fellmann. Traduction française de J. F. Courtine. Collection des travaux de l’Académie internationale d’histoire des sciences n^{o}. Paris (Librairie Philosophique J. Vrin). 1973. 128 pp. Historia Mathematica, 4(1):104–105, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086077900489.

Anonymous:1977:BRHa

[Ano77k] Anonymous. Book review: Histoire des mathématiques: By Jean-Paul Collette. Montréal (Éditions du Renouveau Pédagogique). 1973. 238 p. Historia Mathematica, 4(1):105–106, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900490.

Anonymous:1977:BRIb

[Ano77l] Anonymous. Book review: Istoriya matematiki s drevneishikh vremen do nachala XIX stoletiya. Tom Pervyi: S drevneishikh vremen do nachala novogo vremeni: Edited by A. P. Yushkevich. Moscow (Nauka). 1970. 351 p. Historia Mathematica, 4(1):108–109, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900519.

Anonymous:1977:BRL

[Ano77m] Anonymous. Book review: Leonhard Euler Briefwechsel. Opera omnia. Series quarta A: Commercium epistolicum, vol. I: Herausgeber: A. P. Jushkevich, V. Smirnov, W. Habicht. Basel (Birkhäuser). 1975. xvii + 666 pp. 1 plate. $60. Historia Mathematica, 4(1):97–100, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086077900441.

Anonymous:1977:BRMa

[Ano77n] Anonymous. Book review: Message d’un Mathématicien: Henri Lebesgue: By Lucienne Félix. Paris (Albert Blanchard). 1974. 259 p. 65f. Historia Mathematica, 4(1):100–102, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900453.

Anonymous:1977:BRSb

[Ano77o] Anonymous. Book review: Sets and integration. An outline of the development: By D. van Dalen and A. F. Monna. Groningen (Wolters — Nordhoff). 1972. 162 p. H. F1. 43.30. Historia Mathematica, 4(1):103, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086077900477.

Anonymous:1977:BRSa

[Ano77p] Anonymous. Book review: Stability of motion: Edited by A. T. Fuller. London (Taylor and Francis), New York and Toronto (John Wiley). 1975. Historia Mathematica, 4(1):102–103, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900465.

Anonymous:1977:BRIa

[Ano77q] Anonymous. Book review: The Italian Renaissance of mathematics: By Paul Lawrence Rose. Geneva (Librairie Droz). 1975. xiii + 316 pp. Historia Mathematica, 4(1):106–107, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900507.

Anonymous:1977:BRMb

[Ano77r] Anonymous. Book review: The method of analysis. Its geometrical origin and its general significance: By Hintikka J. & Remes U. Dordrecht-Holland/Boston (D. Reidel Publishing Co.). Boston Studies in the Philosophy of Science vol. XXV/Synthese Library vol. LXXV. 1974. 144 p. Historia Mathematica, 4(3):366–369, August 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900738.

Anonymous:1977:BRN

[Ano77s] Anonymous. Book review: The new elements of mathematics: Volume 1: Arithmetic. Charles S. Peirce. Edited by Carolyn Eisele. The Hague (Mouton Publishers)/Atlantic Highlands, N.J. (Humanities Press). 1976. xl + 250 p. $29.50. Historia Mathematica, 4(3):355– 360, August 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086077900702.

Anonymous:1977:CSM

[Ano77t] Anonymous. Canadian society meeting and Gauss symposium. Historia Mathematica, 4(2):155, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086077901082.

Anonymous:1977:C

[Ano77u] Anonymous. Correspondence. Historia Mathematica, 4(4):453–454, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086077900842.

Anonymous:1977:Ec

[Ano77v] Anonymous. Editorial. Historia Mathematica, 4(3):349–350, August 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900660.

Anonymous:1977:Ea

[Ano77w] Anonymous. Education. Historia Mathematica, 4(1):66, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900374.

Anonymous:1977:Eb

[Ano77x] Anonymous. Education. Historia Mathematica, 4(3):341–343, August 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900623.

Anonymous:1977:Ed

[Ano77y] Anonymous. Education. Historia Mathematica, 4(4):456–457, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086077900878.

Anonymous:1977:GBC

[Ano77z] Anonymous. Gauss bicentennial celebration at Göttingen. Historia Mathematica, 4(2):155, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086077901070.

Anonymous:1977:GBG

[Ano77-27] Anonymous. Gauss bicentennial in Göttingen. Historia Mathematica, 4(3):294, August 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086077900568.

Anonymous:1977:GEF

[Ano77-28] Anonymous. Gauss exhibit in the FRG. Historia Mathematica, 4(3):337–338, August 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508607790060X.

Anonymous:1977:GSO

[Ano77-29] Anonymous. Gauss Symposium in Ottawa, 30 April 1977. Historia Mathematica, 4(1):36, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086077900337.

Anonymous:1977:HMW

[Ano77-30] Anonymous. History of mathematics at the Wheaton College Conference “A Christian Perspective on the Foundations of Mathematics” 28–30 April 1977. Historia Mathematica, 4(4):452, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900830.

Anonymous:1977:IAA

[Ano77-31] Anonymous. Index of authors of articles and reviews in volume IV. Historia Mathematica, 4(4):499, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508607790101X.

Anonymous:1977:IAP

[Ano77-32] Anonymous. Index of authors of publications reviewed and abstracted in volume IV. Historia Mathematica, 4(4):500–504, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086077901021.

Anonymous:1977:M

[Ano77-33] Anonymous. Meetings. Historia Mathematica, 4(4):458, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900891.

Anonymous:1977:NNa

[Ano77-34] Anonymous. News & notices. Historia Mathematica, 4(1):42, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086077900350.

Anonymous:1977:NNb

[Ano77-35] Anonymous. News and notices. Historia Mathematica, 4(1):95–96, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508607790043X.

Anonymous:1977:NNc

[Ano77-36] Anonymous. News and notices. Historia Mathematica, 4(3):352, August 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086077900684.

Anonymous:1977:NNd

[Ano77-37] Anonymous. News and notices. Historia Mathematica, 4(4):457, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508607790088X.

Anonymous:1977:Pa

[Ano77-38] Anonymous. Projects. Historia Mathematica, 4(1):90, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900404.

Anonymous:1977:Pb

[Ano77-39] Anonymous. Projects. Historia Mathematica, 4(3):351–352, August 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900672.

Anonymous:1977:Pc

[Ano77-40] Anonymous. Projects. Historia Mathematica, 4(4):454, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900854.

Anonymous:1977:S

[Ano77-41] Anonymous. Sources. Historia Mathematica, 4(4):455, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900866.

Anonymous:1977:TIL

[Ano77-42] Anonymous. Third International Leibniz Congress. Historia Mathematica, 4(2):204, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086077901124.

Anonymous:1978:ADN

[Ano78a] Anonymous. The 250th anniversary of the death of Newton in Belgrade. Historia Mathematica, 5(1):90, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078901374.

Anonymous:1978:Aa

[Ano78b] Anonymous. Abstracts. Historia Mathematica, 5(1):114–126, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086078901556.

Anonymous:1978:Ac

[Ano78c] Anonymous. Acknowledgement. Historia Mathematica, 5(4):495, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508607890232X.

Anonymous:1978:Ab

[Ano78d] Anonymous. Addendum. Historia Mathematica, 5(2):226–227, May 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078900551.

Anonymous:1978:APA

[Ano78e] Anonymous. AFIPS to publish Annals of the History of Computing. Historia Mathematica, 5(4):472–473, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078902227.

Anonymous:1978:BAC

[Ano78f] Anonymous. Boncompagni Archive at Cornell University. Historia Mathematica, 5(4):470–472, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078902203.

Anonymous:1978:BRE

[Ano78g] Anonymous. Book review: Evolution of mathematical concepts: an elementary study: By Raymond L. Wilder. Paperback reprinting of the 1968 original. London (Transworld). 1974. 216 pp. Historia Mathematica, 5(1):99–105, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078901507.

Anonymous:1978:BRG

[Ano78h] Anonymous. Book review: Galileo, two new sciences: By Stillman Drake. Including Centers of Gravity and Force of Percussion. A new translation with introduction and notes. Madison, Wisconsin (The University of Wisconsin Press). 1974. xxvii + 323 pp. Historia Mathematica, 5(1):107–108, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078901532.

Anonymous:1978:BRH

[Ano78i] Anonymous. Book review: Histoire comparée des numérations écrites: By Geneviève Guitel. Paris (Flammarion). 1975. 857 pp. 48 drawings. 50 plates. 73 tables. Historia Mathematica, 5(1):108–113, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086078901544.

Anonymous:1978:BRV

[Ano78j] Anonymous. Book review: Vorlesungen über neuere geometrie: By M. Pasch, with an appendix by M. Dehn. Berlin (Grundlehren der math. Wissenschaften, Springer-Verlag). 1926. Reprint 1976 (Springer-Verlag). viii + 275 pp. Historia Mathematica, 5(1):107, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086078901520.

Anonymous:1978:CSH

[Ano78k] Anonymous. Canadian society for the history and philosophy of Mathematics/Société canadienne d’histoire et de philosophie des mathématiques. Historia Mathematica, 5(2):220, May 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078900538.

Anonymous:1978:CSA

[Ano78l] Anonymous. Contemporary Scientific Archives Centre, Oxford, England. Historia Mathematica, 5(3):274, August 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078901131.

Anonymous:1978:EBa

[Ano78m] Anonymous. Editorial Board. Historia Mathematica, 5(1):??, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086078901301.

Anonymous:1978:Ea

[Ano78n] Anonymous. Education. Historia Mathematica, 5(1):76, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078901350.

Anonymous:1978:Eb

[Ano78o] Anonymous. Education. Historia Mathematica, 5(1):93, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078901441.

Anonymous:1978:Ec

[Ano78p] Anonymous. Education. Historia Mathematica, 5(3):306, August 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078901155.

Anonymous:1978:FNS

[Ano78q] Anonymous. Five new series in the history of Arabic–Islamic science. Historia Mathematica, 5(2):234, May 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078900605.

Anonymous:1978:GBC

[Ano78r] Anonymous. Gauss bicentennial celebrations at Belgrade. Historia Mathematica, 5(1):90, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078901386.

Anonymous:1978:GL

[Ano78s] Anonymous. Grecheskaya logistika. Historia Mathematica, 5(1): 105–107, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086078901519.

Anonymous:1978:HMA

[Ano78t] Anonymous. History at the meetings of the American Mathematical Society and Mathematical Association of America, Atlanta, January 1978. Historia Mathematica, 5(2):138–166, May 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078900472.

Anonymous:1978:HC

[Ano78u] Anonymous. History of combinatorics. Historia Mathematica, 5(2): 227, May 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086078900563.

Anonymous:1978:IAA

[Ano78v] Anonymous. Index of authors of articles and reviews in volume 5. Historia Mathematica, 5(4):496–500, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078902331.

Anonymous:1978:IHAb

[Ano78w] Anonymous. Institute for the History of Arabic Science. Historia Mathematica, 5(4):472, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078902215.

Anonymous:1978:IHAa

[Ano78x] Anonymous. Institute for the History of Arabic Science, University of Aleppo, Aleppo, Syria. Historia Mathematica, 5(1):95, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078901477.

Anonymous:1978:MRW

[Ano78y] Anonymous. “Mathematics and the real world” a Workshop of the Danish Mathematical Society at Roskilde University Center, Denmark 8–9 August 1978. Historia Mathematica, 5(2):232, May 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078900587.

Anonymous:1978:M

[Ano78z] Anonymous. Meetings. Historia Mathematica, 5(1):94, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078901453.

Anonymous:1978:MUM

[Ano78-27] Anonymous. Monographs in undergraduate mathematics. Historia Mathematica, 5(2):233, May 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078900599.

Anonymous:1978:NSJ

[Ano78-28] Anonymous. New Spanish journal of history of science. Historia Mathematica, 5(1):95, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078901465.

Anonymous:1978:NNS

[Ano78-29] Anonymous. Nihonjin no sugaku. (Japanese) [Japanese School of Residence]. Historia Mathematica, 5(1):99, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078901490.

Anonymous:1978:P

[Ano78-30] Anonymous. Projects. Historia Mathematica, 5(4):469–470, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086078902197.

Anonymous:1978:PNJ

[Ano78-31] Anonymous. Publication of new journal beginning in 1976. Historia Mathematica, 5(2):204, May 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078900502.

Anonymous:1978:STZ

[Ano78-32] Anonymous. Seki Takakazu zenshu: Takakazu Seki’s collected works: Edited and explained by A. Hirayama, K. Shimodaira, and H. Hirose. Osaka (Osaka Kyoiku Tosho). 1974. xviii + 852 pp. Historia Mathematica, 5(1):96–98, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078901489.

Anonymous:1978:S

[Ano78-33] Anonymous. Sources. Historia Mathematica, 5(1):93, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508607890143X.

Anonymous:1978:SSH

[Ano78-34] Anonymous. Spanish society for the history of sciences. Historia Mathematica, 5(1):90, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078901398.

Anonymous:1979:Aa

[Ano79a] Anonymous. Abstracts. Historia Mathematica, 6(1):98–105, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079901216.

Anonymous:1979:AEA

[Ano79b] Anonymous. Abstracts: Edited by Albert C. Lewis. Historia Mathematica, 6(3):338–345, August 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079901411.

Anonymous:1979:Ab

[Ano79c] Anonymous. Acknowledgment. Historia Mathematica, 6(4):485, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086079900478.

Anonymous:1979:ACC

[Ano79d] Anonymous. AFIPS and CBI: a computer history program. Historia Mathematica, 6(4):459, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079900399.

Anonymous:1979:AHC

[Ano79e] Anonymous. Annals of the history of computing. Historia Mathematica, 6(1):72–73, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086079901125.

Anonymous:1979:ALA

[Ano79f] Anonymous. Archive for linear algebra and its applications. Historia Mathematica, 6(2):206–207, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079900909.

Anonymous:1979:ASM

[Ano79g] Anonymous. Archive of the school mathematics study group. Historia Mathematica, 6(2):207–208, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079900910.

Anonymous:1979:BPJ

[Ano79h] Anonymous. Bibliography of Philip Jourdain’s works. Historia Mathematica, 6(3):322–323, August 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079901332.

Anonymous:1979:BPL

[Ano79i] Anonymous. Boncompagni papers in the Library of Trinity College, Dublin. Historia Mathematica, 6(2):209–212, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079900922.

Anonymous:1979:BRG

[Ano79j] Anonymous. Book review: Girolamo Cardano: (1501–1576). By Markus Fierz. Basel/Stuttgart (Birkhaüser). 1977. 140 pp. Historia Mathematica, 6(3):332–334, August 1979. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079901381.

Anonymous:1979:BRH

[Ano79k] Anonymous. Book review: History of the department of mathematics of the University of Kansas, 1866–1970: By G. Baley Price. Lawrence, Kansas (Kansas University Endowment Association, The University of Kansas). 1976. xii + 788 pp. (reproduced typescript). Historia Mathematica, 6(3):334–336, August 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079901393.

Anonymous:1979:BRM

[Ano79l] Anonymous. Book review: Les mathematiques arabes (VIII^{e}–XV^{e} siècles): By Adolf P. Youschkevitch. Translated by M. Cazenave and K. Jaouiche, with preface By R. Taton. Paris (Vrin). 1976. Historia Mathematica, 6(3):337, August 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508607990140X.

Anonymous:1979:BRR

[Ano79m] Anonymous. Book review: Richard Dedekind et les fondements des mathe’matiques: By Pierre Dugac. Préface de Jean Dieudonné. Paris (J. Vrin). 1976. 334. pp. Historia Mathematica, 6(3):328–332, August 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508607990137X.

Anonymous:1979:BAH

[Ano79n] Anonymous. British association for the history of mathematics. Historia Mathematica, 6(2):205–206, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079900880.

Anonymous:1979:CHP

[Ano79o] Anonymous. Colloquium in honor of Professor Dr. Kurt Vogel. Historia Mathematica, 6(2):205, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079900879.

Anonymous:1979:CLS

[Ano79p] Anonymous. Conference of the Leningrad Section of the Soviet Association of the History and Philosophy of Science and Technology. Historia Mathematica, 6(1):67–70, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079901083.

Anonymous:1979:CSAb

[Ano79q] Anonymous. Contemporary scientific archives centre. Historia Mathematica, 6(3):326–327, August 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079901368.

Anonymous:1979:CSAa

[Ano79r] Anonymous. Contemporary Scientific Archives Centre, Oxford, England. Historia Mathematica, 6(1):62, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508607990106X.

Anonymous:1979:CP

[Ano79s] Anonymous. Cube puzzle. Historia Mathematica, 6(2):195–196, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079900788.

Anonymous:1979:DCE

[Ano79t] Anonymous. Dissertation completed: From ENIAC to UNIVAC. Historia Mathematica, 6(2):201, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079900818.

Anonymous:1979:DCJ

[Ano79u] Anonymous. Dissertation completed: Jaina astronomy. Historia Mathematica, 6(1):73–74, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079901149.

Anonymous:1979:DCH

[Ano79v] Anonymous. Dissertation completed: the history of linear optimalization. Historia Mathematica, 6(1):74, February 1979.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079901150.

Anonymous:1979:DPH

[Ano79w] Anonymous. Dissertation in progress: Hensel’s p-adic numbers. Historia Mathematica, 6(1):74, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079901162.

Anonymous:1979:DPR

[Ano79x] Anonymous. Dissertation in progress: Richard Dedekind. Historia Mathematica, 6(2):202, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508607990082X.

Anonymous:1979:HPL

[Ano79y] Anonymous. History and philosophy of logic. Historia Mathematica, 6(3):325–326, August 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086079901356.

Anonymous:1979:HMM

[Ano79z] Anonymous. History of mathematics at Macquarie University. Historia Mathematica, 6(2):206, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079900892.

Anonymous:1979:IAA

[Ano79-27] Anonymous. Index of authors of articles and reviews in volume 6. Historia Mathematica, 6(4):486, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508607990048X.

Anonymous:1979:IAP

[Ano79-28] Anonymous. Index of authors of publications reviewed and abstracted in volume 6. Historia Mathematica, 6(4):487–490, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079900491.

Anonymous:1979:ICS

[Ano79-29] Anonymous. International Christoffel symposium. Historia Mathematica, 6(3):322, August 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086079901320.

Anonymous:1979:ISG

[Ano79-30] Anonymous. International study group on the relation between the history and pedagogy of mathematics. Historia Mathematica, 6(2): 205, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086079900867.

Anonymous:1979:ISL

[Ano79-31] Anonymous. International Symposium on the Life and Work of Christiaan Huygens 1629–1695: Amsterdam 22–25 August 1979. Historia Mathematica, 6(2):204, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079900855.

Anonymous:1979:MCW

[Ano79-32] Anonymous. Mathematics in context — a workshop. Historia Mathematica, 6(1):71–72, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079901113.

Anonymous:1979:NJH

[Ano79-33] Anonymous. New journal in the history of mathematics. Historia Mathematica, 6(2):202–203, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079900831.

Anonymous:1979:NMJ

[Ano79-34] Anonymous. New mathematics journal to accept historical articles. Historia Mathematica, 6(4):458–459, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079900387.

Anonymous:1979:NCa

[Ano79-35] Anonymous. Notes on contributors. Historia Mathematica, 6 (1):106–107, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086079901228.

Anonymous:1979:NCb

[Ano79-36] Anonymous. Notes on contributors. Historia Mathematica, 6 (3):346–348, August 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086079901423.

Anonymous:1979:OHP

[Ano79-37] Anonymous. Oral history program in mathematics. Historia Mathematica, 6(2):203–204, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086079900843.

Anonymous:1979:PAFa

[Ano79-38] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 6(2):231, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086079900971.

Anonymous:1979:PAFb

[Ano79-39] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 6(3):349, August 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086079901435.

Anonymous:1979:PAFc

[Ano79-40] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 6(4):484, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079900466.

Anonymous:1979:PV

[Ano79-41] Anonymous. Peirce at Vienna in 1979. Historia Mathematica, 6(1):70–71, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086079901095.

Anonymous:1979:PDR

[Ano79-42] Anonymous. Professor Dr. Rer. Nat. Habil. Kurt-Reinhard Biermann: On the occasion of his 60th birthday. Historia Mathematica, 6(4):??, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086079900223.

Anonymous:1979:PHM

[Ano79-43] Anonymous. Project on the history of mathematics in Canada. Historia Mathematica, 6(1):74–75, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079901174.

Anonymous:1979:PJM

[Ano79-44] Anonymous. Publications of James mills Peirce. Historia Mathematica, 6(2):201, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086079900806.

Anonymous:1979:SCH

[Ano79-45] Anonymous. Seminar on component of history of mathematics in mathematics curriculum. Historia Mathematica, 6(1):71, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079901101.

Anonymous:1979:SHM

[Ano79-46] Anonymous. Symposium on the history of mathematical methods. Historia Mathematica, 6(4):457–458, November 1979. CODEN HI-MADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0315086079900375.

Anonymous:1979:WSH

[Ano79-47] Anonymous. Workshop on the social history of mathematics. Historia Mathematica, 6(4):453–457, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079900363.

Anonymous:1979:WDH

[Ano79-48] Anonymous. World directory of the history of mathematics. Historia Mathematica, 6(1):73, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079901137.

Anonymous:1980:A

[Ano80a] Anonymous. Acknowledgment. Historia Mathematica, 7(4):469, November 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086080900191.

Anonymous:1980:BRG

[Ano80b] Anonymous. Book review: G. W. Leibniz, Sämtliche Schriften und Briefe. Historia Mathematica, 7(4):448, November 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900105.

Anonymous:1980:BSH

[Ano80c] Anonymous. British Society for the History of Mathematics. Historia Mathematica, 7(2):193–194, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900439.

Anonymous:1980:CDS

[Ano80d] Anonymous. Cahiers du séminaire d’histoire des mathématiques de Paris. Historia Mathematica, 7(4):447, November 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900099.

Anonymous:1980:DTT

[Ano80e] Anonymous. Dissertation on a treatise on the theory of conic sections. Historia Mathematica, 7(2):195–196, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900452.

Anonymous:1980:DCG

[Ano80f] Anonymous. Dissertation on contributions of Gyula. Historia Mathematica, 7(2):196, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086080900464.

Anonymous:1980:DPD

[Ano80g] Anonymous. Dissertation on partial differentiation from Leibniz to Euler. Historia Mathematica, 7(1):90, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900786.

Anonymous:1980:DAV

[Ano80h] Anonymous. Dissertations on Arabic versions of Euclid’s Elements. Historia Mathematica, 7(2):195, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900440.

Anonymous:1980:ECB

[Ano80i] Anonymous. Edition of Charpit’s book review: Mémoire of 1784. Historia Mathematica, 7(1):89, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900750.

Anonymous:1980:E

[Ano80j] Anonymous. Erratum. Historia Mathematica, 7(4):468, November 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608090018X.

Anonymous:1980:FIC

[Ano80k] Anonymous. Fourth International Congress on Mathematical Education. Historia Mathematica, 7(2):193, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900427.

Anonymous:1980:HLC

[Ano80l] Anonymous. History of logic conference. Historia Mathematica, 7(2):191–192, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086080900403.

Anonymous:1980:MHH

[Ano80m] Anonymous. In memoriam: [Helmut Haase (1898–1979)]. Historia Mathematica, 7(2):197, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900488.

Anonymous:1980:IAA

[Ano80n] Anonymous. Index of authors of articles and reviews involume 7. Historia Mathematica, 7(4):473, November 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900221.

Anonymous:1980:IAP

[Ano80o] Anonymous. Index of authors of publications reviewed and abstracted in volume 7. Historia Mathematica, 7(4):474–477, November 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900233.

Anonymous:1980:ISG

[Ano80p] Anonymous. International Study Group on the Relations between the History and the Pedagogy of Mathematics: Activities at I.C.M.E. IV, August 10–16, 1980. Historia Mathematica, 7(1):80–81, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900695.

Anonymous:1980:IBT

[Ano80q] Anonymous. Ivor Bulmer-Thomas: Honorary doctor of science. Historia Mathematica, 7(1):86–88, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900737.

Anonymous:1980:KMF

[Ano80r] Anonymous. The Kenneth O. May Fellowship in the History of Mathematics, University of Toronto. Historia Mathematica, 7(1):86, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086080900725.

Anonymous:1980:MM

[Ano80s] Anonymous. Materia mathematica. Historia Mathematica, 7(1):89, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086080900762.

Anonymous:1980:NHM

[Ano80t] Anonymous. A new history of mathematics journal. Historia Mathematica, 7(2):197, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508608090049X.

Anonymous:1980:NJH

[Ano80u] Anonymous. New journal for the history of mathematics. Historia Mathematica, 7(1):88, February 1980. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900749.

Anonymous:1980:N

[Ano80v] Anonymous. Note. Historia Mathematica, 7(1):89, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900774.

Anonymous:1980:NCa

[Ano80w] Anonymous. Notes on contributors. Historia Mathematica, 7(1): 108–110, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086080900877.

Anonymous:1980:NCb

[Ano80x] Anonymous. Notes on contributors. Historia Mathematica, 7(2):225– 227, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086080900579.

Anonymous:1980:NCc

[Ano80y] Anonymous. Notes on contributors. Historia Mathematica, 7(3):366, August 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086080900294.

Anonymous:1980:NCd

[Ano80z] Anonymous. Notes on contributors. Historia Mathematica, 7(4): 470–471, November 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086080900208.

Anonymous:1980:PAFa

[Ano80-27] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 7(1):111, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900889.

Anonymous:1980:PAFb

[Ano80-28] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 7(2):228, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086080900580.

Anonymous:1980:PAFc

[Ano80-29] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 7(3):367, August 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086080900300.

Anonymous:1980:PAFd

[Ano80-30] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 7(4):472, November 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508608090021X.

Anonymous:1980:RPR

[Ano80-31] Anonymous. Research project “On the relationship between the process of science and the process of education — described by the example of the development of mathematics in the 19th century”. Historia Mathematica, 7(1):90–91, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900798.

Anonymous:1980:SFH

[Ano80-32] Anonymous. Société Française d’Histoire des sciences et des techniques. Historia Mathematica, 7(2):198, May 1980. CODEN HI-MADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0315086080900506.

Anonymous:1980:SAW

[Ano80-33] Anonymous. Study of American women in mathematics. Historia Mathematica, 7(2):196, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900476.

Anonymous:1980:TTD

[Ano80-34] Anonymous. Thesis: a treatise devoted to the theory of conic sections. Historia Mathematica, 7(1):91, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900804.

Anonymous:1980:UMH

[Ano80-35] Anonymous. An unpublished manuscript of Hans Rademacher. Historia Mathematica, 7(4):445–446, November 1980. CODEN HI-MADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0315086080900087.

Anonymous:1980:VBM

[Ano80-36] Anonymous. Viggo Brun in memoriam (1885–1978). Historia Mathematica, 7(1):??, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086080900592.

Anonymous:1980:WGQ

[Ano80-37] Anonymous. A workshop on the growth of quantum mechanics. Historia Mathematica, 7(1):79–80, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900683.

Anonymous:1980:WDH

[Ano80-38] Anonymous. World directory of the history of mathematics. Historia Mathematica, 7(2):198, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900518.

Anonymous:1980:XIC

[Ano80-39] Anonymous. XVIth international Congress of the history of science. Historia Mathematica, 7(2):192–193, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900415.

Anonymous:1981:MHM

[Ano81a] Anonymous. A. I. Markushevich as a historian of mathematics. Historia Mathematica, 8(2):??, May 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086081900240.

Anonymous:1981:AEAa

[Ano81b] Anonymous. Abstracts: Edited by Albert C. Lewis. Historia Mathematica, 8(1):110–118, February 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086081900203.

Anonymous:1981:AEAb

[Ano81c] Anonymous. Abstracts: Edited by Albert C. Lewis. Historia Mathematica, 8(2):220–228, May 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086081900458.

Anonymous:1981:A

[Ano81d] Anonymous. Acknowledgment. Historia Mathematica, 8(4):497, November 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086081900641.

Anonymous:1981:BRBf

[Ano81e] Anonymous. Book review: Chapters in the recent history of mathematics: Special issue of Nieuw Archief voor Wiskunde on the occasion of the bicentennial celebration of the Wiskundig Genootschap, 1778–1978. Amsterdam, Mathematisch Centrum, 1978, vi + 230 pp. illus. Historia Mathematica, 8(1):97–99, February 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608190015X.

Anonymous:1981:BRC

[Ano81f] Anonymous. Book review: Da Cardano a Galois: Momenti di storia dell’Algebra: By Silvio Maracchia. Milano (Feltrinelli), 1979. 240 pp. £6500. Historia Mathematica, 8(2):199–200, May 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900379.

Anonymous:1981:BRBa

[Ano81g] Anonymous. Book review: Didaktik der mathematik: Herausgegeben von Hans-Georg Steiner. (Wege der Forschung, Band CC-CLXI.) Darmstadt (Wissenschaftliche Buchgesellschaft). 1978. xlviii + 478 pp. Historia Mathematica, 8(1):81–85, February 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900094.

Anonymous:1981:BRF

[Ano81h] Anonymous. Book review: Frantsuzskaya shkola teorii funktsiĭ i mnozhestv, na rubezhe XIX–XX VV: [The French School of the Theory of Functions and Sets at the Turn of the XIX–XX Centuries].

By F. A. Medvedev. Moscow (Nauka). 1976. 231 pp. Historia Mathematica, 8(2):208–211, May 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086081900410.

Anonymous:1981:BRGb

[Ano81i] Anonymous. Book review: Geschichte Der Elementarmathematik: Band 1: Arithmetik und Algebra, 4th ed. By Johannes Tropfke. Edited by Kurt Vogel, Karin Reich, Helmuth Gericke. Berlin/New York; Walter de Gruyter, 1979. xiv + 742 pp. DM 184. Historia Mathematica, 8(1):106–109, February 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086081900197.

Anonymous:1981:BRH

[Ano81j] Anonymous. Book review: Histoire du theoreme des accroissements finis: By Pierre Dugac. Paris (Université Pierre et Marie Curie). 1979. 60 pp., reproduced typescript. Historia Mathematica, 8(2):214– 219, May 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086081900434.

Anonymous:1981:BRBd

[Ano81k] Anonymous. Book review: History of analysis: Edited by R. J. Stanton and R. O. Wells, Jr. Houston, Texas (Rice University), 1977. Rice University Studies, Volume 64, Numbers 2 and 3. iii + 228 pp. No price stated. Historia Mathematica, 8(1):92–93, February 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900124.

Anonymous:1981:BRI

[Ano81l] Anonymous. Book review: Imre Lakatos, proofs and refutations. The logic of mathematical discovery: Edited by John Worrall and Elie Zahar. Cambridge University Press, New York and London. 1976. xii + 154 pp. + bibliography + name and subject indexes. Cloth $22.50; paper, $5.25. Historia Mathematica, 8(1):99–104, February 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900161.

Anonymous:1981:BRT

[Ano81m] Anonymous. Book review: La théorie générale de l’équilibre et du mouvement des systems: By Louis Poinsot. Edition critique et commentarires par Patrice Bailhache J. Vrin 1975. Historia Mathematica, 8(2):200–202, May 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086081900380.

Anonymous:1981:BRL

[Ano81n] Anonymous. Book review: Limite, point d’accumulation, compact: By Pierre Dugac. Paris (Université Pierre et Marie Curie). 1980. 80 pp., reproduced typescript. Historia Mathematica, 8(2):214–219, May 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900446.

Anonymous:1981:BRM

[Ano81o] Anonymous. Book review: Matematika XIX veka. Matematicheskaya logika. Algebra. Teoriya chisel. Teoriya veroyatnostei: Edited by A. N. Kolmogorov and A. P. Yushkevich. Moscow (Nauka). 1978. 155 pp. 2 roubles, 40 kopecks. Historia Mathematica, 8(2):212–213, May 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900422.

Anonymous:1981:BRBe

[Ano81p] Anonymous. Book review: Proceedings of the First International Symposium for the History of Arabic Science, April 5–12, 1976. Volume II: Papers in European Languages: Edited by A. Y. al-Hassan, G. Karmi, and N. Namnum. Aleppo (Institute for the History of Arabic Science); 1978. 365 pp. Historia Mathematica, 8(1):95–97, February 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900148.

Anonymous:1981:BRSb

[Ano81q] Anonymous. Book review: Socialist mathematics education: Edited by Frank J. Swetz. 1978. Burgundy Press, Southampton, PA. Historia Mathematica, 8(1):106, February 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086081900185.

Anonymous:1981:BRSa

[Ano81r] Anonymous. Book review: Storia della teoria delle equazioni algebrice: By Raffaela Franci and Laura Toti Rigetelli. U. Mussia, Milan, 1979. 159 pp. Historia Mathematica, 8(1):104–105, February 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900173.

Anonymous:1981:BRBb

[Ano81s] Anonymous. Book review: The bicentennial tribute to American mathematics, 1776–1976: Edited by Dalton Tarwater. The Mathematical Association of America. 1977. viii + 225 pp. Historia Mathematica, 8(1):85–90, February 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086081900100.

Anonymous:1981:BRBc

[Ano81t] Anonymous. Book review: Vito Vol’terra (Vito Volterra): By E. M. Polishchuk. Leningrad (Izdatel’stvo Nauka, Leningradskoe Otdelenie). 1977. 116 pp. Historia Mathematica, 8(1):90–92, February 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900112.

Anonymous:1981:BRW

[Ano81u] Anonymous. Book review: Wittgenstein’s lectures on the foundations of mathematics: (Cambridge, 1939). From the notes of R. G. Bosanquet, Norman Malcolm, Rush Rhees, and Yorick Smythies. Edited by Cora Diamond. Ithaca, N.Y. (Cornell University Press). 1976. 300 pp. Historia Mathematica, 8(2):205–207, May 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086081900409.

Anonymous:1981:BRGa

[Ano81v] Anonymous. Book review: George E. Andrews: Partitions: Yesterday and today: New Zealand Mathematical Society. 1979. 56 pp. Historia Mathematica, 8(1):94–95, February 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086081900136.

Anonymous:1981:CSP

[Ano81w] Anonymous. Centro studi per la matematica medioevale. Historia Mathematica, 8(2):198, May 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086081900367.

Anonymous:1981:BPA

[Ano81x] Anonymous. Die Berliner und die Petersburger Akademie der Wissenschaften im Briefwechsel Leonhard Eulers, Teil III. Historia Mathematica, 8(2):202–205, May 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086081900392.

Anonymous:1981:DPD

[Ano81y] Anonymous. Dissertation in progress: David Hilbert. Historia Mathematica, 8(2):196, May 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086081900331.

Anonymous:1981:DDB

[Ano81z] Anonymous. Doctoral dissertation: Bertrand Russell and the paradoxes of set theory. Historia Mathematica, 8(2):196, May 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900343.

Anonymous:1981:Ea

[Ano81-27] Anonymous. Education. Historia Mathematica, 8(1):61–66, February 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086081900069.

Anonymous:1981:Eb

[Ano81-28] Anonymous. Erratum. Historia Mathematica, 8(1):119, February 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900215.

Anonymous:1981:Ec

[Ano81-29] Anonymous. Erratum. Historia Mathematica, 8(2):229, May 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608190046X.

Anonymous:1981:ECM

[Ano81-30] Anonymous. Ettore Carruccio — the man and the scientist. Historia Mathematica, 8(3):??, August 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086081900677.

Anonymous:1981:FAC

[Ano81-31] Anonymous. The first Australian conference on the history of mathematics. Historia Mathematica, 8(2):189–190, May 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900288.

Anonymous:1981:HMR

[Ano81-32] Anonymous. History of mathematics as a resource for mathematics teaching: a seminar. Historia Mathematica, 8(3):361, August 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900744.

Anonymous:1981:IAA

[Ano81-33] Anonymous. Index of authors of articles and reviews in volume 8. Historia Mathematica, 8(4):498, November 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086081900653.

Anonymous:1981:IAP

[Ano81-34] Anonymous. Index of authors of publications reviewed and abstracted in volume 8. Historia Mathematica, 8(4):499–502, November 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900665.

Anonymous:1981:IGB

[Ano81-35] Anonymous. Isabella Grigoryevna Bashmakova on the occasion of her 60th birthday. Historia Mathematica, 8(4):??, November 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900495.

Anonymous:1981:NJH

[Ano81-36] Anonymous. A new journal in the history of mathematics. Historia Mathematica, 8(3):363, August 1981. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086081900768.

Anonymous:1981:NNa

[Ano81-37] Anonymous. News and notices. Historia Mathematica, 8(1):80, February 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086081900082.

Anonymous:1981:NN

[Ano81-38] Anonymous. News and notices. Historia Mathematica, 8(2):195, May 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508608190032X.

Anonymous:1981:NCa

[Ano81-39] Anonymous. Notes on contributors. Historia Mathematica, 8 (1):120–123, February 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086081900227.

Anonymous:1981:NCb

[Ano81-40] Anonymous. Notes on contributors. Historia Mathematica, 8(2): 230–232, May 1981. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/0315086081900471.

Anonymous:1981:PAFa

[Ano81-41] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 8(1):124, February 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086081900239.

Anonymous:1981:PAFb

[Ano81-42] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 8(2):233, May 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086081900483.

Anonymous:1981:PAFc

[Ano81-43] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 8(3):387, August 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086081900823.

Anonymous:1981:PAFd

[Ano81-44] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 8(4):496, November 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508608190063X.

Anonymous:1981:WDH

[Ano81-45] Anonymous. World directory of historians of mathematics. Historia Mathematica, 8(3):362, August 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086081900756.

Anonymous:1981:WWI

[Ano81-46] Anonymous. World War II history planned. Historia Mathematica, 8 (2):197, May 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086081900355.

Anonymous:1982:Aa

[Ano82a] Anonymous. Abstracts. Historia Mathematica, 9(2):253–259, May 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082900210.

Anonymous:1982:Ab

[Ano82b] Anonymous. Acknowledgment. Historia Mathematica, 9(4):507, November 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086082901185.

Anonymous:1982:BREa

[Ano82c] Anonymous. Book review: Epistemological and social problems of the sciences in the early Nineteenth Century: Edited by H. N. Jahnke and M. Otte. Boston (Reidel). 1981. 420 pp. Historia Mathematica, 9(2):239–245, May 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086082900179.

Anonymous:1982:BRI

[Ano82d] Anonymous. Book review: It seems I am a jew: By Grigori Freiman. Translated, edited, and with an introduction by Melvyn B. Nathanson, Carbondale (Southern Illinois University Press). 1980. 97 pp. $9.95. Historia Mathematica, 9(2):245–247, May 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082900180.

Anonymous:1982:BRP

[Ano82e] Anonymous. Book review: Prolegomena alle “vite dei matematici” di Bernardino Baldi (1587–1596): (Accademia Polacca delle Scienze. Biblioteca e Centro di Studi a Roma: Conferenze e Studi 71). By Bronislaw Bilinski. Wroclaw, 1977. 135 pp. Zloty 38. Historia Mathematica, 9(2):251–252, May 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086082900209.

Anonymous:1982:BREb

[Ano82f] Anonymous. Book review: The emergence of number: By John N. Crossley. Victoria, Australia (Upside Down A Book Company). 1980. 376 pp. Paper. A $14.50 + A $2.50 p. & P. Historia Mathematica, 9(2):247–251, May 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086082900192.

Anonymous:1982:DP

[Ano82g] Anonymous. Dissertation in progress. Historia Mathematica, 9 (4):488, November 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086082901124.

Anonymous:1982:Ea

[Ano82h] Anonymous. Education. Historia Mathematica, 9(2):201–205, May 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082900052.

Anonymous:1982:Eb

[Ano82i] Anonymous. Erratum. Historia Mathematica, 9(4):486, November 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082901100.

Anonymous:1982:FSH

[Ano82j] Anonymous. The first symposium on the history of mathematics in China. Historia Mathematica, 9(2):223–225, May 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608290012X.

Anonymous:1982:HMX

[Ano82k] Anonymous. History of mathematics at the XVIth International Congress of the History of Science, Bucharest, Romania, August 26–September 3, 1981. Historia Mathematica, 9(2):206, May 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082900064.

Anonymous:1982:HMP

[Ano82l] Anonymous. History of mathematics papers presented at the XVIth International Congress of the History of Science, Bucharest, Romania, 1891. Historia Mathematica, 9(2):216–220, May 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082900106.

Anonymous:1982:IAA

[Ano82m] Anonymous. Index of authors of articles and reviews in volume 9. Historia Mathematica, 9(4):508, November 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086082901197.

Anonymous:1982:IAP

[Ano82n] Anonymous. Index of authors of publications reviewed and abstracted in volume 9. Historia Mathematica, 9(4):509–512, November 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082901203.

Anonymous:1982:ICF

[Ano82o] Anonymous. Israel colloquium founded. Historia Mathematica, 9(3): 342, August 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086082901288.

Anonymous:1982:LWC

[Ano82p] Anonymous. The life and work of Charles Babbage. Historia Mathematica, 9(4):485, November 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086082901082.

Anonymous:1982:MMX

[Ano82q] Anonymous. A mathematical manuscript of the XIV century. Historia Mathematica, 9(3):343, August 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508608290129X.

Anonymous:1982:NN

[Ano82r] Anonymous. News and notices. Historia Mathematica, 9(2):227, May 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082900143.

Anonymous:1982:NCa

[Ano82s] Anonymous. Notes on contributors. Historia Mathematica, 9(1): 128–131, February 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086082901549.

Anonymous:1982:NCb

[Ano82t] Anonymous. Notes on contributors. Historia Mathematica, 9(2):260– 261, May 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086082900854.

Anonymous:1982:PAFa

[Ano82u] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 9(1):132, February 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086082901550.

Anonymous:1982:PAFb

[Ano82v] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 9(2):262, May 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086082900866.

Anonymous:1982:PAFc

[Ano82w] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 9(3):371, August 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086082901331.

Anonymous:1982:PAFd

[Ano82x] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 9(4):506, November 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086082901173.

Anonymous:1982:PR

[Ano82y] Anonymous. A plane request. Historia Mathematica, 9(4):487, November 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086082901112.

Anonymous:1982:P

[Ano82z] Anonymous. Projects. Historia Mathematica, 9(2):228, May 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082900155.

Anonymous:1982:SHC

[Ano82-27] Anonymous. Seminario de historia de la ciencia y de la técnica de Aragón (Spanish) [Seminar on the history of science and technology of Aragon]. Historia Mathematica, 9(2):226, May 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082900131.

Anonymous:1982:VAA

[Ano82-28] Anonymous. Volterra Archive at the Accademia nazionale dei lincei. Historia Mathematica, 9(2):229–238, May 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086082900167.

Anonymous:1982:WPH

[Ano82-29] Anonymous. Workshop in the preparation of historical material for classroom use. Historia Mathematica, 9(4):486, November 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082901094.

Anonymous:1983:A

[Ano83a] Anonymous. Acknowledgment. Historia Mathematica, 10(4):494, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086083900265.

Anonymous:1983:ANI

[Ano83b] Anonymous. Australian news and information letter. Historia Mathematica, 10(1):101, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086083900459.

Anonymous:1983:CSD

[Ano83c] Anonymous. Centro studi della matematica medioevale publication series. Historia Mathematica, 10(4):460, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900101.

Anonymous:1983:CWK

[Ano83d] Anonymous. The collected works of Kurt Gödel. Historia Mathematica, 10(1):101–102, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086083900472.

Anonymous:1983:CDG

[Ano83e] Anonymous. Commemoration of the death of Giuseppe Peano. Historia Mathematica, 10(4):458, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086083900083.

Anonymous:1983:CRM

[Ano83f] Anonymous. Conference on Renaissance Mathematics: Keble College, Oxford, September 26–30, 1984. Historia Mathematica, 10(3): 368, August 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086083900848.

Anonymous:1983:DC

[Ano83g] Anonymous. Dissertation completed. Historia Mathematica, 10(1): 101, February 1983. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/0315086083900460.

Anonymous:1983:ECB

[Ano83h] Anonymous. Euler colloquium: Berlin, may 12–14, 1983. Historia Mathematica, 10(1):90–91, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086083900381.

Anonymous:1983:FWD

[Ano83i] Anonymous. First world directory of historians of Chinese mathematics. Historia Mathematica, 10(4):461, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900113.

Anonymous:1983:HLP

[Ano83j] Anonymous. The history of linear perspective. Historia Mathematica, 10(1):100, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086083900447.

Anonymous:1983:HMW

[Ano83k] Anonymous. History of Mathematics Workshop, Summer 1983: University of Toronto, Toronto, Canada July 25–August 2. Historia Mathematica, 10(2):210–211, May 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086083900666.

Anonymous:1983:IAA

[Ano83l] Anonymous. Index of authors of articles and reviews in volume 10. Historia Mathematica, 10(4):495, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900277.

Anonymous:1983:IAP

[Ano83m] Anonymous. Index of authors of publications reviewed and abstracted in volume 10. Historia Mathematica, 10(4):496–499, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086083900289.

Anonymous:1983:ISJ

[Ano83n] Anonymous. International Seminar on Jain Mathematics and Cosmology: November 20–24, 1983. Historia Mathematica, 10(3):369– 371, August 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508608390085X.

Anonymous:1983:JHM

[Ano83o] Anonymous. Journées D’histoire des Mathématiques: Lundi 6 Juin–Vendredi 10 Juin, 1983 Centre International de Rencontres Mathématiques, Marseille, France. Historia Mathematica, 10(2):211– 213, May 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086083900678.

Anonymous:1983:NCa

[Ano83p] Anonymous. Notes on contributors. Historia Mathematica, 10 (1):144–146, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086083900599.

Anonymous:1983:NCb

[Ano83q] Anonymous. Notes on contributors. Historia Mathematica, 10(4): 491–492, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086083900241.

Anonymous:1983:PAFa

[Ano83r] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 10(1):147, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086083900605.

Anonymous:1983:PAFb

[Ano83s] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 10(2):259, May 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086083900769.

Anonymous:1983:PAFc

[Ano83t] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 10(3):382, August 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086083900873.

Anonymous:1983:PAFd

[Ano83u] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 10(4):493, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086083900253.

Anonymous:1983:PP

[Ano83v] Anonymous. Projet de publication. Historia Mathematica, 10(2): 224, May 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086083900691.

Anonymous:1983:RMB

[Ano83w] Anonymous. Recent meetings of the British Society for the History of Mathematics. Historia Mathematica, 10(2):208–209, May 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900642.

Anonymous:1983:SCS

[Ano83x] Anonymous. The second Congress of the Spanish society of history of science. Historia Mathematica, 10(2):209–210, May 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900654.

Anonymous:1983:SHM

[Ano83y] Anonymous. A seminar in the history of mathematics. Historia Mathematica, 10(1):91–92, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086083900393.

Anonymous:1983:UHM

[Ano83z] Anonymous. Use of history of mathematics in schools and colleges. Historia Mathematica, 10(1):92, February 1983. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508608390040X.

Anonymous:1984:AEA

[Ano84a] Anonymous. Abstracts: Edited by Albert C. Lewis. Historia Mathematica, 11(1):112–119, February 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086084900193.

Anonymous:1984:A

[Ano84b] Anonymous. Acknowledgment. Historia Mathematica, 11(4):486, November 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086084900326.

Anonymous:1984:BREb

[Ano84c] Anonymous. Bertrand Russell’s early technical philosophy, 1895– 1922. Historia Mathematica, 11(2):226, May 1984. CODEN HI-MADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/S0315086084800166.

Anonymous:1984:BRSa

[Ano84d] Anonymous. Book review: A source book in Medieval science: Edited by Edward Grant. Cambridge, MA (Harvard Univ. Press). 1974. xviii + 864 pp. Historia Mathematica, 11(1):100–102, February 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086084900144.

Anonymous:1984:BRD

[Ano84e] Anonymous. Book review: Dürer: Kunst und Geometrie: By Eberhard Schröder. Berlin (Akademie-Verlag). 1980. 79 pp. Illustrated. Historia Mathematica, 11(1):97–98, February 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086084900120.

Anonymous:1984:BREa

[Ano84f] Anonymous. Book review: Elie Cartan — Albert Einstein, letters on absolute parallelism 1929–1932: Original text, English translation by Jules Leroy and Jim Ritter; edited by Robert Debever. Princeton, NJ (Princeton Univ. Press). 1979. xvii + 233 pp. Historia Mathematica, 11(1):94–96, February 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086084900119.

Anonymous:1984:BRH

[Ano84g] Anonymous. Book review: History of functional analysis: (Notas de Matemática, No. 77; North-Holland Mathematics Studies, No. 49). By Jean Dieudonné. Amsterdam (North-Holland). 1981. vi + 312 pp. U.S. $29.50/Dfl. 65.00. Historia Mathematica, 11(1):106–108, February 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508608490017X.

Anonymous:1984:BRM

[Ano84h] Anonymous. Book review: Mathematicians and Their Times: History of mathematics and mathematics of history: (North-Holland Mathematics Studies, No. 48). By Laurence Chisholm Young. Amsterdam/New York (North-Holland). 1981. x + 344 pp. $36.50. Historia Mathematica, 11(1):99–100, February 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086084900132.

Anonymous:1984:BRSb

[Ano84i] Anonymous. Book review: Sophie Germain. An essay in the history of the theory of elasticity: Studies in the history of modern science: Vol. 6. By Louis L. Bucciarelli and Nancy Dworsky. Dordrecht/Boston/London (Reidel). 1980. xii + 147 pp. Cloth Dfl. 60/U.S. $30.00 Paper Dfl. 30/U.S. $15.00. Historia Mathematica, 11(1):102–103, February 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086084900156.

Anonymous:1984:BRP

[Ano84j] Anonymous. Book review: The prehistory of the theory of distributions: Studies in the history of mathematics and physical sciences: By Jesper Lützen. New York/Heidelberg/Berlin (Springer-Verlag). 1982. 232 pp. $42.00. Historia Mathematica, 11(1):108–111, February 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086084900181.

Anonymous:1984:BRW

[Ano84k] Anonymous. Book review: Wissenschaft und Bildung in der Mathematik des 19. Jahrhunderts: By Hans Niels Jahnke. Bielefeld, BRD (Institüt für Didaktik der Mathematik, Universität Bielefeld). 1981. Historia Mathematica, 11(1):104–106, February 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086084900168.

Anonymous:1984:DMU

[Ano84l] Anonymous. The department of mathematics: university of algebra (1908–1982). Historia Mathematica, 11(2):221, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086084800130.

Anonymous:1984:DWH

[Ano84m] Anonymous. A directory of women in the history of science, medicine, and technology. Historia Mathematica, 11(2):230, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086084800191.

Anonymous:1984:HDP

[Ano84n] Anonymous. H. Davenport papers and letters. Historia Mathematica, 11(2):231, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086084800208.

Anonymous:1984:IAA

[Ano84o] Anonymous. Index of authors of articles and reviews in volume 11. Historia Mathematica, 11(4):487, November 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086084900338.

Anonymous:1984:IAP

[Ano84p] Anonymous. Index of authors of publications reviewed and abstracted in volume 11. Historia Mathematica, 11(4):488–490, November 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608490034X.

Anonymous:1984:KMF

[Ano84q] Anonymous. Kenneth O. May Fellowship in the History of Mathematics. Historia Mathematica, 11(3):322, August 1984.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086084900375.

Anonymous:1984:NCa

[Ano84r] Anonymous. Notes on contributors. Historia Mathematica, 11 (1):120–122, February 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508608490020X.

Anonymous:1984:NCb

[Ano84s] Anonymous. Notes on contributors. Historia Mathematica, 11(2): 250–251, May 1984. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086084800282.

Anonymous:1984:PAFa

[Ano84t] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 11(1):123, February 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086084900211.

Anonymous:1984:PAFb

[Ano84u] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 11(2):252, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086084800294.

Anonymous:1984:PAFc

[Ano84v] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 11(3):357, August 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508608490048X.

Anonymous:1984:PAFd

[Ano84w] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 11(4):485, November 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086084900314.

Anonymous:1984:RMB

[Ano84x] Anonymous. Recent meetings of the British Society for the History of Mathematics. Historia Mathematica, 11(1):92–93, February 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086084900090.

Anonymous:1984:SMP

[Ano84y] Anonymous. Sociedad Matematica Peruana. Historia Mathematica, 11(2):229, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S031508608480018X.

Anonymous:1984:SWD

[Ano84z] Anonymous. Supplement to world directory. Historia Mathematica, 11(4):481–482, November 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086084900302.

Anonymous:1984:ISI

[Ano84-27] Anonymous. The IVth scientific instrument symposium, October 23–26, 1984, Amsterdam. Historia Mathematica, 11(2):225–226, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086084800154.

Anonymous:1984:TLE

[Ano84-28] Anonymous. A tribute to Leonhard Euler: 1707–1783. Historia Mathematica, 11(2):220, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086084800117.

Anonymous:1984:TLN

[Ano84-29] Anonymous. Tricentenary of Leibniz’s “Nova Methodus”. Historia Mathematica, 11(1):93, February 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086084900107.

Anonymous:1985:AEAb

[Ano85a] Anonymous. Abstract: Edited by Albert C. Lewis. Historia Mathematica, 12(3):300–308, August 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086085900369.

Anonymous:1985:AEAa

[Ano85b] Anonymous. Abstracts: Edited by Albert C. Lewis. Historia Mathematica, 12(2):191–199, May 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086085900163.

Anonymous:1985:A

[Ano85c] Anonymous. Acknowledgment. Historia Mathematica, 12(4):405, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086085900618.

Anonymous:1985:AMS

[Ano85d] Anonymous. Australian mathematical society, 28th annual meeting. Historia Mathematica, 12(1):79, February 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086085900758.

Anonymous:1985:BRG

[Ano85e] Anonymous. Book review: Geschichte der Zahlentheorie im Orient im Mittelalter und zu Beginn der Neuzeit unter besonderer Berücksichtigung persischer Mathematiker: By Alireza Djafari Naini. Verlag Klose (Braunschweig). 1982. 233 pp. DM 38,-. Historia Mathematica, 12(3):295–296, August 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086085900333.

Anonymous:1985:BRI

[Ano85f] Anonymous. Book review: Introduzione all’aritmetica mercantile del medioevo e del rinascimento: By Raffaella Franci and Laura Toti Rigatelli. Published for the University of Siena. Quattri Venti (Urbino). 1982. 126 pp. 8 illustrations. Lire 12,000. Historia Mathematica, 12(3):293–295, August 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086085900321.

Anonymous:1985:BRR

[Ano85g] Anonymous. Book review: La raccolta Carlo Viganò: By Pier Luigi Pizzamiglio. Preface by Carlo Felice Manara. Brescia (Editrice La Scuola), 1979. 219 pp. 20 illustrations. Historia Mathematica, 12 (3):296–297, August 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086085900345.

Anonymous:1985:BRN

[Ano85h] Anonymous. Book review: Numbers: Their history and meaning: By Graham Flegg. New York (Schocken Books). 1983. x + 295 pp. $14.95. (Also in paperback: London (André Deutsch), 1983, £8.95.). Historia Mathematica, 12(3):292–293, August 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508608590031X.

Anonymous:1985:CSH

[Ano85i] Anonymous. Canadian society for history and philosophy of mathematics, annual meeting, June 2–4, 1985. Historia Mathematica, 12 (2):158, May 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086085900060.

Anonymous:1985:CEH

[Ano85j] Anonymous. Comisión de estudios de historia de la Matemática en América latina. Historia Mathematica, 12(4):373–374, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086085900497.

Anonymous:1985:CHP

[Ano85k] Anonymous. Conference on history and philosophy of modern mathematics. Historia Mathematica, 12(1):74, February 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086085900722.

Anonymous:1985:IAA

[Ano85l] Anonymous. Index of authors of articles and reviews in volume 12. Historia Mathematica, 12(4):406, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608590062X.

Anonymous:1985:IAP

[Ano85m] Anonymous. Index of authors of publications reviewed and abstracted in volume 12. Historia Mathematica, 12(4):407–410, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086085900631.

Anonymous:1985:ISJ

[Ano85n] Anonymous. International seminar on Jain mathematics and cosmology: Hastinapur (Meerut), India, April 26–28, 1985. Historia Mathematica, 12(2):172–174, May 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086085900114.

Anonymous:1985:JI

[Ano85o] Anonymous. Jean Itard (1902–1979). Historia Mathematica, 12(1): ??, February 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086085900643.

Anonymous:1985:JHS

[Ano85p] Anonymous. Journées d’histoire des sciences de Marseille: Les 4 et 5 octobre 1985. Historia Mathematica, 12(3):298–299, August 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086085900357.

Anonymous:1985:NEC

[Ano85q] Anonymous. New executive committee for ICHM. Historia Mathematica, 12(4):372, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086085900473.

Anonymous:1985:NUS

[Ano85r] Anonymous. New, updated supplement to the world directory of historians of mathematics. Historia Mathematica, 12(4):372–373, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086085900485.

Anonymous:1985:NCa

[Ano85s] Anonymous. Notes on contributors. Historia Mathematica, 12(1):98– 99, February 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086085900801.

Anonymous:1985:NCb

[Ano85t] Anonymous. Notes on contributors. Historia Mathematica, 12(2): 200–201, May 1985. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/0315086085900175.

Anonymous:1985:NCc

[Ano85u] Anonymous. Notes on contributors. Historia Mathematica, 12 (3):309–311, August 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086085900370.

Anonymous:1985:NCd

[Ano85v] Anonymous. Notes on contributors. Historia Mathematica, 12(4): 401–403, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508608590059X.

Anonymous:1985:PAFa

[Ano85w] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 12(1):100, February 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086085900813.

Anonymous:1985:PAFb

[Ano85x] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 12(2):202, May 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086085900187.

Anonymous:1985:PAFc

[Ano85y] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 12(3):312, August 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086085900382.

Anonymous:1985:PAFd

[Ano85z] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 12(4):404, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086085900606.

Anonymous:1985:PCH

[Ano85-27] Anonymous. Project on the contemporary history of mathematics in Mexico. Historia Mathematica, 12(3):274–275, August 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086085900266.

Anonymous:1985:RMC

[Ano85-28] Anonymous. Renaissance mathematics conference: Keble College, Oxford, September 27–30, 1984. Historia Mathematica, 12(2):174– 175, May 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086085900126.

Anonymous:1985:SDL

[Ano85-29] Anonymous. Symposium on disciplinary limitations in science. Historia Mathematica, 12(4):382, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086085900527.

Anonymous:1985:TIC

[Ano85-30] Anonymous. Third International Conference on the History of Chinese Science: Beijing, China, August 20–25, 1984. Historia Mathematica, 12(1):78, February 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086085900746.

Anonymous:1985:XIC

[Ano85-31] Anonymous. XVIIth international congress of history of science: Symposium on transmission of mathematical sciences. Historia Mathematica, 12(2):159–161, May 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086085900084.

Anonymous:1986:WCD

[Ano86a] Anonymous. 3rd workshop on concept development in Babylonian mathematics: Berlin, December 9–13, 1985. Historia Mathematica, 13(3):282, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086086900960.

Anonymous:1986:AEA

[Ano86b] Anonymous. Abstracts: Edited by Albert C. Lewis. Historia Mathematica, 13(2):196–209, May 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086086900443.

Anonymous:1986:AMH

[Ano86c] Anonymous. African miniconference on history of mathematics. Historia Mathematica, 13(1):60, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508608690234X.

Anonymous:1986:BRCa

[Ano86d] Anonymous. Book review: A convergence of lives — Sofia Kovalevskaia: Scientist, writer, revolutionary: By Ann Hibner Koblitz. Basel/Boston (Birkhäuser). 1983. xx + 305 pp. $19.95. Historia Mathematica, 13(2):191–193, May 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508608690042X.

Anonymous:1986:BRCb

[Ano86e] Anonymous. Book review: Classics of mathematics: Edited by Ronald Calinger. Oak Park, Illinois (Moore Publishing Company). 1982. xxv + 742 pp. $18.00 (paper). Historia Mathematica, 13(2):194– 195, May 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086086900431.

Anonymous:1986:CWS

[Ano86f] Anonymous. Classic works in science, mathematics, and logic. Historia Mathematica, 13(2):170, May 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086086900315.

Anonymous:1986:CIC

[Ano86g] Anonymous. Colloque international CNRS: “La physique” d’Aristote: L’analyse du continu et du mouvement et la possibilité d’une science de la nature. Historia Mathematica, 13(2):173, May 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086086900339.

Anonymous:1986:CCM

[Ano86h] Anonymous. Conference on Concepts of the Mathematical Infinite during the 19th Century: Sonderforschungsbereich “Wissen und Gesellschaft im 19. Jahrhundert” der Ruhr–Universität Bochum, Prof. Dr. Gert König. Historia Mathematica, 13(3):295, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608690100X.

Anonymous:1986:CIM

[Ano86i] Anonymous. Congress on Italian mathematics between the two world wars (1920–1940). Historia Mathematica, 13(3):278, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086086900947.

Anonymous:1986:CNT

[Ano86j] Anonymous. Congress on new trends in the history of science. Historia Mathematica, 13(2):174, May 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086086900340.

Anonymous:1986:CSA

[Ano86k] Anonymous. Contemporary scientific archives centre. Historia Mathematica, 13(1):42, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086086902260.

Anonymous:1986:FAS

[Ano86l] Anonymous. First Austrian symposium on the history of mathematics: Mathematics-stimulating or stimulated? Historia Mathematica, 13(1):58, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086086902302.

Anonymous:1986:FIC

[Ano86m] Anonymous. Fourth International Conference on the History of Chinese Science, May 16–21, 1986. Historia Mathematica, 13(1):57, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086086902296.

Anonymous:1986:HPMa

[Ano86n] Anonymous. History and pedagogy of mathematics study group. Historia Mathematica, 13(1):58–59, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086086902314.

Anonymous:1986:HPMb

[Ano86o] Anonymous. History and philosophy of mathematics meeting. Historia Mathematica, 13(1):59, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086086902326.

Anonymous:1986:IAA

[Ano86p] Anonymous. Index of authors of articles and reviews in volume 13. Historia Mathematica, 13(4):390, November 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086086900625.

Anonymous:1986:IAP

[Ano86q] Anonymous. Index of authors of publications reviewed and abstracted in volumes 1–13. Historia Mathematica, 13(4):391–423, November 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086086900637.

Anonymous:1986:ISH

[Ano86r] Anonymous. International symposium for history of mathematics and mathematical education using Chinese characters. Historia Mathematica, 13(1):60, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086086902338.

Anonymous:1986:NPC

[Ano86s] Anonymous. Newton’s Principia 1687–1987: College Park, Maryland, and Washington, D.C., April 23–25, 1987. Historia Mathematica, 13(4):387, November 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086086900595.

Anonymous:1986:NCa

[Ano86t] Anonymous. Notes on contributors. Historia Mathematica, 13(2): 210–212, May 1986. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/0315086086900455.

Anonymous:1986:NCc

[Ano86u] Anonymous. Notes on contributors. Historia Mathematica, 13(4):388, November 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086086900601.

Anonymous:1986:NCb

[Ano86v] Anonymous. Notes to contributors. Historia Mathematica, 13(3):321– 322, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086086901096.

Anonymous:1986:PCT

[Ano86w] Anonymous. Papers of C. A. Truesdell and R. L. Wilder. Historia Mathematica, 13(2):171–172, May 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086086900327.

Anonymous:1986:PAFa

[Ano86x] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 13(1):97, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086086902442.

Anonymous:1986:PAFb

[Ano86y] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 13(2):213, May 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086086900467.

Anonymous:1986:PAFc

[Ano86z] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 13(3):323, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086086901102.

Anonymous:1986:PAFd

[Ano86-27] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 13(4):389, November 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086086900613.

Anonymous:1986:TWT

[Ano86-28] Anonymous. Two weekends on Teichmüller. Historia Mathematica, 13(2):174–175, May 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086086900352.

Anonymous:1987:AEAb

[Ano87a] Anonymous. Abstracts: Edited by Albert C. Lewis. Historia Mathematica, 14(3):298–307, August 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086087900619.

Anonymous:1987:AEAa

[Ano87b] Anonymous. Abstracts: Edited by Albert C. Lewis and Garry J. Tee. Historia Mathematica, 14(2):220–230, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086087900401.

Anonymous:1987:BCR

[Ano87c] Anonymous. Bicentennial commemoration of R. Boscovich: Osservatorio Astronomico di Brera Milano, September 15–18, 1987. Historia Mathematica, 14(3):288–289, August 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086087900516.

Anonymous:1987:BRC

[Ano87d] Anonymous. Book review: Coloured quadrangles, a guide to the tenth book of Euclid’s Elements: By C. M. Taisbak. Copenhagen (Museum Tusculanum Press). 1982. 78 pp. Opuscula Graeco-Latina (Supplementa Musei Tusculani), Vol. 24. 45 Dkr. Historia Mathematica, 14(1):67–70, February 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086087900115.

Anonymous:1987:BRG

[Ano87e] Anonymous. Book review: George Boole: His life and work: By Desmond MacHale. Profiles of Genius, Vol. 2. Dún Laoghaire (Boole Press Limited). 1985. xiii + 304 pp. $24.95. Historia Mathematica, 14(2):211–214, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086087900383.

Anonymous:1987:BRW

[Ano87f] Anonymous. Book review: I want to be a mathematician: an automathography: By Paul R. Halmos. New York (Springer-Verlag).

1985. 421 pp. with 43 photographs. $41.50. Historia Mathematica, 14(2):200–203, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086087900358.

Anonymous:1987:BRI

[Ano87g] Anonymous. Book review: Ibn al-Haytham’s completion of the conics: By Jan Pieter Hogendijk. Berlin/Heidelberg/New York/Tokyo (Springer-Verlag). 1985. 417 pp. $98.00. Historia Mathematica, 14 (1):65–67, February 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086087900103.

Anonymous:1987:BRR

[Ano87h] Anonymous. Book review: Recherches sur l’oeuvre mathématique de Mei Wending (1633–1721): By Jean-Claude Martzloff. Mémoires de l’Institut des Hautes Etudes Chinoises, Vol. 16. Paris (Collège de France). 1981. 472 pp. Historia Mathematica, 14(1):70–75, February 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086087900127.

Anonymous:1987:BRS

[Ano87i] Anonymous. Book review: Siméon-Denis Poisson et la science de son temps: Edited by Michel Métivier, Pierre Costabel, and Pierre Dugac. Paris (École Polytechnique-Palaisseu avec l’aide de la Délégation aux célébrations nationales). 1981. iv + 285 pp., bibliography, index. Historia Mathematica, 14(2):198–200, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086087900346.

Anonymous:1987:BRHa

[Ano87j] Anonymous. Book review: The history of combinatorial group theory: a case study in the history of ideas: (Studies in the History of Mathematics and Physical Sciences, Vol. 9). By Bruce Chandler and Wilhelm Magnus. (New York/Heidelberg/Berlin: Springer-Verlag). 1982. viii + 234 pp. Historia Mathematica, 14(2):209–211, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086087900371.

Anonymous:1987:BRN

[Ano87k] Anonymous. Book review: The nature of mathematical knowledge: By Philip S. Kitcher. New York/Oxford (Oxford University Press). 1983. xii + 286 pp. $25.00 hardbound; $8.95 paperback. Historia Mathematica, 14(2):204–209, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508608790036X.

Anonymous:1987:BRHb

[Ano87l] Anonymous. Boscovich in Rome, 1725–1759: His scientific life and works: Rome, Italy, April 1988. Historia Mathematica, 14(3):290, August 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086087900541.

Anonymous:1987:DMV

[Ano87m] Anonymous. Deutsche mathematiker vereinigung, Jahrestagung 1986: Universität Marburg, September 14–19, 1986. Historia Mathematica, 14(1):63–64, February 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086087900097.

Anonymous:1987:DMN

[Ano87n] Anonymous. Discussion meeting: Newton’s Principia and its legacy: Organized by D. G. King-Hele and A. R. Hall on behalf of the British National Committee for the History of Science, Medicine and Technology Carlton House Terrace, London, June 30, 1987. Historia Mathematica, 14(3):294–295, August 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086087900589.

Anonymous:1987:E

[Ano87o] Anonymous. Erratum. Historia Mathematica, 14(3):297, August 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086087900607.

Anonymous:1987:IAA

[Ano87p] Anonymous. Index of authors of articles and reviews in volume 14. Historia Mathematica, 14(4):399, November 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086087900747.

Anonymous:1987:IAP

[Ano87q] Anonymous. Index of authors of publications reviewed and abstracted in volume 14. Historia Mathematica, 14(4):401–404, November 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086087900759.

Anonymous:1987:MRD

[Ano87r] Anonymous. Mathesis: Revista de divulgación e información en la historia y filosofía de las matemáticas. Historia Mathematica, 14(1): 54, February 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086087900036.

Anonymous:1987:NLW

[Ano87s] Anonymous. The Nachlass [legacy] of Wilhelm Lorey. Historia Mathematica, 14(1):55–57, February 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086087900048.

Anonymous:1987:NN

[Ano87t] Anonymous. News and notices. Historia Mathematica, 14(1):76, February 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086087900139.

Anonymous:1987:NCa

[Ano87u] Anonymous. Notes on contributors. Historia Mathematica, 14(1):77– 78, February 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086087900140.

Anonymous:1987:NCb

[Ano87v] Anonymous. Notes on contributors. Historia Mathematica, 14(2): 231–233, May 1987. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/0315086087900413.

Anonymous:1987:NCc

[Ano87w] Anonymous. Notes on contributors. Historia Mathematica, 14 (3):308–309, August 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086087900620.

Anonymous:1987:PAFa

[Ano87x] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 14(2):234, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086087900425.

Anonymous:1987:PAFb

[Ano87y] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 14(3):310, August 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086087900632.

Anonymous:1987:PAFc

[Ano87z] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 14(4):399, November 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086087900735.

Anonymous:1987:PRS

[Ano87-27] Anonymous. Pietro Riccardi (1828–1898): e la storiografia delle matematiche in Italia 16–18 marzo 1987. (Italian) [Pietro Riccardi (1828–1898): and historiography of mathematics in Italy 16–18 March 1987]. Historia Mathematica, 14(2):182, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086087900267.

Anonymous:1987:PCI

[Ano87-28] Anonymous. Premier colloque international sur l’histoire des mathématiques arabes. Historia Mathematica, 14(2):194, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086087900310.

Anonymous:1987:PIB

[Ano87-29] Anonymous. Project on interrelations between the development of theoretical mathematics and applications during the 19th century. Historia Mathematica, 14(2):196–197, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086087900334.

Anonymous:1987:RHM

[Ano87-30] Anonymous. Resources in the history of mathematics at the University of Alabama. Historia Mathematica, 14(1):57–59, February 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608790005X.

Anonymous:1987:SLA

[Ano87-31] Anonymous. Second Latin–American Congress on the History of Science and Technology: São Paulo, Brazil, 30 June to 4 July 1988. Historia Mathematica, 14(3):291, August 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086087900553.

Anonymous:1987:SSB

[Ano87-32] Anonymous. Special session on Bertrand Russell: Annual Meeting of the Canadian Society for the History and Philosophy of Mathematics. Historia Mathematica, 14(2):183, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086087900279.

Anonymous:1987:SIP

[Ano87-33] Anonymous. Subject index of publications abstracted in volumes 1–13. Historia Mathematica, 14(1):79–117, February 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086087900152.

Anonymous:1987:DTH

[Ano87-34] Anonymous. The Durham Thomas Harriot Seminar: Thomas Harriot and His Times Durham, December 14–15, 1987. Historia Mathematica, 14(3):289–290, August 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508608790053X.

Anonymous:1987:VIC

[Ano87-35] Anonymous. VIIIth international congress of logic, methodology, and philosophy of science: Moscow, USSR, August 17–22, 1987. Historia Mathematica, 14(1):60, February 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086087900061.

Anonymous:1987:XIC

[Ano87-36] Anonymous. XVIIth international congress of history of science: Symposium No. 18 on Transmission of Mathematical Sciences July 31–August 8, 1985. Historia Mathematica, 14(2):185–191, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086087900292.

Anonymous:1988:SAP

[Ano88a] Anonymous. 2nd school of all Poland of the history of mathematics: Mathematics of the 19th century may 25–29, 1987, miedzyzdroje. Historia Mathematica, 15(2):151–152, May 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900067.

Anonymous:1988:AEAb

[Ano88b] Anonymous. Abstract: Edited by Albert C. Lewis and David E. Zitarelli. Historia Mathematica, 15(4):397–401, November 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900419.

Anonymous:1988:AEAa

[Ano88c] Anonymous. Abstracts: Edited by, Albert C. Lewis and David E. Zitarelli. Historia Mathematica, 15(2):188–194, May 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900183.

Anonymous:1988:A

[Ano88d] Anonymous. Acknowledgment. Historia Mathematica, 15(2):199, May 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086088900225.

Anonymous:1988:BRA

[Ano88e] Anonymous. Book review: Alfred North Whitehead: the man and his work. Volume I: 1861–1910: By Victor Lowe. Baltimore and London (Johns Hopkins University Press). 1985. xi + 352 pp. $27.50. Historia Mathematica, 15(2):176–179, May 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900134.

Anonymous:1988:BRI

[Ano88f] Anonymous. Book review: Intellectual mastery of nature: Theoretical physics from Ohm to Einstein: Vol. 1, The Torch of Mathematics 1800–1870 ($55.00); Vol. 2, The Now Mighty Theoretical Physics 1870–1925 ($65.00). By Christa Jungnickel and Russell Mc-Cormmach. Chicago (University of Chicago Press). 1986. xxiii + 350; xii + 435 pp. Historia Mathematica, 15(4):389–392, November 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900389.

Anonymous:1988:BRM

[Ano88g] Anonymous. Book review: Mathematical perspectives: Essays on mathematics and its historical development: Edited by Joseph W. Dauben. New York (Academic Press). 1981. 259 pp. $49.50. Historia Mathematica, 15(4):386–388, November 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900377.

Anonymous:1988:BRN

[Ano88h] Anonymous. Book review: Native American mathematics: Edited by Michael P. Closs. Austin, Texas (University of Texas Press). 1986. 439 pp. 13 photographs, 55 line drawings, 20 tables. Bibliography. Historia Mathematica, 15(4):384–386, November 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900365.

Anonymous:1988:BRPb

[Ano88i] Anonymous. Book review: Philosophie als Analysis: Studien zur Entwicklung philosophischer Analysiskonzeptionen unter den Einfluss mathematischer Methoden-Modelle im 17. und frühen 18. Jahrhundert: By Hans-Jürgen Engfer. Forschungen und Materialien zur deutschen Aufklärung (Series 2, Vol. 1). Stuttgart-Bad Cannstatt (Fromman-Holzboog). 1982. 293 pp. $50.00. Historia Mathematica, 15(4):392–395, November 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086088900390.

Anonymous:1988:BRPa

[Ano88j] Anonymous. Book review: Ptolemy’s Almagest: Translated and annotated by G. T. Toomer. New York/Berlin/Heidelberg/Tokyo (Springer-Verlag). 1984. 693 pp. Historia Mathematica, 15(2):174– 176, May 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086088900122.

Anonymous:1988:BRSa

[Ano88k] Anonymous. Book review: Science and technology in medieval India — A bibliography of source materials in Sanskrit, Arabic and Persian: By A. Rahman with the Research Assistance of M. A. Alvi, S. A. Khan Ghori, and K. V. Sarma Murthy. New Delhi (Indian National Science Academy). 1982. xxxi + 179 pp. Rs. 200 (foreign price: $70). Historia Mathematica, 15(2):179–181, May 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900146.

Anonymous:1988:BRSc

[Ano88l] Anonymous. Book review: Statistics in Britain, 1865–1930: the social construction of scientific knowledge: By Donald A. MacKenzie. (Edinburgh University Press). 1981. viii + 306 pp. Historia Mathematica, 15(2):185–187, May 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900171.

Anonymous:1988:BRSb

[Ano88m] Anonymous. Book review: The śulbasūtras of Baudhāyana, Āpastamba, Kātyāyana and Mānava with text, English translation and commentary: By S. N. Sen and A. K. Bag. New Delhi (Indian National Science Academy). 1983. viii + 293 pp. Rs. 85, $30. Historia Mathematica, 15(2):183–185, May 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508608890016X.

Anonymous:1988:BRE

[Ano88n] Anonymous. Book review: Über die entstehung von David Hilberts “Grundlagen der Geometrie”: By M.-M. Toepell. Studien zur Wissenschafts-, Sozial- und Bildungsgeschichte der Mathematik, Band 2. Göttingen (Vandenhoeck & Ruprecht). 1986. 293 + xiv pp. DM 78. Historia Mathematica, 15(2):181–183, May 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900158.

Anonymous:1988:CCP

[Ano88o] Anonymous. Colloque “Condorcet”: Paris, 8–11 juin 1988. Historia Mathematica, 15(2):149–151, May 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900055.

Anonymous:1988:Ea

[Ano88p] Anonymous. Erratum. Historia Mathematica, 15(2):195, May 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900195.

Anonymous:1988:Eb

[Ano88q] Anonymous. Erratum. Historia Mathematica, 15(4):396, November 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900407.

Anonymous:1988:FSS

[Ano88r] Anonymous. First Seminar on the Scientific Tradition of Brescia: Francesco Lana Terzi (1631–1687): La scuola gesuitica e la scienza moderna October 8, 1987. Historia Mathematica, 15(2):164, May 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900080.

Anonymous:1988:IAA

[Ano88s] Anonymous. Index of authors of articles and reviews in volume 15. Historia Mathematica, 15(4):406, November 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900432.

Anonymous:1988:IAP

[Ano88t] Anonymous. Index of authors of publications reviewed and abstracted in volume 15. Historia Mathematica, 15(4):407–409, November 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900444.

Anonymous:1988:ISC

[Ano88u] Anonymous. International symposium on Cauchy and the French mathematical world: Paris, France, 27–29 July 1989. Historia Mathematica, 15(4):370–371, November 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900298.

Anonymous:1988:JHS

[Ano88v] Anonymous. Journées d’histoire des sciences: La Mathématisation (1780–1830) 11–12 Mars 1988. Historia Mathematica, 15(3):283, August 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086088900742.

Anonymous:1988:NMI

[Ano88w] Anonymous. Neugebauer microform index of the exact sciences. Historia Mathematica, 15(2):148, May 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900043.

Anonymous:1988:NC

[Ano88x] Anonymous. Notes on contributors. Historia Mathematica, 15(2): 196–197, May 1988. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/0315086088900201.

Anonymous:1988:PAFa

[Ano88y] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 15(1):100, February 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900614.

Anonymous:1988:PAFb

[Ano88z] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 15(2):198, May 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086088900213.

Anonymous:1988:PAFc

[Ano88-27] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 15(3):298, August 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508608890081X.

Anonymous:1988:PAFd

[Ano88-28] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 15(4):405, November 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900420.

Anonymous:1988:SSC

[Ano88-29] Anonymous. Summer school and conference on mathematical logic. Historia Mathematica, 15(1):75–76, February 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900511.

Anonymous:1988:THM

[Ano88-30] Anonymous. Teaching and the history of mathematics. Historia Mathematica, 15(3):287–289, August 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900778.

Anonymous:1988:FIS

[Ano88-31] Anonymous. The Fourth International Symposium on the History of Arabic Science: Aleppo University, April 21–26, 1987. Historia Mathematica, 15(1):85–86, February 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900560.

Anonymous:1989:AEDa

[Ano89a] Anonymous. Abstract: Edited by David E. Zitarelli. Historia Mathematica, 16(2):194–202, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086089900566.

Anonymous:1989:AEDb

[Ano89b] Anonymous. Abstracts: Edited by David E. Zitarelli. Historia Mathematica, 16(3):296–304, August 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086089900311.

Anonymous:1989:BRCb

[Ano89c] Anonymous. Book review: Cantorian set theory and limitation of size: By Michael Hallett. Foreword by M. Dummett. Oxford Logic Guides: 10. Oxford/New York (Clarendon Press-Oxford Univ. Press).

1984. xxii + 343 pp. $25.00. Historia Mathematica, 16(3):291–295, August 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086089900293.

Anonymous:1989:BRCa

[Ano89d] Anonymous. Book review: Chinese mathematics: a concise history: By Li Yan and Du Shiran (translated by John N. Crossley and Anthony W. C. Lun). Oxford (Clarendon). 1987. xiii + 290 pp. $50.00. Historia Mathematica, 16(2):191–193, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086089900554.

Anonymous:1989:BRE

[Ano89e] Anonymous. Book review: Episodes in the mathematics of Medieval Islam: By J. L. Berggren. New York (Springer-Verlag). 1986. xiv + 197 pp. $23.00 (cloth). Historia Mathematica, 16(2):188–190, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086089900542.

Anonymous:1989:BRG

[Ano89f] Anonymous. Book review: Gauß, eine biographische studie: By Walter Kaufmann–Bühler. Berlin/Heidelberg (Springer-Verlag). 1987. 191 S. Kartoniert DM 56. Historia Mathematica, 16(2):178–179, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086089900499.

Anonymous:1989:BRH

[Ano89g] Anonymous. Book review: Histoire de l’École Polytechnique: By Ambroise Fourcy. Introduction and notes by Jean Dhombres. Paris (Belin). 1987. viii + 516 + 198 pp. 140 francs. Historia Mathematica, 16(2):184–187, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086089900530.

Anonymous:1989:BRI

[Ano89h] Anonymous. Book review: Islamic mathematical astronomy: By D. A. King. Collected Studies Series, CS 231. London (Variorum Reprints). 1986. 342 pp. Historia Mathematica, 16(3):295, August 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608990030X.

Anonymous:1989:BRR

[Ano89i] Anonymous. Book review: La république avait besoin de savants: By Janis Langins. Paris (Belin). 1987. 287 pp. 9 illustrations. 90 francs. Historia Mathematica, 16(2):184, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086089900529.

Anonymous:1989:BRN

[Ano89j] Anonymous. Book review: Napoleon und die Naturwissenschaften: By Joachim Fischer. Boethius-Texte und Abhandlungen zur Geschicted der exacten Wissenchaften, Band 16, Stuutgart (Franz Steiner Verlag Wiesbaden GmbH) 1988. xiv + 390 pp. DM. Historia Mathematica, 16(2):180, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086089900505.

Anonymous:1989:BRT

[Ano89k] Anonymous. Book review: Teubner-Archiv zur Mathematik. Teubner Verlagsgesellschaft, Leipzig: Vol. 1: C. F. Gauß, B. Riemann, H. Minkowski. Gaußsche Flächentheorie, Riemannsche Räume und Minkowski-Welt. 1984. 155 pp. DM 28. Vol. 2: G. Cantor. Ü ber unendliche, lineare Punktmannigfalttigkeiten. Arbeiten zur Mengenlehre aus den Jahren 1872–1884. 1984. 180 pp. DM 32. Vol. 3: G. Herglotz, Vorlesungen Über die Mechanik der Kontinua. 1985. 251 pp. DM 48. Vol. 4: H. Riechardt. Gaußunf die anfänge der nichteuklidischen Geometrie. Mit Original- beiten von J. Bolyai, N. I. Lobatschewski und F. Klein. Historia Mathematica, 16(2):180–184, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086089900517.

Anonymous:1989:BRW

[Ano89l] Anonymous. Book review: Women of mathematics: a biobibliographic sourcebook: Edited by Louise S. Grinstein and Paul J. Campbell, with a foreword by Alice Schafer. New York (Greenwood). 1987. 292 pp. $45.00. Historia Mathematica, 16(3):289–291, August 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086089900281.

Anonymous:1989:CIS

[Ano89m] Anonymous. Colloque international 1830–1930: Un siècle de géométrie, de C. F. Gauss et B. Riemann à H. Poincaré et E. Cartan. Epistémologie et histoire: Probablement à l’Institut Henri Poincaré, 11 rue Pierre et Marie Curie, 75230 Paris, Cedex 05, France, 19– 23 September 1989. Historia Mathematica, 16(1):87–88, February 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086089901018.

Anonymous:1989:CMP

[Ano89n] Anonymous. Colloque mathématiques et philosophie à l’age classique: Recherches historiques: Organisé par l’Association “Méditerranée et Histoire des Sciences” et l’équipe REHSEIS. Historia Mathematica, 16(4):386, November 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086089900888.

Anonymous:1989:GSI

[Ano89o] Anonymous. Gauss symposium: an International Conference on Mathematics and Theoretical Physics Guarujá, Brazil, July 22–27, 1989. Historia Mathematica, 16(2):166, May 1989. CODEN HI-MADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0315086089900396.

Anonymous:1989:IAA

[Ano89p] Anonymous. Index of authors of articles and reviews in volume 16. Historia Mathematica, 16(4):409, November 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086089900943.

Anonymous:1989:IAP

[Ano89q] Anonymous. Index of authors of publications reviewed and abstracted in volume 16. Historia Mathematica, 16(4):410–413, November 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086089900955.

Anonymous:1989:ISS

[Ano89r] Anonymous. International symposium on “Structures in mathematical theories”: San Sebastián, Spain, 24–28 September 1990. Historia Mathematica, 16(3):270–271, August 1989. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086089900220.

Anonymous:1989:KRB

[Ano89s] Anonymous. Kurt-Reinhard Biermann zum 70. Geburtstag. (German) [Kurt-Reinhard Biermann’s 70th birthday]. Historia Mathematica, 16(4):??, November 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086089900773.

Anonymous:1989:MIF

[Ano89t] Anonymous. Les mathématiques italiennes et françaises entre les deux guerres: Colloque à l’Institut Culturel Italien Hôtel de Galliffet, 50 rue de Varenne, VIIe, 6–7 décembre 1988. Historia Mathematica, 16(2):173–174, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086089900463.

Anonymous:1989:MPH

[Ano89u] Anonymous. Mario Pieri: His contributions to the foundations and teaching of geometry. Historia Mathematica, 16(3):287–288, August 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608990027X.

Anonymous:1989:NN

[Ano89v] Anonymous. News and notices. Historia Mathematica, 16(3):283– 284, August 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086089900256.

Anonymous:1989:N

[Ano89w] Anonymous. Notes on. Historia Mathematica, 16(4):406–407, November 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508608990092X.

Anonymous:1989:NCa

[Ano89x] Anonymous. Notes on contributors. Historia Mathematica, 16 (1):104–105, February 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086089901080.

Anonymous:1989:NCb

[Ano89y] Anonymous. Notes on contributors. Historia Mathematica, 16(2): 203–204, May 1989. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/0315086089900578.

Anonymous:1989:PAFa

[Ano89z] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 16(1):106, February 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086089901092.

Anonymous:1989:PAFb

[Ano89-27] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 16(2):205, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508608990058X.

Anonymous:1989:PAFc

[Ano89-28] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 16(3):307, August 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086089900323.

Anonymous:1989:PAFd

[Ano89-29] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 16(4):408, November 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086089900931.

Anonymous:1989:RTP

[Ano89-30] Anonymous. Ratio: la théorie des proportions de l’antiquité au XIXème siècle: Trento, Italia, 9–13 Janvier 1989. Historia Mathematica, 16(1):89, February 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508608990102X.

Anonymous:1989:RGR

[Ano89-31] Anonymous. The research group REHSEIS in Paris. Historia Mathematica, 16(2):176–177, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086089900487.

Anonymous:1989:RPC

[Ano89-32] Anonymous. A research program and a course in the history of mathematics: Ethnomathematics. Historia Mathematica, 16(3):285– 287, August 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086089900268.

Anonymous:1989:SAS

[Ano89-33] Anonymous. Science in antiquity. Scientific Centre of Leningrad of the Academy of Sciences of the USSR, March 20–22, 1989: Leningrad Section of the Institute of the History of Science and Technology of the Academy of Sciences of the USSR. Historia Mathematica, 16 (4):381–382, November 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086089900864.

Anonymous:1989:FIL

[Ano89-34] Anonymous. The Fifth International Leibniz Congress: Hanover, Federal Republic of Germany, November 14–19, 1988. Historia Mathematica, 16(2):172–173, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086089900451.

Anonymous:1989:REC

[Ano89-35] Anonymous. The Riccati equation in control, systems, and signals: a Workshop Celebrating Two and a Half Centuries of the Riccati Equation Villa Gallia, Como, Italy, June 26–28, 1989. Historia Mathematica, 16(2):164–165, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086089900384.

Anonymous:1990:MSH

[Ano90a] Anonymous. 3rd Maghrebian Symposium on the History of Arabic Mathematics: Algiers, Algeria, 1–3 December 1990. Historia Mathematica, 17(3):260–261, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609090005X.

Anonymous:1990:A

[Ano90b] Anonymous. Announcement. Historia Mathematica, 17(3):294, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086090900215.

Anonymous:1990:AMT

[Ano90c] Anonymous. Archimede: Mito Tradizione Scienza Convegno Internazionale di Studi: Syracusa-Catania, Italy, October 9–12, 1989. Historia Mathematica, 17(2):165–166, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609090060Q.

Anonymous:1990:AHP

[Ano90d] Anonymous. Association Henri Poincaré pour l’histoire et la philosophie des mathématiques et de la physique modernes. Historia Mathematica, 17(2):167–168, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609090061H.

Anonymous:1990:BCC

[Ano90e] Anonymous. The Bicentennial Conference on Computing: a forthcoming event to commemorate the 200th anniversary of the birth of Charles Babbage London, 1–3 July 1991. Historia Mathematica, 17(3):259, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609090004W.

Anonymous:1990:BRR

[Ano90f] Anonymous. Book review: R. G. Boscovich: His Life and Scientific Work: Istituto della Enciclopedia Italiana, Accademia Nazionale delle Scienze detta dei XL, Piazza Paganica 4, 00186 Roma, May 23–27, 1988. Historia Mathematica, 17(4):378–380, November 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090033A.

Anonymous:1990:CIS

[Ano90g] Anonymous. Colloque International 1830–1930: Un siècle de géometrie, de C. F. Gauss et B. Riemann à H. Poincaré et E. Cartan: Epistémologie, Histoire et mathématiques: Paris, September 18–23, 1989. Historia Mathematica, 17(2):162–164, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090058L.

Anonymous:1990:CIV

[Ano90h] Anonymous. Colloque international sur la vie et l’oeuvre de Girard Desargues. Historia Mathematica, 17(4):374, November 1990.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090028C.

Anonymous:1990:CCM

[Ano90i] Anonymous. Colloque le Continu Mathématique: Cerisy-la-Salle, 11–21 Septembre 1990. Historia Mathematica, 17(2):154–155, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090054H.

Anonymous:1990:COB

[Ano90j] Anonymous. Colloquium on the Occasion of the 450th Birthday of Bartholomäus Scultetus (1540–1614): Museum, Neißstraße 30, 8900 Görlitz, German Democratic Republic, May 26, 1990. Historia Mathematica, 17(3):268, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086090900123.

Anonymous:1990:EQI

[Ano90k] Anonymous. Einstein et les quanta: Institut Henri Poincaré, Paris, France, 9 décembre 1989. Historia Mathematica, 17(3):261, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090006Y.

Anonymous:1990:GWL

[Ano90l] Anonymous. G. W. Leibniz: Mathematiker, Naturwissenschaftler, Techniker: Symposium on the Occasion of the Annual Convention of the Gesellschaft für Angewandte Mathematik und Mechanik Hanover, Auditorium Maximum of the Technische Universität, April 9–10, 1990. Historia Mathematica, 17(4):376–377, November 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086090900318.

Anonymous:1990:GSI

[Ano90m] Anonymous. Gauss Symposium: International Conference on Mathematics and Theoretical Physics Guarujá, Brazil, July 24–27, 1989. Historia Mathematica, 17(1):73, February 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609090085R.

Anonymous:1990:HMS

[Ano90n] Anonymous. History of mathematics symposium in conjunction with the XXIst international Congress of mathematicians. Historia Mathematica, 17(2):153, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609090052F.

Anonymous:1990:IAA

[Ano90o] Anonymous. Index of authors of articles and reviews in volume 17. Historia Mathematica, 17(4):418, November 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090045F.

Anonymous:1990:IAP

[Ano90p] Anonymous. Index of authors of publications reviewed and abstracted in volume 17. Historia Mathematica, 17(4):419–423, November 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090046G.

Anonymous:1990:JHE

[Ano90q] Anonymous. Journées d’Histoire et d’Epistémologie: La gravitation newtonienne: physique et mécanique de Newton à Euler: Observatoire de Meudon, Paris, 22–23 juin 1989. Historia Mathematica, 17(3):267–268, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086090900112.

Anonymous:1990:LNS

[Ano90r] Anonymous. Lubos Nový: Sexagenarian. Historia Mathematica, 17 (2):??, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609090047H.

Anonymous:1990:MHC

[Ano90s] Anonymous. Math-History Conference: University of Wisconsin–La Crosse, October 5 and 6, 1990. Historia Mathematica, 17(2):154, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090053G.

Anonymous:1990:MSP

[Ano90t] Anonymous. Mathematical sciences past and present: 300 years of the Mathematische Gesellschaft in Hamburg Geomatikum, Universität Hamburg, Bundesstr. 55, 2 Hamburg 13, March 18–24, 1990. Historia Mathematica, 17(4):377–378, November 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086090900329.

Anonymous:1990:NEH

[Ano90u] Anonymous. National Endowment for the Humanities Summer Seminar on Plato’s Republic and Euclid’s Elements: June 24– August 2, 1991. Historia Mathematica, 17(4):375–376, November 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090030H.

Anonymous:1990:NSD

[Ano90v] Anonymous. A new series devoted to history and philosophy of science. Historia Mathematica, 17(2):168–169, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090062I.

Anonymous:1990:NYH

[Ano90w] Anonymous. A new yearbook for the history of science. Historia Mathematica, 17(4):384, November 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609090037E.

Anonymous:1990:NCa

[Ano90x] Anonymous. Notes on contributors. Historia Mathematica, 17(1):97, February 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609090090Z.

Anonymous:1990:NCb

[Ano90y] Anonymous. Notes on contributors. Historia Mathematica, 17(2): 198–199, May 1990. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/031508609090071K.

Anonymous:1990:NCc

[Ano90z] Anonymous. Notes on contributors. Historia Mathematica, 17 (3):310–311, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086090900237.

Anonymous:1990:NCd

[Ano90-27] Anonymous. Notes on contributors. Historia Mathematica, 17 (4):415–416, November 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609090043D.

Anonymous:1990:PAFa

[Ano90-28] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 17(1):98, February 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609090091Q.

Anonymous:1990:PAFb

[Ano90-29] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 17(2):200, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609090072L.

Anonymous:1990:PAFc

[Ano90-30] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 17(3):312, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086090900248.

Anonymous:1990:PAFd

[Ano90-31] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 17(4):417, November 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609090044E.

Anonymous:1990:PJS

[Ano90-32] Anonymous. Primeras Jornadas de la Sociedad Leibniz de España: Analogia y expresion en Leibniz Facultad de Filosofia y Ciencias de la Educacion, Universidad Complutense, 28040 Madrid; September 20–22, 1989. Historia Mathematica, 17(2):162, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090057K.

Anonymous:1990:SAG

[Ano90-33] Anonymous. Second annual Göttingen Workshop on the history of modern mathematics 1991: Mathematisches Institut der Universität Göttingen, Bunsenstraße 3–5, 3400 Göttingen, Federal Republic of Germany, July 15–17, 1991. Historia Mathematica, 17(4):374–375, November 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609090029D.

Anonymous:1990:SHM

[Ano90-34] Anonymous. Seminaire d’Histoire des Mathematiques d’Orsay: Départments de Mathématiques, Université de Paris-Sud, Centre d’Orsay, Bt. 425, 1^{e} étage, Salle 121–123, France. Historia Mathematica, 17(3):263, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086090900082.

Anonymous:1990:SIG

[Ano90-35] Anonymous. Symposium International — G. W. Leibniz: Le meilleur des mondes: Domaine de Seillac, France, June 7– 9, 1990. Historia Mathematica, 17(4):380, November 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090034B.

Anonymous:1990:XOM

[Ano90-36] Anonymous. XII. Österreichischer Mathematikerkongreß: Vienna, Austria, September 17–22, 1989. Historia Mathematica, 17(2):164– 165, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609090059M.

Anonymous:1990:XIC

[Ano90-37] Anonymous. XVIIIth international Congress of history of science: Hamburg-munich, 1–9 August 1989. Historia Mathematica, 17(2):158–162, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609090056J.

Anonymous:1991:PYC

[Ano91a] Anonymous. A. P. Yushkevich celebrates 85th birthday. Historia Mathematica, 18(4):311, November 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086091903688.

Anonymous:1991:A

[Ano91b] Anonymous. Abstracts. Historia Mathematica, 18(3):296–306, August 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609190407O.

Anonymous:1991:ADM

[Ano91c] Anonymous. Anniversary of the Deutsche Mathematiker-Vereinigung: Bremen, September 16–22, 1990. Historia Mathematica, 18(2):181– 182, May 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609190505R.

Anonymous:1991:BRH

[Ano91d] Anonymous. Book review: History and philosophy of modern mathematics: Volume XI, Minnesota Studies in the Philosophy of Science. Edited by William Aspray and Philip Kitcher. Minneapolis, (University of Minnesota Press). 1988. 387 pages. Historia Mathematica, 18(3):288–291, August 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609190405M.

Anonymous:1991:BRKa

[Ano91e] Anonymous. Book review: Kepler’s geometrical cosmology: J. V. Field. Chicago, IL (The University of Chicago Press). 1988. xx + 243 pp. $37.50. Historia Mathematica, 18(2):188, May 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609190508U.

Anonymous:1991:BRKb

[Ano91f] Anonymous. Book review: Kepler’s Physical astronomy: (Studies in the History of Mathematics and Physical Sciences 13). Bruce Stephenson. New York, NY (Springer-Verlag). 1987. v + 217 pp. $58.00. Historia Mathematica, 18(2):188–194, May 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609190509V.

Anonymous:1991:BRP

[Ano91g] Anonymous. Book review: Popularizing mathematical methods in the People’s Republic of China: Some personal experiences: By Hua Loo-Keng and Wang Yuan. Mathematical Modeling Series, Vol. 2. Boston (Birkhäuser Boston). 1989. 208 pp. Historia Mathematica, 18(3):284–288, August 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609190404L.

Anonymous:1991:CTH

[Ano91h] Anonymous. Contemporary trends in the historiography of science: Korfu, May 27–June 2, 1991. Historia Mathematica, 18(2):178, May 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609190500W.

Anonymous:1991:E

[Ano91i] Anonymous. Erratum. Historia Mathematica, 18(1):89, February 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086091903644.

Anonymous:1991:FEC

[Ano91j] Anonymous. First European congress of mathematics: Paris, July 6–10, 1992. Historia Mathematica, 18(3):262–263, August 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609190395E.

Anonymous:1991:GSS

[Ano91k] Anonymous. Giornata di studio di storia delle matematiche in onore del prof. Gino Arrighi: Modena, October 10, 1990. (Italian) [Day of study of the history of mathematics in honor of Prof. Gino Arrighi: Modena, October 10, 1990]. Historia Mathematica, 18(2):180, May 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609190503P.

Anonymous:1991:IAA

[Ano91l] Anonymous. Index of authors of articles and reviews in volume 18. Historia Mathematica, 18(4):406, November 1991.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609190388E.

Anonymous:1991:IAP

[Ano91m] Anonymous. Index of authors of publications reviewed and abstracted in volume 18. Historia Mathematica, 18(4):407–410, November 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609190389F.

Anonymous:1991:ICC

[Ano91n] Anonymous. International colloquy: Le concept de vitesse en astronomie, en mathématiques, et en physique d’Archimède à Galilée: Hotel Westminster–Concorde, Nice, June 8–10, 1990. Historia Mathematica, 18(2):181, May 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609190504Q.

Anonymous:1991:ICM

[Ano91o] Anonymous. International Congress of Mathematicians: Kyoto, Japan, August 21–29, 1990. The Tokyo History of Mathematics Symposium 1990 Tokyo, Japan, August 31–September 1, 1990. Historia Mathematica, 18(4):353–356, November 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086091903736.

Anonymous:1991:ISS

[Ano91p] Anonymous. International symposium on structures in mathematical theories: Escuela de Magisterio, Campus of Gipuzkoa, Donostia-San Sebastian, Basque Country, Spain, September 25–29, 1990. Historia Mathematica, 18(3):278–279, August 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609190399I.

Anonymous:1991:JCD

[Ano91q] Anonymous. Journées chaos et déterminisme: Association Henri Poincaré pour l’Histoire et la Philosophie des Mathématiques et de la Physique modernes, Institut Henri Poincaré 11, rue Pierre et Marie Curie, Paris 5è, France, 19–20 October 1990. Historia Mathematica, 18(3):281–282, August 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609190401I.

Anonymous:1991:LHS

[Ano91r] Anonymous. Laboratoire d’histoire des sciences et des techniques CNRS-UPR 21 (Paris, France). Historia Mathematica, 18(2):183– 184, May 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609190506S.

Anonymous:1991:LAH

[Ano91s] Anonymous. L’émergence de l’analyse harmonique abstraite, 1930– 1950: Institut Henri Poincaré, 11, rue Pierre et Marie Curie, 75005 Paris, France, March 13, 1991. Historia Mathematica, 18(3):283, August 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609190403K.

Anonymous:1991:MMC

[Ano91t] Anonymous. Mathematics and Mind: a Conference on the Philosophy of Mathematics at Amherst College, April 5–7, 1991. Historia Mathematica, 18(1):55–56, February 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609190351W.

Anonymous:1991:MPI

[Ano91u] Anonymous. Mathematische Probleme im Mittelalter-Der lateinische und arabische Sprachbereich Wolfenbüttel, June 18–22, 1990. Historia Mathematica, 18(1):62–63, February 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609190353Y.

Anonymous:1991:NCa

[Ano91v] Anonymous. Notes on contributors. Historia Mathematica, 18 (1):104–105, February 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086091903666.

Anonymous:1991:NCb

[Ano91w] Anonymous. Notes on contributors. Historia Mathematica, 18(2): 206–207, May 1991. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/0315086091906133.

Anonymous:1991:NCc

[Ano91x] Anonymous. Notes on contributors. Historia Mathematica, 18 (3):307–308, August 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609190493H.

Anonymous:1991:NCd

[Ano91y] Anonymous. Notes on contributors. Historia Mathematica, 18(4): 403–404, November 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609190386C.

Anonymous:1991:PAFa

[Ano91z] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 18(1):106, February 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086091903677.

Anonymous:1991:PAFb

[Ano91-27] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 18(2):208, May 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086091906144.

Anonymous:1991:PAFc

[Ano91-28] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 18(3):309, August 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609190494I.

Anonymous:1991:PAFd

[Ano91-29] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 18(4):405, November 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609190387D.

Anonymous:1991:PPS

[Ano91-30] Anonymous. The proportions of the Pastos sun: Albis, Víctor Samuel. 1987. Boletín de Matemáticas 21(2–3), 110–134. Historia Mathematica, 18(2):195–205, May 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086091905105.

Anonymous:1991:SSR

[Ano91-31] Anonymous. Summer school: On the relation between mathematics and its applications during the 18th and 19th century: Pfalzakademie Lambrecht (Neustadt/Weinstraße), may 14–18, 1990. Historia Mathematica, 18(1):65–66, February 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086091903552.

Anonymous:1991:TGM

[Ano91-32] Anonymous. Tagung zur Geschichte der Mathematik. (German) [Workshop on the history of mathematics]. Historia Mathematica, 18(3):263, August 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609190396F.

Anonymous:1991:TAG

[Ano91-33] Anonymous. Third Annual Göttingen Workshop on the History of Modern Mathematics: Mathematisches Institut de Universität Göttingen, Göttingen, Federal Republic of Germany, July 13– 15, 1992. Historia Mathematica, 18(4):352–353, November 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086091903725.

Anonymous:1992:ECI

[Ano92a] Anonymous. ρème colloque inter-I.R.E.M. Epistémologie et Histoire des Mathématiques: Histoire d’Infinis Brest, May 22–23, 1992. Historia Mathematica, 19(4):427, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086092900154.

Anonymous:1992:OMG

[Ano92b] Anonymous. 30th Oberwolfach Meeting, Geschichte der Mathematik, May 10–16, 1992: Mathematisches Forschungsinstitut Oberwolfach, Lorenzenhof, 7620 Oberwolfach-Walke, Germany. Historia Mathematica, 19(4):426–427, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086092900143.

Anonymous:1992:AEDa

[Ano92c] Anonymous. Abstracts: Edited by David E. Zitarelli. Historia Mathematica, 19(3):328–337, August 1992. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609290048G.

Anonymous:1992:AEDb

[Ano92d] Anonymous. Abstracts: Edited by David E. Zitarelli. Historia Mathematica, 19(4):444–456, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086092900246.

Anonymous:1992:ARC

[Ano92e] Anonymous. Adam Ries colloquy: Adam Ries — Humanist, Rechenmeister, Bergbeamter, Annaberg-Buchhoz, July 18, 1992. Historia Mathematica, 19(4):429–430, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086092900187.

Anonymous:1992:AMS

[Ano92f] Anonymous. American Mathematical Society Annual Winter Meeting: San Francisco, January 16–19, 1991. Historia Mathematica, 19(1):87, February 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609290065J.

Anonymous:1992:AMI

[Ano92g] Anonymous. Ancient mathematics in Islamic and Occidental cultures: Tradition, transmission, transformation University of Oklahoma, March 6–7, 1992. Historia Mathematica, 19(4):423, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290011Y.

Anonymous:1992:AMD

[Ano92h] Anonymous. Annual meeting of the Deutsche Mathematiker-Vereinigung Section History of Mathematics: Bielefeld, Germany, October 16–20, 1991. Historia Mathematica, 19(2):211, May 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290087R.

Anonymous:1992:BPA

[Ano92i] Anonymous. Bibliography of publications of Aḣmad Salim Sa’idan (1914–1991) on the history of mathematics and astronomy in Islamic civilization, and list of medieval Arabic texts published by him. Historia Mathematica, 19(4):439–443, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086092900235.

Anonymous:1992:BRS

[Ano92j] Anonymous. Book review: Die Streitschriften von Jacob und Johann Bernoulli, variationsrechnung: Bearbeitet und kommentiert von Herman H. Goldstine, mit historischen Anmerkungen von Patricia Radelet-de Grave. Basel, Boston, Berlin (Birkhäuser). 1991. xxv + 604 pp. Historia Mathematica, 19(3):323–324, August 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290046E.

Anonymous:1992:BRF

[Ano92k] Anonymous. Book review: La France mathématique. La société mathématique de France (1870–1914): By H. Gispert, with essays by R. Bkouche, C. Gilain, C. Houzel, J.-P. Kahane, and M. Zerner. Cahiers d’Histoire et de Philosophie des Sciences, new series, No. 34. Paris (Belin). 1991. 425 pp. 95Ffr. Historia Mathematica, 19(3):319– 322, August 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609290044C.

Anonymous:1992:BRM

[Ano92l] Anonymous. Book review: Mathematische schriften: 7th Ser., Vol. 1. By G. W. Leibniz; edited by Eberhard Knobloch and Walter S. Contro. Berlin (Akademie-Verlag). 1990. 954 + xxxix pp. Historia Mathematica, 19(4):435–437, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086092900213.

Anonymous:1992:BRN

[Ano92m] Anonymous. Book review: Norbert Wiener: 1894–1964: By P. R. Masani, Basel, Boston, Berlin (Birkhäuser). 1990. 416 pp. Historia Mathematica, 19(3):316–319, August 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609290043B.

Anonymous:1992:BRI

[Ano92n] Anonymous. Book review: The invisible hand: Economic equilibrium in the history of science: By Bruna Ingrao and Giorgio Israel. Trans. Ian McGilvray. Cambridge, MA (The MIT Press). 1990. xiii + 491 pp. $49.95. Historia Mathematica, 19(3):322–323, August 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290045D.

Anonymous:1992:BRT

[Ano92o] Anonymous. Book review: The taming of chance: Ian Hacking. Cambridge/New York (Cambridge University Press). 1990. xiii + 264 pp. $60.00. Historia Mathematica, 19(4):431–435, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290020C.

Anonymous:1992:CIE

[Ano92p] Anonymous. Colloque international: Les Enfants du Siècle: Sciences et savants dans la France de l’époque romantique (1815–1830). Historia Mathematica, 19(4):420–421, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609290008Y.

Anonymous:1992:CDM

[Ano92q] Anonymous. Colloque: Le développement des mathématiques entre 1900 et 1950: Château de Bourglinster, Luxembourg, June 29–30, 1992. Historia Mathematica, 19(4):429, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086092900176.

Anonymous:1992:CPQ

[Ano92r] Anonymous. Colloque pour le quatrième centenaire de Girard Desargues (1591–1661). Historia Mathematica, 19(3):293–294, August 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290036B.

Anonymous:1992:CID

[Ano92s] Anonymous. Colloquio internazionale: Dalla geometria al calcolo 1637–1704: Palazzo Serra di Cassano, Via Monte di Dio, 14, Naples, Italy, October 29–31, 1991. Historia Mathematica, 19(2):211–212, May 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609290088S.

Anonymous:1992:CHG

[Ano92t] Anonymous. Conference on Hermann G. Grassmann in 1994. Historia Mathematica, 19(1):81, February 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609290060O.

Anonymous:1992:DMC

[Ano92u] Anonymous. Development of mathematics from 1900 to 1950: Château de Bourglinster, June 29–30, 1992. Historia Mathematica, 19(1):79, February 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609290058J.

Anonymous:1992:FMS

[Ano92v] Anonymous. Fourteenth Meeting of the Study Group “History, Philosophy and Foundations of Mathematics”: Gosen (near Berlin), October 14–16, 1991. Historia Mathematica, 19(2):210, May 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290086Q.

Anonymous:1992:FAG

[Ano92w] Anonymous. Fourth Annual Göttingen Workshop on the History of Modern Mathematics: Mathematisches Institut der Georg-August-Universität Göttingen, July 12–14, 1993. Historia Mathematica, 19(4):421, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609290009Z.

Anonymous:1992:FMM

[Ano92x] Anonymous. Fourth Midwest Math History Conference: Miami University, Oxford, Ohio, October 2–3, 1992. Historia Mathematica, 19 (2):204, May 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609290079Q.

Anonymous:1992:IAA

[Ano92y] Anonymous. Index of authors of articles and reviews in volume 19. Historia Mathematica, 19(4):460, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086092900279.

Anonymous:1992:IAP

[Ano92z] Anonymous. Index of authors of publications reviewed and abstracted in volume 19. Historia Mathematica, 19(4):461–464, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290028A.

Anonymous:1992:ITM

[Ano92-27] Anonymous. Innovation et transmission dans les mathématiques arabes (VIII^{e}–XV^{e} siècles): Institut du Monde Arabe, 1, rue des Fossès–Saint-Bernard 75005 Paris, France, 25–27 octobre 1990. Historia Mathematica, 19(1):83, February 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609290062G.

Anonymous:1992:ICT

[Ano92-28] Anonymous. International colloquy: Le travail scientifique dans les correspondances entre savants au tournant des 17e et 18e siècles Auditorium de la Bibliothèque nationale, 2 rue Vivienne, 75002 Paris, France, June 10–13, 1992. Historia Mathematica, 19(4):428–429, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086092900165.

Anonymous:1992:ICC

[Ano92-29] Anonymous. International conference: Contemporary trends in the historiography of science: Korfu, may 27–31, 1991. Historia Mathematica, 19(2):205–206, May 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609290081L.

Anonymous:1992:ISE

[Ano92-30] Anonymous. International Symposium on Early Greek Mathematics: University of Athens, Athens, Greece, August 17–21, 1992. Historia Mathematica, 19(4):430, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086092900198.

Anonymous:1992:ISA

[Ano92-31] Anonymous. International Symposium on the 100th Anniversaries of the Births of Li Yan and Qian Baozong: Beijing, August 21–23, 1992. Historia Mathematica, 19(2):203–204, May 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290078P.

Anonymous:1992:JSE

[Ano92-32] Anonymous. Les jésuites et les sciences dans l’Europe de la Renaissance (XVIe–XVIIe s.). Historia Mathematica, 19(3):292, August 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086092900349.

Anonymous:1992:LSY

[Ano92-33] Anonymous. Luncheon and symposium: 20 years of Historia Mathematica. Historia Mathematica, 19(3):290–291, August 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086092900316.

Anonymous:1992:MMO

[Ano92-34] Anonymous. Mathematics and music: Oxford, May 2–3, 1992. Historia Mathematica, 19(4):425, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086092900132.

Anonymous:1992:MRH

[Ano92-35] Anonymous. Mythes et Réalités historiques de l’Europe mathématique: Paris, April 3–6, 1992. Historia Mathematica, 19(4):424–425, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609290012Z.

Anonymous:1992:NCa

[Ano92-36] Anonymous. Notes on contributors. Historia Mathematica, 19 (3):338–339, August 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609290049H.

Anonymous:1992:NCb

[Ano92-37] Anonymous. Notes on contributors. Historia Mathematica, 19 (4):457–458, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086092900257.

Anonymous:1992:ODS

[Ano92-38] Anonymous. The Oxford dictionary of scientific quotation. Historia Mathematica, 19(3):302, August 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609290039E.

Anonymous:1992:PAFa

[Ano92-39] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 19(1):124, February 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609290072J.

Anonymous:1992:PAFb

[Ano92-40] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 19(2):231, May 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609290093Q.

Anonymous:1992:PAFc

[Ano92-41] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 19(3):340, August 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609290050L.

Anonymous:1992:PAFd

[Ano92-42] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 19(4):459, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086092900268.

Anonymous:1992:RM

[Ano92-43] Anonymous. Renaissance mathematics. Historia Mathematica, 19(3):291–292, August 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086092900338.

Anonymous:1992:RIH

[Ano92-44] Anonymous. Rencontre internationale Histoire des Mathématiques: Universidad Complutense, Facultad de Matématicas (edificio nuevo), Ciudad Universitaria, Madrid, Spain, November 18–22, 1991. Historia Mathematica, 19(2):213–214, May 1992. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609290089T.

Anonymous:1992:RP

[Ano92-45] Anonymous. Research in progress. Historia Mathematica, 19 (3):294–295, August 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609290037C.

Anonymous:1992:SMF

[Ano92-46] Anonymous. Second meeting of the Fachsektion Geschichte der Mathematik der Deutschen Mathematiker-Vereinigung. Historia Mathematica, 19(3):291, August 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086092900327.

Anonymous:1992:TAG

[Ano92-47] Anonymous. Third Annual Göttingen Workshop on the History of Modern Mathematics: July 13–15, 1992. Historia Mathematica, 19 (4):422–423, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086092900109.

Anonymous:1992:TAS

[Ano92-48] Anonymous. Third Austrian Symposion on the History of Mathematics: Neuhofen, November 8–14, 1992. Historia Mathematica, 19 (2):205, May 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609290080U.

Anonymous:1992:TMS

[Ano92-49] Anonymous. Third Maghribian Symposium on the History of Arabic Mathematics: Algiers, December 1–3, 1990. Historia Mathematica, 19(1):83–84, February 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609290063H.

Anonymous:1992:CIT

[Ano92-50] Anonymous. Un colloque international: Le travail scientifique dans les correspondances entre savants au tournant des 17^{e} et 18^{e} siècles: Auditorium de la Bibliothèque National, Paris, 10–13 juin 1992. Historia Mathematica, 19(1):79–80, February 1992. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609290059K.

Anonymous:1992:WDH

[Ano92-51] Anonymous. World directory of historians of mathematics: Third edition in preparation. Historia Mathematica, 19(1):88–89, February 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290066K.

Anonymous:1992:ZNG

[Ano92-52] Anonymous. Zweite Novembertagung zur Geschichte der modernen Mathematik. (German) [Second November meeting on the history of modern mathematics]. Historia Mathematica, 19(3):292–293, August 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290035A.

Anonymous:1993:SSC

[Ano93a] Anonymous. 14th summer school of Czech historians of mathematics: Manetín–BRDO, Czechia, May 31–June 4, 1993. Historia Mathematica, 20(4):434, November 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008800237.

Anonymous:1993:Aa

[Ano93b] Anonymous. Abstracts. Historia Mathematica, 20(1):104–116, February 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008800596.

Anonymous:1993:Ab

[Ano93c] Anonymous. Abstracts. Historia Mathematica, 20(2):223–235, May 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008800511.

Anonymous:1993:Ac

[Ano93d] Anonymous. Abstracts. Historia Mathematica, 20(3):324–339, August 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008800031.

Anonymous:1993:Ad

[Ano93e] Anonymous. Abstracts. Historia Mathematica, 20(4):447–459, November 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008800262.

Anonymous:1993:BRB

[Ano93f] Anonymous. Book review: Before Writing. Vol. 1. From Counting to Cuneiform: By Denise Schmandt-Besserat. Austin, TX (University of Texas Press). 1992. xviii + 269 pp. 175 photos, 43 line drawings, 6 tables, 16 charts. $60.00. Historia Mathematica, 20(2):220–222, May 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086083710190.

Anonymous:1993:BRJ

[Ano93g] Anonymous. Book review: Joseph Lionville 1809–1882. Master of Pure and Applied Mathematics: Jesper Lützen. New York/Berlin/ Heidelberg/London/Paris/Tokyo/Hong Kong (Springer-Verlag). 1990. xix + 884 pp. $98.00. Historia Mathematica, 20(2):205–210, May 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008800481.

Anonymous:1993:BRM

[Ano93h] Anonymous. Book review: The Mathematical Philosophy of Bertrand Russell: Origins and Development: By Francisco Rodriguez-Consuegra. Boston, Basel, Berlin (Birkhäuser). 1991. xvi + 236 pp. Historia Mathematica, 20(2):216–220, May 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508600880050X.

Anonymous:1993:BRS

[Ano93i] Anonymous. Book review: The Space of Mathematics. Philosophical, Epistemological, and Historical Explorations: Edited by Javier Echeverria, Andoni Ibarra, and Thomas Mormann. Berlin, New York (de Gruyter). 1992. xvi + 422 pp. With an index of names. Historia Mathematica, 20(2):210–215, May 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008800493.

Anonymous:1993:CCA

[Ano93j] Anonymous. Christoph Clavius e l’attivitá scientifica dei jesuiti nell’ etá di Galileo. International Conference, April 28–30, 1993. Historia Mathematica, 20(3):317, August 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508600880002X.

Anonymous:1993:DMV

[Ano93k] Anonymous. Deutsche Mathematiker-Vereinigung: Annual meeting: Berlin, September 14–18, 1992. Historia Mathematica, 20(1):96–97, February 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008800572.

Anonymous:1993:E

[Ano93l] Anonymous. Erratum. Historia Mathematica, 20(4):460–463, November 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086083710360.

Anonymous:1993:EMG

[Ano93m] Anonymous. European Mathematics 1848–1939: Gonville and Caius College, Cambridge, September 11–13, 1992. Historia Mathematica, 20(1):95–96, February 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008800559.

Anonymous:1993:FIC

[Ano93n] Anonymous. Fourth international colloquium on the philosophy and history of mathematics. First Colombian colloquium on the history, philosophy and pedagogy of mathematics. Historia Mathematica, 20 (2):198, May 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086083710177.

Anonymous:1993:GNA

[Ano93o] Anonymous. GEHIMAB: a new association for history of mathematics. Historia Mathematica, 20(2):204, May 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508600880047X.

Anonymous:1993:IGA

[Ano93p] Anonymous. Iberia in the Golden Age: Mathematical sciences and their uses, 1500–1700. Historia Mathematica, 20(2):202–203, May 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008800468.

Anonymous:1993:IAA

[Ano93q] Anonymous. Index of authors of articles and reviews in volume 20. Historia Mathematica, 20(4):466, November 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086083710372.

Anonymous:1993:IAP

[Ano93r] Anonymous. Index of authors of publications reviewed and abstracted in volume 20. Historia Mathematica, 20(4):467–471, November 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008800298.

Anonymous:1993:ICA

[Ano93s] Anonymous. International congress for the advancement of science and technology in the Islamic World: Teheran University, Teheran, Iran, May 17–19, 1993. Historia Mathematica, 20(4):435, November 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008800249.

Anonymous:1993:ICH

[Ano93t] Anonymous. International congress Henri Poincaré. Historia Mathematica, 20(2):199, May 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008800420.

Anonymous:1993:KGC

[Ano93u] Anonymous. Kurt Gödel’s collected works. Historia Mathematica, 20(4):436, November 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008800250.

Anonymous:1993:QCM

[Ano93v] Anonymous. Le quatrième colloque maghrébin international sur l’histoire des mathématiques arabes. Historia Mathematica, 20(2): 199–200, May 1993. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086008800432.

Anonymous:1993:MIG

[Ano93w] Anonymous. Mathematics in Iberia in the Golden Age: 31 March– 3 April 1993. Historia Mathematica, 20(1):95, February 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008800547.

Anonymous:1993:MBN

[Ano93x] Anonymous. Mathematik am Beginn der Neuzeit International Colloquy: Sächsische Akademie der Wissenschaften, Leipzig, Germany, October 15, 1992. Historia Mathematica, 20(1):96, February 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008800560.

Anonymous:1993:NCa

[Ano93y] Anonymous. Notes on contributors. Historia Mathematica, 20 (1):117–118, February 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008800602.

Anonymous:1993:NCb

[Ano93z] Anonymous. Notes on contributors. Historia Mathematica, 20(2): 236–237, May 1993. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086008800523.

Anonymous:1993:NCc

[Ano93-27] Anonymous. Notes on contributors. Historia Mathematica, 20 (3):340–341, August 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008800043.

Anonymous:1993:NCd

[Ano93-28] Anonymous. Notes on contributors. Historia Mathematica, 20 (4):464, November 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008800274.

Anonymous:1993:PAFa

[Ano93-29] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 20(1):119, February 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008800614.

Anonymous:1993:PAFb

[Ano93-30] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 20(2):238, May 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008800535.

Anonymous:1993:PAFc

[Ano93-31] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 20(3):342, August 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008800055.

Anonymous:1993:PAFd

[Ano93-32] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 20(4):465, November 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008800286.

Anonymous:1993:PUD

[Ano93-33] Anonymous. Première Université d’Eté Européenne: Histoire des mathématiques. Historia Mathematica, 20(2):199, May 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086083710189.

Anonymous:1993:TAS

[Ano93-34] Anonymous. Third Austrian Symposium on the History of Mathematics. Historia Mathematica, 20(2):201, May 1993. CODEN HI-MADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/S0315086008800444.

Anonymous:1994:GS

[Ano94a] Anonymous. 2nd Gauss symposium. Historia Mathematica, 21(1):75, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008800651.

Anonymous:1994:NGM

[Ano94b] Anonymous. 4. Novembertagung zur Geschichte der modernen Mathematik: Universität Mainz, Deutschland, 31. Oktober-2. November 1993. Historia Mathematica, 21(2):207, May 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008800110.

Anonymous:1994:Aa

[Ano94c] Anonymous. Abstracts. Historia Mathematica, 21(1):97–116, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008800705.

Anonymous:1994:Ab

[Ano94d] Anonymous. Abstracts. Historia Mathematica, 21(2):225–240, May 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008800146.

Anonymous:1994:Ac

[Ano94e] Anonymous. Abstracts. Historia Mathematica, 21(3):386–396, August 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008800304.

Anonymous:1994:Ad

[Ano94f] Anonymous. Abstracts. Historia Mathematica, 21(4):469–482, November 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008800067.

Anonymous:1994:BSH

[Ano94g] Anonymous. Beijing seminar on the history of mathematics: Institute of Mathematics, Academia Sinica, October 11–15, 1993. Historia Mathematica, 21(2):208, May 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008800134.

Anonymous:1994:BPB

[Ano94h] Anonymous. Bibliography of publications by B. L. van der Waerden since 1983 on ancient and medieval science. Historia Mathematica, 21(1):71–72, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086084710093.

Anonymous:1994:FAG

[Ano94i] Anonymous. Fourth Annual Göttingen Workshop on the History of Modern Mathematics. Historia Mathematica, 21(1):85–87, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008800699.

Anonymous:1994:GGB

[Ano94j] Anonymous. George Green: a bicentennial celebration. Historia Mathematica, 21(1):76, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008800663.

Anonymous:1994:IAA

[Ano94k] Anonymous. Index of authors of articles and reviews in volume 21. Historia Mathematica, 21(4):486, November 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086084710366.

Anonymous:1994:IAP

[Ano94l] Anonymous. Index of authors of publications reviewed and abstracted in volume 21. Historia Mathematica, 21(4):487–491, November 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008800092.

Anonymous:1994:Ma

[Ano94m] Anonymous. Meetings. Historia Mathematica, 21(1):73, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008800626.

Anonymous:1994:Mb

[Ano94n] Anonymous. Meetings. Historia Mathematica, 21(2):206, May 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008800109.

Anonymous:1994:NCa

[Ano94o] Anonymous. Notes on contributors. Historia Mathematica, 21 (1):117–118, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008800717.

Anonymous:1994:NCb

[Ano94p] Anonymous. Notes on contributors. Historia Mathematica, 21(2): 241–242, May 1994. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086008800158.

Anonymous:1994:NCc

[Ano94q] Anonymous. Notes on contributors. Historia Mathematica, 21 (3):397–398, August 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008800316.

Anonymous:1994:NCd

[Ano94r] Anonymous. Notes on contributors. Historia Mathematica, 21(4): 483–484, November 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008800079.

Anonymous:1994:PAFa

[Ano94s] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 21(1):119, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008800729.

Anonymous:1994:PAFb

[Ano94t] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 21(2):243, May 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S031508600880016X.

Anonymous:1994:PAFc

[Ano94u] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 21(3):399, August 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008800328.

Anonymous:1994:PAFd

[Ano94v] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 21(4):485, November 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008800080.

Anonymous:1994:SS

[Ano94w] Anonymous. Scientific sections. Historia Mathematica, 21(1):84–85, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008800687.

Anonymous:1994:XIC

[Ano94x] Anonymous. XIXth International Congress of History of Science. Historia Mathematica, 21(1):76–84, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008800675.

Anonymous:1995:Aa

[Ano95a] Anonymous. Abstracts. Historia Mathematica, 22(1):93–109, February 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008800171.

Anonymous:1995:Ab

[Ano95b] Anonymous. Abstracts. Historia Mathematica, 22(2):210–220, May 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008800341.

Anonymous:1995:Ac

[Ano95c] Anonymous. Abstracts. Historia Mathematica, 22(3):331–341, ???? 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008800201.

Anonymous:1995:Ad

[Ano95d] Anonymous. Abstracts. Historia Mathematica, 22(4):449–469, November 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008800377.

Anonymous:1995:AAP

[Ano95e] Anonymous. Adolph Andrei Pavlovich Yushkevich (1906–1993). Historia Mathematica, 22(2):??, May 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086085710129.

Anonymous:1995:BRA

[Ano95f] Anonymous. Book review: Amphora. Festschrift fur Hans Wussing zu seinem 65. Geburtstag/Festschrift for Hans Wussing on the Occasion of His 65th Birthday: Edited by S. S. Demidov, M. Folkerts, D. E. Rowe, and C. J. Scriba. Basel/Boston/Berlin (Birkhäauser Verlag). 1992. xi + 782 pp. U.S. $140.00. Historia Mathematica, 22 (4):425–441, November 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086085710348.

Anonymous:1995:BRB

[Ano95g] Anonymous. Book review: Berkeley’s Philosophy of Mathematics: By Douglas M. Jesseph. Chicago and London (The University of Chicago Press). 1993. xii + 322 pp. Historia Mathematica, 22(2):206–208, May 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508608571021X.

Anonymous:1995:BRE

[Ano95h] Anonymous. Book review: Die Entwicklung des Tensorkalküls. Vom absoluten Differentialkalkül zur Relativitätstheorie: By Karin Reich. Basel/Boston/Berlin (Birkhäuser Verlag). 1994. 331 pp. Historia Mathematica, 22(3):323–326, ???? 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086085710270.

Anonymous:1995:BRMb

[Ano95i] Anonymous. Book review: La mesure de l’état: By Eric Brian. Paris (Albin Michel). 1994. 464 pp. Historia Mathematica, 22(3):327– 330, ???? 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086085710282.

Anonymous:1995:BRL

[Ano95j] Anonymous. Book review: Lakatos’ Philosophy of Mathematics: a Historical Approach: By Tenn Koetsier. Studies in the History and Philosophy of Mathematics. Volume 3. Amsterdam (North-Holland, Elsevier). 1991. 312 pages, including bibliography and index. Historia Mathematica, 22(2):203–206, May 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086085710208.

Anonymous:1995:BRMa

[Ano95k] Anonymous. Book review: Math Worlds: Philosophical and Social Studies of Mathematics and Mathematics Education: Edited by Sal Restivo, Jean Paul Van Bendegem, and Roland Fischer. Albany (State University of New York Press). 1993. 278 pp. Name and subject indexes. Historia Mathematica, 22(2):208–209, May 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086085710221.

Anonymous:1995:BRSb

[Ano95l] Anonymous. Book review: On the Shoulders of Merchants: Exchange and the Mathematical Conception of Nature in Early Modern Europe: By Richard W. Hadden. Albany, NY (State Univ. of New York Press). 1994. xviii + 191 pp. with notes, bibliography, and index. Historia Mathematica, 22(4):446–448, November 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086085710361.

Anonymous:1995:BRSa

[Ano95m] Anonymous. Book review: Sofia Kowalewskaja. Ein Leben für Mathematik and Emanzipation: By Wilderich Tuschmann and Peter Hawig. Basel/Boston/Berlin (Birkhduser). 1993. 185 pp. with 22 Figs. 58 DM. Historia Mathematica, 22(4):442–446, November 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508608571035X.

Anonymous:1995:SIA

[Ano95n] Anonymous. La Sphère, instrument au service de la découverte du monde: D’Autolycos de Pitane à Jean de Sacrobosco: Par Germaine Aujac. Collection Varia. Caen (Paradigme). 1993. 380 pp. Historia Mathematica, 22(2):196–202, May 1995. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086085710191.

Anonymous:1995:GH

[Ano95o] Anonymous. Guy Hirsch (1915–1993). Historia Mathematica, 22 (4):345–346, November 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086085710294.

Anonymous:1995:IAA

[Ano95p] Anonymous. Index of authors of articles and reviews in volume 22. Historia Mathematica, 22(4):474, November 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086085710373.

Anonymous:1995:IAP

[Ano95q] Anonymous. Index of authors of publications reviewed and abstracted in volume 22. Historia Mathematica, 22(4):475–480, November 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008800419.

Anonymous:1995:M

[Ano95r] Anonymous. Meetings. Historia Mathematica, 22(2):194–195, May 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508600880033X.

Anonymous:1995:NCb

[Ano95s] Anonymous. Notes on contributors. Historia Mathematica, 22(2): 221–222, May 1995. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086008800353.

Anonymous:1995:NCc

[Ano95t] Anonymous. Notes on contributors. Historia Mathematica, 22(3): 342, ???? 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008800213.

Anonymous:1995:NCd

[Ano95u] Anonymous. Notes on contributors. Historia Mathematica, 22(4): 471–472, November 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008800390.

Anonymous:1995:NCa

[Ano95v] Anonymous. Notes to contributors. Historia Mathematica, 22 (1):110–111, February 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008800183.

Anonymous:1995:PAFa

[Ano95w] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 22(1):112, February 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008800195.

Anonymous:1995:PAFb

[Ano95x] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 22(2):223, May 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008800365.

Anonymous:1995:PAFc

[Ano95y] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 22(3):343, ???? 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008800225.

Anonymous:1995:PAFd

[Ano95z] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 22(4):473, November 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008800407.

Anonymous:1995:WDH

[Ano95-27] Anonymous. World directory of historians of mathematics. Historia Mathematica, 22(4):470, November 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008800389.

Anonymous:1996:A

[Ano96a] Anonymous. Abstracts. Historia Mathematica, 23(2):217–235, May 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086096900221.

Anonymous:1996:IAA

[Ano96b] Anonymous. Index of authors of articles and reviews in volume 23. Historia Mathematica, 23(4):468, November 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086096900476.

Anonymous:1996:IAP

[Ano96c] Anonymous. Index of authors of publications reviewed and abstracted in volume 23. Historia Mathematica, 23(4):469–473, November 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086096900488.

Anonymous:1996:NCa

[Ano96d] Anonymous. Notes on contributors. Historia Mathematica, 23 (1):114–115, February 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086096900129.

Anonymous:1996:NCb

[Ano96e] Anonymous. Notes on contributors. Historia Mathematica, 23(2): 236–237, May 1996. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086096900233.

Anonymous:1996:NCc

[Ano96f] Anonymous. Notes on contributors. Historia Mathematica, 23 (3):349–350, August 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S031508609690035X.

Anonymous:1996:NCd

[Ano96g] Anonymous. Notes on contributors. Historia Mathematica, 23(4): 465–466, November 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086096900452.

Anonymous:1996:PAFa

[Ano96h] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 23(1):116, February 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086096900130.

Anonymous:1996:PAFb

[Ano96i] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 23(2):238, May 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086096900245.

Anonymous:1996:PAFc

[Ano96j] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 23(3):351, August 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086096900361.

Anonymous:1996:PAFd

[Ano96k] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 23(4):467, November 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086096900464.

Anonymous:1997:A

[Ano97a] Anonymous. Abstracts. Historia Mathematica, 24(4):457–477, November 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086097921753.

Anonymous:1997:IAA

[Ano97b] Anonymous. Index of authors of articles and reviews in volume 24. Historia Mathematica, 24(4):481, November 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086097921807.

Anonymous:1997:IAP

[Ano97c] Anonymous. Index of authors of publications reviewed and abstracted in volume 24. Historia Mathematica, 24(4):482–487, November 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086097921819.

Anonymous:1997:NCa

[Ano97d] Anonymous. Notes on contributors. Historia Mathematica, 24 (1):121–122, February 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086097921534.

Anonymous:1997:NCb

[Ano97e] Anonymous. Notes on contributors. Historia Mathematica, 24(2): 238–239, May 1997. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086097921649.

Anonymous:1997:NCc

[Ano97f] Anonymous. Notes on contributors. Historia Mathematica, 24 (3):356–357, August 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086097921662.

Anonymous:1997:NCd

[Ano97g] Anonymous. Notes on contributors. Historia Mathematica, 24(4): 478–479, November 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086097921789.

Anonymous:1997:PAFa

[Ano97h] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 24(1):123, February 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086097921546.

Anonymous:1997:PAFb

[Ano97i] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 24(3):358, August 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086097921674.

Anonymous:1997:PAFc

[Ano97j] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 24(4):480, November 1997. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086097921790.

Anonymous:1998:Aa

[Ano98a] Anonymous. Abstracts. Historia Mathematica, 25(1):98–119, February 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508609792190X.

Anonymous:1998:Ab

[Ano98b] Anonymous. Abstracts. Historia Mathematica, 25(2):225–241, May 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086098922029.

Anonymous:1998:Ac

[Ano98c] Anonymous. Abstracts. Historia Mathematica, 25(3):340–354, August 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086098922108.

Anonymous:1998:Ad

[Ano98d] Anonymous. Abstracts. Historia Mathematica, 25(4):431–448, November 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086098922194.

Anonymous:1998:IAA

[Ano98e] Anonymous. Index of authors of articles and reviews in volume 25. Historia Mathematica, 25(4):451, November 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086098922248.

Anonymous:1998:IAP

[Ano98f] Anonymous. Index of authors of publications reviewed and abstracted in volume 25. Historia Mathematica, 25(4):452–458, November 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508609892225X.

Anonymous:1998:NCa

[Ano98g] Anonymous. Notes on contributors. Historia Mathematica, 25 (1):120–121, February 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086098922017.

Anonymous:1998:NCb

[Ano98h] Anonymous. Notes on contributors. Historia Mathematica, 25(2): 242, May 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086098922078.

Anonymous:1998:NCc

[Ano98i] Anonymous. Notes on contributors. Historia Mathematica, 25 (3):355–356, August 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086098922157.

Anonymous:1998:NCd

[Ano98j] Anonymous. Notes on contributors. Historia Mathematica, 25(4): 449–450, November 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086098922236.

Anonymous:1998:PAFa

[Ano98k] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 25(1):122, February 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086098922005.

Anonymous:1998:PAFb

[Ano98l] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 25(2):243, May 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S031508609892208X.

Anonymous:1998:PAFc

[Ano98m] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 25(3):357, August 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086098922169.

Anonymous:1999:IAAb

[Ano99a] Anonymous. Index of authors of articles and publications reviewed and abstracted in volume 26. Historia Mathematica, 26(4):413–418, November 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086099922723.

Anonymous:1999:IAAa

[Ano99b] Anonymous. Index of authors of articles and reviews in volume 26. Historia Mathematica, 26(4):412, November 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086099922711.

Anonymous:1999:NCa

[Ano99c] Anonymous. Notes on contributors. Historia Mathematica, 26(1): 104–105, February 1, 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S031508609892239X.

Anonymous:1999:NCb

[Ano99d] Anonymous. Notes on contributors. Historia Mathematica, 26(2): 199–200, May 1999. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086099922449.

Anonymous:1999:NCc

[Ano99e] Anonymous. Notes on contributors. Historia Mathematica, 26 (3):309–310, August 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086099922553.

Anonymous:1999:NCd

[Ano99f] Anonymous. Notes on contributors. Historia Mathematica, 26(4):410, November 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508609992267X.

Anonymous:1999:PAF

[Ano99g] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 26(4):411, November 1999. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086099922681.

Anonymous:2000:IAAb

[Ano00a] Anonymous. Index of authors of articles and publications reviewed and abstracted in volume 27. Historia Mathematica, 27(4):466–472, November 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086000922997.

Anonymous:2000:IAAa

[Ano00b] Anonymous. Index of authors of articles and reviews in volume 27. Historia Mathematica, 27(4):465, November 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086000922973.

Anonymous:2000:NCa

[Ano00c] Anonymous. Notes on contributors. Historia Mathematica, 27 (1):105–106, February 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086000922754.

Anonymous:2000:NCb

[Ano00d] Anonymous. Notes on contributors. Historia Mathematica, 27(2): 222, May 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086000922791.

Anonymous:2000:NCc

[Ano00e] Anonymous. Notes on contributors. Historia Mathematica, 27(3): 330, August 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086000922912.

Anonymous:2000:NCd

[Ano00f] Anonymous. Notes on contributors. Historia Mathematica, 27(4):463, November 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086000922961.

Anonymous:2000:PAF

[Ano00g] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 27(4):464, November 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086000922985.

Anonymous:2001:IAAb

[Ano01a] Anonymous. Index of authors of articles and publications reviewed and abstracted in volume 28. Historia Mathematica, 28(4):337–343, November 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086001923420.

Anonymous:2001:IAAa

[Ano01b] Anonymous. Index of authors of articles and reviews in volume 28. Historia Mathematica, 28(4):336, November 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086002923414.

Anonymous:2001:NCa

[Ano01c] Anonymous. Notes on contributors. Historia Mathematica, 28(1):72, February 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086000923085.

Anonymous:2001:NCb

[Ano01d] Anonymous. Notes on contributors. Historia Mathematica, 28(2): 164–165, May 2001. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S031508600092319X.

Anonymous:2001:NCc

[Ano01e] Anonymous. Notes on contributors. Historia Mathematica, 28 (3):281–282, August 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086001923262.

Anonymous:2001:NCd

[Ano01f] Anonymous. Notes on contributors. Historia Mathematica, 28(4):334, November 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086002923396.

Anonymous:2001:PAF

[Ano01g] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 28(4):335, November 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086002923402.

Anonymous:2002:Aa

[Ano02a] Anonymous. Abstracts. Historia Mathematica, 29(1):74–97, February 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086001923353.

Anonymous:2002:Ab

[Ano02b] Anonymous. Abstracts. Historia Mathematica, 29(2):210–230, May 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086002923505.

Anonymous:2002:Ac

[Ano02c] Anonymous. Abstracts. Historia Mathematica, 29(3):340–360, August 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086002923645.

Anonymous:2002:IAAb

[Ano02d] Anonymous. Index of authors of articles and publications reviewed and abstracted in volume 29. Historia Mathematica, 29(4):516–523, November 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086002923815.

Anonymous:2002:IAAa

[Ano02e] Anonymous. Index of authors of articles and reviews in volume 29. Historia Mathematica, 29(4):515, November 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086002923803.

Anonymous:2002:LE

[Ano02f] Anonymous. Letter to the Editor. Historia Mathematica, 29(3):235– 238, August 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086002923554.

Anonymous:2002:NCa

[Ano02g] Anonymous. Notes on contributors. Historia Mathematica, 29(1):98, February 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086002923475.

Anonymous:2002:NCb

[Ano02h] Anonymous. Notes on contributors. Historia Mathematica, 29(2): 231–232, May 2002. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086002923542.

Anonymous:2002:NCc

[Ano02i] Anonymous. Notes on contributors. Historia Mathematica, 29(3): 361, August 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086002923669.

Anonymous:2002:NCd

[Ano02j] Anonymous. Notes on contributors. Historia Mathematica, 29(4): 511–513, November 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086002923785.

Anonymous:2002:PAFa

[Ano02k] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 29(1):99, February 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086002923487.

Anonymous:2002:PAFb

[Ano02l] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 29(2):233, May 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086002923530.

Anonymous:2002:PAFc

[Ano02m] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 29(3):362, August 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086002923670.

Anonymous:2002:PAFd

[Ano02n] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 29(4):514, November 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086002923797.

Anonymous:2003:AEGa

[Ano03a] Anonymous. Abstracts: Edited by Glen Van Brummelen. Historia Mathematica, 30(1):97–119, February 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086002000320.

Anonymous:2003:AEGb

[Ano03b] Anonymous. Abstracts: Edited by Glen Van Brummelen. Historia Mathematica, 30(2):227–252, May 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086003000272.

Anonymous:2003:AEGc

[Ano03c] Anonymous. Abstracts: Edited by Glen Van Brummelen. Historia Mathematica, 30(3):381–403, August 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508600300034X.

Anonymous:2003:AEGd

[Ano03d] Anonymous. Abstracts: Edited by Glen Van Brummelen. Historia Mathematica, 30(4):522–551, November 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086003000752.

Anonymous:2003:Ca

[Ano03e] Anonymous. Contents. Historia Mathematica, 30(1):??, February 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086003000107.

Anonymous:2003:Cb

[Ano03f] Anonymous. Contents. Historia Mathematica, 30(2):??, May 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086003000375.

Anonymous:2003:Cc

[Ano03g] Anonymous. Contents. Historia Mathematica, 30(3):??, August 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508600300065X.

Anonymous:2003:Cd

[Ano03h] Anonymous. Contents. Historia Mathematica, 30(4):??, November 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086003000867.

Anonymous:2003:EBa

[Ano03i] Anonymous. Editorial Board. Historia Mathematica, 30(1):??, February 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086003000090.

Anonymous:2003:EBb

[Ano03j] Anonymous. Editorial Board. Historia Mathematica, 30(2):??, May 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086003000363.

Anonymous:2003:EBc

[Ano03k] Anonymous. Editorial Board. Historia Mathematica, 30(3):??, August 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086003000648.

Anonymous:2003:EBd

[Ano03l] Anonymous. Editorial Board. Historia Mathematica, 30(4):??, November 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086003000855.

Anonymous:2003:IAAb

[Ano03m] Anonymous. Index of author of articles and publications reviewed and abstracted in volume 30. Historia Mathematica, 30(4):556–563, November 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086003000910.

Anonymous:2003:IAAa

[Ano03n] Anonymous. Index of author of articles and reviews in volume 30. Historia Mathematica, 30(4):555, November 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086003000909.

Anonymous:2003:NCa

[Ano03o] Anonymous. Notes on contributors. Historia Mathematica, 30 (1):120–121, February 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086003000120.

Anonymous:2003:NCb

[Ano03p] Anonymous. Notes on contributors. Historia Mathematica, 30(2): 253–254, May 2003. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086003000533.

Anonymous:2003:NCc

[Ano03q] Anonymous. Notes on contributors. Historia Mathematica, 30(3): 404, August 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086003000673.

Anonymous:2003:NCd

[Ano03r] Anonymous. Notes on contributors. Historia Mathematica, 30(4): 552–553, November 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086003000880.

Anonymous:2003:PAFa

[Ano03s] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 30(1):122, February 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086003000132.

Anonymous:2003:PAFb

[Ano03t] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 30(2):255, May 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086003000405.

Anonymous:2003:PAFc

[Ano03u] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 30(3):405, August 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086003000685.

Anonymous:2003:PAFd

[Ano03v] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 30(4):554, November 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086003000983.

Anonymous:2003:PN

[Ano03w] Anonymous. Publisher’s note. Historia Mathematica, 30(1):iii–iv, February 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086003000065.

Anonymous:2004:AEGa

[Ano04a] Anonymous. Abstracts: Edited by Glen Van Brummelen. Historia Mathematica, 31(1):87–114, February 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086003000995.

Anonymous:2004:AEGb

[Ano04b] Anonymous. Abstracts: Edited by Glen Van Brummelen. Historia Mathematica, 31(2):231–258, May 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086004000163.

Anonymous:2004:AEGc

[Ano04c] Anonymous. Abstracts: Edited by Glen Van Brummelen. Historia Mathematica, 31(3):376–404, August 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508600400045X.

Anonymous:2004:AEGd

[Ano04d] Anonymous. Abstracts: Edited by Glen Van Brummelen. Historia Mathematica, 31(4):502–530, November 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086004000618.

Anonymous:2004:Ca

[Ano04e] Anonymous. Contents. Historia Mathematica, 31(1):??, February 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086004000059.

Anonymous:2004:Cb

[Ano04f] Anonymous. Contents. Historia Mathematica, 31(3):??, August 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086004000485.

Anonymous:2004:EBa

[Ano04g] Anonymous. Editorial Board. Historia Mathematica, 31(1):??, February 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086004000047.

Anonymous:2004:EBb

[Ano04h] Anonymous. Editorial Board. Historia Mathematica, 31(2):??, May 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086004000254.

Anonymous:2004:EBc

[Ano04i] Anonymous. Editorial Board. Historia Mathematica, 31(3):??, August 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086004000473.

Anonymous:2004:EBd

[Ano04j] Anonymous. Editorial Board. Historia Mathematica, 31(4):??, November 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086004000643.

Anonymous:2004:IAAb

[Ano04k] Anonymous. Index of authors of articles and publications reviewed and abstracted in volume 31. Historia Mathematica, 31(4):534–540, November 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086004000709.

Anonymous:2004:IAAa

[Ano04l] Anonymous. Index of authors of articles and reviews in volume 31. Historia Mathematica, 31(4):533, November 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086004000692.

Anonymous:2004:NCa

[Ano04m] Anonymous. Notes on contributors. Historia Mathematica, 31 (1):125–126, February 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086004000102.

Anonymous:2004:NCb

[Ano04n] Anonymous. Notes on contributors. Historia Mathematica, 31(2): 259–260, May 2004. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086004000394.

Anonymous:2004:NCc

[Ano04o] Anonymous. Notes on contributors. Historia Mathematica, 31 (3):405–406, August 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086004000515.

Anonymous:2004:NCd

[Ano04p] Anonymous. Notes on contributors. Historia Mathematica, 31(4):531, November 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086004000771.

Anonymous:2004:PAFa

[Ano04q] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 31(1):127, February 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086004000114.

Anonymous:2004:PAFb

[Ano04r] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 31(2):261, May 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086004000291.

Anonymous:2004:PAFc

[Ano04s] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 31(3):407, August 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086004000527.

Anonymous:2004:PAFd

[Ano04t] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 31(4):532, November 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086004000680.

Anonymous:2004:TCa

[Ano04u] Anonymous. Table of contents. Historia Mathematica, 31(2):??, May 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086004000266.

Anonymous:2004:TCb

[Ano04v] Anonymous. Table of contents. Historia Mathematica, 31(4):??, November 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086004000655.

Anonymous:2005:AESa

[Ano05a] Anonymous. Abstracts: Edited by Sloan Evans Despeaux and Glen Van Brummelen. Historia Mathematica, 32(1):107–125, February 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086004000904.

Anonymous:2005:AESb

[Ano05b] Anonymous. Abstracts: Edited by Sloan Evans Despeaux and Glen Van Brummelen. Historia Mathematica, 32(2):241–267, May 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005000339.

Anonymous:2005:AESc

[Ano05c] Anonymous. Abstracts: Edited by Sloan Evans Despeaux and Glen Van Brummelen. Historia Mathematica, 32(3):368–393, August 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005000418.

Anonymous:2005:AESd

[Ano05d] Anonymous. Abstracts: Edited by Sloan Evans Despeaux and Kim Plofker. Historia Mathematica, 32(4):495–515, November 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508600500073X.

Anonymous:2005:Ca

[Ano05e] Anonymous. Contents. Historia Mathematica, 32(1):??, February 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005000054.

Anonymous:2005:Cb

[Ano05f] Anonymous. Contents. Historia Mathematica, 32(2):??, May 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005000169.

Anonymous:2005:Cc

[Ano05g] Anonymous. Contents. Historia Mathematica, 32(3):??, August 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005000558.

Anonymous:2005:Cd

[Ano05h] Anonymous. Contents. Historia Mathematica, 32(4):??, November 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005000844.

Anonymous:2005:EBa

[Ano05i] Anonymous. Editorial Board. Historia Mathematica, 32(1):??, February 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086005000042.

Anonymous:2005:EBb

[Ano05j] Anonymous. Editorial Board. Historia Mathematica, 32(2):??, May 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005000157.

Anonymous:2005:EBc

[Ano05k] Anonymous. Editorial Board. Historia Mathematica, 32(3):??, August 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086005000546.

Anonymous:2005:EBd

[Ano05l] Anonymous. Editorial Board. Historia Mathematica, 32(4):??, November 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086005000832.

Anonymous:2005:IAAb

[Ano05m] Anonymous. Index of authors of articles and publications reviewed and abstracted in volume 32. Historia Mathematica, 32(4):520–526, November 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086005000959.

Anonymous:2005:IAAa

[Ano05n] Anonymous. Index of authors of articles and reviews in volume 32. Historia Mathematica, 32(4):519, November 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005000947.

Anonymous:2005:NCa

[Ano05o] Anonymous. Notes on contributors. Historia Mathematica, 32(1):126, February 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508600500011X.

Anonymous:2005:NCb

[Ano05p] Anonymous. Notes on contributors. Historia Mathematica, 32(2): 268, May 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086005000182.

Anonymous:2005:NCc

[Ano05q] Anonymous. Notes on contributors. Historia Mathematica, 32 (3):394–395, August 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086005000662.

Anonymous:2005:NCd

[Ano05r] Anonymous. Notes on contributors. Historia Mathematica, 32(4): 516–517, November 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086005000868.

Anonymous:2005:PAFa

[Ano05s] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 32(1):127, February 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086005000078.

Anonymous:2005:PAFb

[Ano05t] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 32(2):269, May 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086005000194.

Anonymous:2005:PAFc

[Ano05u] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 32(3):396, August 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086005000583.

Anonymous:2005:PAFd

[Ano05v] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 32(4):518, November 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508600500087X.

Anonymous:2006:A

[Ano06a] Anonymous. Abstracts. Historia Mathematica, 33(3):377–394, August 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086006000589.

Anonymous:2006:AESa

[Ano06b] Anonymous. Abstracts: Edited by Sloan Evans Despeaux and Kim Plofker. Historia Mathematica, 33(1):132–145, February 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005001096.

Anonymous:2006:AESb

[Ano06c] Anonymous. Abstracts: Edited by Sloan Evans Despeaux and Kim Plofker. Historia Mathematica, 33(2):261–270, May 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086006000243.

Anonymous:2006:AESc

[Ano06d] Anonymous. Abstracts: Edited by Sloan Evans Despeaux and Kim Plofker. Historia Mathematica, 33(4):509–526, November 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086006000899.

Anonymous:2006:EBa

[Ano06e] Anonymous. Editorial Board. Historia Mathematica, 33(1):??, February 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086006000036.

Anonymous:2006:EBb

[Ano06f] Anonymous. Editorial Board. Historia Mathematica, 33(2):??, May 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086006000279.

Anonymous:2006:EBc

[Ano06g] Anonymous. Editorial Board. Historia Mathematica, 33(3):??, August 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086006000619.

Anonymous:2006:EBd

[Ano06h] Anonymous. Editorial Board. Historia Mathematica, 33(4):??, November 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086006001005.

Anonymous:2006:IAAb

[Ano06i] Anonymous. Index of authors of articles and publications reviewed and abstracted in volume 33. Historia Mathematica, 33(4):530–533, November 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086006001078.

Anonymous:2006:IAAa

[Ano06j] Anonymous. Index of authors of articles and reviews in volume 33. Historia Mathematica, 33(4):529, November 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086006001066.

Anonymous:2006:NCa

[Ano06k] Anonymous. Notes on contributors. Historia Mathematica, 33(1):146, February 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086006000061.

Anonymous:2006:NCb

[Ano06l] Anonymous. Notes on contributors. Historia Mathematica, 33(2): 271–272, May 2006. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086006000310.

Anonymous:2006:NCc

[Ano06m] Anonymous. Notes on contributors. Historia Mathematica, 33 (3):395–396, August 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086006000644.

Anonymous:2006:NCd

[Ano06n] Anonymous. Notes on contributors. Historia Mathematica, 33(4):527, November 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086006001042.

Anonymous:2006:PAFa

[Ano06o] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 33(1):147, February 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086006000073.

Anonymous:2006:PAFb

[Ano06p] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 33(2):273, May 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086006000322.

Anonymous:2006:PAFc

[Ano06q] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 33(3):397, August 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086006000760.

Anonymous:2006:PAFd

[Ano06r] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 33(4):528, November 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086006001054.

Anonymous:2007:EBa

[Ano07a] Anonymous. Editorial Board. Historia Mathematica, 34(1):??, February 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086006001340.

Anonymous:2007:EBb

[Ano07b] Anonymous. Editorial Board. Historia Mathematica, 34(2):??, May 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086007000109.

Anonymous:2007:EBc

[Ano07c] Anonymous. Editorial Board. Historia Mathematica, 34(3):??, August 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086007000407.

Anonymous:2007:EBd

[Ano07d] Anonymous. Editorial Board. Historia Mathematica, 34(4):??, November 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508600700064X.

Anonymous:2007:CEH

[Ano07e] Anonymous, editor. La correspondance entre Henri Poincaré et des physiciens, chimistes, et ingenieurs. (French) [The correspondence betweeen Henri Poincaré and physicians, chemists, and engineers]. Publications des archives Henri-Poincaré, publications of the Henri Poincaré archives. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2007. ISBN 3-7643-7136-6 (hardcover). ???? pp. LCCN ????

Anonymous:2007:NCa

[Ano07f] Anonymous. Notes on contributors. Historia Mathematica, 34 (1):134–135, February 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S031508600600139X.

Anonymous:2007:NCb

[Ano07g] Anonymous. Notes on contributors. Historia Mathematica, 34(2): 259–260, May 2007. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086007000134.

Anonymous:2007:NCc

[Ano07h] Anonymous. Notes on contributors. Historia Mathematica, 34 (3):375–376, August 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086007000456.

Anonymous:2007:NCd

[Ano07i] Anonymous. Notes on contributors. Historia Mathematica, 34(4): 474–475, November 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086007000687.

Anonymous:2007:PAF

[Ano07j] Anonymous. Papers to appear in forthcoming issues. Historia Mathematica, 34(1):136, February 2007. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086006001406.

Anonymous:2008:EBa

[Ano08a] Anonymous. Editorial Board. Historia Mathematica, 35(1):??, February 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508600700105X.

Anonymous:2008:EBb

[Ano08b] Anonymous. Editorial Board. Historia Mathematica, 35(2):??, May 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008000116.

Anonymous:2008:EBc

[Ano08c] Anonymous. Editorial Board. Historia Mathematica, 35(3):??, August 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008000372.

Anonymous:2008:EBd

[Ano08d] Anonymous. Editorial Board. Historia Mathematica, 35(4):??, November 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008000736.

Anonymous:2008:KMP

[Ano08e] Anonymous. Kenneth O. May Prize of the International Commission on the History of Mathematics. Historia Mathematica, 35(3):169, August 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008000359.

Anonymous:2008:NCa

[Ano08f] Anonymous. Notes on contributors. Historia Mathematica, 35(1):80– 81, February 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086007001097.

Anonymous:2008:NCb

[Ano08g] Anonymous. Notes on contributors. Historia Mathematica, 35(2): 148–149, May 2008. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086008000153.

Anonymous:2008:NCc

[Ano08h] Anonymous. Notes on contributors. Historia Mathematica, 35 (3):254–255, August 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008000414.

Anonymous:2008:NCd

[Ano08i] Anonymous. Notes on contributors. Historia Mathematica, 35(4): 360–361, November 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008000773.

Anonymous:2009:CCE

[Ano09a] Anonymous. Cover 2 (C2): Editorial board. Historia Mathematica, 36(4):??, November 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086009000779.

Anonymous:2009:EBa

[Ano09b] Anonymous. Editorial Board. Historia Mathematica, 36(1):??, February 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008001134.

Anonymous:2009:EBb

[Ano09c] Anonymous. Editorial Board. Historia Mathematica, 36(2):??, May 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009000123.

Anonymous:2009:EBc

[Ano09d] Anonymous. Editorial Board. Historia Mathematica, 36(3):??, August 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086009000342.

Anonymous:2009:NCa (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008001171.

Anonymous:2009:NCb

[Ano09f] Anonymous. Notes on contributors. Historia Mathematica, 36(2): 211–212, May 2009. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086009000160.

Anonymous:2009:NCc

[Ano09g] Anonymous. Notes on contributors. Historia Mathematica, 36 (3):315–316, August 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086009000378.

Anonymous:2009:NCd

[Ano09h] Anonymous. Notes on contributors (2pp). Historia Mathematica, 36 (4):471–472, November 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086009000809.

Anonymous:2010:A

[Ano10a] Anonymous. Abstracts. Historia Mathematica, 37(1):129–153, February 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086009000755.

Anonymous:2010:IEBa

[Ano10b] Anonymous. IFC: Editorial board. Historia Mathematica, 37(1):??, February 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086009001049.

Anonymous:2010:IEBb

[Ano10c] Anonymous. IFC: Editorial board. Historia Mathematica, 37(2): ??, May 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508601000008X.

Anonymous:2010:IEBc

[Ano10d] Anonymous. IFC: Editorial board. Historia Mathematica, 37(3):??, August 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508601000039X.

Anonymous:2010:IEBd

[Ano10e] Anonymous. IFC: Editorial board. Historia Mathematica, 37(4):??, November 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086010000571.

Anonymous:2010:MP

[Ano10f] Anonymous. The Montucla Prize, 2009. Historia Mathematica, 37(2):157–158, May 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086010000054.

Anonymous:2010:NCa

[Ano10g] Anonymous. Notes on contributors (2pp). Historia Mathematica, 37(1):154–155, February 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086009001074.

Anonymous:2010:NCb

[Ano10h] Anonymous. Notes on contributors (2pp). Historia Mathematica, 37(2):339–340, May 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S031508601000011X.

Anonymous:2010:NCc

[Ano10i] Anonymous. Notes on contributors (2pp). Historia Mathematica, 37(3):563–564, August 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S031508601000042X.

Anonymous:2010:NCd

[Ano10j] Anonymous. Notes on contributors (2pp). Historia Mathematica, 37 (4):752–753, November 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086010000601.

Anonymous:2010:SAK

[Ano10k] Anonymous. The sixth award of the Kenneth O. May Medal and Prize. Historia Mathematica, 37(1):4–7, February 2010.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009001244.

Anonymous:2011:A

[Ano11a] Anonymous. Abstracts. Historia Mathematica, 38(4):565–589, November 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086011000474.

Anonymous:2011:IEBa

[Ano11b] Anonymous. IFC: Editorial board. Historia Mathematica, 38(1):??, February 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086010000789.

Anonymous:2011:IEBb

[Ano11c] Anonymous. IFC: Editorial board. Historia Mathematica, 38(2): ??, May 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086011000073.

Anonymous:2011:IEBc

[Ano11d] Anonymous. IFC: Editorial board. Historia Mathematica, 38(3):??, August 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508601100036X.

Anonymous:2011:IEBd

[Ano11e] Anonymous. IFC: Editorial board. Historia Mathematica, 38(4):??, November 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086011000528.

Anonymous:2011:NCb

[Ano11f] Anonymous. Notes on contributors (1pp). Historia Mathematica, 38 (2):334, May 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086011000103.

Anonymous:2011:NCa

[Ano11g] Anonymous. Notes on contributors (2pp). Historia Mathematica, 38(1):160–161, February 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086010000819.

Anonymous:2011:NCc

[Ano11h] Anonymous. Notes on contributors (2pp). Historia Mathematica, 38(3):453, August 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086011000395.

Anonymous:2011:NCd

[Ano11i] Anonymous. Notes on contributors (2pp). Historia Mathematica, 38(4):590, November 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086011000553.

Anonymous:2012:IEBa

[Ano12a] Anonymous. IFC: Editorial board. Historia Mathematica, 39(1):??, February 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086011000796.

Anonymous:2012:IEBb

[Ano12b] Anonymous. IFC: Editorial board. Historia Mathematica, 39(2): ??, May 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086012000225.

Anonymous:2012:IEBc

[Ano12c] Anonymous. IFC: Editorial board. Historia Mathematica, 39(3):??, August 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086012000407.

Anonymous:2012:IEBd

[Ano12d] Anonymous. IFC: Editorial board. Historia Mathematica, 39(4):??, November 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508601200064X.

Anonymous:2012:NCa (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086011000826.

Anonymous:2012:NCb

[Ano12f] Anonymous. Notes on contributors (2pp). Historia Mathematica, 39(2):263–264, May 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086012000250.

Anonymous:2012:NCc

[Ano12g] Anonymous. Notes on contributors (2pp). Historia Mathematica, 39(3):385–386, August 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086012000432.

Anonymous:2012:NCd

[Ano12h] Anonymous. Notes on contributors (2pp). Historia Mathematica, 39 (4):504–505, November 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086012000675.

Anonymous:2013:EBa

[Ano13a] Anonymous. Editorial Board. Historia Mathematica, 40(1):??, February 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086012000857.

Anonymous:2013:EBb

[Ano13b] Anonymous. Editorial Board. Historia Mathematica, 40(2):??, May 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086013000074.

Anonymous:2013:EBc

[Ano13c] Anonymous. Editorial Board. Historia Mathematica, 40(3):??, August 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086013000268.

Anonymous:2013:NCa

[Ano13d] Anonymous. Notes on contributors. Historia Mathematica, 40(1):125, February 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086012000882.

Anonymous:2013:NCb

[Ano13e] Anonymous. Notes on contributors. Historia Mathematica, 40(2): 232–233, May 2013. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086013000104.

Anonymous:2013:NCc

[Ano13f] Anonymous. Notes on contributors. Historia Mathematica, 40 (3):351–352, August 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086013000293.

Anonymous:2013:NCd

[Ano13g] Anonymous. Notes on contributors. Historia Mathematica, 40 (4):481–482, October 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086013000669.

Anonymous:2014:EBa

[Ano14a] Anonymous. Editorial Board. Historia Mathematica, 41(3):IFC, August 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086014000597.

Anonymous:2014:EBb

[Ano14b] Anonymous. Editorial Board. Historia Mathematica, 41(4):IFC, November 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086014000871.

Anonymous:2014:NCa

[Ano14c] Anonymous. Notes on contributors. Historia Mathematica, 41 (1):137–138, February 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086014000159.

Anonymous:2014:NCb

[Ano14d] Anonymous. Notes on contributors. Historia Mathematica, 41(2): 258–259, May 2014. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086014000378.

Anonymous:2014:NCc

[Ano14e] Anonymous. Notes on contributors. Historia Mathematica, 41 (3):365–366, August 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086014000627.

Anonymous:2014:NCd

[Ano14f] Anonymous. Notes on contributors. Historia Mathematica, 41(4): 518–519, November 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086014000937.

Anonymous:2015:EBa

[Ano15a] Anonymous. Editorial Board. Historia Mathematica, 42(1):IFC, February 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086015000038.

Anonymous:2015:EBb

[Ano15b] Anonymous. Editorial Board. Historia Mathematica, 42(2):ifc, May 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086015000221.

Anonymous:2015:EBc

[Ano15c] Anonymous. Editorial Board. Historia Mathematica, 42(3):ifc, August 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508601500049X.

Anonymous:2015:EBd

[Ano15d] Anonymous. Editorial Board. Historia Mathematica, 42(4):ifc, November 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086015000804.

Anonymous:2015:NCa (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086015000129.

Anonymous:2015:NCb

[Ano15f] Anonymous. Notes on contributors. Historia Mathematica, 42(2): 260–261, May 2015. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086015000324.

Anonymous:2015:NCc

[Ano15g] Anonymous. Notes on contributors. Historia Mathematica, 42 (3):382–383, August 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086015000579.

Anonymous:2015:NCd

[Ano15h] Anonymous. Notes on contributors. Historia Mathematica, 42(4): 516–517, November 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086015000932.

Anonymous:2016:EBa

[Ano16a] Anonymous. Editorial Board. Historia Mathematica, 43(1):ifc, February 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086016000033.

Anonymous:2016:EBb

[Ano16b] Anonymous. Editorial Board. Historia Mathematica, 43(2):ifc, ???? 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086016300039.

Anonymous:2016:EBc

[Ano16c] Anonymous. Editorial Board. Historia Mathematica, 43(3):ifc, August 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086016300258.

Anonymous:2016:EBd

[Ano16d] Anonymous. Editorial Board. Historia Mathematica, 43(4):ifc, November 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086016300532.

Anonymous:2016:NCa

[Ano16e] Anonymous. Notes on contributors. Historia Mathematica, 43 (1):127–128, February 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086016000100.

Anonymous:2016:NCb

[Ano16f] Anonymous. Notes on contributors. Historia Mathematica, 43 (2):239–240, ???? 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086016300064.

Anonymous:2016:NCc

[Ano16g] Anonymous. Notes on contributors. Historia Mathematica, 43(3): 368, August 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086016300283.

Anonymous:2016:NC

[Ano16h] Anonymous. Notes on contributors. Historia Mathematica, 43(4): 477–478, November 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086016300568.

Anonymous:2017:EBa

[Ano17a] Anonymous. Editorial Board. Historia Mathematica, 44(1):ifc, February 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086017300022.

Anonymous:2017:EBb

[Ano17b] Anonymous. Editorial Board. Historia Mathematica, 44(2):ifc, May 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508601730037X.

Anonymous:2017:EBc

[Ano17c] Anonymous. Editorial Board. Historia Mathematica, 44(3):ifc, August 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086017300551.

Anonymous:2017:NCa

[Ano17d] Anonymous. Notes on contributors. Historia Mathematica, 44 (1):102–103, February 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086017300058.

Anonymous:2017:NCb

[Ano17e] Anonymous. Notes on contributors. Historia Mathematica, 44(2): 198–199, May 2017. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S031508601730040X.

Anonymous:2017:NCc

[Ano17f] Anonymous. Notes on contributors. Historia Mathematica, 44 (3):313–314, August 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086017300587.

Anonymous:2017:NCd

[Ano17g] Anonymous. Notes on contributors. Historia Mathematica, 44(4): 457–458, November 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086017300915.

Anonymous:2018:EBa

[Ano18a] Anonymous. Editorial board. Historia Mathematica, 45(1):ifc, February 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086018300107.

Anonymous:2018:EBb

[Ano18b] Anonymous. Editorial board. Historia Mathematica, 45(2): ifc, May 2018. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0315086018300740.

Anonymous:2018:EBc

[Ano18c] Anonymous. Editorial Board. Historia Mathematica, 45(3):217–322, August 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086018301137.

Anonymous:2018:EBd

[Ano18d] Anonymous. Editorial Board. Historia Mathematica, 45(4):ifc, November 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086018301617.

Anonymous:2018:NCa

[Ano18e] Anonymous. Notes on contributors. Historia Mathematica, 45(1):109, February 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086018300132.

Anonymous:2018:NCb

[Ano18f] Anonymous. Notes on contributors. Historia Mathematica, 45(2): 214–215, May 2018. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0315086018300776.

Anonymous:2018:NCc

[Ano18g] Anonymous. Notes on contributors. Historia Mathematica, 45 (3):306–320, August 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086018301162.

Anonymous:2018:NCd

[Ano18h] Anonymous. Notes on contributors. Historia Mathematica, 45(4): 450–451, November 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086018301642.

Anonymous:2019:EBa

[Ano19a] Anonymous. Editorial Board. Historia Mathematica, 46(??):ifc, February 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086019300035.

Anonymous:2019:EBb

[Ano19b] Anonymous. Editorial Board. Historia Mathematica, 47(??):ifc, ???? 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086019300242.

Anonymous:2019:EBc

[Ano19c] Anonymous. Editorial Board. Historia Mathematica, 48(??):ifc, August 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508601930045X.

Anonymous:2019:EBd

[Ano19d] Anonymous. Editorial Board. Historia Mathematica, 49(??):ifc, November 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508601930062X.

Anonymous:2019:NCa

[Ano19e] Anonymous. Notes on contributors. Historia Mathematica, 46 (??):131–132, February 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086019300163.

Anonymous:2019:NCb

[Ano19f] Anonymous. Notes on contributors. Historia Mathematica, 47 (??):144–145, ???? 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086019300394.

Anonymous:2019:NCc

[Ano19g] Anonymous. Notes on contributors. Historia Mathematica, 48 (??):122–123, August 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086019300552.

Anonymous:2019:NCd

[Ano19h] Anonymous. Notes on contributors. Historia Mathematica, 49(??): 116–117, November 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086019300734.

Anonymous:2020:EBa

[Ano20a] Anonymous. Editorial Board. Historia Mathematica, 50(??):ifc, February 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086020300082.

Anonymous:2020:EBb

[Ano20b] Anonymous. Editorial Board. Historia Mathematica, 51(??):ifc, May 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086020300276.

Anonymous:2020:EBc

[Ano20c] Anonymous. Editorial Board. Historia Mathematica, 52(??):ifc, August 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086020300525.

Anonymous:2020:EBd

[Ano20d] Anonymous. Editorial Board. Historia Mathematica, 53(??):ifc, November 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086020300768.

Anonymous:2020:NCa

[Ano20e] Anonymous. Notes on contributors. Historia Mathematica, 50 (??):113, February 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086020300203.

Anonymous:2020:NCb

[Ano20f] Anonymous. Notes on contributors. Historia Mathematica, 51(??): 116, May 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086020300409.

Anonymous:2020:NCc

[Ano20g] Anonymous. Notes on contributors. Historia Mathematica, 52 (??):125–126, August 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086020300689.

Anonymous:2020:NCd

[Ano20h] Anonymous. Notes on contributors. Historia Mathematica, 53(??): 147–148, November 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086020300872.

Anonymous:2020:NNE

[Ano20i] Anonymous. Notes on new Editorial Board members. Historia Mathematica, 50(??):111–112, February 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086020300197.

Anonymous:2021:AMP

[Ano21a] Anonymous. Awarding the Montucla Prize for 2021. Historia Mathematica, 56(??):4–5, August 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086021000458.

Anonymous:2021:EBa

[Ano21b] Anonymous. Editorial Board. Historia Mathematica, 54(??):ifc, February 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086021000082.

Anonymous:2021:EBb

[Ano21c] Anonymous. Editorial Board. Historia Mathematica, 55(??):ifc, May 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508602100032X.

Anonymous:2021:EBc

[Ano21d] Anonymous. Editorial Board. Historia Mathematica, 56(??):ifc, August 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086021000471.

Anonymous:2021:EBd

[Ano21e] Anonymous. Editorial Board. Historia Mathematica, 57(??):ifc, November 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086021000707.

Anonymous:2021:NCa

[Ano21f] Anonymous. Notes on contributors. Historia Mathematica, 54 (??):143–144, February 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086021000173.

Anonymous:2021:NCb

[Ano21g] Anonymous. Notes on contributors. Historia Mathematica, 55 (??):115–116, May 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086021000306.

Anonymous:2021:NCc

[Ano21h] Anonymous. Notes on contributors. Historia Mathematica, 56(??): 107, August 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086021000598.

Anonymous:2021:NCd

[Ano21i] Anonymous. Notes on contributors. Historia Mathematica, 57(??): 117–118, November 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086021000811.

Anonymous:2021:PA

[Ano21j] Anonymous. Pages 1–108 (August 2021). Historia Mathematica, 56 (??):1–108, August 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Anonymous:2021:PM

[Ano21k] Anonymous. Pages 1–116 (May 2021). Historia Mathematica, 55 (??):1–116, May 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Anonymous:2021:PN

[Ano21l] Anonymous. Pages 1–118 (November 2021). Historia Mathematica, 57(??):1–118, November 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Anonymous:2021:PF

[Ano21m] Anonymous. Pages 1–144 (February 2021). Historia Mathematica, 54(??):1–144, February 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Anonymous:2022:EBa

[Ano22a] Anonymous. Editorial Board. Historia Mathematica, 58(??):ifc, February 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508602200012X.

Anonymous:2022:EBb

[Ano22b] Anonymous. Editorial Board. Historia Mathematica, 59(??):ifc, May 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086022000349.

Anonymous:2022:EBc

[Ano22c] Anonymous. Editorial Board. Historia Mathematica, 60(??):ifc, August 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086022000507.

Anonymous:2022:EBd

[Ano22d] Anonymous. Editorial Board. Historia Mathematica, 61(??):ifc, November 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086022000799.

Anonymous:2022:NCa

[Ano22e] Anonymous. Notes on contributors. Historia Mathematica, 58 (??):133–134, February 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086022000155.

Anonymous:2022:NCb

[Ano22f] Anonymous. Notes on contributors. Historia Mathematica, 59 (??):215–217, May 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086022000428.

Anonymous:2022:NCc

[Ano22g] Anonymous. Notes on contributors. Historia Mathematica, 60(??): 97, August 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086022000647.

Anonymous:2022:NCd

[Ano22h] Anonymous. Notes on contributors. Historia Mathematica, 61 (??):97–98, November 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086022000866.

Anonymous:2022:PF

[Ano22i] Anonymous. Pages 1–134 (February 2022). Historia Mathematica, 58(??):1–134, February 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Anonymous:2022:PA

[Ano22j] Anonymous. Pages 1–98 (August 2022). Historia Mathematica, 60 (??):1–98, August 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Anonymous:2022:PN

[Ano22k] Anonymous. Pages 1–98 (November 2022). Historia Mathematica, 61(??):1–98, November 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Anonymous:2022:TC

[Ano22l] Anonymous. Table of contents. Historia Mathematica, 59(??):1, May 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086022000258.

Anonymous:2023:EBa

[Ano23a] Anonymous. Editorial Board. Historia Mathematica, 62(??):ifc, February 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086023000058.

Anonymous:2023:EBb

[Ano23b] Anonymous. Editorial Board. Historia Mathematica, 63(??):ifc, May 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086023000204.

Anonymous:2023:EBc

[Ano23c] Anonymous. Editorial Board. Historia Mathematica, 64(??):ifc, August 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508602300054X.

Anonymous:2023:EBd

[Ano23d] Anonymous. Editorial Board. Historia Mathematica, 65(??):ii, November 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086023000800.

Anonymous:2023:NCa

[Ano23e] Anonymous. Notes on contributors. Historia Mathematica, 62(??): 92, February 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086023000113.

Anonymous:2023:NCb

[Ano23f] Anonymous. Notes on contributors. Historia Mathematica, 63(??): 92, May 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086023000332.

Anonymous:2023:NCc

[Ano23g] Anonymous. Notes on contributors. Historia Mathematica, 64 (??):76–77, August 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086023000642.

Anonymous:2023:PN

[Ano23h] Anonymous. Pages 1–54 (November 2023). Historia Mathematica, 65(??):1–54, November 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Anonymous:2023:PA

[Ano23i] Anonymous. Pages 1–78 (August 2023). Historia Mathematica, 64 (??):1–78, August 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Anonymous:2023:PF

[Ano23j] Anonymous. Pages 1–92 (February 2023). Historia Mathematica, 62(??):1–92, February 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Anonymous:2023:PM

[Ano23k] Anonymous. Pages 1–92 (May 2023). Historia Mathematica, 63(??): 1–92, May 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Anonymous:2024:Aa

[Ano24a] Anonymous. Abstracts. Historia Mathematica, 66(??):50–59, March 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086023000678.

Anonymous:2024:EBa

[Ano24b] Anonymous. Editorial Board. Historia Mathematica, 66(??):ii, March 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086024000156.

Anonymous:2024:EBb

[Ano24c] Anonymous. Editorial Board. Historia Mathematica, 67(??):ii, June 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086024000661.

Anonymous:2024:EBc

[Ano24d] Anonymous. Editorial Board. Historia Mathematica, 68(??):ii, September 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086024000818.

Anonymous:2024:EBd

[Ano24e] Anonymous. Editorial Board. Historia Mathematica, 69(??):ii, December 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086024000958.

Anonymous:2024:PJ

[Ano24f] Anonymous. Pages 1–54 (June 2024). Historia Mathematica, 67(??): 1–54, June 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Anonymous:2024:PMa

[Ano24g] Anonymous. Pages 1–60 (March 2024). Historia Mathematica, 66 (??):1–60, March 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Anonymous:2024:PD

[Ano24h] Anonymous. Pages 1–72 (December 2024). Historia Mathematica, 69(??):1–72, December 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Anonymous:2024:PS

[Ano24i] Anonymous. Pages 1–76 (September 2024). Historia Mathematica, 68(??):1–76, September 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Antropov:1995:EPF

[Ant95] A. A. Antropov. On Euler’s partition of forms into genera. Historia Mathematica, 22(2):188–193, May 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508608571018X.

Agargun:2001:HSF

[AÖ01] A. Göksel Agargün and E. Mehmet Özkan. A historical survey of the fundamental theorem of arithmetic. Historia Mathematica, 28(3):207–214, August 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086001923183.

Altmann:2005:MSU

[AO05] Simon L. Altmann and Eduardo L. Ortiz, editors. Mathematics and social utopias in France: Olinde Rodrigues and his times, volume 28 of History of mathematics, 0899-2428. American Mathematical Society, Providence, RI, USA, 2005. ISBN 0-8218-3860-1. xii + 168 pp.

LCCN QA29.R58 M38 2005. URL http://www.loc.gov/catdir/
toc/fy0612/2005048180.html.

Ahmad:1976:SRA

[AR76] Salah Ahmad and Roshdi Rashed. A. S. Saidan’s review of Al-Bahir en algèbre d’as-samaw’al. Historia Mathematica, 3(1):78–80, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086076900136.

Artin:2007:EEA

[AR07] Emil Artin and Michael I. (Michael Ira) Rosen, editors. Exposition by Emil Artin: a selection, volume 30 of History of mathematics, 0899-2428. American Mathematical Society, Providence, RI, USA, 2007. ISBN 0-8218-4172-6. x + 346 pp. LCCN QA247 .A782 2007.

Arthur:2024:UIL

[AR24] Richard Arthur and David Rabouin. “On the unviability of interpreting Leibniz’s infinitesimals through non-standard analysis”. Historia Mathematica, 66(??):26–42, March 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508602300085X.

Arboleda:1986:LAS

[Arb86] Luis Carlos Arboleda. Latin American Seminar on the teaching of History of Science: Cali, Colombia, November 4–10, 1984. Historia Mathematica, 13(2):176–177, May 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086086900376.

Archibald:1996:BRH

[Arc96] Tom Archibald. Book review: A History of Mathematics: an Introduction. By Victor J. Katz. New York (Harper Collins). Historia Mathematica, 23(1):89–92, February 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086096900099.

Archibald:1999:BRV

[Arc99] Thomas Archibald. Book review: Va Pensiero: Immagini della matematica nell’Italia dell’ottocento. By Umberto Bottazzini. Historia Mathematica, 26(1):79–81, February 1, 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086099922218.

Archibald:2009:BRI

[Arc09] Thomas Archibald. Book review: Italian Mathematics Between the Two World Wars, Angelo Guerraggio, Pietro Nastasi. Birkhäuser, Basel (2006), ISBN 978-3-7643-6555-4. Historia Mathematica, 36 (3):289–290, August 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008001067.

Archibald:2017:BRP

[Arc17] T. Archibald. Book review: La patria ci vuole eroi: matematici e vita politica nell’Italia del Resorgimento, Umberto Bottazzini, Pietro Nastasi. Zanichelli, Bologna (2013). Historia Mathematica, 44(1):76– 79, February 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086016300386.

Arney:1988:RHM

[Arn88a] David C. Arney. Resources in the history of mathematics at the United States Military Academy. Historia Mathematica, 15(4): 368–369, November 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086088900286.

Arnold:1988:KZC

[Arn88b] Bernd J. Arnold. Das Kontinuitätsprogramm zwischen Cantor und der Nonstandard-Analysis. (German) [The continuity program between Cantor and nonstandard analysis]. Historia Mathematica, 15(3):250–263, August 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086088900651.

Aronson:1983:MMM

[Aro83] Charles D. Aronson. Marx’s mathematical manuscripts. Historia Mathematica, 10(1):103–104, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086083900484.

Abardia:2012:WDG

[ARR12] Judit Abardia, Agustí Reventós, and Carlos J. Rodríguez. What did Gauss read in the appendix? Historia Mathematica, 39(3):292–323, August 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086012000122.

Artmann:1991:MGM

[Art91] Benno Artmann. A meeting on Greek mathematics: Mathematisches Forschungsinstitut Oberwolfach, August 19–25, 1990. Historia Mathematica, 18(2):180, May 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609190502O.

Artin:2013:EAI

[Art13] Tom Artin. Emil Artin’s Iceland Journal 1925. Free Scholar Press, ????, 2013. ISBN 1-4825-5168-3. ???? pp. LCCN ????

Artmann:1993:PFT

[AS93] Benno Artmann and Lothar Schäfer. On Plato’s “fairest triangles” (Timaeus 54a). Historia Mathematica, 20(3):255–264, August 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086083710220.

Ascher:1984:MIN

[Asc84] Marcia Ascher. Mathematical ideas in non-Western cultures. Historia Mathematica, 11(1):76–80, February 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086084900065.

Ascher:1988:GCS

[Asc88] Marcia Ascher. Graphs in cultures: a study in ethnomathematics. Historia Mathematica, 15(3):201–227, August 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900626.

Ascher:1995:MMM

[Asc95] Marcia Ascher. Models and maps from the Marshall Islands: a case in ethnomathematics. Historia Mathematica, 22(4):347–370, November 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086085710300.

Ascher:1997:MSC

[Asc97] Marcia Ascher. Malagasy Sikidy: a case in ethnomathematics. Historia Mathematica, 24(4):376–395, November 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086097921595.

Aspray:1983:BRE

[Asp83] William Aspray. Book review: From Eniac to Univac: an appraisal of the Eckert–Mauchly computers: By Nancy Stern. Bedford, MA (Digital Press). 1981. x + 286 pp. Appendix: “First Draft of a Report on the EDVAC” by John von Neumann. Bibliography. Historia Mathematica, 10(2):228–233, May 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508608390071X.

Asper:2019:PPM

[Asp19] Markus Asper. Personae at play. ‘Men of Mathematics’ in commentary. Historia Mathematica, 47(??):4–15, ???? 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508601830137X.

Archibald:2016:E

[ASS16] Tom Archibald and Reinhard Siegmund-Schultze. Editorial. Historia Mathematica, 43(1):1–2, February 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086015001019.

Atzema:2004:BRC

[Atz04] Eisso J. Atzema. Book review: Caspar Wessel. On the Analytical Representation of Direction. An Attempt Chiefly to Solving Plane and Spherical Polygons (1797): Bodil Branner and Jesper Lützen, Eds. Copenhagen (Matematik-fysiske Meddelser 46:1), 1999. Historia Mathematica, 31(1):116–119, February 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086003000600.

Atzema:2006:BRL

[Atz06] Eisso J. Atzema. Book review: Lenses and Waves. Christiaan Huygens and the Mathematical Science of Optics in the Seventeenth Century, vol. 9, Fokko Jan Dijksterhuis, in: Archimedes. New Studies in the History and Philosophy of Science and Technology. Kluwer Academic, Dordrecht (2004), ISBN 1-4020-2697-8. Historia Mathematica, 33(3):367–369, August 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086005000637.

Aubin:2002:SSM

[Aub02] David Aubin. Stephen Smale: the mathematician who broke the dimensional barrier. Historia Mathematica, 29(4):490–491, November 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086002923700.

Aubin:2014:BRP

[Aub14] David Aubin. Book review: “Principles of Mechanics that are Susceptible of Application to Society”: an unpublished notebook of Adolphe Quetelet at the root of his social physics. Historia Mathematica, 41(2):204–223, May 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086014000020.

Audin:2011:FJM

[Aud11] Michèle Audin. Fatou, Julia, Montel: the great prize of mathematical sciences of 1918, and beyond, volume 2014 of Lecture notes in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 3-642-17853-7 (paperback), 3-642-17854-5 (e-book). ISSN 0075-8434. viii + 332 pp. LCCN QA26 .A8313 2011.

Audin:2012:JFT

[Aud12] Michèle Audin. Jacques Feldbau, Topologe : das Schicksal eines jüdischen Mathematikers (1914–1945). Mathematik im Kontext. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-25803-4, 3-642-25804-2. xii + 96 pp. LCCN ????

Aujac:1976:SSG

[Auj76] Germaine Aujac. Sphérique et sphéropée en Grèce ancienne. Historia Mathematica, 3(4):441–447, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086076900744.

Aujac:1986:RIE

[Auj86] Germaine Aujac. Le rapport di isou (Euclide V, définition 17): Définition, utilisation, transmission. Historia Mathematica, 13(4): 370–386, November 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086086900583.

Aupetit:1976:BRF

[Aup76] Bernard Aupetit. Book review: Functional analysis in historical perspective: By A. F. Monna. Utrecht (Oosthoek Publishing Co). 1973. 17S p. Historia Mathematica, 3(1):93–95, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900215.

Ausejo:2013:MP

[Aus13] Elena Ausejo. Montucla Prize 2013. Historia Mathematica, 40 (4):357–358, October 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086013000487.

Ausejo:2016:BRI

[Aus16] Elena Ausejo. Book review: Infinitos, paradojas y principios. Escritos históricos en torno a los fundamentos de las matemáticas, Alejandro R. Garciadiego. Plaza y Valdés Editores, Madrid, México DF (2014), ISBN: 978-84-15271-88-8. Historia Mathematica, 43 (4):449–451, November 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086016000161.

Auvinet:2013:CAL

[Auv13] Jérôme Auvinet. Charles-Ange Laisant: itinéraires et engagements d’un mathématicien de la Troisième République. Histoire des sciences. Hermann, Paris, France, 2013. ISBN 2-7056-8697-5 (paperback). 241 pp. LCCN QA29.L354 A98 2013.

Axworthy:2018:DBP

[Axw18] Angela Axworthy. The debate between Peletier and Clavius on superposition. Historia Mathematica, 45(1):1–38, February 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086017300745.

Ayoub:2004:MMA

[Ayo04] Raymond George Ayoub, editor. Musings of the masters: an anthology of mathematical reflections. Spectrum series. Mathematical Association of America, Washington, DC, USA, 2004. ISBN 0-88385-549-6. xvi + 277 pp. LCCN QA7 .M87 2004. URL

http://www.loc.gov/catdir/enhancements/fy0632/2004104239-
d.html;   http://www.loc.gov/catdir/enhancements/fy0733/
2004104239-b.html;   http://www.loc.gov/catdir/toc/fy0605/
2004104239.html.

Ayoub:2006:BRS

[Ayo06] Raymond Ayoub. Book review: Sherlock Holmes in Babylon and Other Tales of Mathematical History. Edited by Marlow Anderson, Victor Katz, and Robin Wilson. Washington, DC (The Mathematical Association of America). 2004. ISBN 0-88385-546-1. 398 pp. $51.95. Historia Mathematica, 33(3):359–361, August 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005000790.

Babini:1977:LDC

[Bab77] J. Babini. La labor del comité international pour la métrologie historique. Historia Mathematica, 4(3):344–345, August 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900635.

Bagheri:1997:MPF

[Bag97a] Mohammad Bagheri. Mathematical problems of the famous Iranian poet Nās.er-e Khosrow. Historia Mathematica, 24(2):193–196, May 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086096921448.

Bagheri:1997:NFL

[Bag97b] Mohammad Bagheri. A newly found letter of Al-Kāshī on scientific life in Samarkand. Historia Mathematica, 24(3):241–256, August 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508609692145X.

Bagheri:2002:AG

[Bag02] Mohammad Bagheri. Abolghassem Ghorbani (1912–2001). Historia Mathematica, 29(3):244–246, August 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508600292361X.

Baker:1975:SHR

[Bak75] Alan Baker. Some historical remarks on number theory. Historia Mathematica, 2(4):549–553, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075901196.

Baltus:2004:DPF

[Bal04] Christopher Baltus. D’Alembert’s proof of the fundamental theorem of algebra. Historia Mathematica, 31(4):414–428, November 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086003001083.

Baltus:2023:PDH

[Bal23] Christopher Baltus. Poncelet’s discovery of homology. Historia Mathematica, 63(??):1–20, May 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086022000878.

Barbeau:1974:BRC

[Bar74] E. J. Barbeau. Book review: Classical and modern integration theories: By Ivan N. Pesin. New York (Academic Press, Probability and Mathematical Statistics Series #8), 1970. 215 pp. $13.25. Historia Mathematica, 1(1):97–98, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508607490189X.

Barbeau:1977:BRL

[Bar77] E. J. Barbeau. Book review: Leonhardi euleri opera omnia. Commentationes mechanicae et astronomicae ad scientiam navalem pertinentes. Volumen prius: Edited by Walter Habicht. Series 2, Volume 20 of the Collected Works of L. Euler. Basel (Orell Füssli Turici). 1974. lx + 275 pp. Historia Mathematica, 4(2):221, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077901227.

Barbeau:1978:EA

[Bar78a] E. J. Barbeau. Errata and additions. Historia Mathematica, 5(3): 332, August 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086078901180.

Baron:1978:BRG

[Bar78b] Margaret E. Baron. Book review: Gottfried Wilhelm Leibniz. Mathematischer naturwissenschaftlicher und technischer Briefwechsel. Erster Band: 1672–1676: Edited by J. E. Hofmann and the Leibniz-Archiv der Niedersächsichen Landesbibliothek, Hannover. Berlin, G.D.R. (Akademie-Verlag). 1976. lxxv + 871 pp. (Sämtliche Schriften und Briefe, herausgegeben von der Akademie der Wissenschaften der DDR, Dritte Reihe, Erster Band). Historia Mathematica, 5(2):234– 237, May 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086078900617.

Baron:1982:BRMa

[Bar82a] Margaret E. Baron. Book review: Die mathematischen studien von G. W. Leibniz zur kombinatorik: Studia leibnitiana supplementa: Vol. XI. By Eberhard Knobloch. Wiesbaden (Franz Steiner Verlag). 1973. 277 + xvi pp., 2 plates. Historia Mathematica, 9(1):109–113, February 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086082901501.

Baron:1982:BRMb

[Bar82b] Margaret E. Baron. Book review: Die mathematischen studien von G. W. Leibniz zur kombinatorik: Textband: Studia leibnitiana supplementa: Vol. XVI. By Eberhard Knobloch. Wiesbaden (Franz Steiner Verlag). 1976. 339 + xii pp., 2 plates. Historia Mathematica, 9 (1):109–113, February 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086082901513.

Barany:2011:GKG

[Bar11] Michael J. Barany. God, King, and geometry: revisiting the introduction to Cauchy’s Cours d’analyse. Historia Mathematica, 38(3):368– 388, August 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086010001114.

Bashmakova:1981:AAC

[Bas81] Isabella G. Bashmakova. Arithmetic of algebraic curves from Diophantus to Poincaré. Historia Mathematica, 8(4):393–416, November 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900501.

Batterson:2006:PGF

[Bat06] Steve Batterson. Pursuit of genius: Flexner, Einstein, and the early faculty at the Institute for Advanced Study. A. K. Peters, Ltd., Wellesley, MA, USA, 2006. ISBN 1-56881-259-0. xi + 301 pp. LCCN QA13.5.N383 I583 2006.

Baumwell:1974:AYA

[Bau74] Jeffrey H. Baumwell. Anniversaries and 100 years ago. Historia Mathematica, 1(3):290, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074900688.

Bayley:2011:BRF

[Bay11] Melanie Bayley. Book review: Flatland by Edwin A. Abbott: an Edition with Notes and Commentary. By William F. Lindgren and Thomas F. Banchoff. New York (Cambridge University Press). 2010. ISBN 978-0-521-75994-6. ix + 294. $14.00. Historia Mathematica, 38(1):131–135, February 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086010000522.

Brack-Bernsen:1997:MOS

[BB97] Lis Brack-Bernsen. In memoriam: Olaf Schmidt (1913–1996). Historia Mathematica, 24(2):131–134, May 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086097921340.

Brack-Bernsen:2004:BRA

[BB04] Lis Brack-Bernsen. Book review: Ägyptische Algorithmen. Eine Untersuchung zu den mittelägyptischen mathematischen Aufgabentexten: By Annette Imhausen. Wiesbaden (Harrassowitz Verlag). 2003. ISBN 3-447-04644-9. 387 pp. 58 EUR . Historia Mathematica, 31(2):228– 230, May 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508600300079X.

Bender:2021:WCM

[BB21] Andrea Bender and Sieghard Beller. Ways of counting in Micronesia. Historia Mathematica, 56(??):40–72, August 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086021000239.

Brigaglia:2004:RRB

[BC04] Aldo Brigaglia and Ciro Ciliberto. Remarks on the relations between the Italian and American schools of algebraic geometry in the first decades of the 20th century. Historia Mathematica, 31(3):310–319, August 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086003000831.

Brigaglia:2011:LCA

[BD11] Aldo Brigaglia and Simonetta Di Sieno. The Luigi Cremona Archive of the Mazzini Institute of Genoa. Historia Mathematica, 38(1):96– 110, February 2011. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086010000248.

Bottazzini:1988:FIC

[BDD^{+}88] U. Bottazzini, A. Dahan, J. Dhombres, E. Giusti, and S. Roero. Franco-Italian colloquy ars analytica: September 28–October 2, 1987. Historia Mathematica, 15(2):170–171, May 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900109.

Bashmakova:1986:APY

[BDG^{+}86] I. G. Bashmakova, S. S. Demidov, A. T. Grigorian, A. N. Kolmogorov, Yu. V. Prokhorov, and B. A. Rosenfeld. Adolph Pavlovich Yushkevich on the occasion of his eightieth birthday. Historia Mathematica, 13(3):215–220, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508608690087X.

Bradley:2007:EA

[BDS07] Robert E. Bradley, Lawrence A. D’Antonio, and Charles Edward Sandifer, editors. Euler at 300: an appreciation, volume 5 of The MAA tercentenary Euler celebration; Spectrum series. Mathematical Association of America, Washington, DC, USA, 2007. ISBN 0-88385-565-8. xvi + 298 pp. LCCN QA29.E8 E95 2007. URL

http://www.loc.gov/catdir/enhancements/fy0916/2007930320-
b.html;   http://www.loc.gov/catdir/enhancements/fy0916/
2007930320-d.html;   http://www.loc.gov/catdir/toc/fy0805/
2007930320.html.

Bebiano:1996:MHL

[Beb96] Natália Bebiano. Mathematical horizons: Light and darkness in Portugal in the 18th century. Historia Mathematica, 23(3):239–245, August 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086096900257.

Becher:1980:WBP

[Bec80] Harvey W. Becher. Woodhouse, Babbage, Peacock, and modern algebra. Historia Mathematica, 7(4):389–400, November 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900038.

Beckers:1999:LND

[Bec99] Danny J. Beckers. Lagrange in The Netherlands: Dutch attempts to obtain rigor in calculus, 1797–1840. Historia Mathematica, 26(3):224– 238, August 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086099922462.

Beckers:2001:ULO

[Bec01] Danny J. Beckers. “Untiring labor overcomes all!” the history of the Dutch mathematical society in comparison to its various counterparts in Europe. Historia Mathematica, 28(1):31–47, February 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086000922948.

Becker:2019:EEU

[Bec19] Holger Becker. An even earlier (1717) usage of the expression “golden section”. Historia Mathematica, 49(??):82–83, November 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086019300588.

Beery:2006:BRD

[Bee06] Janet L. Beery. Book review: A Discourse Concerning Algebra: English Algebra to 1685. By Jacqueline A. Stedall. Oxford (Oxford University Press). 2002. ISBN 0-19-852495-1. 294 pp. $116. Historia Mathematica, 33(2):250–253, May 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086005000704.

Beeley:2009:BRD

[Bee09] Philip Beeley. Book review: Dal movimento alla forza. Leibniz: l’infinitesimo tra logica e metafisica, Raffaele Cirino. Rubbettino, Soveria Mannelli (2006), ISBN 88 498 1593-X. Historia Mathematica, 36(1):80–82, February 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S031508600800061X.

Beeley:2014:CJS

[Bee14] Philip Beeley. Christoph J. Scriba (6 October 1929–26 July 2013). Historia Mathematica, 41(1):6–12, February 2014. CODEN HI-MADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/S0315086014000068.

Boyer:1975:RAP

[BEJ^{+}75] Carl B. Boyer, Churc’hill Eisenhart, Phillip S. Jones, Kenneth O. May, David Rosenblatt, and Charles Weiner. A renewed appeal for the preservation of archival materials by the CBMS Advisory Committee on History. Historia Mathematica, 2(1):72–73, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900439.

Bekemeyer:1988:RGH

[Bek88] Bernd Bekemeyer. Research group on the history of mathematical textbooks (GRIMM). Historia Mathematica, 15(1):87–88, February 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900572.

Bellhouse:2002:SRD

[Bel02] David Bellhouse. On some recently discovered manuscripts of Thomas Bayes. Historia Mathematica, 29(4):383–394, November 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508600292344X.

Bellhouse:2005:DCL

[Bel05] David Bellhouse. Decoding Cardano’s Liber de Ludo Aleae. Historia Mathematica, 32(2):180–202, May 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086004000400.

Bellhouse:2007:BRA

[Bel07a] David Bellhouse. Book review: The Art of Conjecturing, together with Letter to a Friend on Sets in Court Tennis. The Johns Hopkins University Press, Baltimore (2005), ISBN 0-8018-8235-4. Historia Mathematica, 34(4):435–436, November 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086006001327.

Bellhouse:2007:LSP

[Bel07b] David R. Bellhouse. Lord Stanhope’s papers on the Doctrine of Chances. Historia Mathematica, 34(2):173–186, May 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086006000929.

Bellhouse:2008:BRD

[Bel08] David Bellhouse. Book review: Disciplining Statistics: Demography and Vital Statistics in France and England, 1830–1885, Libby Schweber. Duke University Press, Durham and London (2006), ISBN 0-8223-3814-9. Historia Mathematica, 35(3):249–252, August 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086007000602.

Bellhouse:2012:BRH

[Bel12] David Bellhouse. Book review: A History of the Central Limit Theorem: From Classical to Modern Probability Theory. By Hans Fischer. Dordrecht, Heidelberg, London, New York (Springer). 2010. ISBN 978-0-387-87856-0. 418 pp. Historia Mathematica, 39(4):465–467, November 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086012000134.

Bellhouse:2015:AEM

[Bel15] David R. Bellhouse. An analysis of errors in mathematical tables. Historia Mathematica, 42(3):280–295, August 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086014001104.

Benzi:1988:PNF

[Ben88] Margherita Benzi. Un “Probabilista neoclassico”: Francesco Paolo Cantelli. Historia Mathematica, 15(1):53–72, February 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900493.

Benzi:1995:DCI

[Ben95] Margherita Benzi. Dubbiezze e controversie: Il dibattito su logica e probabilitá in Italia nei primi anni del Novecento. (Italian) [Doubts and controversies: the debate on logic and probability in Italy in the early years of the twentieth century]. Historia Mathematica, 22(1):43– 63, February 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508608571004X.

Berggren:1974:BRM

[Ber74] J. L. Berggren. Book review: Mathematics in civilization: By H. L. Resnikoff and R. O. Wells, Jr. New York (Holt, Rinehart and Winston), 1973. 384 pp. $12.50. Historia Mathematica, 1(1):102–106, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508607490192X.

Berggren:1975:ICM

[Ber75] J. L. Berggren. International Congress of mathematicians 1974. Historia Mathematica, 2(1):83, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075900518.

Berggren:1977:LBA

[Ber77] J. L. Berggren. A lacuna in Book I of Archimedes’ sphere and cylinder. Historia Mathematica, 4(1):1–6, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900283.

Berggren:1984:HGM

[Ber84a] J. L. Berggren. History of Greek mathematics: a survey of recent research. Historia Mathematica, 11(4):394–410, November 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086084900247.

Berggren:1984:AJS

[Ber84b] John L. Berggren. Abraham J. Sachs (1914–1983): In memoriam. Historia Mathematica, 11(2):124–125, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086084800026.

Berry:1985:CSH

[Ber85] John W. Berry. Canadian society for the history and philosophy of mathematics: Annual meeting-1984. Historia Mathematica, 12(2):161–164, May 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086085900096.

Berggren:1987:BRD

[Ber87] J. L. Berggren. Book review: Diophante: Les arithmétiques: Volume III: Book IV, ccvi + 162 numbered pp. (261 total pp.), illus., bibl.; Volume IV: Books V–VII, cxxxiv + 197 numbered pp. (253 total pp.), app., index. Edited by Roshdi Rashed. Collection des Universités de France. Paris (Société d’Édition “Les Belles Lettres”). 1984. Fr. 500 (cloth). Historia Mathematica, 14(4):386–389, November 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086087900711.

Bergmann:1990:NEA

[Ber90] Werner Bergmann. Noch einmal zu den Ausnahmeregeln der Gauss’schen Osterformel. (German) [Once again, the exception rules of the Gaussian Easter formula]. Historia Mathematica, 17(3):256– 258, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609090003V.

Berggren:1997:MHS

[Ber97] J. Lennart Berggren. Mathematics and her sisters in Medieval Islam: a selective review of work done from 1985 to 1995. Historia Mathematica, 24(4):407–440, November 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508609792173X.

Berggren:2004:SFR

[Ber04] J. L. Berggren. On the sector-figure and related texts, vol. 108. Institute for the History of Arabic–Islamic Science at the Johann Wolfgang Goethe University, Frankfurt am Main (2001). Historia Mathematica, 31(4):498–499, November 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086004000588.

Berggren:2007:BRA

[Ber07] J. L. Berggren. Book review: Al-Farghani on the Astrolabe, Band 52, Richard Lorch (Ed.), in: Boethius. Franz Steiner Verlag, Stuttgart (2005), ISBN 3-515-08713-3. Historia Mathematica, 34(4):433–434, November 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086006001303.

Bottazzini:1989:MVL

[BF89] Umberto Bottazzini and Stefano Francesconi. Manuscript volumes and lecture notes of Salvatore Pincherle. Historia Mathematica, 16 (4):379–380, November 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086089900852.

Bottazzini:2000:TMN

[BF00] Umberto Bottazzini and Craig Fraser. At the turn of the millennium: New challenges for the history of mathematics and for Historia Mathematica. Historia Mathematica, 27(1):1–3, February 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086000922778.

Belcastro:2023:DCN

[BF23] Alessandro Belcastro and Giuseppina Fenaroli. A determination of Catalan numbers in 18th century Italy by Giovanni Rizzetti (1675–1751). Historia Mathematica, 64(??):34–47, August 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086023000368.

Belcastro:1994:CGR

[BFG94] A. Belcastro, G. Fenaroli, and A. C. Garibaldi. Considerazioni di Giovanni Rizzetti sul calcolo delle probabilità e sul teorema di Jakob Bernoulli. Historia Mathematica, 21(4):420–452, November 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086084710329.

Birkhoff:1975:B

[BG75a] Garrett Birkhoff and Sue Ann Garwood. Bibliographies. Historia Mathematica, 2(4):605–615, November 1975. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075901305.

Birkhoff:1975:P

[BG75b] Garrett Birkhoff and Sue Ann Garwood. Preface. Historia Mathematica, 2(4):429–430, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508607590097X.

Barrow-Green:2000:R

[BG00] June Barrow-Green. Reviews. Historia Mathematica, 27(1):82–85, February 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086099922656.

Barrow-Green:2010:DES

[BG10] June Barrow-Green. The dramatic episode of Sundman. Historia Mathematica, 37(2):164–203, May 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086009001360.

Barrow-Green:2021:AMP

[BG21a] June Barrow-Green. Awarding of the May Prizes for 2021. Historia Mathematica, 56(??):1–3, August 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086021000446.

Barrow-Green:2021:KGE

[BG21b] June Barrow-Green. “Knowledge gained by experience”: Olaus Henrici — engineer, geometer and maker of mathematical models. Historia Mathematica, 54(??):41–76, February 2021. CODEN HI-MADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/S0315086020300744.

Barrow-Green:2006:GUP

[BGG06a] June Barrow-Green and Jeremy Gray. Geometry and its uses in physics, 1900–1930: Four papers from an Open University Conference, Milton Keynes, September 2002. Historia Mathematica, 33 (3):275–276, August 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086006000504.

Barrow-Green:2006:GC

[BGG06b] June Barrow-Green and Jeremy Gray. Geometry at Cambridge, 1863–1940. Historia Mathematica, 33(3):315–356, August 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005000819.

Barrow-Green:2010:E

[BGG10] June Barrow-Green and Niccolò Guicciardini. Editorial. Historia Mathematica, 37(1):1–2, February 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086009001232.

Bashmakova:1976:APY

[BGM^{+}76] Isabella G. Bashmakova, A. T. Grigoryan, A. I. Markushevich, F. A. Medvedev, and B. A. Rozenfeld. Adolph Pavlovich Yushkevich (on the occasion of his seventieth anniversary). Historia Mathematica, 3(3):259–278, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076900963.

Barrow-Green:2016:FMS

[BGSS16] June Barrow-Green and Reinhard Siegmund-Schultze. “The first man on the street” — tracing a famous Hilbert quote (1900) back to Gergonne (1825). Historia Mathematica, 43(4):415–426, November 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086016300465.

Brentjes:1989:NTI

[BH89] Sonja Brentjes and Jan P. Hogendijk. Notes on Thābit ibn Qurra and his rule for amicable numbers. Historia Mathematica, 16(4):373–378, November 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086089900840.

Belhoste:2014:RMM

[BH14] Bruno Belhoste and Denise Hazebrouck. Récréations et mathématiques mondaines au XVIII^{e} siècle: le cas de Guyot. Historia Mathematica, 41(4):490–505, November 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086014000779.

Bidwell:1974:BRM

[Bid74] James K. Bidwell. Book review: The Mathematical Association of America: Its first fifty years: Edited by Kenneth O. May. The Mathematical Association of American, 1972. 180 p. $8.00. Historia Mathematica, 1(3):342–344, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074900822.

Biermann:1974:BRM

[Bie74a] Kurt-R. Biermann. Book review: Mathematicians from antiquity to today. A preliminary edition. Vol. I: By J. Fang, in collaboration with U. Dudley. Hauppauge, New York (Paideia Press), 1972. 341 p. $12.80. Historia Mathematica, 1(3):337–339, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074900792.

Biermann:1974:UDM

[Bie74b] Kurt-R. Biermann. Übersiedlung eines deutschen Mathematikers von Braunschweig nach Kaza’n im Jahre 1807/08: Zur Biographie von M. Bartels, Des Lehrers von Gauss und Lobacevskij. (German) [Settlement of a German mathematician of Braunschweig in Kaza’n in the years 1807/08: Biography of M. Bartels, the teacher of Gauss and Lobacevskij]. Historia Mathematica, 1(1):65–77, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074901682.

Biermann:1978:CIJ

[Bie78] K.-R Biermann. Colloque International J. H. Lambert la vie, l’oeuvre et l’influence de Jean-Henri Lambert 26–30 Septembre 1977. Historia Mathematica, 5(3):348–351, August 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078901234.

Biermann:1980:ABV

[Bie80] Kurt-R. Biermann. Apropos Briefwechsel A. v. Humboldts mit C. F. Gauss. (German) [On correspondence of A. v. Humboldt with C. F. Gauss]. Historia Mathematica, 7(2):189–190, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900397.

Biermann:1983:CFG

[Bie83] Kurt-R. Biermann. C. F. Gauss als mathematik- und astronomiehistoriker. Historia Mathematica, 10(4):422–434, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900046.

Biggs:1979:RC

[Big79] N. L. Biggs. The roots of combinatorics. Historia Mathematica, 6(2):109–136, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086079900740.

Biggs:2019:JRM

[Big19] Norman Biggs. John Reynolds of the Mint: a mathematician in the service of King and Commonwealth. Historia Mathematica, 48(??): 1–28, August 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086018300041.

Billoux:1985:CMC

[Bil85] Claudine Billoux. La “correspondance mathématique” de Camille Jordan dans les archives de l’École Polytechnique. Historia Mathematica, 12(1):80–81, February 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508608590076X.

Binder:1983:HMO

[Bin83] Christa Binder. History of mathematics, Oberwolfach, 1981. Historia Mathematica, 10(2):214–223, May 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508608390068X.

Binder:1989:IOS

[Bin89] Christa Binder. II. Österreichisches Symposium zur Geschichte der Mathematik. Historia Mathematica, 16(2):166, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086089900402.

Binder:1990:RFA

[Bin90] Christa Binder. Report on the first Austrian symposium on the history of mathematics: Mathematics-Stimulating or stimulated?: Neuhofen an der Ybbs, Lower Austria, 9 to 15 November 1986. Historia Mathematica, 17(2):155–157, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609090055I.

Binder:1991:RSA

[Bin91] Christa Binder. Report on the Second Austrian Symposium on the History of Mathematics: Mathematics-à la mode?: Neuhofen an der Ybbs, Lower Austria, 22 to 28 October 1989. Historia Mathematica, 18(4):361–362, November 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086091903769.

Birkhoff:1974:BRO

[Bir74] Garrett Birkhoff. Book review: Origins of modern algebra: By Lubos Novy. Prague (Academia), 1973. 260 p. Historia Mathematica, 1(3):333–337, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074900780.

Birkhoff:1975:Ia

[Bir75a] Garrett Birkhoff. Introduction. Historia Mathematica, 2(4):437–438, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086075901007.

Birkhoff:1975:Ib

[Bir75b] Garrett Birkhoff. Introduction. Historia Mathematica, 2(4):473–474, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086075901081.

Birkhoff:1975:Ic

[Bir75c] Garrett Birkhoff. Introduction. Historia Mathematica, 2(4):503–505, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086075901123.

Birkhoff:1975:Id

[Bir75d] Garrett Birkhoff. Introduction. Historia Mathematica, 2(4):535–536, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086075901172.

Birkhoff:1975:Ie

[Bir75e] Garrett Birkhoff. Introduction. Historia Mathematica, 2(4):573–574, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086075901238.

Birkhoff:1977:BRA

[Bir77] Garrett Birkhoff. Book review: Adventures of a mathematician: By S. M. Ulam. New York (Scribners). 1976. 317 pp. $17.50. Historia Mathematica, 4(2):233–236, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086077901318.

Birkhoff:1978:BRH

[Bir78] Garrett Birkhoff. Book review: A History of Numerical Analysis from the 16th Through the 19th Century, by Herman H. Goldstine. Historia Mathematica, 5(4):479–482, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078902264.

Birkhoff:1980:BRE

[Bir80] Garrett Birkhoff. Book review: Die entstehung der verbandsteorie: By Herbert Mehrtens. Hildesheim (Gerstenberg Verlag). 1979. ix + 363 pp. 42 DM. Historia Mathematica, 7(4):454–457, November 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900154.

Birkhoff:1983:BRJ

[Bir83] Garrett Birkhoff. Book review: John von Neumann and Norbert Wiener: By Steve Heims. Cambridge, Mass., and London (M.I.T. Press). 1981. Historia Mathematica, 10(2):243–248, May 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900745.

Birkeland:1996:LSL

[Bir96] Bent Birkeland. Ludvig Sylow’s lectures on algebraic equations and substitutions, Christiania (Oslo), 1862: an introduction and a summary. Historia Mathematica, 23(2):182–199, May 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086096900178.

Birkhoff:1997:BRJ

[Bir97] Garrett Birkhoff. Book review: John von Neumann. By Norman Macrae. New York (Random House). 1992. Historia Mathematica, 24(1):98–100, February 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086097921376.

Bishop:1975:CCM

[Bis75] Errett Bishop. The crisis in contemporary mathematics. Historia Mathematica, 2(4):507–517, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075901135.

Bekken:1989:HMG

[BJ89a] Otto B. Bekken and Bengt Johansson. History of mathematics Gimlekollen Mediasenter, Kristiansand, Norway, August 7–13, 1988. Historia Mathematica, 16(1):90–92, February 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086089901031.

Butzer:1989:PCH

[BJ89b] Paul L. Butzer and François Jongmans. P. L. Chebyshev (1821– 1894) and his contacts with Western European scientists. Historia Mathematica, 16(1):46–68, February 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086089900980.

Berggren:2000:PGA

[BJ00] J. L. Berggren and Alexander Jones. Ptolemy’s Geography: an annotated translation of the theoretical chapters. Princeton University Press, Princeton, NJ, USA, 2000. ISBN 0-691-01042-0, 0-691-09259-1 (paperback). xiii + 192 + 8 pp. LCCN GA881.1 .B47 2000.

Birkhoff:1984:EFA

[BK84] Garrett Birkhoff and Erwin Kreyszig. The establishment of functional analysis. Historia Mathematica, 11(3):258–321, August 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086084900363.

Bhaaskara:2006:EMS

[BK06] Bhāaskara and Agathe Keller. Expounding the mathematical seed: a translation of Bhāaskara I on the mathematical chapter of the āAryabhaòtāiya, volume 30–31 of Science networks. Historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2006. ISBN 3-7643-7299-0 (set), 3-7643-7594-9 (set: e-), 3-7643-7291-5 (v. 1), 3-7643-7592-2 (v. 1: e-), 3-7643-7292-3 (v. 2), 3-7643-7593-0 (v. 2: e-). ???? pp. LCCN QA27.I4 A7723 2006. URL

http://www.loc.gov/catdir/enhancements/fy0663/2006040242-
d.html;   http://www.loc.gov/catdir/enhancements/fy0812/
2006040242-t.html.

Bashmakova:1982:BRA

[BKM^{+}82] Isabella G. Bashmakova, A. N. Kolmogorov, A. I. Markushevitch, A. N. Parshin, and A. P. Youshkevitch. Book review: Abrégé d’histoire des mathématiques, 1700–1900: Edited by J. Dieudonné. I. Algèbre, analyse classique, théorie des nombres. II. Fonctions elliptiques, analyse fonctionnelle, géométrie différentielle, topologie, probabilités, logique mathématique. Sous la direction de J. Dieudonné, de l’Institut. Avec la collaboration de P. Dugac, W. J. et Fern Ellison, J. Guérindon, M. Guillaume, G. Hirsch, C. Houzel, Paulette Libermann, M. Loève, et J.-L. Verley. Paris (Hermann), 1978. x + 385, vii + 472. Historia Mathematica, 9(3):346–360, August 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082901318.

Barbeau:1976:EPD

[BL76] E. J. Barbeau and P. J. Leah. Euler’s 1760 paper on divergent series. Historia Mathematica, 3(2):141–160, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086076900306.

Blackmore:1976:BRE

[Bla76] John T. Blackmore. Book review: The Einstein decade: By Cornelius Lanczos. New York and London (Academic Press), 1974. $12.50.

Historia Mathematica, 3(2):236, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086076900598.

Blay:2005:BRG

[Bla05] Michel Blay. Book review: Gottfried Wilhelm Leibniz. Sämtliche Schriften und Briefe, VII. Reihe. Mathematische Schriften, Band 3: 1672–1676. Differenzen, Folgen, Reihen. Siegmund Probst, Wolfgang Knobloch, Nora Gädeke (Eds.), Akademie Verlag, Munich, 2003, 880 pp. ISBN 3-05-004003-3. 258 EUR . Historia Mathematica, 32(3):362– 363, August 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086004000850.

Blaasjo:2012:RQC

[Blå12] Viktor Blåsjö. The rectification of quadratures as a central foundational problem for the early Leibnizian calculus. Historia Mathematica, 39(4):405–431, November 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086012000365.

Blanco:2014:TSW

[Bla14] Mónica Blanco. Thomas Simpson: Weaving fluxions in 18thcentury London. Historia Mathematica, 41(1):38–81, February 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086013000438.

Blaasjo:2017:WBC

[Blå17a] Viktor Blåsjö. On what has been called Leibniz’s rigorous foundation of infinitesimal geometry by means of Riemannian sums. Historia Mathematica, 44(2):134–149, May 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086017300307.

Blaasjo:2017:RK

[Blå17b] Viktor Blåsjö. Reply to Knobloch. Historia Mathematica, 44(4): 420–422, November 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086017300794.

Blaasjo:2023:NCG

[Blå23] Viktor Blåsjö. Newton on constructions in geometry. Historia Mathematica, 65(??):14–29, November 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086023000666.

Butzer:1992:AAS

[BLF92] Paul L. Butzer, Dietrich Lohrmann, and Ludwig Falkenstein. 4. Aachen Alcuin Symposium Science and History in Western and Eastern Civilization, 8th and 9th Centuries: Aachen, Germany, September 25–28, 1991. Historia Mathematica, 19(2):209–210, May 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290085P.

Biggs:1977:CPM

[BLW77] Norman L. Biggs, E. Keith Lloyd, and Robin J. Wilson. C. S. Peirce and de Morgan on the Four-Colour Conjecture. Historia Mathematica, 4(2):215–216, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086077901185.

Brock:1974:LJT

[BM74] W. H. Brock and R. M. Macleod. The life and journals of Thomas Archer Hirst, FRS (1830–1892). Historia Mathematica, 1(2):181– 183, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074900081.

Bos:1977:IMS

[BM77] H. J. M. Bos and H. Mehrtens. The interactions of mathematics and society in history some exploratory remarks. Historia Mathematica, 4(1):7–30, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086077900313.

Batog:1996:SPB

[BM96] Tadeusz Batóg and Roman Murawski. Stanisław Piatkiewicz and the beginnings of mathematical logic in Poland. Historia Mathematica, 23(1):68–73, February 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086096900051.

Bottcher:2008:SMG

[BMW08] Karl-Heinz Böttcher, Bertram Maurer, and Klaus Wendel. Stuttgarter Mathematiker: Geschichte der Mathematik an der Universität Stuttgart von 1829 bis 1945 in Biographien. (German) [Stuttgart mathematician: History of Mathematics the University of Stuttgart from 1829 to 1945 in biographies], volume 2 of Veröffentlichungen des Universitätsarchivs, 1436–2880. Universität Stuttgart, Stuttgart, Germany, 2008. ISBN 3-926269-34-0. 245 pp. LCCN QA28 .B68 2008.

Bogoljubov:2007:EMS

[BMY07a] N. N. (Nikolaĭ Nikolaevich) Bogoljubov, G. K. Mikhaĭlov, and A. P. (Adolf Pavlovich) Yushkevich, editors. Euler and modern science, volume 4 of The MAA tercentenary Euler celebration; Spectrum series. Mathematical Association of America, Washington, DC, USA, 2007. ISBN 0-88385-564-X. xiv + 425 pp.

LCCN Q143.E84 R3913 2007. URL http://www.loc.gov/catdir/
enhancements/fy0903/2007929594-d.html;   http://www.loc.
gov/catdir/enhancements/fy0903/2007929594-t.html.

Bogolyubov:2007:EMS

[BMY07b] N. N. Bogolyubov, G. K. Mikhaĭlov, and A. P. Yushkevich, editors. Euler and Modern Science, volume 4 of MAA tercentenary Euler celebration: Spectrum series. Mathematical Association of America, Washington, DC, USA, English edition, 2007. ISBN 0-88385-564-X. xiv + 425 pp. LCCN Q143.E84 R3913 2007. Translated by Robert Burns from the 1988 Russian original, Razvitie ideĭ Leonarda Eulera i sovremennaya nauka.

Birkeland:2022:LMS

[BN22] Lene Birkeland and Rolf Nossum. A letter from Malevich to Semevsky about Kovalevskaya. Historia Mathematica, 58(??):92–119, February 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086021000641.

Bo:2024:MGD

[Bo24] Álvaro José Campillo Bo. Book review: Motion and Genetic Definitions in the Sixteenth-Century Euclidean Tradition, Angela Axworthy. (2021). Historia Mathematica, 69(??):67–68, December 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086024000909.

Boi:1989:SGC

[Boi89] Luciano Boi. 1830–1930: Un siècle de géométrie, de C. F. Gauss et B. Riemann à H. Poincaré et E. Cartan; épistémologie, histoire, et mathématiques: Colloque International, 18–23 septembre 1989. Historia Mathematica, 16(3):269–270, August 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086089900219.

Bolling:1993:EMB

[Böl93] Reinhard Bölling. Zum ersten Mal: Blick in einen Brief Kowalewskajas an Weierstraß. (German) [For the first time: A view of a letter of Kowalewskaja to Weierstrass]. Historia Mathematica, 20(2):126– 150, May 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086083710128.

Booker:1986:ISG

[Boo86] George Booker. International study group for the relationship between the history and pedagogy of mathematics (HPM). Historia Mathematica, 13(2):184–190, May 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086086900418.

Borowczyk:1989:UEH

[Bor89] Jacques Borowczyk. Université d’été d’Histoire des Mathématiques, 28 août au 2 septembre 1988: Organisée par la Commission INTER-IREM d’épistémologie et l’IREM de Poitiers. Historia Mathematica, 16(2):167–168, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086089900414.

Borowczyk:1991:VOF

[Bor91] Jacques Borowczyk. Sur la vie et l’oeuvre de François Budan (1761–1840). Historia Mathematica, 18(2):129–157, May 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609190496K.

Borrelli:2014:BRD

[Bor14] Arianna Borrelli. Book review: Le destin douloureux de Walther Ritz (1878–1909), physicien théoricien de génie, Jean-Claude Pont (Ed.). Vallesia, Archive de l’État du Valais, Sion (2012), ISBN 978-2-9700636-5-0. Historia Mathematica, 41(1):107–110, February 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086013000396.

Borrelli:2018:BRZ

[Bor18] Arianna Borrelli. Book review: Zwischen zwei Disziplinen. B. L. van der Waerden und die Entwicklung der Quantenmechanik, Martina R. Schneider. Springer-Verlag, Berlin, Heidelberg (2011), e-ISBN 978-3-642-21825-5. 34,99 EUR (e-book). 411. 44,99 EUR (softcover)., ISBN: 978-3-642-21824-8. Historia Mathematica, 45(2): 189–193, May 2018. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0315086018300260.

Bos:1977:BRP

[Bos77] H. J. M. Bos. Book review: Die Principien der höheren Analysis in ihrer Entwickelung von Leibniz bis auf Lagrange als ein historischkritischer Beitrag zur Geschichte der Mathematik Dargestellt: By Hermann Weissenborn. Halle (H. W. Schmidt). 1856. photomechanic reprint (Leipzig, Zentralantiquariat der Deutschen Demokratischen Republik). 1972. Historia Mathematica, 4(2):236–239, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508607790132X.

Bos:1992:HF

[Bos92] Henk J. M. Bos. Hans Freudenthal (1905–1990). Historia Mathematica, 19(1):106–108, February 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609290070R.

Bos:2002:ITS

[Bos02] Henk J. M. Bos. Out of the Ivory Tower: the significance of Dirk Struik as historian of mathematics. Historia Mathematica, 29(4): 363–368, November 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086002923773.

Bos:2003:BRC

[Bos03] Henk J. M. Bos. Book review: Clifford Truesdell: Eine wissenschaftliche Biographie des Dichters, Mathematikers und Naturphilosophen: By Yurie A. Ignatieff and Heike Willig. Aachen (Shaker-Verlag). 1999. ISBN 3-8265-4804-3. Historia Mathematica, 30(4): 519–521, November 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086003000569.

Bos:2010:PEB

[Bos10] Erik-Jan Bos. Princess Elizabeth of Bohemia and Descartes’ letters (1650–1665). Historia Mathematica, 37(3):485–502, August 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009001256.

Bottazzini:1977:MPE

[Bot77] Umberto Bottazzini. The mathematical papers of Enrico Betti in the Scuola Normale Superiore of Pisa. Historia Mathematica, 4(2):207– 209, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086077901148.

Bottazzini:1978:RPT

[Bot78] Umberto Bottazzini. Ricerche di P. Tardy sui differenziali di indice qualunque, 1844–1868. (Italian) [Research of P. Tardy on differentials of arbitary index, 1844–1868]. Historia Mathematica, 5(4):411–418, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086078902094.

Bottazzini:1980:AUI

[Bot80a] Umberto Bottazzini. Algebraische Untersuchungen in Italien, 1850–1863. Historia Mathematica, 7(1):24–37, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900610.

Bottazzini:1980:TLL

[Bot80b] Umberto Bottazzini. Tardy’s letters and library in Genoa. Historia Mathematica, 7(1):84–85, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900713.

Bottcher:2012:MNL

[Böt12] Frauke Böttcher. Das mathematische und naturphilosophische Lernen und Arbeiten der Marquise du Châtelet (1706–1749) Wissenszugänge einer Frau im 18. Jahrhundert. (German) [The mathematical and philosophical learning and work of the Marquise du Châtelet (1706– 1749). Access to knowledge of a woman in the 18th Century]. Mathematik im Kontext. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-32486-X (paperback), 3-642-32487-8 (e-book). ISSN 2191-074X. xii + 343 pp. LCCN PQ1981 .D55; QA29.D75 B68 2013.

Bowen:1984:BRB

[Bow84] Alan C. Bowen. Book review: The beginnings of Greek mathematics: By Árpád Szabó (translated by A. M. Ungar from Anfänge der Griechischen Mathematik). Synthese Historical Library 17. Dordrecht/Boston (Reidel). 1978. 353 pp. + index (5 pp.). Dfl. 100,- /U.S. $47.50. Historia Mathematica, 11(3):335–345, August 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086084900454.

Bowen:1987:ISP

[Bow87] Alan C. Bowen. The interaction of science and philosophy in fifth and fourth century Greece, May 30–June 1, 1986. Historia Mathematica, 14(3):291–292, August 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086087900565.

Boyer:1974:BRD

[Boy74] Carl B. Boyer. Book review: Differential equations: with applications and historical notes: By George F. Simmons. New York (McGraw-Hill), 1972, 482 pp. Historia Mathematica, 1(1):100–102, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074901918.

Barany:2017:NCW

[BPL17] Michael J. Barany, Anne-Sandrine Paumier, and Jesper Lützen. From Nancy to Copenhagen to the World: The internationalization of Laurent Schwartz and his theory of distributions. Historia Mathematica, 44(4):367–394, November 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086017300320.

Bolzano:2004:MWB

[BR04] Bernard Bolzano and Steve Russ. The mathematical works of Bernard Bolzano. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2004. ISBN 0-19-853930-4. xxx + 698 + 1 pp.

LCCN QA3 .B5775213 2004. URL http://www.loc.gov/catdir/
enhancements/fy0627/2005295416-d.html;   http://www.loc.
gov/catdir/enhancements/fy0627/2005295416-t.html.

Brioschi:2016:EBC

[BR16] Alessio Brioschi and Riccardo Rosso. Eugenio Beltrami’s courses on the analytic and mechanical theory of heat. I. Analytic theory of heat. Historia Mathematica, 43(4):399–414, November 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086016300234.

Brackenridge:1989:NMD

[Bra89] J. Bruce Brackenridge. Newton’s mature dynamics and the Principia: a simplified solution to the Kepler problem. Historia Mathematica, 16(1):36–45, February 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086089900979.

Bradley:2007:BRM

[Bra07] Robert E. Bradley. Book review: Mathematics and the Historian’s Craft: the Kenneth O. May Lectures, Glen Van Brummelen, Michael Kinyon (Eds.). Springer-Verlag, New York (2005), ISBN 0-387-25284-3. Historia Mathematica, 34(2):238–240, May 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086006000528.

Bradley:2009:BRLb

[Bra09a] Robert E. Bradley. Book review: Lacroix and the Calculus. By João Caramalho Domingues. Basel (Birkhäuser). 2008. ISBN 978-3-7643-8637-5. 468 pp. US$169.00. Historia Mathematica, 36(4):437–439, November 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086009000676.

Bradley:2009:BRLa

[Bra09b] Robert E. Bradley. Book review: Leonhard Euler ≪ incomparable géomètregg, Philippe Henry. Editions Médicine et Hygiène, Geneva (2007), ISBN 978-2-88049-241-0. Historia Mathematica, 36(3):281– 283, August 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008001055.

Bradley:2012:BRR

[Bra12] Robert E. Bradley. Book review: Routes of Learning: Highways, Pathways and Byways in the History of Mathematics. By Ivor Grattan-Guinness. Baltimore (The Johns Hopkins University Press). 2009. xii + 372 pp. Historia Mathematica, 39(2):234–236, May 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086011000711.

Bradley:2017:BRB

[Bra17] Robert E. Bradley. Book review: A Brief History of Numbers, Leo Corry. Oxford University Press (2015), 309 + xiii pages. Historia Mathematica, 44(4):423–424, November 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086017300733.

Breger:1991:ICC

[Bre91] Herbert Breger. International conference on le continu mathématique: Cerisy-la-Salle, France, September 11–21, 1990. Historia Mathematica, 18(3):277, August 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609190398H.

Brentjes:2007:BRD

[Bre07] Sonja Brentjes. Book review: Le Développement de la Géométrie aux IXe-XIe Siècles. Abu Sahl al-Quhi, Philippe Abgrall, in: Librairie Scientifique et Technique, Collection Sciences dans l’Histoire. Albert Blanchard, Paris (2004), ISBN 2-85367-221-2. Historia Mathematica, 34(3):346–348, August 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086006000978.

Breard:2012:BRC

[Bré12] Andrea Bréard. Book review: Le calendrier chinois: structure et calculs (104 av. J.-C.-1644). Indétermination céleste et réforme permanente. La construction chinoise officielle du temps quotidien discret à partir d’un temps mathématique caché, linéaire et continu. By Jean-Claude Martzloff. Paris (Honoré Champion). 2009. ISBN 978-2-7453-1911-1. 453 pp. 50 EUR. Historia Mathematica, 39(3):335–338, August 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508601200002X.

Brentjes:2019:MCA

[Bre19] Sonja Brentjes. Mathematical commentaries in Arabic and Persian — purposes, forms, and styles. Historia Mathematica, 47(??):54– 66, ???? 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508601830140X.

Browder:1975:RFA

[Bro75] Felix E. Browder. The relation of functional analysis to concrete analysis in 20th century mathematics. Historia Mathematica, 2 (4):577–590, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086075901251.

Brush:1977:BRJ

[Bru77] Stephen G. Brush. Book review: Joseph Fourier: the man and the physicist: By John Herivel. Oxford (Clarendon Press). 1975. xii + 350 pp. $30.95. Historia Mathematica, 4(2):219–221, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077901215.

Bruneau:2007:BRN

[Bru07] Olivier Bruneau. Book review: Newton et les origines de l’Analyse: 1664–1666, Marco Panza. Albert Blanchard, Paris (2005), ISBN 2-85367-230-1. Historia Mathematica, 34(2):230–231, May 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086006000255.

Biermann:1977:GBA

[BS77a] K.-R Biermann and Horst Sachs. Gauss bicentennial activities in the GDR. Historia Mathematica, 4(3):346–347, August 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900647.

Bruckheimer:1977:SCR

[BS77b] M. Bruckheimer and Y. Salomon. Some comments on R. J. Gillings’ analysis of the 2n table in the Rhind Papyrus. Historia Mathematica, 4(4):445–452, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086077900829. See response [Gil78].

Belanger:2005:SVS

[BS05] Jay Belanger and Dorothy Stein. Shadowy vision: spanners in the mechanization of mathematics. Historia Mathematica, 32(1):76–93, February 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086004000217.

Bernoulli:2006:ACT

[BS06] Jakob Bernoulli and Edith Dudley Sylla. The art of conjecturing, together with letter to a friend on sets in court tennis. The Johns Hopkins University Press, Baltimore, MD, USA, 2006. ISBN 0-8018-8235-4. xx + 430 pp. LCCN QA3 .B48513 2006. URL

http://www.loc.gov/catdir/enhancements/fy0664/2005009296-
b.html;   http://www.loc.gov/catdir/enhancements/fy0664/
2005009296-d.html;   http://www.loc.gov/catdir/toc/ecip0510/
2005009296.html.

Boskoff:2007:BSH

[BS07] Wladimir G. Boskoff and Bogdan D. Suceava. Barbilian spaces: the history of a geometric idea. Historia Mathematica, 34(2):221– 224, May 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086006000851.

Beery:2009:THD

[BS09] Janet Beery and Jacqueline A. Stedall, editors. Thomas Harriot’s doctrine of triangular numbers: the ‘Magisteria magna’. Heritage of European mathematics. European Mathematical Society, Zürich, Switzerland, 2009. ISBN 3-03719-059-0 (hardcover). 135 pp. LCCN QA281 .T45 2009.

Brueckler:2013:TAH

[BS13] Franka Miriam Brueckler and Vladimir Stilinović. Teaching arithmetic in the Habsburg Empire at the end of the 18th Century — a textbook example. Historia Mathematica, 40(3):309–323, August 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086013000177.

Bulmer-Thomas:1981:BRG

[BT81] Ivor Bulmer-Thomas. Book review: Greek mathematics and the architects to Justinian: by John Warren. London (Coach Publishing, 102 St. Paul’s Road, London Nl; Exclusive distributor, ISBS Inc., P. O. Box 555, Forest Grove, OR 97116, U.S.A.). 1976. Paper cover. 14 pp. 11 plans and illustrations. £1.00. Historia Mathematica, 8 (4):482–485, November 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086081900598.

Bulmer-Thomas:1984:BRM

[BT84] Ivor Bulmer-Thomas. Book review: The mathematical science of Christopher Wren: By J. A. Bennett. New York/Cambridge/ London (Cambridge Univ. Press). 1983. viii + 148 pp. 35 figures in text. $29.95. Historia Mathematica, 11(3):330–333, August 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086084900430.

Becchi:2003:EHM

[BTB^{+}03] Antonio Becchi, C. (Clifford) Truesdell, Edoardo Benvenuto, et al., editors. Essays on the history of mechanics: in memory of Clifford Ambrose Truesdell and Edoardo Benvenuto, Between mechanics and architecture. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2003. ISBN 3-7643-1476-1.

LCCN TA350.3 .E87 2003. URL http://www.loc.gov/catdir/
enhancements/fy0812/2003051924-d.html;   http://www.loc.
gov/catdir/enhancements/fy0812/2003051924-t.html.

Buchwald:1977:BRK

[Buc77] Jed Z. Buchwald. Book review: The kind of motion we call heat, a history of the kinetic theory of gases in the Nineteenth Century: By Stephen G. Brush. Vol. I: Physics and the Atomists. Vol. II: Statistical Physics and Irreversible Processes. Amsterdam-New York-Oxford (North Holland). 1976. xxix + 769 pp. $83.95. Historia Mathematica, 4(4):466–467, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086077900933.

Bucciarelli:1980:BRW

[Buc80] Louis L. Bucciarelli. Book review: Women and mathematics, science and engineering. A partially annotated bibliography with emphasis on mathematics and with references on related topics: By Else Høyrop. Roskilde University Library, Denmark. 1978. Historia Mathematica, 7(4):451, November 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086080900129.

Buckle:2023:HEY

[Buc23] David Buckle. How the estimate of \sqrt{2} on YBC 7289 may have been calculated. Historia Mathematica, 62(??):3–18, February 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086022000477.

Bullynck:2009:DPT

[Bul09a] Maarten Bullynck. Decimal periods and their tables: a German research topic (1765–1801). Historia Mathematica, 36(2):137–160, May 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508600800092X.

Bullynck:2009:MAB

[Bul09b] Maarten Bullynck. Modular arithmetic before C. F. Gauss: Systematizations and discussions on remainder problems in 18th-Century Germany. Historia Mathematica, 36(1):48–72, February 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008000670.

Bullynck:2016:BRN

[Bul16a] Maarten Bullynck. Book review: Newton and the Origin of Civilization, Jed Z. Buchwald, Mordechai Feingold. Princeton University Press, Princeton, New Jersey (2013), ISBN: 978-0-691-15478-7. Historia Mathematica, 43(2):213–215, ???? 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086015000476.

Bullynck:2016:HAP

[Bul16b] Maarten Bullynck. Histories of algorithms: Past, present and future. Historia Mathematica, 43(3):332–341, August 2016.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086015000725.

Bullynck:2022:SJH

[Bul22] Maarten Bullynck. Supplementband: Johann Heinrich Lamberts Monatsbuch. Neu herausgegeben, eingeleitet, kommentiert und mit Verzeichnissen zu Lamberts Schriften, Briefen und nachgelassenen Manuskripten versehen (in two volumes), Niels W. Bokhove, Armin Emmel (Eds.). Olms, Hildesheim u.a. (2020), Cloth. clvii + 925 pages, ISBN: 978-3-487-15822-8. Historia Mathematica, 60(??):64– 66, August 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086022000064.

Bullynck:2024:BRE

[Bul24] Maarten Bullynck. Book review: Edouardo Noble “The Rise and Fall of the German Combinatorial Analysis”. Basel, Birkhäuser, 2022. Historia Mathematica, 69(??):69–70, December 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086024000910.

Bunge:1974:BRP

[Bun74] Mario Bunge. Book review: Philosophy and mathematics: From Plato to the present: By Robert J. Baum. San Francisco, California (Freeman, Cooper & Co.), 1973. 332 pp. Historia Mathematica, 1(3):350, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508607490086X.

Bunge:1976:BRMa

[Bun76a] Mario Bunge. Book review: From mathematics to philosophy: By Hao Wang. London (Routledge and Kegan Paul), New York (Humanities Press) 1974. 442 p. U.S. $23.90. Historia Mathematica, 3(1):100–102, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086076900240.

Bunge:1976:BRMb

[Bun76b] Mario Bunge. Book review: La mathématisation des doctrines informes: Colloque tenu à l’Institut d’histoire des sciences de l’Université de Paris sous la direction de George Canguilhem. Paris (Hermann). 1972. 237 pp. Historia Mathematica, 3(2):235, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900574.

Burckhardt:1975:ABJ

[Bur75] Johann Jakob Burckhardt. Address on the 65th birthday of J. E. Hofmann at Oberwolfach. Historia Mathematica, 2(2):137–146, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508607590138X.

Burckhardt:1986:EWN

[Bur86] J. J. Burckhardt. Euler’s work on number theory: a concordance for A. Weil’s number theory. Historia Mathematica, 13(1):28–35, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086086902235.

Burchsted:1987:AAM

[Bur87] Frederic F. Burchsted. Archives of American mathematics. Historia Mathematica, 14(4):366–374, November 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508608790067X.

Burn:1992:INE

[Bur92] R. P. Burn. Irrational numbers in English language textbooks, 1890–1915: Constructions and postulates for the completeness of the real numbers. Historia Mathematica, 19(2):158–176, May 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290074L.

Burchsted:1993:NAA

[Bur93] Frederic F. Burchsted. New archives of American mathematics collections. Historia Mathematica, 20(1):92–94, February 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086083710098.

Burn:2001:AAS

[Bur01] R. P. Burn. Alphonse Antonio de Sarasa and logarithms. Historia Mathematica, 28(1):1–17, February 2001. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508600092295X.

Burnett:2008:BRC

[Bur08] Charles Burnett. Book review: Campanus of Novara and Euclid’s Elements, 2 vols, Band 51 (1–2), H. L. L. Busard, in: Boethius. Franz Steiner-Verlag, Stuttgart (2005). Historia Mathematica, 35(1):50–51, February 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086007000043.

Burdick:2009:MWP

[Bur09] Bruce Stanley Burdick. Mathematical works printed in the Americas, 1554–1700. Johns Hopkins studies in the history of mathematics. The Johns Hopkins University Press, Baltimore, MD, USA, 2009. ISBN 0-8018-8823-9 (hardcover). x + 373 pp. LCCN Z6652 .B87 2009. URL

http://www.loc.gov/catdir/toc/ecip0819/2008024009.html.

Burgess:2010:PN

[Bur10] Tyge Burgess. Publisher’s note. Historia Mathematica, 37(1):3, February 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086009001220.

Burn:2012:ATC

[Bur12] R. P. Burn. Another theorem of Cauchy which ‘admits exceptions’. Historia Mathematica, 39(2):206–210, May 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086011000772.

Busard:1976:UEE

[Bus76] H. L. L. Busard. Über die Überlieferung der Elemente Euklids über die Länder des Nahen Ostens nach West-Europa. (German) [On the lore of the Elements of Euclid in the countries of the nearby East to West Europe]. Historia Mathematica, 3(3):279–290, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900975.

Butzer:1981:OLW

[But81] Paul L. Butzer. An outline of the life and work of E. B. Christoffel (1829–1900). Historia Mathematica, 8(3):243–276, August 1981.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900689.

Berggren:2005:AKR

[BV05] J. L. Berggren and Glen Van Brummelen. Al-Kuhi’s revision of Book I of Euclid’s Elements. Historia Mathematica, 32(4):426–452, November 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086005000327.

Calinger:1975:BRM

[Cal75] Ronald Calinger. Book review: Die mathematik und ihre Dozenten an der Berliner Universität 1810–1920: By Kurt-R. Biermann. Berlin (Akademie-Verlag), SCV2 1973. 265 p. with 14 portraits. Historia Mathematica, 2(3):367–369, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075900932.

Calinger:1983:BRH

[Cal83] Ronal Calinger. Book review: A history of the calculus of variations from the 17th through the 19th century: By Herman H. Goldstine. New York (Springer-Verlag). 1980. Studies in the History of Mathematics and Physical Sciences 5, xvii + 410 pp. Historia Mathematica, 10(1):118–121, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086083900538.

Calinger:1992:BRI

[Cal92] Ronald Calinger. Book review: Introduction to analysis of the infinite, book I: By Leonhard Euler; translated by John D. Blanton. New York (Springer-Verlag). 1988. xiii + 327 pp. Historia Mathematica, 19(2):214–216, May 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609290090X.

Calinger:1996:LEF

[Cal96] Ronald Calinger. Leonhard Euler: the first St. Petersburg years (1727–1741). Historia Mathematica, 23(2):121–166, May 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086096900154.

Calinger:2008:BRE

[Cal08] Ronald Calinger. Book review: The Early Mathematics of Leonhard Euler, C. Edward Sandifer. The Mathematical Association of America, Washington, DC (2007), ISBN 0-88385-559-3. Historia Mathematica, 35(2):140–141, May 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086007000900.

Cambell:1976:ECA

[Cam76] Paul J. Cambell. An experimental course of African mathematics. Historia Mathematica, 3(4):477–478, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900823.

Campbell:1977:GEQ

[Cam77] Paul J. Campbell. Gauss and the Eight Queens Problem: a study in miniature of the propagation of historical error. Historia Mathematica, 4(4):397–404, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086077900763.

Campbell:1978:OZL

[Cam78] Paul J. Campbell. The origin of “Zorn’s Lemma”. Historia Mathematica, 5(1):77–89, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086078901362.

Caparrini:2013:BRD

[Cap13] Sandro Caparrini. Book review: La via delle acque (1550–1700): Appropriazione delle arti e trasformazione delle matematiche, Cesare S. Maffioli. Olschki, Firenze (2010), ISBN: 978-88-222-6008-6. Historia Mathematica, 40(1):89–91, February 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086012000341.

Carolino:2007:CBE

[Car07] Luís Miguel Carolino. Cristoforo Borri and the epistemological status of mathematics in seventeenth-century Portugal. Historia Mathematica, 34(2):187–205, May 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086006000577.

Carman:2018:AOI

[Car18] Christián C. Carman. Accounting for overspecification and indifference to visual accuracy in manuscript diagrams: a tentative explanation based on transmission. Historia Mathematica, 45(3):ifc, August 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086018300648.

Carman:2022:FVM

[Car22] Christián C. Carman. François Viète’s method for calculating the eccentricity in a bisected model and its possible application to Kepler’s Vicarious Hypothesis. Historia Mathematica, 58(??):71–91, February 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086021000665.

Carman:2024:FTM

[Car24] Christián C. Carman. Fabricius’s theory for Mars: the model that shocked Kepler. Historia Mathematica, 69(??):1–21, December 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086024000338.

Cassedy:1977:BRH

[Cas77] James H. Cassedy. Book review: On the history of statistics and probability: Edited by D. B. Owen. Proceedings of a Symposium on the American Mathematical Heritage to Celebrate the Bicentennial of the United States of America, held at Southern Methodist University, 27–29 May 1974. New York and Basel (Marcel Dekker). 1976. 466 p. Illustrations. Bibliographies. Index. Historia Mathematica, 4(2):231– 232, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508607790129X.

Cassedy:1980:BRH

[Cas80] James H. Cassedy. Book review: The history of statistics in the 17th and 18th centuries against the changing background of intellectual, scientific and religious thought: Lectures by Karl Pearson given at University College, London, during the academic sessions 1921–1933. Edited by E. S. Pearson. New York, Macmillan, 1978. xix, 744 pp. Historia Mathematica, 7(2):210–212, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900543.

Caveing:1985:TBA

[Cav85] Maurice Caveing. La tablette babylonienne AO 17264 du musée du Louvre et le problème des six frères. Historia Mathematica, 12(1):6– 24, February 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086085900655.

Caveing:1995:BRM

[Cav95] Maurice Caveing. Book review: Mathématiques Égyptiennes. Recherches sur les connaissances mathématiques de l’Égypte pharaonique: By Sylvia Couchoud. Paris: Éditions Le Léopard d’Or. 1993. 208 pp. 150 FF. Historia Mathematica, 22(1):80–83, February 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086085710099.

Cerroni:2021:CMP

[CB21] Cinzia Cerroni and Aldo Brigaglia. The “Circolo Matematico di Palermo” and the First World War: the crisis of scientific internationalism: a view through the unedited correspondence of De Franchis with Edmund Landau and other mathematicians. Historia Mathematica, 55(??):64–94, May 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086021000252.

Cerroni:2004:NDG

[Cer04] Cinzia Cerroni. Non-Desarguian geometries and the foundations of geometry from David Hilbert to Ruth Moufang. Historia Mathematica, 31(3):320–336, August 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086003000491.

Cerroni:2017:TCS

[Cer17] Cinzia Cerroni. From the theory of “congeneric surd equations” to “Segrè’s bicomplex numbers”. Historia Mathematica, 44(3):232–251, August 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086017300241.

Chabas:1992:NHH

[CG92] José Chabás and Bernard R. Goldstein. Nicholaus de Heybech and his table for finding true syzygy. Historia Mathematica, 19(3):265– 289, August 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609290030F.

Crossley:1990:TSA

[CH90] John N. Crossley and Alan S. Henry. Thus spake al-Khāwarizmi: a translation of the text of Cambridge University Library Ms. II.vi.5. Historia Mathematica, 17(2):103–131, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609090048I.

Chang:1976:BRL

[Cha76] Wei-Ching Chang. Book review: The life and times of the Central Limit Theorem: By William J. Adams. New York (Kaedmon). 1974. 119 p. Historia Mathematica, 3(3):353–354, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508607690118X.

Chandler:1977:PTM

[Cha77] Philip P. Chandler. The Principia ’s theory of the motion of the lunar apse. Historia Mathematica, 4(4):405–410, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900775.

Charbonneau:1981:CAC

[Cha81] Louis Charbonneau. Congrès annuel, Canadian Society for the History and Philosophy of Mathematics/Société canadienne d’histoire et de philosophie des mathematiques. Historia Mathematica, 8(2):191– 192, May 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086081900306.

Charbonneau:1987:CAC

[Cha87] Louis Charbonneau. Congrès annuel-1986: Canadian society for the history and philosophy of mathematics/Société canadienne d’histoire et de philosophie des mathématiques. Historia Mathematica, 14 (4):375–380, November 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086087900681.

Chabert:1990:DSF

[Cha90] Jean-Luc Chabert. Un demi-siècle de fractales: 1870–1920. (French) [A half-century of fractals: 1870–1920]. Historia Mathematica, 17 (4):339–365, November 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609090026A.

Chern:1984:BRG

[Che84] S. S. Chern. Book review: Geschichte des mannigfaltigkeitsbegriffs von Riemann bis Poincaré: By Erhard Scholz. Boston (Birkhauser). 1980. 430 pp. Historia Mathematica, 11(3):330, August 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086084900429.

Chemla:1987:CHM

[Che87] Karine Chemla. Conference on the history of Mediterranean mathematics: Centre International de Rencontres Mathématiques, Luminy, Marseille April 16–21, 1984. Historia Mathematica, 14(2):191– 193, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086087900309.

Chemla:1994:BRM

[Che94] Karine Chemla. Book review: Les Méthodes rapides pour la trigonométrie et le rapport précis du cercle (1774). Tradition chinoise et apport occidental en mathématiques: By Catherine Jami, with a foreword by Jacques Gernet. Memoires de l’Institut des Hautes Études Chinoises, Vol. XXXII. Paris (Collége de France, Institut des Hautes Études Chinoises). 1990. 230 pp. Historia Mathematica, 21(2):220–224, May 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086084710226.

Chemla:2009:MPH

[Che09] Karine Chemla. On mathematical problems as historically determined artifacts: Reflections inspired by sources from ancient China. Historia Mathematica, 36(3):213–246, August 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008001109.

Chen:2010:ETM

[Che10] Jiang-Ping Jeff Chen. The evolution of transformation media in spherical trigonometry in 17th- and 18th-Century China, and its relation to “Western learning”. Historia Mathematica, 37(1):62–109, February 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086009000329.

Chemla:2012:MEM

[Che12] Karine Chemla. The mathematics of Egypt, Mesopotamia, China, India, and Islam. A sourcebook. Historia Mathematica, 39(3):324– 334, August 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086012000316.

Chemla:2014:EHM

[Che14a] Karine Chemla. Explorations in the history of mathematical recreations: an introduction. Historia Mathematica, 41(4):367–376, November 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086014000792.

Chemla:2014:TGK

[Che14b] Karine Chemla. Translated into German by Klaus Volkert as Cantor und die Franzosen. Mathematik, Philosophie und das Unendliche Mathematik im Kontext. Berlin (Springer). 2011. ISBN 978-3-642-14868-2. 230 pp. 35 EUR. Historia Mathematica, 41(3):342–344, August 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086014000093.

Chen:2017:SRT

[Che17] Zhihui Chen. Scholars’ recreation of two traditions of mathematical commentaries in late eighteenth-century China. Historia Mathematica, 44(2):105–133, May 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086017300253.

Chin:2022:MLA

[Chi22] Siu A. Chin. Modern light on ancient feud: Robert Hooke and Newton’s graphical method. Historia Mathematica, 60(??):1–14, August 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508602200009X.

Choquet:1975:NTS

[Cho75] Gustave Choquet. Notice sur les travaux scientifiques. Historia Mathematica, 2(2):153–160, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508607590141X.

Chowdhury:1986:END

[Cho86] M. R. Chowdhury. Emmy Noether Day at ICTP. Historia Mathematica, 13(2):175–176, May 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086086900364.

Chowdhury:1992:WHY

[Cho92] M. R. Chowdhury. W. H. Young at Aberystwyth. Historia Mathematica, 19(1):71–75, February 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609290056H.

Christian:1984:GCJ

[Chr84] Curt C. Christian. Gödel’s contribution to the justification of Leibniz’ notion of the infinitesimals. Historia Mathematica, 11(2):215– 219, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086084800105.

Christianidis:1991:TPD

[Chr91] Jean Christianidis. ’Aρíμητικη Στoιχϵιωσις: Un traité perdu de Diophante d’Alexandrie? Historia Mathematica, 18(3):239–246, August 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609190392B.

Christianidis:1998:IBD

[Chr98] Jean Christianidis. Une interprétation byzantine de Diophante. (French) [A Byzantine interpretation of Diophantus]. Historia Mathematica, 25(1):22–28, February 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086097921728.

Christianidis:2007:WDS

[Chr07] Jean Christianidis. The way of Diophantus: Some clarifications on Diophantus’ method of solution. Historia Mathematica, 34(3):289– 305, August 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086006001285.

Cibrario:1976:BRS

[Cib76] Maria Cinquini Cibrario. Book review: Selected works of Giuseppe Peano: Translated and edited by Hubert C. Kennedy. Toronto (University of Toronto Press). 1973. 249 p. $12.50. Historia Mathematica, 3(2):230–232, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076900550.

Cifoletti:2000:BRJ

[Cif00] Giovanna C. Cifoletti. Book review: Johannes Faulhaber, 1580–1635, Rechenmeister in einer Welt des Umbruchs. By Ivo Schneider. Historia Mathematica, 27(4):440–442, November 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086000922882.

Chase:1984:BCM

[CJ84] Gene B. Chase and Calvin Jongsma. Bibliography of Christianity and mathematics: 1910–1983. Historia Mathematica, 11(2):220–221, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086084800129.

Christianidis:1992:ES

[CK92] Jean P. Christianidis and Nikos Kastanis. Evangelos S. Stamatis (1898–1990). Historia Mathematica, 19(1):99–105, February 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290069N.

Clulee:1976:BRM

[Clu76] Nicholas H. Clulee. Book review: The mathematical praeface to the Elements of Geometrie of Euclid of Megara (1570): By John Dee. Introduction by Allen G. Debus. New York (Science History Publication). 1975. 94 p. US$8.95. Historia Mathematica, 3(3):355–357, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086076901191.

Cogliati:2018:CSS

[CM18] A. Cogliati and P. Mastrolia. Cartan, Schouten and the search for connection. Historia Mathematica, 45(1):39–74, February 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508601730071X.

Christianidis:2019:CTE

[CM19a] Jean Christianidis and Athanasia Megremi. Corrigendum to “Tracing the early history of algebra: Testimonies on Diophantus in the Greek–speaking world (4th-7th century CE)” [Historia Mathematica 47 (2019) 16–38]. Historia Mathematica, 49(??):115, November 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508601930059X. See [CM19b].

Christianidis:2019:TEH

[CM19b] Jean Christianidis and Athanasia Megremi. Tracing the early history of algebra: Testimonies on Diophantus in the Greek-speaking world (4th–7th century CE). Historia Mathematica, 47(??):16–38, ???? 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/

pii/S0315086018301411. See corrigendum [CM19a].

Clark:2022:FSP

[CM22] David Clark and Todd Munson. On five sangaku problems appearing in Yamaguchi’s travel diary. Historia Mathematica, 61(??):1–2, November 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086022000696.

Calude:1979:FER

[CMT79] Cristian Calude, Solomon Marcus, and Ionel Tevy. The first example of a recursive function which is not primitive recursive. Historia Mathematica, 6(4):380–384, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079900247.

Christianidis:2013:PAL

[CO13] Jean Christianidis and Jeffrey Oaks. Practicing algebra in late antiquity: the problem-solving of Diophantus of Alexandria. Historia Mathematica, 40(2):127–163, May 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086012000742.

Cochell:1998:EHC

[Coc98] Gary G. Cochell. The early history of the Cornell mathematics department: a case study in the emergence of the American mathematical research community. Historia Mathematica, 25(2):133–153, May 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086097921832.

Cogliati:2012:BRN

[Cog12] Alberto Cogliati. Book review: The Noether Theorems. Invariance and Conservation Laws in the Twentieth Century. By Yvette Kosmann-Schwarzbach. Translated by Bertram E. Schwarzbach. Dordrecht, Heidelberg, London, New York (Springer). 2011. ISBN 978-0-387-87868-3. xiii + 205 p. 8 illus.; 83,15 EUR. Historia Mathematica, 39(3):353–356, August 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086011000942.

Cogliati:2014:EHI

[Cog14] Alberto Cogliati. Early history of infinite continuous groups, 1883–1898. Historia Mathematica, 41(3):291–332, August 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086014000305.

Cogliati:2015:BRG

[Cog15] Alberto Cogliati. Book review: Les Groupes de Lie dans l’oeuvre de Hermann Weyl. Traduction et commentaire de l’article: Théorie de la répresentation des groupes continus semi-simples par des transformations linéaires (1925–1926), with the collaboration of Amaury Thuillier, Christophe Eckes (Ed.). Presses Universitaires de Nancy, Nancy (2013), ISBN: 978-2-8143-0180-1. Historia Mathematica, 42 (4):491–494, November 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086014001037.

Cogliati:2016:SLC

[Cog16] Alberto Cogliati. Schouten, Levi-Civita and the notion of parallelism in Riemannian geometry. Historia Mathematica, 43(4):427–443, November 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086016300441.

Cohen:1982:BRL

[Coh82] I. Bernard Cohen. Book review: The library of Isaac Newton: By John Harrison. Cambridge, London, New York, Melbourne (Cambridge University Press), 1978. xiv + 286 pp. $62.50. Historia Mathematica, 9(1):106–109, February 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086082901495.

Contro:1985:HMO

[Con85] Walter S. Contro. History of mathematics, Oberwolfach, 1982. Historia Mathematica, 12(2):164–172, May 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086085900102.

Connon:2006:BRA

[Con06] Emma Connon. Book review: Abel’s Proof: an Essay on the Sources and Meaning of Mathematical Unsolvability, Peter Pesic. MIT Press, Cambridge, MA/London (2003), ISBN 0-262-16216-4. Historia Mathematica, 33(3):371–372, August 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086005000625.

Cooke:1986:PSK

[Coo86] Roger Cooke. A portrait of Sophia Kovalevsky. Historia Mathematica, 13(1):74–77, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086086902375.

Cooney:1988:SWM

[Coo88] Miriam P. Cooney. A seminar on women and mathematics. Historia Mathematica, 15(4):380–383, November 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900353.

Cooke:1995:BRB

[Coo95] Roger Cooke. Book review: Briefwechsel zwischen Karl Weierstraß and Sofja Kowalewskaja: Edited with an introduction and commentary by Reinhard Bölling. Berlin (Akademie-Verlag). 1993. 504 pp. Historia Mathematica, 22(1):73–77, February 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086085710075.

Cooke:1996:BRG

[Coo96] Roger Cooke. Book review: The Golden Years of Moscow Mathematics. Edited by Smilka Zdravkovska and Peter Duren. London and Providence (London Mathematical Society and American Mathematical Society). 1993. Historia Mathematica, 23(2):200–203, May 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508609690018X.

Cooke:1999:BRL

[Coo99] Roger Cooke. Book review: L’Europe mathématique/Mathematical Europe. Edited by Catherine Goldstein, Jeremy Gray, and Jim Ritter. Historia Mathematica, 26(2):173–175, May 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086099922358.

Cooper:2010:NIP

[Coo10] Leon Cooper. A new interpretation of problem 10 of the Moscow mathematical papyrus. Historia Mathematica, 37(1):11–27, February 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009000305.

Cooper:2011:DES

[Coo11] Leon Cooper. Did Egyptian scribes have an algorithmic means for determining the circumference of a circle? Historia Mathematica, 38 (4):455–484, November 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086011000449.

Corry:1996:BRO

[Cor96] Leo Corry. Book review: oCorps et modèles-Essai sur l’histoire de l’algèbre réelle. By Hourya Sinaceur. Paris: Librairie Philosophique J. Vrin. Historia Mathematica, 23(3):323–327, August 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086096900312.

Corry:2004:DHA

[Cor04] Leo Corry. David Hilbert and the Axiomatization of Physics (1898– 1918): From Grundlagen der Geometrie to Grundlagen der Physik, volume 10 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2004. ISBN 1-4020-2777-X (hardcover), 1-4020-2778-8 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xvii + 513

pp. LCCN QA29.H5 C67 2004. URL http://www.loc.gov/catdir/
enhancements/fy0823/2006280837-d.html;   http://www.loc.
gov/catdir/enhancements/fy0823/2006280837-t.html.

Corry:2016:SDL

[Cor16] Leo Corry. Some distributivity-like results in the medieval arithmetic of Jordanus Nemorarius and Campanus de Novara. Historia Mathematica, 43(3):310–331, August 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086016300222.

Coxeter:1975:STC

[Cox75] H. S. M. Coxeter. The space-time continuum. Historia Mathematica, 2(3):289–298, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086075900671.

Coxeter:1977:GG

[Cox77] H. S. M. Coxeter. Gauss as a geometer. Historia Mathematica, 4 (4):379–396, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086077900751.

Cozzoli:2012:BRL

[Coz12] Daniele Cozzoli. Book review: “L’opere stupende dell’arti più ingegnose.” La recezione degli Pneumatiká di Erone Alessandrino nella cultura italiana del Cinquecento. By Oreste Trabucco. Firenze (Olschki). 2010. ISBN 978-88-222-5990-5. 184 pp. 20 EUR. Historia Mathematica, 39(2):224–226, May 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086011000486.

Cogliati:2022:OFT

[CR22] Alberto Cogliati and Rachele Rivis. The origins of the fundamental theorem of surface theory. Historia Mathematica, 61(??):45–79, November 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086022000660.

Craik:1998:GAB

[Cra98] Alex D. D. Craik. Geometry, analysis, and the baptism of slaves: John West in Scotland and Jamaica. Historia Mathematica, 25(1):29–74, February 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086097921741.

Craik:1999:CAE

[Cra99] Alex D. D. Craik. Calculus and analysis in early 19th-Century Britain: the work of William Wallace. Historia Mathematica, 26 (3):239–267, August 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086099922504.

Craik:2000:GVA

[Cra00] Alex D. D. Craik. Geometry versus analysis in early 19th-Century Scotland: John Leslie, William Wallace, and Thomas Carlyle. Historia Mathematica, 27(2):133–163, May 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086099922644.

Craik:2003:LTE

[Cra03] Alex D. D. Craik. The logarithmic tables of Edward Sang and his daughters. Historia Mathematica, 30(1):47–84, February 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086002000186.

Craik:2007:MHM

[Cra07] Alex D. D. Craik. Mr Hopkins’ men: Cambridge reform and British mathematics in the 19th Century. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-84628-790-1 (hardcover), 1-84628-791-X (ebk.). xiv + 405 + 24

pp. LCCN QA14.G7 C73 2007. URL http://www.loc.gov/catdir/
enhancements/fy0825/2006940325-b.html;   http://www.loc.

gov/catdir/enhancements/fy0825/2006940325-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0825/2006940325-t.html.

Craik:2009:PVM

[Cra09] Alex D. D. Craik. A proportional view: the mathematics of James glenie (1750–1817). Historia Mathematica, 36(3):247–272, August 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008001110.

Craik:2012:FBA

[Cra12] Alex D. D. Craik. A forgotten British analyst: Nicolas vilant (1737–1807). Historia Mathematica, 39(2):174–205, May 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086011000759.

Craik:2013:PFE

[Cra13] Alex D. D. Craik. Polylogarithms, functional equations and more: the elusive essays of William Spence (1777–1815). Historia Mathematica, 40(4):386–422, October 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086013000426.

Crepel:1987:PMC

[Cré87] Pierre Crépel. Le premier manuscrit de Condorcet sur le calcul des probabilités (1772). Historia Mathematica, 14(3):282–283, August 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086087900474.

Cretney:2014:OEE

[Cre14] Rosanna Cretney. The origins of Euler’s early work on continued fractions. Historia Mathematica, 41(2):139–156, May 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086014000044.

Cretney:2016:ERE

[Cre16] Rosanna Cretney. Editing and reading early modern mathematical texts in the digital age. Historia Mathematica, 43(1):87–97, February 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086015000452.

Crilly:1978:CAG

[Cri78] Tony Crilly. Cayley’s anticipation of a generalised Cayley– Hamilton theorem. Historia Mathematica, 5(2):211–219, May 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078900526.

Crilly:1982:BSH

[Cri82] A. J. Crilly. The British Society for the History of Mathematics Pembroke College, Oxford, September 1981. Historia Mathematica, 9(1):59–60, February 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086082901380.

Crilly:1986:RCI

[Cri86] Tony Crilly. The rise of Cayley’s invariant theory (1841–1862). Historia Mathematica, 13(3):241–254, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086086900911.

Crilly:1988:DCI

[Cri88] Tony Crilly. The decline of Cayley’s invariant theory (1863– 1895). Historia Mathematica, 15(4):332–347, November 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900250.

Crilly:1999:ACS

[Cri99] Tony Crilly. Arthur Cayley as Sadleirian Professor: a glimpse of mathematics teaching at 19th-Century Cambridge. Historia Mathematica, 26(2):125–160, May 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086099922334.

Crilly:2004:CMJ

[Cri04] Tony Crilly. The Cambridge Mathematical Journal and its descendants: the linchpin of a research community in the early and mid-Victorian age. Historia Mathematica, 31(4):455–497, November 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086004000205.

Crilly:2006:ACM

[Cri06] A. J. Crilly. Arthur Cayley: mathematician laureate of the Victorian age. The Johns Hopkins University Press, Baltimore, MD, USA, 2006. ISBN 0-8018-8011-4. xxi + 609 pp. LCCN QA29

.C39 C75 2006. URL http://www.loc.gov/catdir/bios/jhu052/
2004005682.html;   http://www.loc.gov/catdir/description/
jhu051/2004005682.html;   http://www.loc.gov/catdir/toc/
fy0612/2004005682.html.

Crippa:2017:BRM

[Cri17] Davide Crippa. Book review: Mathematische Schriften, Reihe 7, sechster Band, 1673-1676, Arithmetische Kreisquadratur, Gottfried Wilhelm Leibniz. Akademie-Verlag, Berlin (2012). Historia Mathematica, 44(1):73–76, February 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086016300428.

Crialesi:2024:BLB

[Cri24a] Clelia V. Crialesi. Brill, Leiden-Boston (2022), Review of Di Liscia, Daniel A., and Sylla, Edith D. (eds.) with the collaboration of Bakker, Paul J.J.M. Quantifying Aristotle (‘Medieval and Early Modern Philosophy and Science’, 34), pp. 479. Historia Mathematica, 67(??): 20–22, June 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086024000296.

Crippa:2024:BRR

[Cri24b] Davide Crippa. Book review: Reactionary Mathematics: a Genealogy of Purity, Massimo Mazzotti. University of Chicago Press (2023). Historia Mathematica, 66(??):47–49, March 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086023000708.

Crowe:1975:GAA

[Cro75a] Donald W. Crowe. The geometry of African art II. A catalog of Benin patterns. Historia Mathematica, 2(3):253–271, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900634.

Crowe:1975:TLCb

[Cro75b] Michael J. Crowe. Ten “laws” concerning conceptual change in mathematics. Historia Mathematica, 2(4):469–470, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075901056.

Crowe:1975:TLCa

[Cro75c] Michael J. Crowe. Ten “laws” concerning patterns of change in the history of mathematics. Historia Mathematica, 2(2):161–166, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075901421.

Crowe:1976:BWC

[Cro76] Michael J. Crowe. Bibliography of the writings of Carl B. Boyer. Historia Mathematica, 3(4):397–401, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900665.

Crosland:1977:BRC

[Cro77] Maurice P. Crosland. Book review: Condorcet. From natural philosophy to social mathematics: By Keith Michael Baker. Chicago (University of Chicago Press). 1975. 552 p. Historia Mathematica, 4(2):217– 218, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086077901197.

Cohen:1981:JWM

[CS81] G. L. Cohen and A. G. Shannon. John Ward’s method for the calculation of pi [π ]. Historia Mathematica, 8(2):133–144, May 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900252.

Cullen:2007:SSS

[Cul07] Christopher Cullen. The Suàn shù shu, “Writings on reckoning”: Rewriting the history of early Chinese mathematics in the light of an excavated manuscript. Historia Mathematica, 34(1):10–44, February 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005001084.

Chambris:2022:HUF

[CV22] Christine Chambris and Jana Visnovska. On the history of units in French elementary school arithmetic: the case of proportionality. Historia Mathematica, 59(??):99–118, May 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086021000240.

Crilly:2017:ACR

[CWW17] Tony Crilly, Steven H. Weintraub, and Paul R. Wolfson. Arthur Cayley, Robert Harley and the quintic equation: newly discovered letters 1859–1863. Historia Mathematica, 44(2):150–169, May 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086016300490.

DAmbrosio:1979:HIA

[D’A79] Ubiratan D’Ambrosio. History of Ibero–American mathematics. Historia Mathematica, 6(4):452–453, November 1979. CODEN HI-MADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0315086079900351.

DAntonio:2010:BRS

[D’A10] Lawrence D’Antonio. Book review: The St. Petersburg School of Number Theory. By B. N. Delone. Translated from the Russian by Robert Burns. History of Mathematics, vol. 26. Providence/ London (American Mathematical Society/London Mathematical Society). 2005. ISBN 978-0-8218-3457-2. xv + 278 pp. US$62.00. Historia Mathematica, 37(1):120–124, February 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086009000652.

DAmbrosio:2016:PG

[D’A16] Ubiratan D’Ambrosio. Paulus Gerdes (1952–2014). Historia Mathematica, 43(2):129–132, ???? 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086016000173.

Dadic:1984:SMA

[Dad84] Zarko Dadić. Some methodological aspects of Getaldić’s mathematical works. Historia Mathematica, 11(2):207–214, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086084800099.

Dale:1985:SSE

[Dal85] A. I. Dale. A study of some early investigations into exchangeability. Historia Mathematica, 12(4):323–336, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086085900424.

Dale:1988:BTI

[Dal88] A. I. Dale. On Bayes’ theorem and the inverse Bernoulli theorem. Historia Mathematica, 15(4):348–360, November 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900262.

Dale:1991:TBW

[Dal91] A. I. Dale. Thomas Bayes’s work on infinite series. Historia Mathematica, 18(4):312–327, November 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086091903699.

Dale:1992:ACC

[Dal92] A. I. Dale. On the authorship of “A Calculation of the Credibility of Human Testimony”. Historia Mathematica, 19(4):414–417, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290006W.

Darrigol:2015:MRC

[Dar15] Olivier Darrigol. The mystery of Riemann’s curvature. Historia Mathematica, 42(1):47–83, February 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086014000299.

Daston:1979:DCP

[Das79] L. J. Daston. D’Alembert’s critique of probability theory. Historia Mathematica, 6(3):259–279, August 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079901253.

Daston:1980:PER

[Das80] Lorraine J. Daston. Probabilistic expectation and rationality in classical probability theory. Historia Mathematica, 7(3):234–260, August 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900257.

Daston:1981:BRJ

[Das81] Lorraine J. Daston. Book review: I. J. Bienayme. Statistical theory anticipated: By C. C. Heyde and E. Seneta. Studies in the History of Mathematics and the Physical Sciences (3). Berlin (Springer-Verlag). 1977. ix + 172 pp. $20.80. Historia Mathematica, 8(4):487–488, November 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086081900616.

Daston:1994:HPC

[Das94] Lorraine Daston. How probabilities came to be objective and subjective. Historia Mathematica, 21(3):330–344, August 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086084710287.

Dauben:1975:IDP

[Dau75] Joseph W. Dauben. The invariance of dimension: Problems in the early development of set theory and topology [1]. Historia Mathematica, 2(3):273–288, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508607590066X.

Dauben:1979:NE

[Dau79] Joseph W. Dauben. Note from the editor. Historia Mathematica, 6(1):3–4, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086079900995.

Dauben:1980:MWW

[Dau80] Joseph W. Dauben. Mathematicians and World War I: the international diplomacy of G. H. Hardy and Gösta Mittag-Leffler as reflected in their personal correspondence. Historia Mathematica, 7(3):261– 288, August 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086080900269.

Dauben:1982:ICH

[Dau82a] Joseph W. Dauben. International commission on the history of mathematics: Report of the Editor of historia Mathematica to the international commission on the history of mathematics. Historia Mathematica, 9(2):211–215, May 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508608290009X.

Dauben:1982:PPM

[Dau82b] Joseph W. Dauben. Peirce’s place in mathematics. Historia Mathematica, 9(3):311–325, August 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086082901252.

Dauben:1983:E

[Dau83] Joseph W. Dauben. Editorial. Historia Mathematica, 10(1):1–2, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086083900290.

Dauben:1984:E

[Dau84] Joseph W. Dauben. Editorial. Historia Mathematica, 11(1):1–2, February 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086084900016.

Dauben:1985:E

[Dau85a] Joseph W. Dauben. Editorial. Historia Mathematica, 12(4):316–320, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086085900400.

Dauben:1985:HMA

[Dau85b] Joseph Warren Dauben. The history of mathematics from antiquity to the present: a selective bibliography, volume 6 / 313 of Bibliographies of the history of science and technology; Garland reference library of the humanities. Garland, New York, NY, USA, 1985. ISBN 0-8240-9284-8. xxxix + 467 pp. LCCN Z6651 .D38 1985.

Dauben:1986:MIC

[Dau86] Joseph W. Dauben. A message from the international commission on the history of mathematics. Historia Mathematica, 13(1):1, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086086902193.

Dauben:1990:NYG

[Dau90] Joseph W. Dauben. New Year’s greetings from the ICHM. Historia Mathematica, 17(1):1–3, February 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609090073M.

Dauben:1992:MRH

[Dau92] Joseph W. Dauben. A message to readers of Historia Mathematica. Historia Mathematica, 19(1):1–3, February 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609290051C.

Dauben:1993:AMI

[Dau93a] Joseph W. Dauben. A 20th anniversary message from the international commission on the history of mathematics. Historia Mathematica, 20(1):1–4, February 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086083710013.

Dauben:1993:BRH

[Dau93b] Joseph W. Dauben. Book review: Histoire des mathématiques chinoises: Jean-Claude Martzloff. Paris (Masson). 1988. xx + 375 pp. Historia Mathematica, 20(4):437–446, November 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086083710359.

Dauben:1994:MRH

[Dau94] Joseph W. Dauben. A message to readers of Historia Mathematica from the Past Chairman of the International Commission on History of Mathematics. Historia Mathematica, 21(1):1–2, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086084710019.

Dauben:1999:HMY

[Dau99] Joseph W. Dauben. Historia Mathematica: 25 years/context and content. Historia Mathematica, 26(1):1–28, February 1, 1999.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086099922279.

Dauben:2012:BRN

[Dau12] Joseph W. Dauben. Book review: Naming Infinity. A True Story of Religious Mysticism and Mathematical Creativity. By Loren Graham and Jean-Michel Kantor. Cambridge (The Belknap Press of Harvard University Press). 2009. Historia Mathematica, 39(3):341–346, August 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086011000723.

David:1975:BRP

[Dav75] F. N. David. Book review: Probability and statistical inference in ancient and medieval Jewish literature: By N. L. Rabinovitch. University of Toronto Press. 1973. 218 p. $12.50. Historia Mathematica, 2(1):101, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086075900610.

Davis:1976:BRD

[Dav76] Chandler Davis. Book review: Development of mathematics in the Moldavian SSR. Short sketch: By V. P. Bychkov and K. S. Sibirskii. Kishinev (Izd. Shtiinstsa). 1974. 47 p. Historia Mathematica, 3(2):235–236, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076900586.

Davitt:1990:BRM

[Dav90] Richard M. Davitt. Book review: Mathematics: People, problems, results: Edited by Douglas M. Campbell and John C. Higgins. Belmont, California (Wadsworth International). 1984. 3 Vols. xxiv + 871 pp. Original price was $39.95 (cloth); $27.95 (paper). Historia Mathematica, 17(3):289–291, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609090019A.

David:1991:A

[Dav91] E. Zitarelli David. Abstracts. Historia Mathematica, 18(1):90–103, February 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086091903655.

Davison:2019:GHH

[Dav19] Brenda Davison. Book review: The G. H. Hardy Reader, Donald J. Albers, Gerald L. Alexanderson, William Dunham. Cambridge University Press, Cambridge (2015), pp. xii + 395, illus. (Softcover), ISBN: 978-1-107-59464-7. Historia Mathematica, 46(??):102–103, February 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086018301320.

Dawson:1985:CGZ

[Daw85] John W. Dawson, Jr. Completing the Gödel–Zermelo correspondence. Historia Mathematica, 12(1):66–70, February 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086085900709.

Dawson:1986:PKG

[Daw86] John W. Dawson, Jr. The papers of Kurt Gödel. Historia Mathematica, 13(3):277, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086086900935.

Dawes:2016:ALD

[Daw16a] Jonathan H. P. Dawes. After 1952: the later development of Alan Turing’s ideas on the mathematics of pattern formation. Historia Mathematica, 43(1):49–64, February 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086015000397.

Dawson:2016:BRD

[Daw16b] John W. Dawson, Jr. Book review: David Hilbert’s Lectures on the Foundations of Arithmetic and Logic, 1917–1933, William Ewald, Wilfried Sieg (Eds.). Springer (2013). Historia Mathematica, 43 (1):105–107, February 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086015000464.

Debuiche:2023:SPL

[DB23] Valérie Debuiche and Mattia Brancato. Scientia Perspectiva. Leibniz and geometric perspective. Historia Mathematica, 63(??):47–69, May 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086023000150.

Robinson:1976:GBS

[dBR76] G. de B. Robinson. Gauss Bicentennial Symposium Ontario Science Centre, Toronto, Ontario June 3, 4, 1977. Historia Mathematica, 3(4):448, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076900756.

Dahan-Dalmedico:1987:MTS

[DD87] Amy Dahan-Dalmédico. Mécanique et théorie des surfaces: les travaux de Sophie Germain. (French) [Mechanics and the theory of surfaces: the works of Sophie Germain]. Historia Mathematica, 14 (4):347–365, November 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086087900668.

DeYoung:1984:ATT

[De 84] Gregg De Young. The Arabic textual traditions of Euclid’s Elements. Historia Mathematica, 11(2):147–160, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086084800063.

DeYoung:1995:EGM

[De 95] Gregg De Young. Euclidean geometry in the mathematical tradition of Islamic India. Historia Mathematica, 22(2):138–153, May 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086085710142.

DeYoung:2005:DAE

[De 05] Gregg De Young. Diagrams in the Arabic Euclidean tradition: a preliminary assessment. Historia Mathematica, 32(2):129–179, May 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086004000424.

DeYoung:2009:DAE

[De 09] Gregg De Young. Diagrams in ancient Egyptian geometry: Survey and assessment. Historia Mathematica, 36(4):321–373, November 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009000081.

DeYoung:2015:THU

[De 15] Gregg De Young. Two hitherto unknown Arabic Euclid manuscripts. Historia Mathematica, 42(2):132–154, May 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086014000858.

DeYoung:2019:CCB

[De 19] Gregg De Young. The commentator’s craft: Building a mathematical commentary from quotations. Historia Mathematica, 47(??):67– 86, ???? 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086018301423.

DeRisi:2020:BRF

[De 20] Vincenzo De Risi. Book review: Fermat et les débuts modernes de la géométrie, Roshdi Rashed. Olms, Hildesheim (2018), viii + 320., ISBN: 978-3-487-15685-9. Historia Mathematica, 50(??):84–87, February 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508601930076X.

dEnfert:2001:EAC

[d’E01] Renaud d’Enfert. Épures d’architecture. de la coupe des pierres à la géométrie descriptive, XVIe–XIXe siècles. Historia Mathematica, 28(2):127–129, May 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086001923043.

Debuiche:2013:PLI

[Deb13] Valérie Debuiche. Perspective in Leibniz’s invention of Characteristica Geometrica : the problem of Desargues’ influence. Historia Mathematica, 40(4):359–385, October 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086013000463.

Decaillot:2010:CFM

[Déc10] Anne-Marie Décaillot. Cantor und die Franzosen: Mathematik, Philosophie und das Unendliche. (German) [Cantor and the French: Mathematics, philosophy and the infinite]. Mathematik im Kontext. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-14869-7, 3-642-14868-9. xiv + 230 pp.

LCCN QA29.C34 D43 2010eb. URL http://site.ebrary.com/id/
10430300.

Decaillot:2014:RME

[Déc14] Anne-Marie Décaillot. Les récréations mathématiques d’Édouard Lucas: quelques éclairages. Historia Mathematica, 41(4):506–517, November 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086014000561.

DelCentina:2002:MAP

[Del02] Andrea Del Centina. The manuscript of Abel’s Parisian memoir found in its entirety. Historia Mathematica, 29(1):65–69, February 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/

pii/S031508600192333X. See correction [Del03a].

DelCentina:2003:CMA

[Del03a] Andrea Del Centina. Corrigendum to: “The manuscript of Abel’s Parisian memoir found in its entirety”: [Historia Mathematica 29 (2002) 65–69]. Historia Mathematica, 30(1):94–95, February 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086002000228. See [Del02].

Delve:2003:CPE

[Del03b] Janet Delve. The College of Preceptors and the Educational Times: Changes for British mathematics education in the mid-Nineteenth Century. Historia Mathematica, 30(2):140–172, May 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086003000260.

DelCentina:2005:LSG

[Del05] Andrea Del Centina. Letters of Sophie Germain preserved in Florence. Historia Mathematica, 32(1):60–75, February 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086003001022.

DelCentina:2006:ASM

[Del06] Andrea Del Centina. Abel’s surviving manuscripts including one recently found in London. Historia Mathematica, 33(2):224–233, May 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005000728.

Demidov:1974:SCH

[Dem74] S. S. Demidov. Summer Conference on the History of Mathematics and Mechanics in the XIX and XX Centuries Tartu, Estonian S.S.R., 3–10 July 1973. Historia Mathematica, 1(1):91–92, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074901839.

Demidov:1986:APY

[Dem86] Sergei Demidov. Adolph Pavlovich Yushkevich: Selected publications (1976–1986). Historia Mathematica, 13(3):221–223, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086086900881.

Despeaux:2007:LMR

[Des07] Sloan Evans Despeaux. Launching mathematical research without a formal mandate: the role of university-affiliated journals in Britain, 1837–1870. Historia Mathematica, 34(1):89–106, February 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086006000218.

Despeaux:2009:BRM

[Des09] Sloan Evans Despeaux. Book review: Mr Hopkins’ Men: Cambridge Reform and British Mathematics in the 19th Century, Alex D. D. Craik. Springer, London (2007), ISBN 978-1-84628-790-9. Historia Mathematica, 36(2):183–185, May 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008000645.

Descotes:2010:UMM

[Des10] Dominique Descotes. An unknown mathematical manuscript by Blaise Pascal. Historia Mathematica, 37(3):503–534, August 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009001323.

DelCentina:2010:GLM

[DF10] Andrea Del Centina and A. Fiocca. Guglielmo Libri matematico e storico della matematica: l’irresistibile ascesa dall’Ateneo pisano all’Institut de France. (Italian) [Guglielmo Libri, mathematician and historian of mathematics: the irresistible rise of the University of Pisa at the Institute of France], volume 47; 2 of Cultura e memoria / Provincia di Firenze; Fondi scientifici della Biblioteca Moreniana. L. S. Olschki, Firenze, Italia, 2010. ISBN 88-222-6003-1. xxi + 320 + xxii pp. LCCN QA29.L589 D45 2010.

DelCentina:2018:BGP

[DF18] Andrea Del Centina and Alessandra Fiocca. Boscovich’s geometrical principle of continuity, and the “mysteries of the infinity”. Historia Mathematica, 45(2):131–175, May 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0315086018300363.

DelCentina:2021:CTR

[DF21] Andrea Del Centina and Alessandra Fiocca. The chords theorem recalled to life at the turn of the eighteenth century. Historia Mathematica, 56(??):6–39, August 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086021000203.

Dhombres:1990:CRT

[DG90] Jean Dhombres and Enrico Giusti. Colloque: Ratio-la théorie des proportions de l’Antiquité au XIXème siècle: Centro internazionale per la ricerca matematica, trento, 9–13 janvier 1989. Historia Mathematica, 17(1):73–75, February 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609090086S.

Dunnington:2004:CFG

[DGD04] Guy Waldo Dunnington, Jeremy Gray, and Fritz-Egbert Dohse. Carl Friedrich Gauss: titan of science. Mathematical Association of America, Washington, DC, USA, 2004. ISBN 0-88385-547-X. xxix + 537 + 16 pp. LCCN QA29.G3 D82 2004. URL http://www.loc.

gov/catdir/enhancements/fy0632/2003113540-d.html;   http:
//www.loc.gov/catdir/toc/cam051/2003113540.html.

Dauben:2015:IGG

[DGL^{+}15] Joseph W. Dauben, Niccolò Guicciardini, Albert C. Lewis, Karen Hunger Parshall, and Adrian C. Rice. Ivor Grattan-Guinness (June 23, 1941–December 12, 2014). Historia Mathematica, 42(4): 385–406, November 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086015000750.

Dhombres:1983:FMS

[Dho83] Jean Dhombres. French Mathematical Society: a week on the history of mathematics (Marseilles, CIRM, 6–10 June, 1983). Historia Mathematica, 10(4):459, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086083900095.

Dhombres:1985:BRO

[Dho85] Jean Dhombres. Book review: The origins of Cauchy’s rigorous calculus: By Judith V. Grabiner. Cambridge, Mass. (MIT Press). 1981. Historia Mathematica, 12(1):86–90, February 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086085900783.

Dhombres:1987:AAV

[Dho87] J. Dhombres. Ars analytica (from Viète (1591) to Cauchy (1821)). Historia Mathematica, 14(3):289, August 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086087900528.

Dhombres:1988:LEB

[Dho88a] Jean Dhombres. Leonhard Euler 1707–1783: Beiträge zu leben und werk: Gedenkband des Kantons Basel-Stadt. Basel/Boston/Stuttgart (Birkhäuser Verlag). 1983. 555 pp. Historia Mathematica, 15(3):289– 290, August 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508608890078X.

Dhombres:1988:FST

[Dho88b] Jean Dhombres. La France scientifique et technique: Les années 1816– 1824: Journée de la société française d’histoire des sciences et des techniques, Conservatoire des arts et métiers, Paris, 7 Mai 1988. Historia Mathematica, 15(4):375, November 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900328.

Dhombres:1990:CMM

[Dho90] Jean Dhombres. Cauchy et le monde mathématique français: Paris, 27–29 juillet 1989. (French) [Cauchy and the French mathematical world: 27–29 July 1989]. Historia Mathematica, 17(1):70–72, February 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609090084Q.

Dhombres:1991:PEE

[Dho91] Jean Dhombres. Première École d’été d’Histoire des Mathématiques: Herdade da Mitra, Evora, Portugal, 16–21 Juillet 1990 L’histoire des mathématiques: La méthode des textes. Historia Mathematica, 18(3):279–281, August 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609190400R.

Dhombres:2011:BRR

[Dho11] Jean Dhombres. Book review: Le retour du réfoulé. Der Wiederaufstieg der synthetischen Methode an der École Polytechnique. By Gert Schubring. Augsburg (Dr Erwin Rauner Verlag). 2004. ISBN 3-936905-07-X. [10], 150 pp. No price given. Historia Mathematica, 38(4):563–564, November 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086011000255.

Diaz:1975:HMP

[Dia75] Joaquin B. Diaz. How mathematics progresses. Historia Mathematica, 2(4):601–602, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075901287.

Diacu:1999:BRP

[Dia99] Florin Diacu. Book review: Poincaré and the Three-Body Problem. By June Barrow-Green. Historia Mathematica, 26(2):175–178, May 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508609992236X.

Dieudonne:1975:IRA

[Die75] Jean Dieudonné. Introductory remarks on algebra, topology and analysis. Historia Mathematica, 2(4):537–548, November 1975.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075901184.

Dijksterhuis:2004:LWC

[Dij04] Fokko Jan Dijksterhuis. Lenses and waves: Christiaan Huygens and the mathematical science of optics in the Seventeenth Century, volume 9 of Archimedes. Kluwer Academic Press, Dordrecht, The Netherlands; Boston, MA, USA, 2004. ISBN 1-4020-2697-8 (hardcover). vii + 287 pp. LCCN QC352 .D555 2004. URL http://

www.loc.gov/catdir/enhancements/fy0814/2006284750-d.html;
 http://www.loc.gov/catdir/toc/fy0702/2006284750.html.

Dillon:1981:BRP

[Dil81] John Dillon. Book review: Die Pythagoreer: Religiöse Bruderschaft und Schule der Wissenschaft: By B. L. van der Waerden. Zurich and Munich (Artemis-Verlag). 1979. 505 pp. Historia Mathematica, 8(3):377–378, August 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508608190080X.

Djebbar:1997:RLA

[Dje97] Ahmed Djebbar. La rédaction de L’istikmal d’al-Mu’taman (XI^{e} s.) par Ibn Sartaq, un mathématicien des XIII^{e}–XIV^{e} siècles. Historia Mathematica, 24(2):185–192, May 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086096921357.

Djebbar:1998:RAH

[Dje98] Ahmed Djebbar. Raf‘ al-h·ijāb ‘an wujūh a‘māl al-h·isāb li Ibn al-Bannā al-Murrākušī (†721/1321). Historia Mathematica, 25(1):86–88, February 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086097921881.

Demidov:1986:TSC

[DK86] S. S. Demidov and S. N. Kiro. The Third Summer Conference on the History of Mathematics, Odessa, USSR. Historia Mathematica, 13(2):177–179, May 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086086900388.

deLorenzo:2001:QSP

[dL01] Javier de Lorenzo. ‘?Qué son y para qué sirven los números? y otros escritos sobre los fundamentos de la Matemática. Historia Mathematica, 28(2):139–140, May 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086001923134.

Dobrovolskii:1998:MNL

[DLS98] Viacheslav Alekseevich Dobrovolśkii, Natalia Vasilevna Lokot’, and Jean-Marie Strelcyn. Mikhail Nikolaevich Lagutinskii (1871–1915): Un mathématicien méconnu. Historia Mathematica, 25(3):245–264, August 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086098921942.

Charles:1986:ICS

[dMC86a] Maria Sol de Mora Charles. III Congrès de la societé Espagnole d’histoire des sciences: San Sebastián, Pays Basque, 1–6 Octobre, 1984. Historia Mathematica, 13(2):182–184, May 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086086900406.

deMora-Charles:1986:LPP

[dMC86b] Maria Sol de Mora-Charles. Leibniz et le problème des partis. quelques papiers inédits. Historia Mathematica, 13(4):352–369, November 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086086900571.

deMora-Charles:1992:QJH

[dMC92] S. de Mora-Charles. Quelques jeux de hazard selon Leibniz: (manuscrits inédits). (French) [Some games of chance according to Leibniz: unpublished manuscripts]. Historia Mathematica, 19(2):125– 157, May 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609290073K.

Despeaux:2007:Aa

[DMP07a] Sloan Evans Despeaux, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 34(1):116–133, February 2007.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086006001315.

Despeaux:2007:Ab

[DMP07b] Sloan Evans Despeaux, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 34(2):241–258, May 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086007000079.

Despeaux:2007:Ac

[DMP07c] Sloan Evans Despeaux, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 34(3):359–374, August 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086007000353.

Despeaux:2007:Ad

[DMP07d] Sloan Evans Despeaux, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 34(4):455–473, November 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086007000614.

Despeaux:2008:Aa

[DMP08a] Sloan Evans Despeaux, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 35(1):61–79, February 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086007001036.

Despeaux:2008:Ab

[DMP08b] Sloan Evans Despeaux, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 35(2):150–168, May 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008000098.

Despeaux:2008:Ac

[DMP08c] Sloan Evans Despeaux, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 35(3):256–272, August 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008000347.

Despeaux:2008:Ad

[DMP08d] Sloan Evans Despeaux, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 35(4):342–358, November 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008000578.

DAlessandro:2012:ALI

[DN12] Paolo D’Alessandro and Pier Daniele Napolitani. Archimede latino: Iacopo da San Cassiano e il Corpus Archimedeo alla metà del Quattrocento; con edizione della “Circuli dimensio” e della “Quadratura parabolae”, volume 1 of Sciences et savoirs. Les Belles Lettres, Paris, France, 2012. ISBN 2-251-22001-1. 371 + xxviii pp. LCCN QA29.A7 .A73 2012.

Dodgson:2013:EHM

[Dod13] Charles L. Dodgson. Euclid and His Modern Rivals. Cambridge University Press, Cambridge, UK; New York, NY, USA, 2013. ISBN 1-108-00100-9. 336 (est.) pp. LCCN ???? URL http://www.

cambridge.org/us/knowledge/isbn/item2708824/Euclid%20and%
20His%20Modern%20Rivals.

Domingues:2004:VLI

[Dom04] João Caramalho Domingues. Variables, limits, and infinitesimals in Portugal in the late 18th century. Historia Mathematica, 31(1):15–33, February 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086003000041.

Domingues:2008:LC

[Dom08] João Caramalho Domingues. Lacroix and the calculus, volume 35 of Science networks historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2008. ISBN 3-7643-8637-1. xiv + 468 pp. LCCN QA303.2 .D66 2008. URL http://www.loc.

gov/catdir/toc/fy1001/2008920493.html.

Donoghue:1998:SMT

[Don98] Eileen F. Donoghue. In search of mathematical treasures: David Eugene Smith and George Arthur Plimpton. Historia Mathematica, 25(4):359–365, November 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086098922030.

Donahue:2005:BRJ

[Don05] William H. Donahue. Book review: Johannes Kepler Gesammelte Werke, vol. 21.1, Manuscripta Astronomica III; De Calendario Gregoriano; Manuscripta Mathematica: Edited by Volker Bialas, Friederike Boockmann, and Eberhard Knobloch. Munich (Verlag C. H. Beck). 2002. ISBN 3-406-47427-6. 699 pp. EUR 102. Historia Mathematica, 32(2):237–238, May 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086004000023.

Donahue:2008:BRM

[Don08] W. H. Donahue. Book review: Miscellanea Kepleriana. Festschrift für Volker Bialas zum 65. Geburtstag, F. Boockmann, D. A. Di Liscia, H. Kothmann (Eds.). Dr. Erwin Rauner Verlag, Augsburg (2005). Historia Mathematica, 35(4):340–341, November 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008000086.

Dorier:1995:GOG

[Dor95] Jean-Luc Dorier. A general outline of the genesis of vector space theory. Historia Mathematica, 22(3):227–261, ???? 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086085710245.

Dhombres:1988:ERP

[DP88] J. Dhombres and M. Pensivy. Esprit de rigueur et présentation mathématique au XVIII^{ème} siècle: le cas d’une démonstration d’Aepinus. Historia Mathematica, 15(1):9–31, February 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900468.

Demidov:2002:IGB

[DPS^{+}02] Sergey S. Demidov, Alexei N. Parshin, Igor R. Shafarevich, Svetlana S. Petrova, Galina S. Smirnova, Vladimir M. Tikhomirov, and Ioannis M. Vandoulakis. Isabella Grigoryevna Bashmakova on the eightieth anniversary of her birth. Historia Mathematica, 29(4): 370–382, November 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086002923633.

Demidov:1991:IGB

[DPY91] S. S. Demidov, S. S. Petrova, and A. P. Yushkevich. Isabella Grigoryevna Bashmakova on the 70th anniversary of her birth. Historia Mathematica, 18(3):209–211, August 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609190390J.

Dauben:1978:KM

[DR78] Joseph W. Dauben and Laura Roebuck. Kenneth O. May, 1915–1977. Historia Mathematica, 5(1):1–2, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078901313.

Durand-Richard:2003:BRN

[DR03] Marie-José Durand-Richard. Book review: Le nombre, une hydre à n visages. Entre nombres complexes et vecteurs: Edited by Dominique Flament. Paris (Editions de la Maison des sciences de l’homme). 1997. 301 pp. 190 FF. Historia Mathematica, 30(2):217–220, May 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086002000290.

Despeaux:2016:AMA

[DR16] Sloan Evans Despeaux and Adrian C. Rice. Augustus De Morgan’s anonymous reviews for The Athenæum: a mirror of a Victorian mathematician. Historia Mathematica, 43(2):148–171, ???? 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086015000749.

Durand-Richard:2019:IRL

[DR19] Marie-José Durand-Richard. Book review: Iris Runge: a Life at the Crossroads of Mathematics, Science, and Industry, Renate Tobies. Springer/Birkhäuser (2012), 442 pp. + 16 Plates. 139.99$, ISBN: 978-3-0348-0251-2. Historia Mathematica, 46(??):99–102, February 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086018301101.

Drake:1974:MDG

[Dra74] Stillman Drake. Mathematics and discovery in Galileo’s physics. Historia Mathematica, 1(2):129–150, May 1974. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074900020.

Drake:1978:MTK

[Dra78] Stillman Drake. A memorial tribute to Kenneth O. May. Historia Mathematica, 5(1):3–12, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078901325.

Drake:1984:BRN

[Dra84] Stillman Drake. Book review: Nicolas Chuquet: La géométrie. Premiere géométrie algébraique en langue française (1484): Introduction, text, and notes by Hervé l’Huillier. Paris (J. Vrin). 491 pp. fr. 75.90. Historia Mathematica, 11(3):345–346, August 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086084900466.

Drucker:1985:SHL

[Dru85] Thomas L. Drucker. Session on the History of Logic 817th Meeting of the American Mathematical Society: Chicago, Illinois, 22– 23 March 1985. Historia Mathematica, 12(4):378–382, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086085900515.

Drucker:1986:FDP

[Dru86] Thomas Drucker. Frege, Dedekind, and Peano on the foundations of arithmetic. Methodology and science foundation. Historia Mathematica, 13(1):77–78, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086086902387.

Drucker:1991:BRZ

[Dru91] Thomas Drucker. Book review: Zermelo’s axiom of choice: Its origins, development and influence: By Gregory H. Moore. New York: Springer-Verlag. 1982. xiv + 410 pp. Historia Mathematica, 18 (4):364–369, November 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609190378B.

Drucker:2007:BRP

[Dru07] Thomas Drucker. Book review: The Political Pamphlets and Letters of Charles Lutwidge Dodgson and Related Pieces: a Mathematical Approach, Francine F. Abeles (Ed.). Lewis Carroll Society of North America, New York (2001), ISBN 0-930326-14-8. Historia Mathematica, 34(2):231–233, May 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086006000942.

Drucker:2012:BRL

[Dru12] Thomas Drucker. Book review: The Logic Pamphlets of Charles Lutwidge Dodgson and Related Pieces. Edited by Francine F. Abeles. New York (Lewis Carroll Society of North America). 2010. ISBN 978-0-930326-25-8. xx + 271 pp. Historia Mathematica, 39(3):338–341, August 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086012000109.

Dry:2014:NPS

[Dry14] Sarah Dry. The Newton Papers: the Strange and True Odyssey of Isaac Newton’s Manuscripts. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-995104-7 (hardcover). xi + 238 pp. LCCN QC16.N7 D79 2014.

daSilva:1990:SHD

[dS90] Clóvis Pereira da Silva. A social history of the development of university mathematics in Brazil from 1810 to 1920. Historia Mathematica, 17(3):271, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086090900145.

Dold-Samplonius:1997:MBL

[DS97] Yvonne Dold-Samplonius. In memoriam: Bartel Leendert van der Waerden (1903–1996). Historia Mathematica, 24(2):125–130, May 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086097921522.

daSilva:1999:IPT

[dS99a] Circe Mary Silva da Silva. The influence of positivism on the teaching of mathematics in Brazil: 1870–1930. Historia Mathematica, 26 (4):327–343, November 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086099922516.

Dold-Samplonius:1999:BRA

[DS99b] Yvonne Dold-Samplonius. Book review: Al-Sijzi’s Treatise on Geometrical Problem Solving. Translated and annotated by Jan P. Hogendijk. Historia Mathematica, 26(4):382–383, November 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086099922528.

Dold-Samplonius:1999:BRT

[DS99c] Yvonne Dold-Samplonius. Book review: The Topkapi Scroll — Geometry and Ornament in Islamic Architecture: Topkapi Palace Library MS. H. 1956; With an Essay on the Geometry of the Muqarnas by Mohammad al-Asad. By Gülru Necipoğlu. Historia Mathematica, 26(2):166–172, May 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086099922346.

Dold-Samplonius:2002:CPY

[DS^{+}02] Yvonne Dold-Samplonius et al., editors. From China to Paris: 2000 years transmission of mathematical ideas, volume 46 of Boethius. Franz Steiner Verlag, Stuttgart, Germany, 2002. ISBN 3-515-08223-9. LCCN QA21 .F765 2002.

Durnova:2019:VHI

[DS19] Helena Durnová and Tilman Sauer. Václav Hlavatý on intuition in Riemannian space. Historia Mathematica, 49(??):60–79, November 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086018301587.

deSaint-Gervais:2010:USR

[dSG10] Henri Paul de Saint-Gervais. Uniformisation des surfaces de Riemann: retour sur un théorème centenaire. ENS éditions, Lyon, France, 2010. ISBN 2-84788-233-2. 544 pp. LCCN QA333 .S22 2010.

Dubbey:1977:BPM

[Dub77] J. M. Dubbey. Babbage, Peacock and modern algebra. Historia Mathematica, 4(3):295–302, August 1977. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508607790057X.

Ducheyne:2016:BRN

[Duc16] Steffen Ducheyne. Book review: The Newton Papers. The Strange and True Odyssey of Isaac Newton’s Manuscripts, Sarah Dry. Oxford University Press, New York (2014), ISBN: 978-0-19-995104-8. Historia Mathematica, 43(3):342–345, August 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086015000762.

Dugac:1976:PHA

[Dug76] Pierre Dugac. Problèmes de l’histoire de l’analyse mathématique au XIXème siècle. Cas de Karl Weierstrass et de Richard Dedekind. Historia Mathematica, 3(1):5–19, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086076900021.

Dugac:1982:HPD

[Dug82] Pierre Dugac. Hommage à Poisson du congrés de Bucarest et de l’École Polytechnique de Paris. Historia Mathematica, 9(2):221– 222, May 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086082900118.

Dumbaugh:2015:BRE

[Dum15] Della Dumbaugh. Book review: Emil Artin’s Iceland Journal 1925, Tom Artin (Ed.). Free Scholar Press (2013), ISBN: 978-1-4825-5168-6. Historia Mathematica, 42(1):105–107, February 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086014000317.

Dumbaugh:2018:BRH

[Dum18] Della Dumbaugh. Book review: A History in Sum: 150 Years of Mathematics at Harvard, 1825–1875, Steve Nadis, Shing-Tung Yang (Eds.). Harvard University Press, Cambridge (2013), 249 pp. Historia Mathematica, 45(1):85–90, February 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086017300721.

Dun:1989:ISC

[Dun89] Liu Dun. International symposium for the commemoration of mei wending and the Chinese third annual meeting on the history of mathematics. Historia Mathematica, 16(3):281–282, August 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086089900244.

Durand:2021:LSM

[Dur21] Antonin Durand. Book review: L’élite sous la mitraille. Les normaliens, les mathématiques et la Grande Guerre, 1900–1925, David Aubin. Éditions Rue d’Ulm, Paris (2018), coll. <<Figures normaliennes>>, 360 pp., preface by Claude Viterbo, 25 EUR. Historia Mathematica, 55(??):95–97, May 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086021000215.

Dutka:1984:RER

[Dut84] Jacques Dutka. Richardson extrapolation and Romberg integration. Historia Mathematica, 11(1):3–21, February 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086084900028.

Demidov:1980:SSC

[DV80] S. S. Demidov and A. I. Volodarsky. Second Summer Conference on the History of Mathematics in the Soviet Union. Historia Mathematica, 7(1):71–75, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508608090066X.

Demidov:2007:IGB

[DV07] Sergei S. Demidov and Ioannis M. Vandoulakis. Isabella Grigoryevna Bashmakova (1921–2005). Historia Mathematica, 34(2):137–139, May 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086006001248.

Devreese:2008:MNM

[DV08] J. T. (Jozef T.) Devreese and Guido Vanden Berghe. ’Magic is no magic’: the wonderful world of Simon Stevin. WIT, Southampton, UK, 2008. ISBN 1-84564-092-6 (hardcover). xxxvi + 310 pp. LCCN QA29.S68 D462 2008.

deVarent:2022:SNV

[dV22] Charlotte de Varent. Small numerical variations in a set of similar problems from Nippur on the area of the square. Historia Mathematica, 58(??):35–70, February 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086021000653.

deWreede:2010:DUA

[dW10] Liesbeth C. de Wreede. A dialogue on the use of arithmetic in geometry: Van Ceulen’s and Snellius’s Fundamenta Arithmetica et Geometrica. Historia Mathematica, 37(3):376–402, August 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086010000236.

deYoung:1992:IIH

[dY92] Gregg de Young. Isḣāq ibn Hunayn, Hunayn ibn Isḣāq, and the third Arabic translation of Euclid’s Elements. Historia Mathematica, 19(2):188–199, May 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609290076N.

Dyck:1975:BRE

[Dyc75a] Martin Dyck. Book review: Einstein. Zhizn, smert, bessmertie: [Einstein: Life, Death, Immortality]. By B. G. Kuznetsov. Moskva (Izdatelstvo “Nauka”), 1972. 607 pp. Historia Mathematica, 2(3):344–347, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086075900828.

Dyck:1975:BRG

[Dyc75b] Martin Dyck. Book review: Goethe kak estestvoispytatel: [Goethe as a Natural Scientist]. By I. I. Kanaev. Leningrad (Izdatelstvo “Nauka”). 1970. 466 pp. Historia Mathematica, 2(1):92–95, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900579.

De:1997:RCM

[DZ97] Mo De and Jiang Zhenhua. The recent Chinese and Mongolian translations of Euclid’s Elements. Historia Mathematica, 24(2):197– 199, May 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086096921497.

Ebbinghaus:2004:ZMB

[Ebb04] Heinz-Dieter Ebbinghaus. Zermelo in the mirror of the Baer correspondence, 1930–1931. Historia Mathematica, 31(1):76–86, February 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086003000788.

Ebbinghaus:2007:ZHC

[Ebb07] Heinz-Dieter Ebbinghaus. Zermelo and the Heidelberg Congress 1904. Historia Mathematica, 34(4):428–432, November 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086006001236.

Eberhard:1996:BRM

[Ebe96] Andrea Eberhard. Book review: Les mathématiques japonaises à l’époque d’Edo (1600–1868): Une étude des travaux de Seki Takakazu (?–1708) et de Takebe Katahiro (1664–1793) By Annick Horiuchi. Paris (Librairie Philosophique J. Vrin). Historia Mathematica, 23 (3):327–329, August 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086096900324.

Eckert:2013:ASA

[Eck13] Michael Eckert. Arnold Sommerfeld: Atomphysiker und Kulturbote, 1868–1951: eine Biografie. (German) [Arnold Sommerfeld: nuclear physicist and messenger of culture, 1868–1951: a biography], volume 29 of Abhandlungen und Berichte (Deutsches Museum (Germany)). Wallstein, Göttingen, Germany, 2013. ISBN 3-8353-1206-5 (hardcover). 604 pp. LCCN QC16.S76 E25 2013.

Edson:1984:CA

[Eds84] W. A. Edson. Chebyshev’s approximations. Historia Mathematica, 11 (2):233, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086084800221.

Edwards:1977:BRC

[Edw77a] Harold M. Edwards. Book review: Courant in Göttingen and New York: By Constance Reid. New York, Heidelberg, Berlin (Springer-Verlag). 1976. 314 pp. Historia Mathematica, 4(4):479–480, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900982.

Edwards:1977:BRE

[Edw77b] Harold M. Edwards. Book review: Ernst Eduard Kummer. Collected papers: Volume I. Contributions to Number Theory. Edited by André Weil. Berlin-Heidelberg-New York (Springer-Verlag). 1975. 957 p. Historia Mathematica, 4(4):475–478, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900970.

Edwards:1978:KNL

[Edw78] Harold M. Edwards. On the Kronecker Nachlass [legacy]. Historia Mathematica, 5(4):419–426, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078902100.

Edwards:1980:BRV

[Edw80] Harols M. Edwards. Book review: Variationen über ein zahlentheoretisches Thema von C. F. Gauss: By H. Pieper. Basel and Stuttgart (Birkhäuser). 1978. Historia Mathematica, 7(1):95–96, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608090083X.

Edwards:1991:NWR

[Edw91] A. W. F. Edwards. A note on Waring’s Rule. Historia Mathematica, 18(2):177, May 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609190499N.

Edwards:1992:CCB

[Edw92] A. W. F. Edwards. Celebration of the centenary of the birth of Sir Ronald Aylmer Fisher, 1990. Historia Mathematica, 19(1):81–82, February 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609290061F.

Edwards:2004:CMS

[Edw04] A. W. F. (Anthony William Fairbank) Edwards. Cogwheels of the mind: the story of Venn diagrams. The Johns Hopkins University Press, Baltimore, MD, USA, 2004. ISBN 0-8018-7434-3 (hardcover). xvi + 110 pp. LCCN QA248 .E28 2004. URL http://www.

e-streams.com/es0801/es0801\_3843.html;   http://www.loc.
gov/catdir/bios/jhu051/2003010633.html;   http://www.loc.
gov/catdir/description/jhu051/2003010633.html;   http://
www.loc.gov/catdir/toc/ecip043/2003010633.html.

Enea:2020:AAM

[EF20] Maria Rosaria Enea and Giovanni Ferraro. Analytic and arithmetic methods in Liouville’s identities. Historia Mathematica, 53(??):48–70, November 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508602030046X.

DeGandt:1988:CAC

[eFD88] Organisépar Amy Dahan et François De Gandt. Colloque-atelier (CIRM Luminy 6–10 Juillet 1987): Physique et analyse de Newton à Gauss. Historia Mathematica, 15(4):371–372, November 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900304.

Engelberg:1981:MNM

[EG81] Don Engelberg and Michael Gertner. A marginal note of Mersenne concerning the “Galileian spiral”. Historia Mathematica, 8(1):1–14, February 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508608190001X.

Ehrhardt:2011:QBJ

[Ehr11] Caroline Ehrhardt. A quarrel between Joseph Liouville and Guillaume Libri at the French Academy of Sciences in the middle of the Nineteenth Century. Historia Mathematica, 38(3):389–414, August 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508601100005X.

Ehrhardt:2015:TSL

[Ehr15] Caroline Ehrhardt. Tactics: In search of a long-term mathematical project (1844–1896). Historia Mathematica, 42(4):436–467, November 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086014001098.

Ehrhardt:2018:LTE

[Ehr18] Caroline Ehrhardt. A locus for transnational exchanges: European mathematical journals for students and teachers, 1860s–1914. Historia Mathematica, 45(4):376–394, November 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086018300910.

Eden:2012:GME

[EI12] Alp Eden and Gürol Irzik. German mathematicians in exile in Turkey: Richard von Mises, William Prager, Hilda Geiringer, and their impact on Turkish mathematics. Historia Mathematica, 39(4):432–459, November 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086012000377.

Eisenhart:1975:BRS

[Eis75] Churchill Eisenhart. Book review: Statistical papers in honor of George W. Snedecor: Edited by T. A. Bancroft assisted by Susan Alice Brown. Iowa State University Press, 1972. 345 p. US $10.50. Historia Mathematica, 2(2):211–218, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075901561.

Eisele:1978:CCV

[Eis78] Carolyn Eisele. A comment on C. V. Jones’ review of the New Elements. Historia Mathematica, 5(1):91–92, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078901404.

Eisenberg:1980:AAH

[Eis80] Sheldon M. Eisenberg. Another approach to a history of mathematics course for undergraduates. Historia Mathematica, 7(1):82–83, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086080900701.

Eisele:1982:MMT

[Eis82] Carolyn Eisele. Mathematical methodology in the thought of Charles S. Peirce. Historia Mathematica, 9(3):333–341, August 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082901276.

Elior:2018:ATE

[Eli18] Ofer Elior. The Arabic tradition of Euclid’s Elements preserved in the Latin translation by Adelard of Bath and the Hebrew translation by Rabbi Jacob. Historia Mathematica, 45(2):111– 130, May 2018. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL https://www.sciencedirect.com/
science/article/pii/S031508601830003X.

Elior:2024:HTE

[Eli24] Ofer Elior. The Hebrew translation of Euclid’s Elements ascribed to Rabbi Jacob: a new analysis following the “Discovery” of the Arabic version in MS Paris, BULAC ARA. 606. Historia Mathematica, 68(??):1–21, September 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086024000090.

Enea:2018:CEM

[Ene18] Maria Rosaria Enea. Circulation of an editorial model: the casestudy of the short-lived Le Matematiche Pure ed Applicate. Historia Mathematica, 45(4):395–413, November 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086018300909.

Engels:1977:QCA

[Eng77] Hermann Engels. Quadrature of the circle in ancient Egypt. Historia Mathematica, 4(2):137–140, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086077901045.

Engels:1979:GRE

[Eng79] Hermann Engels. Zur Geschichte der Richardson-Extrapolation. (German) [On the history of Richardson extrapolation]. Historia Mathematica, 6(3):280–293, August 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079901265.

Engelsman:1980:LEC

[Eng80] S. B. Engelsman. Lagrange’s early contributions to the theory of firstorder partial differential equations. Historia Mathematica, 7(1):7–23, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086080900609.

Engelsman:1981:TGMa

[Eng81a] S. B. Engelsman. Tagung “Geschichte des Mathemetik” im mathematischen Forschungsinstitut Oberwolfach. (German) [conference on “History of mathematics” in the Mathematical Research Institute in Oberwolfach]. Historia Mathematica, 8(1):67–79, February 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900070.

Engelsman:1981:TGM

[Eng81b] S. B. Engelsman. Tagung “Geschichte des Mathemetik” im mathematischen Forschungsinstitut Oberwolfach. (German) [conference on “History of mathematics” in the Mathematical Research Institute in Oberwolfach]. Historia Mathematica, 8(4):460–472, November 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900550.

Enros:1976:BRR

[Enr76] Philip C. Enros. Book review: Rapport historique sur les progrès des sciences mathématiques depuis 1789, et sur leur état actuel, etc.: By J. B. J. Delambre. Reprint of the original Paris edition of 1810. Amsterdam (B. M. Israël). 1966. viii + 362 pp. Historia Mathematica, 3(3):342–344, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076901154.

Enros:1980:BRM

[Enr80] Philip C. Enros. Book review: The mathematical work of Charles Babbage: By John Michael Dubbey. New York/London (Cambridge University Press). 1978. viii + 235 pp. $26.50. Historia Mathematica, 7(1):98–100, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086080900853.

Enros:1983:ASP

[Enr83] Philip C. Enros. The analytical society (1812–1813): Precursor of the renewal of Cambridge mathematics. Historia Mathematica, 10(1):24– 47, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086083900319.

Enros:1984:KMB

[Enr84] Philip Enros. Kenneth O. May — bibliography. Historia Mathematica, 11(4):380–393, November 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086084900235.

Enros:1985:BRGb

[Enr85a] Philip C. Enros. Book review: Great moments in mathematics (after 1650): By Howard Eves. Dolciani Mathematical Expositions No. 7. Washington, D.C. (The Mathematical Association of America). 1981. xii + 263 pp. $23.50 hardcover; $13 paperback; $22 for the 2-volume set, paperback. Historia Mathematica, 12(4):391–393, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086085900576.

Enros:1985:BRGa

[Enr85b] Philip C. Enros. Book review: Great moments in mathematics (before 1650): By Howard Eves. Dolciani Mathematical Expositions No. 5. Washington, D.C. (The Mathematical Association of America). 1980. xiv + 270 pp. $22 hardcover; $13 paperback. Historia Mathematica, 12(4):391–393, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086085900564.

Epple:1995:BPA

[Epp95] Moritz Epple. Branch points of algebraic functions and the beginnings of modern knot theory. Historia Mathematica, 22(4):371–401, November 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086085710312.

Erlichson:1992:NSE

[Erl92] Herman Erlichson. Newton’s solution to the equiangular spiral problem and a new solution using only the equiangular property. Historia Mathematica, 19(4):402–413, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609290005V.

Erlichson:1994:VQM

[Erl94] Herman Erlichson. The visualization of quadratures in the mystery of corollary 3 to proposition 41 of Newton’s Principia. Historia Mathematica, 21(2):148–161, May 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086084710147.

Erlichson:1997:HSE

[Erl97] Herman Erlichson. Hooke’s September 1685 ellipse vertices construction and Newton’s instantaneous impulse construction. Historia Mathematica, 24(2):167–184, May 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086096921333.

Erlichson:2003:PLP

[Erl03] Herman Erlichson. Passage to the limit in Proposition I, Book I of Newton’s Principia. Historia Mathematica, 30(4):432–440, November 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086002000083.

Ewald:2013:DHL

[ESH^{+}13] William Bragg Ewald, Jr., Wilfried Sieg, Michael Hallett, Ulrich Majer, and Dirk Schlimm, editors. David Hilbert’s lectures on the foundations of arithmetic and logic, 1917–1933, volume 3 of David Hilbert’s lectures on the foundations of mathematics and physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 3-540-20578-0 (hard cover), 3-540-69444-7. xxv + 1062 pp. LCCN ????

Esty:1979:APM

[Est79] Warren W. Esty. Ancient portraits of mathematicians. Historia Mathematica, 6(4):437–441, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079900302.

Esteve:2006:AGP

[Est06] Ma. Rosa Massa Esteve. Algebra and geometry in Pietro Mengoli (1625–1686). Historia Mathematica, 33(1):82–112, February 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005000029.

Euclides:2003:DED

[ET03] Euclides and Christian Marinus Taisbak. Dedomena: Euclid’s data or the importance of being given, volume 45 of Acta historica scientiarum naturalium et medicinalium. Museum Tusculanum Press, Copenhagen, Denmark, 2003. ISBN 87-7289-815-1. 271 pp. LCCN QA31. EUR 45,00. URL http://www.gbv.de/dms/

goettingen/362001758.pdf;   http://www.zentralblatt-math.
org/zmath/en/search/?an=1032.01005.

Eckes:2014:GLO

[ET14] Christophe Eckes and Amaury Thuillier. Les groupes de Lie dans l’oeuvre de Hermann Weyl: traduction et commentaire de l’article Théorie de la représentation des groupes continus semi-simples par des transformations linéaires (1925–1926). Histoires de géométries. Presses universitaires de Nancy, Nancy, France, 2014. ISBN 2-8143-0180-2 (paperback). ISSN 2115-6492. 401 pp. LCCN ???? URL http:

//catalogue.bnf.fr/ark:/12148/cb437736943. English preface by

Erhard Scholz.

Elstrodt:1999:RSC

[EU99] Jürgen Elstrodt and Peter Ullrich. A real sheet of complex Riemannian function theory: a recently discovered sketch in Riemann’s own hand. Historia Mathematica, 26(3):268–288, August 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086099922413.

Evans:1994:BRL

[Eva94] James Evans. Book review: L’Oeuvre de Torricelli: Science galiléenne et nouvelle géométrie. Publications de la Faculté des Lettres et Sciences Humaines de Nice, François De Gandt (Ed.). Les Belles Lettres, Paris (1987), 247. Historia Mathematica, 21(4): 463–464, November 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086084710342.

Evans:1998:VMS

[Eva98] G. R. Evans. Vestigia Mathematica, Studies in Medieval and Early Modern Mathematics in Honour of H. L. L. Busard. Historia Mathematica, 25(1):85–86, February 1998. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086097921686.

Evans:2013:BRM

[Eva13] James Evans. Book review: The Mathematics of the Heavens and the Earth: the Early History of Trigonometry, Glen Van Brummelen. Princeton University Press, Princeton and Oxford (2009), ISBN: 978-0-691-12973-0. Historia Mathematica, 40(3):324–325, August 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086012000791.

Eves:1976:BRH

[Eve76] Howard Eves. Book review: History of mathematics: Topics for schools: By Walter Popp. Translated from the German by Maxim Bruckheimer. London, England (Transworld Student Library). 1974. 150 pp. Historia Mathematica, 3(4):495–496, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900902.

Ewing:1992:BRH

[Ewi92] John Ewing. Book review: The history of modern mathematics: Edited by David E. Rowe and John McCleary. San Diego (Academic Press). 2 vols., xvi + 453 and xvi + 325 pp, $39.50 each. Historia Mathematica, 19(1):93–98, February 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609290068M.

Fukagawa:2022:ASG

[FC22] Hidetoshi Fukagawa and David Clark. Ajima’s solution to the Gion shrine problem: a modern interpretation. Historia Mathematica, 61(??):3–10, November 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S031508602200026X.

Feffer:1997:MPP

[Fef97] Loren Butler Feffer. Mathematical physics and the planning of American mathematics: Ideology and institutions. Historia Mathematica, 24(1):66–85, February 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086097921248.

Feferman:1998:LL

[Fef98] Solomon Feferman. In the light of logic. Logic and computation in philosophy. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1998. ISBN 0-19-508030-0. xii + 340 pp. LCCN QA9.2 .F44 1998. URL http://www.loc.gov/

catdir/enhancements/fy0604/97051336-d.html;   http://www.
loc.gov/catdir/enhancements/fy0604/97051336-t.html.

Feke:2014:MMR

[Fek14] Jacqueline Feke. Meta-mathematical rhetoric: Hero and Ptolemy against the philosophers. Historia Mathematica, 41(3):261–276, August 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086014000287.

Fellmann:1978:BGW

[Fel78] E. A. Fellmann. Über eine Bemerkung von G. W. Leibniz zu einem Theorem in Newtons Principia mathematica. (German) [On a remark of G. W. Leibniz on a theorem in Newton’s Principia mathematica]. Historia Mathematica, 5(3):267–273, August 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508607890112X.

Fellmann:1983:LEM

[Fel83] Emil A. Fellmann. Leonhard Euler, 1707–1783: Memorial volume on the 200th anniversary of his death. Historia Mathematica, 10(2):225– 227, May 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086083900708.

Fellmann:1985:RRC

[Fel85] Emil A. Fellmann. Review of a review: Critical remarks on S. S. Chern’s review of Erhard Scholz’s Geschichte des Mannigfaltigkeitsbegriffs von Riemann bis Poincaré. Historia Mathematica, 12(3):276– 277, August 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086085900278.

Fellmann:2007:LE

[Fel07] Emil Alfred Fellmann. Leonhard Euler. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2007. ISBN 3-7643-7538-8. xv + 179 pp. LCCN QA29.E8 F452 2007. URL http://www.

loc.gov/catdir/toc/fy0709/2006937470.html. Translated by E.

Gautschi and W. Gautschi from the 1995 German original of the same title.

Fenton:1989:EPH

[Fen89] P. C. Fenton. An extremal problem in Harriot’s mathematics. Historia Mathematica, 16(2):154–163, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086089900372.

Fenster:1997:RMM

[Fen97] Della Dumbaugh Fenster. Role modeling in mathematics: the case of Leonard Eugene Dickson (1874–1954). Historia Mathematica, 24(1):7–24, February 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086097921200.

Fenster:2003:FMC

[Fen03] Della Dumbaugh Fenster. Funds for mathematics: Carnegie Institution of Washington support for mathematics from 1902 to 1921. Historia Mathematica, 30(2):195–216, May 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086003000053.

Fenster:2008:BRE

[Fen08] Della D. Fenster. Book review: Exposition by Emil Artin: a Selection, vol. 30, M. Rosen (Ed.), in: History of Mathematics. American Mathematical Society/London Mathematical Society, Providence/London (2006), ISBN 978-0-8218-4172-3. Historia Mathematica, 35(1):53–56, February 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086007000584.

Fenster:2009:PGM

[Fen09] Della D. Fenster. Plato’s Ghost: the modernist transformation of mathematics. Historia Mathematica, 36(4):444–447, November 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009000536.

Ferreiros:1993:RBG

[Fer93] José Ferreirós. On the relations between Georg Cantor and Richard Dedekind. Historia Mathematica, 20(4):343–363, November 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508608371030X.

Ferreiros:1995:WFM

[Fer95] José Ferreirós. “What fermented in me for years”: Cantor’s discovery of transfinite numbers. Historia Mathematica, 22(1):33–42, February 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086085710038.

Fernandez:1998:BRN

[Fer98a] Max Fernández. Book review: La Notion de nombre chez Dedekind, Cantor, Frege. By Jean-Pierre Belna. Historia Mathematica, 25(2): 219–221, May 1998. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S031508609792187X.

Ferraro:1998:SAE

[Fer98b] Giovanni Ferraro. Some aspects of Euler’s theory of series: Inexplicable functions and the Euler–Maclaurin summation formula. Historia Mathematica, 25(3):290–317, August 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086098921954.

Ferraro:2000:FFR

[Fer00] Giovanni Ferraro. Functions, functional relations, and the laws of continuity in Euler. Historia Mathematica, 27(2):107–132, May 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508600092278X.

Ferraro:2004:DDC

[Fer04a] Giovanni Ferraro. Differentials and differential coefficients in the Eulerian foundations of the calculus. Historia Mathematica, 31(1):34–61, February 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086003000302.

Ferreiros:2004:BRF

[Fer04b] José Ferreirós. Book review: From Frege to Gödel. A Source Book in Mathematical Logic, 1879–1931: By Jean van Heijenoort. Cambridge, MA (Harvard University Press). 1967; new paperback edn., 2002. 664 pages, 1 halftone. ISBN 0-674-32449-8. $27.95. Historia Mathematica, 31(1):119–124, February 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086003000582.

Ferraro:2007:CFM

[Fer07] Giovanni Ferraro. Convergence and formal manipulation in the theory of series from 1730 to 1815. Historia Mathematica, 34(1):62–88, February 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086005000777.

Ferraro:2008:RDT

[Fer08] Giovanni Ferraro. The rise and development of the theory of series up to the early 1820s. Sources and studies in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 0-387-73467-8. xv + 389

pp. LCCN QA295 .F474 2008. URL http://www.loc.gov/catdir/
enhancements/fy0826/2007939827-d.html;   http://www.loc.
gov/catdir/enhancements/fy0826/2007939827-t.html.

Ferraro:2011:BRC

[Fer11] Giovanni Ferraro. Book review: La construction tractionnelle des équations différentielles. By D. Tournès. Paris (Albert Blanchard). 2009. ISBN 978-2-85367-247-4. viii, 406 pp. No price given. Historia Mathematica, 38(3):425–428, August 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086011000243.

Ferraro:2020:ESM

[Fer20] Giovanni Ferraro. Euler and the structure of mathematics. Historia Mathematica, 50(??):2–24, February 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086019300400.

Fauvel:2000:OFY

[FFW00] John Fauvel, Raymond Flood, and Robin J. Wilson, editors. Oxford figures: 800 years of the mathematical sciences. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-852309-2. 296 pp. LCCN QA14.G73 O947 2000. URL http://

www.loc.gov/catdir/enhancements/fy0606/99027797-d.html;
http://www.loc.gov/catdir/enhancements/fy0606/99027797-t.
html.

Fauvel:1990:ASC

[FG90] John Fauvel and Paulus Gerdes. African slave and calculating prodigy: Bicentenary of the death of Thomas Fuller. Historia Mathematica, 17(2):141–151, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609090050N.

FerreirosDominguez:2006:AMM

[FG06] José Ferreirós Domínguez and Jeremy Gray, editors. The architecture of modern mathematics: essays in history and philosophy. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-856793-6. xii + 442 pp. LCCN QA21

.A73 2006. (alk. paper). URL http://www.loc.gov/catdir/
enhancements/fy0636/2005036586-d.html;   http://www.loc.
gov/catdir/toc/ecip065/2005036586.html.

Fery:2012:AAF

[FG12] Suzanne Féry and Christian Gilain. Aventures de l’analyse de Fermat à Borel: mélanges en l’honneur de Christian Gilain. Histoires de géométries. Presses Universitaires de Nancy, Éditions Universitaires de Lorraine, Nancy, France, 2012. ISBN 2-8143-0125-X. 728 pp. LCCN ????

Fibonacci:2008:FPG

[FH08] Leonardo Fibonacci and Barnabas Hughes, editors. Fibonacci’s De practica geometrie. Sources and studies in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 0-387-72930-5. xxxv + 408

pp. LCCN QA32 .F4613 2008. URL http://www.loc.gov/catdir/
enhancements/fy0826/2007934985-d.html;   http://www.loc.
gov/catdir/enhancements/fy0826/2007934985-t.html.

Fiette:2022:MCS

[Fie22] Baptiste Fiette. Measuring crops with the sukunnûm-number. Historia Mathematica, 59(??):71–98, May 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086020300483.

Findlen:2011:CFM

[Fin11] Paula Findlen. Calculations of faith: mathematics, philosophy, and sanctity in 18th-century Italy (new work on Maria Gaetana Agnesi). Historia Mathematica, 38(2):248–291, May 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086010000315.

Fiocca:2001:DST

[Fio01] Alessandra Fiocca. Desargues en son temps. (French) [Desargues in his time]. Historia Mathematica, 28(4):312–314, November 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086001923250.

Fiocca:2017:BBS

[Fio17] Alessandra Fiocca. The Bullettino di Bibliografia e di Storia delle Scienze Matematiche e Fisiche (1868–1887), an example of the internationalisation of research. Historia Mathematica, 44(1):1–30, February 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086016300519.

Fisher:1978:CIS

[Fis78] Gordon M. Fisher. Cauchy and the infinitely small. Historia Mathematica, 5(3):313–331, August 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078901179.

Fischler:1979:REI

[Fis79] Roger Fischler. A remark on Euclid II, 11. Historia Mathematica, 6(4):418–422, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086079900272.

Fisch:1982:PPA

[Fis82] Max H. Fisch. Peirce’s place in American life. Historia Mathematica, 9(3):265–289, August 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086082901227.

Fischer:1994:DCM

[Fis94] Hans Fischer. Dirichlet’s contributions to mathematical probability theory. Historia Mathematica, 21(1):39–63, February 1994.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508608471007X.

Fischer:2011:HCL

[Fis11] Hans Fischer. A history of the Central Limit Theorem: from classical to modern probability theory. Sources and studies in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 0-387-87856-4 (hardcover). xvi + 402 pp. LCCN QA273.67 .F57 2011.

Folkerts:1989:CJS

[FK89] Menso Folkerts and Eberhard Knobloch. Christoph J. Scriba — 60 Jahre. Historia Mathematica, 16(3):207–212, August 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086089900177.

Farwell:1990:MLR

[FK90] Ruth Farwell and Christopher Knee. The missing link: Riemann’s “Commentatio,” differential geometry and tensor analysis. Historia Mathematica, 17(3):223–255, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609090002U.

Farghaanai:2005:A

[FL05] Farghāanāi and Richard Lorch, editors. On the astrolabe, volume 52 of Boethius. Steiner, Stuttgart, Germany, 2005. ISBN 3-515-08713-3 (hardcover). 447 pp. LCCN QB85 .F37 2005.

Flament:2003:HNC

[Fla03] D. (Dominique) Flament. Histoire des nombres complexes: entre algèbre et géométrie. (French) [History of complex numbers: between algebra and geometry]. CNRS histoire des sciences. CNRS, Paris, France, 2003. ISBN 2-271-06128-8. 501 pp. LCCN QA255 .F53 2003.

Flegg:1975:HMC

[Fle75] Graham Flegg. History of mathematics course (AM289) at the open university, great Britain. Historia Mathematica, 2(3):332–333, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086075900774.

Fletcher:1986:AFM

[Fle86a] Colin R. Fletcher. Ampère, Faraday, and Maxwell: Three founders of electromagnetism. Historia Mathematica, 13(3):296, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086086901011.

Fletcher:1986:RMG

[Fle86b] Colin R. Fletcher. Refugee mathematicians: a German crisis and a British response, 1933–1936. Historia Mathematica, 13(1):13–27, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086086902223.

Fletcher:1989:FT

[Fle89] Colin R. Fletcher. Fermat’s theorem. Historia Mathematica, 16(2): 149–153, May 1989. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/0315086089900360.

Fletcher:1991:RFF

[Fle91] Colin R. Fletcher. A reconstruction of the Frenicle–Fermat correspondence of 1640. Historia Mathematica, 18(4):344–351, November 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086091903714.

Folkerts:2016:JBM

[FLT16] Menso Folkerts, Dieter Launert, and Andreas Thom. Jost Bürgi’s method for calculating sines. Historia Mathematica, 43(2):133–147, ???? 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086016000185.

Flament:2011:JM

[FN11] Dominique. Flament and Philippe Nabonnand. Justifier en mathématiques. Éditions de la Maison des Sciences de l’Homme, Paris, France, 2011. ISBN 2-7351-1414-7 (paperback). 371 pp. LCCN ????

Folkerts:1976:RCH

[Fol76] M. Folkerts. Report on the 19th colloquium on the history of mathematics at the Mathematical Research Institute Oberwolfach (Black Forest). Historia Mathematica, 3(1):81–85, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900148.

Folkerts:1978:NAA

[Fol78] M. Folkerts. Der Nachlass Axel Anthon Björnbos. (German) [The legacy of Axel Anthon Björnbos]. Historia Mathematica, 5(3):333– 339, August 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086078901192.

Folkerts:1983:BUA

[Fol83a] Menso Folkerts. Eine bisher unbekannte Abhandlung über das Rechenbrett aus dem Beginnenden 14. Jahrhundert. (German) [A previously unknown treatise on the abacus from the early 14th Century]. Historia Mathematica, 10(4):435–447, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900058.

Folkerts:1983:KVB

[Fol83b] Menso Folkerts. Kurt Vogel: Biographie und bibliographie. Historia Mathematica, 10(3):261–273, August 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086083900770.

Folkerts:1983:MHW

[Fol83c] Menso Folkerts. Materials for the history of Western European mathematics in the Middle Ages and Renaissance. Historia Mathematica, 10(1):98–99, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086083900423.

Folkerts:1984:HGO

[Fol84] Menso Folkerts. Helmuth Gericke: On the occasion of his 75th birthday. Historia Mathematica, 11(3):253–257, August 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086084900351.

Folkerts:1986:KV

[Fol86] Menso Folkerts. Kurt Vogel (1888–1985). Historia Mathematica, 13(3):224–228, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086086900893.

Folkerts:1987:BRH

[Fol87] Menso Folkerts. Book review: Humanistische Naturwissenschaft. Zur Geschichte der Wiener mathematischen Schulen des 15. und 16. Jahrhunderts: By Helmuth Grössing. (Saecula Spiritalia, herausgegeben von Dieter Wuttke, Band 8). Baden-Baden (Verlag Valentin Koerner). 1983. 355 Seiten. Historia Mathematica, 14(4):381–383, November 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086087900693.

Folkerts:1994:SKM

[Fol94] Menso Folkerts. Second Kenneth O. May Prize awarded in Zaragoza. Historia Mathematica, 21(1):8–9, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086084710044.

Folkerts:2009:BRC

[Fol09a] Menso Folkerts. Book review: Circa inicium algorismi. Die Algorismus-Vorlesung von Nicolaus Matz. Mit dem Faksimile des Autographen aus der Handschrift D 692 der Nicolaus-Matz-Bibliothek (Kirchenbibliothek) Michelstadt, Martin Hellmann (Ed.). Stadt Michelstadt, Michelstadt (2006). Historia Mathematica, 36(1):78–79, February 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008000591.

Folkerts:2009:HB

[Fol09b] Menso Folkerts. Hubertus L. L. Busard: (1923–2007). Historia Mathematica, 36(4):317–320, November 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086009000585.

Folkerts:2012:BRM

[Fol12] Menso Folkerts. Book review: Mathématiques et connaissance du monde réel avant Galilée. Edited by Sabine Rommeveaux. Montreuil (Omnisciene). 2010. ISBN 978-2-916097-26-8. 350 pp. 45. EUR. Historia Mathematica, 39(4):463–465, November 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086012000079.

Folkerts:2014:BRA

[Fol14] Menso Folkerts. Book review: Archimede Latino. Iacopo da San Cassiano e il corpus Archimedeo alla metà del quattrocento. Con edizione della Circuli dimensio e della Quadratura parabolae, Paolo d’Alessandro, Pier Daniele Napolitani (Eds.), in: Sciences et savoirs. Les Belles Lettres, Paris (2012), ISBN 978-2-251-22001-7. Historia Mathematica, 41(2):233–235, May 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508601400010X.

Folkerts:2015:BRL

[Fol15] Menso Folkerts. Book review: The Liber mahameleth. A 12thcentury mathematical treatise. 3 vols., Jacques Sesiano (Ed.), in: Sources and Studies in the History of Mathematics and Physical Sciences. Springer, Cham, Heidelberg, New York, Dordrecht, London (2014), ISBN: 978-3-319-03939-8. Historia Mathematica, 42(3):352– 354, August 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086014001074.

Forti:1975:BRB

[For75] Umberto Forti. Book review: Babbage, la macchina analitica [(Italian) [Babbage: the Analytical Engine]]: By Mario G. Losano. Milan (Etas Kompas), 1973, 191 p. Historia Mathematica, 2(3):350–353, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086075900853.

Forbes:1977:DBA

[For77] Eric G. Forbes. Descartes and the birth of analytic geometry. Historia Mathematica, 4(2):141–151, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086077901057.

Forbes:1978:AWC

[For78] Eric G. Forbes. The astronomical work of Carl Friedrich Gauss (1777–1855). Historia Mathematica, 5(2):167–181, May 1978.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078900484.

Ford:1998:IPF

[For98] Charles E. Ford. The influence of P. A. Florensky on N. N. Luzin. Historia Mathematica, 25(3):332–339, August 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086097921820.

Forcada:2010:BRM

[For10] Miquel Forcada. Book review: Les mathématiques infinitésimales du IXe au XIe siècle. Volume V: Ibn al-Haytham, Astronomie, géométrie sphérique et trigonométrie. By Roshdi Rashed. London (Al-Furqan Islamic Heritage Foundation). 2006. ISBN 1-905122-09-8, vol. 5. xiv, 972, v pp. No price given. Historia Mathematica, 37(2):302–305, May 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009000548.

Fowler:1984:EGT

[Fow84] D. H. Fowler. Eric Gardner Turner (1911–1983): In memoriam. Historia Mathematica, 11(2):126–130, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086084800038.

Fowler:1992:IRB

[Fow92a] D. H. Fowler. An invitation to read Book X of Euclid’s Elements. Historia Mathematica, 19(3):233–264, August 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609290029B.

Fowler:1992:DMS

[Fow92b] David H. Fowler. Dýnamis, mithartum, and square. Historia Mathematica, 19(4):418–419, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609290007X.

Fowler:1998:WRK

[Fow98] David Fowler. Wilbur Richard Knorr (1945–1997): an appreciation. Historia Mathematica, 25(2):123–132, May 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086098921991.

Ferraro:2003:DSR

[FP03] Giovanni Ferraro and Marco Panza. Developing into series and returning from series: a note on the foundations of Eighteenth-Century analysis. Historia Mathematica, 30(1):17–46, February 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086002000174.

Franci:1982:UWG

[FR82] Raffaella Franci and Laura Toti Rigatelli. Unpublished writings of Giuseppe S. Verzaglia. Historia Mathematica, 9(1):76–77, February 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082901410.

Franci:1983:MBF

[FR83] Raffaella Franci and Laura Toti Rigatelli. Maestro Benedetto da Firenze e la storia dell’algebra. (Italian) [Maestro Benedetto of Florence and the history of algebra]. Historia Mathematica, 10(3):297– 317, August 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086083900800.

Fowler:1998:SRA

[FR98] David Fowler and Eleanor Robson. Square root approximations in old Babylonian mathematics: YBC 7289 in context. Historia Mathematica, 25(4):366–378, November 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086098922091.

Folkerts:2008:HG

[FR08] Menso Folkerts and Karin Reich. Helmuth Gericke (1909–2007). Historia Mathematica, 35(3):170–172, August 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008000281.

Fraser:1987:JLL

[Fra87] Craig G. Fraser. Joseph Louis Lagrange’s algebraic vision of the calculus. Historia Mathematica, 14(1):38–53, February 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086087900024.

Franci:1988:AMA

[Fra88] Raffaella Franci. Antonio de’ Mazzinghi: an algebraist of the 14th Century. Historia Mathematica, 15(3):240–249, August 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608890064X.

Fraser:1992:IPV

[Fra92] Craig G. Fraser. Isoperimetric problems in the variational calculus of Euler and Lagrange. Historia Mathematica, 19(1):4–23, February 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290052D.

Franchella:1995:JBT

[Fra95] Miriam Franchella. L. E. J. Brouwer: Toward intuitionistic logic. Historia Mathematica, 22(3):304–322, ???? 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086085710269.

Fraser:2004:BRJ

[Fra04a] Craig Fraser. Book review: Jean Le Rond d’Alembert. OEuvres complètes Série I: Traités et mémoires mathématiques, 1736– 1756. Volume 6: Premiers textes de mécanique céleste, 1747–1749: Edited by Michelle Chapront–Touzé. Paris (CNRS Editions). 2002. 531 pp. EUR 60. Historia Mathematica, 31(3):368–369, August 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086003001009.

Fraser:2004:EN

[Fra04b] Craig Fraser. Editor’s note. Historia Mathematica, 31(1):1, February 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086004000151.

Fraser:2006:BRE

[Fra06] Craig Fraser. Book review: Essays on the History of Mechanics: In Memory of Clifford Ambrose Truesdell and Edoardo Benvenuto, Antonio Becchi, Massimo Corradi, Frederico Foce, Orietta Pedemonte (Eds.). Birkhäuser Verlag, Boston/Basel (2003), ISBN 3-7643-1476-1. Historia Mathematica, 33(2):253–254, May 2006. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086004000898.

Fraser:2009:SCF

[Fra09] Craig G. Fraser. Sufficient conditions, fields and the calculus of variations. Historia Mathematica, 36(4):420–427, November 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009000056.

Fraser:2010:BRP

[Fra10] Craig G. Fraser. Book review: The Princeton Companion to Mathematics. Edited by Timothy Gowers. Associate Editors June Barrow-Green and Imre Leader. Princeton (Princeton University Press). 2008. ISBN 978-0-691-11880-2. 1008 pp. $99.00. Historia Mathematica, 37 (1):110–112, February 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086009000664.

Fraser:2013:AMP

[Fra13] Craig Fraser. Awarding of the May Prizes for 2013. Historia Mathematica, 40(4):353–356, October 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086013000499.

Franchella:2016:FJK

[Fra16] Miriam Franchella. In the footsteps of Julius König’s paradox. Historia Mathematica, 43(1):65–86, February 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086015000737.

Fraser:2017:AMP

[Fra17a] Craig Fraser. Awarding of the May Prizes for 2017. Historia Mathematica, 44(4):315–317, November 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508601730085X.

Fraser:2017:BRH

[Fra17b] Craig Fraser. Book review: Heavenly Mathematics: the Forgotten Art of Spherical Trigonometry, Glen Van Brummelen. Princeton University Press (2013). Historia Mathematica, 44(1):71–73, February 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086016300477.

Freudenthal:1975:CMT

[Fre75] Hans Freudenthal. The cradle of modern topology, according to Brouwer’s inedita. Historia Mathematica, 2(4):495–502, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075901111.

Freudenthal:1976:BRT

[Fre76] H. Freudenthal. Book review: La topologie algébrique des origins à Poincaré: By Jean-Claude Pont. Bibliothéque de Philosophie contemporaine. (Presse Universitaire de France). 1974. Historia Mathematica, 3(3):350–352, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076901178.

Freudenthal:1977:BRN

[Fre77] Hans Freudenthal. Book review: Notice sur les travaux scientifiques, suivi de le parallelisme absolu et la théorie unitaire du champ: By Élie Cartan. Paris (Gauthier-Villars). 1974. Collection “Discours de la Méthode”. 128 p. Historia Mathematica, 4(2):218, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077901203.

Freudenthal:1980:HFP

[Fre80] Hans Freudenthal. Huygens’ foundations of probability. Historia Mathematica, 7(2):113–117, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900312.

Freeman:1982:P

[Fre82] James B. Freeman. Preface. Historia Mathematica, 9(3):263–264, August 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086082901215.

Freguglia:1989:SCS

[Fre89] Paolo Freguglia. Study centre for scientific thought between 1500 and 1600: Perugia university, 4th seminar of study. Historia Mathematica, 16(2):170, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086089900438.

Frei:1992:BRH

[Fre92] Günther Frei. Book review: Hermann Weyl 1885–1985: Edited by K. Chandrasekharan. Berlin/Heidelberg/New York (Springer-Verlag). 1986. vi + 109 pp. No price given. Historia Mathematica, 19(2):217– 229, May 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609290091O.

Frei:1993:BLV

[Fre93] Günther Frei. Bartel Leendert van der Waerden: Zum 90. Geburtstag. (German) [Bartel Leendert van der Waerden’s 90th birthday]. Historia Mathematica, 20(1):5–11, February 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086083710025.

Friday:1974:NMP

[Fri74a] J. R. Friday. A note on mathematical papers in the archives of the Royal institution of great Britain. Historia Mathematica, 1(1):85–86, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508607490175X.

Frisinger:1974:MHM

[Fri74b] H. Howard Frisinger. Mathematicians in the history of meteorology: the pressure-height problem from Pascal to Laplace. Historia Mathematica, 1(3):263–286, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074900664.

Friberg:1981:MTB

[Fri81a] Jöran Friberg. Methods and traditions of Babylonian mathematics: Plimpton 322, Pythagorean triples, and the Babylonian triangle parameter equations. Historia Mathematica, 8(3):277–318, August 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900690.

Frisinger:1981:SFT

[Fri81b] H. Howard Frisinger. The solution of a famous two-centuries-old problem: the Leonhard Euler–Latin square conjecture. Historia Mathematica, 8(1):56–60, February 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086081900057.

Fritzsche:1991:EAS

[Fri91] Bernd Fritzsche. Einige Anmerkungen zu Sophus Lies Krankheit. (German) [Some notes on Sophus Lie’s disease]. Historia Mathematica, 18(3):247–252, August 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609190393C.

Friedman:2021:MGC

[Fri21] Michael Friedman. On Mascheroni’s La geometria del compasso at the beginning of the 19th century. Historia Mathematica, 57(??):55–79, November 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086021000409.

Fried:2024:SU

[Fri24] Michael N. Fried. Sabetai Unguru 1931–2024. Historia Mathematica, 67(??):28–31, June 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086024000326.

Folkerts:2003:KRB

[FS03] Menso Folkerts and Christoph J. Scriba. Kurt-Reinhard Biermann (1919–2002). Historia Mathematica, 30(3):258–262, August 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086003000314.

Fenster:2010:EAH

[FS10] Della D. Fenster and Joachim Schwermer. Emil Artin und Helmut Hasse: Die korrespondenz 1923–1934. Historia Mathematica, 37(2):312–314, May 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086009000524.

Ferreira:2022:CNP

[FS22] Débora Ferreira and Gert Schubring. “Complex numbers” and the problem of multiplication between quantities. Historia Mathematica, 59(??):119–145, May 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086020300495.

Fowler:1983:HPE

[FT83] D. H. Fowler and E. G. Turner. Hibeh papyrus i 27: an early example of Greek arithmetical notation. Historia Mathematica, 10(3):344–359, August 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086083900824.

Fowler:1999:DEC

[FT99] David Fowler and Chr. Marinus Taisbak. Did Euclid’s circles have two kinds of radius? Historia Mathematica, 26(4):361–364, November 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086099922541.

Fu:1991:WDL

[Fu91] Daiwie Fu. Why did Liu hui fail to derive the volume of a sphere? Historia Mathematica, 18(3):212–238, August 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609190391A.

Fried:2001:APC

[FU01] Michael N. Fried and Sabetai Unguru. Apollonius of Perga’s Conica: text, context, subtext, volume 222 of Mnemosyne, bibliotheca classica Batava. Supplementum, 0169-8958. E. J. Brill, Leiden, The Netherlands, 2001. ISBN 90-04-11977-9. xii + 499 pp. LCCN QA31 .F75 2001.

Fukagawa:1987:ACJ

[Fuk87] Hidetosi Fukagawa. Algebraic curves in Japan during the Edo period. Historia Mathematica, 14(3):235–242, August 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086087900437.

Fuller:1999:HVH

[Ful99] A. Thomas Fuller. Horner versus Holdred: an episode in the history of root computation. Historia Mathematica, 26(1):29–51, February 1, 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086098922145.

Goldstein:2001:R

[GA01] Catherine Goldstein and Tom Archibald. Review. Historia Mathematica, 28(1):54–57, February 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086000922924.

Guillen:2020:NVQ

[GA20] Elías Fuentes Guillén and Carmen Martínez Adame. The notion of variable quantities ω in Bolzano’s early works. Historia Mathematica, 50(??):25–49, February 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086019300436.

Gabbey:1976:BRI

[Gab76] W. Allan Gabbey. Book review: Isaac Newton’s philosophiae naturalis principia mathematica: the Third Edition (1726) with variant readings. Assembled and edited by Alexandre Koyré and I. Bernard Cohen, with the assistance of Anne Whitman. 2 volumes. Cambridge, Mass. (Harvard University Press) and Cambridge (Cambridge University Press). 1972. 916 pp. U.S.$60. £25. Historia Mathematica, 3(2):237–243, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076900604.

Gabriel:1979:GFN

[Gab79] Gottfried Gabriel. Über einen Gedankenstrich bei Frege, eine Nachlese zur Edition seines wissenschaftlichen Nachlasses. (German) [A dash at Frege, a pickings for the edition of his scientific legacy]. Historia Mathematica, 6(1):34–35, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079901022.

Gaiduk:1975:FRW

[Gai75] Yu. M. Gaiduk. Further references on Wilbraham and Gibbs phenomenon. Historia Mathematica, 2(1):85, February 1975.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900543.

Galuzzi:2010:NAC

[Gal10] Massimo Galuzzi. Newton’s attempt to construct a unitary view of mathematics. Historia Mathematica, 37(3):535–562, August 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086010000303.

Galuzzi:2012:BRG

[Gal12] Massimo Galuzzi. Book review: Gerolamo Saccheri: Euclide vendicato da ogni neo. Edited by Vincenzo De Risi. 2 vols. Pisa (Edizioni della Normale). 2011. ISBN 978-88-7642-404-5. 148 + 252 pp. 40 EUR. Historia Mathematica, 39(1):107–108, February 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086011000747.

Gamelin:1996:BRH

[Gam96] T. W. Gamelin. Book review: A History of Complex Dynamics, from Schröder to Fatou and Julia. By Daniel S. Alexander. Wiesbaden (Vieweg). Historia Mathematica, 23(1):74–84, February 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086096900063.

Gana:1985:PDD

[Gan85] Francesco Gana. Peirce e Dedekind: La definizione di insieme finito. (Italian) [Peirce and Dedekind: The definition of the finite set]. Historia Mathematica, 12(3):203–218, August 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086085900199.

Gana:1986:DUN

[Gan86] Francesco Gana. Dio e l’uomo nella matematica di Kronecker. (Italian) [God and man in Kronecker’s mathematics]. Historia Mathematica, 13(3):255–276, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086086900923.

Garwood:1975:ADQ

[Gar75] Sue Ann Garwood. Additional discussion questions. Historia Mathematica, 2(4):603–604, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075901299.

Garland:1979:CCF

[Gar79] G. D. Garland. The contributions of Carl Friedrich Gauss to geomagnetism. Historia Mathematica, 6(1):5–29, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079901009.

Garciadiego:1985:ESN

[Gar85] Alejandro R. Garciadiego. The emergence of some of the nonlogical paradoxes of the theory of sets, 1903–1908. Historia Mathematica, 12(4):337–351, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086085900436.

Garciadiego:1986:FIC

[Gar86a] Alejandro R. Garciadiego. First international colloquium on the philosophy and history of mathematics. Historia Mathematica, 13 (3):292–293, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086086900984.

Garciadiego:1986:RHF

[Gar86b] Alejandro R. Garciadiego. On rewriting the history of the foundations of mathematics at the turn of the century. Historia Mathematica, 13(1):39–41, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086086902259.

Garciadiego:1992:IIC

[Gar92] Alejandro R. Garciadiego. II international colloquium on the history and philosophy of mathematics: December 10–13, 1990. Historia Mathematica, 19(1):84–87, February 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609290064I.

Garciadiego:1998:BRN

[Gar98] Alejandro R. Garciadiego. Book review: El nacimiento de la teoría de conjuntos, 1854–1908.By José Ferreirós Domínguez. Colección de Estudios, n^{∘} 39. Madrid (Ediciones de la Universidad Autónoma de Madrid). 1991 . Historia Mathematica, 25(3):318–320, August 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086098921930.

Gario:2001:GCD

[Gar01] Paola Gario. Guido Castelnuovo: Documents for a biography. Historia Mathematica, 28(1):48–53, February 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086000922936.

Garber:2005:BRG

[Gar05] Daniel Garber. Book review: Gottfried Wilhelm Leibniz, Sämtliche Schriften und Briefe, III. Reihe. Mathematischer, naturwissenschaftlicher und technischer Briefwechsel, Band 5: 1691–1693. Heinz–Jürgen Heß, James G. O’Hara (Eds.), Akademie Verlag, Munich, ISBN 3-05-003471-82003, 2003, 735 pp. 258 EUR. Historia Mathematica, 32(3):360–362, August 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086005000376.

Gauthier:2013:BRR

[Gau13] Sébastien Gauthier. Book review: Regards sur les mathématiques en France entre les deux guerres. Revue d’histoire des sciences 62 (1). Paris (Armand Colin). January-June 2009. ISBN 978-2-200-92598-7. 342 pp. 45 EUR. Historia Mathematica, 40(2):206–209, May 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086012000304.

Gowers:2008:PCM

[GBGL08] Timothy Gowers, June Barrow-Green, and Imre Leader, editors. The Princeton Companion to Mathematics. Princeton University Press, Princeton, NJ, USA, 2008. ISBN 0-691-11880-9 (hardcover). xx + 1034 pp. LCCN QA11.2 .P745 2008. URL http://www.loc.gov/

catdir/toc/ecip0818/2008020450.html.

Grinstein:1982:AJP

[GC82] Louise S. Grinstein and Paul J. Campbell. Anna Johnson Pell Wheeler: Her life and work. Historia Mathematica, 9(1):37–53, February 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086082901367.

Guillen:2021:ECE

[GC21] Elías Fuentes Guillén and Davide Crippa. The 1804 examination for the chair of Elementary Mathematics at the University of Prague. Historia Mathematica, 57(??):24–54.e18, November 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086021000422.

Gergonne:1974:AML

[Ger74] J. D. Gergonne. The application of the method of least squares to the interpolation of sequences. Historia Mathematica, 1(4):439–447, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074900342.

Gerdes:1985:TAM

[Ger85] Paulus Gerdes. Three alternate methods of obtaining the ancient Egyptian formula for the area of a circle. Historia Mathematica, 12(3):261–268, August 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086085900242.

Gerdes:1994:MHS

[Ger94] Paulus Gerdes. On mathematics in the history of Sub-Saharan Africa. Historia Mathematica, 21(3):345–376, August 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086084710299.

Gessner:2018:BRO

[Ges18] Samuel Gessner. Book review: Obras Matemáticas, Francisco de Melo, Vol. 1: Edição crítica e tradução, Bernardo Mota, Henrique Leitão. Centro de Estudos Clássicos, Biblioteca Nacional de Portugal, Lisboa (2014), 505 pages, 25 EUR., ISBN: 978-972-9376-31-3. Historia Mathematica, 45(2):182–185, May 2018. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0315086018300272.

Grattan-Guinness:1975:MBW

[GG75a] I. Grattan-Guinness. Mathematical bibliography for W. H. and G. C. Young. Historia Mathematica, 2(1):43–58, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900361.

Grattan-Guinness:1975:MPL

[GG75b] I. Grattan-Guinness. The mathematical papers and library of Sir Edward Collingwood in the University of Durham. Historia Mathematica, 2(2):200–202, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086075901524.

Grattan-Guinness:1975:PNH

[GG75c] I. Grattan-Guinness. Preliminary notes on the historical significance of quantification and of the axioms of choice in the development of mathematical analysis. Historia Mathematica, 2(4):475–488, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086075901093.

Grattan-Guinness:1975:RLP

[GG75d] I. Grattan-Guinness. Russell’s logical progress: Some new light from manuscript sources. Historia Mathematica, 2(4):489–493, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508607590110X.

Grattan-Guinness:1976:BRO

[GG76a] I. Grattan-Guinness. Book review: Oeuvres complètes d’Augustin Cauchy. II^{e} série — tome XV: Edited under the direction of the Paris Academy of Sciences and with the assistance of the Centre National de la Recherche Scientifique. Paris, Brussels and Montréal (Gauthier-Villars), 1974. vii + 640 pp. Historia Mathematica, 3(4):481–485, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086076900847.

Grattan-Guinness:1976:CWP

[GG76b] I. Grattan-Guinness. Corrections to the workshop proceedings. Historia Mathematica, 3(2):211, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086076900410.

Grattan-Guinness:1976:HMH

[GG76c] Ivor Grattan-Guinness. History of mathematics and histories of other sciences. Historia Mathematica, 3(3):325–327, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076901038.

Grattan-Guinness:1977:BRD

[GG77] I. Grattan-Guinness. Book review: Dirichlet’s principle. A mathematical comedy of errors and its influence on the development of analysis: By A. F. Monna. Utrecht (Oosthoek, Scheltema and Holkema). 1975. vii + 138 pp. Historia Mathematica, 4(2):222–225, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077901239.

Grattan-Guinness:1978:HBR

[GG78] I. Grattan-Guinness. How Bertrand Russell discovered his paradox. Historia Mathematica, 5(2):127–137, May 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078900460.

Grattan-Guinness:1979:MKG

[GG79] I. Grattan-Guinness. In memoriam Kurt Gödel: His 1931 correspondence with Zermelo on his incompletability theorem. Historia Mathematica, 6(3):294–304, August 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079901277.

Grattan-Guinness:1980:BRP

[GG80] I. Grattan-Guinness. Book review: The paradoxes of the theory of sets and dialectic: By I. N. Burova. Moscow (Nauka). 1976. 176pp. 62 Kopecks. Historia Mathematica, 7(1):92–93, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900816.

Grattan-Guinness:1986:BRW

[GG86] I. Grattan-Guinness. Book review: Writings of Charles S. Peirce: a chronological edition, volume 2, 1867–1871: Editor in chief: Edward C. Moore. Bloomington, Indiana (Indiana University Press). 1984. xlviii + 649 pp. $35.00. Historia Mathematica, 13(3):301–303, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086086901047.

Grattan-Guinness:1987:BSH

[GG87a] I. Grattan-Guinness. The British society for the history of mathematics. Historia Mathematica, 14(3):296, August 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086087900590.

Grattan-Guinness:1987:MMA

[GG87b] Ivor Grattan-Guinness. Mathematics, mechanics, and astronomy: Newton (1687), Lagrange (1788), Poincaré (1889), Gonville and Caius College, Cambridge, England, 17–20 September 1987. Historia Mathematica, 14(2):181–182, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086087900255.

Grattan-Guinness:1988:MMA

[GG88] Ivor Grattan-Guinness. Mathematics, mechanics, and astronomy: Newton (1687), Lagrange (1788), Poincaré (1889) September 17–20, 1987. Historia Mathematica, 15(2):165–170, May 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900092.

Grattan-Guinness:1990:BRL

[GG90a] I. Grattan-Guinness. Book review: A list of certain letters inserted in books from the Library of Augustus de Morgan (1806–1871) now in the University of London Library: By Maxine Merrington. London (University of London Library). 1990. 34 pp. £2.00. Historia Mathematica, 17(4):391, November 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609090040K.

Grattan-Guinness:1990:BRK

[GG90b] I. Grattan-Guinness. Book review: Kurt Gödel. Collected works. Volume II. Publications 1938–1974: Editor-in-chief Solomon Feferman. New York and Oxford: Oxford University Press. 1990. xv + 407 pp. No price stated. Historia Mathematica, 17(4):392–393, November 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090041B.

Grattan-Guinness:1990:BRE

[GG90c] I. Grattan-Guinness. Book review: The examiner. James Booth and the origins of common examinations: By Frank Foden. Leeds Studies in Adult and Continuing Education. Leeds (University of Leeds School of Continuing Education). 1989. vii + 221 pp. No price stated. Historia Mathematica, 17(3):291–293, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090020E.

Grattan-Guinness:1990:FFG

[GG90d] I. Grattan-Guinness. From Fourier to fractals: Gonville and Caius College, Cambridge, September 13–16, 1989. Historia Mathematica, 17(3):263–267, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086090900093.

Grattan-Guinness:1990:VM

[GG90e] Ivor Grattan-Guinness. The varieties of mechanics by 1800. Historia Mathematica, 17(4):313–338, November 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086090900259.

Grattan-Guinness:1992:NET

[GG92a] I. Grattan-Guinness. A note on the educational times and mathematical questions. Historia Mathematica, 19(1):76–78, February 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290057I.

Grattan-Guinness:1992:BFB

[GG92b] I. Grattan-Guinness. The Babbage/Faraday Bicentenary Conference: Cambridge University, England, July 5–7, 1991. Historia Mathematica, 19(2):207, May 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609290083N.

Golland:1993:ETS

[GG93] Louise Ahrndt Golland and Ronald William Golland. Euler’s troublesome series: an early example of the use of trigonometric series. Historia Mathematica, 20(1):54–67, February 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086083710062.

Grattan-Guinness:1994:CET

[GG94] I. Grattan-Guinness. Contributing to the educational times: Letters to W. J. C. Miller. Historia Mathematica, 21(2):204–205, May 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086084710196.

Grattan-Guinness:1996:NMR

[GG96] Ivor Grattan-Guinness. Numbers, magnitudes, ratios, and proportions in Euclid’s Elements: How did he handle them? Historia Mathematica, 23(4):355–375, November 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086096900385.

Grattan-Guinness:1997:NMR

[GG97] Ivor Grattan-Guinness. Numbers, magnitudes, ratios, and proportions in Euclid’s Elements: How did he handle them?: Volume 23, number 4 (1996), pages 355–375. Historia Mathematica, 24(2): 213, May 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086097921558.

Grattan-Guinness:1998:RRV

[GG98] Ivor Grattan-Guinness. A response to the review by Victor J. Katz (Historia Mathematica 24 (1997), 445–452) of Ivor Grattan-Guinness, Ed., Companion Encyclopedia of the History and Philosophy of the Mathematical Sciences. Historia Mathematica, 25(1):94, February 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086098921966.

Grattan-Guinness:2002:APG

[GG02] I. Grattan-Guinness. Algebras, projective geometry, mathematical logic, and constructing the world: Intersections in the philosophy of mathematics of A. N. Whitehead. Historia Mathematica, 29(4):427– 462, November 2002. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086002923566. See erratum [GG03].

Grattan-Guinness:2003:EAP

[GG03] I. Grattan-Guinness. Erratum to “Algebras, projective geometry, mathematical logic, and constructing the world: Intersections in the philosophy of mathematics of A. N. Whitehead”: [Historia Mathematica 29 (2002) 427–462]. Historia Mathematica, 30(1):96, February 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086003000077. See [GG02].

Grattan-Guinness:2004:MPD

[GG04] I. Grattan-Guinness. The mathematics of the past: distinguishing its history from our heritage. Historia Mathematica, 31(2):163–185, May 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086003000326.

Grattan-Guinnes:2005:BRL

[GG05a] I. Grattan-Guinnes. Book review: L’Archivio Giuseppe Peano, 2002. ISBN 88-9007876-1-7. EUR 30. L’opera omnia di Giuseppe Peano, 2002. ISBN 88-9007876-2-5. EUR 50 Le riviste di Giuseppe Peano, 2003. ISBN 88-9007876-7-6. EUR 30 Edited by Clara Silvia Roero [L’Archivio also with Natalia Nervo and Tiziana Armano]. Turin (Department of Mathematics, University of Turin [L’Archivio also with the Comune di Cuneo]) [three CD-ROMs]. Historia Mathematica, 32 (4):489–492, November 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086004000886.

Grattan-Guinness:2005:BRF

[GG05b] I. Grattan-Guinness. Book review: Francesco Faà di Bruno. Ricerca scientifica insegnamento e divulgazione, vol. 12, Livia Giacardi (Ed.), in: Studi e fonti per la storia della Università di Torino. Palazzo Carignano, Turin (2004). Historia Mathematica, 32(4):488–489, November 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086004000874.

Grattan-Guinness:2005:BRL

[GG05c] I. Grattan-Guinness. Book review: In the Light of Logic, Solomon Feferman. Oxford Univ. Press, New York, Oxford (1998), ISBN 0-19-508030-0. Historia Mathematica, 32(3):366–367, August 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508600400093X.

Grattan-Guinness:2005:JG

[GG05d] I. Grattan-Guinness. John Greenberg (1945–2004). Historia Mathematica, 32(1):1–3, February 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086004000813.

Grattan-Guinness:2007:BRK

[GG07] I. Grattan-Guinness. Book review: Kardinalität und Kardinäle. Wissenschaftshistorische Aufarbeitung der Korrespondenz zwischen Georg Cantor und katholischen Theologen seiner Zeit, vol. 53, Christian Tapp, in: Boethius: Texte und Abhandlungen zur Geschichte der Mathematik und der Naturwissenschaften. Franz Steiner Verlag, Stuttgart (2005), ISBN 3-515-08620-X. Historia Mathematica, 34 (1):111–112, February 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086006000206.

Grattan-Guinness:2008:BRE

[GG08] I. Grattan-Guinness. Book review: Euler at 300. An appreciation, R. E. Bradley, L. A. D’Antonio, C. E. Sandifer (Eds.). The Mathematical Association of America, Washington, DC (2007), ISBN 978-0-88385-565-2. Historia Mathematica, 35(2):143–144, May 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508600700095X.

Grattan-Guinness:2010:BRH

[GG10] I. Grattan-Guinness. Book review: Hermann Graßmann. Biography. By Hans-Joachim Petsche. Translated by Mark Minnes. Basel (Birkhäuser). 2009. ISBN 978-3-7643-8859-1. xx + 306 pp. No price stated. Historia Mathematica, 37(4):714–716, November 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009001293.

Grattan-Guinness:2013:BRW

[GG13] Ivor Grattan-Guinness. Book review: Wolfgang Sartorius von Waltershausen, Gaußzum Gedächtniss: Biographie Carl Friedrich Gauß, Leipzig 1856. Mit dem von Karin Reich verfassten Essay “Wolfgang Sartorius von Waltershausen (1809–1876)”, Karin Reich (Ed.). Edition Am Gutenbergplatz, Leipzig (2012), ISBN: 978-3-937219-57-8. Historia Mathematica, 40(3):329–330, August 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086012000754.

Grattan-Guinness:2014:AFL

[GG14] I. Grattan-Guinness. From anomaly to fundament: Louis Poinsot’s theories of the couple in mechanics. Historia Mathematica, 41 (1):82–102, February 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086013000475.

Grattan-Guinness:2015:WC

[GG15] I. Grattan-Guinness. Working on Coriolis. Historia Mathematica, 42(2):223–229, May 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086014000731.

Grattan-Guinness:2005:LWW

[GGC^{+}05] I. Grattan-Guinness, Roger Cooke, et al., editors. Landmark writings in Western mathematics 1640–1940. Elsevier, Amsterdam, The Netherlands, 2005. ISBN 0-444-50871-6 (hardcover), 0-08-045744-4 (electronic bk.). xvii + 1022 pp. LCCN QA21. URL http://www.

myilibrary.com?id=63373.

Grattan-Guinness:1982:MPC

[GGE82] I. Grattan-Guinness and S. Engelsman. The manuscripts of Paul Charpit. Historia Mathematica, 9(1):65–75, February 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082901409.

Giacardi:1990:MDH

[Gia90] Livia Giacardi. Memorial days in honour of Angelo Genocchi: Turin, Italy, May 12–13, 1989. Historia Mathematica, 17(1):66–67, February 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090081N.

Giacardi:2001:CSA

[Gia01] Livia Giacardi. The Corrado Segre Archive. Historia Mathematica, 28(4):296–301, November 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086001923274.

Gillispie:1975:BRD

[Gil75] Charles C. Gillispie. Book review: Diderot: By Arthur M. Wilson. Oxford University Press, 1972. 935 p. Historia Mathematica, 2(3):342–344, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086075900816.

Gillon:1977:ITD

[Gil77] Brendan S. Gillon. Introduction, translation, and discussion of chao Chün–Ch’ing’s “Notes to the diagrams of short legs and long legs and of squares and circles”. Historia Mathematica, 4(3):253–293, August 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086077900556.

Gillings:1978:RSC

[Gil78] R. J. Gillings. Response to “Some comments on R. J. Gillings’ analysis of the 2/n table in the Rhind Papyrus”. Historia Mathematica, 5(2):221–226, May 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508607890054X. See [BS77b].

Gillings:1979:RRM

[Gil79] R. J. Gillings. The recto of the Rhind Mathematical Papyrus and the Egyptian mathematical leather roll. Historia Mathematica, 6 (4):442–447, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086079900314.

Gillings:1981:EML

[Gil81] Richard J. Gillings. The Egyptian mathematical leather role — Line 8. How did the scribe do it? Historia Mathematica, 8(4):456–457, November 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086081900537.

Gillies:1987:WBB

[Gil87] Donald A. Gillies. Was Bayes a Bayesian? Historia Mathematica, 14(4):325–346, November 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086087900656.

Gillispie:1989:SBH

[Gil89] Charles C. Gillispie. Salomon Bochner as historian of mathematics and science. Historia Mathematica, 16(4):316–323, November 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086089900785.

Gilain:2015:BRC

[Gil15] Christian Gilain. Book review: Correspondance de Pierre Simon Laplace (1749–1827), Roger Hahn (Ed.). Brepols, Turnhout (2013), ISBN: 978-2-503-54129-7. Historia Mathematica, 42(1):102–105, February 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086014000494.

Gingerich:1983:BRJ

[Gin83] Owen Gingerich. Book review: Johannes Kepler — Mysterium cosmographicum — The secret of the universe: Translation by A. M. Duncan. Introduction and Commentary by E. J. Aiton. New York (Abaris Books). 1981. 268 pp. $20. Historia Mathematica, 10 (1):116–117, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086083900514.

Gispert:1996:BRH

[Gis96] Hélène Gispert. Book review: Histoire de fractions, fractions d’histoire. Coordinated by Paul Benoit, Karine Chemla, and Jim Ritter. Science Networks 10. Basel (Birkhäuser). Historia Mathematica, 23(3):319–323, August 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086096900300.

Gispert:1999:BRE

[Gis99a] Hélène Gispert. Book review: Eugène Catalan: Géomètre sans patrie, républicain sans république. By François Jongmans. Historia Mathematica, 26(1):77–79, February 1, 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086099922206.

Gispert:1999:DHM

[Gis99b] Hélène Gispert. Les débuts de l’histoire des mathématiques sur les scènes internationales et le cas de l’entreprise encyclopédique de Felix Klein et Jules Molk. Historia Mathematica, 26(4):344–360, November 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086099922607.

Gispert:2014:BRC

[Gis14] Hélène Gispert. Book review: Charles-Ange Laisant. Itinéraires et engagements d’un mathématicien de la Troisième République, Jérôme Auvinet. Hermann, Paris (2013), ISBN: 978-2-7056-8697-0. Historia Mathematica, 41(3):333–337, August 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086014000329.

Gispert:2018:AEU

[Gis18] Hélène Gispert. Une ambition éditoriale “universelle et confraternelle” : le bulletin bibliographique de L’Enseignement mathématique (1899–1920). (French) [A “universal and confraternal” editorial ambition: the bibliographic bulletin of mathematics teaching (1899– 1920)]. Historia Mathematica, 45(4):414–432, November 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086018300879.

Gittleman:1974:BRD

[Git74] Arthur P. Gittleman. Book review: Diofant i diofantovy uravneniya: By I. G. Bashmakova. Moscow (Nauka), 1972, 68 pp. 12 k. Historia Mathematica, 1(1):96–97, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074901876.

Giusti:1986:CEH

[Giu86] Enrico Giusti. Critical editions and history of mathematics. Historia Mathematica, 13(1):72–73, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086086902363.

Giusti:1999:BRM

[Giu99] Enrico Giusti. Book review: Les mathématiques infinitésimales du IXe au XIe siècle. Edited by Roshdi Rashed. Historia Mathematica, 26(2):161–164, May 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086099922309.

Glushkov:1976:AML

[Glu76] Stanisław Glushkov. On approximation methods of Leonardo Fibonacci. Historia Mathematica, 3(3):291–296, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900987.

Glushkov:1977:IVC

[Glu77] Stanislav Glushkov. An interpretation of Viète’s “Calculus of triangles” as a precursor of the algebra of complex numbers. Historia Mathematica, 4(2):127–136, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086077901033.

Glushkov:1978:A

[Glu78] S. Glushkov. An acknowledgement. Historia Mathematica, 5(1):92, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086078901416.

Gluchoff:2005:PMA

[Glu05] Alan Gluchoff. Pure mathematics applied in early Twentieth-Century America: the case of T. H. Gronwall, consulting mathematician. Historia Mathematica, 32(3):312–357, August 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086004000448.

Gluchoff:2011:AMF

[Glu11] Alan Gluchoff. Artillerymen and mathematicians: Forest ray Moulton and changes in American exterior ballistics, 1885–1934. Historia Mathematica, 38(4):506–547, November 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086011000279.

Gluchoff:2012:BRH

[Glu12] Alan Gluchoff. Book review: The History and Development of Nomography. By H. A. Evesham. (Docent Press). 2011 (original copyright 1982). 267 pp., paperback. Historia Mathematica, 39 (4):469–475, November 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086012000092.

Guerraggio:2004:OQC

[GM04] Angelo Guerraggio and Elena Molho. The origins of quasi-concavity: a development between mathematics and economics. Historia Mathematica, 31(1):62–75, February 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086003000570.

Guerraggio:1992:IMC

[GN92] Angelo Guerraggio and Pietro Nastasi. Italian mathematics of the 19th and 20th centuries. publication of correspondence. Historia Mathematica, 19(1):91–92, February 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609290067L.

Guerraggio:2005:IMB

[GN05] Angelo Guerraggio and Pietro Nastasi. Italian mathematics between the two World Wars, volume 29 of Science networks historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2005. ISBN 3-7643-6555-2. x + 299 pp.

LCCN QA13.5.I8 G84 2005. URL http://www.loc.gov/catdir/
enhancements/fy0663/2005045206-d.html;   http://www.loc.
gov/catdir/enhancements/fy0812/2005045206-t.html.

Gray:1999:TRI

[GO99] Jeremy Gray and Eduardo L. Ortiz. On the transmission of Riemann’s ideas to Portugal. Historia Mathematica, 26(1):52–67, February 1, 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086098922182.

Goethe:2008:MRW

[Goe08] Norma B. Goethe. Modes of representation, working tools, and the history of mathematics. Historia Mathematica, 35(3):242–247, August 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008000074. See erratum [Goe09].

Goethe:2009:EMR

[Goe09] Norma B. Goethe. Erratum to “Modes of representation, working tools, and the history of mathematics” [Historia Math. 35 (3) (2008) 242–247]. Historia Mathematica, 36(1):109, February 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508600800102X. See [Goe08].

Goldstein:1976:BRA

[Gol76] Bernard R. Goldstein. Book review: Der Almagest: Die Syntaxis mathematica des Claudius Ptolemäus in arabisch-lateinscher Überlieferung: By Paul Kunitzsch. Wiesbaden (Otto Harrassowitz). 1974. xvi + 384 p.; 10 plates. DM 160. Historia Mathematica, 3 (4):488–489, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076900860.

Goldstein:1980:SLV

[Gol80] Bernard R. Goldstein. Solar and lunar velocities in the Alfonsine Tables. Historia Mathematica, 7(2):134–140, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900348.

Goldstein:1982:BRI

[Gol82] Bernard R. Goldstein. Book review: Les instruments de la théorie des planetes selon Ptolemee: Équatoires et horologerie planétaire du XIII^{e} au XVI^{e} siècle: By Emmanuel Poulle. Geneva (Librairie Droz). 1980. 2 vol. 1162 pp. 208 figures. 70 plates. 400 FF. Historia Mathematica, 9(1):97–99, February 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086082901458.

Goldstein:1984:EME

[Gol84] Bernard R. Goldstein. Eratosthenes on the “measurement” of the earth. Historia Mathematica, 11(4):411–416, November 1984.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086084900259.

Goldstein:1992:SCV

[Gol92] Catherine Goldstein. On a seventeenth century version of the “fundamental theorem of arithmetic”. Historia Mathematica, 19(2):177– 187, May 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609290075M.

Goldstein:2000:MGM

[Gol00] Joel A. Goldstein. A matter of great magnitude: the conflict over arithmetization in 16th-, 17th-, and 18th-Century English editions of Euclid’s Elements Books I through VI (1561–1795). Historia Mathematica, 27(1):36–53, February 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086099922632.

Goldstein:2013:BRE

[Gol13] Catherine Goldstein. Book review: Étienne Bézout (1730–1783) Mathématicien des Lumières, Liliane Alfonsi. L’Harmattan, Paris (2011), ISBN: 978-2-296-56553-1. Historia Mathematica, 40(1):93–96, February 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508601200033X.

Goncalves:2008:APR

[Gon08] Carlos H. B. Gonçalves. An alternative to the Pythagorean rule? Reevaluating Problem 1 of cuneiform tablet BM 34 568. Historia Mathematica, 35(3):173–189, August 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086007001000.

Goodstein:1977:RLC

[Goo77] Judith R. Goodstein. Ricci and Levi-Civita’s tensor analysis paper: Lie Groups: History, Frontiers, and Applications, Volume 2. Translated and edited by Robert Hermann. Brookline, Massachusetts (Math. Sci. Press). 1975. 261 p. paper. Historia Mathematica, 4(2): 228, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086077901264.

Goulding:2010:DHR

[Gou10] Robert Goulding. Defending Hypatia: Ramus, Savile, and the Renaissance rediscovery of mathematical history, volume 25 of Archimedes, 1385-0180. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 90-481-3541-9 (hardcover), 90-481-3542-7 (e-book). xx + 201 pp. LCCN QA24 .G68 2010. URL

http://www.loc.gov/catdir/enhancements/fy1304/2009941801-
d.html;   http://www.loc.gov/catdir/enhancements/fy1304/
2009941801-t.html.

Gowing:1995:HCN

[Gow95] Ronald Gowing. Halley, Cotes, and the nautical meridian. Historia Mathematica, 22(1):19–32, February 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086085710026.

Garibaldi:1991:IVE

[GP91] U. Garibaldi and M. A. Penco. Intensional vs extensional probabilities from their origins to Laplace. Historia Mathematica, 18(1):16–35, February 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086091903493.

Gray:2007:EHM

[GP07] Jeremy Gray and Karen Hunger Parshall, editors. Episodes in the history of modern algebra (1800–1950), volume 32 of History of mathematics. American Mathematical Society, Providence, RI, USA, 2007. ISBN 0-8218-4343-5. viii + 336 pp. LCCN QA151 .E65 2007.

Grabiner:1975:MHH

[Gra75a] Judith V. Grabiner. The mathematician, the historian, and the history of mathematics. Historia Mathematica, 2(4):439–447, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075901019.

Gray:1975:AMJ

[Gra75b] J. D. Gray. Archives of mathematical journals. Historia Mathematica, 2(2):202, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086075901536.

Granger:1976:BRC

[Gra76a] Gilles Gaston Granger. Book review: Condorcet mathématique et société: Choix de textes et commentaire. By Roshdi Rashed. Paris (Hermann). 1974. ISBN 2 7056 1. 218 pp. Historia Mathematica, 3 (4):496–497, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076900914.

Gray:1976:CCW

[Gra76b] J. D. Gray. Comments on collected works, in particular those of Émile Borel and Jacques Hadamard. Historia Mathematica, 3(2):203– 206, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086076900380.

Gray:1976:AGM

[Gra76c] J. J. Gray. The 1975 annual general meeting of the British society for the history of mathematics. Historia Mathematica, 3(2):217–218, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508607690046X.

Grabiner:1977:BRM

[Gra77] Judith V. Grabiner. Book review: Men and institutions in American mathematics: Edited by J. Dalton Tarwater, John T. White, and John D. Miller. Texas Tech University (Graduate Studies, No. 13). 1976. $5.00. Historia Mathematica, 4(4):464–465, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900921.

Grabiner:1978:OCT

[Gra78a] Judith V. Grabiner. The origins of Cauchy’s theory of the derivative. Historia Mathematica, 5(4):379–409, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078902082.

Gray:1978:BSH

[Gra78b] J. J. Gray. The British Society for the History of Mathematics. Historia Mathematica, 5(4):467–468, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078902185.

Gray:1978:JHL

[Gra78c] J. J. Gray. Johann Heinrich Lambert, mathematician and scientist, 1728–1777. Historia Mathematica, 5(1):13–41, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078901337.

Gray:1978:GEG

[Gra78d] Jeremy Gray. Gauss und die nicht-euklidische Geometrie. Historia Mathematica, 5(4):482–483, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078902276.

Gray:1979:NEG

[Gra79] Jeremy Gray. Non-Euclidean geometry — a re-interpretation. Historia Mathematica, 6(3):236–258, August 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079901241.

Gray:1980:HCF

[Gra80a] Jeremy Gray. The history of the concept of a finite-dimensional vector space. Historia Mathematica, 7(1):65–70, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900658.

Gray:1980:NKM

[Gra80b] Jeremy Gray. A note on Karl M. Peterson. Historia Mathematica, 7(4):444, November 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086080900075.

Grant:1983:BRJ

[Gra83] Edward Grant. Book review: Jordanus de nemore. De numeris datis: a critical edition and translation by Barnabas Bernard Hughes, O.F.M. Berkeley (University of California Press). 1981. xi + 212 pp. Historia Mathematica, 10(1):125–129, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900551.

Grabiner:1984:AID

[Gra84a] Judith V. Grabiner. Artificial intelligence: Debates about its use and abuse. Historia Mathematica, 11(4):471–480, November 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086084900296.

Gray:1984:WHM

[Gra84b] Jeremy Gray. Workshop on the history of mathematics, 14–15 April 1984. Historia Mathematica, 11(3):323–324, August 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086084900387.

Gray:1986:NTA

[Gra86] Jeremy Gray. Number theory: an approach through history; from Hammurapi to Legendre. Historia Mathematica, 13(1):86–88, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086086902429.

Gray:1987:RSH

[Gra87] Jeremy Gray. A report on the Symposium on the History of Recent Mathematics: Sandbjerg, Denmark, August 21–25, 1986. Historia Mathematica, 14(1):62–63, February 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086087900085.

Gray:1990:BRG

[Gra90] Jeremy Gray. Book review: Die Geschichte der geometrischen Mechanik im 19. Jahrhundert: By Renatus Ziegler. Franz Steiner Verlag Wiesbaden GmbH Stuttgart. Boethius XIII. 1985. vii + 260 pp. DM 48. Historia Mathematica, 17(1):76–77, February 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090087T.

Gray:1991:BRHb

[Gra91a] Jeremy Gray. Book review: A history of algebraic and differential topology, 1900–1960: By Jean Dieudonné. Boston and Basel (Birkhäuser). 1989. xxi + 648 pp. Historia Mathematica, 18(4): 374–377, November 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086091903817.

Gray:1991:BRHa

[Gra91b] Jeremy Gray. Book review: A history of non-Euclidean geometry: Evolution of the concept of a geometric space: By B. A. Rosenfeld. Translated by Abe Shenitzer. With the editorial assistance of Hardy Grant. New York (Springer-Verlag). 1988. xi + 471 pp. Historia Mathematica, 18(4):373–374, November 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609190380G.

Grabiner:1994:BRA

[Gra94a] Judith V. Grabiner. Book review: Augustin-Louis Cauchy: a Biography: By Bruno Belhoste. Translated by Frank Ragland. New York, Berlin, Heidelberg, London, Paris, Tokyo, Hong Kong, Barcelona (Springer-Verlag). 1991. $79.00. Historia Mathematica, 21(2):215– 220, May 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086084710214.

Gray:1994:BRK

[Gra94b] Jeremy Gray. Book review: A. N. Kolmogorov, A. P. Yushkevich, Editors, Mathematics of the 19th century: Mathematical logic, algebra, number theory, probability Theory (1992) Birkhäuser,Boston and Basel 308 $149.50. Historia Mathematica, 21(1):93–94, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Gray:1994:BRM

[Gra94c] Jeremy Gray. Book review: Mathematics of the 19th century: Mathematical logic, algebra, number theory, probability theory, A. N. Kolmogorov, A. P. Yushkevich (Eds.). Birkhäuser, Boston and Basel (1992), 308. Historia Mathematica, 21(1):93–94, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086084710111.

Gray:1996:BRD

[Gra96] Jeremy Gray. Book review: Development of Mathematics, 1900–1950. Edited by Jean-Paul Pier. Historia Mathematica, 23(4):437–440, November 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086096900415.

Gray:1997:BRM

[Gra97a] Jeremy Gray. Book review: Mathematics of the 19th Century; Geometry, Analytic Function Theory. Edited by A. N. Kolmogorov and A. P. Yushkevich, and translated from the Russian by Roger Cooke. Basel/Boston/Berlin (Birkhäuser). 1996. ISBN 3-7643-5048-2. x + 290 pp., sFr 98. Historia Mathematica, 24(3):332–334, August 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086097921431.

Gray:1997:RML

[Gra97b] Jeremy Gray. Reply to “Is mathematical logic a part of Mathematics?”. Historia Mathematica, 24(4):454–456, November 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086097921765.

Gray:1998:BRP

[Gra98a] Jeremy Gray. Book review: Psychologism: a Case Study in the Sociology of Philosophical Knowledge. By Martin Kusch. London and New York (Routledge). 1995. Historia Mathematica, 25(3):325–326, August 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086098922054.

Gray:1998:RRV

[Gra98b] Jeremy Gray. A response to the review by Victor J. Katz (Historia Mathematica 24 (1997), 445–452) of Ivor Grattan-Guinness, Ed., Companion Encyclopedia of the History and Philosophy of the Mathematical Sciences. Historia Mathematica, 25(1):95–97, February 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086098921978.

Gray:1999:BRD

[Gra99a] Jeremy Gray. Book review: La découverte de la géométrie non euclidienne sur la pseudosphère: Les lettres d’Eugenio Beltrami à Jules Hoüel (1868–1881). Introduction, notes and critical commentaries by Luciano Boi, Livia Giacardi, Rossana Tazzioli. Historia Mathematica, 26(2):164–166, May 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086099922322.

Gray:1999:BRR

[Gra99b] Jeremy Gray. Book review: The Riemann Legacy: Riemannian Ideas in Mathematics and Physics. Edited by K. Maurin. Historia Mathematica, 26(3):291, August 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086099922371.

Gray:2001:MPH

[Gra01] Jeremy Gray. Matériaux pour l’Histoire des Mathématiques au XX^{e} Siècle. actes du colloque à la mémoire de Jean Dieudonné (Nice, 1996). Historia Mathematica, 28(2):123–125, May 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508600192302X.

Gray:2002:JFJ

[Gra02] Jeremy Gray. John Fauvel (21 July 1947, Glasgow–12 May 2001, Leamington Spa). Historia Mathematica, 29(3):239–243, August 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086002923591.

Gray:2006:BRM

[Gra06] Jeremy Gray. Book review: The Mathematical Works of Bernard Bolzano. By Steve Russ. Oxford (Oxford University Press). 2004. ISBN 0-19-853930-4. xxx + 698 pp. $229.50. Historia Mathematica, 33(3):369–371, August 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086005000716.

Gray:2008:BRT

[Gra08] Jeremy Gray. Book review: Les Théorèmes de Noether. Invariance et lois de conservation au XXe siècle, avec une traduction de l’article original, “Invariante Variationsprobleme”, 2nd, revised edition, Y. Kosmann-Schwarzbach, L. Meersseman. Éditions de l’École Polytechnique, Palaiseau (2006), ISBN 2-7302-1138-1. Historia Mathematica, 35(3):252–253, August 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086007000365.

Grant:2009:WWS

[Gra09a] Hardy Grant. What’s in a word? symmetry through the centuries. Historia Mathematica, 36(2):171–177, May 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008000992.

Gray:2009:BRG

[Gra09b] Jeremy Gray. Book review: The Geometry of an Art: the History of the Mathematical Theory of Perspective from Alberti to Monge, Kirsti Andersen. Springer, New York (2007). Historia Mathematica, 36(2):182–183, May 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008000657.

Grant:2010:BRO

[Gra10a] Hardy Grant. Book review: The Oxford Handbook of the History of Mathematics. Edited by Eleanor Robson and Jacqueline Stedall. Oxford (Oxford University Press). 2009. ISBN 978-0-19-921312-2. vii + 918 pp. £85.00. Historia Mathematica, 37(1):112–118, February 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009000706.

Gray:2010:BRE

[Gra10b] Jeremy Gray. Book review: Erhabene Welten: das Leben Rolf Nevanlinnas. By Olli Lehto. Translated from Finnish by M. Stern. Boston (Birkhäuser). 2008. ISBN 978-7643-7701-4. ix + 299 pp. No price given. Historia Mathematica, 37(2):314–316, May 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508600900055X.

Grant:2011:BRD

[Gra11] Hardy Grant. Book review: Duel at Dawn: Heroes, Martyrs and the Rise of Modern Mathematics. By Amir Alexander. Cambridge, Mass. and London (Harvard University Press). 2010. ISBN 978-0-674-04661-0. 307 pp. US$28.95. Historia Mathematica, 38(1):125–131, February 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086010000339.

Gray:2012:BRE

[Gra12] Jeremy Gray. Book review: Eléments d’une biographie de l’espace projectif. By Lise Bioesmat-Martagon. Collection Histoires de géométries. Nancy (Presses Universitaires de Nancy). 2010. ISBN 978-2-8143-0032-3. 315 pp. 25 EUR. Historia Mathematica, 39 (1):105–106, February 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086011000735.

Grant:2013:BRH

[Gra13] Hardy Grant. Book review: Hidden Harmonies: the Lives and Times of the Pythagorean Theorem, Robert Kaplan, Ellen Kaplan. Bloomsbury Press, New York, Berlin, London, and Sydney (2011), ISBN: 978-1-59691-522-0. Historia Mathematica, 40(2):203–205, May 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086012000158.

Gray:2018:BRE

[Gra18] Jeremy Gray. Book review: Emil Artin and Beyond — Class Field Theory and L-Functions, Della Dumbaugh, Joachim Schwermer. European Mathematical Society (2015), 231 pp. $76.00, ISBN: 978-3-03719-146-0. Historia Mathematica, 45(2):193–195, May 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S0315086018300399.

Grcar:2011:HOE

[Grc11] Joseph F. Grcar. How ordinary elimination became Gaussian elimination. Historia Mathematica, 38(2):163–218, May 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086010000376.

Greenwood:1977:KHM

[Gre77] Robert E. Greenwood. The kinship of E. H. Moore and R. L. Moore. Historia Mathematica, 4(2):153–155, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086077901069.

Gregory:1983:NKF

[Gre83] Frederick Gregory. Neo-Kantian foundations of geometry in the German Romantic Period. Historia Mathematica, 10(2):184–201, May 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900629.

Greenberg:1984:AFR

[Gre84] John L. Greenberg. Alexis Fontaine’s route to the calculus of several variables. Historia Mathematica, 11(1):22–38, February 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608490003X.

Greenberg:1985:BRP

[Gre85] John L. Greenberg. Book review: Penser les mathématiques. Séminaire de philosophie et mathématiques de L’École normale supérieure: Edited by J. Dieudonné, M. Loi, and R. Thom. Paris (Éditions du Seuil). 1982. 273 pp. Historia Mathematica, 12(4): 388–389, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086085900540.

Greenberg:1988:BVC

[Gre88] John L. Greenberg. Breaking a “viscous circle”: Unscrambling A.-C. Clairaut’s iterative method of 1743. Historia Mathematica, 15(3):228– 239, August 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086088900638.

Gridgeman:1976:BRB

[Gri76] N. T. Gridgeman. Book review: A Boolean anthology: Selected writings of Mary Boole. By D. G. Tahta. London (Association of Teachers of Mathematics.) 1972. 79 p. 0.50. Historia Mathematica, 3(1):102–103, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076900252.

Gridgeman:1982:BRR

[Gri82] N. T. Gridgeman. Book review: R. A. Fisher. The life of a scientist: By Joan Fisher Box. New York (John Wiley & Sons), 1978. xii + 512 pp. Illustrated. Historia Mathematica, 9(1):104–106, February 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082901483.

Grossing:1986:BRN

[Grö86] Helmut Grössing. Book review: Nicolaus Cusanus und die Entstechung der exakten Wissenschaften: By Fritz Nagel. Buchreihe der Cusanus-Gesellschaft, Band IX. Münster (Aschendorff). 1984. vii + 192 pp. DM 48. Historia Mathematica, 13(3):308–310, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086086901072.

Grosholz:1987:TLM

[Gro87] Emily Grosholz. Two Leibnizian manuscripts of 1690 concerning differential equations. Historia Mathematica, 14(1):1–37, February 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086087900012.

Gray:1977:BSH

[GS77] J. J. Gray and Hon Secretary. The British Society for the History of Mathematics. Historia Mathematica, 4(2):211–213, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077901161.

Grunbaum:1991:BRS

[GS91] Branko Grünbaum and Marjorie Senechal. Book review: Symmetrie– Gruppe–Dualität: By Erhard Scholz. Basel (Birkhäuser). 1980. 406 pp. Historia Mathematica, 18(4):377–380, November 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086091903828.

Guerraggio:1988:IMB

[Gue88] Angelo Guerraggio. Italian mathematics between the two world wars (1920–1940). Historia Mathematica, 15(1):76–77, February 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900523.

Guitard:1987:EHI

[Gui87] Thierry Guitard. On an episode in the history of the integral calculus. Historia Mathematica, 14(2):215–219, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086087900395.

Guicciardini:1996:EHD

[Gui96] Niccolò Guicciardini. An episode in the history of dynamics: Jakob Hermann’s proof (1716–1717) of proposition 1, book 1, of Newton’s Principia. Historia Mathematica, 23(2):167–181, May 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086096900166.

Guicciardini:1999:LE

[Gui99] Niccolò Guicciardini. Letter to the Editor. Historia Mathematica, 26(3):292–294, August 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086099922425.

Guicciardini:2001:SCW

[Gui01] Niccolò Guicciardini. Squaring the circle. The war between Hobbes and Wallis. Historia Mathematica, 28(2):132–135, May 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086001923067.

Guicciardini:2003:CCR

[Gui03] Niccolo Guicciardini. Conceptualism and contextualism in the recent historiography of Newton’s Principia. Historia Mathematica, 30 (4):407–431, November 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S031508600300051X.

Guicciardini:2009:CDT

[Gui09a] Niccolò Guicciardini. Corrigendum to “Derek Thomas Whiteside (1932–2008)” [Historia Math. 36 (1) (2009) 4–9]. Historia Mathematica, 36(3):314, August 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086009000470. See [Gui09b].

Guicciardini:2009:DTW

[Gui09b] Niccolò Guicciardini. Derek Thomas Whiteside (1932–2008). Historia Mathematica, 36(1):4–9, February 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008000712.

See corrigendum [Gui09a].

Guicciardini:2024:HJM

[Gui24] Niccolò Guicciardini. Henk J. M. Bos (1940–2024): a first assessment of his legacy in the field of history of mathematics. Historia Mathematica, 68(??):40–48, September 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086024000727.

Gupta:1974:IFT

[Gup74] Radha Charan Gupta. An Indian form of third order Taylor series approximation of the sine. Historia Mathematica, 1(3):287–289, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074900676.

Gupta:1975:HMS

[Gup75a] R. C. Gupta. History of mathematics at a summer school in India. Historia Mathematica, 2(1):74, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075900464.

Gupta:1975:SJS

[Gup75b] R. C. Gupta. Seminar on Jaina Studies. Historia Mathematica, 2(1): 83, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508607590052X.

Gupta:1976:AAP

[Gup76a] R. C. Gupta. 1500th anniversary of Āryabhata I at Patna (India). Historia Mathematica, 3(4):468–469, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900781.

Gupta:1976:IDT

[Gup76b] R. C. Gupta. Indian doctoral theses on history of mathematics 1950– 1975. Historia Mathematica, 3(2):215, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086076900446.

Gupta:1977:BAA

[Gup77a] R. C. Gupta. 1500th birth anniversary of Āryabhata at Delhi. Historia Mathematica, 4(1):91–93, February 1977. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086077900416.

Gupta:1977:SJB

[Gup77b] R. C. Gupta. Silver jubilee of Bhārata Gaṅita Pariṡad. Historia Mathematica, 4(2):214–215, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086077901173.

Gupta:1977:PRC

[Gup77c] Radha Charan Gupta. Parame’svara’s rule for the circumradius of a cyclic quadrilateral. Historia Mathematica, 4(1):67–74, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900386.

Gupta:1978:ISH

[Gup78] R. C. Gupta. Indian society for the history of mathematics. Historia Mathematica, 5(4):465–466, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078902173.

Gupta:1979:SHP

[Gup79] R. C. Gupta. Symposium on history and philosophy of mathematical sciences. Historia Mathematica, 6(3):320–322, August 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079901319.

Gupta:1980:BDH

[Gup80] R. C. Gupta. Bibhutibhusan datta (1888–1958), historian of Indian mathematics. Historia Mathematica, 7(2):126–133, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900336.

Gupta:1981:FAC

[Gup81] R. C. Gupta. Fourth Annual Conference of the Indian Society for History of Mathematics. Historia Mathematica, 8(3):351–355, August 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900719.

Gupta:1985:HMI

[Gup85] R. C. Gupta. History of mathematics at the Indian science congress-1985. Historia Mathematica, 12(4):375–378, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086085900503.

Gupta:1986:SRI

[Gup86] R. C. Gupta. Symposium on the relevance and importance of the history of mathematics in present times in India. Historia Mathematica, 13(2):180–182, May 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508608690039X.

Gupta:1988:NSS

[Gup88] R. C. Gupta. National seminar on the scientific heritage of India. Historia Mathematica, 15(1):77–81, February 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900535.

Gupta:1991:SAM

[Gup91] R. C. Gupta. Seminar on astronomy and mathematics in Ancient and Medieval India: a dialogue between traditional scholars and university-trained scientists. Historia Mathematica, 18(1):56–62, February 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609190352X.

Gupta:1993:IDT

[Gup93] R. C. Gupta. Indian doctoral theses in the field of history of mathematics. Historia Mathematica, 20(3):310–314, August 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508608371027X.

Gurevitch:2024:BRM

[Gur24] Eric Moses Gurevitch. Book review: Misra, Anuj, 2022. Learning With Spheres: the Goladhyaya in Nityananda’s Sarvasiddhantaraja. Routledge, London 400 pp. DOI: 10.4324/9780429506680. Historia Mathematica, 69(??):62–64, December 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086024000879.

Goldstein:1994:BRA

[GW94a] Catherine Goldstein and Paul R. Wolfson. Book review: Amy Dahan Dalmedico, Editor, Mathématisations. Augustin-Louis Cauchy et l’école française (1993) 460 Argenteuil (Editions du Choix) and Paris (Albert Blanchard), 260 Ftr. Historia Mathematica, 21(1):88– 93, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Goldstein:1994:BRM

[GW94b] Catherine Goldstein and Paul R. Wolfson. Book review: Mathématisations. Augustin-Louis Cauchy et l’école française, Amy Dahan Dalmedico (Ed.) (1993), 460. Historia Mathematica, 21(1):88–93, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508608471010X.

Grigoryan:1977:BRA

[GY77] A. T. Grigoryan and A. P. Yushkevich. Boris A. Rosenfeld on the 60th anniversary of his birth. Historia Mathematica, 4(4):411–414, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086077900787.

Hecht:2008:KZB

[H^{+}08] Hartmut Hecht et al. Kosmos und Zahl: Beiträge zur Mathematikund Astronomiegeschichte, zu Alexander von Humboldt und Leibniz. (German) [Cosmos and number: contributions to mathematics and the history of astronomy, from Alexander von Humboldt and Leibniz], volume 58 of Boethius. Steiner, Stuttgart, Germany, 2008. ISBN 3-515-09176-9. 510 pp. LCCN Q125 .K667 2008.

Haffner:2016:BRB

[Haf16] Emmylou Haffner. Book review: Der Briefwechsel Richard Dedekind– Heinrich Weber, vol. 5, Katrin Scheel, in: Abhandlungen der Akademie der Wissenschaften in Hamburg. De Gruyter, Oldenbourg (2014). Historia Mathematica, 43(1):101–103, February 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086015000427.

Haffner:2017:SUA

[Haf17] Emmylou Haffner. Strategical use(s) of arithmetic in Richard Dedekind and Heinrich Weber’s Theorie der algebraischen Funktionen einer Veränderlichen. Historia Mathematica, 44(1):31–69, February 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086016300489.

Hahn:2011:CPS

[Hah11] Roger Hahn, editor. Correspondance de Pierre-Simon Laplace (1749– 1827), volume 87 of De Diversis Artibus. Brepols, Turnhout, Belgium, 2011. ISBN 2-503-54129-1. 1200 pp. LCCN Q124.6-127.2.

Hairetdinova:1986:STM

[Hai86] N. G. Hairetdinova. On spherical trigonometry in the medieval near east and in Europe. Historia Mathematica, 13(2):136–146, May 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086086900261.

Hall:1978:BRM

[Hal78] Tord Hall. Book review: Mathematisches Tagebuch, 1796–1814 von C. F. Gauss. Historia Mathematica, 5(4):486–487, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078902306.

Hall:1984:BRC

[Hal84] A. Rupert Hall. Book review: The collected letters of Colin Maclaurin: Edited by Stella Mills. 4 Church Lane, Nantwich, Cheshire, England (Shiva Pub. Ltd.). 1982. xx + 496 pp. Historia Mathematica, 11(2):237–239, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086084800245.

Halleux:1986:SST

[Hal86] Robert Halleux. Sluse et son temps. Historia Mathematica, 13 (3):278–281, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086086900959.

Halperin:1987:JNM

[Hal87] Israel Halperin. John von Neumann memorabilia. Historia Mathematica, 14(3):287, August 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086087900504.

Halperin:1990:LE

[Hal90] Israel Halperin. Letter to the editor. Historia Mathematica, 17(2): 186, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609090069P.

Hamadanizadeh:1980:TTA

[Ham80] Javad Hamadanizadeh. The trigonometric tables of al-Kashi in his Zij-i Khaqani. Historia Mathematica, 7(1):38–45, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900622.

Hannah:2007:FPL

[Han07] John Hannah. False position in Leonardo of Pisa’s Liber abbaci. Historia Mathematica, 34(3):306–332, August 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086006001297.

Harary:1977:BRG

[Har77a] Frank Harary. Book review: Graph theory 1736–1936: By N. L. Biggs, E. K. Lloyd, and R. J. Wilson. Oxford (Clarendon Press). 1976. 239 pp. $27.85. Historia Mathematica, 4(4):480–481, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900994.

Harary:1977:TII

[Har77b] Frank Harary. On “The Tantalizer” and “Instant Insanity”. Historia Mathematica, 4(2):205–206, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086077901136.

Hartshorne:2006:BRH

[Har06] Robin Hartshorne. Book review: Histoire des nombres complexes: Entre algèbre et géométrie, Dominique Flament. CNRS Editions, Paris (2003), ISBN 2-271-06128-8. Historia Mathematica, 33(4): 493–496, November 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S031508600600053X.

Hart:2011:CRL

[Har11] Roger (Roger Preston) Hart. The Chinese roots of linear algebra. The Johns Hopkins University Press, Baltimore, MD, USA, 2011. ISBN 0-8018-9755-6 (hardcover). xiii + 286 pp. LCCN QA27.C5 H37 2011.

Hashagen:2004:BRK

[Has04] Ulf Hashagen. Book review: Karl Culmann (1821–1881) und die graphische Statik: By Bertram Maurer. Berlin/Diepholz/Stuttgart (GNT-Verlag). 1998. 550 pp. ISBN 3-928186-41-8. EUR 40. Historia Mathematica, 31(3):369–371, August 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086004000230.

Hashagen:2010:HJN

[Has10a] Ulf Hashagen. Die Habilitation von John von Neumann an der Friedrich-Wilhelms-Universität in Berlin: Urteile über einen ungarisch-jüdischen Mathematiker in Deutschland im Jahr 1927. (German) [The habilitation of John von Neumann at the Friedrich Wilhelm University in Berlin: Judgments about a Hungarian Jewish mathematician in Germany in 1927]. Historia Mathematica, 37(2): 242–280, May 2010. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086009000317. See erratum [Has10b].

Hashagen:2010:EHJ

[Has10b] Ulf Hashagen. Erratum to “Die Habilitation von John von Neumann an der Friedrich-Wilhelms-Universität in Berlin: Urteile über einen ungarisch-jüdischen Mathematiker in Deutschland im Jahr 1927” [Historia Mathematica 37 (2) (2010) 242–280] [MR2609069]. Historia Mathematica, 37(4):723, November 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086010000352.

See [Has10a].

Hawkins:1975:CST

[Haw75a] Thomas Hawkins. Cauchy and the spectral theory of matrices. Historia Mathematica, 2(1):1–29, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075900324.

Hawkins:1975:MPF

[Haw75b] Thomas Hawkins. Mathematical progress without fusion. Historia Mathematica, 2(4):563–566, November 1975. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075901214.

Hawkins:1976:FEA

[Haw76] Thomas Hawkins. The Friedrich Engel archive in Giessen. Historia Mathematica, 3(2):213–214, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086076900434.

Hawkins:1977:BRS

[Haw77a] Thomas Hawkins. Book review: Sophus Lie’s 1880 transformation group paper: By Robert Hermann with translations by Michael Ackerman. Brookline, Mass. (Math. Sci. Press). 1975. 587 pp. US $28. Historia Mathematica, 4(2):229–230, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086077901276.

Hawkins:1977:SLD

[Haw77b] Thomas Hawkins. Sophus Lie’s 1884 differential invariant paper: By Robert Hermann with translations by Michael Ackerman. Brookline, Mass. (Math. Sci. Press). 1976. 281 pp. US $20. Historia Mathematica, 4(2):229–230, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086077901288.

Hawkins:1980:NEG

[Haw80] Thomas Hawkins. Non–Euclidean geometry and Weierstrassian mathematics: the background to Killing’s work on Lie algebras. Historia Mathematica, 7(3):289–342, August 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900270.

Hawkins:1981:BRG

[Haw81] Thomas Hawkins. Book review: Georg Cantor. His mathematics and philosophy of the infinite: By Joseph Warren Dauben. Cambridge, Mass./London (Harvard University Press). 1979. ix + 404 pp. $27.50. Historia Mathematica, 8(3):368–375, August 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900781.

Hawkins:1984:EPF

[Haw84] Thomas Hawkins. The Erlanger Programm of Felix Klein: Reflections on its place in the history of mathematics. Historia Mathematica, 11 (4):442–470, November 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086084900284.

Hawkins:1996:BRE

[Haw96] Thomas Hawkins. Book review: Élie Cartan (1869–1951). By M. A. Akivis and B. A. Rosenfeld. Translated from a Russian manuscript by V. V. Goldberg. Providence (American Mathematical Society). Historia Mathematica, 23(1):92–95, February 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086096900105.

Hawkins:2003:BRP

[Haw03] Thomas Hawkins. Book review: Pioneers of Representation Theory: Frobenius, Burnside, Schur, and Brauer: By Charles W. Curtis. History of Mathematics, Vol. 15. Providence, RI (American Mathematical Society). 1999. 287 pp. $49 (hardcover), $39 (soft-cover). Historia Mathematica, 30(1):91–93, February 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086002000125.

Hayashi:1987:VPM

[Hay87] Takao Hayashi. Varāhamihira’s pandiagonal magic square of the order four. Historia Mathematica, 14(2):159–166, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608790019X.

Hayashi:1997:ATR

[Hay97] Takao Hayashi. Aryabha t·a’s rule and table for sine-differences. Historia Mathematica, 24(4):396–406, November 1997. CODEN HI-MADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/S0315086097921601.

Hess:2011:GWF

[HB11] Heinz-Jürgen Hess and Malte-Ludolf Babin, editors. Gottfried Wilhelm Freiherr von Leibniz: Die mathematischen Zeitschriftenartikel. Georg Olms Verlag, Hildesheim, Germany, 2011. ISBN 3-487-14534-0. xviii + 547 pp. LCCN QA29.L46 A25 2011.

Hausdorff:2001:GWE

[HB^{+}13] Felix Hausdorff, Egbert Brieskorn, et al. Gesammelte Werke: einschliesslich der under dem Pseudonym Paul Mongré erschienenen philosophischen und literarischen Schriften und ausgewählter Texte aus dem Nachlass. (German) [Collected Works: including under the pseudonym Paul Mongré published philosophical and literary writings and selected texts from the estate]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001–2013. ISBN 3-540-42224-2 (v. 2), 3-540-41760-5 (v. 4), 3-540-30624-2 (v. 5), 3-540-20836-4 (v. 7è). ???? pp. LCCN QA3 .H315 2001.

Hedman:1999:EDC

[Hed99] Bruce A. Hedman. An earlier date for “Cramer’s Rule”. Historia Mathematica, 26(4):365–368, November 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086099922474.

Heeffer:2014:HAS

[Hee14] Albrecht Heeffer. How algebra spoiled recreational problems: a case study in the cross-cultural dissemination of mathematics. Historia Mathematica, 41(4):400–437, November 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086014000755.

Hilbert:2004:DHL

[HEH^{+}04] David Hilbert, William Bragg Ewald, Michael Hallett, Ulrich Majer, Wilfried Sieg, et al., editors. David Hilbert’s lectures on the foundations of mathematics and physics, 1891–1933. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-64373-7 (v. 1). ???? pp. LCCN QA7 .H53 2004.

Heibert:1981:BRS

[Hei81] Erwin Heibert. Book review: Sadi Carnot et l’essor de la thermodynamique: Table Ronde du Centre National de la Recherche Scientifique. Paris, École Polytechnique, 11–13 juin 1974. Paris (Éditions du Centre National de la Recherche Scientifique). 1976. 435 pp. Historia Mathematica, 8(3):375–377, August 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086081900793.

Heinzmann:2009:BRO

[Hei09] Gerhard Heinzmann. Book review: Oskar Becker und die Philosophie der Mathematik, Volker Peckhaus (Ed.). Wilhelm Fink Verlag, München (2005), 352, ISBN 3-7705-4126-X. Historia Mathematica, 36(2):189–191, May 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008001043.

Helena:1983:BRS

[Hel83] M. Pycior Helena. Book review: Sir William Rowan Hamilton: By Thomas L. Hankins. Baltimore/London (Johns Hopkins University Press). 1980. xxi + 474 pp. $32.50. Historia Mathematica, 10 (1):129–132, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086083900563.

Heller:2023:FKT

[Hel23] Henning Heller. Felix Klein’s teaching of Galois theory. Historia Mathematica, 63(??):21–46, May 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086023000162.

Hendy:1975:EFT

[Hen75] M. D. Hendy. Euclid and the fundamental theorem of arithmetic. Historia Mathematica, 2(2):189–191, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075901469.

Herivel:1975:BRJ

[Her75a] J. W. Herivel. Book review: Joseph Fourier 1768–1830: a Survey of His Life and Work Based on a Critical Edition of His Monograph on the Propagation of Heat, Presented to the Institut de France in 1807. By I. Grattan-Guinness in collaboration with J. R. Ravetz. Cambridge, Mass. & London (MIT Press). 1972 528 p. £9. Historia Mathematica, 2(2):222–223, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086075901597.

Hermelink:1975:ERB

[Her75b] Heinrich Hermelink. The earliest reckoning books existing in the Persian language. Historia Mathematica, 2(3):299–303, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900683.

Herbst:2014:BRG

[Her14] Alexander Herbst. Book review: Gottfried Wilhelm Leibniz: Die mathematischen Zeitschriftenartikel. Edited and translated by Heinz-Jürgen Heß and Malte-Ludolf Babin. Hildesheim (Georg Olms Verlag). 2011. ISBN 978-3-487-14534-1. 547 pp. 1 CD. 74 EUR. Historia Mathematica, 41(2):235–237, May 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086013000700.

Hess:1990:BRG

[Hes90] Heinz-Jürgen Hess. Book review: Die gesammelten Werke der Mathematiker und Physiker der Familie Bernoulli: Daniel Bernoulli: Die Werke. Band 3: Mechanik. Hrsg. von D. Speiser, A. de Baenst-Vandenbroucke, J. L. Pietenpol und R. Radelet-de Grave. Mit Beiträgen von H. Straub. Basel/Boston/Stuttgart (Birkhäuser). 1987. xxvii + 457 S. DM 230. Johann (I) Bernoulli: Der Briefwechsel. Band 2: Der Briefwechsel mit Pierre Varignon. Erster Teil: 1692–1702. Hrsg. von P. Costabel und J. Peiffer unter Benutzung von Vorarbeiten von J. O. Fleckenstein. Basel/Boston/Berlin (Birkhäuser). 1988. xviii + 442 S. DM 198. Historia Mathematica, 17(3):279–283, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086090900167.

Herz-Fischler:1984:WPE

[HF84] Roger Herz-Fischler. What are propositions 84 and 85 of Euclid’s data all about? Historia Mathematica, 11(1):86–91, February 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086084900089.

Herz-Fischler:2019:EUE

[HF19] Roger Herz-Fischler. An early usage of the expression “golden section”. Historia Mathematica, 49(??):80–81, November 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086019300217.

Hilton:1983:BRM

[Hil83] Peter Hilton. Book review: The mathematical experience: By Philip J. Davis and Reuben Hersh. Boston et al.: Birkhauser, 1981. xix + 440 pp. Historia Mathematica, 10(1):122–125, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608390054X.

Hirose:1974:BRC

[Hir74] Hideo Hirose. Book review: Chinese science: Exploration of an ancient tradition: Edited by Shigeru Nakayama and Nathan Sivin. Cambridge, Mass. (MIT Press). 1973. 378 p. $12.50. Historia Mathematica, 1(3):344–346, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074900834.

Hayashi:1998:TOA

[HK98] Takao Hayashi and Takanori Kusuba. Twenty-one algebraic normal forms of Citrabhanu. Historia Mathematica, 25(1):1–21, February 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086097921716.

Hollings:2017:LMM

[HMR17] Christopher Hollings, Ursula Martin, and Adrian Rice. The Lovelace– De Morgan mathematical correspondence: a critical re-appraisal. Historia Mathematica, 44(3):202–231, August 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086017300319.

Hurtado:2021:GOP

[HNE21] Ricardo Rodríguez Hurtado, Juan A. Nicolás, and Javier Echeverría Ezponda. The geometric origin of perspectivist science in G. W. Leibniz. Analysis based on unpublished manuscripts. Historia Mathematica, 55(??):1–22, May 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086020300902.

Hochkirchen:1996:BRC

[Hoc96] Thomas Hochkirchen. Book review: Creating Modern Probability. Its Mathematics, Physics and Philosophy in Historical Perspective. By Jan von Plato. Cambridge/New York/Melbourne (Cambridge Univ. Press). 1994. Historia Mathematica, 23(2):203–207, May 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086096900191.

Hodgson:1975:BRC

[Hod75] P. E. Hodgson. Book review: Causality and scientific explanation: By William A. Wallace. Vol. I. Medieval and Early Classical Science. Vol. II. Classical and Contemporary Science. University of Michigan Press. 1972/4. $13.20 and $14.00. Historia Mathematica, 2(1):97–100, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086075900592.

Hodes:1977:BRT

[Hod77] Elizabeth Hodes. Book review: Thinkers and tinkers: Early American men of science: By Silvio A. Bedini. New York (Charles Scribner’s Sons). 1975. xix + 520 pp. 102 plates. Historia Mathematica, 4(4):461–463, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508607790091X.

Hofmann:1987:WWD

[Hof87] Joseph Ehrenfried Hofmann. Wie ist wohl Demokrit zum Rauminhalt der Pyramide gekommen?. (German) [How did Democritus find the cubic volume of the pyramid?]. Historia Mathematica, 14(2):173– 174, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086087900218.

Hofmann:1992:GHG

[Hof92] Karl H. Hofmann. Zur Geschichte des Halbgruppenbegriffs. (German) [On the history of the concept of semigroups]. Historia Mathematica, 19(1):40–59, February 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609290054F.

Hogan:1974:BMH

[Hog74] Edward R. Hogan. The beginnings of mathematics in a howling wilderness. Historia Mathematica, 1(2):151–166, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074900044.

Hogan:1976:GBM

[Hog76] Edward R. Hogan. George Baron and the mathematical correspondent. Historia Mathematica, 3(4):403–415, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900689.

Hogan:1977:RAA

[Hog77] Edward R. Hogan. Robert Adrian: American mathematician. Historia Mathematica, 4(2):157–172, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086077901094.

Hogan:1981:TSA

[Hog81a] Edward R. Hogan. Theodore Strong and ante-bellum American mathematics. Historia Mathematica, 8(4):439–455, November 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900525.

Hogendijk:1981:HTA

[Hog81b] Jan P. Hogendijk. How trisections of the angle were transmitted from Greek to Islamic geometry. Historia Mathematica, 8(4):417–438, November 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086081900513.

Hogan:1985:MM

[Hog85a] Edward R. Hogan. The mathematical miscellany (1836–1839). Historia Mathematica, 12(3):245–257, August 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086085900229.

Hogendijk:1985:BRB

[Hog85b] Jan P. Hogendijk. Book review: Books IV to VII of Diophantus’ Arithmetica in the Arabic translation attributed to Qustā ibn Lūqā: By J. Sesiano. New York/Heidelberg/Berlin (Springer-Verlag). 1982. xii + 502 pp. Sources in the History of Mathematics and Physical Sciences, Vol. 3. $72.00. Historia Mathematica, 12(1):82–85, February 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086085900771.

Hogendijk:1985:TIQ

[Hog85c] Jan P. Hogendijk. Thābit ibn Qurra and the pair of amicable numbers 17296, 18416. Historia Mathematica, 12(3):269–273, August 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086085900254.

Hogendijk:1986:DCG

[Hog86] Jan P. Hogendijk. Discovery of an 11th-century geometrical compilation: the Istikmal of Yusuf al-M^{⊃} taman ibn Hud, King of Saragossa. Historia Mathematica, 13(1):43–52, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086086902272.

Hogendijk:1987:OIE

[Hog87] Jan P. Hogendijk. Observations on the icosahedron in Euclid’s Elements. Historia Mathematica, 14(2):175–177, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608790022X.

Hogendijk:1989:SAA

[Hog89] Jan P. Hogendijk. Sharaf al-Dīn al-Tūsī on the number of positive roots of cubic equations. Historia Mathematica, 16(1):69–85, February 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086089900992.

Hogan:1990:FAW

[Hog90] Edward R. Hogan. The fading amateur: William Lenhart and 19thcentury American mathematics. Historia Mathematica, 17(1):6–15, February 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609090075O.

Hogan:1991:PSA

[Hog91] Edward Hogan. “A proper spirit is abroad:” Peirce, Sylvester, Ward, and American mathematics, 1829–1843. Historia Mathematica, 18(2):158–172, May 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609190497L.

Hogendijk:1995:AMI

[Hog95] Jan P. Hogendijk. Al-Mu’taman ibn Hud, 11th Century King of Saragossa and brilliant mathematician. Historia Mathematica, 22 (1):1–18, February 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086085710014.

Hogendijk:2001:IIS

[Hog01a] Jan P. Hogendijk. Ibrahim ibn Sinan: logique et géométrie au X^{e} siècle. Historia Mathematica, 28(2):136–139, May 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086001923122.

Hogendijk:2001:IMA

[Hog01b] Jan P. Hogendijk. Islamic mathematics and astronomy. Historia Mathematica, 28(2):135–136, May 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086001923110.

Hogendijk:2002:SAB

[Hog02a] Jan P. Hogendijk. The surface area of the bicylinder and Archimedes’ method. Historia Mathematica, 29(2):199–203, May 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086002923499.

Hogendijk:2002:TEI

[Hog02b] Jan P. Hogendijk. Two editions of Ibn al-Haytham’s completion of the conics. Historia Mathematica, 29(3):247–265, August 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086002923529.

Hogendijk:2003:BRW

[Hog03] Jan P. Hogendijk. Book review: World-Maps for Finding the Direction and Distance to Mecca: Innovation and Tradition in Islamic Science: By David A. King. Leiden (Brill). 1999. ISBN 90-04-11363. xxix + 638 pp. Islamic Philosophy, Theology and Science, Vol. 36, Dfl 99,17, Euro 45, US $56. Historia Mathematica, 30(1):85–87, February 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086002000150.

Hogan:2008:HHB

[Hog08] Edward R. Hogan. Of the human heart: a biography of Benjamin Peirce. Lehigh University Press, Bethlehem, PA, USA, 2008. ISBN 0-934223-93-9. 429 pp. LCCN QA29.P66 .H56 2008. URL http://

www.loc.gov/catdir/toc/ecip0715/2007015983.html.

Hogendijk:2010:SFM

[Hog10] Jan P. Hogendijk. The scholar and the fencing master: the exchanges between Joseph Justus Scaliger and Ludolph van Ceulen on the circle quadrature (1594–1596). Historia Mathematica, 37(3):345–375, August 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086010000224.

Hogendijk:2015:YDS

[Hog15] Jan P. Hogendijk. Yvonne Dold-Samplonius (20 May 1937–16 June 2014). Historia Mathematica, 42(1):1–4, February 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086014000846.

Hogendijk:2024:BRJ

[Hog24] Jan P. Hogendijk. Book review: Julio Samsó, On both sides of the strait of Gibraltar: Studies in the history of medieval astronomy in the Iberian Peninsula and the Maghrib. Leiden: Brill, 2020. XXI + 1005 pp, ISBN 978-90-04-43656-5. Handbook of Oriental Studies, Section 1, The Near and Middle East, Vol. 144. Price 207 Euro. Historia Mathematica, 67(??):15, June 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086024000302.

Hollings:2015:AAA

[Hol15] Christopher Hollings. The acceptance of abstract algebra in the USSR, as viewed through periodic surveys of the progress of Soviet mathematical science. Historia Mathematica, 42(2):193–222, May 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086014000834.

Hormigon:1991:GGW

[Hor91] Mariano Hormigon. García de Galdeano’s works on algebra. Historia Mathematica, 18(1):1–15, February 1991. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086091903482.

Houser:1987:NKZ

[Hou87] Nathan Houser. A note on the Koenigsberger–Zeuner Repertorium. Historia Mathematica, 14(2):178–179, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086087900231.

Houzel:1989:CHM

[Hou89] Christian Houzel. Colloque d’Histoire des mathématiques sous le patronage de la société mathématique de France: Au Centre International de Rencontres Mathématiques, Marseille, du 17 au 22 juin 1985. Historia Mathematica, 16(2):169, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086089900426.

Howell:2018:MTG

[How18] Russell W. Howell. Book review: Mathematicians and Their Gods: Interactions Between Mathematics and Religious Beliefs, Snezana Lawrence, Mark McCartney (Eds.). Oxford University Press (2015), 320 pp. $44.00, ISBN: 978-0-19-870305-1. Historia Mathematica, 45(3):277–299, August 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086018301071.

Hoyrup:1982:IES

[Høy82] Jens Høyrup. Investigations of an early Sumerian division problem, c. 2500 B.C. Historia Mathematica, 9(1):19–36, February 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082901355.

Hoyrup:1983:SCI

[Høy83] Jens Høyrup. Sixth-century intuitive probability: the statistical significance of a miracle. Historia Mathematica, 10(1):80–84, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900368.

Hoyrup:1990:BRM

[Høy90a] Jens Høyrup. Book review: A mathematical history of division in extreme and mean ratio: By Roger Herz-Fischler. Waterloo, Ontario (Wilfrid Laurier Univ. Press). 1987. xvi + 191 pp. Historia Mathematica, 17(2):175–178, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609090065L.

Hoyrup:1990:BRW

[Høy90b] Jens Høyrup. Book review: Was ist ein Tonsystem? Eine historischsystematische Theorie der abendländischen Tonsysteme, gegründet auf die antiken Theoretiker Aristoxenos, Eukleides und Ptolemaios, dargestellt mit Mitteln der modernen Algebra: By Wilfried Neumaier. (Quellen und Studien zur Musikgeschichte von der Antike bis in die Gegenwart, Band 9). Frankfurt a.M. etc. (Peter Lang). 1986. Historia Mathematica, 17(2):172–174, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609090064K.

Hoyrup:1990:DBT

[Høy90c] Jens Høyrup. Dýnamis, the Babylonians, and Theaetetus 147c7– 148d7. Historia Mathematica, 17(3):201–222, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090001T.

Hoyrup:1993:RJY

[Høy93] Jens Høyrup. Reviewed by Jens Høyrup. Historia Mathematica, 20(1):99–103, February 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086083710104.

Hoyrup:1995:BRP

[Høy95] Jens Høyrup. Book review: Peri Ton Mathematon (Apeiron, Vol. XXIX, No. 4): Edited by Ian Mueller. Edmonton (Academic Printing and Publishing). 1991. xii + 251 pages. Index by name and subject. Index of passages cited. Historia Mathematica, 22(1):84–87, February 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086085710105.

Hoyrup:2002:LWS

[Høy02a] Jens Høyrup. Lengths, widths, surfaces: a portrait of old Babylonian algebra and its kin. Sources and studies in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 0-387-95303-5. xiv + 459 pp. LCCN QA22 .H83 2002. URL http://www.loc.

gov/catdir/enhancements/fy0816/2001032839-d.html;   http:
//www.loc.gov/catdir/toc/fy022/2001032839.html.

Hoyrup:2002:NOB

[Høy02b] Jens Høyrup. A note on old Babylonian computational techniques. Historia Mathematica, 29(2):193–198, May 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086002923438.

Hoyrup:2004:CDC

[Høy04] Jens Høyrup. Conceptual divergence-canons and taboos-and critique: reflections on explanatory categories. Historia Mathematica, 31(2):129–147, May 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086003000247.

Hoyrup:2006:JFB

[Høy06] Jens Høyrup. Jacopo da Firenze and the beginning of Italian vernacular algebra. Historia Mathematica, 33(1):4–42, February 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005000340.

Hoyrup:2007:BRL

[Høy07] Jens Høyrup. Book review: L’Algèbre arabe. Genèse d’un art By Ahmed Djebbar. Préface de Bernard Maitte. Paris (≪Inflexions≫, collection dirigée par Jean Rosmorduc; Vuibert/ADAPT-SNES). 2005. ISBN 2 7117 5381 6 (Vuibert), 2 9096 8065 7 (ADAPT). viii + 212 pp. Historia Mathematica, 34(2):225–227, May 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086006000814.

Hoyrup:2009:BRA

[Høy09a] Jens Høyrup. Book review: Arithmetische und algebraische Schriften der frühen Neuzeit, vol. 17, Rainer Gebhardt (Ed.), in:

Schriften des Adam-Ries-Bundes Annaberg-Buchholz. Adam-Ries-Bund, Annaberg-Buchholz (2005), ISBN 3-930430-68-1. Historia Mathematica, 36(3):278–279, August 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086009000020.

Hoyrup:2009:BRF

[Høy09b] Jens Høyrup. Book review: Fibonacci’s De Practica Geometrie, Barnabas Hughes (Ed.), in: Sources and Studies in the History of Mathematics and Physical Sciences. Springer, New York (2008), ISBN 978-0-387-72930-5. Historia Mathematica, 36(1):73–77, February 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008000566.

Hoyrup:2023:BRA

[Høy23] Jens Høyrup. Book review: Adam Ries, Coß 1. 2 vols., Bernd Rüdiger, Gebhardt Rainer, Menso Folkerts (Eds.). Adam-Ries-Bund, Annaberg-Buchholz (2023), 320 + 278 pp. Historia Mathematica, 64(??):48–56, August 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086023000356.

Hoyrup:2024:JRA

[Høy24a] Jens Høyrup. Johannes Regiomontanus, Aufgaben und Übungen zur Algebrass, Menso Folkerts, Martin Hellmann. Erwin Rauner Verlag, Augsburg (2023), Essay review byed., trans. Historia Mathematica, 66(??):43–46, March 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086023000691.

Hoyrup:2024:RMM

[Høy24b] Jens Høyrup. Review of: Marc Moyon, Le Liber augmenti et diminutionis: Contribution à l’histoire des mathématiques médiévales. XVI, 204 pp. (Boethius, 73). Stuttgart: Franz Steiner Verlag. ISBN 978-3-515-13573-3 (print), ISBN 978-3-515-13581-8 (E-Book). Historia Mathematica, 67(??):16–19, June 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086024000089.

Hausdorff:2005:HOS

[HP05] Felix Hausdorff and J. M. (Jacob M.) Plotkin, editors. Hausdorff on ordered sets, volume 25 of History of mathematics, 0899-2428.

American Mathematical Society, Providence, RI, USA, 2005. ISBN 0-8218-3788-5. xviii + 322 pp. LCCN QA171.48 .H38 2005.

Hollings:2020:TLO

[HP20] Christopher D. Hollings and R. B. Parkinson. Two letters from Otto Neugebauer to Thomas Eric Peet on ancient Egyptian mathematics. Historia Mathematica, 52(??):66–98, August 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086020300434.

Hug:2015:LEA

[HS15] Vanja Hug and Thomas Steiner. Une lettre d’Euler à d’Alembert retrouvée. Historia Mathematica, 42(1):84–94, February 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086014000822.

Hudecek:2012:BRC

[Hud12] Jiri Hudecek. Book review: The Chinese Roots of Linear Algebra. By Roger Hart. Baltimore, (Johns Hopkins University Press). 2011. ISBN 978-0-8018-9755-9. xiii + 286 pp. US$65 (hardcover). Historia Mathematica, 39(2):226–231, May 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508601100067X.

Hughes:1974:EKR

[Hug74] Barnabas B. Hughes. The earliest known record of California Indian numbers. Historia Mathematica, 1(1):79–82, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074901724.

Hughes:1975:CHM

[Hug75] Barnabas Hughes. A course in history of mathematics for mathematics teachers. Historia Mathematica, 2(3):333–335, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900786.

Hughes:1976:BRH

[Hug76] Barnabas Hughes. Book review: History of mathematics: By Arthur Gittleman. Columbus, Ohio (Charles E. Merrill). 1975. 291 p. US$14.95. Historia Mathematica, 3(3):359–361, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508607690121X.

Hughes:1987:ECA

[Hug87] Barnabas Hughes. An early 15th-century algebra codex: a description. Historia Mathematica, 14(2):167–172, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086087900206.

Hughes:1989:ATJ

[Hug89] Barnabas Hughes. The arithmetical triangle of Jordanus de Nemore. Historia Mathematica, 16(3):213–223, August 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086089900189.

Hughes:2005:BRM

[Hug05] Barnabas Hughes. Book review: Maestro Dardi (sec. xiv) Aliabraa Argibra dal manoscrito I.VII.17 della Biblioteca Comunale di Siena: Quaderno del Centro Studi della Matematica Medioevale, No. 26. Edited by Raffaelle Franci. Siena (Centro Studi della Matematica Medioevale). 2001. 299 pp. Gratis. Historia Mathematica, 32(1):97– 98, February 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086003000478.

Hughes:2007:BRG

[Hug07] Barnabas B. Hughes. Book review: Giochi matematici alla corte di Carlomagno. Problemi per rendere acuta la mente dei giovani, Raffaella Franci. Edizioni ETS, Pisa (2005), ISBN 88-467-1351-6. Historia Mathematica, 34(3):344–345, August 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086006000954.

Hughes:2010:EAF

[Hug10] Barnabas Hughes. An early abridgement of Fibonacci’s de practica geometrie. Historia Mathematica, 37(4):615–640, November 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009000743.

Hunter:1996:DBM

[Hun96] Patti W. Hunter. Drawing the boundaries: Mathematical statistics in 20th-Century America. Historia Mathematica, 23(1):7–30, February 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086096900026.

Hunter:2007:BRK

[Hun07] Patti W. Hunter. Book review: Karl Pearson: the Scientific Life in a Statistical Age, Theodore M. Porter. Princeton University Press, Princeton, NJ (2004), ISBN 0-691-11445-5. Historia Mathematica, 34(2):233–236, May 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086006000838.

Hunter:2012:BRH

[Hun12] Patti W. Hunter. Book review: How to Read Historical Mathematics. By Benjamin Wardhaugh. Princeton and Oxford (Princeton University Press). 2010. ISBN 978-0-691-14014-8. 113 pp. US$23.95. Historia Mathematica, 39(2):232–233, May 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086011000681.

Hyman:1978:BRA

[Hym78] Anthony Hyman. Book review: Ada, Countess of Lovelace: Byron’s legitimate daughter: By Doris Langley Moore. New York (Harper & Row). 1977. $25.00. Historia Mathematica, 5(3):366–367, August 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078901283.

Hamadani-Zadeh:1985:SOI

[HZ85] Javad Hamadani-Zadeh. A second-order interpolation scheme described in the Zij-i Ilkhani. Historia Mathematica, 12(1):56–59, February 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086085900680.

Israel:1998:ETW

[IM98] Giorgio Israel and Marta Menghini. The “Essential Tension” at work in qualitative analysis: a case study of the opposite points of view of Poincaré and Enriques on the relationships between analysis and geometry. Historia Mathematica, 25(4):379–411, November 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508609892211X.

Imhausen:2003:AAU

[Imh03a] Annette Imhausen. Ägyptische Algorithmen: eine Untersuchung zu den mittelägyptischen mathematischen Aufgabentexten. (German) [Egyptian algorithms: a study of the Middle Egyptian mathematical problem texts], volume 65 of èAgyptologische Abhandlungen, 0568-0476. Harrassowitz, Wiesbaden, Germany, 2003. ISBN 3-447-04644-9 (paperback). xi + 387 pp. LCCN QA27.E3 I44 2003.

Imhausen:2003:CDB

[Imh03b] Annette Imhausen. Calculating the daily bread: Rations in theory and practice. Historia Mathematica, 30(1):3–16, February 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086002000095.

Imhausen:2021:QVH

[Imh21] Annette Imhausen. Quo vadis History of Ancient Mathematics who will you take with you, and who will be left behind? Essay review prompted by a recent publication. Historia Mathematica, 57 (??):80–93, November 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086021000197.

Israel:1983:CMR

[IN83] Giorgio Israel and Laura Nurzia. Correspondence and manuscripts recovered at the Istituto Matematico “G. Castelnuovo” of the University of Rome. Historia Mathematica, 10(1):93–97, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900411.

Israel:1984:PVT

[IN84] Giorgio Israel and Laura Nurzia. The Poincaré-Volterra theorem: a significant event in the history of the theory of analytic functions. Historia Mathematica, 11(2):161–192, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086084800075.

Indorato:1989:RFD

[IN89] Luigi Indorato and Pietro Nastasi. The 1740 resolution of the Fermat– Descartes controversy. Historia Mathematica, 16(2):137–148, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086089900359.

Ineichen:1988:DKV

[Ine88] Robert Ineichen. Dante-Kommentare und die Vorgeschichte der Stochastik. (German) [Dante commentaries and the prehistory of stochastics]. Historia Mathematica, 15(3):264–269, August 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900663.

Ineichen:1991:SWS

[Ine91] Robert Ineichen. Der schlechte Würfel — Ein selten behandeltes Problem in der Geschichte der Stochastik. (German) [The bad die — a rarely treated problem in the history of stochastics]. Historia Mathematica, 18(3):253–261, August 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609190394D.

Ineichen:1994:VCL

[Ine94] Robert Ineichen. Der “Vierfeldertest” von Carl Liebermeister (Bemerkungen zur Entwicklung der medizinischen Statistik im 19. Jahrhundert). (German) [The “Four-Fields Test” of Carl Liebermeister (comments on the development of medical statistics in the 19th Century)]. Historia Mathematica, 21(1):28–38, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086084710068.

Ingaliso:2011:MPA

[Ing11] Luigi Ingaliso. Mario Pieri’s address at the University of Catania. Historia Mathematica, 38(2):232–247, May 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086010000327.

Inonu:2006:MNA

[Inö06] Erdal Inönü. Mehmet Nadir: an amateur mathematician in Ottoman Turkey. Historia Mathematica, 33(2):234–242, May 2006.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005000698.

Oya:1975:BRM

[iÔ75] Shin ichi Ôya. Book review: Ming Kan no Syuzan Syo [a]: [Chinese Abacus Books Published in the Ming Age] By Akihito Kodama. Part 1: On the Chinese abacus books published in the Ming age (in Japanese), 115 p. Part 2: Chinese abacus books published in the last half of the 16th century of the Ming age (in Chinese), photocopy, 127 p. Tokyo (Fuji Junior College Press). 1972 242 p. = Y1500. Historia Mathematica, 2(2):218–219, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086075901573.

Israel:1985:BRI

[Isr85] Giorgio Israel. Book review: II Circolo matematico di Palermo: By Aldo Brigaglia and Guido Masotto. Bari (Edizioni Dedalo). 1982. 444 pp. Lire It. 10,000. Historia Mathematica, 12(4):383–388, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086085900539.

Israel:1990:CBB

[Isr90] Giorgio Israel. On correspondence between B. Boncompagni and A. Genocchi. Historia Mathematica, 17(1):48–54, February 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090078R.

Jaeck:2019:CDI

[Jaë19] Frédéric Jaëck. Calcul différentiel et intégral adapté aux substitutions par Volterra. (French) [Differential and integral calculation for substitutions by Volterra]. Historia Mathematica, 48(??):29–68, August 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086018300582.

Jahnke:1993:AAG

[Jah93] Hans Niels Jahnke. Algebraic analysis in Germany, 1780–1840: Some mathematical and philosophical issues. Historia Mathematica, 20 (3):265–284, August 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086083710232.

Jahnke:2012:BRC

[Jah12] Hans Niels Jahnke. Book review: Cauchy’s Cours d’analyse. An Annotated Translation. By Robert E. Bradley, and C. Edward Sandifer. Sources and Studies in the History of Mathematics and Physical Sciences. Dordrecht (Springer). 2009. Historia Mathematica, 39 (4):467–469, November 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086012000146.

Jami:1988:WIC

[Jam88] Catherine Jami. Western influence and Chinese tradition in an Eighteenth Century Chinese mathematical work. Historia Mathematica, 15(4):311–331, November 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086088900249.

Jami:1997:MJN

[Jam97] Catherine Jami. In memoriam: Joseph Needham (December 9, 1900–March 24, 1995). Historia Mathematica, 24(1):1–5, February 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086097921261.

James:2013:BRR

[Jam13] Kathryn James. Book review: Robert Recorde: the Life and Times of a Tudor Mathematician, Gareth Roberts, Fenny Smith (Eds.). University of Wales Press, distributed by University of Chicago Press, Swansea (2012), ISBN: 978-0-7083-2526-1. Historia Mathematica, 40(4):455–457, October 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086013000220.

Jankvist:2011:CME

[Jan11] Uffe Thomas Jankvist. A century of mathematics education: ICMI’s first hundred years. Historia Mathematica, 38(2):292–302, May 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086010000066.

Jayawardene:1983:BRH

[Jay83] S. A. Jayawardene. Book review: Über die Handschrift CLM 26639 der bayerischen Staatsbibliothek Müchen: Eine mögliche Quelle zu Widmanns deutschem Rechenbuch von 1489. Ein Beitrag zur Geschichte der Mathematik im ausgehenden Mittelalter. By Wolfgang Kaunzner. (Arbor scientiarum, Beiträge zur Wissenschaftsgeschichte, Reihe B: Texte, Band I). Hildesheim (Gerstenberg Verlag). 1978. vii + 125 pp. 114 DM. Historia Mathematica, 10(4):478–479, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900186.

Jayawardana:2024:MCR

[Jay24a] Chandana Jayawardana. Mathematical cognition related to the large numbers in early societies: a study based on 5th-century Buddhist Commentaries in Sri Lanka. Historia Mathematica, 67(??):1–12, June 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/

pii/S0315086024000041. See withdrawal notice [Jay24b].

Jayawardana:2024:WNM

[Jay24b] Chandana Jayawardana. Withdrawal notice to: Mathematical cognition related to the large numbers in early societies: a study based on 5th-century Buddhist Commentaries in Sri Lanka. [Hist. Math. 67 (2024) 1–12]. Historia Mathematica, 69(??):71, December 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086024000776. See [Jay24a].

Jentsch:1976:HMD

[Jen76] Werner Jentsch. Über ein hallenser manuskript der dissertation Georg Cantors. (German) [On a Halle manuscript of the dissertation of Georg Cantor]. Historia Mathematica, 3(4):449–462, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900768.

Jentsch:1986:AUK

[Jen86] Werner Jentsch. Auszüge aus einer unveröffentlichten Korrespondenz von Emmy Noether und Hermann Weyl mit Heinrich Brandt. (German) [Excerpts from unpublished correspondence of Emmy Noether and Hermann Weyl with Henry Brandt]. Historia Mathematica, 13(1):5–12, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086086902211.

Jones:1984:KMH

[JET84] Charles V. Jones, Philip C. Enros, and Henry S. Tropp. Kenneth O. May, 1915–1977: His early life to 1946. Historia Mathematica, 11(4):359–379, November 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086084900223.

Jozwik:2019:SK

[JMT19] Izabela Jó’zwik, Lech Maligranda, and Małgorzata Terepeta. Stefan Kempisty (1892–1940). Historia Mathematica, 48(??):69–86, August 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086018301551.

Johnson:1977:CCJ

[Joh77] Dale M. Johnson. The correspondence of Camille Jordan. Historia Mathematica, 4(2):210, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508607790115X.

Johnsen:1982:BTC

[Joh82] Karsten Johnsen. Bemerkungen zu einer Tagebuchnotiz von Carl Friedrich Gauss. (German) [Comments on a diary of Carl Friedrich Gauss]. Historia Mathematica, 9(2):191–194, May 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082900039.

Johnsen:1984:BCF

[Joh84] Karsten Johnsen. Zum Beweis von C. F. Gauss für die irreduzibilität des p-ten kreisteilungspolynoms. (German) [On the proof of C. F. Gauss of the irreducibility index of the p-th circular distribution polynomials]. Historia Mathematica, 11(2):131–141, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508608480004X.

Johnsen:1986:RTE

[Joh86] Karsten Johnsen. Remarks to the third entry in Gauss’ diary. Historia Mathematica, 13(2):168–169, May 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086086900303.

Johansson:2011:CRE

[Joh11] Bo Göran Johansson. Cube root extraction in medieval mathematics. Historia Mathematica, 38(3):338–367, August 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086010000546.

Jones:1975:ROB

[Jon75a] Charles V. Jones. Reviews of old books. Historia Mathematica, 2(2): 199, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086075901500.

Jones:1975:RHM

[Jon75b] Phillip S. Jones. A role for the history of mathematics. Historia Mathematica, 2(4):471–472, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075901068.

Jones:1976:BRM

[Jon76a] Phillip S. Jones. Book review: Mathematics and mathematicians: By P. Dedron and J. Itard, translated by J. V. Field. London (Transworld Publishers Ltd.) 1974. Vol. I, 326 p.; Vol. II, 222 p. £1.50, £1.00. Historia Mathematica, 3(3):357–359, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086076901208.

Jones:1976:LCK

[Jon76b] Phillip S. Jones. Louis Charles Karpinski, historian of mathematics and cartography. Historia Mathematica, 3(2):185–202, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900367.

Jones:1976:BD

[Jon76c] Phillips S. Jones. Bibhutibhusan datta. Historia Mathematica, 3 (1):77–78, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076900124.

Jones:1977:AMC

[Jon77a] Charles V. Jones. Annual meeting of the Canadian society for the history and philosophy of Mathematics/Société canadienne d’Histoire et de philosophie des Mathématiques. Historia Mathematica, 4(3):353– 354, August 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086077900696.

Jones:1977:MTC

[Jon77b] Charles V. Jones. A missing table in C. S. Peirce’s New Elements of Mathematics, vol. 1. Historia Mathematica, 4(4):441–442, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900805.

Jones:1977:BRR

[Jon77c] Phillip S. Jones. Book review: The Renaissance rediscovery of linear perspective: By Samuel Y. Edgerton, Jr. New York (Basic Books, Inc.). 1975. xvii + 206 pp. $16.50. Historia Mathematica, 4(4):472–474, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086077900969.

Jones:1981:AMH

[Jon81] Charles V. Jones. Annual meeting of the History of Science Society, 1980. Historia Mathematica, 8(2):192–194, May 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900318.

Jongsma:1985:BRB

[Jon85] Calvin Jongsma. Book review: The Boole–de Morgan correspondence: 1842–1864: By Gordon C. Smith. London, New York (Oxford University Press (Clarendon)). 1982. 156 pp. $44.00. Historia Mathematica, 12(2):186–190, May 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086085900151.

Jordan:1981:AAS

[Jor81] Steven L. Jordan. American attempts to solve Newton’s pasturage problem. Historia Mathematica, 8(2):145–160, May 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900264.

Jurkowitz:2007:BRA

[Jur07] Edward Jurkowitz. Book review: Arnold Sommerfeld. Wissenschaftlicher Briefwechsel. Band I: 1892–1918, Band II: 1919– 1951, CD-ROM: 1892–1951, Michael Eckert, Karl Märker (Eds.). Verlag für Geschichte der Naturwissenschaften und der Technik, Diepholz (2000) (2004), ISBN 978-3-928186-44-5. Historia Mathematica, 34(4):446–448, November 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086007000316.

Kahane:1975:COM

[Kah75] J. P. Kahane. A central object in mathematics: the Brownian motion. Historia Mathematica, 2(4):595–600, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075901275.

Kidwell:2008:TAM

[KAHR08] Peggy Aldrich Kidwell, Amy Ackerberg-Hastings, and David Lindsay Roberts. Tools of American mathematics teaching, 1800– 2000. Johns Hopkins studies in the history of mathematics. The Johns Hopkins University Press, Baltimore, MD, USA, 2008. ISBN 0-8018-8814-X (hardcover). xviii + 418 pp.

LCCN QA71 .K46 2008. URL http://www.loc.gov/catdir/
enhancements/fy0805/2007040065-b.html;   http://www.loc.
gov/catdir/enhancements/fy0805/2007040065-d.html;   http:
//www.loc.gov/catdir/toc/ecip081/2007040065.html.

Kangsheng:1988:MSA

[Kan88] Shen Kangsheng. Mutual-subtraction algorithm and its applications in ancient China. Historia Mathematica, 15(2):135–147, May 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900031.

Karpinska:2022:DNM

[Kar22] Karolina Karpińska. “Denominate numbers” in mathematics school textbooks by Stefan Banach. Historia Mathematica, 59(??):164– 196, May 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508602100001X.

Kastanis:1990:FGC

[Kas90a] Nikos Kastanis. The first Greek conference on history and philosophy of mathematics-theme: History and philosophy of ancient Greek mathematics: Athens, March 2–4, 1989. Historia Mathematica, 17(1):65–66, February 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609090080W.

Kastanis:1990:FRM

[Kas90b] Nikos Kastanis. French revolution and mathematics: Thessaloniki, November 8–10, 1989. Historia Mathematica, 17(3):267, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090010B.

Kastanis:2003:TMG

[Kas03] Andreas Kastanis. The teaching of mathematics in the Greek Military Academy during the first years of its foundation (1828–1834). Historia Mathematica, 30(2):123–139, May 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508600200023X.

Katz:1981:HDF

[Kat81] Victor J. Katz. The history of differential forms from Clairaut to Poincaré. Historia Mathematica, 8(2):161–188, May 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900276.

Katz:1987:CTF

[Kat87] Victor J. Katz. The calculus of the trigonometric functions. Historia Mathematica, 14(4):311–324, November 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086087900644.

Katz:1992:BRC

[Kat92] Victor J. Katz. Book review: The crest of the peacock: Non-European roots of mathematics: By George Gheverghese Joseph. London (I. B. Tauris & Co., Ltd) and New York (St. Martin’s Press). 1991. xvi + 368 pp. including Bibliography and Index. Historia Mathematica, 19(3):310–315, August 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609290042A.

Katz:1997:CEH

[Kat97] Victor J. Katz. Companion encyclopedia of the history and philosophy of the mathematical sciences. Historia Mathematica, 24 (4):445–452, November 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086097921613.

Katz:2007:BRC

[Kat07] Victor J. Katz. Book review: From Calculus to Computers: Using the Last 200 Years of Mathematics History in the Classroom, Amy Shell-Gellasch, Dick Jardine (Eds.). The Mathematical Association of America, Washington, DC (2005), ISBN 0-88385-178-4. Historia Mathematica, 34(2):236–238, May 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086006000826.

Katz:2009:MES

[Kat09] Victor J. Katz. Mathematics emerging: a sourcebook 1540– 1900. Historia Mathematica, 36(4):433–436, November 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009000561.

Katz:2011:BRD

[Kat11] Victor J. Katz. Book review: “Dig Where You Stand”: Proceedings of a Conference on On-going Research in the History of Mathematics Education, Gardabaer, Iceland, June 20–24, 2009. Edited by Kristín Bjarnadóttir, Fulvia Furinghetti and Gert Schubring. Reykjavik (University of Iceland). 2009. ISBN 978-9979-793-99-1. vi + 254 pp. No price given. Historia Mathematica, 38(2):303–308, May 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086011000036.

Katz:2014:MAA

[Kat14] Victor J. Katz. Marcia Ascher (23 April 1935–11 June 2013). Historia Mathematica, 41(1):3–5, February 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086014000056.

Katz:2020:BRM

[Kat20] Victor J. Katz. Book review: The Mathematical World of Charles L. Dodgson, Robin Wilson, Amirouche Moktefi (Eds.). Oxford University Press (2019), 266 + xiv pages. Historia Mathematica, 53 (??):128–131, November 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086020300033.

Knobloch:1974:MSG

[KB74] Eberhard Knobloch and West Berlin. The mathematical studies of G. W. Leibniz on combinatorics. Historia Mathematica, 1(4):409–430, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074900329.

Kennedy:1974:LAB

[Kc74] E. S. Kennedy and Yusuf ^{c}Id. A letter of al-bīrūnī H*abash al-H*āsib’s analemma for the qibla. Historia Mathematica, 1(1):3–11, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508607490161X.

Kolmogorov:1996:MXV

[Kë96] A. N. (Andreĭ Nikolaevich) Kolmogorov and A. P. (Adol§f Pavlovich) ëIìUshkevich, editors. Matematika XIX veka. Mathematics of the 19th Century: geometry, analytic function theory. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1996. ISBN 0-8176-5048-2 (Boston), 3-7643-5048-2 (Basel). x + 291 pp. LCCN QA443.5

.M3813 1996. URL http://www.loc.gov/catdir/enhancements/
fy0813/95052794-d.html.

Keller:2005:MDS

[Kel05] Agathe Keller. Making diagrams speak, in Bhāskara I’s commentary on the book review: Aryabhat.īya. Historia Mathematica, 32(3):275– 302, August 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508600400062X.

Kennedy:1974:PCN

[Ken74a] Hubert C. Kennedy. Peano’s concept of number. Historia Mathematica, 1(4):387–408, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074900317.

Kennedy:1974:PBG

[Ken74b] Hubert C. Kennedy. A prospective biography of Giuseppe Peano (1858–1932). Historia Mathematica, 1(1):87–89, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074901785.

Kenenedy:1975:BRI

[Ken75] E. S. Kenenedy. Book review: Iz istorii tochnykh nauk na srednevekovom blizhnem i srednem vostoke: [From the History of the Exact Sciences in the Medieval Near and Middle East], By S. K. Sirazhdinov. Tashkent (Academy of Sciences of the Uzbek S.S.R.). 1972. 247 p. Historia Mathematica, 2(1):86–90, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900555.

Kendall:1976:BRH

[Ken76a] Maurice Kendall. Book review: The heritage of Copernicus: Theories “pleasing to the mind”: Edited by Jerzy Neyman. Cambridge, Mass. and London (The M. I. T. Press). 552 p. Historia Mathematica, 3(1):95–96, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076900227.

Kennedy:1976:BRKb

[Ken76b] H. C. Kennedy. Book review: Karl Marx. Manoscritti matematici: Translated and edited by Francesco Matarrese and Augusto Ponzio. Bari, Italy (Dedalo Libri). 1975. 184 p. 3,000 lire. Historia Mathematica, 3(4):490–494, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086076900896.

Kennedy:1976:BRKa

[Ken76c] H. C. Kennedy. Book review: Karl Marx. Mathematische manuskripte: Edited, with an introduction and commentary, by Wolfgang Endemann. Kronberg Taunus, BRD (Spriptor Verlag). 1974. 178 p. Historia Mathematica, 3(4):490, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900884.

Kennedy:1976:BRI

[Ken76d] Hubert C. Kennedy. Book review: Istoriya matematiki s drevneishikh vremen do nachala XIX stoletiya: By A. P. Yushkevich et al. Tom tretii: Matematika XVIII stoletiya [History of Mathematics from Ancient Times to the Beginning of the 19th Century. Volume 3: Mathematics of the 18th century]. Moscow (Nauka). 1972. 495 p. Historia Mathematica, 3(1):87–90, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086076900185.

Kennedy:1977:KMF

[Ken77] Hubert C. Kennedy. Karl Marx and the foundations of differential calculus. Historia Mathematica, 4(3):303–318, August 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900581.

Kennedy:1978:C

[Ken78a] Hubert Kennedy. Correction. Historia Mathematica, 5(1):92, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086078901428.

Kennedy:1978:BRM

[Ken78b] Hubert C. Kennedy. Book review: La matematica come sistema ipotetico-deduttivo, profilo storico: by Silvio Maracchia. Firenze (Felice Le Monnier). 1975. 219 pp. L. 6000. Historia Mathematica, 5(2):237–239, May 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086078900629.

Kennedy:1979:JMP

[Ken79a] Hubert Kennedy. James Mills Peirce and the cult of quaternions. Historia Mathematica, 6(4):423–429, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079900284.

Kennedy:1979:TBJ

[Ken79b] Hubert Kennedy. Towards a biography of James mills Peirce. Historia Mathematica, 6(2):196–201, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508607990079X.

Kennedy:1982:BRM

[Ken82] Hubert Kennedy. Book review: Matematica: Un profilo storico: By Dirk J. Struik. With an appendix by Umberto Bottazzini. Translated by Umberto Bottazzini and Virginio Sella. Bologna (Il Mulino), 1981. 363 pp. L. 12,000. Historia Mathematica, 9(4):495–496, November 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608290115X.

Kennedy:1984:AMT

[Ken84] E. S. Kennedy. Applied mathematics in the Tenth Century: Abu’lwafa’ calculates the distance Baghdad–Mecca. Historia Mathematica, 11(2):193–206, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086084800087.

Kent:2008:MMC

[Ken08] Deborah Kent. The Mathematical Miscellany and The Cambridge Miscellany of Mathematics: Closely connected attempts to introduce research-level mathematics in America, 1836–1843. Historia Mathematica, 35(2):102–122, May 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086007001012.

Ketner:1982:CEP

[Ket82] Kenneth Laine Ketner. Carolyn Eisele’s place in Peirce studies. Historia Mathematica, 9(3):326–332, August 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086082901264.

Kichenassamy:2006:BRQ

[Kic06] Satyanad Kichenassamy. Baudhayana’s rule for the quadrature of the circle. Historia Mathematica, 33(2):149–183, May 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005000479.

Kichenassamy:2010:BDA

[Kic10] Satyanad Kichenassamy. Brahmagupta’s derivation of the area of a cyclic quadrilateral. Historia Mathematica, 37(1):28–61, February 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009000688.

Kichenassamy:2012:BPP

[Kic12] Satyanad Kichenassamy. Brahmagupta’s propositions on the perpendiculars of cyclic quadrilaterals. Historia Mathematica, 39(4): 387–404, November 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086012000559.

Kichenassamy:2015:CPT

[Kic15] Satyanad Kichenassamy. Continued proportions and Tartaglia’s solution of cubic equations. Historia Mathematica, 42(4):407–435, November 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086015000403.

Kidwell:2009:CDM

[Kid09] Peggy Aldrich Kidwell. Computing devices, mathematics education and mathematics: Sexton’s omnimetre in its time. Historia Mathematica, 36(4):395–404, November 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086009000299.

Kilpatrick:1983:BRM

[Kil83] Jeremy Kilpatrick. Book review: Mathematical discoveries: 1600– 1750: By P. L. Griffiths. Ilfracombe, England (Stockwell). 1977. 121 pp. £2.75. Historia Mathematica, 10(1):117–118, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900526.

Kimberling:1975:MAM

[Kim75] Clark Kimberling. Matheematical aspects of music. Historia Mathematica, 2(1):65–66, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075900403.

Kimble:1979:PP

[Kim79] Gerald W. Kimble. From pebbles to position. Historia Mathematica, 6(3):323–325, August 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086079901344.

King:1974:MIM

[Kin74a] David A. King. On medieval Islamic multiplication tables. Historia Mathematica, 1(3):317–323, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508607490072X.

King:1974:SIP

[Kin74b] David A. King. Smithsonian institution project in medieval Islamic astronomy. Historia Mathematica, 1(2):183–184, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074900093.

King:1979:SNM

[Kin79] David A. King. Supplementary notes on medieval Islamic multiplication tables. Historia Mathematica, 6(4):405–417, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079900260.

King:1986:BRM

[Kin86a] David A. King. Book review: Matematiki i astronomi musulmanskogo srednevekovya i Ikh trudy (VIII–XVII vv.): [Mathematicians and Astronomers of the Islamic Middle Ages (VIII–XVII Centuries) and Their Works]. By G. P. Matvievskaya and B. A. Rosenfeld. Moscow (Nauka). 1983. 3 vols., 479 + 650 + 372 pp. Historia Mathematica, 13(3):306–308, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086086901060.

King:1986:BRS

[Kin86b] David A. King. Book review: Studies in the exact sciences in medieval Islam: By Ali A. Al-Daffa and John S. Stroyls. Dhahran (University of Petroleum and Minerals); New York (John Wiley & Sons). 1984. $39.95. Historia Mathematica, 13(3):303–306, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086086901059.

Kiss:1999:NJB

[Kis99] Elemér Kiss. Notes on János Bolyai’s researches in number theory. Historia Mathematica, 26(1):68–76, February 1, 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086098922121.

Kitcher:1978:BRG

[Kit78] Philip Kitcher. Book review: Grössenlehre II. Reine zahlenlehre: Bernard Bolzano: Edited by Jan Berg. Gesamtausgabe. Series II, A8. Stuttgart-Bad Cannstatt (Friedrich Fromm Verlag). 1976. Historia Mathematica, 5(3):358–363, August 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508607890126X.

Kitcher:1981:BRF

[Kit81] Philip Kitcher. Book review: Sur les fondements de l’analyse au XIX^{e} siècle: By Pierre Dugac. Université Catholique de Louvain. 1980. 179 pp. Historia Mathematica, 8(4):473–476, November 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900562.

Kjeldsen:1993:EHM

[Kje93] Tinne Hoff Kjeldsen. The early history of the moment problem. Historia Mathematica, 20(1):19–44, February 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086083710049.

Kjeldsen:2000:CHA

[Kje00] Tinne Hoff Kjeldsen. A contextualized historical analysis of the Kuhn–Tucker theorem in nonlinear programming: the impact of World War II. Historia Mathematica, 27(4):331–361, November 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086000922894.

Kjeldsen:2017:BRT

[Kje17] Tinne Hoff Kjeldsen. Book review: Taming the Unknown: a History of Algebra from Antiquity to the Early Twentieth Century, Victor J. Katz, Karen Hunger Parshall. Princeton University Press, Princeton/Oxford (2014). Historia Mathematica, 44(1):70–71, February 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086016300453.

Kracht:1990:WTH

[KK90] Manfred Kracht and Erwin Kreyszig. E. W. von Tschirnhaus: His role in early calculus and his work and impact on algebra. Historia Mathematica, 17(1):16–35, February 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609090076P.

Kaplan:2011:HHL

[KK11] Robert Kaplan and Ellen Kaplan. Hidden harmonies: the lives and times of the Pythagorean theorem. Bloomsbury Press, New York, NY, USA, 2011. ISBN 1-59691-522-6 (hardcover). xii + 290 pp. LCCN QA460.P8 K37 2011. URL http://catdir.loc.

gov/catdir/enhancements/fy1010/2010019959-b.html;   http:
//catdir.loc.gov/catdir/enhancements/fy1010/2010019959-d.
html.

Katz:2018:FCC

[KKS18] Boris Katz, Mikhail G. Katz, and Sam Sanders. A footnote to The crisis in contemporary mathematics. Historia Mathematica, 45(2): 176–181, May 2018. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0315086018300466.

Klamkin:1984:BRS

[Kla84] M. S. Klamkin. Book review: The Scottish book: Mathematics from the Scottish cafe: Edited by R. Daniel Mauldin. Boston (Birkhauser). 1981. xiii + 268 pp. $24.95. Historia Mathematica, 11(3):329, August 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086084900417.

Kleiner:2001:MAR

[Kle01] Israel Kleiner. Modern algebra and the rise of mathematical structures. Historia Mathematica, 28(4):304–312, November 2001.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086001923249.

Kleiner:2007:HAA

[Kle07] Israel Kleiner. A history of abstract algebra. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2007. ISBN 0-8176-4684-1 (paperback), 0-8176-4685-X (ebook), 0-387-34050-5 (hardcover). xiii + 168 pp. LCCN QA162 .K595 2007. URL

http://www.loc.gov/catdir/enhancements/fy0812/2007932362-
d.html;   http://www.loc.gov/catdir/enhancements/fy0812/
2007932362-t.html.

Kleinert:2017:BRJ

[Kle17] Andreas Kleinert. Book review: Jean le Rond D’Alembert : OEuvres complètes. Série V : Correspondance générale. Vol. 2 : Correspondance générale 1741–1752, Irène Passeron (Ed.). CNRS Éditions, Paris (2015), ISBN: 978-2-271-08749-2. Historia Mathematica, 44(2): 172–174, May 2017. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086017300344.

Kline:1975:OR

[Kli75] Morris Kline. Opening remarks. Historia Mathematica, 2(4):575, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508607590124X.

Kline:1976:CBB

[Kli76] Morris Kline. Carl B. Boyer — in memoriam. Historia Mathematica, 3(4):387–394, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508607690063X.

Knapp:1988:ZZT

[Kna88] Hans Georg Knapp. Zahl als Zeichen — zur “Technisierung” der Arithmetik im Mittelalter. (German) [Number as a sign — “mechanization” of arithmetic in the Middle Ages]. Historia Mathematica, 15(2):114–134, May 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508608890002X.

Knipping:2001:AMT

[Kni01] Christine Knipping. “Aller Männerkultur zum Trotz.” Frauen in Mathematik und Naturwissenschaften. (German) [“All male culture notwithstanding.” Women in mathematics and natural sciences]. Historia Mathematica, 28(3):253–254, August 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086001923201.

Knobloch:1978:BBM

[Kno78a] Eberhard Knobloch. Biographien bedeutender mathematiker, eine sammlung von biographien. (German) [Biographies of important mathematician, a collection of biographies]. Historia Mathematica, 5(4):483–485, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086078902288.

Knorr:1978:ASH

[Kno78b] Wilbur R. Knorr. Archimedes and the spirals: the heuristic background. Historia Mathematica, 5(1):43–75, February 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078901349.

Knobloch:1980:BRI

[Kno80] Eberhard Knobloch. Book review: Ian Hacking: the emergence of probability: Cambridge. 1975. 209 pp. Historia Mathematica, 7(2):212–216, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086080900555.

Knorr:1982:TFA

[Kno82] Wilbur Knorr. Techniques of fractions in ancient Egypt and Greece. Historia Mathematica, 9(2):133–171, May 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086082900015.

Knobloch:1983:RLE

[Kno83a] Eberhard Knobloch. Von Riemann zu Lebesgue — Zur Entwicklung der Integrationstheorie. (German) [From Riemann to Lebesgue — on the development of the theory of integration]. Historia Mathematica, 10(3):318–343, August 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086083900812.

Knorr:1983:TGG

[Kno83b] Wilbur Knorr. On the transmission of geometry from Greek into Arabic. Historia Mathematica, 10(1):71–78, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900344.

Knobloch:1988:BRL

[Kno88a] Eberhard Knobloch. Book review: En lisant le mémoire de Seidel de 1847: Par Pierre Dugac. Université Pierre et Marie Curie (Paris VI), DEUGSSM. 2ème Année, Option d’Histoire des Sciences (Février– Avril 1984). Paris. 1984. 29 pp. Historia Mathematica, 15(1):90–97, February 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086088900596.

Knoche:1988:HAW

[Kno88b] Norbert Knoche. History of algebra: Workshop, June 25–27, 1987, Essen. Historia Mathematica, 15(3):285–286, August 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900766.

Knoblocho:1990:A

[Kno90] Eberhard Knoblocho. Announcement. Historia Mathematica, 17(1): 4–5, February 1990. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/031508609090074N.

Knobloch:1991:EMB

[Kno91a] Eberhard Knobloch. Evert Marie Bruins (1909–1990). Historia Mathematica, 18(4):381–389, November 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086091903839.

Knorr:1991:MCQ

[Kno91b] Wilbur R. Knorr. On a medieval circle quadrature: De circulo quadrando. Historia Mathematica, 18(2):107–128, May 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609190495J.

Knobloch:1993:MSH

[Kno93a] Eberhard Knobloch. For Matthias Schramm on his sixtyfifth birthday. Historia Mathematica, 20(2):121–125, May 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086083710116.

Knorr:1993:ASD

[Kno93b] Wilbur R. Knorr. Arithmêtikê stoicheiôsis: On Diophantus and Hero of Alexandria. Historia Mathematica, 20(2):180–192, May 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086083710153.

Knorr:1993:AEA

[Kno93c] Wilbur R. Knorr. On an alleged error in Archimedes’ Conoids, prop. 1. Historia Mathematica, 20(2):193–197, May 1993. CODEN HI-MADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/S0315086083710165.

Knobloch:1994:MIC

[Kno94] Eberhard H. Knobloch. A message from the international commission on the history of mathematics. Historia Mathematica, 21(1):3, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086084710020.

Knobloch:2006:BCL

[Kno06] Eberhard Knobloch. Beyond Cartesian limits: Leibniz’s passage from algebraic to “transcendental” mathematics. Historia Mathematica, 33(1):113–131, February 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086004000199.

Knobloch:2007:BRC

[Kno07] Eberhard Knobloch. Book review: Clavius, une clé pour Euclide au XVIe siècle, Sabine Rommevaux. Librairie Philosophique J. Vrin, Paris (2005), ISBN 2-7116-1787-4. Historia Mathematica, 34(2):229– 230, May 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086006000565.

Knobloch:2017:LEJ

[Kno17] Eberhard Knobloch. Letter to the editors of the journal Historia Mathematica. Historia Mathematica, 44(3):280–282, August 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086017300514.

Knuth:1974:BRO

[Knu74] Donald E. Knuth. Book review: The origins of digital computers: Selected papers: Edited by Brian Randell. New York (Springer-Verlag), 1973. 464 p. Historia Mathematica, 1(2):204–207, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508607490024X.

Knuth:1983:BRH

[Knu83a] Donald E. Knuth. Book review: History of binary and other nondecimal numeration: By Anton Glaser. Los Angeles, CA (Tomash Publishers). 1981. 218 + xiii pp. Historia Mathematica, 10(2):236– 243, May 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086083900733.

Knuth:1983:RHB

[Knu83b] Donald E. Knuth. Review of History of Binary and other Nondecimal Numeration, by Anton Glaser. Historia Mathematica, 10(??):236– 243, 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Knudsen:2011:BRG

[Knu11] Toke L. Knudsen. Book review: Gaṅitasarakaumudī: the Moonlight of the Essence of Mathematics by T .hakkura Pheru. Edited with Introduction, Translation, and Mathematical Commentary by SaKHYa. New Delhi (Manohar). 2009. ISBN 978-81-7304-809-8. viii + 278 pp. No price given. Historia Mathematica, 38(3):423–425, August 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086011000048.

Kranz:2009:MMC

[KO09] Horst Kranz and Walter Oberschelp. Mechanisches Memorieren und Chiffrieren um 1430: Johannes Fontanas Tractatus de instrumentis artis memorie. (German) [Mechanical memorization and ciphering in 1430: John Fontana’s Tractatus de instrumentis artis memorie], volume 59 of Boethius (Series); Wissenschaftsgeschichte. Steiner, Stuttgart, Germany, 2009. ISBN 3-515-09296-X (hardcover). 167 pp. LCCN BF383 .K77 2009.

Koelblen:1994:JTP

[Koe94] Sabine Koelblen. Journée sur la théorie des proportions: Paris, Centre Koyré, May 22, 1993. Historia Mathematica, 21(1):74–75, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508600880064X.

Koppelman:1975:PM

[Kop75] Elaine Koppelman. Progress in mathematics. Historia Mathematica, 2(4):457–463, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086075901032.

Kotz:1979:BRF

[Kot79] Samuel Kotz. Book review: Felix Klein and Sophus Lie: By I. M. Yaglom. (In Russian.) Mathematics and Cybernetics, No. 11. Moscow (Znanie). 1977. Historia Mathematica, 6(4):468–469, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079900430.

Knobloch:1985:NEN

[KP85] Eberhard Knobloch and Helena Pycior. New editors named for Historia Mathematica. Historia Mathematica, 12(4):313, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086085900394.

Knobloch:1986:E

[KP86] Eberhard H. Knobloch and Helena M. Pycior. Editorial. Historia Mathematica, 13(1):3–4, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508608690220X.

Kennedy:1985:TPG

[KQS85] John W. Kennedy, Louis V. Quintas, and Maciej M. Syslo. The theorem on planar graphs. Historia Mathematica, 12(4):356–368, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508608590045X.

Kremer:2003:TJS

[Kre03] Richard L. Kremer. Thoughts on John of Saxony’s method for finding times of true syzygy. Historia Mathematica, 30(3):263–277, August 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086002000253.

Kromer:2006:BRH

[Krö06] Ralf Krömer. Book review: Das Homöomorphieproblem, insbesondere der 3-Mannigfaltigkeiten, in der Topologie 1892–1935: By Klaus Volkert. Philosophia scientiæ, Cahier spécial 4. Paris (Kimé). 2002. ISBN 2-84174-275-X. 343 pp. EUR 29. Historia Mathematica, 33(2):259–260, May 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086005000765.

Kromer:2008:GTE

[Krö08] Ralf Krömer. A German translation of Euler’s articles on the polyhedron formula now available on the world wide Web. Historia Mathematica, 35(4):327–328, November 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008000293.

Kromer:2014:BRJ

[Krö14] Ralf Krömer. Book review: Jacques Feldbau, Topologe. Das Schicksal eines jüdischen Mathematikers (1914–1945), Michèle Audin, in: Mathematik im Kontext. Springer, Berlin, Heidelberg (2012), ISBN 978-3-642-25803-9. Historia Mathematica, 41(1):112–115, February 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086013000694.

Kromer:2015:BRJ

[Krö15] Ralf Krömer. Book review: Justifier en mathématiques, Dominique Flament, Philippe Nabonnand (Eds.). Éditions de la Maison des sciences de l’homme, Paris (2011), ISBN: 978-2-7351-1414-6. Historia Mathematica, 42(2):233–237, May 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086013000414.

Kromer:2017:BRG

[Krö17] Ralf Krömer. Book review: Géométrie et Topologie différentielles (1918–1932). Renaud Chorlay. Paris (Hermann) (2015), ISBN: 978-2-7056-8106-7. Historia Mathematica, 44(3):285–287, August 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086017300216.

Kopell:1975:CBT

[KS75] Nancy Kopell and Gabriel Stolzenberg. Commentary on Bishop’s talk. Historia Mathematica, 2(4):519–521, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL

http://www.sciencedirect.com/science/article/pii/0315086075901147.

Krafft:2008:NNS

[KS08] Fritz A. Krafft and Christoph J. Scriba. Nikolai Nikolajevich Stuloff (1914–2006). Historia Mathematica, 35(4):273–276, November 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508600800058X.

Kaufholz-Soldat:2012:BRM

[KS12] Eva Kaufholz-Soldat. Book review: “Morgen möchte ich wieder 100 herrliche Sachen ausrechnen.” Iris Runge bei Osram und Telefunken. By Renate Tobies. Boethius, vol. 61. Stuttgart (Steiner). 2010. ISBN 978-3-515-09638-6. 412 pp. 72 EUR. Historia Mathematica, 39(3):351–353, August 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086012000031.

Kaufholz-Soldat:2013:BRR

[KS13a] Eva Kaufholz-Soldat. Book review: Remembering Sofya Kovalevskaya, Michèle Audin. Springer, London (2011), ISBN: 978-0-85729-928-4. Historia Mathematica, 40(4):458–460, October 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086013000244.

Kosmann-Schwarzbach:2013:SDP

[KS13b] Yvette Kosmann-Schwarzbach. Siméon-Denis Poisson: les mathématiques au service de la science. Histoire des mathématiques et des sciences physiques. Les Éditions de l’École polytechnique, Palaiseau, France, 2013. ISBN 2-7302-1584-0. 522 pp. LCCN QA29.P64 S55 2012.

Kullman:1976:THY

[Kul76] David E. Kullman. Two hundred years of mathematics in America. Historia Mathematica, 3(4):463–468, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508607690077X.

Labarthe:2001:SAA

[Lab01] Marie-Hélène Labarthe. Summa de l’art d’Aritmètica. Francesc Santcliment. Historia Mathematica, 28(4):302–304, November 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086001923237.

Lai:1975:DNR

[Lai75] Tyrone Lai. Did Newton renounce infinitesimals? Historia Mathematica, 2(2):127–136, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086075901378.

Laird:2010:MTD

[Lai10] W. R. Laird. Mathématiques et théorie du mouvement (XIVe — XVIe siècles). Historia Mathematica, 37(2):305–308, May 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009000573.

Lamarra:1987:LLP

[Lam87] Antonio Lamarra. L’infinito in Leibniz. Problemi e terminologia das unendliche bei Leibniz. Probleme und Terminologie: Simposio Internazionale, November 6–8, 1986, Rome. Historia Mathematica, 14(3):293–294, August 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086087900577.

Lamande:2019:COM

[Lam19] Pierre Lamandé. Sur la conception des objets et des méthodes mathématiques dans les textes philosophiques de d’Alembert. Historia Mathematica, 49(??):20–59, November 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086018300600.

Lane:1975:PMS

[Lan75] Saunders Mac Lane. A proposed “mini-study” of mathematical developments. Historia Mathematica, 2(1):59–61, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900373.

Langermann:1996:MHT

[Lan96] Y. Tzvi Langermann. Medieval Hebrew texts on the quadrature of the lune. Historia Mathematica, 23(1):31–53, February 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086096900038.

Langton:2007:BRD

[Lan07] Stacy G. Langton. Book review: Dr. Euler’s Fabulous Formula: Cures Many Mathematical Ills, Paul J. Nahin. Princeton University Press, Princeton (2006), ISBN 0-691-11822-1. Historia Mathematica, 34(4):451–454, November 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S031508600700002X.

Laugwitz:1987:ISQ

[Lau87] Detlef Laugwitz. Infinitely small quantities in Cauchy’s textbooks. Historia Mathematica, 14(3):258–274, August 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086087900450.

Laugwitz:1992:LZI

[Lau92] Detlef Laugwitz. “Das letzte Ziel ist immer die Darstellung einer Funktion”: Grundlagen der analysis bei Weierstraß 1886, historische Wurzeln und Parallelen. (German) [“The ultimate goal is always the representation of a function”: Fundamentals of analysis by Weierstraß, 1886, historical roots and parallels]. Historia Mathematica, 19 (4):341–355, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609290001R.

Laugwitz:1999:BRT

[Lau99] Detlef Laugwitz, editor. Bernhard Riemann, 1826–1866: turning points in the conception of mathematics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1999. ISBN 0-8176-4040-1. xvi + 357 pp. LCCN QA29.R425 L3813 1999.

Lawrence:1998:BRE

[Law98] Snezana Lawrence. Book review: Entre Mécanique et Architecture — Between Mechanics and Architecture. Edited by Patricia Radeletde-Grave and Edoardo Benvenuto. Historia Mathematica, 25(2):221– 224, May 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086097921893.

Lawvere:2005:BRN

[Law05] F. William Lawvere. Book review: A new branch of mathematics, “The Ausdehnungslehre of 1844,” and other works. Open Court (1995). Translated by Lloyd C. Kannenberg, with foreword by Albert C. Lewis. Extension Theory, “The Ausdehnungslehre of 1862”, vol. 19. American Mathematical Society… London Mathematical Society (2000). Translated and with a foreword and notes by Lloyd C. Kannenberg. Historia Mathematica, 32(1):99–106, February 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508600400059X.

Lebesgue:2004:LIL

[LBB^{+}04] Henri Léon Lebesgue, Emile Borel, Bernard Bru, Pierre Dugac, et al. Les lendemains de l’intégrale: lettres à Émile Borel. (French) [The tomorrows of the integral: letters to Émile Borel]. Vuibert, Paris, France, 2004. ISBN 2-7117-5309-3. 239 pp. LCCN QA29.L44 A4 2004.

Lugaresi:2022:FEC

[LD22] M. G. Lugaresi and C. D’Alterio. Federigo Enriques (1871–1946): a critical study of Lezioni di geometria proiettiva. Historia Mathematica, 60(??):15–52, August 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086022000222.

Le:2015:GEF

[Lê15] François Lê. “Geometrical equations”: Forgotten premises of Felix Klein’s Erlanger Programm. Historia Mathematica, 42(3):315–342, August 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086014001050.

Le:2018:RCT

[Lê18] François Lê. The recognition and the constitution of the theorems of closure. Historia Mathematica, 45(3):217–236, August 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086018300545.

Le:2020:GGO

[Lê20] François Lê. “Are the genre and the Geschlecht one and the same number?” an inquiry into Alfred Clebsch’s Geschlecht. Historia Mathematica, 53(??):71–107, November 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086020300239.

Le:2021:TFA

[Lê21] François Lê. Book review: Théorie des fonctions algébriques d’une variable, Richard Dedekind, Heinrich Weber. Vrin (2020), Texte introduit, traduit et annoté par Emmylou Haffner. Mathesis. Historia Mathematica, 57(??):94–96, November 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508602100063X.

Le:2024:BRR

[Lê24] François Lê. Book review: The Richness of the History of Mathematics: a Tribute to Jeremy Gray. Karine Chemla, José Ferreirós, Lizhen Ji, Erhard Scholz, Chang Wang (eds.). Springer, 2023. Historia Mathematica, 68(??):38–39, September 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086024000740.

Lefevre:2010:BRM

[Lef10] Wolfgang Lefèvre. Book review: Mechanisches Memorieren und Chiffrieren um 1430: Johannes Fontanas Tractatus de instrumentis artis memorie. Edited by H. Kranz and W. Oberschelp. Stuttgart (Franz Steiner Verlag). 2007. ISBN 978-3-515-09296-8. 167 pp. 38.00 EUR. Historia Mathematica, 37(4):710–712, November 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009000718.

Leighton:1977:SEH

[Lei77] Albert C. Leighton. Some examples of historical cryptanalysis. Historia Mathematica, 4(3):319–337, August 1977. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086077900593.

Leighton:1979:BRA

[Lei79] Albert C. Leighton. Book review: An annotated bibliography of cryptography: By David Shulman. Garland Reference Library of the Humanities, Vol. 37 New York/London (Garland Publishing Inc.). 1976. xvi + 372 pp. illus. $35.00. Historia Mathematica, 6(2):213–218, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079900934.

Leitao:2009:BRA

[Lei09] Henrique Leitão. Book review: Análise Histórica de Livros de Matemática. Notas de Aula, Gert Schubring. Autores Associados, Campinas, SP (2003), ISBN 85-7496-061-6. Historia Mathematica, 36(2):178–179, May 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S031508600900007X.

Lemaire:1974:BRA

[Lem74] Claude Lemaire. Book review: Abriss der Geschichte der Mathematik. Fünfte, erweiterte Auflage. Mit einem Anhang über die Mathematik des 20. Jahrhunderts von I. Pogrebysski: By Dirk J. Struik. Berlin (VEB Deutscher Verlag der Wissenschaften), 1972. 255 pp. DM 12.80. Historia Mathematica, 1(1):95–96, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074901864.

Lenoir:1979:DGT

[Len79] Timothy Lenoir. Descartes and the geometrization of thought: the methodological background of Descartes’ géométrie. Historia Mathematica, 6(4):355–379, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079900235.

Levitt:1997:BRM

[Lev97] Norman Levitt. Book review: In Measure, Number, and Weight: Studies in Mathematics and Culture. By Jens Høyrup. Albany (State University of New York Press). 1994, xviii + 430 pp. $19.95. Historia Mathematica, 24(3):339–342, August 1997. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086097921625.

Levy:1999:BRG

[Lév99] Tony Lévy. Book review: Germanische Zahlwörter: Sprach- und kulturgeschichtliche Untersuchungen insbesondere zur Zahl 12. By George Schuppener. Historia Mathematica, 26(1):81–82, February 1, 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508609992222X.

Lewis:1975:BMM

[Lew75] Albert C. Lewis. Bringing modern mathematics to bear on historical research. Historia Mathematica, 2(1):84–85, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900531.

Lewis:1976:HMA

[Lew76] Albert C. Lewis. History of mathematics at the American association for the advancement of science annual meeting, Boston 18–24 February 1976. Historia Mathematica, 3(2):218–219, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900471.

Lewis:1978:Aa

[Lew78a] Albert C. Lewis. Abstracts. Historia Mathematica, 5(2):246–251, May 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086078900654.

Lewis:1978:Ab

[Lew78b] Albert C. Lewis. Abstracts. Historia Mathematica, 5(3):368–377, August 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086078901295.

Lewis:1978:Ac

[Lew78c] Albert C. Lewis. Abstracts. Historia Mathematica, 5(4):488–494, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086078902318.

Lewis:1978:EAA

[Lew78d] Albert C. Lewis. Establishment of “The archives of American mathematics”. Historia Mathematica, 5(3):340–341, August 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078901209.

Lewis:1979:Aa

[Lew79a] Albert C. Lewis. Abstracts. Historia Mathematica, 6(2):229–230, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508607990096X.

Lewis:1979:Ab

[Lew79b] Albert C. Lewis. Abstracts. Historia Mathematica, 6(4):473–479, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086079900454.

Lewis:1980:Aa

[Lew80a] Albert C. Lewis. Abstracts. Historia Mathematica, 7(1):101–107, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086080900865.

Lewis:1980:Ab

[Lew80b] Albert C. Lewis. Abstracts. Historia Mathematica, 7(2):217–224, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086080900567.

Lewis:1980:Ac

[Lew80c] Albert C. Lewis. Abstracts. Historia Mathematica, 7(4):459–467, November 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086080900178.

Lewis:1981:Aa

[Lew81a] Albert C. Lewis. Abstracts. Historia Mathematica, 8(3):379–383, August 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086081900811.

Lewis:1981:Ab

[Lew81b] Albert C. Lewis. Abstracts. Historia Mathematica, 8(4):489–493, November 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086081900628.

Lewis:1982:Aa

[Lew82a] Albert C. Lewis. Abstracts. Historia Mathematica, 9(1):124–127, February 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086082901537.

Lewis:1982:Ab

[Lew82b] Albert C. Lewis. Abstracts. Historia Mathematica, 9(3):361–367, August 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508608290132X.

Lewis:1982:Ac

[Lew82c] Albert C. Lewis. Abstracts. Historia Mathematica, 9(4):497–504, November 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086082901161.

Lewis:1983:Aa

[Lew83a] Albert C. Lewis. Abstracts. Historia Mathematica, 10(1):136–143, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086083900587.

Lewis:1983:Ab

[Lew83b] Albert C. Lewis. Abstracts. Historia Mathematica, 10(2):249– 256, May 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086083900757.

Lewis:1983:Ac

[Lew83c] Albert C. Lewis. Abstracts. Historia Mathematica, 10(3):372–379, August 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086083900861.

Lewis:1983:Ad

[Lew83d] Albert C. Lewis. Abstracts. Historia Mathematica, 10(4):485–490, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508608390023X.

Lewis:1984:Aa

[Lew84a] Albert C. Lewis. Abstracts. Historia Mathematica, 11(2):242– 249, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086084800270.

Lewis:1984:Ab

[Lew84b] Albert C. Lewis. Abstracts. Historia Mathematica, 11(3):347–354, August 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086084900478.

Lewis:1985:Aa

[Lew85a] Albert C. Lewis. Abstracts. Historia Mathematica, 12(1):91–97, February 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086085900795.

Lewis:1985:Ab

[Lew85b] Albert C. Lewis. Abstracts. Historia Mathematica, 12(4):394–400, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086085900588.

Lewis:1985:BRC

[Lew85c] Albert C. Lewis. Bertrand Russell Conference. Historia Mathematica, 12(1):74–78, February 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086085900734.

Lewis:1986:Aa

[Lew86a] Albert C. Lewis. Abstracts. Historia Mathematica, 13(1):89–94, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086086902430.

Lewis:1986:Ab

[Lew86b] Albert C. Lewis. Abstracts. Historia Mathematica, 13(3):311–320, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086086901084.

Lewis:1987:A

[Lew87] Albert C. Lewis. Abstracts. Historia Mathematica, 14(4):390–396, November 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086087900723.

Lewis:1988:Aa

[Lew88] Albert C. Lewis. Abstracts. Historia Mathematica, 15(1):98–99, February 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086088900602.

Lewis:1990:BRM

[Lew90] Albert C. Lewis. Book review: Mathematical visions: the pursuit of geometry in Victorian England: By Joan L. Richards. Boston/San Diego (Academic Press). 1988. xiii + 266 pp. $34.95. Historia Mathematica, 17(3):272–278, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086090900156.

Lewis:1995:BRS

[Lew95] Albert C. Lewis. Book review: The search for E. T. Bell, also known as John Taine: By Constance Reid. Washington, D.C. (The Mathematical Association of America). 1993. x + 372 pp. $35.00. Historia Mathematica, 22(1):78–80, February 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086085710087.

Lewis:1998:BRP

[Lew98] Albert C. Lewis. Book review: The Problem of the Earth’s Shape from Newton to Clairaut: the Rise of Mathematical Science in Eighteenth-Century Paris and the Fall of “Normal” Science. By John L. Greenberg. Cambridge/New York (Cambridge University Press). 1995. Historia Mathematica, 25(3):320–325, August 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086098922042.

Lewis:1999:BRQ

[Lew99] Albert C. Lewis. Book review: The Quest for Longitude: the Proceedings of the Longitude Symposium, Harvard University, Cambridge, Massachusetts, November 4–6, 1993. Edited by William J. H. Andrewes. Historia Mathematica, 26(2):178–180, May 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086099922383.

Lewis:2004:BRM

[Lew04a] Albert C. Lewis. The beginnings of the R. L. Moore school of topology. Historia Mathematica, 31(3):279–295, August 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086003000508.

Lewis:2004:KMI

[Lew04b] Albert C. Lewis. Kenneth O. May and information retrieval in mathematics. Historia Mathematica, 31(2):186–195, May 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086003001034.

Lewis:2009:BRG

[Lew09] Albert C. Lewis. Book review: Graßmann, Hans-Joachim Petsche. Birkhäuser, Basel, Boston, Berlin (2006), ISBN 978-3-7643-7257-6. Historia Mathematica, 36(1):82–85, February 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008000633.

Letter:1974:RD

[LGG74] A. Letter and Ivor Grattan-Guinness. Recording duplication. Historia Mathematica, 1(4):456, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074900433.

Langermann:1984:HUH

[LH84] Y. T. Langermann and J. P. Hogendijk. A hitherto unknown Hellenistic treatise on the regular polyhedra. Historia Mathematica, 11(3):325–326, August 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086084900399.

Lichtenberg:1997:IGS

[Lic97] Heiner Lichtenberg. Zur Interpretation der Gaußschen Osterformel und ihrer Ausnahmeregeln. (German) [Gauss’s interpretation of the Easter formula its exclusion rules]. Historia Mathematica, 24(4): 441–444, November 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086097921704.

Linis:1974:BRA

[Lin74] V. Linis. Book review: Akademik V. A. Steklov: By V. S. Vladimirov and I. I. Markush. Moscow (Znanie), 1973. 63 p. Russ. 10 kop. Historia Mathematica, 1(3):339–341, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074900809.

Lloyd:1979:FMV

[Llo79] E. Keith Lloyd. On the forgotten Mr. Vincent. Historia Mathematica, 6(4):448–450, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086079900338.

Lloyd:1990:RPM

[Llo90] E. Keith Lloyd. Redfield’s proofs of MacMahon’s conjecture. Historia Mathematica, 17(1):36–47, February 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609090077Q.

Lim:2013:GAP

[lLW13] Tina Su lyn Lim and Donald B. Wagner. The Grand Astrologer’s platform and ramp: Four problems in solid geometry from Wang Xiaotong’s ‘Continuation of ancient mathematics’ (7th Century AD). Historia Mathematica, 40(1):3–35, February 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086012000596.

Laubenbacher:2001:LSN

[LMP01] Reinhard Laubenbacher, Gary McGrath, and David Pengelley. Lagrange and the solution of numerical equations. Historia Mathematica, 28(3):220–231, August 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S031508600192316X.

Laugwitz:1994:RCH

[LN94] Detlef Laugwitz and Erwin Neuenschwander. Riemann and the Cauchy–Hadamard formula for the convergence of power series. Historia Mathematica, 21(1):64–70, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086084710081.

LoBello:1983:AMM

[Lo 83] Anthony Lo Bello. Albertus Magnus and mathematics: a translation with annotations of those portions of the commentary on Euclid’s Elements published by Bernhard Geyer. Historia Mathematica, 10(1):3–23, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086083900307.

LoBello:1984:BRM

[Lo 84] Anthony Lo Bello. Book review: Medieval mensuration: quadrans vetus and geometrie due sunt partes principales…: Edited by Nan L. Hahn. Transactions of the American Philosophical Society, Vol. 72 Part 8. Philadelphia (The American Philosophical Society). 1982. 204 pp. + lxxxv. Historia Mathematica, 11(2):239–240, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086084800257.

LoBello:1985:BRA

[Lo 85] Anthony Lo Bello. Book review: Albertus (Magnus)’ commentaar op Euclides’ Elementen der geometrie: (The Commentary of Albertus Magnus on Euclid’s Elements of Geometry). By Paul M. J. E. Tummers. 2 vols. (Vol. I: 329 pp.; Vol. II: iv + 370 pp.). Nijmegen (Ingenium Publishers). 1984. Historia Mathematica, 12(2):185–186, May 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608590014X.

Lokken:1980:DNI

[Lok80] Roy N. Lokken. Discussions on Newton’s infinitesimals in Eighteenth-Century Anglo–America. Historia Mathematica, 7(2):141–155, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608090035X.

Longa:2021:PPM

[Lon21] Gianluca Longa. Book review: Ptolemy’s Philosophy: Mathematics as a Way of Life, Jacqueline Feke. Princeton University Press (2018), 256 pp., ISBN: 978-0-691-17958-2. Historia Mathematica, 54 (??):117–119, February 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086020300690.

Lorenat:2015:FRI

[Lor15] Jemma Lorenat. Figures real, imagined, and missing in Poncelet, Plücker, and Gergonne. Historia Mathematica, 42(2):155–192, May 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086014000767.

Lorenat:2016:BRS

[Lor16] Jemma Lorenat. Book review: Sciences Mathématiques 1750– 1850: continuités et ruptures, Christian Gilain, Alexandre Guilbaud (Eds.). CNRS Éditions, Paris (2015), ISBN: 0-8018-7434-3. Historia Mathematica, 43(4):446–449, November 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086016300210.

Lorenat:2018:PLW

[Lor18] Jemma Lorenat. Book review: Pushing Limits: From West Point to Berkeley and Beyond, Ted Hill. American Mathematical Society and MAA Press, Providence (2017), 294 pages. 45 dollars, ISBN: 978-1-4704-3584-4. Historia Mathematica, 45(3):302–304, August 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086018300697.

Laubenbacher:2010:VCQ

[LP10] Reinhard Laubenbacher and David Pengelley. “Voici ce que j’ai trouvé:” Sophie Germain’s grand plan to prove Fermat’s Last Theorem. Historia Mathematica, 37(4):641–692, November 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009001347.

Luciano:2018:CIL

[Luc18] Erika Luciano. Constructing an international library: the collections of journals in Turin’s Special Mathematics Library. Historia Mathematica, 45(4):433–449, November 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086018300880.

Luciano:2021:FGT

[Luc21] Erika Luciano. On Francesco G. Tricomi’s heritage: Archive and miscellany. Historia Mathematica, 56(??):73–84, August 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086021000392.

Lumpkin:1980:EPT

[Lum80] Beatrice Lumpkin. The Egyptians and Pythagorean triples. Historia Mathematica, 7(2):186–187, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900373.

Lumiste:1997:MBR

[Lum97] Ülo Lumiste. Martin Bartels as researcher: His contribution to analytical methods in geometry. Historia Mathematica, 24(1):46–65, February 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086097921236.

Lutzen:1982:JLC

[Lüt82] Jesper Lützen. Joseph Liouville’s contribution to the theory of integral equations. Historia Mathematica, 9(4):373–391, November 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082901045.

Lutzen:2005:MIG

[Lüt05] Jesper Lützen. Mechanistic images in geometric form: Heinrich Hertz’s Principles of Mechanics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-856737-5. xiii + 318

pp. LCCN QA805 .L85 2005. URL http://www.loc.gov/catdir/
enhancements/fy0638/2005001534-d.html;   http://www.loc.
gov/catdir/enhancements/fy0725/2005001534-b.html;   http:
//www.loc.gov/catdir/toc/ecip056/2005001534.html.

Lutzen:2008:BRP

[Lüt08] Jesper Lützen. Book review: Perspectives on Scandinavian Science in the Early Twentieth Century, Reinhard Siegmund-Schultze, Sørensen (Eds.). The Norwegian Academy of Sciences and Letters. Novus Forlag, Oslo (2006), ISBN 978-82-7099-442-7. Historia Mathematica, 35 (4):334–336, November 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086007000985.

Lutzen:2009:BRA

[Lüt09a] Jesper Lützen. Book review: Axiomatik und Empirie. Eine wissenschaftstheoriegeschichtliche Untersuchung zur Mathematischen Naturphilosophie von Newton bis Neumann (Axiomatics and empiricism. An investigation from the point of view of the history of the theory of science of mathematical natural philosophy from Newton through Neumann), Helmut Pulte. Wissenschaftliche Buchgesellschaft, Darmstadt (2005), ISBN 3-534-15894-6. Historia Mathematica, 36(1):87–89, February 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008001079.

Lutzen:2009:WWW

[Lüt09b] Jesper Lützen. Why was wantzel overlooked for a century? the changing importance of an impossibility result. Historia Mathematica, 36 (4):374–394, November 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S031508600900010X.

Lutzen:2015:BRA

[Lüt15] Jesper Lützen. Book review: Aventures de l’analyse de Fermat à Borel. Mélanges en l’honneur de Christian Gilain, Susanne Féry (Ed.), in: Histoires de géométries. Presses Universitaires de Nancy, Editions Universitaires de Lorraine, Nancy (2012), ISBN: 978-2-8143-0125-2. Historia Mathematica, 42(2):230–233, May 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086014000780.

Lutzen:2019:HMI

[Lüt19] Jesper Lützen. How mathematical impossibility changed welfare economics: a history of Arrow’s impossibility theorem. Historia Mathematica, 46(??):56–87, February 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086018300508.

Lutzen:2020:BRI

[Lüt20] Jesper Lützen. Book review: The Impossibility of Squaring the Circle in the 17th Century. A Debate Among Gregory, Huygens and Leibniz, Davide Crippa. Birkhäuser (2019), viii + 184 pp. Historia Mathematica, 52(??):99–103, August 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086020300045.

Lutzen:2021:HGR

[Lüt21] Jesper Lützen. Hjelmslev’s geometry of reality. Historia Mathematica, 54(??):95–116, February 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086020300677.

Lay-Yong:1979:BRW

[LY79] Lam Lay-Yong. Book review: Was Pythagoras Chinese? An examination of right triangle theory in ancient China: by Frank J. Swetz and T. I. Kao. Pennsylvania (Pennsylvania State University Press). 1977. 75pp. $3.95. Historia Mathematica, 6(2):221–228, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079900958.

Lay-Yong:1980:CCB

[LY80] Lam Lay-Yong. The Chinese connection between the Pascal triangle and the solution of numerical equations of any degree. Historia Mathematica, 7(4):407–424, November 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900051.

Lay-Yong:1985:CCC

[LYK85] Lam Lay-Yong and Shen Kangsheng. The Chinese concept of Cavalieri’s principle and its applications. Historia Mathematica, 12(3):219– 228, August 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086085900205.

Lay-Yong:1989:MSL

[LYK89] Lam Lay-Yong and Shen Kangshen. Methods of solving linear equations in traditional China. Historia Mathematica, 16(2):107–122, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086089900335.

Lay-Yong:1986:CMA

[LYTS86] Lam Lay-Yong and Ang Tian-Se. Circle measurements in ancient China. Historia Mathematica, 13(4):325–340, November 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086086900558.

Lewis:1988:Ab

[LZ88] Albert C. Lewis and David E. Zitarelli. Abstract. Historia Mathematica, 15(3):295–297, August 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086088900808.

Mancosu:2010:DCH

[MA10] Paolo Mancosu and Andrew Arana. Descartes and the cylindrical helix. Historia Mathematica, 37(3):403–427, August 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009000949.

Maat:2007:BRC

[Maa07] Jaap Maat. Book review: The Correspondence of John Wallis (1616–1703): Volume II (1660–September 1668), Philip Beeley, Christoph J. Scriba (Eds.). Oxford University Press, Oxford (2005), ISBN 0-19-856601-8. Historia Mathematica, 34(1):108–111, February 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005001011.

Mackey:1975:SRH

[Mac75] George Mackey. Some remarks on harmonic analysis. Historia Mathematica, 2(4):555–562, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075901202.

MacLachlan:1976:BRN

[Mac76] James MacLachlan. Book review: Les nouvelles pensées de Galilée: Par Marin Mersenne. Edition critique avec introduction et notes. Par Pierre Costabel et Michel-Paul Lerner. Avant-propos de Bernard Rochot. Paris (Librairie Philosophique J. Vrin). 1973. Vol. I, 130 pp.; Vol. II (Notes et Index), pp. 131–319. Historia Mathematica, 3(2):223–224, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076900525.

Maclachlan:1977:MSG

[Mac77] James Maclachlan. Mersenne’s solution for Galileo’s problem of the rotating Earth. Historia Mathematica, 4(2):173–182, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077901100.

Mack:1978:RCS

[Mac78a] J. M. Mack. Report on a conference for secondary teachers on the history of mathematics. Historia Mathematica, 5(4):461–463, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508607890215X.

MacKenzie:1978:WSM

[Mac78b] Donald MacKenzie. Workshop on the sociology of mathematical knowledge and the social history of mathematics. Historia Mathematica, 5(2):228–232, May 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086078900575.

Maeyama:1989:SAG

[Mae89] Y. Maeyama. A simple, accurate, geometrical approximation to the Keplerian motion. Historia Mathematica, 16(4):353–359, November 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086089900827.

Maffioli:2010:DAA

[Maf10] Cesare Maffioli. La via delle acque, 1500–1700: appropriazione delle arti e trasformazione delle matematiche. (Italian) [The way of the waters, 1500–1700: appropriation of arts and transformation of mathematics], volume 4 of Hyperchen. L. S. Olschki, Firenze, Italia, 2010. ISBN 88-222-6008-2. xxii + 394 + 1 pp. LCCN TC145 .M3155 2010.

43.00 EUR. URL http://www.loc.gov/catdir/toc/casalini09/
10482326.pdf.

Mahoney:1978:BRD

[Mah78] Micheal S. Mahoney. Book review: Doctrine de l’angle solide: By Florimond de Beaune (1601–1652). Edited by Pierre Costabel, transcribed from the original manuscript by Bernard Barbiche, typed by Arlettte Chancrin. Paris (Librairie Philosophique J. Vrin). (Collections des travaux de l’Académie Internationale d’Histoire des Sciences, No. 19). 1975. 157 pp. Historia Mathematica, 5(3):363–366, August 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086078901271.

Mahoney:1984:CCM

[Mah84] Michael S. Mahoney. Changing canons of mathematical and physical intelligibility in the later 17th century. Historia Mathematica, 11 (4):417–423, November 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086084900260.

Mainzer:1987:SST

[Mai87] Klaus Mainzer. Symmetrie symposion: Technische Hochschule Darmstadt 13.–17. Juni 1986. Historia Mathematica, 14(2):183–185, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086087900280.

Maieru:2007:JWV

[Mai07] Luigi Maierù. John Wallis: una vita per un progetto. (Italian) [John Wallis: a life for a project], volume 1 of Historia scientiae in Mediterraneo. Rubbettino, Soveria Mannelli, Italia, 2007. ISBN 88-498-1904-8. 601 pp. LCCN QA29.W3 M37 2007. 40.00 EUR. URL http://

www.loc.gov/catdir/toc/casalini07/07275021.pdf.

Maieru:2008:SGG

[Mai08] Luigi Maierù. Scienza, geometria, geometrie: un percorso storicodidattico. (Italian) [Science, geometry, geometry: a historicaleducational path]. Università. Rubbettino, Soveria Mannelli, Italia, 2008. ISBN 88-498-2156-5. 334 pp. LCCN QA443.5 .M33 2008.

13.00 EUR. URL http://www.loc.gov/catdir/toc/casalini07/
08342571.pdf.

Maieru:2009:SCN

[Mai09] Luigi Maierù. Le sezioni coniche nel Seicento. (Italian) [Conic sections in the Seventeenth Century]. Università. Rubbettino, Soveria Mannelli (Catanzaro), Italia, 2009. ISBN 88-498-2461-0. 403 pp. LCCN QA385 .M354 2009.

Maidment:2021:EML

[Mai21] Alison Maidment. The Edinburgh Mathematical Laboratory and Edmund Taylor Whittaker’s role in the early development of numerical analysis in Britain. Historia Mathematica, 55(??):39–63, May 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086020300719.

Malik:1979:HMC

[Mal79] M. A. Malik. A history of mathematics course for teachers. Historia Mathematica, 6(4):450–451, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508607990034X.

Malet:1997:BRM

[Mal97] Antoni Malet. Book review: La méthode de Fermat: Son statut et sa diffusion. By Giovanna Cleonice Cifoletti. Cahiers d’Histoire et de Philosophie des Sciences, No. 33. Paris (Société Française d’Histoire des Sciences et des Techniques. Belin Diffusion). 1990. Pp. 243. ISSN 0221-3664. FF60. Historia Mathematica, 24(2):208–209, May 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086097921510.

Malet:2006:BRG

[Mal06a] Antoni Malet. Book review: Gino Arrighi. La matematica dell’Età di Mezzo. Scritti scelti, F. Barberini, R. Franci, L. Toti Rigatelli (Eds.). Edizioni ETS, Pisa (2004), ISBN 88-467-1043-6. Historia Mathematica, 33(3):364–365, August 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086005000455.

Malet:2006:RNN

[Mal06b] Antoni Malet. Renaissance notions of number and magnitude. Historia Mathematica, 33(1):63–81, February 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086004000928.

Malpangotto:2013:LAS

[Mal13] Michela Malpangotto. L’univers auquel s’est confronté Copernic: La sphère de Mercure dans les Theoricae novae planetarum de Georg Peurbach. Historia Mathematica, 40(3):262–308, August 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086013000165.

Mancosu:1989:MCF

[Man89] Paolo Mancosu. The metaphysics of the calculus: a foundational debate in the Paris Academy of Sciences, 1700–1706. Historia Mathematica, 16(3):224–248, August 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086089900190.

Manders:2006:ARF

[Man06] Kenneth Manders. Algebra in Roth, Faulhaber, and Descartes. Historia Mathematica, 33(2):184–209, May 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086005000480.

Martzloff:1990:DCM

[Mar90a] Jean-Claude Martzloff. Deuxième colloque maghrébin sur l’histoire des mathématiques arabes: Tunis, 1–3 décembre 1988. Historia Mathematica, 17(1):67–69, February 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609090082O.

Martzloff:1990:SGF

[Mar90b] Jean-Claude Martzloff. Séminaire sur la géometrie des figures à travers les âges: Université d’Oran es-senia, institut des sciences exactes, Algérie, 1–3 juillet 1989. Historia Mathematica, 17(1):69–70, February 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609090083P.

Martzloff:1990:SJP

[Mar90c] Jean-Claude Martzloff. A survey of Japanese publications on the history of Japanese traditional mathematics (wasan) from the last 30 years. Historia Mathematica, 17(4):366–373, November 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090027B.

Marchisotto:1993:MPH

[Mar93a] Elena Anne Marchisotto. Mario Pieri and his contributions to geometry and foundations of mathematics. Historia Mathematica, 20 (3):285–303, August 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086083710244.

Martzloff:1993:ERR

[Mar93b] Jean-Claude Martzloff. Eléments de réflexion sur les réactions chinoises à la géométrie euclidienne à la fin du XVlle siècle — le Jihe lunyue à de Du Zhigeng {b} vu principalement à partir de la préface de l’auteur et deux notices bibliographiques rédigées par des lettres illustrés. Historia Mathematica, 20(2):160–179, May 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086083710141.

Martzloff:1995:BRF

[Mar95] Jean-Claude Martzloff. Book review: Fleeting footsteps: Tracing the conception of arithmetic and algebra in ancient China: By Lam Lay Yong and Ang Tian Se. Singapore (World Scientific). 1992. xvi + 199 pp. 15 £. Historia Mathematica, 22(1):67–73, February 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086085710063.

Martzloff:1997:NRC

[Mar97] Jean-Claude Martzloff. Note on the recent Chinese and Mongolian translations of Euclid’s Elements. Historia Mathematica, 24(2):200– 202, May 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086096921485.

Martini:1999:FLI

[Mar99] Laura Martini. The first lectures in Italy on Galois theory: Bologna, 1886–1887. Historia Mathematica, 26(3):201–223, August 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086099922498.

Martens:2000:KPN

[Mar00] Rhonda Martens. Kepler’s philosophy and the new astronomy. Princeton University Press, Princeton, NJ, USA, 2000. ISBN 0-691-05069-4. xii + 201 pp. LCCN QB355.3 .M37 2000. URL http://www.

loc.gov/catdir/description/prin022/00036681.html;   http:
//www.loc.gov/catdir/toc/fy02/00036681.html.

Martini:2004:ARS

[Mar04] Laura Martini. Algebraic research schools in Italy at the turn of the Twentieth Century: the cases of Rome, Palermo, and Pisa. Historia Mathematica, 31(3):296–309, August 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508600300082X.

Marchisotto:2006:PGM

[Mar06] Elena Anne Corie Marchisotto. The projective geometry of Mario Pieri: a legacy of Georg Karl Christian von Staudt. Historia Mathematica, 33(3):277–314, August 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086005000753.

Martini:2009:BRH

[Mar09] Laura Martini. Book review: A History of Abstract Algebra, Israel Kleiner. Birkhäuser, Boston (2007), ISBN 978-0-8176-4684-4. Historia Mathematica, 36(3):273–274, August 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008000700.

Maronne:2010:BRM

[Mar10a] Sébastien Maronne. Book review: Mathématiciens français du XVIIe siècle. Descartes, Fermat, Pascal. Edited by Michel Serfati, and Dominique Descotes. Clermont-Ferrand (Presses Universitaires Blaise Pascal). 2008. ISBN 978-2-84516-354-6. 284 pp. 20 EUR. Historia Mathematica, 37(4):712–714, November 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508600900072X.

Maronne:2010:CEI

[Mar10b] Sébastien Maronne. Contexts, emergence and issues of Cartesian geometry: In honour of Henk Bos’s 70th birthday. Historia Mathematica, 37(3):341–344, August 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086010000297.

Maronne:2010:OEM

[Mar10c] Sébastien Maronne. The ovals in the Excerpta Mathematica and the origins of Descartes’ method of normals. Historia Mathematica, 37(3):460–484, August 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086010000261.

Maronne:2014:BRS

[Mar14] Sébastien Maronne. Book review: Le sezione coniche nel Seicento, Luigi Maierù. Rubbettino Editore, Soveria Mannelli (2009), ISBN 978-88-498-2461-2. Historia Mathematica, 41(1):110–112, February 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086013000402.

Martins:2020:OHA

[Mar20] Ana Patrícia Martins. An overview on the history of actuarial calculus in Portugal until the late 19th century. Historia Mathematica, 51(??): 49–90, May 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508602030001X.

Massa:1997:MQP

[Mas97] Ma. Rosa Massa. Mengoli on “Quasi Proportions”. Historia Mathematica, 24(3):257–280, August 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086096921473.

MassaEsteve:2008:SLE

[Mas08] Ma. Rosa Massa Esteve. Symbolic language in early modern mathematics: the algebra of Pierre Hérigone (1580–1643). Historia Mathematica, 35(4):285–301, November 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508600800030X.

Matvievskaya:1974:BRA

[Mat74] G. P. Matvievskaya. Book review: Arabic arithmetic. The arithmetic of Abu al-Wafa’ al-Buzajani, 10th century: Mss. OR. 103 Leiden and 42 M Cairo. Edited with introduction, commentaries, and ample reference to the arithmetic of Al-Karaji (11th century), Ms. 855 Istanbul. By A. S. Saidan. Amman (Jordanian University), 1971. 468 pp. (Entirely in Arabic). Historia Mathematica, 1(1):108–111, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074901943.

Matvievskaya:1975:BRB

[Mat75] G. P. Matvievskaya. Book review: ‘Boethius’ Geometrie II: By Menso Folkerts. Boethius Texte und Abhandlungen zur Geschichte der Exakten Wissenschaften, Bank IX. Wiesbaden (Franz Steiner). 287 pp. DM.80. Historia Mathematica, 2(3):339–341, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900804.

Matvievskaya:1993:HMI

[Mat93] G. P. Matvievskaya. History of medieval Islamic mathematics: Research in Uzbekistan. Historia Mathematica, 20(3):239–246, August 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086083710207.

Mattmuller:2014:DBS

[Mat14] Martin Mattmüller. The difficult birth of stochastics: Jacob Bernoulli’s Ars Conjectandi (1713). Historia Mathematica, 41(3):277– 290, August 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086014000482.

Mawhin:2006:BRH

[Maw06] Jean Mawhin. Book review: Henri Lebesgue. Les lendemains de l’intégrale. Lettres à Emile Borel, Bernard Bru, Pierre Dugac (Eds.). Vuibert, Paris (2004), ISBN 2-7117-5309-3. Historia Mathematica, 33(4):496–497, November 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086005001023.

May:1975:HVL

[May75a] Kenneth O. May. Historiographic vices I. Logical attribution. Historia Mathematica, 2(2):185–187, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075901445.

May:1975:HVI

[May75b] Kenneth O. May. Historiographic vices II. Priority chasing. Historia Mathematica, 2(3):315–317, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075900725.

May:1975:RCH

[May75c] Kenneth O. May. Report of the commission on the history of mathematics to the division of history of the international union for the history and philosophy of science. Historia Mathematica, 2(1):77–82, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086075900488.

May:1975:WGH

[May75d] Kenneth O. May. What is good history and who should do it? Historia Mathematica, 2(4):449–455, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075901020.

Mayer:1976:DRT

[May76] Jean Mayer. Développements récents de la théorie des graphies. Historia Mathematica, 3(1):55–62, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086076900070.

May:1977:Aa

[May77a] Kenneth O. May. Abstracts. Historia Mathematica, 4(2):240–252, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086077901331.

May:1977:Ab

[May77b] Kenneth O. May. Abstracts. Historia Mathematica, 4(4):482–498, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086077901008.

May:1977:RCH

[May77c] Kenneth O. May. Report of the commission on the history of mathematics to the division of history of the international union of the history and philosophy of science August 1977. Historia Mathematica, 4(4):459–460, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086077900908.

Maziarz:1977:BRA

[Maz77] Edward A. Maziarz. Book review: Arts and sciences at Padua: the studium of Padua before 1350: By Nancy G. Sirasi. Toronto (Pontifical Institute of Mediaeval Studies). 1973. 199 p. Historia Mathematica, 4(2):227, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086077901252.

Mazliak:2016:BRP

[Maz16] Laurent Mazliak. Book review: Principes et Formules du Calcul des Probabilités pour assigner les limites des variations des événements naturels, Emmanuel-Etienne Duvillard de Durand. Editions de l’INED, Paris (2010). Historia Mathematica, 43(1):98–100, February 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086015000373.

Mazliak:2021:BRE

[Maz21] Laurent Mazliak. Book review: Émile Borel. À l’aube de la théorie des quanta : notes inédites d’Émile Borel sur un cours de Paul Langevin au Collège de France (1912–1913), Martha Cecilia Bustamante De La Ossa (Ed.). Brepols Publishers, Turnhout (2019), 20×26 cm. 65,00 EUR, ISBN: 978-2-503-58638-0. Historia Mathematica, 54 (??):120–122, February 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086020300707.

Mazliak:2024:BRM

[Maz24] Laurent Mazliak. Book review: Le Monde des mathématiques, Pierre-Michel Menger, Pierre Verschueren. Éditions du Seuil, Paris (2023), 832 pages, ISBN-10 2-02-153070-1, ISBN-13 978-2-02-153070-4. Historia Mathematica, 67(??):23–24, June 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086024000028.

Murray:2020:WFS

[MB20] Lori L. Murray and David R. Bellhouse. W. F. Sheppard’s correspondence with Karl Pearson and the development of his tables and moment estimates. Historia Mathematica, 53(??):108–117, November 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086020300458.

McCleary:1990:BRR

[McC90] John McCleary. Book review: Riemann, topology, and physics: By Michael Monastyrsky. Boston, Basel, Stuttgart (Birkhäuser). 1987. xiii + 158 pp. Historia Mathematica, 17(4):389–390, November 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090039G.

McCleary:1994:JRC

[McC94] John McCleary. On Jacobi’s remarkable curve theorem. Historia Mathematica, 21(3):377–385, August 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086084710305.

McGuigan:1974:BRH

[McG74] Robert McGuigan. Book review: A history of set theory: By Philip E. Johnson. Boston (Prindle, Weber & Schmidt), 1972. 118 pp. Historia Mathematica, 1(1):106–108, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074901931.

McKirahan:1982:BRT

[McK82] Richard D. McKirahan, Jr. Book review: Theon of Smyrna. Mathematics useful for understanding Plato: Translated from the 1892 Greek/French edition of J. Dupuis by Robert and Deborah Lawlor and edited and annotated by Christos Toulis and others. Secret Doctrine Reference Series. San Diego, CA (Wizards Bookshelf). 1970. ISBN 0-913510-24-6. xvi + 174 pp. $11.95. Historia Mathematica, 9(1):100–104, February 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086082901471.

McLarty:2012:BRG

[McL12] Colin McLarty. Book review: From a Geometrical Point of View: a Study of the History and Philosophy of Category Theory. By Jean-Pierre Marquis. Logic, Epistemology, and the Unity of Science. (Springer). 2010. ISBN 978-90-481-8117-9. 320 pp. $199. Historia Mathematica, 39(4):475–479, November 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086012000110.

Mehrtens:1976:KTM

[Meh76] Herbert Mehrtens. T. S. Kuhn’s theories and mathematics: a discussion paper on the “new historiography” of mathematics. Historia Mathematica, 3(3):297–320, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086076901002.

Mehrtens:1980:WSH

[Meh80] Herbert Mehrtens. Workshop on the social history of mathematics. Historia Mathematica, 7(1):75–79, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900671.

Melville:2002:WSA

[Mel02] Duncan J. Melville. Weighing stones in ancient Mesopotamia. Historia Mathematica, 29(1):1–12, February 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086001923286.

Melville:2004:PWM

[Mel04] Duncan J. Melville. Poles and walls in Mesopotamia and Egypt. Historia Mathematica, 31(2):148–162, May 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086003000776.

Melville:2006:BRL

[Mel06] Duncan J. Melville. Book review: Lengths, Widths, Surfaces. A Portrait of Old Babylonian Algebra and Its Kin, Jens Høyrup. Springer-Verlag, New York (2002), ISBN 0-387-95303-5. Historia Mathematica, 33(2):243–248, May 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086005000789.

Melville:2009:MAI

[Mel09] Duncan J. Melville. Mathematics in Ancient Iraq: a social history. Historia Mathematica, 36(4):428–433, November 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009000639.

Menghini:1986:NCB

[Men86] Marta Menghini. Notes on the correspondence between Luigi Cremona and Max Noether. Historia Mathematica, 13(4):341–351, November 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508608690056X.

Meo:2004:MLC

[Meo04] M. Meo. The mathematical life of Cauchy’s group theorem. Historia Mathematica, 31(2):196–221, May 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508600300003X.

Meskens:1994:WGL

[Mes94] Ad Meskens. Wine gauging in late 16th- and early 17th-Century Antwerp. Historia Mathematica, 21(2):121–147, May 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086084710135.

Masia-Fornos:2010:LPA

[MF10] Ramon Masià-Fornos. A “lacuna” in Proposition 9 of Archimedes’ On the Sphere and the Cylinder, Book I. Historia Mathematica, 37 (4):568–578, November 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086009000937.

Moore:1981:BFP

[MG81] Gregory H. Moore and Alejandro Garciadiego. Burali-Forti’s paradox: a reappraisal of its origins. Historia Mathematica, 8(3):319–350, August 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086081900707.

Miatello:2008:DPR

[Mia08] Luca Miatello. The difference \frac{51}{2} in a problem of rations from the Rhind mathematical papyrus. Historia Mathematica, 35(4):277–284, November 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008000311.

Michiwaki:1988:ISH

[Mic88] Yoshimasa Michiwaki. The International Symposium on the History and Education of Mathematics using Chinese Characters (ISHEM): Kiryu, Japan, August 7–9, 1987. Historia Mathematica, 15(3):280, August 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086088900717.

Michel:2023:RJL

[Mic23] Nicolas Michel. Review of Jesper Lützen: A History of Mathematical Impossibility. Oxford University Press (2022). Historia Mathematica, 65(??):33–35, November 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S031508602300068X.

Mikolas:1975:SHA

[Mik75] M. Mikolás. Some historical aspects of the development of mathematical analysis in Hungary. Historia Mathematica, 2(3):304–308, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900695.

Millan:1993:ISM

[Mil93] Ana Millán. International symposium mathematical journals. Centenary of El Progreso Matematico (1891–1991) Zaragoza, Spain, September 19–21, 1991. Historia Mathematica, 20(3):315–317, August 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008800018.

Melville:2009:Aa

[MMP09a] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 36(1):90–108, February 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008001018.

Melville:2009:Ab

[MMP09b] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 36(2):192–210, May 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086009000275.

Melville:2009:Ac

[MMP09c] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 36(3):293–313, August 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086009000287.

Melville:2009:Ad

[MMP09d] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 36(4):448–470, November 2009.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009000512.

Melville:2010:Aa

[MMP10a] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 37(2):317–338, May 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086010000029.

Melville:2010:Ab

[MMP10b] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 37(4):724–751, November 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086010000534.

Melville:2011:Aa

[MMP11a] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 38(1):136–159, February 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086010000765.

Melville:2011:Ab

[MMP11b] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 38(2):309–333, May 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086011000024.

Melville:2011:Ac

[MMP11c] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 38(3):429–452, August 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086011000292.

Melville:2012:Aa

[MMP12a] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 39(1):114–139, February 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086011000760.

Melville:2012:Ab

[MMP12b] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 39(2):237–262, May 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086012000067.

Melville:2012:Ac

[MMP12c] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 39(3):357–384, August 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086012000328.

Melville:2012:Ad

[MMP12d] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 39(4):480–503, November 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086012000389.

Melville:2013:Aa

[MMP13a] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 40(1):97–124, February 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086012000778.

Melville:2013:Ab

[MMP13b] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 40(2):210–231, May 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086013000049.

Melville:2013:Ac

[MMP13c] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 40(3):331–350, August 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086013000219.

Melville:2013:A

[MMP13d] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 40(4):461–480, October 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508601300044X.

Melville:2014:Aa

[MMP14a] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 41(1):116–136, February 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086013000505.

Melville:2014:Ab

[MMP14b] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 41(2):238–257, May 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086014000081.

Melville:2014:A

[MMP14c] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 41(3):345–364, August 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086014000500.

Melville:2015:Aa

[MMP15a] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 42(1):108–128, February 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086014000809.

Melville:2015:Ab

[MMP15b] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 42(2):238–259, May 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086014001025.

Melville:2015:Ac

[MMP15c] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 42(3):358–381, August 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086015000415.

Melville:2015:Ad

[MMP15d] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 42(4):495–515, November 2015.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086015000713.

Melville:2016:Aa

[MMP16a] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 43(1):108–126, February 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086015000786.

Melville:2016:Ab

[MMP16b] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 43(2):216–238, ???? 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508601600015X.

Melville:2016:Ac

[MMP16c] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 43(3):346–367, August 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086016300192.

Melville:2016:Ad

[MMP16d] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 43(4):452–476, November 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086016300416.

Melville:2017:Aa

[MMP17a] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 44(1):83–101, February 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086016300660.

Melville:2017:Ab

[MMP17b] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 44(2):179–197, May 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086017300125.

Melville:2017:Ac

[MMP17c] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 44(3):293–312, August 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086017300356.

Melville:2017:Ad

[MMP17d] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 44(4):436–456, November 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086017300526.

Melville:2018:Aa

[MMP18a] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 45(1):91–108, February 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086017301003.

Melville:2018:Ab

[MMP18b] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 45(2):196–213, May 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0315086018300053.

Melville:2018:A

[MMP18c] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 45(3):304–305, August 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086018300673.

Melville:2019:Aa

[MMP19a] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 46(??):104–130, February 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086018301484.

Melville:2019:Ab

[MMP19b] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 47(??):122–143, ???? 2019. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086019300011.

Melville:2019:Ac

[MMP19c] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 48(??):101–121, August 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086019300229.

Melville:2019:A

[MMP19d] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 49(??):94–114, November 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086019300424.

Melville:2020:Aa

[MMP20a] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 50(??):91–110, February 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086019300758.

Melville:2020:Ab

[MMP20b] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 51(??):94–115, May 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086020300057.

Melville:2020:Ac

[MMP20c] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 52(??):104–124, August 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086020300227.

Melville:2020:Ad

[MMP20d] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 53(??):132–146, November 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086020300501.

Melville:2021:Aa

[MMP21a] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 54(??):123–142, February 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086020300732.

Melville:2021:Ab

[MMP21b] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 55(??):98–114, May 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086021000021.

Melville:2021:Ac

[MMP21c] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 56(??):85–106, August 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086021000227.

Melville:2021:Ad

[MMP21d] Duncan J. Melville, Laura Martini, and Kim Plofker. Abstracts. Historia Mathematica, 57(??):97–116, November 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086021000434.

Malet:2023:E

[MO23] Antoni Malet and Jeffrey Oaks. Editorial. Historia Mathematica, 62(??):1–2, February 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086023000022.

Molland:1976:SFD

[Mol76] A. G. Molland. Shifting the foundations: Descartes’s transformation of ancient geometry. Historia Mathematica, 3(1):21–49, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900045.

Mollan:1981:PHP

[Mol81] Charles Mollan. Perspectives in the history and philosophy of science. Historia Mathematica, 8(2):190, May 1981. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508608190029X.

Mona:1975:BRG

[Mon75] A. F. Mona. Book review: Giuseppe Peano: By Hubert C. Kennedy. Basel (Birkhäuser Verlag). 1974. 31 p. (Beiheft 14 of Elemente der Mathematik). Historia Mathematica, 2(1):90–92, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900567.

Monna:1983:BRF

[Mon83] A. F. Monna. Book review: Felix Hausdorff: Ein mathematiker seiner Zeit: By Herbert Mehrtens, with forewords by Professor E. Brieskorn and E. Voigt. Bonn (Mathematisches Institut). 1980. 40 pp. Illustrated. Historia Mathematica, 10(4):483–484, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900228.

Montelle:2011:HAW

[Mon11] Clemency Montelle. ‘Having the answers’: Writing the history of mathematics in India. Historia Mathematica, 38(1):111–122, February 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508601000025X.

Montelle:2017:BRS

[Mon17] Clemency Montelle. Book review: The Siddhantasundara of Jñanaraja: an English Translation with Commentary, Toke Lindegaard Knudsen. Johns Hopkins University Press, Baltimore (2014). Historia Mathematica, 44(3):283–285, August 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086017300290.

Moore:1975:BRM

[Moo75a] Gregory H. Moore. Book review: Men and discoveries in mathematics: By Bryan Morgan. London (John Murray), 1972. 235 p. Photos, figures. U.K. £2. Historia Mathematica, 2(3):358–359, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900890.

Moore:1975:PBE

[Moo75b] Gregory H. Moore. A prospective biography of Ernst Zermelo (1871–1953). Historia Mathematica, 2(1):62–63, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900385.

Moore:1976:EZH

[Moo76] Gregory H. Moore. Ernst Zermelo, A. E. Harward, and the axiomatization of set theory. Historia Mathematica, 3(2):206–209, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900392.

Moore:1977:BRF

[Moo77] Gregory H. Moore. Book review: From Frege to Gödel: a source book in mathematical logic, 1879–1931: Edited by Jean van Heijenoort. Cambridge (Harvard University Press). 1971. xi + 660 pp. Second corrected printing. Historia Mathematica, 4(4):468–471, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900945.

Moore:1982:BRH

[Moo82a] Gregory H. Moore. Book review: A half century of Polish mathematics: Remembrances and reflections: By Kazimierz Kuratowski. Preface by Stanislaw Ulam. Oxford (Pergamon) and Warsaw (PWN, Polish Scientific Publishers). 1980. viii + 200 pp. 72 photographs. Historia Mathematica, 9(4):489–495, November 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086082901136.

Moore:1982:BRW

[Moo82b] Gregory H. Moore. Book review: Waclaw Sierpiński: Oeuvres choisies: Edited by Stanislaw Hartman, Kazimierz Kuratowski, Edward Marczewski, Andrezj Mostowski, Andrezj Schinzel, Roman Sikorski, and Marceli Stark. 3 vols. Warsaw (PWN). 1974–1976. 3 photographs. Historia Mathematica, 9(4):489–495, November 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082901148.

Moore:1995:ALA

[Moo95] Gregory H. Moore. The axiomatization of linear algebra: 1875–1940. Historia Mathematica, 22(3):262–303, ???? 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086085710257.

Moore:1997:MLP

[Moo97] Gregory H. Moore. Is mathematical logic a part of mathematics? Historia Mathematica, 24(2):210–212, May 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086097921571.

Moore:2002:HIR

[Moo02] Gregory H. Moore. Hilbert on the infinite: the role of set theory in the evolution of Hilbert’s thought. Historia Mathematica, 29(1):40– 64, February 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086001923328.

Moore:2003:BRK

[Moo03] Gregory H. Moore. Book review: From Kant to Hilbert: a Source Book in the Foundations of Mathematics: Selected and edited by William Ewald. Oxford (Clarendon Press). 1996. Two volumes. xviii + 1340 pp. with a personal name index. $300 US. Historia Mathematica, 30(4):516–519, November 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086002000101.

Moore:2007:ECH

[Moo07] Gregory H. Moore. The evolution of the concept of homeomorphism. Historia Mathematica, 34(3):333–343, August 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086006000863.

Moore:2008:EOS

[Moo08] Gregory H. Moore. The emergence of open sets, closed sets, and limit points in analysis and topology. Historia Mathematica, 35(3):220– 241, August 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008000050.

Morelon:1996:BRI

[Mor96] Régis Morelon. Book review: Islamic Astronomy and Medieval Spain. By Julio Samsó. Collected Studies Series CS 428. Aldershot, VT (Variorum). Historia Mathematica, 23(3):330–332, August 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086096900336.

Morgan:2019:BRH

[Mor19] Daniel Patrick Morgan. Book review: Heavenly Numbers: Astronomy and Authority in Early Imperial China, Christopher Cullen. Oxford University Press (2017), 426 + xiv pages. Historia Mathematica, 46(??):96–99, February 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086018301769.

Moss:1979:BLA

[Mos79] Barbara Moss. Beppo Levi and the axiom of choice. Historia Mathematica, 6(1):54–56, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079901046.

Mosley:2016:BRC

[Mos16] Derrick Mosley. Book review: The Correspondence of John Wallis, Volumes III and IV, Phillip Beeley, Christoph Scriba (Eds.). Oxford University Press (2012) (2014), ISBN: 978-0-19-856947-3. Historia Mathematica, 43(4):444–446, November 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086015000774.

Moyon:2017:BRA

[Moy17] Marc Moyon. Book review: From Alexandria, Through Baghdad. Surveys and Studies in the Ancient Greek and Medieval Islamic Mathematical Sciences in Honor of J.L. Berggren, Nathan Sidoli, Glen Van Brummelen (Eds.). Springer (2014). Historia Mathematica, 44(2):174–178, May 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086017300289.

Moyon:2019:LRN

[Moy19] Marc Moyon. The Liber Restauracionis: a newly discovered copy of a mediæval algebra in Florence. Historia Mathematica, 46(??):1–37, February 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086018300715.

Mozaffari:2013:WPC

[Moz13] S. Mohammad Mozaffari. Wabkanawi’s prediction and calculations of the annular solar eclipse of 30 January 1283. Historia Mathematica, 40(3):235–261, August 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086013000207.

Malet:2006:OAA

[MP06] Antoni Malet and Marco Panza. The Origins of Algebra: From al-Khwarizmi to Descartes: International Workshop held at Barcelona, 27–29 March 2003. Historia Mathematica, 33(1):1–3, February 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005000972.

Melville:2022:Aa

[MP22a] Duncan J. Melville and Kim Plofker. Abstracts. Historia Mathematica, 58(??):120–132, February 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086021000677.

Melville:2022:Ab

[MP22b] Duncan J. Melville and Kim Plofker. Abstracts. Historia Mathematica, 59(??):197–214, May 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086022000052.

Melville:2022:Ac

[MP22c] Duncan J. Melville and Kim Plofker. Abstracts. Historia Mathematica, 60(??):69–96, August 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086022000453.

Melville:2022:Ad

[MP22d] Duncan J. Melville and Kim Plofker. Abstracts. Historia Mathematica, 61(??):80–96, November 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086022000684.

Melville:2023:Aa

[MP23a] Duncan J. Melville and Kim Plofker. Abstracts. Historia Mathematica, 62(??):73–91, February 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086023000010.

Melville:2023:Ab

[MP23b] Duncan J. Melville and Kim Plofker. Abstracts. Historia Mathematica, 63(??):75–91, May 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086023000137.

Melville:2023:Ac

[MP23c] Duncan J. Melville and Kim Plofker. Abstracts. Historia Mathematica, 64(??):57–75, August 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086023000149.

Melville:2023:Ad

[MP23d] Duncan J. Melville and Kim Plofker. Abstracts. Historia Mathematica, 65(??):36–53, November 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508602300037X.

Melville:2024:Aa

[MP24a] Duncan J. Melville and Kim Plofker. Abstracts. Historia Mathematica, 67(??):32–54, June 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086024000314.

Melville:2024:Ab

[MP24b] Duncan J. Melville and Kim Plofker. Abstracts. Historia Mathematica, 68(??):51–76, September 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086024000752.

Mrozik:2016:BRP

[Mro16] Dagmar Mrozik. Book review: Pearls from a Lost City: the Lvov School of Mathematics, vol. 40, Roman Duda, in: History of Mathematics. American Mathematical Society (2014), ISBN: 978-1-4704-1076-6. Historia Mathematica, 43(2):207–209, ???? 2016.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086015000622.

Malcolm:2005:JPH

[MS05] Noel Malcolm and Jacqueline A. Stedall. John Pell (1611– 1685) and his correspondence with Sir Charles Cavendish: the mental world of an early mathematician. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-856484-8 (hardcover). 657 pp. LCCN QA29.P6725 A4

2005. URL http://www.h-net.org/review/hrev-a0e7r9-aa;
http://www.loc.gov/catdir/enhancements/fy0627/2005297429-
d.html;   http://www.loc.gov/catdir/enhancements/fy0627/
2005297429-t.html.

Malet:2022:E

[MS22] Antoni Malet and Nathan Sidoli. Editorial. Historia Mathematica, 58(??):1–2, February 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086022000015.

Mueller:1980:BRE

[Mue80] Ian Mueller. Book review: Von Eudoxos zu Aristoteles. Das Fortwirken der eudoxischen Proportionentheorie in der aristotelischen Lehre vom Kontinuum: (Studien zur Antiken Philosophie, Bd. 8). By Hans-Joachim Waschkies. Verlag B. R. Grüner, 1977. 453 pp. Hfl. 90. Historia Mathematica, 7(2):200–210, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900531.

Mueller:1996:BRE

[Mue96] Ian Mueller. Book review: Euclide d’Alexandrie, Les Éléments. Traduction et commentaires par Bernard Vitrac. Paris (Presses universitaires de France). vol. I, livres I–IV avec une introduction générale par Maurice Caveing. 1990. Historia Mathematica, 23(1):85–87, February 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086096900075.

Mugnai:1987:IWL

[Mug87] Massimo Mugnai. International workshop on the Leibniz Renaissance: the reasons for contemporary interest in Leibniz’ thought. Historia Mathematica, 14(1):61–62, February 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086087900073.

Murdoch:2007:MC

[Mur07] John E. Murdoch. Marshall Clagett (1916–2005). Historia Mathematica, 34(3):261–265, August 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S031508600600125X.

Muusepp:1975:DLE

[Müü75] Peeter Müüsepp. D’Alembert’s letter to Euler of 3 March 1766. Historia Mathematica, 2(3):309–311, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075900701.

Muursepp:1977:GLF

[Müü77] Peeter Müürsepp. Gauss’ letter to Fuss of 4 April 1803. Historia Mathematica, 4(1):37–41, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086077900349.

Muursepp:1978:GTU

[Muü78] Peeter Muür̈sepp. Gauss and Tartu University. Historia Mathematica, 5(4):455–459, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086078902148.

Mansfield:2017:PBE

[MW17] Daniel F. Mansfield and N. J. Wildberger. Plimpton 322 is Babylonian exact sexagesimal trigonometry. Historia Mathematica, 44 (4):395–419, November 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086017300691.

Matvievskaya:1999:MAM

[MY99] Galina Pavlovna Matvievskaya and Gulnara Eskenderovna Yusupova. The mathematical and astronomical manuscripts in the collection of the Institute of Oriental Studies of the Academy of Sciences of Uzbekistan at Tashkent. Historia Mathematica, 26(4):378–381, November 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086099922437.

Netz:2011:AP

[N^{+}11] Reviel Netz et al., editors. The Archimedes palimpsest. The Archimedes palimpsest publications. Cambridge University Press, Cambridge, UK; New York, NY, USA, 2011. ISBN 1-107-01684-3 (hardcover: set), 1-107-01457-3 (hardcover: vol. 1), 1-107-01437-9 (hardcover: vol. 2). ???? pp. LCCN QA31 .A832 2011. URL http:/

/assets.cambridge.org/97811070/14572/cover/9781107014572.

jpg. Published for the Walters Art Museum.

Nabonnand:2004:BRH

[Nab04] Philippe Nabonnand. Book review: Hermann Weyl’s Raum– Zeit–Materie and a General Introduction to His Scientific Work: Edited by Erhard Scholz. DMV-Seminar 30. Basel/Boston/Berlin (Birkhäuser). 2001. 403 pp. Historia Mathematica, 31(3):371–375, August 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086004000436.

Nagel:1982:CPP

[Nag82] Ernest Nagel. Charles Peirce’s place in philosophy. Historia Mathematica, 9(3):302–310, August 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086082901240.

Nagel:1988:ACD

[Nag88] Fritz Nagel. 1687/1987: Der Ausbau des Calculus durch Leibniz und die Brüder Bernoulli: Symposion, Basel, June 15–17, 1987. Historia Mathematica, 15(1):81–82, February 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900547.

Nagel:1991:CWJ

[Nag91a] Fritz Nagel. A catalog of the works of Jacob Hermann (1678–1733). Historia Mathematica, 18(1):36–54, February 1991. CODEN HI-MADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0315086091903507.

Nagel:1991:LWJa

[Nag91b] Fritz Nagel. The life and work of Jacob Hermann: a Working Conference in Pisa, March 5–6, 1990. Historia Mathematica, 18(1):64–65, February 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609190354Z.

Nagel:1991:LWJb

[Nag91c] Fritz Nagel. The Life and Work of Jacob Hermann: a Working Conference, part 2 (Ravello, September 28–29, 1990), and part 3 (Pisa, February 25–26, 1991). Historia Mathematica, 18(4):357, November 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086091903747.

Nahin:2006:DEF

[Nah06] Paul J. Nahin. Dr. Euler’s Fabulous Formula: Cures Many Mathematical Ills. Princeton University Press, Princeton, NJ, USA, 2006. ISBN 0-691-11822-1 (hardcover). xx + 380 pp.

LCCN QA255 .N339 2006. URL http://www.loc.gov/catdir/
enhancements/fy0654/2005056550-d.html;   http://www.loc.
gov/catdir/enhancements/fy0734/2005056550-b.html. The Eu-

ler formula, e^{iπ} + 1 = 0, relates five important mathematical constants, and the digits of the binary number system.

Nakane:2003:BRM

[Nak03] Michiyo Nakane. Book review: The Mathematical Papers of Sir William Rowan Hamilton: Edited by Brendan K. P. Scaife. Cambridge (Cambridge University Press). 2000. pp. viii + 842. Historia Mathematica, 30(4):514–516, November 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086003000594.

Nauenberg:1998:HMO

[Nau98] Michael Nauenberg. On Hooke’s 1685 manuscript on orbital mechanics. Historia Mathematica, 25(1):89–93, February 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086097921911.

Nauenberg:2003:KAL

[Nau03] Michael Nauenberg. Kepler’s area law in the Principia: filling in some details in Newton’s proof of Proposition 1. Historia Mathematica, 30 (4):441–456, November 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086002000277.

Novikoff:1977:BLN

[NB77] Albert Novikoff and Jack Barone. The Borel Law of Normal Numbers, the Borel Zero–One Law, and the work of Van Vleck. Historia Mathematica, 4(1):43–65, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086077900362.

Needham:1979:BRC

[Nee79] Joseph Needham. Book review: A critical study of the Yang Hui Suan Fa: a thirteenth-century Chinese mathematical treatise: By Lam Lay Yong (L. Y. Lam). Singapore (University Press). 1977. xx + 360 pp. Historia Mathematica, 6(4):466–468, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079900429.

Netz:2002:BRR

[Net02] Reviel Netz. Book review: La rivoluzione dimenticata. Il pensiero scientifico greco e la scienza moderna. By Lucio Russo. Historia Mathematica, 29(1):72–73, February 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086001923109.

Netz:2024:BRF

[Net24] Reviel Netz. Book review: Francisci Maurolyci Archimedea, curaverunt Riccardo Bellé, Pier Daniele Napolitani, Beatrice Sisana. Adnotationes anglice interpretatus est Walter Roy Laird. Tomus A. Opera Archimedis ex traditione Maurolyci, curaverunt Riccardo Bellé, Pier Daniele Napolitani, Beatrice Sisana, Pisa, Fabrizio Serra Editore 2022, pp. LXXIV-474, Tomus B. De momentis, curaverunt Riccardo Bellé, Pier Daniele Napolitani, Beatrice Sisana, adiuvante Roberta Tucci, Pisa, Fabrizio Serra Editore 2022, pp. XIV–186, ISBN: Tomus A: 978-88-3315-476-3, Tomus B: 978-88-3315-478-7. Historia Mathematica, 69(??):65–66, December 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086024000880.

Neuenschawander:1986:BRC

[Neu86] Erwin Neuenschawander. Book review: Carl Friedrich Gauss: a bibliography: Compiled by Uta C. Merzbach. Wilmington, Delaware (Scholarly Resources Inc.). 1984. xxv + 551 pp. $95.00. Historia Mathematica, 13(3):297–299, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086086901023.

Neuenschwander:1988:BRN

[Neu88] Erwin Neuenschwander. A brief report on a number of recently discovered sets of notes on Riemann’s lectures and on the transmission of the Riemann Nachlass [legacy]. Historia Mathematica, 15(2):101– 113, May 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086088900018.

Neuenschwander:1989:UPJ

[Neu89] Erwin Neuenschwander. The unpublished papers of Joseph Liouville in Bordeaux. Historia Mathematica, 16(4):334–342, November 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086089900803.

Neuenschwander:1993:NEB

[Neu93] Erwin Neuenschwander. Der Nachlass von Erich Bessel-Hagen im archiv der Universität Bonn. (German) [The legacy of Erich Bessel-Hagen in the archive of the University of Bonn]. Historia Mathematica, 20(4):382–414, November 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086083710323.

Neumann:2008:BRE

[Neu08] Peter M. Neumann. Book review: Episodes in the History of Modern Algebra, vol. 32, Jeremy J. Gray, Karen Hunger Parshall (Eds.), in: History of Mathematics. American Mathematical Society, London Mathematical Society, Providence and London (2007), ISBN 978-0-8218-4343-7. Historia Mathematica, 35(4):337–338, November 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508600800027X.

Neumann:2012:EEW

[Neu12] Peter M. Neumann. The editors and editions of the writings of Évariste Galois. Historia Mathematica, 39(2):211–221, May 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086012000043.

Neumann:2015:JAS

[Neu15] Peter M. Neumann. Jacqueline Anne Stedall (4 August 1950– 27 September 2014). Historia Mathematica, 42(1):5–13, February 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086014001049.

Neyman:1978:ATP

[Ney78] Jerzy Neyman. About the theory of probability and mathematical statistics. Historia Mathematica, 5(4):485–486, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508607890229X.

Nicholson:1993:DUC

[Nic93] Julia Nicholson. The development and understanding of the concept of quotient group. Historia Mathematica, 20(1):68–88, February 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086083710074.

Nickerson:2013:BRT

[Nic13] Sylvia Nickerson. Book review: Turbulent Times in Mathematics. The Life of J. C. Fields and the History of the Fields Medal, Elaine McKinnon Riehm, Frances Hoffman. American Mathematical Society and The Fields Institute, Providence, Rhode Island and Toronto, Ontario (2011), ISBN: 978-0-8218-6914-7. Historia Mathematica, 40(3):325–329, August 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S031508601200078X.

Nikolic:2009:SMK

[Nik09] Aleksandar Nikolić. The story of majorizability as Karamata’s condition of convergence for Abel summable series. Historia Mathematica, 36(4):405–419, November 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086009000500.

Nikolic:2018:KFD

[Nik18] Aleksandar Nikolić. Karamata functions and differential equations: Achievements from the 20th century. Historia Mathematica, 45 (3):237–276, August 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086017301155.

Nissen:1989:WBB

[Nis89] Hans J. Nissen. Workshop über Begriffsbildungen in der Babylonischen Mathematik. (German) [Workshop on concepts in Babylonian mathematics]. Historia Mathematica, 16(1):92–93, February 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086089901043.

Navarro-Loidi:2008:ILS

[NLL08] Juan Navarro-Loidi and José Llombart. The introduction of logarithms into Spain. Historia Mathematica, 35(2):83–101, May 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086007000626.

Norton:1975:HBS

[Nor75] B. J. Norton. History of the biometric school of statistics and its influences on the social sciences. Historia Mathematica, 2(2):194– 196, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086075901482.

North:2003:BRJ

[Nor03] J. D. North. Book review: James Joseph Sylvester. Life and Work in Letters: By Karen Hunger Parshall. Oxford (Clarendon). 1998. 337 pp. $98. Historia Mathematica, 30(1):88–90, February 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086002000137.

Nordmann:2007:BRM

[Nor07] Alfred Nordmann. Book review: Mechanistic Images in Geometric Form: Heinrich Hertz’s Principles of Mechanics, Jesper Lützen. Oxford University Press, Oxford (2005), ISBN 0-19-856737-5. Historia Mathematica, 34(3):349–352, August 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086006000887.

Nossum:2012:EMO

[Nos12] Rolf Nossum. Emigration of mathematicians from outside Germanspeaking academia 1933–1963, supported by the Society for the Protection of Science and Learning. Historia Mathematica, 39(1):84–104, February 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086011000498.

Nastasi:1994:PMS

[NS94] P. Nastasi and A. Scimone. Pietro Mengoli and the six-square problem. Historia Mathematica, 21(1):10–27, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086084710056.

Nastasi:2005:TSP

[NT05] Pietro Nastasi and Rossana Tazzioli. Toward a scientific and personal biography of Tullio Levi-Civita (1873–1941). Historia Mathematica, 32(2):203–236, May 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086004000229.

Nurzia:1992:RGH

[Nur92] Laura Nurzia. The research group on history of mathematics at the university of Rome “La Sapienza”. Historia Mathematica, 19(3):296– 302, August 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609290038D.

Oaks:2005:MEP

[OA05] Jeffrey A. Oaks and Haitham M. Alkhateeb. Mal, enunciations, and the prehistory of Arabic algebra. Historia Mathematica, 32 (4):400–425, November 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086005000352.

Oaks:2007:SEA

[OA07] Jeffrey A. Oaks and Haitham M. Alkhateeb. Simplifying equations in Arabic algebra. Historia Mathematica, 34(1):45–61, February 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508600600022X.

Oaks:2012:BRL

[Oak12] Jeffrey A. Oaks. Book review: L’Eredit arabo-islamica nelle scienze e nelle arti del calcolodell’Europa medievale (The Arabic-Islamic heritage in the sciences and arts of calculation in Medieval Europe). By Nadia Ambrosetti. Milano (Edizioni Universitarie di Lettere Economia Diritto). 2008. ISBN 978-88-7916-388-0. 407 pp. 37 EUR. Historia Mathematica, 39(4):460–462, November 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086012000286.

Oaks:2019:PAA

[Oak19] Jeffrey A. Oaks. Proofs and algebra in al-Farisi’s commentary. Historia Mathematica, 47(??):106–121, ???? 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086018301393.

Oberle:1978:CIJ

[Obe78] R. Oberlé. Le Colloque International Jean Henri Lambert 26– 30 Septembre 1977. Historia Mathematica, 5(3):346–348, August 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078901222.

Ohrstrom:1985:WRH

[Øhr85] Peter Øhrstrøm. W. R. Hamilton’s view of algebra as the science of pure time and his revision of this view. Historia Mathematica, 12(1):45–55, February 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086085900679.

Oliveri:2008:BRA

[Oli08] Gianluigi Oliveri. Book review: The Architecture of Modern Mathematics: Essays in History and Philosophy, José Ferreirós, Jeremy Gray. Oxford University Press, Oxford (2006), ISBN 978-0-19-856793-6. Historia Mathematica, 35(1):57–60, February 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086007000286.

Olson:1976:BRT

[Ols76] Richard G. Olson. Book review: Thomas Reid’s inquiry: the geometry of visibles and the case for realism: By Norman Daniels with a foreward by Hilary Putnam. New York (Burt Franklin and Co.). 1974. xix + 147 pp. Historia Mathematica, 3(4):486–487, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900859.

Pherau:2009:GME

[òP09] òThakkura Pherāu, editor. Gaònitasāarakaumudāi = The moonlight of the essence of mathematics. Manohar Publishers and Distributors, New Delhi, India, 2009. ISBN 81-7304-809-6. xlvi + 278 pp. LCCN QA27.I4 T4613 2009.

Oravas:1982:BRG

[Ora82] Gunhard A. E. Oravas. Book review: Geschichte der mechanischen Prinzipien und ihrer wichtigsten Anwendungen: By István Szabó. Basel and Stuttgart (Birkhäuser Verlag), 1977. xv + 491 pp. Historia Mathematica, 9(1):81–94, February 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086082901434.

Orlova:2019:LLR

[OS19] Nadezda Orlova and Sergei Soloviev. Logic and logicians in Russia before 1917: Living in a wider world. Historia Mathematica, 46(??):38–55, February 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086018300612.

Theon:1979:MUU

[oSTT79] of Smyrna Theon and Christos Toulis. Mathematics useful for understanding Plato. Secret doctrine reference series. Wizards Bookshelf, San Diego, CA, USA, 1979. ISBN 0-913510-24-6. xvi + 174 pp. LCCN QA31 .T4813 1979.

Oswald:2017:UPE

[Osw17] Nicola M. R. Oswald. An unpublished paper ‘Über einige durch unendliche Reihen definirte Functionen eines complexen Argumentes’ by Adolf Hurwitz. Historia Mathematica, 44(3):252–279, August 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086017300265.

Otte:1988:SSC

[Ott88] Michael Otte. Symposium on “ ‘Style’ as a Category of the History and Philosophy of Science”: Brakel, Federal Republic of Germany, June 17–19, 1987. Historia Mathematica, 15(1):83–84, February 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900559.

Otte:1989:IHG

[Ott89] Michael Otte. The ideas of Hermann Grassmann in the context of the mathematical and philosophical tradition since Leibniz. Historia Mathematica, 16(1):1–35, February 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086089900967.

Ozdural:2000:MAC

[Özd00] Alpay Özdural. Mathematics and arts: Connections between theory and practice in the medieval Islamic world. Historia Mathematica, 27(2):171–201, May 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086099922747.

Panteki:1987:WWI

[Pan87] M. Panteki. William Wallace and the introduction of continental calculus to Britain: a letter to George Peacock. Historia Mathematica, 14(2):119–132, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086087900164.

Panza:1988:DFD

[Pan88] Marco Panza. Le débat sur les fondements du calcul différentiel en France dans la seconde moitié du XVIIIe siècle: L’hypothèse réductionniste et la naissance de l’algèbre opératoire. Historia Mathematica, 15(2):172–173, May 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900110.

Panteki:2003:FLB

[Pan03] Maria Panteki. French “logique” and British “logic”: on the origins of Augustus de Morgan’s early logical inquiries, 1805–1835. Historia Mathematica, 30(3):278–340, August 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086003000259.

Panza:2005:NOA

[Pan05] Marco Panza. Newton et les origines de l’analyse, 1664–1666. (French) [Newton and the origins of analysis, 1664–1666]. Collection Sciences dans l’histoire, 1258-0996. A. Blanchard, Paris, France, 2005. ISBN 2-85367-230-1. xvii + 548 pp. LCCN QA300 .P3464

Panza:2008:BRI

[Pan08] Marco Panza. Book review: Incógnitas, variables y otros fantasmas matemáticos. El lenguaje de la variación y la covariación numéricas, Angel Marín. Universidad Pública de Navarra, Pamplona (2006), ISBN 84-9769-173-3. Historia Mathematica, 35(4):338–339, November 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008000232.

Panza:2009:BRJ

[Pan09] Marco Panza. Book review: John Wallis. Una vita per un progetto, Luigi Maierù. Rubettino Editore, Soveria Mannelli (2007), ISBN 978-88-498-1904-5. Historia Mathematica, 36(3):279–281, August 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009000032.

Panza:2011:RGE

[Pan11] Marco Panza. Rethinking geometrical exactness. Historia Mathematica, 38(1):42–95, February 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086010000558.

Panza:2018:BRM

[Pan18] Marco Panza. Book review: Meccanicismo. Trionfi e miserie della visione meccanica del mondo, Giorgio Israel. Zaniclelli, Bologna (2015), pp. xiv + 322; 28 EUR, ISBN: 978-88-08-52129-3. Historia Mathematica, 45(2):186–189, May 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL https://www.sciencedirect.

com/science/article/pii/S0315086018300351.

Parshall:1989:BRA

[Par89] Karen Hunger Parshall. Book review: Ada: a life and a legacy: (MIT Press Series in the History of Computing, Vol. 3). By Dorothy Stein. Cambridge, MA/London (The MIT Press). 1985. xix + 321 pp. $9.95. Historia Mathematica, 16(1):94–95, February 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086089901055.

Parshall:1991:BRN

[Par91] Brian J. Parshall. Book review: Nathan Jacobson: Collected mathematical papers: By Nathan Jacobson. Boston, Basel, Berlin (Birkhäuser). 1989. 3 Vols. xviii + 1606 pp. Historia Mathematica, 18(4):370–372, November 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609190379C.

Parshall:1992:BRC

[Par92a] Karen Hunger Parshall. Book review: Convolutions in French mathematics, 1800–1840: By Ivor Grattan-Guinness. Science Networks: Historical Studies, Vol. 2. Basel, Boston, Berlin (Birkhäuser). 1990. 3 Vols. 1602 pp. DM 398. Historia Mathematica, 19(3):303–310, August 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290040I.

Parshall:1992:BRE

[Par92b] Karen Hunger Parshall. Book review: Ethnomathematics: a multicultural view of mathematical ideas: By Marcia Ascher. Pacific Grove, CA (Brooks/Cole Publishing Co.). 1991. xii + 203 pp. including Index. Historia Mathematica, 19(3):310, August 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086092900419.

Parshall:1995:NME

[Par95] Karen V. H. Parshall. Note from the managing editor. Historia Mathematica, 22(3):225, ???? 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086085710233.

Parshall:1997:BRA

[Par97] Karen Hunger Parshall. Book review: Algebra e geometria (1860– 1940): Il contributo italiano. Edited by Aldo Brigaglia, Ciro Ciliberto, and E. Sernesi. Supplemento ai Rendiconti del Circolo matematico di Palermo, Ser. 2, No. 36 (1994). Palermo (Circolo matematico di Palermo). 1994. 277 pp. Historia Mathematica, 24(3):334–339, August 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086096921424.

Parshall:2002:GBS

[Par02] Karen Hunger Parshall. George Boole: Selected manuscripts on logic and its philosophy. Historia Mathematica, 29(4):491–493, November 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086002923724.

Parshall:2003:ICH

[Par03] Karen Hunger Parshall. The international commission for the history of mathematics (ICHM): an introduction and a call to the international community of historians of mathematics. Historia Mathematica, 30(1):1–2, February 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086002000265.

Parshall:2004:DMR

[Par04] Karen Hunger Parshall. Defining a mathematical research school: the case of algebra at the University of Chicago, 1892–1945. Historia Mathematica, 31(3):263–278, August 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508600300048X.

Parker:2005:RMM

[Par05a] John Parker. R. L. Moore: mathematician and teacher. Spectrum series. Mathematical Association of America, Washington, DC, USA, 2005. ISBN 0-88385-550-X. xiv + 387 pp. LCCN QA29.M66 P37 2005.

Parshall:2005:FAK

[Par05b] Karen Parshall. The fifth award of the Kenneth O. May Medal and Prize. Historia Mathematica, 32(4):397–399, November 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005000741.

Parshall:2007:BRR

[Par07] Karen Hunger Parshall. Book review: R. L. Moore: Mathematician and Teacher, John Parker. Mathematical Association of America (2005), ISBN 0-88385-550-X. Historia Mathematica, 34(4):448–451, November 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508600600098X.

Patterson:1975:LJS

[Pat75] Andrew M. Patterson. The Leonard J. Savage papers. Historia Mathematica, 2(1):70–72, February 1975. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075900427.

Patterson:1990:EQE

[Pat90] S. J. Patterson. Eisenstein and the quintic equation. Historia Mathematica, 17(2):132–140, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609090049J.

Patton:2014:HO

[Pat14] Lydia Patton. Hilbert’s objectivity. Historia Mathematica, 41(2):188– 203, May 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086014000032.

Polo-Blanco:2008:ABS

[PB08] Irene Polo-Blanco. Alicia Boole Stott, a geometer in higher dimension. Historia Mathematica, 35(2):123–139, May 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086007000973.

Peckhaus:2020:BRK

[Pec20] Volker Peckhaus. Book review: Kurt Gödel: Collected Works. Vol. IV: Correspondence A-G. Vol. V: Correspondence H-Z, Solomon Feferman (Ed.). Clarendon Press, Oxford (2003), Paperback 2014. Vol. IV (pbk). xxii + 662 pp. + 12 ill. Vol. V: ISBN 978-0-19-968962-0 (pbk). xxvi + 664 pp. + 15 ill., ISBN: 978-0-19-968961-3. Historia Mathematica, 50(??):87–90, February 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086019300746.

Pedoe:1975:ADW

[Ped75a] Dan Pedoe. Ausz disem wirdt vil dings gemacht [(German) [for this is done much recently]]: a Dürer construction for tangent circles. Historia Mathematica, 2(3):312–314, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075900713.

Pedoe:1975:BRE

[Ped75b] Daniel Pedoe. Book review: Euclid and his modern rivals: By Lewis Carroll (Charles L. Dodgson). New York (Dover Publications). 1973 275 p. U.S. $3.50. Historia Mathematica, 2(2):219–222, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075901585.

Pedersen:1977:BRT

[Ped77] Kurt Møller Pedersen. Book review: Thomas Hobbes: De homine. Traité de l’homme: Traduction et commentaire par Paul-Marie Maurin. Preface par Vasco Ronchi. Paris (Albert Blanchard). 1974. 204 pp. 30 Frs. Historia Mathematica, 4(4):471–472, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900957.

Peetre:1995:OSB

[Pee95] Jaak Peetre. Outline of a scientific biography of Ernst Meissel (1826–1895). Historia Mathematica, 22(2):154–178, May 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086085710154.

Peiffer:2015:BRM

[Pei15] Jeanne Peiffer. Book review: Das mathematische und naturphilosophische Lernen und Arbeiten der Marquise du Châtelet (1706-1749). Wissenszugänge einer Frau im 18. Jahrhundert, Frauke Böttcher, in: Mathematik im Kontext. Springer Spektrum, Berlin, Heidelberg (2013), ISBN: 978-3-642-32486-4. Historia Mathematica, 42(1):95–98, February 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086014000524.

Peiffer:2017:BRP

[Pei17] Jeanne Peiffer. Book review: Principes mathématiques de la philosophie naturelle. La traduction française des Philosophiae naturalis principia mathematica. Centre international d’étude du XVIII ^{e} siècle, Ferney-Voltaire (2015), ISBN: 978-2-84559-115-8. Historia Mathematica, 44(2):170–172, May 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086017300332.

Peiffer:2022:LEO

[Pei22] Jeanne Peiffer. Leonhardi Euleri Opera omnia, Series IV A : Commercium Epistolicum, Vol. 8 Ediderunt Andreas Kleinert et Thomas Steiner, auxilio Gisela Kleinert et Martin Mattmüller. Basel (Birkhäuser). 2018. Historia Mathematica, 60(??):66–68, August 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086021000847.

Pejlare:2007:TBW

[Pej07] Johanna Pejlare. Torsten Brodén’s work on the foundations of Euclidean geometry. Historia Mathematica, 34(4):402–427, November 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086007000055.

Pepper:1975:PEL

[Pep75] Jon V. Pepper. The proposed edition of the life and works of Thomas Harriot (1560–1621). Historia Mathematica, 2(1):64–65, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900397.

Pepe:1990:EMD

[Pep90] Luigi Pepe. European mathematics during the Napoleonic Period: Istituto Nazionale di Alta Matematica, Cortona, Italy, October 9–14, 1989. Historia Mathematica, 17(3):261–262, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090007Z.

Pepe:1992:EMD

[Pep92] Luigi Pepe. European mathematics during the XVth Century: Istituto Nazionale di alta matematica, Cortona, Italy, June 24–28, 1991. Historia Mathematica, 19(2):206–207, May 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609290082M.

Perl:1979:LDW

[Per79] Teri Perl. The Ladies Diary or Woman’s Almanack, 1704–1841. Historia Mathematica, 6(1):36–53, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079901034.

Perl:1980:WML

[Per80] Teri Hoch Perl. Women and mathematics in the late 18th century: a query. Historia Mathematica, 7(2):188–189, May 1980.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900385.

PereiraDaSilva:1993:RGH

[Per93] Clóvis Pereira Da Silva. The research group on history of mathematics at the Federal University of Paraná. Historia Mathematica, 20(3):318–319, August 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086083710281.

Perry:2010:BRM

[Per10] Andrew Perry. Book review: Mathematical Works Printed in the Americas, 1554–1700. By Bruce Stanley Burdick. Baltimore (The Johns Hopkins University Press). 2009. ISBN 978-0-8018-8823-6. 392 pp. $55.00. Historia Mathematica, 37(2):308–309, May 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508600900130X.

Pesic:2001:VNL

[Pes01] Peter Pesic. The validity of Newton’s lemma 28. Historia Mathematica, 28(3):215–219, August 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086001923213.

Pesic:2003:APE

[Pes03] Peter Pesic. Abel’s proof: an essay on the sources and meaning of mathematical unsolvability. MIT Press, Cambridge, MA, USA, 2003. ISBN 0-262-16216-4. viii + 213 pp. LCCN QA212 .P47 2003. URL

http://www.loc.gov/catdir/toc/fy037/2002031991.html.

Petrova:1974:HDA

[Pet74] S. S. Petrova. Sur l’histoire des démonstrations analytiques du théorème fondamental de l’algèbre. Historia Mathematica, 1(3):255– 261, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074900627.

Peters:1978:GWL

[Pet78a] J. F. Peters. Gottfried Wilhelm Leibniz. Historia Mathematica, 5(4):474, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086078902239.

Peters:1978:ITR

[Pet78b] J. F. Peters. With informal thoughts regarding the use and improvement of the German language. Historia Mathematica, 5(4):474–478, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086078902240.

Peters:1978:HMC

[Pet78c] James V. Peters. A history of mathematics course for undergraduates. Historia Mathematica, 5(4):463–464, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078902161.

Peters:1982:BRL

[Pet82] J. F. Peters. Book review: Leibniz à Paris (1672–1676): Symposium of the G. W. Leibniz-Society (Hannover) and of the National Center for Scientific Research (Paris), at Chantilly (France), 14–18 November 1976. Studia Leibnitiana Supplementa, Vol. XVII. (t. 1): Les sciences. Edited by A. Heinekamp and D. Metter. Wiesbaden, Franz Steiner Verlag, 1978. 242 pp. Historia Mathematica, 9(1):113–123, February 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086082901525.

Petruso:1985:APP

[Pet85] Karl M. Petruso. Additive progression in prehistoric mathematics: a conjecture. Historia Mathematica, 12(2):101–106, May 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086085900011.

Petry:1991:HN

[Pet91] M. J. Petry. Hegel and Newtonianism. Historia Mathematica, 18 (4):357–361, November 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086091903758.

Petrunic:2009:KLL

[Pet09a] Josipa G. Petrunić. Kelvin: Life, labours and legacy. Historia Mathematica, 36(4):439–444, November 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086009000597.

Petsche:2009:HGB

[Pet09b] Hans-Joachim Petsche. Hermann Grassmann: biography. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2009. ISBN 3-7643-8859-5 (hardcover). ???? pp. LCCN ????

Petrie:2012:LEU

[Pet12] Bruce J. Petrie. Leonhard Euler’s use and understanding of mathematical transcendence. Historia Mathematica, 39(3):280–291, August 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086012000353.

Panza:2024:MGO

[PG24] Marco Panza and Niccolò Guicciardini. Massimo Galuzzi (1943– 2023). obituary. Historia Mathematica, 68(??):49–50, September 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086024000764.

Peiffer:2018:IBM

[PGN18] Jeanne Peiffer, Hélène Gispert, and Philippe Nabonnand. Interplay between mathematical journals on various scales 1850–1950. Historia Mathematica, 45(4):323–333, November 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086018301277.

Parshall:1996:HMW

[PH96a] Karen Hunge Parshall and Jan P. Hogendijk. Historia Mathematica and the World Wide Web. Historia Mathematica, 23(4):353–354, November 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086096900373.

Parshall:1996:HMH

[PH96b] Karen Hunger Parshall and Jan P. Hogenduk. The history of mathematics, the history of science, mathematics, and Historia Mathematica. Historia Mathematica, 23(1):1–5, February 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086096900014.

Parshall:1997:CUH

[PH97] Karen Hunger Parshall and Jan P. Hogendijk. Changes underway at Historia Mathematica. Historia Mathematica, 24(4):359, November 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086097921777.

Parshall:1999:TTO

[PH99] Karen Hunger Parshall and Jan P. Hogendijk. Thoughts and thanks from the outgoing Editors. Historia Mathematica, 26(4):311–313, November 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508609992270X.

Phillips:1978:NNL

[Phi78] Esther R. Phillips. Nicolai Nicolaevich Luzin and the Moscow school of the theory of functions. Historia Mathematica, 5(3):275–305, August 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086078901143.

Phillips:1979:BRN

[Phi79a] Esther R. Phillips. Book review: Nikolai Ivanovich Fuss: 1755–1826: By V. I. Lysenko. (In Russian.) Moscow (Nauka). 1975. 119 pp. Historia Mathematica, 6(4):469–472, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079900442.

Phillips:1979:KMP

[Phi79b] Esther R. Phillips. Karl M. Peterson: the earliest derivation of the Mainardi–Codazzi equations and the Fundamental Theorem of Surface Theory. Historia Mathematica, 6(2):137–163, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079900752.

Phillips:1979:RPI

[Phi79c] Esther R. Phillips. Razvitie ponyattiya integrala. (Russian) [The development of the concept of integral]. Historia Mathematica, 6(1):85– 91, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086079901198.

Picutti:1989:PHS

[Pic89] Ettore Picutti. Pour l’histoire des sept premiers nombres parfaits. Historia Mathematica, 16(2):123–136, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086089900347.

Pier:1980:HNC

[Pie80] Jean-Paul Pier. Historique de la notion de compacité. Historia Mathematica, 7(4):425–443, November 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900063.

Piegorsch:1983:JGM

[Pie83] W. W. Piegorsch. Has J. G. Mendel been “Too accurate” in his experiments? The χ^{2} test and its significance to the evaluation of genetic segregation. Historia Mathematica, 10(1):99–100, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900435.

Pier:1986:MIL

[Pie86] Jean-Paul Pier. Mesures invariantes — de Lebesgue à nos jours. (French) [Invariant measures — from Lebesgue to our days]. Historia Mathematica, 13(3):229–240, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508608690090X.

Piegorsch:1989:DRA

[Pie89] Walter W. Piegorsch. Durand’s rules for approximate integration. Historia Mathematica, 16(4):324–333, November 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086089900797.

Pieper:1993:ECF

[Pie93] Herbert Pieper. On Euler’s contributions to the four-squares theorem. Historia Mathematica, 20(1):12–18, February 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086083710037.

Pingree:1976:BRS

[Pin76] David Pingree. Book review: Science awakening II. The birth of astronomy: By Bartel L. van der Waerden, with contributions by Peter Huber. Leyden (Noordhoff International Publishing) and New York (Oxford University Press). 1974. 362 p. 31 plates. Historia Mathematica, 3(1):90–91, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076900197.

Pingree:1980:BRA

[Pin80] David Pingree. Book review: The arithmetic of al-Uqlidisi: By A. S. Saidan. Dordrecht/Boston (Reidel). 1978. xvii + 492 pp. Historia Mathematica, 7(1):97–98, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900841.

Piotrowski:1985:WKH

[Pio85] Walerian Piotrowski. Wladysław Kretkowski and Hilbert’s Third Problem. Historia Mathematica, 12(3):258–260, August 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086085900230.

Pedersen:2011:SA

[PJ11] Olaf Pedersen and Alexander Jones. A survey of the Almagest. Sources and studies in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 0-387-84825-8 (hardcover), 0-387-84826-6 (e-). 480 pp. LCCN ????

Peckhaus:2002:HP

[PK02] Volker Peckhaus and Reinhard Kahle. Hilbert’s Paradox. Historia Mathematica, 29(2):157–175, May 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086002923451.

Plackett:1976:BRP

[Pla76] R. L. Plackett. Book review: Probability theory: a historical sketch: By L. E. Maistrov. Translated and edited by Samuel Kotz. New York and London (Academic Press). 1974. 294 pp. Historia Mathematica, 3(2):233–234, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076900562.

Pledger:1978:BRS

[Ple78] K. E. Pledger. Book review: Les systèmes d’équations polynômes dans le Siyuan Yujian (1303): by Jock Hoe. Paris, Collège de France. Mémoires de l’Institut des hautes études chinoises vol. VI. 1977. 338 pp. Distributed by presses universitaires de France. Historia Mathematica, 5(2):239–241, May 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086078900630.

Plofker:1996:ESM

[Plo96] Kim Plofker. An example of the secant method of iterative approximation in a Fifteenth-Century Sanskrit text. Historia Mathematica, 23(3):246–256, August 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086096900269.

Plofker:2001:EIT

[Plo01] Kim Plofker. The “Error” in the Indian “Taylor series approximation” to the sine. Historia Mathematica, 28(4):283–295, November 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086001923316.

Plofker:2006:BRP

[Plo06a] Kim Plofker. Book review: Ptolemy’s Geography: an Annotated Translation of the Theoretical Chapters, J. Lennart Berggren, Alexander Jones. Princeton University Press, Princeton, NJ (2001), ISBN 0-691-09259-1. Historia Mathematica, 33(3):362–364, August 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005001059.

Plotkin:2006:BRF

[Plo06b] J. M. Plotkin. Book review: Felix Hausdorff. Gesammelte Werke. Band II: Grundzüge der Mengenlehre, Egbert Brieskorn, Srishti D.

Chatterji, Moritz Epple, Ulrich Felgner, Horst Herrlich, Mirek Husek, Vladimir Kanovei, Peter Koepke, Gerhard Preuß, Walter Purkert, Erhard Scholz (Eds.). Springer-Verlag, Berlin/Heidelberg/New York (2002), ISBN 3-540-42224-2. Historia Mathematica, 33(3):372–374, August 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086006000541.

Plofker:2007:DP

[Plo07] Kim Plofker. David pingree (1933–2005). Historia Mathematica, 34(1):3–6, February 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086006000917.

Pont:1978:JHR

[Pon78] Jean-Claude Pont. Johann-Heinrich Rehbein: Une voie philosophique vers la topologie. Historia Mathematica, 5(4):443–454, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078902136.

Portnoy:1975:BRG

[Por75] Esther Portnoy. Book review: Georgii Nikolaevich Nikoladze: By A. N. Bogolyubov. Moscow (“Nauka”), 1973. 166 p. Historia Mathematica, 2(3):349, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086075900841.

Portnoy:1982:RCD

[Por82] Esther Portnoy. Riemann’s contribution to differential geometry. Historia Mathematica, 9(1):1–18, February 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086082901343.

Porter:2004:KPS

[Por04] Theodore M. Porter. Karl Pearson: the scientific life in a statistical age. Princeton University Press, Princeton, NJ, USA, 2004. ISBN 0-691-11445-5. viii + 342 pp. LCCN QA276.P64 P67 2004. URL http:

//www.loc.gov/catdir/description/prin051/2003054820.html;
 http://www.loc.gov/catdir/toc/prin051/2003054820.html.

Porter:2005:BRH

[Por05] T. M. Porter. Book review: The History of Mathematical Tables, From Sumer to Spreadsheets, Martin Campbell-Kelly, Mary Croarken, Raymond Flood, Eleanor Robson. Oxford University Press, Oxford (2003). Historia Mathematica, 32(1):98–99, February 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086004000564.

Porter:2019:TBC

[Por19] Christopher P. Porter. Three books on computability, with a special focus on Turing’s legacy. Historia Mathematica, 46(??):88–95, February 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086018301265.

Pourciau:2001:NNL

[Pou01] Bruce Pourciau. Newton and the notion of limit. Historia Mathematica, 28(1):18–30, February 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086000923012.

Pourciau:2007:FDT

[Pou07] Bruce Pourciau. Force, deflection, and time: Proposition VI of Newton’s Principia. Historia Mathematica, 34(2):140–172, May 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086006000930.

Powell:1976:AOB

[Pow76] Marvin A. Powell, Jr. The antecedents of old Babylonian place notation and the early history of Babylonian mathematics. Historia Mathematica, 3(4):417–439, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086076900719.

Phillips:1978:BRM

[PP78] Esther R. Phillips and George W. Phillips. Book review: Memoirs [Vospominaniya]: By P. Ya. Kochina. Moscow (Science-Nauka). 1974. 299 pp. One plate. Historia Mathematica, 5(2):241–245, May 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078900642.

Pont:1991:POF

[PP91] Jean-Claude Pont and Marco Panza. La philosophie ouverte de F. Gonseth à l’épreuve de la géométrie: Geneva, Switzerland, December 14–15, 1990. Historia Mathematica, 18(3):282, August 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609190402J.

Pont:2012:DDW

[PR12] Jean-Claude Pont and Christophe Rossel. Le destin douloureux de Walther Ritz (1878–1909), physicien théoricien de génie, volume 24 of Cahiers de Vallesia. Archives de l’état du Valais, Vallesia, France, 2012. ISBN 2-9700636-5-4 (hardcover). 264 + 41 pp. LCCN ????

Pisati:2022:FCR

[PR22] Arrigo Pisati and Riccardo Rosso. Felice Casorati and the reception of Gaussian optics in Italy. Historia Mathematica, 61(??):25–44, November 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086022000489.

Preveraud:2024:ICM

[Pre24] Thomas Preveraud. The International Commission on Mathematical Instruction, 1908–2008: People, events and challenges in mathematics education. Historia Mathematica, 67(??):25–27, June 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086024000119.

Price:1980:BRH

[Pri80] G. Baley Price. Book review: A history of the mathematics departments of the University of Colorado: By Burton W. Jones and Wolfgang Thron. Copyright 1979 by the Mathematics Department, University of Colorado, Boulder, Colorado. viii + 102 pages. Paperbound. Historia Mathematica, 7(1):93–95, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900828.

Proust:2022:SPV

[Pro22] Christine Proust. The sexagesimal place-value notation and abstract numbers in mathematical cuneiform texts. Historia Mathematica, 59(??):54–70, May 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086022000088.

Papamastorakis:1983:PMG

[PS83a] Emmanuil Papamastorakis and Roland Stowasser. Problemgeschichite der mathematik. I. (German) [Problem history of mathematics. I]. Historia Mathematica, 10(4):480–482, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900198.

Papamastorakis:1983:PMI

[PS83b] Emmanuil Papamastorakis and Roland Stowasser. Problemgeschichite der mathematik. II. Historia Mathematica, 10(4):480–482, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086083900204.

Papamastorakis:1983:PNM

[PS83c] Emmanuil Papamastorakis and Roland Stowasser. Problemgeschichte der neueren mathematik. (1800–1950). Historia Mathematica, 10 (4):480–482, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086083900216.

Petrova:1997:OMS

[PS97] Svetlana S. Petrova and Alexander D. Solov’ev. The origin of the method of steepest descent. Historia Mathematica, 24(4):361–375, November 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086096921461.

Pulte:1998:JCL

[Pul98] Helmut Pulte. Jacobi’s criticism of Lagrange: the changing role of mathematics in the foundations of classical mechanics. Historia Mathematica, 25(2):154–184, May 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086097921868.

Pulte:2005:AEW

[Pul05] Helmut Pulte. Axiomatik und Empirie: eine wissenschaftstheoriegeschichtliche Untersuchung zur Mathematischen Naturphilosophie von Newton bis Neumann. (German) [Axiomatics and empiricism: a historical approach to science theory to mathematical natural philosophy of Newton to Neumann]. Edition Universität. Wissenschaftliche Buchgesellschaft, Darmstadt, Germany, 2005. ISBN 3-534-15894-6 (paperback). 502 pp. LCCN QA8.4 .P85 2005.

Purkert:1986:GMB

[Pur86a] W. Purkert. Geschichte der Mathematik: 8.12. bis 14.12. 1985 Mathematisches Forschungsinstitut Oberwolfach. (German) [History of mathematics: 12–14 August 1985, Mathematical Research Institute in Oberwolfach]. Historia Mathematica, 13(3):283–292, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086086900972.

Purkert:1986:HMO

[Pur86b] Walter Purkert. History of mathematics, Oberwolfach, 1984. Historia Mathematica, 13(1):61–72, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086086902351.

Purkert:1988:TGM

[Pur88] Walter Purkert. 28. Tagung zur Geschichte der Mathematik: Mathematisches Forschungsinstitut Oberwolfach, 27.4–2.5.1987. (German) [28th meeting on the history of mathematics: Mathematical Research Institute in Oberwolfach, 27 April–2 May 1987]. Historia Mathematica, 15(2):152–164, May 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086088900079.

Purkert:1990:TGM

[Pur90] Walter Purkert. 29. Tagung zur Geschichte der mathematik: Mathematisches Forschungsinstitu Oberwolfach, 20.11.–26.11. 1988. (German) [29th meeting on the history of mathematics: Mathematical Research Institute in Oberwolfach, 20–26 November, 1988]. Historia Mathematica, 17(1):55–64, February 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609090079S.

Purkert:1991:TGM

[Pur91] Walter Purkert. 30. Tagung zur Geschichte der mathematik: Mathematisches Forschungsinstitut Oberwolfach, 6.5-12.5. 1990. (German) [30th meeting on the history of mathematics: Mathematical Research Institute in Oberwolfach, 6–12 May, 1990]. Historia Mathematica, 18(3):264–276, August 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609190397G.

Putnam:1975:WMT

[Put75] Hilary Putnam. What is mathematical truth? Historia Mathematica, 2(4):529–533, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086075901160.

Putnam:1982:PL

[Put82] Hilary Putnam. Peirce the logician. Historia Mathematica, 9(3):290– 301, August 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086082901239.

Proust:2022:ICA

[PV22] Christine Proust and Eric Vandendriessche. Introduction — a critical approach to the opposition between “concrete” and “abstract” numbers. Historia Mathematica, 59(??):3–11, May 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086022000234.

Pycior:1981:GPB

[Pyc81] Helena M. Pycior. George Peacock and the British origins of symbolical algebra. Historia Mathematica, 8(1):23–45, February 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900033.

Pycior:1982:ECS

[Pyc82] Helena M. Pycior. Early criticism of the symbolical approach to algebra. Historia Mathematica, 9(4):392–412, November 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082901057.

Pycior:1984:HMS

[Pyc84a] Helena M. Pycior. History of Mathematics summer workshop University of Toronto, July 25–August 2, 1983. Historia Mathematica, 11(2):227–229, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086084800178.

Pycior:1984:IEB

[Pyc84b] Helena M. Pycior. Internalism, externalism, and beyond: 19thcentury British algebra. Historia Mathematica, 11(4):424–441, November 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086084900272.

Pycior:1997:SIN

[Pyc97] Helena M. (Helena Mary) Pycior. Symbols, impossible numbers, and geometric entanglements: British algebra through the commentaries on Newton’s Universal arithmetick. Cambridge University Press, Cambridge, UK; New York, NY, USA, 1997. ISBN 0-521-48124-4. xi + 328 pp. LCCN QA151 .P94 1997. URL http://www.

loc.gov/catdir/description/cam027/96025469.html;   http:/
/www.loc.gov/catdir/toc/cam026/96025469.html.

Pyenson:1974:JAC

[Pye74] Lewis Pyenson. Joint Atlantic Conference on the History of Physics (JAS), Montréal, Canada, 22–23 March 1973. Historia Mathematica, 1(2):185–186, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074900111.

Pyenson:1975:BRH

[Pye75] Lewis Pyenson. Book review: The history of quantum theory: By Friedrich Hund. Translated by Gordon Reece. London (Harrap), 1974. Illustrations, index. 260 p. Can. $15.00. Historia Mathematica, 2 (3):370–375, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086075900944.

Pyenson:1976:BRP

[Pye76] Lewis Pyenson. Book review: Physics at Seventeenth and Eighteenth Century Leiden: Edward G. Ruestow. The Hague (Martinus Nijhoff.) 1973. 177 p. Gld. 24.50. Historia Mathematica, 3(1):96–100, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900239.

Pysior:1981:BRS

[Pys81] Helena Pysior. Book review: Studies in the scientific and mathematical philosophy of Charles S. Peirce: Essays by Carolyn Eisele: Edited by R. M. Martin. Studies in Philosophy, No. 29. The Hague/Paris/New York (Mouton Publishers), 1979. xii + 386 pp. 96 DM. Historia Mathematica, 8(4):477–481, November 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900574.

Pambuccian:2011:PPF

[PZ11] Victor Pambuccian and Tudor Zamfirescu. Paolo Pizzetti: the forgotten originator of triangle comparison geometry. Historia Mathematica, 38(3):415–422, August 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086010000911.

Rabinovitch:1974:ATI

[Rab74] Nachum L. Rabinovitch. An Archimedean tract of Immanuel Tovelem (14th Cent.). Historia Mathematica, 1(1):13–27, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074901633.

Rabouin:2008:BRV

[Rab08] David Rabouin. Book review: Verità matematiche e forme della natura da Galileo a Newton, Emilio Sergio. Aracne, Rome (2006), ISBN 88-548-0626-9. Historia Mathematica, 35(3):248–249, August 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086007000377.

Rabouin:2010:WDK

[Rab10] David Rabouin. What Descartes knew of mathematics in 1628. Historia Mathematica, 37(3):428–459, August 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086010000285.

Rabouin:2024:NFC

[Rab24] David Rabouin. Negatives as fictions in 16^{th} and 17^{th} century mathematics. Historia Mathematica, 69(??):41–61, December 2024.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086024000892.

Radu:2000:JGC

[Rad00] Mircea Radu. Justus Grassmann’s contributions to the foundations of mathematics: Mathematical and philosophical aspects. Historia Mathematica, 27(1):4–35, February 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086099922668.

Radloff:2002:EGP

[Rad02] Ivo Radloff. Évariste Galois: Principles and applications. Historia Mathematica, 29(2):114–137, May 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086001923377.

Radu:2003:DAA

[Rad03] Mircea Radu. A debate about the axiomatization of arithmetic: Otto Hölder against Robert Graßmann. Historia Mathematica, 30(3):341– 377, August 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086002000241.

Raina:1992:MFC

[Rai92] Dhruv Raina. Mathematical foundations of a cultural project or Ramchandra’s treatise “through the unsentimentalised light of mathematics”. Historia Mathematica, 19(4):371–384, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290003T.

Ramskov:2000:DMS

[Ram00a] Kurt Ramskov. The Danish mathematical society through 125 years. Historia Mathematica, 27(3):223–242, August 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086000922833.

Ramskov:2000:SDM

[Ram00b] Kurt Ramskov. Sources for Danish mathematics. Historia Mathematica, 27(2):164–170, May 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086099922735.

Rashed:1989:IAH

[Ras89] Roshdi Rashed. Ibn al-Haytham et les nombres parfaits. (French) [Ibn al-Haytham and perfect numbers]. Historia Mathematica, 16 (4):343–352, November 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086089900815.

Raynaud:2012:AAW

[Ray12] Dominique Raynaud. Abu al-Wafa’ Latinus? A study of method. Historia Mathematica, 39(1):34–83, February 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086011000693.

Raynaud:2023:VCA

[Ray23] Dominique Raynaud. Da Vinci’s Codex Atlanticus, fols. 395r and 686r-686v, refers to Leonardo Pisano volgarizzato, not to Giorgio Valla. Historia Mathematica, 64(??):1–18, August 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086023000174.

Robertson:1985:PDB

[RC85] Debra Robertson and J. J. Cross. Paul du Bois-Reymond. Historia Mathematica, 12(2):157, May 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086085900059.

Rechenberg:2007:BRJ

[Rec07] Helmut Rechenberg. Book review: John von Neumann: Selected Letters, vol. 27, Miklós Rédei (Ed.), in: History of Mathematics. American Mathematical Society, London Mathematical Society, Providence, RI (2005), ISBN 0-8218-3776-1. Historia Mathematica, 34(3):354–356, August 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S031508600600084X.

Redei:2005:JNS

[Réd05] Miklós Rédei, editor. John von Neumann: selected letters, volume 27 of History of Mathematics. American Mathematical Society, Providence, RI, USA, 2005. ISBN 0-8218-3776-1. ISSN 0899-2428. xxvi + 301 pp. LCCN QA29.V66 A4 2005. With a foreword by P. Lax and an introduction by Marina von Neumann Whitman, Edited and with a preface and introductory comments by Miklós Rédei.

Reizins:1976:BRE

[Rei76] L. Reizins. Book review: Essays on the development of the analytical theory of differential equations: By V. A. Dobrovolskii. Kiev (“Vishcha Shkola,” Kiev State University). 455 p. Historia Mathematica, 3(2):221–223, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076900513.

Reingold:1986:CPS

[Rei86] Nathan Reingold. A calculating people: the spead of numeracy in early America. Historia Mathematica, 13(1):81–82, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086086902405.

Reich:2015:BRR

[Rei15] Ulrich Reich. Book review: Récréations mathématiques au moyen âge, Jacques Sesiano. Presses Polytechniques et Universitaires Romandes, Lausanne (2014), ISBN: 978-2-88074-498-4. Historia Mathematica, 42(4):488–491, November 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086014001062.

Remmert:2012:BRG

[Rem12] Volker R. Remmert. Book review: Guglielmo Libri, matematico e storico della matematica. L’irresitibile ascesa dall’Ateneo pisano all’Institut de France. By Andrea Del Centina and Alessandra Fiocca. Florence (Olschki). 2010. ISBN 978-88-222-6003-1. xxii + 342 pp. + CD-ROM. 39 EUR. Historia Mathematica, 39(1):111–113, February 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086011000450.

Rey:2006:RLP

[Rey06] Javier Docampo Rey. Reading Luca Pacioli’s Summa in Catalonia: an early 16th-Century Catalan manuscript on algebra and arithmetic. Historia Mathematica, 33(1):43–62, February 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005000364.

Rey:2009:ADE

[Rey09] Javier Docampo Rey. Algebraic diagrams in an early sixteenthcentury Catalan manuscript and their possible sources. Historia Mathematica, 36(2):113–136, May 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008001006.

Reich:1989:SEF

[RFS89] Karin Reich, Menso Folkerts, and Christoph J. Scriba. Das schriftenverzeichnis von Ewald Fettweis (1881–1967) samt einer Würdigung von Olindo Falsirol. (German) [The index of writings of Ewald Fettweis (1881–1967), together with an appreciation by Olindo Falsirol]. Historia Mathematica, 16(4):360–372, November 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086089900839.

Riehm:2011:TTM

[RH11] Elaine McKinnon Riehm and Frances Hoffman. Turbulent times in mathematics: the life of J. C. Fields and the history of the Fields Medal. American Mathematical Society, Providence, RI, USA, 2011. ISBN 0-8218-6914-0. xii + 257 + 16 pp. LCCN QA29.F54 R53 2011.

Richards:1980:ASB

[Ric80] Joan L. Richards. The art and the science of British algebra: a study in the perception of mathematical truth. Historia Mathematica, 7(3):343–365, August 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086080900282.

Richards:1982:BRI

[Ric82] Joan L. Richards. Book review: Ideas of space: By Jeremy Gray. New York/London (Oxford University Press (Clarendon)) 1980. 216 pp. $49.50. Historia Mathematica, 9(1):95–97, February 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082901446.

Richards:1984:BRG

[Ric84] Joan Livingston Richards. Book review: Geschichte der geometrie: By Klaus Mainzer. Mannheim, Wein, Zurich (Bibliographisches Institut). 1980. 207 pp. Bibliography. Name and subject indexes. Historia Mathematica, 11(2):240–241, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086084800269.

Richards:1991:BRB

[Ric91] Joan L. Richards. Book review: Boole: L’oiseau de nuit en plein jour: By Souleymane Bachir Diagne. Paris: Belin, 1989. Historia Mathematica, 18(1):79, February 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609190360A.

Rice:1996:MMS

[Ric96] Adrian Rice. Mathematics in the Metropolis: a survey of Victorian London. Historia Mathematica, 23(4):376–417, November 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086096900397.

Rice:1998:BRM

[Ric98] Adrian Rice. Book review: Messengers of Mathematics: European Mathematical Journals (1800–1946). Edited by Elena Ausejo and Mariano Hormigón. Madrid (Siglo XXI de España Editores, S. A.). 1993. Historia Mathematica, 25(3):326–331, August 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086098922066.

Rider:1983:BRLa

[Rid83a] Robin E. Rider. Book review: Leonhardi Euleri commentationes mechanicae ad theoriam machinarum pertinentes volumen posterius: Series secunda. Opera mechanica et astronomica. Vol.

16. Bern (Societatis Scientiarum Naturalium Helveticae). 1979. xvii + 327 pp. Historia Mathematica, 10(4):476–477, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900162.

Rider:1983:BRLb

[Rid83b] Robin E. Rider. Book review: Leonhardi Euleri commentationes mechanicae et astronomicae ad scientiam navalem pertinentes volumen posterius: Series secunda. Opera mechanica et astronomica. Vol. 21. Bern (Societatis Scientiarum Naturalium Helveticae). 1978. ccxlii + 241 pp. Historia Mathematica, 10(4):476–477, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900174.

Rider:1988:BRH

[Rid88] Robin E. Rider. Book review: The history of mathematics from antiquity to the present: a selective bibliography: Edited by Joseph W. Dauben. New York (Garland). 1985. 508 pp. $80.00. ISBN 0-8240-9284-8. Historia Mathematica, 15(1):88–89, February 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900584.

Rider:1994:BRO

[Rid94] Robin E. Rider. Book review: Operations Analysis in the U.S. Army Eighth Air Force in World War II, 4, Charles W. McArthur, in: History of Mathematics (1990), xxiv + 349. Historia Mathematica, 21(1):95–96, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086084710123.

Ritter:1988:HFF

[Rit88] Jim Ritter. History of fractions or fractions of a history. Historia Mathematica, 15(4):372–375, November 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900316.

Rusnock:2005:BUC

[RKL05] Paul Rusnock and Angus Kerr-Lawson. Bolzano and uniform continuity. Historia Mathematica, 32(3):303–311, August 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086004000849.

Roitman:2016:SCP

[RL16] Pedro Roitman and Hervé Le Ferrand. The strange case of Paul Appell’s last memoir on Monge’s problem: “Sur les déblais et remblais”. Historia Mathematica, 43(3):288–309, August 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086016300209.

Robinson:1980:GGI

[Rob80] Donald W. Robinson. Gauss and generalized inverses. Historia Mathematica, 7(2):118–125, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900324.

Roberts:1996:AHW

[Rob96] David L. Roberts. Albert Harry Wheeler (1873–1950): a case study in the stratification of American mathematical activity. Historia Mathematica, 23(3):269–287, August 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086096900282.

Robson:2001:NSH

[Rob01] Eleanor Robson. Neither Sherlock Holmes nor Babylon: a reassessment of Plimpton 322. Historia Mathematica, 28(3):167–206, August 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086001923171.

Rodrigues:1988:ACM

[Rod88] JoséFrancisco Rodrigues. Anastácio da Cunha (1744–1787), o matemático e o poeta: Lisbon, Portugal, October 8–10, 1987. Historia Mathematica, 15(3):282, August 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900730.

Roero:1989:GSD

[Roe89] C. S. Roero. Giornate di storia della matematica: Cetraro (Cosenza) Italy, September 8–12, 1988. (Italian) [Days of the history of mathematics: Cetraro (Cosenza) Italy, September 8–12, 1988]. Historia Mathematica, 16(2):171–172, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508608990044X.

Roero:2015:MGA

[Roe15] Clara Silvia Roero. M. G. Agnesi, R. Rampinelli and the Riccati family: a cultural fellowship formed for an important scientific purpose, the Instituzioni analitiche. Historia Mathematica, 42(3):296– 314, August 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086014001086.

Rogers:1974:HMT

[Rog74] Leo Rogers. History of Mathematics at Teachers Conferences in England, Easter 1974. Historia Mathematica, 1(3):325–326, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074900755.

Rogers:1977:HMT

[Rog77] L. F. Rogers. History of Mathematics at the Third International Congress for Mathematical Education Karlsruhe, 16–21 August 1976. Historia Mathematica, 4(1):94–95, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900428.

Romano:1997:BRT

[Rom97] Antonella Romano. Book review: Tra scienza e immaginazione: Le matematiche presso il Collegio Gesuitico Napoletano (1552–1670 ca.). By Romano Gatto. Firenze (Olschki, collana “Biblioteca di Nuncius”). 1994, 392 pp. Historia Mathematica, 24(2):203–205, May 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508609692128X.

Rommevaux:1999:BRA

[Rom99] Sabine Rommevaux. Book review: Die älteste lateinische Schrift über das indische Rechnen nach al-Hwarizmi. Edited by Menso Folkerts. Historia Mathematica, 26(3):289–290, August 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086099922310.

Rommevaux:2004:BRA

[Rom04] Sabine Rommevaux. Book review: Abu l-‘Abbas an-Nayrīzīs Exzerpte aus (Ps.- ?) Simplicius’ Kommentar zu den Definitionen, Postulaten und Axiomen in Euclids Elementa I: Eingeleitet, ediert und mit arabischen und lateinischen Glossaren versehen von Rüdiger Arnzen. Köln (Essen). 2002. Historia Mathematica, 31(1):115–116, February 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086003000521.

Romano:2005:BRG

[Rom05] Antonella Romano. Book review: Giambattista Riccioli e il merito scientifico dei gesuiti nell’età barocca: a cura di Maria Teresa Borgato. Biblioteca di Nuncius XLIV. Florence (Olschki). 2002. xviii + 483 pp. Historia Mathematica, 32(2):239–240, May 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086004000412.

Rommevaux:2010:MCD

[Rom10] Sabine Rommevaux. Mathématiques et connaissance du monde réel avant Galilée. (French) [Mathematics and knowledge of the real world before Galileo]. Histoire des savoirs. Omniscience, Montreuil, France, 2010. ISBN 2-916097-26-0. 348 pp. LCCN QA611.28 .M28 2010.

Rommevaux:2013:TPT

[Rom13] Sabine Rommevaux. A treatise on proportion in the tradition of Thomas Bradwardine: the De proportionibus libri duo (1528) of Jean Fernel. Historia Mathematica, 40(2):164–182, May 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086013000037.

Rosen:1974:BRI

[Ros74] Saul Rosen. Book review: The IBM Watson Laboratory at Columbia University: a history: By Jean Ford Brennan. International Business Machines Corp., 1971. 75 p. Historia Mathematica, 1(3):341–342, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074900810.

Ross:1976:OFS

[Ros76] Richard P. Ross. Oronce Fine’s De speculo ustorio: a heretofore ignored early French Renaissance printed treatise on mathematical optics. Historia Mathematica, 3(1):63–70, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900082.

Ross:1977:DFC

[Ros77] Bertram Ross. The development of fractional calculus 1695–1900. Historia Mathematica, 4(1):75–89, February 1977. CODEN HI-MADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0315086077900398.

Rosinska:1981:TTG

[Ros81] Grazyna Rosińska. Tables trigonométriques de Giovanni Bianchini. Historia Mathematica, 8(1):46–55, February 1981. CODEN HI-MADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0315086081900045.

Ross:1989:LC

[Ros89] George Ross. Lagrange conference. Historia Mathematica, 16(2):174– 175, May 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086089900475.

Rosenfeld:1992:ASS

[Ros92] Boris A. Rosenfeld. Aḣmad Salim Sa’idan (1914–1991). Historia Mathematica, 19(4):438–439, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086092900224.

Rosso:2017:EBC

[Ros17] Riccardo Rosso. Eugenio Beltrami’s courses on the analytic and mechanical theory of heat, II. Mechanical theory of heat. Historia Mathematica, 44(4):320–341, November 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086017300708.

Rosso:2020:PEW

[Ros20] Riccardo Rosso. Probability and exams: the work of Antonio Bordoni. Historia Mathematica, 53(??):33–47, November 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086020300215.

Rowe:1983:BREa

[Row83a] David E. Rowe. Book review: Emmy Noether, 1882–1935: By Auguste Dick. Translated by H. I. Blocher. Boston/Basel/Stuttgart (Birkhäuser). 1981. 193 pp. Historia Mathematica, 10(4):462–466, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086083900125.

Rowe:1983:BREb

[Row83b] David E. Rowe. Book review: Emmy Noether: a tribute to her life and work: Edited by James W. Brewer and Martha K. Smith. New York and Basel (Marcel Dekker). 1981. 180 pp. Historia Mathematica, 10(4):462–466, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086083900137.

Rowe:1983:FCH

[Row83c] David E. Rowe. A forgotten chapter in the history of Felix Klein’s Erlanger Programm. Historia Mathematica, 10(4):448–454, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608390006X.

Rowe:1985:BRY

[Row85a] David E. Rowe. Book review: 100 Years of mathematics. A personal viewpoint: By George Temple. New York (Springer). 1981. xvi + 316 pp. Bibliography. Author and subject indexes. Historia Mathematica, 12(4):390–391, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086085900552.

Rowe:1985:BRF

[Row85b] David E. Rowe. Book review: Felix Klein, Biographien hervorragender Naturwissenschaftler, Techniker, und Mediziner, Band 50: By Renate Tobies with Fritz König. BSB B. G. Teubner (Leipzig). 1981. 104 pp. Historia Mathematica, 12(3):278–291, August 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608590028X.

Rowe:1985:BRN

[Row85c] David E. Rowe. Book review: Neohumanism and the Persistence of Pure Mathematics in Wilhelmian Germany: By Lewis Pyenson. American Philosophical Society (Philadelphia). 1983. 136 pp. Historia Mathematica, 12(3):278–291, August 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086085900308.

Rowe:1985:BRU

[Row85d] David E. Rowe. Book review: Universität, Technische Hochschule und Industrie. Ein Beitrag zur Emanzipation der Technik im 19. Jahrhundert unter besonderer Berücksichtigung der Bestrebungen Felix Kleins: By Karl-Heinz Manegold. Duncker & Humblot (Berlin). 1970. 330 pp. Historia Mathematica, 12(3):278–291, August 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086085900291.

Rowe:1985:FKE

[Row85e] David E. Rowe. Felix Klein’s “Erlanger Antrittsrede”: a transcription with English translation and commentary. Historia Mathematica, 12(2):123–141, May 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086085900035.

Rowe:1989:SHM

[Row89] David E. Rowe. Symposium on the history of modern mathematics: Vassar College, Poughkeepsie, New York, June 20–24, 1988. Historia Mathematica, 16(3):271–280, August 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086089900232.

Rowe:1990:BRF

[Row90a] David E. Rowe. Book review: Felix Klein and Sophus Lie: Evolution of the idea of symmetry in the Nineteenth Century: By I. M. Yaglom. Translated by S. Sossinsky. Boston/Basel (Birkhäuser). 1988. vii + 237 pp. $40.00. Historia Mathematica, 17(4):385–389, November 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090038F.

Rowe:1990:WHM

[Row90b] David E. Rowe. Workshop on the history of modern mathematics Göttingen, July 16–17, 1990. Historia Mathematica, 17(4):381, November 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609090035C.

Rowe:1993:BRR

[Row93] David E. Rowe. Book review: Revolutions in mathematics: Edited by Donald Gillies. Oxford (Clarendon Press). 1992. 353 pp. £55. Historia Mathematica, 20(3):320–323, August 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086083710293.

Rowe:1994:DJS

[Row94] David E. Rowe. Dirk Jan Struik and his contributions to the history of mathematics. Historia Mathematica, 21(3):245–273, August 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086084710238.

Rowe:2000:R

[Row00] David E. Rowe. Reviews. Historia Mathematica, 27(4):432–440, November 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086000922870.

Rowe:2006:BRZ

[Row06a] David E. Rowe. Book review: “Zwei wirkliche Kerle”: Neues zur Entdeckung der Gravitationsgleichungen der Allgemeinen Relativitätstheorie durch Albert Einstein und David Hilbert, Wuensch Daniela. Termessos, Göttingen (2005), ISBN 3-938016-04-3. Historia Mathematica, 33(4):500–508, November 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086005000509.

Rowe:2006:TGG

[Row06b] David E. Rowe. “Two genuine guys”: New information on the discovery of the gravitation equations of the General Relativity Theory through Albert Einstein and David Hilbert. Historia Mathematica, 33(4):500–508, November 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Rowe:2007:BRD

[Row07] David E. Rowe. Book review: David Hilbert’s Lectures on the Foundations of Geometry, 1891–1902, Michael Hallett, Ulrich Majer (Eds.). Springer-Verlag, Berlin, Heidelberg, New York (2004), ISBN 978-3-540-64373-9. Historia Mathematica, 34(4):440–443, November 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086007000080.

Rowe:2008:BRW

[Row08] David E. Rowe. Book review: Walther von Dyck (1856–1934). Mathematik, Technik und Wissenschaftsorganisation an der TH München, Ulf Hashagen. Franz Steiner Verlag, Stuttgart (2003). Historia Mathematica, 35(4):333–334, November 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008000608.

Rowe:2015:BRU

[Row15] David E. Rowe. Book review: Uniformisation des surfaces de Riemann. Retour sur un théorème centenaire, Henri Paul de Saint-Gervais. ENS Éditions, Lyon (2010), ISBN: 978-2-84788-233-9. Historia Mathematica, 42(1):98–102, February 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086014000512.

Rowe:2019:BRF

[Row19] David E. Rowe. Book review: Felix Klein. Visionen für Mathematik, Anwendungen und Unterricht, Renate Tobies. Springer, Heidelberg (2019), 574 pp. 55 EUR. Historia Mathematica, 49(??):84–93, November 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086019300606.

Rowe:2021:RSP

[Row21] David E. Rowe. On resolving singularities of plane curves via a theorem attributed to Alfred Clebsch. Historia Mathematica, 55(??): 23–38, May 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086021000185.

Rowe:2023:KFP

[Row23] David E. Rowe. “Kann eine Frau Privatdozentin werden?” — die Umfrage des Preußischen Kultusministeriums zur Habilitation von Frauen 1907, Cordula Tollmien. tredition, Hamburg (2021), 296 S., 19,99 EUR, ISBN: 978-3-347-05156-0. Historia Mathematica, 63(??): 70–74, May 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086023000034.

Royle:2017:IWT

[Roy17] Tony Royle. The impact of the women of the Technical Section of the Admiralty Air Department on the structural integrity of aircraft during World War One. Historia Mathematica, 44(4):342–366, November 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086017300502.

Rozenfeld:1975:MPM

[Roz75] Boris A. Rozenfeld. A medieval physico-mathematical manuscript newly discovered in the Kuibyshev Regional Library. Historia Mathematica, 2(1):67–69, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075900415.

Rozenfeld:1976:LPM

[Roz76] Boris A. Rozenfeld. The list of physico-mathematical works of Ibl Al-Haytham written by himself. Historia Mathematica, 3(1):75–76, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086076900100.

Rozenfeld:1990:BRT

[Roz90] B. A. Rozenfeld. Book review: La théorie des parallèles en pays d’Islam. Contribution a la préhistoire des géométries noneuclidiennes: By Khalil Jaouiche. Paris (Vrin). 1986. 266 pp. 198 F. Historia Mathematica, 17(2):170–172, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609090063J.

Rowe:1994:E

[RP94] David E. Rowe and Karen Hunger Parshall. Editorial. Historia Mathematica, 21(1):5–7, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086084710032.

Raugh:2019:LCA

[RP19] Michael Raugh and Siegmund Probst. The Leibniz catenary and approximation of e — an analysis of his unpublished calculations. Historia Mathematica, 49(??):1–19, November 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086018301290.

Robins:1985:MBA

[RS85] Gay Robins and Charles C. D. Shute. Mathematical bases of ancient Egyptian architecture and graphic art. Historia Mathematica, 12(2):107–122, May 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086085900023.

Robson:2009:OHH

[RS09] Eleanor Robson and Jacqueline A. Stedall, editors. The Oxford Handbook of the History of Mathematics. Oxford handbooks. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-921312-7 (hardcover). vii + 918 pp.

LCCN QA21 .O94 2009. URL http://www.loc.gov/catdir/
enhancements/fy0907/2008031793-b.html;   http://www.loc.
gov/catdir/enhancements/fy0907/2008031793-d.html;   http:
//www.loc.gov/catdir/toc/ecip0823/2008031793.html.

Rang:1981:ZDR

[RT81] B. Rang and W. Thomas. Zermelo’s discovery of the “Russell Paradox”. Historia Mathematica, 8(1):15–22, February 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900021.

Rossi:2002:WFS

[RT02] Corinna Rossi and Christopher A. Tout. Were the Fibonacci series and the Golden Section known in Ancient Egypt? Historia Mathematica, 29(2):101–113, May 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086001923341.

Russell:1975:BRO

[Rus75] Colin A. Russell. Book review: The origins of feedback control: By Otto Mayr. Cambridge, Mass. and London (MIT Press). 1970. 158 p. $7.95. Historia Mathematica, 2(1):100, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900609.

Russ:1980:TBP

[Rus80] S. B. Russ. A translation of Bolzano’s paper on the Intermediate Value Theorem. Historia Mathematica, 7(2):156–185, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900361.

Rice:1998:NIS

[RW98] Adrian C. Rice and Robin J. Wilson. From national to international society: the London Mathematical Society, 1867–1900. Historia Mathematica, 25(2):185–217, May 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508609892198X.

Rice:2003:RBA

[RW03] Adrian C. Rice and Robin J. Wilson. The rise of British analysis in the early 20th century: the role of G. H. Hardy and the London Mathematical Society. Historia Mathematica, 30(2):173–194, May 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086003000028.

Rice:1995:SCN

[RWG95] Adrian C. Rice, Robin J. Wilson, and J. Helen Garner. From student club to national society: the founding of the London Mathematical Society in 1865. Historia Mathematica, 22(4):402–421, November 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086085710324.

Sacks:1975:RAF

[Sac75] Gerald Sacks. Remarks against foundational activity. Historia Mathematica, 2(4):523–528, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075901159.

Sacchetti:2020:FCK

[Sac20] Andrea Sacchetti. Francesco Carlini: Kepler’s equation and the asymptotic solution to singular differential equations. Historia Mathematica, 53(??):1–32, November 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086020300446.

Sageng:1991:BRD

[Sag91] Erik Sageng. Book review: The development of Newtonian calculus in Britain, 1700–1800: By Niccolò Guicciardini. Cambridge/New York (Cambridge University Press). 1989. xii + 228 pp. $54.50. Historia Mathematica, 18(1):85–88, February 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086091903633.

Saidan:1974:BRA

[Sai74] A. S. Saidan. Book review: Al-B a hir in algebra: By as-Samaw’al. Edited with notes and introduction by Salah Ahmad and Roshdi Rashed. Damas, Syria (Ministry of Higher Education of the Syrian Arab Republic. Printed at the University of Damas). 1972. 335 p. Fr. 15. $3.00. Historia Mathematica, 1(3):346–347, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074900846.

Saito:1994:BRE

[Sai94] Ken Saito. Book review: Euclides reformatus: La teoria delle proporzioni nella scuola galileiana, Enrico Giusti. Bollati Boringhieri, Turin (1993), 348. Historia Mathematica, 21(4):465–468, November 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086084710354.

Saito:1995:DCN

[Sai95] Ken Saito. Doubling the cube: a new interpretation of its significance for early Greek geometry. Historia Mathematica, 22(2):119–137, May 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086085710130.

Saito:2005:BRA

[Sai05] Ken Saito. Book review: Apollonius of Perga’s Conica: Text, Context, Subtext, Michael N. Fried, Sabetai Unguru. Brill, Leiden (2001), ISBN 90-04-11977-9. Historia Mathematica, 32(4):481–494, November 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086005000121.

Saito:2019:RED

[Sai19] Ken Saito. Re-examination of the different origins of the arithmetical books of Euclid’s Elements. Historia Mathematica, 47(??):39–53, ???? 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086018301435.

Saliba:1978:BRM

[Sal78] George Saliba. Book review: The Muslim contribution to mathematics, Ali Abdullah Al-Daffa’. Croom Helm and Humanities Press, London and New Jersey (1977), 121. Historia Mathematica, 5(4):478–479, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086078902252.

Saliba:1987:BRD

[Sal87] George Saliba. Book review: Diophante: Les arithmétiques: Volume III: Book IV, ccvi + 162 pp., illus., bibl.; Volume IV: Books V–VII, cxxxiv + 197 pp., app., index. Edited by Roshdi Rashed. Collection des Universités de France. Paris (Société d’Edition “les belles Lettres”). 1984. Fr 500 (cloth). Historia Mathematica, 14(4):383–386, November 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508608790070X.

Sammarchi:2024:ASR

[Sam24] Eleonora Sammarchi. Additive and subtractive as relational entities in the algebra of al-Zanjani (and his predecessors). Historia Mathematica, 69(??):22–40, December 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086024000867.

Sandifer:2003:BRS

[San03] C. Edward Sandifer. Book review: A Station Favorable to the Pursuits of Science: Primary Materials in the History of Mathematics at the United States Military Academy: By Joe Albree, David C. Arney, and V. Frederick Rickey. Providence, RI (American Mathematical Society) and London (London Mathematical Society). History of Mathematics, Vol. 18. 2000. 272 pp. $59. Historia Mathematica, 30(1):87–88, February 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086002000149.

Sandifer:2005:BRS

[San05] Edward Sandifer. Book review: The Science of Conjecture: Evidence and Probability before Pascal, James Franklin. Johns Hopkins University Press, Baltimore (2002), ISBN 0-8018-7109-3. Historia Mathematica, 32(1):94–97, February 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086004000862.

Sandifer:2006:BRP

[San06] C. Edward Sandifer. Book review: Pensions and Insurance before 1800: a Social History. By Christopher G. Lewin. East Lothian, Scotland (Tuckwell Press). 2003. ISBN 1-86232-211-2. xvi + 471 pp. £25. Historia Mathematica, 33(3):365–367, August 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005000807.

Sandifer:2007:EML

[San07] Charles Edward Sandifer. The Early Mathematics of Leonhard Euler, volume 1 of Spectrum series; MAA tercentenary Euler celebration. Mathematical Association of America, Washington, DC, USA, 2007. ISBN 0-88385-559-3. xix + 391 pp. LCCN QA29.E8 S26 2007. URL

http://www.loc.gov/catdir/enhancements/fy0743/2006933948-
d.html;   http://www.loc.gov/catdir/enhancements/fy0743/
2006933948-t.html.

Sandifer:2009:BRR

[San09] Edward Sandifer. Book review: The Rise and Development of the Theory of Series up to the Early 1820s, Giovanni Ferraro. Springer, New York (2008), ISBN 978-0-387-73467-5. Historia Mathematica, 36(2):179–182, May 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008000669.

Saraiva:1988:PSH

[Sar88] Luis Saraiva. Portuguese seminar of history of mathematics. Historia Mathematica, 15(3):283–284, August 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900754.

Saraiva:1993:FHP

[Sar93] Luis Manuel Ribeiro Saraiva. On the first history of Portuguese mathematics. Historia Mathematica, 20(4):415–427, November 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086083710335.

Saraiva:1997:HMW

[Sar97] Luis M. R. Saraiva. Historiography of mathematics in the works of Rodolfo Guimarães. Historia Mathematica, 24(1):86–97, February 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086097921364.

Saraiva:2008:MML

[Sar08] Luis Manuel Ribeiro Saraiva. Mathematics in the Memoirs of the Lisbon Academy of Sciences in the 19th Century. Historia Mathematica, 35(4):302–326, November 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008000268.

Saraiva:2015:EBP

[Sar15] Luís Manuel Ribeiro Saraiva. Étienne Bézout in Portugal: the reform of the Portuguese University and beyond (1772–1838). Historia Mathematica, 42(1):14–46, February 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086014000536.

Sauer:2006:BRD

[Sau06a] Tilman Sauer. Book review: David Hilbert and the Axiomatization of Physics (1898–1918). From Grundlagen der Geometrie to Grundlagen der Physik, vol. 10, Leo Corry, in: Archimedes. New Studies in the History and Philosophy of Science and Technology. Kluwer Academic, Dordrecht/Boston/London (2004), ISBN 1-4020-2777-X. Historia Mathematica, 33(4):497–500, November 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086006000188.

Sauer:2006:FET

[Sau06b] Tilman Sauer. Field equations in teleparallel space–time: Einstein’s Fernparallelismus approach toward unified field theory. Historia Mathematica, 33(4):399–439, November 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086005001060.

Sauer:2009:BRC

[Sau09] Tilman Sauer. Book review: La correspondance entre Henri Poincaré et les physiciens, chimistes, et ingénieurs. Birkhäuser, Basel, Boston, Berlin (2007), ISBN 978-3-7643-7136-4. Historia Mathematica, 36 (2):185–188, May 2009. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086008000621.

Sauer:2010:LBH

[Sau10] Tilman Sauer. Ludwig Boltzmann (1844–1906): Zum hundertsten Todestag [hundredth anniversary of his death]. Historia Mathematica, 37(2):310–311, May 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086009000093.

Sauer:2014:BRA

[Sau14] Tilman Sauer. Book review: Arnold Sommerfeld. Atomphysiker und Kulturbote 1868-1951. Eine Biografie, vol. 29, Michael Eckert, in: Deutsches Museum. Abhandlungen und Berichte Neue Folge. Wallstein-Verlag, Göttingen (2013), ISBN: 978-3-8353-1206-7. Historia Mathematica, 41(3):338–341, August 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086013000712.

Scriba:2003:CJW

[SB12] Christoph J. Scriba and Philip Beeley, editors. The correspondence of John Wallis. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2003–2012. ISBN 0-19-851066-7 (v. 1), 0-19-856601-8 (v. 2), 0-19-856947-5 (v. 3). ???? pp. LCCN QA29.W3 A4 2003. URL

http://www.loc.gov/catdir/enhancements/fy0620/2004270866-
d.html;   http://www.loc.gov/catdir/enhancements/fy0620/
2004270866-t.html.

Scanlan:1994:BRP

[Sca94] Michael Scanlan. Book review: Perspectives on the History of Mathematical Logic: Edited by Thomas Drucker. Boston, Basel, Berlin (Birkhauser). 1991. xxiii + 195 pp. With an index of names. Historia Mathematica, 21(2):209–214, May 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086084710202.

Schaaf:1976:BRL

[Sch76a] William L. Schaaf. Book review: L’arithmétique amusante: [Arithmetical recreations]. By Edouard Lucas. Facsimile reprint. Paris (Albert Blanchard). 1974. 266 p. fr. 22. Historia Mathematica, 3(1):91– 93, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086076900203.

Schneider:1976:IPM

[Sch76b] Ivo Schneider. The introduction of probability into mathematics. Historia Mathematica, 3(2):135–140, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508607690029X.

Schlissel:1977:IDW

[Sch77] Arthur Schlissel. The initial development of the WKB solutions of linear second order ordinary differential equations and their use in the connection problem. Historia Mathematica, 4(2):183–204, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077901112.

Schein:1979:HNE

[Sch79] Boris M. Schein. History of non-Euclidean geometry. The development of the concept of geometric space. Historia Mathematica, 6 (4):460–464, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086079900405.

Scholz:1982:HIB

[Sch82a] Erhard Scholz. Herbart’s influence on Bernhard Riemann. Historia Mathematica, 9(4):413–440, November 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086082901069.

Schubring:1982:ABT

[Sch82b] Gert Schubring. Ansätze zur begründung theoretischer terme in der mathematik-Die theorie des unendlichen bei Johann Schultz. Historia Mathematica, 9(4):441–484, November 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086082901070.

Schneider:1983:PIJ

[Sch83] Ivo Schneider. Potenzsummenformeln im 17. jahrhundert. Historia Mathematica, 10(3):286–296, August 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086083900794.

Scharlau:1986:MCR

[Sch86a] Winfried Scharlau. The mathematical correspondence of Rudolf Lipschitz. Historia Mathematica, 13(2):165–167, May 1986.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086086900297.

Schubring:1986:TPD

[Sch86b] Gert Schubring. The three parts of the Dirichlet Nachlass [legacy]. Historia Mathematica, 13(1):52–56, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086086902284.

Scharlau:1988:ESS

[Sch88] Winfried Scharlau. Die Entdeckung der Sylow–Sätze. (German) [Discovery of the Sylow theorems]. Historia Mathematica, 15(1):40–52, February 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086088900481.

Scholz:1989:BRS

[Sch89] Erhard Scholz. Book review: Studies in the history of mathematics: Edited by Esther R. Phillips. Studies in Mathematics, Vol. 26. Mathematical Association of America. 1987. 308 pp. Historia Mathematica, 16(4):390–392, November 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086089900906.

Schlote:1990:SGM

[Sch90a] Karl-Heinz Schlote. “Seminar zur Geschichte der Mathematik” in Leipzig. Historia Mathematica, 17(3):269–270, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086090900134.

Schubring:1990:BRK

[Sch90b] Gert Schubring. Book review: Die Krise der Anschauung: Studien zur Wissenschafts-, Sozial- und Bildungsgeschichte der Mathematik,

3. By Klaus Th. Volkert. Göttingen (Vandenhoeck & Ruprecht). 1986. xxxii + 420 pp. 98 DM. Historia Mathematica, 17(2):178–180, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090066M.

Schneider:1991:BRC

[Sch91a] Ivo Schneider. Book review: Classical probability in the enlightenment: By Lorraine Daston. Princeton, NJ (Princeton Univ. Press). 1988. xviii + 423 pp. including bibliography and index. Historia Mathematica, 18(1):67, February 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086091903563.

Schneider:1991:BRH

[Sch91b] Ivo Schneider. Book review: The history of statistics — The measurement of uncertainty before 1900: By Stephen M. Stigler. Cambridge, MA (Harvard University Press). 1986. xviii + 410 pp. including two appendices, suggested readings, bibliography, and index. Historia Mathematica, 18(1):67–74, February 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086091903585.

Schneider:1991:BRR

[Sch91c] Ivo Schneider. Book review: The rise of statistical thinking 1820– 1900: By Theodore M. Porter. Princeton, NJ (Princeton Univ. Press). 1986. xii + 322 pp. including index. Historia Mathematica, 18(1):67, February 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086091903574.

Schot:1992:EYS

[Sch92] Steven H. Schot. Eighty years of Sommerfeld’s radiation condition. Historia Mathematica, 19(4):385–401, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290004U.

Schreiber:1993:NCP

[Sch93a] Peter Schreiber. A note on the cattle problem of Archimedes. Historia Mathematica, 20(3):304–306, August 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086083710256.

Schubring:1993:BBU

[Sch93b] Gert Schubring. Bernard Bolzano — not as unknown to his contemporaries as is commonly believed? Historia Mathematica, 20(1):45–53, February 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086083710050.

Schreiber:1995:SJS

[Sch95] Peter Schreiber. A supplement to J. Shallit’s paper “Book Review: Origins of the Analysis of the Euclidean Algorithm”. Historia Mathematica, 22(4):422–424, November 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086085710336.

Schreiber:1996:BRL

[Sch96] Peter Schreiber. Book review: Leonardo Pisani Liber abbaci oder Lesevergnügen eines Mathematikers. By Heinz Lüneburg. Mannheim/ Leipzig/Wien/Zürich (BI Wissenschaftsverlag). Historia Mathematica, 23(1):88–89, February 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086096900087.

Schubring:1998:UPW

[Sch98] Gert Schubring. An unknown part of Weierstraß’s Nachlaß [Estate]. Historia Mathematica, 25(4):423–430, November 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086098922170.

Schappacher:1999:BRR

[Sch99a] Norbert Schappacher. Book review: Riposte Armonie: Lettere di Federigo Enriques a Guido Castelnuovo. Edited by Umberto Bottazini, Alberto Conte, and Paola Gario. Historia Mathematica, 26 (4):385–388, November 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086099922589.

Schreiber:1999:NHD

[Sch99b] Peter Schreiber. A new hypothesis on Dürer’s enigmatic polyhedron in his copper engraving “Melencolia I”. Historia Mathematica, 26 (4):369–377, November 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086099922450.

Schreiber:2000:R

[Sch00a] Peter Schreiber. Reviews. Historia Mathematica, 27(1):80–81, February 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086000922626.

Schuller:2000:BRL

[Sch00b] Volkmar Schüller. Book review: Leonhardi Euleri opera omnia, secunda, vol. 31: Commentationes mechanicae et astronomicae ad physicam cosmicam pertinentes. Edited by Eric J. Aiton. Historia Mathematica, 27(3):305–309, August 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086000922869.

Schirrmacher:2002:BRE

[Sch02] Arne Schirrmacher. Book review: Die Entstehung der Knotentheorie. Kontexte und Konstruktionen einer modernen mathematischen Theorie. By Moritz Epple. Historia Mathematica, 29(1):70–72, February 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086001923092.

Schubring:2005:BRS

[Sch05] Gert Schubring. Book review: Symbols, Impossible Numbers, and Geometric Entanglements: British Algebra Through the Commentaries on Newton’s Universal Arithmetick, Helena Mary Pycior. Cambridge Univ. Press, Cambridge, UK (1997), ISBN 0-521-48124-4. Historia Mathematica, 32(3):363–366, August 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086004000825.

Scholz:2006:IGQ

[Sch06a] Erhard Scholz. Introducing groups into quantum theory (1926– 1930). Historia Mathematica, 33(4):440–490, November 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005001072.

Schweber:2006:DSD

[Sch06b] Libby Schweber. Disciplining statistics: demography and vital statistics in France and England, 1830/1885. Politics, history, and culture. Duke University Press, Durham, NC, USA, 2006. ISBN 0-8223-3825-4 (hardcover), 0-8223-3814-9 (paperback). 277 pp.

LCCN HB853.F8 S38 2006. URL http://www.loc.gov/catdir/
toc/ecip0611/2006010426.html.

Scholz:2007:BRH

[Sch07] Erhard Scholz. Book review: Hausdorff on Ordered Sets, vol. 25, J. M. Plotkin (Ed.), in: History of Mathematics. American Mathematical Society, London Mathematical Society, Providence, RI (2005), ISBN 0-8218-3788-5. Historia Mathematica, 34(3):352–354, August 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086006000875.

Schubring:2009:BRS

[Sch09] Gert Schubring. Book review: Stuttgarter Mathematiker: Geschichte der Mathematik an der Universität Stuttgart von 1829 bis 1945 in Biographien, Karl-Heinz Böttcher, Bertram Maurer. Universität Stuttgart, Stuttgart (2008), ISBN 978-3-926269-34-8. Historia Mathematica, 36(3):285–287, August 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008001031.

Schneider:2011:BRK

[Sch11] Ivo Schneider. Book review: Kosmos und Zahl — Beiträge zur Mathematik- und Astronomiegeschichte, zu Alexander von Humboldt und Leibniz. Edited by Hartmut Hecht, Regina Mikosch, Ingo Schwarz, Harald Siebert, and Romy Werther. Boethius, vol.

58. Stuttgart (Steiner). 2008. ISBN 978-3-515-09176-3. 510 pp. 72 EUR. Historia Mathematica, 38(4):561–563, November 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086011000462.

Schubring:2012:BRM

[Sch12] Gert Schubring. Book review: Mathematicians at War. Volterra and his French colleagues in World War I. By Laurent Mazliak and Rossana Tazzioli. Dordrecht, Heidelberg, London, New York (Springer). 2009. ISBN 978-90-481-2739-9. ix + 194 p. 106,95 EUR. Historia Mathematica, 39(3):347–350, August 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086011000930.

Schlimm:2013:CBM

[Sch13a] Dirk Schlimm. The correspondence between Moritz Pasch and Felix Klein. Historia Mathematica, 40(2):183–202, May 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086013000050.

Schneider:2013:BRA

[Sch13b] Ivo Schneider. Book review: The Archimedes Palimpsest. Edited in two volumes by Reviel Netz, William Noel, Natalia Tchernetska, and Nigel Wilson. Volume 1: Catalogue and Commentary; volume 2: Images and Transcriptions. Cambridge (Cambridge University Press). 2011. ISBN 978-1-107-01684-2, Price: £150 ($215). Historia Mathematica, 40(1):84–89, February 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086012000705.

Schubring:2013:BRS

[Sch13c] Gert Schubring. Book review: Scienza, Geometria, Geometrie. Un percorso storico-didattico [Science, Geometry, Geometries. A historical-educational itinerary], Luigi Maierú. Rubbettino Editore, Soveria Mannelli (2008), ISBN: 978-88-498-2156-7. Historia Mathematica, 40(2):205–206, May 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086012000298.

Scheel:2014:BRD

[Sch14] Katrin Scheel. Der Briefwechsel, Richard Dedekind–Heinrich Weber. (German) [Correspondence, Richard Dedekind–Heinrich Weber], volume 5 of Abhandlungen, Akademie der Wissenschaften, Hamburg. Walter de Gruyter, New York, NY, USA and Berlin, Germany, 2014. ISBN 3-11-037366-1. ???? pp. LCCN QA155.2 .S34 2014.

Schneider:2016:BRM

[Sch16] Martina R. Schneider. Book review: Mathematik bei den Juden. Band II: 1551-1840. Georg Olms, Hildesheim, Zürich, New York (2014), ISBN: 978-3-487-13585-4. Historia Mathematica, 43(2):209–213, ???? 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086015000609.

Schneider:2017:BRE

[Sch17] Martina R. Schneider. Book review: Emmy Noether, die Noether-Schule und die moderne Algebra. Zur Geschichte einer kulturellen Bewegung, Mechthild Koreuber, in: Mathematik im Kontext. Springer Spektrum, Berlin, Heidelberg (2015), ISBN: 978-3-662-44149-7. Historia Mathematica, 44(1):79–82, February 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086016300507.

Schubring:2019:BRK

[Sch19] Gert Schubring. Book review: Karl Weierstraß (1815–1897). Aspekte seines Lebens und Wirkens — Aspects of his Life and Work, Wolfgang König, Jürgen Sprekels (Eds.). Springer Fachmedien, Wiesbaden (2016), 259 pp. Historia Mathematica, 48(??):96–100, August 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086019300412.

Sialaros:2019:EN

[SCM19] Michalis Sialaros, Jean Christianidis, and Athanasia Megremi. Editorial note. Historia Mathematica, 47(??):1–3, ???? 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086018301381.

Scoth:2016:HED

[Sco16] Roberto Scoth. Higher education, dissemination and spread of the mathematical sciences in Sardinia (1720–1848). Historia Mathematica, 43(2):172–193, ???? 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086015000993.

Scott:2022:UD

[Sco22] Patrick Scott. Ubiratan D’Ambrosio (1932–2021). Historia Mathematica, 58(??):3–6, February 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086022000027.

Scriba:1975:CJH

[Scr75a] C. J. Scriba. Chronology of J. E. Hofmann, biobibliographic note, and supplementary bibliography of his publications. Historia Mathematica, 2(2):147–152, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086075901391.

Scriba:1975:PFH

[Scr75b] Christoph J. Scriba. The place and function of a ‘historical introduction’ in the curriculum for mathematics students. Historia Mathematica, 2(3):327–331, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086075900762.

Scriba:1979:HHM

[Scr79a] Christoph J. Scriba. Heinrich Hermelink in memoriam. Historia Mathematica, 6(3):233–235, August 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508607990123X.

Scriba:1979:MRH

[Scr79b] Christoph J. Scriba. A message to readers of Historia Mathematica from the Chairman of the International Commission on the History of Mathematics. Historia Mathematica, 6(1):1–2, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079900983.

Scriba:1981:WKN

[Scr81] Christoph J. Scriba. Wie kommt “Napoleons Satz” zu seinem namen?. (German) [how does “Napoleon’s set” get its name?]. Historia Mathematica, 8(4):458–459, November 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086081900549.

Scriba:1982:ICHa

[Scr82a] Christoph J. Scriba. International commission on the history of mathematics: Business Meeting Bucharest, Romania, September 1, 1981. Historia Mathematica, 9(2):207–208, May 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086082900076.

Scriba:1982:ICHb

[Scr82b] Christoph J. Scriba. International commission on the history of mathematics: Report to the division of history of the international union of history and philosophy of science August 1981. Historia Mathematica, 9(2):209–210, May 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086082900088.

Scriba:1983:GCD

[Scr83] Christoph J. Scriba. Gregory’s converging double sequence: a new look at the controversy between Huygens and Gregory over the “analytical” quadrature of the circle. Historia Mathematica, 10(3):274– 285, August 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086083900782.

Scriba:1984:STM

[Scr84] C. J. Scriba. Symposium on transmission of mathematical science: International commission on the history of Mathematics (International Union on the history and philosophy of Science, division of the history of Science) XVIIth International Congress on the history of Science Berkeley, 1–8 August 1985. Historia Mathematica, 11(2):222– 225, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086084800142.

Scriba:1985:ICH

[Scr85a] C. J. Scriba. International Commission on the History of Mathematics: meeting at the Berkeley Congress. Historia Mathematica, 12(2):158–159, May 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086085900072.

Scriba:1985:FPC

[Scr85b] Christoph J. Scriba. A farewell from the parting Chairman of the International Commission on the History of Mathematics. Historia Mathematica, 12(4):321–322, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086085900412.

Scriba:1985:TYH

[Scr85c] Christoph J. Scriba. Thirty years of the “History of mathematics” at Oberwolfach: In remembrance of J. E. Hofmann (1900–1973). Historia Mathematica, 12(4):369–371, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086085900461.

Scriba:1987:XIC

[Scr87] C. J. Scriba. XVIIIth International Congress of the History of Science: Hamburg and Munich, Federal Republic of Germany 1– 9 August 1989. Historia Mathematica, 14(2):180–181, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086087900243.

Scriba:1988:ICH

[Scr88a] C. J. Scriba. 18th international Congress on the history of science. Historia Mathematica, 15(3):275–278, August 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900699.

Scriba:1988:WI

[Scr88b] Christoph J. Scriba. Who is it? Historia Mathematica, 15(1):73–74, February 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508608890050X.

Scriba:1992:COS

[Scr92] Chirstoph J. Scriba. Clas-Olof Selenius (1922–1991). Historia Mathematica, 19(3):325–327, August 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609290047F.

Scriba:1996:BTR

[Scr96] Christoph J. Scriba. A birthday tribute to R. C. Gupta. Historia Mathematica, 23(2):117–120, May 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086096900142.

Scriba:2003:BRK

[Scr03] Christoph J. Scriba. Book review: Die kubischen Gleichungen bei Nicolo Tartaglia. Die relevanten Textstellen aus seinen Quesiti et inventioni diversi auf Deutsch übersetzt und kommentiert: By Friedrich Katscher. Veröffentlichungen der Kommission für Geschichte der Naturwissenschaften, Mathematik und Medizin No. 53. Vienno (Österreichische Akademie der Wissenschaften, Mathematisch-Naturwissenschaftliche Klasse). 2001. iv+ 76 S. ISBN 3-7001-2984-X. Historia Mathematica, 30(2):220–221, May 2003.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086002000307.

Scriba:2004:AP

[Scr04] Christoph J. Scriba. Adolf Prag (1906–2004). Historia Mathematica, 31(4):409–413, November 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086004000606.

Scriba:2007:BRF

[Scr07] Christoph J. Scriba. Book review: Felix Hausdorff. Gesammelte Werke. Band VII: Philosophisches Werk, Werner Stegmaier (Ed.). Springer-Verlag, Berlin/Heidelberg/New York (2004), ISBN 3-540-20836-4. Historia Mathematica, 34(1):112–113, February 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086006000516.

Scriba:2008:BRL

[Scr08] Christoph J. Scriba. Book review: Leonhard Euler. Birkhäuser Verlag, Basel/Boston/Berlin (2007), ISBN 3-7643-7538-8. Historia Mathematica, 35(2):142–143, May 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086007000948.

Schlote:1994:AGF

[SD94] Karl-Heinz Schlote and Uwe Dathe. Die Anfänge von Gottlob Freges wissenschaftlicher Laufbahn. (German) [The beginnings of Gottlob Frege’s scientific career]. Historia Mathematica, 21(2):185–195, May 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086084710172.

Serfati:2008:MFD

[SD08] Michel Serfati and Dominique Descotes. Mathématiciens français du XVIIe siècle: Descartes, Fermat, Pascal. (French) [French mathematicians of the 17th Century: Descartes, Fermat, Pascal]. Centre d’études sur les réformes, l’humanisme et l’âge classique. Presses universitaires Blaise-Pascal, Clermont-Ferrand, France, 2008. ISBN 2-84516-354-1. 280 pp. LCCN QA28 .M3795 2008.

Segal:1976:BRD

[Seg76] Sanford L. Segal. Book review: Development of mathematics in the German Democratic Republic: Edited by Horst Sachs. Berlin (VEB Deutscher Verlag der Wissenschaften). 776 p. Historia Mathematica, 3(2):225–226, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076900537.

Segal:1980:HH

[Seg80] S. L. Segal. Helmut Hasse in 1934. Historia Mathematica, 7(1):46–56, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086080900634.

Segal:1986:MGP

[Seg86] Sanford L. Segal. Mathematics and German politics: the national socialist experience. Historia Mathematica, 13(2):118–135, May 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608690025X.

Segal:1990:BRM

[Seg90] S. L. Segal. Book review: Die Mathematik und ihre Dozenten an der Berliner Universität, 1810–1933: 2nd ed. By Kurt Biermann. Berlin (Akademie-Verlag). 1988. Historia Mathematica, 17(3):287– 289, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086090900189.

Selenius:1974:BRR

[Sel74] Clas-Olof Selenius. Book review: Razvitie teorii chisel v rossii: By E. P. Ozhigova. Leningrad (Izd Nauka), 1972. 360 pp. 1r, 80k. Historia Mathematica, 1(1):97, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074901888.

Selenius:1975:RCP

[Sel75] Clas-Olof Selenius. Rationale of the Chakravala process of Jayadeva and Bhaskara II. Historia Mathematica, 2(2):167–184, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075901433.

Selenius:1990:BRH

[Sel90] Clas-Olof Selenius. Book review: The history of mathematics in Finland 1828–1918: By Gustav Elfving. The History of Learning and Science in Finland 1828–1918, Vol. 4. Helsingfors (Societas Scientiarum Fennica). 1981. 195 pp. Historia Mathematica, 17(2):180–182, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090067N.

Seligman:1991:BRC

[Sel91] George B. Seligman. Book review: A century of mathematics in America: 3 vols. Peter A. Duren, Ed. Providence, Rhode Island (American Mathematical Society). 1988. Historia Mathematica, 18 (1):82–85, February 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086091903622.

Selles:2006:IFN

[Sel06] Manuel A. Sellés. Infinitesimals in the foundations of Newton’s mechanics. Historia Mathematica, 33(2):210–223, May 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005000443.

Sommerfeld:2000:WBG

[SEM04] Arnold Sommerfeld, Michael Eckert, and Karl Märker. Wissenschaftlicher Briefwechsel. (German) [Scientific Correspondence]. Deutsches Museum, Berlin, Germany, 2000–2004. ISBN 3-928186-44-2 (set), 3-928186-49-3 (Vol. 1), 3-928186-53-1 (Vol. 2). 694 (vol. 1), 735 (vol. 2) pp. LCCN QC16.S76 A4 2000.

Senechal:1984:BRD

[Sen84] Marjorie L. Senechal. Book review: Descartes on polyhedra: By P. J. Federico. New York (Academic Press). 1983. Historia Mathematica, 11(3):333–334, August 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086084900442.

Seneta:1987:CFE

[Sen87] E. Seneta. Chuprov on finite exchangeability, expectation of ratios, and measures of association. Historia Mathematica, 14(3):243–257, August 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086087900449.

Seneta:1991:BRH

[Sen91] Eugene Seneta. Book review: A history of probability and statistics and their applications before 1750: By Anders Hald. Wiley Series in Probability and Mathematical Statistics. New York (Wiley). 1990. xii + 586 pp. Historia Mathematica, 18(1):75–78, February 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086091903596.

Seneta:1992:HSL

[Sen92] E. Seneta. On the history of the Strong Law of Large Numbers and Boole’s inequality. Historia Mathematica, 19(1):24–39, February 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290053E.

Seneta:1994:CLH

[Sen94] E. Seneta. Carl Liebermeister’s hypergeometric tails. Historia Mathematica, 21(4):453–462, November 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086084710330.

Seneta:1999:BRC

[Sen99] Eugene Seneta. Book review: Condorcet. Arithmétique politique. Textes rares ou inédits (1767–1789). Critical edition commented by Bernard Bru and Pierre Crépel. Historia Mathematica, 26(4): 383–385, November 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086099922577.

Seneta:2004:MRM

[Sen04] Eugene Seneta. Mathematics, religion, and Marxism in the Soviet Union in the 1930s. Historia Mathematica, 31(3):337–367, August 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086003000466.

Seneta:2006:BRS

[Sen06] Eugene Seneta. Book review: Statistics on the Table: the History of Statistical Concepts and Methods, Paperback edition, M. Stigler Stephen (2002). Harvard Univ. Press, Cambridge, MA (1999), ISBN 0-674-00979-7. Historia Mathematica, 33(2):255–258, May 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005000522.

Seneta:2010:BRM

[Sen10] Eugene Seneta. Book review: Mathematicians Fleeing from Nazi Germany. Individual Fates and Global Impact. By Reinhard Siegmund-Schultze. Princeton and Oxford (Princeton University Press). 2009. ISBN 978-0-691-14041-4. xxv + 471 pp. US$49.50. Historia Mathematica, 37(4):716–722, November 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086009001335.

Sepkoski:2005:NCS

[Sep05] David Sepkoski. Nominalism and constructivism in seventeenthcentury mathematical philosophy. Historia Mathematica, 32(1):33– 59, February 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086003000818.

Serge:1990:BRC

[Ser90] Michael Serge. Book review: Cultura, scienze e tecniche nella Venezia del cinquecento. Atti del convegno internazionale di studio: Giovan Battista Benedetti e il suo tempo: Edited by Antonio Manno. Venice (Istituto Veneto di Scienze, Lettere ed Arti). 1987. 503 pp. Historia Mathematica, 17(1):78–80, February 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609090088U.

Sergio:2006:VMF

[Ser06] Emilio Sergio. Verità matematiche e forme della natura da Galileo a Newton. (Italian) [Mathematical truths and forms of nature from Galileo to Newton]. Aracne, Roma, Italia, 2006. ISBN 88-548-0626-9. 424 pp. LCCN QA8.4 .S4162 2006. 23.00 EUR. URL http://www.

loc.gov/catdir/toc/casalini05/06929067.pdf.

Sereni:2018:SCJ

[Ser18] Andrea Sereni. Book review: The Selected Correspondence of L. E. J. Brouwer. Sources and Studies in the History of Mathematics and Physical Science, Dirk Van Dalen (Ed.). Springer-Verlag, London (2011), p. VIII, 532. Hardcover ISBN: 978-0-85729-537-8, Softcover ISBN: 978-1-4471-2691-1. Historia Mathematica, 45(3):300– 302, August 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086018300454.

Sesiano:2013:RMA

[Ses13] Jacques Sesiano. Récréations mathématiques au Moyen Age. (French) [Mathematical recreations in the Middle Ages]. Presses polytechniques et universitaires romandes, Lausanne, Suisse, 2013. ISBN 2-88074-498-9 (paperback). 289 pp. LCCN ????

Sesiano:2014:LM

[Ses14] Jacques Sesiano. Liber Mahameleth. Sources and Studies in the History of Mathematics and Physical Sciences. Springer International Publishing AG, Cham, Switzerland, 2014. ISBN 3-319-03939-3 (hardcover). ???? pp. LCCN QA32 .S47 2014.

Shell-Gellasch:2004:BRH

[SG04] Amy Shell-Gellasch. Book review: A History of Mathematics: By Jeff Suzuki. Prentice-Hall (New Jersey). 2002. 815 pp. Historia Mathematica, 31(2):227–228, May 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086003000624.

Shell-Gellasch:2007:HHR

[SG07] Amy Shell-Gellasch, editor. Hands on history: a resource for teaching mathematics, volume 72 of MAA notes. Mathematical Association of America, Washington, DC, USA, 2007. ISBN 0-88385-182-2 (paperback). xii + 177 pp. LCCN QA11.2 .H367 2007. URL

http://www.loc.gov/catdir/enhancements/fy0827/2007937009-
b.html;   http://www.loc.gov/catdir/enhancements/fy0827/
2007937009-d.html;   http://www.loc.gov/catdir/enhancements/
fy0827/2007937009-t.html.

Shell-Gellasch:2009:BRT

[SG09] Amy Shell-Gellasch. Book review: Tools of American Mathematics Teaching, 1800–2000, Peggy Aldrich Kidwell, Amy Ackerberg-Hastings, David Lindsay Roberts. Johns Hopkins University Press, Baltimore (2008), ISBN 978-0-8018-8814-4. Historia Mathematica, 36(3):283–285, August 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008001092.

Shell-Gellasch:2005:CCU

[SGJ05] Amy Shell-Gellasch and Dick Jardine, editors. From calculus to computers: using the last 200 years of mathematics history in the classroom, volume #68 of MAA notes. Mathematical Association of America, Washington, DC, USA, 2005. ISBN 0-88385-178-4. xii + 255

pp. LCCN QA21 .F758 2005. URL http://www.loc.gov/catdir/
enhancements/fy0659/2005908422-d.html;   http://www.loc.
gov/catdir/enhancements/fy0659/2005908422-t.html.

Shallit:1994:OAE

[Sha94] Jeffrey Shallit. Origins of the analysis of the Euclidean algorithm. Historia Mathematica, 21(4):401–419, November 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086084710317.

Sheynin:1993:CSC

[She93] Oscar B. Sheynin. Chuprov, Slutsky, and Chetverikov: Some comments. Historia Mathematica, 20(3):247–254, August 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086083710219.

Sheynin:1994:ITP

[She94] Oscar Sheynin. Ivory’s treatment of pendulum observations. Historia Mathematica, 21(2):174–184, May 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086084710160.

Sheynin:1996:MMT

[She96] Oscar Sheynin. Mendeleev and the mathematical treatment of observations in natural science. Historia Mathematica, 23(1):54–67, February 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508609690004X.

Shirley:1986:HMA

[Shi86] Lawrence H. Shirley. History of mathematics in Africa: a section of the Second Pan-African Congress of Mathematicians of the African Mathematical Union. Historia Mathematica, 13(3):293–295, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086086900996.

Shirley:1988:SPA

[Shi88] Lawrence Shirley. Second Pan-African Congress of Mathematicians: University of Jos, Jos, Nigeria, March 23–29, 1986. Historia Mathematica, 15(3):281, August 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086088900729.

Shirong:1993:SIS

[Shi93] Guo Shirong. The second international symposium on the history of mathematics and mathematical education in the regions using Chinese characters. Historia Mathematica, 20(1):97–98, February 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008800584.

Shirong:1994:SHS

[Shi94] Guo Shirong. The symposium on history of science on the occasion of the founding of the Institute for the History of Science at Inner Mongolia Normal University. Historia Mathematica, 21(2):207–208, May 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008800122.

Shoesmith:1983:EEP

[Sho83] Eddie Shoesmith. Expectation and the early probabilists. Historia Mathematica, 10(1):78–80, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086083900356.

Shoesmith:1985:TSA

[Sho85] Eddie Shoesmith. Thomas Simpson and the arithmetic mean. Historia Mathematica, 12(4):352–355, November 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086085900448.

Shoesmith:1986:HSG

[Sho86] Eddie Shoesmith. Huygens’ solution to the Gambler’s Ruin Problem. Historia Mathematica, 13(2):157–164, May 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086086900285.

Shoesmith:1987:CCA

[Sho87] Eddie Shoesmith. The continental controversy over Arbuthnot’s argument for divine providence. Historia Mathematica, 14(2):133– 146, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086087900176.

Shore:2002:BSH

[Sho02] Steven N. Shore. Blue sky and hot piles: the evolution of radiative transfer theory from atmospheres to nuclear reactors. Historia Mathematica, 29(4):463–489, November 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086002923608.

Shore:2003:MMC

[Sho03] Steven N. Shore. Macrocosmos/microcosmos: celestial mechanics and old quantum theory. Historia Mathematica, 30(4):494–513, November 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086002000289.

Shulman:1980:BRU

[Shu80] David Shulman. Book review: United States diplomatic codes and ciphers 1775–1938: By Ralph E. Weber. Chicago. xviii + 633 pp. Illus. $49.95. Historia Mathematica, 7(4):452–454, November 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900142.

Shuchun:1992:GSE

[Shu92] Guo Shuchun. Guo Shuchun’s edition of the Jiu Zhang Suan Shu: (nine chapters on the mathematical art). Historia Mathematica, 19(2):200–202, May 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609290077O.

Sidoli:2019:IAD

[SI19] Nathan Sidoli and Yoichi Isahaya. Nas.ir al-Din al-T .usi’s comments on Euclid’s data. Historia Mathematica, 47(??):87–105, ???? 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086019300205.

Sialaros:2023:BRN

[Sia23] Michalis Sialaros. Book review: A New History of Greek Mathematics, Reviel Netz. Cambridge University Press, Cambridge (2022). Historia Mathematica, 65(??):30–32, November 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086023000654.

Sidoli:2004:UTD

[Sid04] Nathan Sidoli. On the use of the term diastema in ancient Greek constructions. Historia Mathematica, 31(1):2–10, February 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086003000764.

Sidoli:2006:BRC

[Sid06a] Nathan Sidoli. Book review: Classics in the History of Greek Mathematics, Jean Christianidis (Ed.). Kluwer Academic, Dordrecht (2004), ISBN 1-4020-0081-2. Historia Mathematica, 33(4):491–493, November 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086005001035.

Sidoli:2006:BRE

[Sid06b] Nathan Sidoli. Book review: Euclid’s Data: the Importance of Being Given. Museum Tusculanum Press, Copenhagen (2003), ISBN 87-7289-815-1. Historia Mathematica, 33(2):248–250, May 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086005000649.

Sidoli:2012:BRT

[Sid12] Nathan Camillo Sidoli. Book review: Tenbin no Majutsushi: Arukimedesu no Sugaku (Sorcerer of the Scales: Archimedes’ Mathematics). By Eiji Hayashi and Ken Saito. Tokyo (Kyoritsu Shuppan). 2009. ISBN 978-4-320-01910-2. x + 249 pp. + 8 color plates. = Y3300.

Historia Mathematica, 39(2):222–224, May 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086011000280.

Sidoli:2013:BRS

[Sid13] Nathan Sidoli. Book review: A Survey of the Almagest. Springer, New York (2010), ISBN: 978-0-387-84825-9. Historia Mathematica, 40(1):92–93, February 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086012000717.

Sidoli:2014:MTP

[Sid14] Nathan Sidoli. Mathematical tables in Ptolemy’s Almagest. Historia Mathematica, 41(1):13–37, February 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086013000530.

Sidoli:2024:BRA

[Sid24] Nathan Sidoli. Book review: Aristarchus of Samos: On the Sizes and Distances of the Sun and the Moon by Christián C. Carman and Rodolfo P. Buzón. London (Routledge) 2023. Historia Mathematica, 67(??):13–14, June 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086024000107.

Silva:2008:ACB

[Sil08] M. Céu Silva. The algebraic content of Bento Fernandes’s Tratado da arte de arismetica (1555). Historia Mathematica, 35(3):190–219, August 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086008000062.

Simon:1996:BRG

[Sim96] Gérard Simon. Book review: Géométrie et dioptrique au X^{e} siècle. Ibn Sahl, al–Qûhî et Ibn al-Haytham. By Roshdi Rashed. Paris (Les Belles Lettres). 1993. Historia Mathematica, 23(2):210–216, May 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508609690021X.

Simonson:2000:MPGa

[Sim00a] Shai Simonson. The missing problems of Gersonides — a critical edition, I. Historia Mathematica, 27(3):243–302, August 2000.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086000922808.

Simonson:2000:MPGb

[Sim00b] Shai Simonson. The missing problems of Gersonides — a critical edition, II. Historia Mathematica, 27(4):384–431, November 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508600092281X.

Singh:1985:CFN

[Sin85a] Parmanand Singh. The so-called Fibonacci numbers in Ancient and Medieval India. Historia Mathematica, 12(3):229–244, August 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086085900217.

Sinha:1985:SPE

[Sin85b] Kripa Nath Sinha. ’Srîpati: an eleventh-century Indian mathematician. Historia Mathematica, 12(1):25–44, February 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086085900667.

Sisma:2002:GHR

[Sis02] Pavel Sisma. Georg Hamel and Richard von Mises in Brno. Historia Mathematica, 29(2):176–192, May 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086002923463.

Sreeram:2024:SYM

[SK24] G. Sreeram and Aditya Kolachana. The sam .yoga-meru: a combinatorial tool in the Saṅgita-ratnakara. Historia Mathematica, 66(??): 14–25, March 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086024000053.

Stoyan:2018:JWP

[SM18] Dietrich Stoyan and Thomas Morel. Julius Weisbach’s pioneering contribution to orthogonal linear regression (1840). Historia Mathematica, 45(1):75–84, February 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086017301167.

Smadja:2016:TCS

[Sma16] Ivahn Smadja. On two conjectures that shaped the historiography of indeterminate analysis: Strachey and Chasles on Sanskrit sources. Historia Mathematica, 43(3):241–287, August 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086016300015.

Smillie:1979:HAC

[Smi79a] K. W. Smillie. An historical approach to computing science. Historia Mathematica, 6(1):63–66, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079901071.

Smithies:1979:BRO

[Smi79b] F. Smithies. Book review: Ocherki istorii teorii funktsii deistvitel’nogo peremennogo: (Essays on the history of the theory of functions of a real variable). By F. A. Medvedev. Moscow (Izdatel’stvo Nauka). 1975. 248 pp. Historia Mathematica, 6(2):218–220, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079900946.

Smith:1980:TBF

[Smi80] G. C. Smith. Thomas Bayes and fluxions. Historia Mathematica, 7 (4):379–388, November 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086080900026.

Smith:1982:MOA

[Smi82] G. C. Smith. Matthew O’Brien’s anticipation of vectorial mathematics. Historia Mathematica, 9(2):172–190, May 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082900027.

Soifer:2009:MCB

[Soi09] Alexander Soifer. The Mathematical Coloring Book: Mathematics of Coloring and the Colorful Life of its Creators. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009.

ISBN 0-387-74640-4 (hardcover). xxx + 607 pp. LCCN QA166 .S65 2009.

Sorensen:2005:ECU

[Sør05] Henrik Kragh Sørensen. Exceptions and counterexamples: Understanding Abel’s comment on Cauchy’s theorem. Historia Mathematica, 32(4):453–480, November 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086004000916.

Sorrenson:2013:PMI

[Sor13] Richard Sorrenson. Perfect mechanics: instrument makers at the Royal Society of London in the eighteenth century. Docent Press, Boston, MA, USA, 2013. ISBN 0-9887449-2-9. ix + 240 pp. LCCN Q185 .S67 2013.

Sorensen:2014:CAE

[Sør14] Henrik Kragh Sørensen. Confluences of agendas: Emigrant mathematicians in transit in Denmark, 1933–1945. Historia Mathematica, 41(2):157–187, May 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086013000517.

Sorensen:2017:MP

[Sør17] Henrik Kragh Sørensen. Montucla Prize 2017. Historia Mathematica, 44(4):318–319, November 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086017300861.

Spalt:1989:RGI

[Spa89] Detlef D. Spalt. Resultatismus oder geneseologie? — Impulse der Nichtstandard-Analysis für die Geschichtsschreibung: Arbeitstagung vom 9.–11. Januar 1989 in Darmstadt, Technische Hochschule. Historia Mathematica, 16(4):382–385, November 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086089900876.

Spalt:2011:WFG

[Spa11] Detlef D. Spalt. Welche Funktionsbegriffe gab Leonhard Euler?. (German) [Which function terms did Leonhard Euler give?]. Historia Mathematica, 38(4):485–505, November 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086011000309.

Spira:1976:BRR

[Spi76] Robert Spira. Book review: Riemann’s zeta function: By Harold M. Edwards. New York. (Academic Press). 1974. 315 pp. Historia Mathematica, 3(4):489–490, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086076900872.

Spiesser:2000:PLB

[Spi00a] Maryvonne Spiesser. Problèmes linéaires dans le book review: Compendy de la praticque des nombres de Barthélemy de Romans et Mathieu Préhoude (1471): Une approche nouvelle basée sur des sources proches du Liber abbaci de Léonard de Pise. Historia Mathematica, 27(4):362–383, November 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086000922857.

Spiesser:2000:PLC

[Spi00b] Maryvonne Spiesser. Problèmes linéaires dans le Compendy de la praticque des nombres de Barthélemy de Romans et Mathieu Préhoude (1471): une approche nouvelle basée sur des sources proches du Liber abbaci de Léonard de Pise. (French) [Linear problems in the Compendium of the number practice by Barthélemy de Romans and Mathieu Préhoude (1471): an new approach based on sources near Liber abbaci of Leonardo of Pisa]. Historia Mathematica, 27(4):362–383, 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Sprott:1978:GCS

[Spr78] D. A. Sprott. Gauss’s contributions to statistics. Historia Mathematica, 5(2):183–203, May 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086078900496.

Swetz:1984:BCB

[SS84] Frank J. Swetz and Ang Tian Se. A brief chronological and bibliographic guide to the history of Chinese mathematics. Historia Mathematica, 11(1):39–56, February 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086084900041.

Se:1986:CMC

[SS86] Ang Tian Se and Frank J. Swetz. A Chinese mathematical classic of the third century: the sea island mathematical manual of Liu hui.

Historia Mathematica, 13(2):99–117, May 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086086900248.

Siegmund-Schultze:1988:BWS

[SS88] Reinhard Siegmund-Schultze. Der Beweis des Weierstraßschen Approximationssatzes 1885 vor dem Hintergrund der Entwicklung der Fourieranalysis. (French) [On the evidence for the Weierstrass Approximation Theorem in 1885 against the background of Fourier analysis]. Historia Mathematica, 15(4):299–310, November 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086088900237.

Siegmund-Schultze:1989:BRR

[SS89] Reinhard Siegmund-Schultze. Book review: Rudolf Lipschitz: Briefwechsel mit Cantor, Dedekind, Helmholtz, Kronecker, Weierstrass und Anderen: Bearbeitet von Winfried Scharlau (Dokumente zur Geschichte der Mathematik, im Auftrag der Deutschen Mathematiker-Vereinigung herausgegeben von Winfried Scharlau, Band 2). Braunschweig/Wiesbaden (Vieweg). 1986. xviii + 253 Seiten. Historia Mathematica, 16(4):387–389, November 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608990089X.

Siegmund-Schultze:1993:HGM

[SS93] Reinhard Siegmund-Schultze. Hilda Geiringer-von Mises, Charlier series, ideology, and the human side of the emancipation of applied mathematics at the University of Berlin during the 1920s. Historia Mathematica, 20(4):364–381, November 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086083710311.

Siegmund-Schultze:1994:SCM

[SS94] Reinhard Siegmund-Schultze. “Scientific control” in mathematical reviewing and German–U.S.–American relations between the two World Wars. Historia Mathematica, 21(3):306–329, August 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086084710275.

Siegmund-Schultze:1997:BRF

[SS97a] Reinhard Siegmund-Schultze. Book review: Das Fotoalbum für Weierstraß. A Photo Album for Weierstrass. Edited by Reinhard Bölling. Braunschweig and Wiesbaden (Friedrich Vieweg & Sohn). 1994. Historia Mathematica, 24(1):100–101, February 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086097921388.

Siegmund-Schultze:1997:EMR

[SS97b] Reinhard Siegmund-Schultze. The emancipation of mathematical research publishing in the United States from German dominance (1878–1945). Historia Mathematica, 24(2):135–166, May 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086096921321.

Siegmund-Schultze:2007:BRL

[SS07] Reinhard Siegmund-Schultze. Book review: Landmark Writings in Western Mathematics 1640—1940, I. Grattan-Guinness (Ed.). Elsevier, Amsterdam (2005), ISBN 0-444-50871-6. Historia Mathematica, 34(1):114–115, February 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086005001047.

Siegmund-Schultze:2008:BRB

[SS08] Reinhard Siegmund-Schultze. Book review: Biographisches Lexikon in Mathematik promovierter Personen an deutschen Universitäten und Technischen Hochschulen WS 1907/08 bis WS 1944/45, R. Tobies. Erwin Rauner Verlag, Augsburg (2006), ISBN 3-936905-21-5. Historia Mathematica, 35(1):56–57, February 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086007000328.

Siegmund-Schultze:2009:MFN

[SS09] R. (Reinhard) Siegmund-Schultze. Mathematicians fleeing from Nazi Germany: individual fates and global impact. Princeton University Press, Princeton, NJ, USA, 2009. ISBN 0-691-12593-7 (hardcover),0-691-14041-3 (paperback). xxv + 471 pp. LCCN QA27.G4 S53 2009.

Saito:2010:FDA

[SS10a] Ken Saito and Nathan Sidoli. The function of diorism in ancient Greek analysis. Historia Mathematica, 37(4):579–614, November 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009001359.

Siegmund-Schultze:2010:MOC

[SS10b] Reinhard Siegmund-Schultze. On a missed opportunity for collaboration between historians and mathematicians: a biographical avalanche triggered by Professor Ioan James, FRS. Historia Mathematica, 37(4):693–707, November 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086010000042.

Siegmund-Schultze:2010:SVT

[SS10c] Reinhard Siegmund-Schultze. Sets versus trial sequences, Hausdorff versus von Mises: “Pure” mathematics prevails in the foundations of probability around 1920. Historia Mathematica, 37(2):204–241, May 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086009001311.

Sidoli:2012:CAG

[SS12a] Nathan Sidoli and Ken Saito. Comparative analysis in Greek geometry. Historia Mathematica, 39(1):1–33, February 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508601100070X.

Siegmund-Schultze:2012:HWI

[SS12b] Reinhard Siegmund-Schultze. Hans Wußing (1927–2011) and the blooming of the history of mathematics and sciences in the German Democratic Republic — a biographical essay. Historia Mathematica, 39(2):143–173, May 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086012000055.

Siegmund-Schultze:2017:RJM

[SS17] Reinhard Siegmund-Schultze. Book review: Recollections of a Jewish Mathematician in Germany, by Abraham A. Fraenkel. Edited by Jiska Cohen-Mansfield, translated by Allison Brown. Springer/ Birkhäuser, Basel (2016), ISBN: 978-3-319-30845-6. Historia Mathematica, 44(3):288–292, August 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086017300277.

Siegmund-Schultze:2018:IVS

[SS18] Reinhard Siegmund-Schultze. The interplay of various Scandinavian mathematical journals (1859–1953) and the road towards internationalization. Historia Mathematica, 45(4):354–375, November 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S031508601830096X.

Siegmund-Schultze:2022:FHC

[SS22a] Reinhard Siegmund-Schultze. Felix Hausdorff’s collected works — a meta-review. Historia Mathematica, 60(??):53–64, August 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086022000441.

Siegmund-Schultze:2022:LSH

[SS22b] Reinhard Siegmund-Schultze. From lattices via social history to theories of modernity in mathematics: a biographical essay for Herbert Mehrtens (1946–2021). Historia Mathematica, 58(??):17–34, February 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086022000040.

Sidoli:2017:EN

[SSS17] Nathan Camillo Sidoli and Reinhard Siegmund-Schultze. Editorial note. Historia Mathematica, 44(3):201, August 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086017300538.

Sidoli:2020:EN

[SSS20] Nathan Sidoli and Reinhard Siegmund-Schultze. Editorial note. Historia Mathematica, 50(??):1, February 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086020300069.

Siegmund-Schultze:2007:RMP

[SSZ07] Reinhard Siegmund-Schultze and Sandy Zabell. Richard von Mises and the “problem of two races”: a statistical satire in 1934. Historia Mathematica, 34(2):206–220, May 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086006000590.

Schappacher:2016:ENH

[ST16] Norbert Schappacher and Cordula Tollmien. Emmy Noether, Hermann Weyl, and the Göttingen Academy: a marginal note. Historia Mathematica, 43(2):194–197, ???? 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086015001007.

Steneck:1980:BRW

[Ste80] Nicholas H. Steneck. Book review: Witelonis perspectivae liber primus: (Book I of Witelo’s Perspectiva). An English Translation with Introduction and Commentary and Latin Edition of the Mathematical Book of Witelo’s Perspectiva. By Sabetai Unguru. Wroclaw (The Polish Academy of Sciences). 1977. Studia Copernicana, Vol.

15. 330 pp., 12 plates. Historia Mathematica, 7(2):199–200, May 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608090052X.

Stedall:2001:OON

[Ste01] Jacqueline A. Stedall. Of our own nation: John Wallis’s account of mathematical learning in medieval England. Historia Mathematica, 28(2):73–122, May 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086000923073.

Stedall:2002:DCA

[Ste02] Jacqueline A. Stedall. A discourse concerning algebra: English algebra to 1685. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-852495-1. xii + 294 pp. LCCN QA151 .S74

2002. URL http://www.loc.gov/catdir/enhancements/fy0613/
2002029001-d.html;   http://www.loc.gov/catdir/toc/fy037/
2002029001.html.

Stedall:2005:BRO

[Ste05] Jackie Stedall. Book review: Oxford Figures: 800 Years of the Mathematical Sciences, John Fauvel, Raymond Flood, Robin Wilson. Oxford University Press, Oxford (2000), ISBN 0-19-852309-2. Historia Mathematica, 32(4):492–494, November 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086005000133.

Stedall:2007:SCV

[Ste07a] Jacqueline Stedall. Symbolism, combinations, and visual imagery in the mathematics of Thomas Harriot. Historia Mathematica, 34 (4):380–401, November 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S031508600700033X.

Steele:2007:AA

[Ste07b] John Steele. Asger Aaboe (1922–2007). Historia Mathematica, 34 (4):377–379, November 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086007000596.

Stedall:2009:BRM

[Ste09] Jacqueline Stedall. Book review: ‘Magic is no magic’: the wonderful world of Simon Stevin, J. T. Devreese, G. Vanden Berghe. WIT Press, Southampton, Boston (2008), ISBN 978-1-84564-092-7. Historia Mathematica, 36(1):79–80, February 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008000323.

Steele:2011:JPB

[Ste11] John Steele. John P. Britton (1939–2010). Historia Mathematica, 38(3):335–337, August 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086011000267.

Stedall:2012:JWF

[Ste12] Jacqueline Stedall. John Wallis and the French: his quarrels with Fermat, Pascal, Dulaurens, and Descartes. Historia Mathematica, 39(3):265–279, August 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086012000080.

Stenhouse:2020:MSE

[Ste20] Brigitte Stenhouse. Mary Somerville’s early contributions to the circulation of differential calculus. Historia Mathematica, 51(??):1– 25, May 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086019300771.

Stigler:1974:GPD

[Sti74] Stephen M. Stigler. Gergonne’s 1815 paper on the design and analysis of polynomial regression experiments. Historia Mathematica, 1 (4):431–439, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074900330.

Stigler:1976:APG

[Sti76a] Stephen M. Stigler. The anonymous Professor Gergonne. Historia Mathematica, 3(1):71–74, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086076900094.

Stipanic:1976:VFM

[Sti76b] Ernest Stipanić. Vzaimosvyaz filosofii i matematiki v protsesse istoricheskogo razvitiya ot epokhi vosrozhdeniya do nachala XX veka (Russian) [The relationship between philosophy and mathematics in the process of historical development from the Renaissance to the beginning of the XXth century]. Historia Mathematica, 3(3):344–350, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086076901166.

Stigler:1977:AGP

[Sti77] Stephen M. Stigler. An attack on Gauss, published by Legendre in 1820. Historia Mathematica, 4(1):31–35, February 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900325.

Stipanic:1978:HPM

[Sti78] Ernest Stipanić. History and philosophy of mathematics as a subject of the regular curriculum at the Scientific–Mathematical University of Belgrade. Historia Mathematica, 5(3):342–345, August 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086078901210.

Stone:1976:REW

[Sto76] Marshall Stone. A reminiscence on the extension of the Weierstrass Approximation Theorem. Historia Mathematica, 3(3):328, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076901051.

Stolze:1978:HDT

[Sto78] Charles H. Stolze. A history of the divergence theorem. Historia Mathematica, 5(4):437–442, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078902124.

Struik:1978:BRB

[Str78] D. J. Struik. Book review: Briefwechsel zwischen Alexander von Humboldt und Carl Friedrich Gauss: Zum 200. Geburtstag von C. F. Gauss im Auftrage des Gauss-Komitees bei der Akademie der Wissenschaften der DDR, neu herausgegeben durch Kurt-R. Biermann. Berlin (Akademie-Verlag). 1977. 202 p. 36 mark. Historia Mathematica, 5(3):352–353, August 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086078901246.

Struik:1981:EBC

[Str81] Dirk J. Struik. Elwin Bruno Christoffel: 10 November 1829, Monschau; 5 März 1900, Strassburg; Professor den Mathematik in Zürich, Berlin und Strassburg. Gedenkschrift zur 150. Wiederkehr des Geburtstages. Heimatblätter des Kreises Aachen 34/35 Jahrgang. 1978-3/4, 1979–1. 20 pp. Historia Mathematica, 8(4):481–482, November 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086081900586.

Struik:1982:BRT

[Str82a] Dirk J. Struik. Book review: Two decades of mathematics in the Netherlands 1920–1940. A retrospection on the occasion of the bicentennial of the Wiskundig Genootschap: Edited by E. M. J. Bertin, H. J. M. Bos, and A. W. Grootendorst. Part I, xxiv + 193 pp.; Part II, pp. 195-379. Amsterdam (Mathematical Centre). 1978. Historia Mathematica, 9(1):99–100, February 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508608290146X.

Struik:1982:MMN

[Str82b] Dirk J. Struik. Minoan and Mycenaean numerals. Historia Mathematica, 9(1):54–58, February 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086082901379.

Strobl:1985:WAH

[Str85] Walter Strobl. Aus den wissenschaftlichen anfängen Hermann Minkowskis. (German) [From the scientific beginnings of Hermann Minkowski]. Historia Mathematica, 12(2):142–156, May 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086085900047.

Struik:1994:MTE

[Str94] Dirk J. Struik. On my trip to Europe, August 1–9, 1989. Historia Mathematica, 21(3):276–279, August 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086084710251.

Strickland:2023:HLT

[Str23] Lloyd Strickland. How Leibniz tried to tell the world he had squared the circle. Historia Mathematica, 62(??):19–39, February 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086022000465.

Stuewer:1980:EH

[Stu80] Roger H. Stuewer. Erwin N. Hiebert. Historia Mathematica, 7 (3):229–233, August 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086080900245.

Swetz:1974:IMH

[Swe74] Frank J. Swetz. The introduction of mathematics in higher education in China, 1865–1887. Historia Mathematica, 1(2):167–179, May 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074900056.

Swerdlow:1979:BRH

[Swe79] N. M. Swerdlow. Book review: A history of ancient mathematical astronomy: Studies in the History of Mathematics and Physical Sciences, Vol. 1, in 3 parts. By O. Neugebauer. Berlin/ Heidelberg/New York (Springer-Verlag). 1975. 1456 + xxi pp. (incl. figs., plates). $124.70. Historia Mathematica, 6(1):76–85, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079901186.

Swetz:1986:GAA

[Swe86] Frank J. Swetz. Geometry and algebra in ancient civilizations. Historia Mathematica, 13(1):83–85, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086086902417.

Swetz:1992:ICH

[Swe92] Frank J. Swetz. The International Conference on the History of Mathematics and Education in History of Mathematics: Baotou, Inner Mongolia, The People’s Republic of China, August 9–14, 1991. Historia Mathematica, 19(2):208–209, May 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609290084O.

Swetz:2000:BRM

[Swe00] Frank Swetz. Book review: The Measure of Reality: Quantification and Western Society, 1250–1600. By Alfred W. Crosby. Cambridge. Historia Mathematica, 27(3):303–305, August 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086000922845.

Swerdlow:2001:APO

[Swe01] N. M. Swerdlow. Astronomical papyri from Oxyrhynchus. Historia Mathematica, 28(2):129–132, May 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086001923055.

Struik:1990:RKJ

[SY90] Dirk Struik and Adolf P. Yushkevich. Remarks of Kirk J. Struik and Adolf P. Yushkevich on receiving the K. O. May Prize in Hamburg, August 3, 1990 (see Historia Mathematica 17(1) (1990), 1–3). Historia Mathematica, 17(4):382–384, November 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609090036D.

Sylla:1981:BRP

[Syl81] Edith Sylla. Book review: Practical geometry in the high middle ages: Artis culuslibet consummatio and the pratike de geometrie: Edited with translation and commentary by Stephen K. Victor. Philadelphia (The American Philosophical Society). 1979. Memoirs, Vol. 134. xii + 638 pp., illustrated, indices of technical terms. Historia Mathematica, 8(3):364–367, August 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508608190077X.

Sylla:1983:BRM

[Syl83] Edith Dudley Sylla. Book review: The medieval Latin translation of the Data of Euclid: By Shuntaro Ito. Tokyo/Boston, Basel, Stuttgart (University of Tokyo Press/Birkhaüser). 1980. ix + 256 pp. $34.00. Historia Mathematica, 10(4):471–475, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900150.

Sylla:2003:BRW

[Syl03] Edith Dudley Sylla. Book review: Die Werke von Jakob Bernoulli. Vol. 5, Differentialgeometrie: Edited by André Weil and Martin Mattmüller. Basel/Boston/Berlin (Birkhäuser Verlag). 1999. xxv + 445 pp. EUR 211. Historia Mathematica, 30(3):378–380, August 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086003000454.

Sylla:2007:BRQ

[Syl07] Edith Dudley Sylla. Book review: Questiones circa Tractatum Proportionum Magistri Thome Braduardini By Blaise de Parme. Edited by Joel Biard and Sabine Rommevaux. Textes Philosophiques du Moyen Age XXII. Paris (Vrin). 2005. Historia Mathematica, 34(2):227–229, May 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086006000553.

Synowiec:1983:DEM

[Syn83] John Synowiec. Distributions: the evolution of a mathematical theory. Historia Mathematica, 10(2):149–183, May 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900617.

Syta:1991:VMH

[Syt91] G. Syta. Voronoy memorial house. Historia Mathematica, 18(4):363, November 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609190377A.

Szabo:1974:MPF

[Sza74] Árpád Szabó. “Die Muse der Pythagoreer” [1] Zur Frühgeschichte der Geometrie. (German) [The Muse of the Pythagoreans [1] Toward an early history of geometry]. Historia Mathematica, 1(3):291–316, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074900706.

Taisbak:2014:CRI

[Tai14] Christian Marinus Taisbak. Cube roots of integers. A conjecture about Heron’s method in Metrika III. 20. Historia Mathematica, 41(1):103–106, February 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086013000529.

Tanner:1974:BRH

[Tan74] R. C. H. Tanner. Book review: Humanism and the rise of science in Tudor England: By Antonia McLean. New York (Neale Watson Academic Publications) and London (Heineman Educational Books). 1972. 258 pp. U.S. $10.50, U.K. £3.75. Historia Mathematica, 1 (3):348–350, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074900858.

Tankard:1979:HGW

[Tan79] James W. Tankard, Jr. The H. G. Wells quote on statistics: a question of accuracy. Historia Mathematica, 6(1):30–33, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079901010.

Taton:1975:BRL

[Tat75] René Taton. Book review: Lazare Carnot savant: By Charles Coulston Gillispie. A monograph treating Carnot’s scientific work, with facsimile reproduction of his unpublished writings on mechanics and on the calculus, and an essay concerning the latter by A. P. Youschkevitch. Princeton (Princeton University Press). 1971. 370 p. U.S. $17.50. Historia Mathematica, 2(2):204–207, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075901548.

Taton:1986:FAF

[Tat86] René Taton. Sur les fondements de l’analyse à la fin du XVIII^{e} d’après le traité de Lacroix: Par Pierre Dugac. Université Pierre et Marie Curie (Paris). DEUGSSM, 2^{e} année, option d’histoire des sciences (oct.–déc. 1982). 39 pp. Historia Mathematica, 13(3):299–301, August 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086086901035.

Tattersall:1987:TJD

[Tat87] James J. Tattersall. Thomas Jefferson and Douwes’ method for determining latitude. Historia Mathematica, 14(3):275–281, August 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086087900462.

Tattersall:1988:WPC

[Tat88] James J. Tattersall. Who put the “C” in A.-T. Vandermonde? Historia Mathematica, 15(4):361–367, November 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900274.

Taton:1991:PC

[Tat91] René Taton. Pierre Costabel (1912–1989). Historia Mathematica, 18(3):292–295, August 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609190406N.

Tattersall:1992:NSB

[Tat92] J. J. Tattersall. Nicholas Saunderson: the blind Lucasian professor. Historia Mathematica, 19(4):356–370, November 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290002S.

Tattersall:2011:BRT

[Tat11] James J. Tattersall. Book review: Thomas Harriot’s Doctrine of Triangular Numbers: the ‘Magisteria Magna’. Edited by Janet Beery and Jacqueline Stedall. Zürich (European Mathematical Society). 2008. ISBN 978-3-03719-059-3. 144 pages. EUR 64.00. Historia Mathematica, 38(1):123–125, February 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086010000340.

Tazzioli:1997:RDP

[Taz97] Rossana Tazzioli. The role of differential parameters in Beltrami’s work. Historia Mathematica, 24(1):25–45, February 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086097921212.

Tazzioli:2001:GFS

[Taz01] Rossana Tazzioli. Green’s function in some contributions of 19th century mathematicians. Historia Mathematica, 28(3):232–252, August 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086001923158.

Tazzioli:2009:BRD

[Taz09] Rossana Tazzioli. Book review: Dalla “moderna geometria” alla “nuova geometria italiana.” Viaggiando per Napoli, Torino e dintorni. Lettere di: Sannia, Segre, Peano, Castelnuovo, D’Ovidio, Del Pezzo, Pascal e altri a Federico Amodeo. By F. Palladino and N. Palladino, with the collaboration of F. S. Tortoriello (Eds.). Firenze (Leo S. Olschki). 2006. ISBN 88-222-5529-1. lx, 570 pp. EUR 60. Historia Mathematica, 36(2):188–189, May 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008000682.

Tazzioli:2015:BRS

[Taz15] Rossana Tazzioli. Book review: Siméon-Denis Poisson. Les mathématiques au service de la science, Yvette Kosmann-Schwarzbach (Ed.). École polytechnique-Palaiseau, Paris (2013), ISBN: 978-2-7302-1584-8. Historia Mathematica, 42(3):354–357, August 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086014000548.

Tazzioli:2018:IBL

[Taz18] Rossana Tazzioli. Interplay between local and international journals: the case of Sicily, 1880–1920. Historia Mathematica, 45(4):334–353, November 2018. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086018300703.

Tapp:2005:KKW

[TC05] Christian Tapp and Georg Cantor. Kardinalität und Kardinäle: wissenschaftshistorische Aufarbeitung der Korrespondenz zwischen Georg Cantor und katholischen Theologen seiner Zeit. (German) [Cardinality and Cardinals: scientific historical analysis of the correspondence between Georg Cantor and Catholic theologians of his time], volume 53 of Boethius. Steiner, Stuttgart, Germany, 2005. ISBN 3-515-08620-X (hardcover). 607 pp. LCCN QA29.C34 T37 2005.

Taisbak:2003:DED

[TE03] C. M. (Christian Marinus) Taisbak and Euclid. Dedomena: Euclid’s Data, or, The importance of being given, volume 45 of Acta historica scientiarum naturalium et medicinalium. Museum Tusculanum Press, Copenhagen, Denmark, 2003. ISBN 87-7289-815-1. ISSN 0065-1311. 271 pp. LCCN QA31 .T349 2003.

Tee:1978:BRE

[Tee78] G. J. Tee. Book review: The exhaustive treatise on shadows: By Abu al-Rayhān Muhammad b. Ahmad al-Bīrūnī. Translation and Commentary by E. S. Kennedy. Aleppo (Institute for the History of Arabic Science, University of Aleppo). 1976. Volume 1, Translation (xviii + 281 pp.); Volume 2, Commentary (xviii + 223 pp.). Historia Mathematica, 5(3):354–357, August 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078901258.

Tee:1979:HGW

[Tee79] Garry J. Tee. H. G. Wells and statistics. Historia Mathematica, 6 (4):447–448, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086079900326.

Tee:1982:CBM

[Tee82] Garry J. Tee. Charles Babbage materials in New Zealand and Australia. Historia Mathematica, 9(3):344–345, August 1982.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082901306.

Tee:1984:BRS

[Tee84] G. J. Tee. Book review: Śiṡyadhīvṙddhida Tantra of Lalla with the commentary of Mallikārjuna Sūri, Part 2, translation and mathematical notes: Critical edition with introduction, English translation, notes, and indices by Bina Chatterjee. New Delhi (Indian National Science Academy). 1981. xxviii + 340 pp. Historia Mathematica, 11(3):327–329, August 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086084900405.

Terrall:1989:BRL

[Ter89] Mary Terrall. Book review: Leonhard Euler, Opera Omnia: Series IVA: Commercium epistolicum. Vol. VI: Correspondance de Leonhard Euler avec P.-L. M. de Maupertuis et Frédéric II. Edited by Pierre Costabel, Eduard Winter, Asot T. Grigorijan, and Adolph P. Juskevic. Basel (Birkhäuser). 1986. xi + 454 pp. SwF 192. Historia Mathematica, 16(1):96–97, February 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086089901067.

Tsaban:1998:RA

[TG98] Boaz Tsaban and David Garber. On the Rabbinical approximation of π. Historia Mathematica, 25(1):75–84, February 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086097921856.

Theisen:1984:ERS

[The84] Wilfred Theisen. Euclid, relativity, and sailing. Historia Mathematica, 11(1):81–85, February 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086084900077.

Thiem:1975:RMV

[Thi75] LêVăn Thiêm. Report on mathematics in Viet Nam. Historia Mathematica, 2(1):41–42, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508607590035X.

Thiele:1997:SCF

[Thi97] Rüdiger Thiele. On some contributions to field theory in the calculus of variations from Beltrami to Carathéodory. Historia Mathematica, 24(3):281–300, August 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086097921583.

Thiele:2000:R

[Thi00] Rüdiger Thiele. Reviews. Historia Mathematica, 27(1):77–79, February 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086099922565.

Thiele:2005:MS

[Thi05] Rüdiger Thiele. Matthias Schramm, 1928–2005. Historia Mathematica, 32(3):271–274, August 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086005000467.

Thiele:2006:BRG

[Thi06] Rüdiger Thiele. Book review: Gauss: Titan of Science, Revised edition, with contributions by Jeremy Gray and Fritz-Egbert Dohse, G. Waldo Dunnington. Mathematical Association of America, Washington, DC (2003), ISBN 0-88385-547-X. Historia Mathematica, 33(2):258–259, May 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086005000510.

Thiele:2007:BTK

[Thi07] Rüdiger Thiele. A birthday tribute to Karin Reich, with a selected bibliography. Historia Mathematica, 34(1):7–9, February 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086006000905.

Thiele:2008:BRE

[Thi08] Rüdiger Thiele. Book review: Euler and Modern Science, N. N. Bogolyubov, G. K. Mikhailov, A. P. Yushkevich (Eds.). The Mathematical Association of America, Washington, DC (2007), ISBN 978-0-88385-564-5. Historia Mathematica, 35(2):144–147, May 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086007001024.

Thomas:1976:BRR

[Tho76] Norman L. Thomas. Book review: On a research program in early modern physics: By Aant Elzinga. Stockholm (Kungl. Boktryckeriet P. A. Norstedt and Soner). 1972. 144p. Historia Mathematica, 3(2):226–229, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086076900549.

Thoren:1988:PR

[Tho88] Victor E. Thoren. Prosthaphaeresis revisited. Historia Mathematica, 15(1):32–39, February 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508608890047X.

Thomaidis:2011:SRM

[Tho11] Yannis Thomaidis. Some remarks on the meaning of equality in Diophantos’s arithmetica. Historia Mathematica, 38(1):28–41, February 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086010000364.

Timmons:2004:PAE

[Tim04] Todd Timmons. A prosopographical analysis of the early American mathematics publication community. Historia Mathematica, 31 (4):429–454, November 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086003000806.

Timmons:2009:BRH

[Tim09] Todd Timmons. Book review: Of the Human Heart: a Biography of Benjamin Peirce, Edward R. Hogan. Lehigh University Press, Bethlehem (2008), ISBN 978-0-934223-93-5. Historia Mathematica, 36(1):85–87, February 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008000335.

Toader:2021:PPP

[Toa21] Iulian D. Toader. Permanence as a principle of practice. Historia Mathematica, 54(??):77–94, February 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086020300628.

Tobies:1981:ITF

[Tob81] R. Tobies. IV. Tagung der Fachsektion Geschichte, Philosophie und Grundlagen der Mathematik der Mathematischen Gesellschaft der DDR. (German) [IVth Meeting of the History Section, Philosophy and the Mathematical Foundations of Mathematics Society of DDR]. Historia Mathematica, 8(3):358–360, August 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900732.

Tobies:1990:BRE

[Tob90] Renate Tobies. Book review: Die Entstehung des Mathematiklehrerberufs im 19. Jahrhundert: Studien und materialien zum Prozeßder professionalisierung in Preußen (1810–1870): By Gert Schubring. Bielefelder Beiträge zur Ausbildungsforschung und Studienreform, Bd. 2. Weinheim/Basel (Beltz). 1983. 325 S. Historia Mathematica, 17(2):182–185, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609090068O.

Tobies:2006:BLM

[Tob06] Renate Tobies. Biographisches Lexikon in Mathematik promovierter Personen: an deutschen Universitäten und Technischen Hochschulen, WS 1907/08 bis WS 1944/45. (German) [Biographical dictionary of mathematics Ph.D.’s at German academic technical universities, WS 1907/08 to WS 1944/45], volume Heft 58 of Algorismus. E. Rauner, Augsburg, Germany, 2006. ISBN 3-936905-21-5 (paperback). 402 pp. LCCN QA30 .T63 2006.

Tobies:2010:MMI

[Tob10] Renate Tobies. “Morgen möchte ich wieder 100 herrliche Sachen ausrechnen”: Iris Runge bei Osram und Telefunken. (German) [“Tomorrow I would like to find 100 beautiful things calculate”: Iris Runge at Osram and Telefunken], volume 61 of Boethius; Wissenschaftsgeschichte. Steiner, Stuttgart, Germany, 2010. ISBN 3-515-09638-8 (hardcover). 396 + 16 pp. LCCN QA29.R86 T63 2010.

URL http://bvbr.bib-bvb.de:8991/F?func=service\%26doc\
_library=BVB01\%26doc\_number=018976984\%26line\_number=
0001\%26func\_code=DB\_RECORDS\%26service\_type=MEDIA.

Toepell:1996:BRG

[Toe96] Michael Toepell. Book review: Geschichte der Algebra. Eine Einführung. Edited by Erhard Scholz. Unter Mitarbeit von Kirsti Andersen et al. Lehrbücher und Monographien zur Didaktik der Mathematik Bd. 16. Wien/Zürich (B.I.-Wissenschafts Verlag Mannheim). 1990. Historia Mathematica, 23(4):441–445, November 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086096900427.

Tonietti:1982:RPS

[Ton82] Tito Tonietti. A research proposal to study the formalist and intuitionist mathematicians of the Weimar Republic. Historia Mathematica, 9(1):61–64, February 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086082901392.

Toomer:1980:BRT

[Too80] G. J. Toomer. Book review: The translation of the Elements of Euclid from the Arabic into Latin by Hermann of Carinthia: Books VII–XII. Edited by H. L. L. Busard. (Mathematical Centre Tracts 84). Amsterdam (Mathematische Centrum). 1977. 198 pp. Historia Mathematica, 7(4):457–458, November 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086080900166.

Toth:1979:BRL

[Tot79] Imre Toth. Book review: Lobatchevski: By V. F. Kagan. Moscow (Edition Mir). 1974. Historia Mathematica, 6(1):91–97, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079901204.

Tournes:2003:LGE

[Tou03] Dominique Tournès. L’intégration graphique des équations différentielles ordinaires. Historia Mathematica, 30(4):457–493, November 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086003000338.

Trabucco:2010:LSD

[Tra10] Oreste Trabucco. L’opere stupende dell’arti più ingegnose: la recezione degli Pneumatika di Erone Alessandrino nella cultura italiana del Cinquecento. (Italian) [The wonderful and most ingenious works of art: the reception of Pneumatika of Heron of Alexandria in Italian culture of the Sixteenth Century], volume 47 of Quaderni di “Rinascimento” / Istituto nazionale di studi sul Rinascimento. L. S. Olschki, Firenze, Italia, 2010. ISBN 88-222-5990-4. 184 pp. LCCN QC142.H53 T73 2010. 20.00 EUR. URL http://www.loc.

gov/catdir/toc/casalini08/00925467.pdf.

Tracey:2021:DET

[Tra21] Kevin Gerard Tracey. ‘Disturbed’ by Euclid: Thomas Fincke and the reading of Ramist mathematics in sixteenth-century Germany. Historia Mathematica, 54(??):1–40, February 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086020300720.

Tropp:1974:BRC

[Tro74] Henry S. Tropp. Book review: A computer perspective: By the office of Charles and Ray Eames. Cambridge, Mass. (Harvard University Press), 1973, 174 pp., illustrated. $15. Historia Mathematica, 1(1):99–100, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086074901906.

Tropp:1975:ROC

[Tro75] Henry Tropp. The role of oral communication. Historia Mathematica, 2(4):465–468, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086075901044.

Tropp:1976:OHF

[Tro76] Henry S. Tropp. The origins and history of the fields medal. Historia Mathematica, 3(2):167–181, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086076900331.

Tropp:1983:BRH

[Tro83] Henry S. Tropp. Book review: A history of computing in the twentieth century: Edited by N. Metropolis, J. Howlett, and Gian-Carlo Rota. New York (Academic Press). 1980. xix + 659 pp. $29.50. Historia Mathematica, 10(2):233–236, May 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086083900721.

Truesdell:1975:BRL

[Tru75] C. Truesdell. Book review: Leibniz and dynamics: By Pierre Costabel. Translated by R. E. W. Maddison. Paris (Hermann) and London (Methuen) and Ithaca, N. Y. (Cornell University Press), 1973. 141 p. Historia Mathematica, 2(3):360–361, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900907.

Tian-Se:1978:CIR

[TS78] Ang Tian-Se. Chinese interest in right-angled triangles. Historia Mathematica, 5(3):253–266, August 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078901118.

Tuominen:2023:PBS

[Tuo23] Antti J. V. Tuominen. Pygmies, Bushmen, and savage numbers — a case study in a sequence of bad citations. Historia Mathematica, 62(??):51–72, February 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086022000672.

Turner:1981:ESP

[Tur81] Steven Turner. Epistemological and social problems of the development of the sciences in the early 19th century. Historia Mathematica, 8(3):355–357, August 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086081900720.

Turner:2008:BRM

[Tur08] Laura E. Turner. Book review: Mathematics and Social Utopias in France: Olinde Rodrigues and His Times, Simon Altmann, Eduardo L. Ortiz (Eds.). American Mathematical Society, Providence (2005), ISBN 978-0-8218-3860-0. Historia Mathematica, 35(1):52–53, February 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086007000067.

Turner:2013:MLT

[Tur13] Laura E. Turner. The Mittag-Leffler Theorem: the origin, evolution, and reception of a mathematical result, 1876–1884. Historia Mathematica, 40(1):36–83, February 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086012000766.

Turner:2020:CRI

[Tur20] Laura E. Turner. Cultivating a research imperative: Mentoring mathematics at Stockholms Högskola, 1882–1887. Historia Mathematica, 50(??):50–83, February 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086018301356.

Ullrich:2003:BRR

[Ull03] Peter Ullrich. Book review: Riemanns Einführung in die Funktionentheorie. Eine quellenkritische Edition seiner Vorlesungen mit einer Bibliographie zur Wirkungsgeschichte der Riemannschen Funktionentheorie: By Erwin Neuenschwander. Abhandlungen der Akademie der Wissenschaften in Göttingen. Göttingen (Vandenhoeck & Ruprecht). 1996. 232 pp. Historia Mathematica, 30(2):221–223, May 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086002000319.

Umezawa:1976:BRG

[Ume76] Toshio Umezawa. Book review: Giuseppe Peano: By Hubert C. Kennedy. Basel (Birkhäuser Verlag) 31 p. (Beiheft 14 of Elemente der Mathematik). Historia Mathematica, 3(3):341–342, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076901142.

Unguru:1978:TCP

[Ung78] Sabetai Unguru. A thirteenth century ‘proof’ of the parallel postulate. Historia Mathematica, 5(2):205–210, May 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086078900514.

Unguru:1985:BRD

[Ung85] Sabetai Unguru. Book review: Digging for structure into the Elements: Euclid, Hilbert, and Mueller Philosophy of mathematics and deductive structure in Euclid’s Elements: By Ian Mueller. Cambridge, Mass., London (MIT Press). 1981. xv + 378 pp. $45.00. Historia Mathematica, 12(2):176–184, May 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086085900138.

Unger:2020:ATW

[Ung20] J. Marshall Unger. On the acceptance of trigonometry in wasan: Evidence from a text of Aida Yasuaki. Historia Mathematica, 52(??):51– 65, August 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086020300471.

Unger:2022:MPE

[Ung22] J. Marshall Unger. Mathematics and philology: an example from wasan. Historia Mathematica, 61(??):18–24, November 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086022000246.

Unger:2023:CQS

[Ung23] J. Marshall Unger. Cyclic quadrilaterals: Solutions of two Japanese problems and their proofs. Historia Mathematica, 65(??):1–13, November 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086023000381.

Ustina:1974:HWG

[Ust74] Fred Ustina. Henry Wilbraham and Gibbs Phenomenon in 1848. Historia Mathematica, 1(1):83–84, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086074901736.

Vaccaro:2020:HON

[Vac20] Maria Alessandra Vaccaro. Historical origins of the nine-point conic. The contribution of Eugenio Beltrami. Historia Mathematica, 51(??): 26–48, May 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086019300783.

Vagaggini:1982:MTG

[Vag82] Rossella Vagaggini. A manuscript by Teofilo Gallaccini in the Siena municipal library. Historia Mathematica, 9(1):78–80, February 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082901422.

VanStigt:1979:RPB

[Van79] Walter P. Van Stigt. The rejected parts of Brouwer’s dissertation on the foundations of mathematics. Historia Mathematica, 6(4):385–404, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086079900259.

VanEgmond:1980:BRP

[Van80] Warren Van Egmond. Book review: Piero dell Francesca’s mathematical treatises: the “trattato d’abaco” and “libellus de quinque corporibus regularibus”: By Margaret Daly Davis. Ravenna (Longo Editore). 1977 xxii + 135 pp., 36 plates. Lit 8000. Historia Mathematica, 7(4):449–451, November 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086080900117.

VanEgmond:1983:AMD

[Van83a] Warren Van Egmond. The algebra of master Dardi of Pisa. Historia Mathematica, 10(4):399–421, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086083900034.

VanOss:1983:DFD

[Van83b] Rosine G. Van Oss. D’Alembert and the fourth dimension. Historia Mathematica, 10(4):455–457, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086083900071.

VanLooy:1984:CHA

[Van84] Herman Van Looy. A chronology and historical analysis of the mathematical manuscripts of Gregorius a Sancto Vincentio (1584–1667). Historia Mathematica, 11(1):57–75, February 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086084900053.

VanBrummelen:1998:MMT

[Van98] Glen Van Brummelen. Mathematical methods in the tables of planetary motion in Kushyar ibn Labban’s Jami^{c} Zīj. Historia Mathematica, 25(3):265–280, August 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086097921844.

VanBrummelen:2000:Aa

[Van00a] Glen Van Brummelen. Abstracts. Historia Mathematica, 27(1):86– 104, February 2000. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086099922693.

VanBrummelen:2000:Ab

[Van00b] Glen Van Brummelen. Abstracts. Historia Mathematica, 27(2):202– 221, May 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086000922766.

VanBrummelen:2000:Ac

[Van00c] Glen Van Brummelen. Abstracts. Historia Mathematica, 27(3):310– 329, August 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086000922821.

VanBrummelen:2000:Ad

[Van00d] Glen Van Brummelen. Abstracts. Historia Mathematica, 27(4): 443–462, November 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086000922900.

VanBrummelen:2001:Aa

[Van01a] Glen Van Brummelen. Abstracts. Historia Mathematica, 28(1):58– 71, February 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086000923000.

VanBrummelen:2001:Ab

[Van01b] Glen Van Brummelen. Abstracts. Historia Mathematica, 28(2):141– 163, May 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086001923146.

VanBrummelen:2001:Ac

[Van01c] Glen Van Brummelen. Abstracts. Historia Mathematica, 28(3):255– 280, August 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086001923225.

VanBrummelen:2001:Ad

[Van01d] Glen Van Brummelen. Abstracts. Historia Mathematica, 28(4): 315–333, November 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086001923304.

VanBrummelen:2002:A

[Van02] Glen Van Brummelen. Abstracts. Historia Mathematica, 29(4): 494–510, November 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086002923761.

VanBrummelen:2005:BRM

[Van05] Glen Van Brummelen. Book review: The Melon-Shaped Astrolabe in Arabic Astronomy, E. S. Kennedy, P. Kunitzsch, R. P. Lorch (Eds.). Franz Steiner Verlag, Stuttgart (1999). Historia Mathematica, 32(3):358–360, August 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086004000837.

VanEgmond:2006:BRC

[Van06] Warren Van Egmond. Book review: From China to Paris: 2000 Years Transmission of Mathematical Ideas. Edited by Yvonne Dold-Samplonius, Joseph W. Dauben, Menso Folkerts, and Benno van Dalen. Boethius, Band 46. Stuttgart (Franz Steiner Verlag). 2002. ISBN 3-515-08223-9. ix + 467 pp. Euro 68. Historia Mathematica, 33(3):361–362, August 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086005000960.

VanBrummelen:2009:MHE

[Van09] Glen Van Brummelen. The mathematics of the heavens and the earth: the early history of trigonometry. Princeton University Press, Princeton, NJ, USA, 2009. ISBN 0-691-12973-8 (hardcover). xvii + 329 pp.

LCCN QA24 .V36 2009. URL http://www.loc.gov/catdir/toc/
ecip0824/2008032521.html.

VanBrummelen:2010:AAA

[Van10] Glen Van Brummelen. Aspects of the astrolabe: ‘architectonica ratio’ in tenth- and eleventh-century Europe. Historia Mathematica, 37 (1):118–120, February 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086009000627.

Vandendriessche:2014:WWR

[Van14] Eric Vandendriessche. W. W. Rouse Ball and the mathematics of string figures. Historia Mathematica, 41(4):438–462, November 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086014000743.

Vandendriessche:2022:CNP

[Van22] Eric Vandendriessche. The concrete numbers of “primitive” societies: a historiographical approach. Historia Mathematica, 59(??): 12–34, May 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086022000076.

VanBrummelen:2024:DFN

[Van24] Glen Van Brummelen. Decimal fractional numeration and the decimal point in 15th-century Italy. Historia Mathematica, 66(??):1– 13, March 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086024000016.

vanBevern:2020:HNA

[vBS20] René van Bevern and Viktoriia A. Slugina. A historical note on the 3/2-approximation algorithm for the metric traveling salesman problem. Historia Mathematica, 53(??):118–127, November 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086020300240.

vanDalen:2009:JUT

[vD09] Benno van Dalen. Journals under threat: a joint response from history of science, technology and medicine editors. Historia Mathematica, 36(1):1–3, February 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086008000955.

vanDalen:2010:EK

[vD10] Benno van Dalen. Edward S. Kennedy: 1912–2009. Historia Mathematica, 37(2):159–163, May 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086010000030.

vanDalen:2007:E

[vDBG07] Benno van Dalen and June Barrow-Green. Editorial. Historia Mathematica, 34(1):1–2, February 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086006001443.

vanderWaerden:1975:SMB

[vdW75] B. L. van der Waerden. On the sources of my book Moderne algebra. Historia Mathematica, 2(1):31–40, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900348.

vanderWaerden:1976:BRE

[vdW76a] B. L. van der Waerden. Book review: The evolution of the Euclidean Elements: By Wilbur Richard Knorr. A Study of the Theory of Incommensurable Magnitudes and Its Significance for Early Greek Geometry. Dordrecht-Holland/ Boston (D. Reidel Publishing Co.). 1975. xii + 374 p. $49. Historia Mathematica, 3(4):497–499, November 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086076900926.

vanderWaerden:1976:DPR

[vdW76b] B. L. van der Waerden. D. Pingree’s review of my book: Science awakening II. Historia Mathematica, 3(2):209–211, May 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900409.

vanderWaerden:1980:BRL

[vdW80] B. L. van der Waerden. Book review: Vom logos zur analogie. Die Geschichte eines mathematischen Terminus: Von Thekal Horovitz. Zürich (Hans Rohr). 1978. Historia Mathematica, 7(4):452, November 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086080900130.

Vekerdi:1975:BRV

[Vek75] Laszlo Vekerdi. Book review: The vortex theory of planetary motions: By E. J. Aiton. London (Macdonald) and New York (American Elsevier). 1972. 282 p. $18.50. Historia Mathematica, 2(1):95–97, February 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086075900580.

Venezia:2002:CSI

[Ven02] Antonio Venezia. Un caso storico irrisolto: Le critiche di Popper alla logica quantistica di Birkhoff e von Neumann. (Italian) [a case of unsolved history: The criticism of Popper of the quantum logic of Birkhoff and von Neumann]. Historia Mathematica, 29(??):138–156, ???? 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Verin:2001:GAI

[Vér01] Hélène Vérin. Giambattista Aleotti e gli ingegneri del Rinascimento. (Italian) [Giambattista Aleotti and the engineers of the Renaissance]. Historia Mathematica, 28(2):125–127, May 2001. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086001923031.

Verburgt:2013:RLE

[Ver13] Lukas Verburgt. Robert Leslie Ellis’s work on philosophy of science and the foundations of probability theory. Historia Mathematica, 40(4):423–454, October 2013. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086013000451.

Verburgt:2015:OSM

[Ver15] Lukas M. Verburgt. The objective and the subjective in midnineteenth-century British probability theory. Historia Mathematica, 42(4):468–487, November 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086015000191.

Valerio:2022:NDW

[VF22] Miguel Valério and Silvia Ferrara. Numeracy at the dawn of writing: Mesopotamia and beyond. Historia Mathematica, 59(??):35–53, May 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086020300665.

vanHeijenoort:1967:FGS

[vH67] Jean van Heijenoort. From Frege to Gödel: a source book in mathematical logic, 1879–1931. Harvard University Press, Cambridge, MA, USA, 1967. ISBN 0-674-32450-1 (hardcover), 0-674-32449-8 (paperback). viii + 664 pp. LCCN QA9 .V35 1967. Reprinted 2002.

Verburgt:2016:YKP

[VHK16] Lukas M. Verburgt and Olga Hoppe-Kondrikova. On A. Ya. Khinchin’s paper ‘Ideas of intuitionism and the struggle for a subject matter in contemporary mathematics’ (1926): a translation with introduction and commentary. Historia Mathematica, 43(4):369–398, November 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508601630043X.

Vilain:1996:BRR

[Vil96] Christiane Vilain. Book review: La réforme de la dynamique, De corporum concursu (1678) et autres textes inédits. By G. W. Leibniz. Edition, introduction, French translations and commentaries by Michel Fichant. Paris (Librairie philosophique J. Vrin). 1994. Historia Mathematica, 23(2):207–210, May 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086096900208.

Vilain:1997:BRD

[Vil97] Christiane Vilain. Book review: La dynamique de Leibniz. By François Duchesneau. Collection Mathesis. Paris (Vrin). 1994. 367 pp. 195 FF. Historia Mathematica, 24(2):206–207, May 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086097921509.

Vilkko:1998:RFB

[Vil98] Risto Vilkko. The reception of Frege’s Begriffsschrift. Historia Mathematica, 25(4):412–422, November 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086098922133.

Vitrac:1996:BRA

[Vit96] Bernard Vitrac. Book review: Anaritus’ Commentary on Euclid. The Latin Translation, I–IV. Edited by P. M. J. E. Tummers. Artistarium, Supplementa, IX. Historia Mathematica, 23(4):445–446, November 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086096900439.

Villarino:1987:SED

[VJRZ87] Mark Villarino, Michael Josephy, and Angel Ruiz-Zúñiga. A Spanish edition of Disquisitiones Arithmeticae. Historia Mathematica, 14(2):195–196, May 1987. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086087900322.

VanBrummelen:2005:MHC

[VK05] Glen Van Brummelen and Michael Kinyon, editors. Mathematics and the historian’s craft: the Kenneth O. May lectures, volume 21 of CMS books in mathematics; Ouvrages de mathematiques de la SMC; 21. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-25284-3 (hardcover). xvi + 357 pp. LCCN QA21 .M3613 2005. URL http://www.loc.

gov/catdir/enhancements/fy0662/2005923503-d.html;   http:
//www.loc.gov/catdir/toc/fy0604/2005923503.html.

vanMaanen:1984:BRL

[vM84] Jan A. van Maanen. Book review: The land of Stevin and Huygens: Studies in the history of modern science: Vol. 7. By Dirk J. Struik. Dordrecht/Boston/London (Reidel). 1981. xx + 162 pp., 24 plates. Bibliography. Index. Bound $39.50, paperback $19.50. Translation (by D. J. Struik) of: Het Land van Stevin en Huygens: Sunschrift, Vol. 134. By D. J. Struik. 3rd rev. ed. Nijmegen (Sun). 1979. 156 pp., 16 plates. Bibliography. Glossary. Index. Hfl 18.50. Historia Mathematica, 11(2):234–237, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086084800233.

vanMaanen:1985:UMM

[vM85] Jan A. van Maanen. Unknown manuscript material of Christiaan Huygens. Historia Mathematica, 12(1):60–65, February 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086085900692.

vanMaanen:2007:BRJ

[vM07] Jan van Maanen. Book review: John Pell (1611–1685) and his correspondence with Sir Charles Cavendish. The mental world of an early modern mathematician, Noel Malcolm, Jacqueline Stedall. Oxford University Press, Oxford (2005), ISBN 0-19-856484-8. Historia Mathematica, 34(1):107–108, February 2007. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508600600019X.

Volterra:2009:MWV

[VMT09] Vito Volterra, Laurent Mazliak, and Rossana Tazzioli, editors. Mathematicians at war: Volterra and his French colleagues in World War I, volume 22 of Archimedes, 1385-0180. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 90-481-2739-4, 90-481-2740-8 (e). ix + 194 pp.

LCCN QA29.V64 A4 2009. URL http://www.loc.gov/catdir/
enhancements/fy1111/2009927019-d.html;   http://www.loc.
gov/catdir/enhancements/fy1111/2009927019-t.html.

Vogel:1983:VRC

[Vog83] Kurt Vogel. Ein Vermessungsproblem reist von China nach Paris. (German) [A measurement problem travels from China to Paris]. Historia Mathematica, 10(3):360–367, August 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086083900836.

Vogt:1994:SHM

[Vog94] Annette Vogt. Symposium on history of mathematics and mathematics teaching: BRDO/Manetin, Czechia, April 2–4, 1993. Historia Mathematica, 21(1):74, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086008800638.

Vogt:2010:BAR

[Vog10] Annette B. Vogt. Boris Abramovich Rosenfeld: (1917–2008). Historia Mathematica, 37(1):8–10, February 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086009000640.

Volodarsky:1977:BRC

[Vol77] Alexander Volodarsky. Book review: Census of the exact sciences in Sanskrit: By David Pingree. Series A. Philadelphia (American Philosophical Society). Vol. 1, 1970, 60 p. Vol. 2, 1971, 147 p. Vol. 3, 1976, 208 p. Historia Mathematica, 4(2):226–227, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077901240.

Volk:1981:BRB

[Vol81] Otto Volk. Book review: Briefwechsel zwischen Alexander von Humboldt und Heinrich Christian Schumacher: Von Kurt R. Biermann. Zum 200. Geburtstag von H. Ch. Schumacher. Beiträge zur Alexander-von-Humboldt-Forschung, Schriftenreihe der Alexandervon-Humboldt-Forschungsstelle der Akademie der Wissenschaften der DDR, Band 6, 192 S. mit 4 Tafeln. Berlin (Akademie-Verlag). 1979. Historia Mathematica, 8(4):485–487, November 1981. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086081900604.

Volkov:1991:DMV

[Vol91] Alexei Volkov. Dissertation on “Matematika v drevnem Kitaie (3-7 vv.)”. Historia Mathematica, 18(2):185–187, May 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609190507T.

Volkert:1993:CHW

[Vol93a] Klaus Volkert. Colloquium Heinrich Weber: Heidelberg, May 16, 1992. Historia Mathematica, 20(2):201–202, May 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086008800456.

Volkert:1993:HPT

[Vol93b] Klaus Volkert. On Helmholtz’ paper “Über die thatsächlichen Grundlagen der Geometrie”. Historia Mathematica, 20(3):307–309, August 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086083710268.

Volkov:1997:ZYH

[Vol97] Alexeüı Volkov. Zhao Youqin and his calculation of π. Historia Mathematica, 24(3):301–331, August 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086097921637.

Volkert:2002:BRC

[Vol02] Klaus Volkert. Book review: La correspondance entre Henri Poincaré et Gösta Mittag-Leffler. Avec en annexe les lettres échangées par Poincaré avec Fredholm, Gyldén et Phragmén. Presentation and notes by Philippe Nabonnand. Historia Mathematica, 29(2):204–209, May 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086001923365.

Volkert:2006:GKD

[Vol06] Klaus Volkert. Georg Klügel’s dissertation (1763) now accessible on the World-Wide Web. Historia Mathematica, 33(3):357–358, August 2006. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086006000231.

Volkov:2010:CUC

[Vol10] Alexei Volkov. Commentaries upon commentaries: the translation of the Jiu zhang suan shu by Karine Chemla and Guo Shuchun. Historia Mathematica, 37(2):281–301, May 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086009000603.

Volkert:2016:BRT

[Vol16] Klaus Thomas Volkert. Book review: La théorie des lignes parallèles de Johann Heinrich Lambert, A. Papadoupoulos, G. Théret. Blanchard, Paris (2004). Historia Mathematica, 43(2):198–206, ???? 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086015000439.

Volkert:2019:NM

[Vol19] Klaus Volkert. Note on models. Historia Mathematica, 48(??):87–95, August 2019. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086018301113.

Vonneuman:1988:BJN

[Von88] Nicholas A. Vonneuman. Biography of John von Neumann. Historia Mathematica, 15(3):275, August 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086088900687.

VonPlato:1993:OPD

[Von93] Jan Von Plato. Oresme’s proof of the density of rotations of a circle through an irrational angle. Historia Mathematica, 20(4):428–433, November 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086083710347.

Voss:1975:F

[Vos75] John Voss. Foreword. Historia Mathematica, 2(4):425–427, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075900968.

Vucinich:1999:MDS

[Vuc99] Alexander Vucinich. Mathematics and dialectics in the Soviet Union: the Pre-Stalin Period. Historia Mathematica, 26(2):107–124, May 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086099922292.

Vucinich:2000:SMD

[Vuc00] Alexander Vucinich. Soviet mathematics and dialectics in the Stalin era. Historia Mathematica, 27(1):54–76, February 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086099922619.

Vucinich:2002:SMD

[Vuc02] Alexander Vucinich. Soviet mathematics and dialectics in the Post-Stalin Era: New horizons. Historia Mathematica, 29(1):13–39, February 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086001923298.

VanDyck:2014:VWJ

[VV14] Maarten Van Dyck and Koen Vermeir. Varieties of wonder: John Wilkins’ mathematical magic and the perpetuity of invention. Historia Mathematica, 41(4):463–489, November 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508601400055X.

Wagner:1978:BWB

[Wag78] Ascher Wagner. A bibliography of William Burnside (1852–1927). Historia Mathematica, 5(3):307–312, August 1978. CODEN HI-MADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0315086078901167.

Wagner:1979:ECD

[Wag79] Donald Blackmore Wagner. An early Chinese derivation of the volume of a pyramid: Liu hui, third century A.D. Historia Mathematica, 6(2):164–188, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086079900764.

Wagner:1983:EII

[Wag83] Robert J. Wagner. Euclid’s intended interpretation of superposition. Historia Mathematica, 10(1):63–70, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900332.

Wagner:1985:PPT

[Wag85] Donald B. Wagner. A proof of the Pythagorean Theorem by Liu Hui (Third Century A.D.). Historia Mathematica, 12(1):71–73, February 1985. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086085900710.

Wagner:2015:CTO

[Wag15] Roy Wagner. Citrabhanu’s twenty-one algebraic problems in Malayalam and Sanskrit. Historia Mathematica, 42(3):263–279, August 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086015000178.

Wallis:1974:BBM

[Wal74] P. J. Wallis. A biobibliography of British mathematics and its applications up to 1850. Historia Mathematica, 1(4):449–454, November 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508607490038X.

Wallis:1975:BSL

[Wal75] Peter J. Wallis. The book subscription lists project — its relevance for historians of mathematics. Historia Mathematica, 2(3):321–326, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086075900750.

Walther:1994:RFM

[Wal94] Helen May Walther. Reminiscences of a former M.I.T. student. Historia Mathematica, 21(3):274–275, August 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S031508608471024X.

Walter:2017:HPS

[Wal17] Scott A. Walter. Book review: Henri Poincaré: A Scientific Biography, Jeremy Gray. Princeton University Press, Princeton (2013), 608 pp., $39.95. Historia Mathematica, 44(4):425–435, November 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086017300496.

Wang:2022:HJB

[Wan22] Xiaofei Wang. How Jean-Baptiste Delambre read ancient Greek arithmetic on the basis of the arithmetic of “complex numbers” at the turn of the 19th century. Historia Mathematica, 59(??):146–163, May 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086020300914.

Wardhaugh:2008:MLS

[War08] Benjamin Wardhaugh. Musical logarithms in the seventeenth century: Descartes, Mercator, Newton. Historia Mathematica, 35(1):19– 36, February 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086007000341.

Wardhaugh:2010:HRH

[War10] Benjamin Wardhaugh. How to read historical mathematics. Princeton University Press, Princeton, NJ, USA, 2010. ISBN 0-691-14014-6 (hardcover). xii + 116 pp. LCCN QA21 .W324 2010.

Wardhaugh:2015:LCC

[War15] Benjamin Wardhaugh. A ‘lost’ chapter in the calculation of π: Baron Zach and MS Bodleian 949. Historia Mathematica, 42(3):343–351, August 2015. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508601500018X.

Wardhaugh:2024:ETE

[War24] Benjamin Wardhaugh. Euclidean terms in European languages, 1482–1703. Historia Mathematica, 68(??):22–37, September 2024. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086024000739.

Waszek:2020:BRN

[Was20] David Waszek. Book review: New Light on George Boole, Desmond MacHale, Yvonne Cohen. Cork University Press (2018), 492 pp. Historia Mathematica, 51(??):91–93, May 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086019300795.

Waterhouse:1979:GI

[Wat79] William C. Waterhouse. Gauss on infinity. Historia Mathematica, 6(4):430–436, November 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086079900296.

Waterhouse:1983:QPB

[Wat83] William C. Waterhouse. Quantifier problems beyond uniformity. Historia Mathematica, 10(2):202–207, May 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086083900630.

Waterhouse:1984:NGR

[Wat84] William C. Waterhouse. A note by Gauss referring to ideals? Historia Mathematica, 11(2):142–146, May 1984. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086084800051.

Waterhouse:1986:NNS

[Wat86] William C. Waterhouse. A neglected note showing Gauss at work. Historia Mathematica, 13(2):147–156, May 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086086900273.

Waterhouse:1991:EIE

[Wat91] William C. Waterhouse. An elementary interpretation of entry 12 in Gauss’s Tagebuch. Historia Mathematica, 18(2):173–176, May 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609190498M.

Waterhouse:1993:HMD

[Wat93] William C. Waterhouse. Harmonic means and Diophantus I.39. Historia Mathematica, 20(1):89–91, February 1993. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086083710086.

Waterhouse:1995:CPA

[Wat95] William C. Waterhouse. On the cattle problem of Archimedes. Historia Mathematica, 22(2):186–187, May 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086085710178.

Watkins:2009:BRM

[Wat09] John J. Watkins. Book review: The Mathematical Coloring Book: Mathematics of Coloring and the Colorful Life of Its Creators, Alexander Soifer. Springer Science + Business Media, LLC, New York (2009), ISBN 978-0-387-74640-1. Historia Mathematica, 36(3):275– 277, August 2009. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086009000068.

Wagner-Dobler:1996:NCM

[WDB96] Roland Wagner-Döbler and Jan Berg. Nineteenth-Century mathematics in the mirror of its literature: a quantitative approach. Historia Mathematica, 23(3):288–318, August 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086096900294.

Weaver:1977:ESM

[Wea77] Warren Weaver. Early support for mathematics from Rockefeller agencies. Historia Mathematica, 4(4):443–444, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077900817.

Weeks:1991:LMG

[Wee91] Dennis Weeks. The life and mathematics of George Campbell, F.R.S. Historia Mathematica, 18(4):328–343, November 1991.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609190370D.

Wefelscheid:1989:ELM

[Wef89] Heinrich Wefelscheid. Edmund Landau memorabilia. Historia Mathematica, 16(1):86, February 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086089901006.

Weinstock:1992:NPI

[Wei92] Robert Weinstock. Newton’s Principia and inverse-square orbits: the flaw reexamined. Historia Mathematica, 19(1):60–70, February 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290055G.

Weinstock:1998:NPI

[Wei98] Robert Weinstock. Newton’s Principia and inverse-square orbits in a resisting medium: a spiral of twisted logic. Historia Mathematica, 25(3):281–289, August 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086098921929.

Weitzel:2004:FHP

[Wei04] Hans Weitzel. A further hypothesis on the polyhedron of A. Dürer’s engraving Melencolia I. Historia Mathematica, 31(1):11–14, February 2004. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086003000296.

Wiescher:2023:JPP

[Wie23] Michael Wiescher. Julius Plücker — a path from geometry to optics. Historia Mathematica, 64(??):19–33, August 2023. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086023000186.

Wilder:1974:HSC

[Wil74] R. L. Wilder. Hereditary stress as a cultural force in mathematics. Historia Mathematica, 1(1):29–46, February 1974.

CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086074901669.

Wilder:1975:BRA

[Wil75a] R. L. Wilder. Book review: Africa counts: By Claudia Zaslavsky. Boston (Prindle, Weber and Schmidt). 1973, 338 p, U.S. $12.50. Historia Mathematica, 2(2):207–210, May 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508607590155X.

Williams:1975:BRW

[Wil75b] Mary E. Williams. Book review: Women in mathematics: By Lynn M. Osen. Cambridge, Massachusetts (MIT Press), 1974. 185 pp. $5.95. Historia Mathematica, 2(3):348–349, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508607590083X.

Wilson:1976:NLO

[Wil76] John Wilson. New light on the origin of the four-color conjecture. Historia Mathematica, 3(3):329–330, August 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086076901063.

Wilkes:1977:BCP

[Wil77a] M. V. Wilkes. Babbage as a computer pioneer. Historia Mathematica, 4(4):415–440, November 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086077900799.

Williams:1977:BRG

[Wil77b] Pearce Williams. Book review: George Green, Miller, Snienton: By R. M. Bowley, L. J. Challis, F. W. Sheard, F. M. Wilkins Jones, and David Phillips. Nottingham (Museum and Art Gallery, The Castle). 1976. 96 p. Historia Mathematica, 4(2):232, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086077901306.

Wilder:1979:SCM

[Wil79a] R. L. Wilder. Some comments on M. J. Crowe’s review of evolution of mathematical concepts. Historia Mathematica, 6(1):57–62, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086079901058.

Williamson:1979:HMC

[Wil79b] Bruce Williamson. History of mathematics — a course outline. Historia Mathematica, 6(3):318–320, August 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086079901307.

Williamson:1980:RCF

[Wil80] Frank Williamson, Jr. Richard Courant and the finite element method: a further look. Historia Mathematica, 7(4):369–378, November 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086080900014.

William:1983:BRB

[Wil83] F. Aspray William. Book review: Brouwer’s Cambridge lectures on intuitionism: Edited by D. van Dalen. Cambridge (Cambridge University Press). 1981. xii + 109 pp. Appendix: Selections from Brouwer’s private notes. Editor’s preface and notes. Bibliography. Historia Mathematica, 10(1):132–135, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900575.

Wilkes:1990:BAM

[Wil90] Maurice V. Wilkes. The British Association Mathematical Tables Committee. Historia Mathematica, 17(2):152, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609090051E.

Wilson:1991:EJA

[Wil91] Curtis Wilson. Eric John Aiton (1920–1991). Historia Mathematica, 18(4):390–392, November 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508609190384A.

Wilson:1994:NES

[Wil94] Curtis Wilson. Newton on the equiangular spiral: an addendum to Erlichson’s account. Historia Mathematica, 21(2):196–203, May 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086084710184.

Wilson:1997:FFI

[Wil97] Robin Wilson. From five fingers to infinity: a journey through the history of mathematics. Historia Mathematica, 24(4):452–453, November 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086097921698.

Wilson:2005:BRK

[Wil05] Curtis Wilson. Book review: Kepler’s Philosophy and the New Astronomy, Rhonda Martens. Princeton University Press, Princeton, NJ (2000), ISBN 0-691-05069-4. Historia Mathematica, 32(4):485–487, November 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086004000941.

Wilson:2007:BRA

[Wil07] Robin Wilson. Book review: Arthur Cayley: Mathematician Laureate of the Victorian Age, Tony Crilly. The Johns Hopkins University Press, Baltimore (2006), ISBN 0-8018-8011-4. Historia Mathematica, 34(4):436–439, November 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086006001261.

Williams:2008:BRE

[Wil08a] Clemency Williams. Book review: Expounding the Mathematical Seed: a translation of Bhaskara I on the Mathematical Chapter of the Aryabhat.iya, Agathe Keller. Birkhaüser, Basel, Boston, Berlin (2006), ISBN 978-3-7643-7299-6. Historia Mathematica, 35(1):47–50, February 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086007000298.

Wilson:2008:BRJ

[Wil08b] Curtis Wilson. Book review: Jean le Rond d’Alembert, Oeuvres Complètes, Série I, Traités et mémoires mathématiques, 1736–1756, Volume 7, Précession et nutation (1749–1752), Michelle Chapront– Touzé, Jean Souchay (Eds.). CNRS Éditions, Paris (2006). Historia Mathematica, 35(4):329–332, November 2008. CODEN HIMADS.

ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086007000997.

Wilson:2008:LCN

[Wil08c] Robin J. Wilson. Lewis Carroll in numberland: his fantastical mathematical logical life: an agony in eight fits. W. W. Norton & Co., New York, NY, USA, 2008. ISBN 0-393-06027-6 (hardcover). xi + 237 pp. LCCN QA29.C353 W55 2008. URL http://www.loc.gov/

catdir/toc/ecip0826/2008037342.html.

Winterbourne:1982:APT

[Win82] Anthony T. Winterbourne. Algebra and pure time: Hamilton’s affinity with Kant. Historia Mathematica, 9(2):195–200, May 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086082900040.

Wisan:1976:BRG

[Wis76] Winifred L. Wisan. Book review: Galileo Revisited: an essay review of Galileo’s intellectual revolution: Middle period, 1610–1632: By William R. Shea. New York (Science History Publications). 1972. 204 pp. $15. Historia Mathematica, 3(1):103–109, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900264.

Wate-Mizuno:2014:MRD

[WM14] Mitsuko Wate-Mizuno. Mathematical recreations of Dénes Kőnig and his work on graph theory. Historia Mathematica, 41(4):377–399, November 2014. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086014000573.

Wolfson:1991:BRM

[Wol91] Paul R. Wolfson. Book review: Mathematics and its history: By John Stillwell. New York (Springer-Verlag). 1989. 371 pp. $49.80. Historia Mathematica, 18(1):80–81, February 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/031508609190361Z.

Wolfson:2003:BRB

[Wol03] Paul R. Wolfson. Book review: Bernhard Riemann, 1826–1866: Turning Points in the Conception of Mathematics: By Detlef Laugwitz. Tr. Abe Shenitzer with the editorial assistance of the author, Hardy Grant, and Sarah Shenitzer. Boston (Birkhäuser). 1999. xvii + 357 pp. ISBN 0-8176-4040-1, 3-7643-4040-1. Historia Mathematica, 30(2):223–226, May 2003. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086002000113.

Wolfson:2008:GBO

[Wol08] Paul R. Wolfson. George Boole and the origins of invariant theory. Historia Mathematica, 35(1):37–46, February 2008. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086007000389.

Wong:2022:SSS

[Won22] Peter Wong. On a sangaku of Sugino’o Shrine (Yamagata) and Yamaguchi Kanzan’s second trip. Historia Mathematica, 61(??):11–17, November 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086022000106.

Wuensch:2005:ZWK

[Wue05] Daniela Wuensch. “Zwei wirkliche Kerle”: neues zur Entdeckung der Gravitationsgleichungen der allgemeinen Relativitätstheorie durch Albert Einstein und David Hilbert. (German) [“Two real guys”: the discovery by Albert Einstein and David Hilbert of the gravitational equations of General Relativity]. Termessos, Göttingen, Germany, 2005. ISBN 3-938016-04-3. 125 pp. LCCN QC173.6 .W84 2005.

Wustner:2002:HRS

[Wüs02] Michael Wüstner. Historical remarks on the surjectivity of the exponential function of Lie groups. Historia Mathematica, 29(3):266–272, August 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086002923578.

Wallis:1980:FP

[WW80] Ruth Wallis and Peter Wallis. Female philomaths. Historia Mathematica, 7(1):57–64, February 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086080900646.

Xu:2005:FCT

[Xu05] Yibao Xu. The first Chinese translation of the last nine books of Euclid’s Elements and its source. Historia Mathematica, 32(1):4–32, February 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086003001095.

Xu:2007:L

[Xu07] Yibao Xu. Li Di (1927–2006). Historia Mathematica, 34(3):266–270, August 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086007000304.

Xu:2010:CM

[XX10] Yibao Xu and Fengxian Xu. Chen Meidong: (1942–2008). Historia Mathematica, 37(4):565–567, November 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086009000731.

Yadegari:1980:BTW

[Yad80] Mohammad Yadegari. The binomial theorem: a widespread concept in medieval Islamic mathematics. Historia Mathematica, 7 (4):401–406, November 1980. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/031508608090004X.

Yano:1983:BRG

[Yan83] Michio Yano. Book review: Geometry in Ancient & Medieval India: By T. A. Sarasvati Amma. Delhi/Varanasi/Patna (Motilal Banarsidass). 1979. xi + 280 pp. Rs. 60. Historia Mathematica, 10 (4):467–470, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086083900149.

Ycart:2016:JBD

[Yca16] Bernard Ycart. Jakob Bielfeld (1717–1770) and the diffusion of statistical concepts in eighteenth century Europe. Historia Mathematica, 43(1):26–48, February 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086015000440.

Youschkevitch:1983:FMS

[YDD83] A. P. Youschkevitch, S. S. Demidov, and Pierre Dugac. F. A. Medvedev et son apport à l’histoire de la théorie des fonctions. (French) [F. A. Medvedev and his contribution to the history of the theory of functions]. Historia Mathematica, 10(4):397–398, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900022.

Ying:2011:KSN

[Yin11] Jia-Ming Ying. The Kujang sulhae: Nam Pyǒng-Gil’s reinterpretation of the mathematical methods of the book review: Jiuzhang suanshu. Historia Mathematica, 38(1):1–27, February 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086010000273.

Yoke:1975:BRM

[Yok75] Ho Peng Yoke. Book review: Mathematics education in China: Its growth and development: By Frank Swetz. Cambridge, Mass. (MIT Press), 1974. 364 p. Appendices, Index. $15.00. Historia Mathematica, 2(3):365–366, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086075900920.

Yong:1974:YHC

[Yon74] Lam Lay Yong. Yang Hui’s commentary on the Ying Nu chapter of the Chiu Chang Suan Shu. Historia Mathematica, 1(1):47–64, February 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086074901670.

Yong:1975:BRC

[Yon75] Lam Lay Yong. Book review: Chinese mathematics in the Thirteenth Century: the Shu-shu chiu-chang of Ch’in Chiu-Shao: by Ulrich Libbrecht. Cambridge, Massachusetts (MIT Press), 1973. 586 p. $25. Historia Mathematica, 2(3):353–355, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086075900865.

Yoshida:1975:QGM

[Yos75] Katsuhiko Yoshida. Quantitative growth of modern mathematics in Japan prior to 1930. Historia Mathematica, 2(4):472, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508607590107X.

Yoshida:1976:QGM

[Yos76] Katsuhiko Yoshida. Quantitative growth of modern mathematics in Japan prior to 1930. Historia Mathematica, 3(1):51–54, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086076900069.

Youschkevitch:1983:KHP

[You83] A. P. Youschkevitch. A. N. Kolmogorov: Historian and philosopher of mathematics on the occasion of his 80th birthday. Historia Mathematica, 10(4):383–395, November 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086083900010.

Youschkevitch:1986:CMS

[You86] A. P. Youschkevitch. Chinese mathematics: Some bibliographic comments. Historia Mathematica, 13(1):36–38, February 1986. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086086902247.

Yshkevick:1990:BRI

[Ysh90] Adolf P. Yshkevick. Book review: II Circolo matematico di Palermo: By Aldo Brigaglia and Guido Masotto. Bari (Edizioni Dedalo). 1982. 444 pp. Lire H. 10,000. Historia Mathematica, 17(3):284–287, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086090900178.

Yussupova:1995:ZMA

[Yus95] Gulnava Yussupova. Zwei mittelalterliche arabische Ausgaben der “Sphaerica” des Menelaos von Alexandria. (German) [Two medieval Arabic editions of the “Sphaerica” of Menelaus of Alexandria]. Historia Mathematica, 22(1):64–66, February 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086085710051.

Yuste:2010:BRT

[Yus10] Piedad Yuste. Book review: Tablettes Mathématiques de Nippur. By Christine Proust. Istanbul (Institut Francais d’Etudes Anatoliennes Georges Dumézil). Paris (De Boccard Ed). 2007. ISBN 978-2-906053-92-2. 356, xlix pp. No price given. Historia Mathematica, 37 (4):708–710, November 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086009000615.

Zaitsev:1995:FAM

[Zai95] Evgeny A. Zaitsev. Fedor Andreevich Medvedev (1923–1993). Historia Mathematica, 22(1):88–92, February 1995. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086085710117.

Zaslavsky:1979:TEM

[Zas79] Claudia Zaslavsky. Teacher education: Mathematics in other cultures. Historia Mathematica, 6(2):189–194, May 1979. CODEN HI-MADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/0315086079900776.

Zaslavsky:1983:BRN

[Zas83a] Claudia Zaslavsky. Book review: A Negro history compendium: By Ethel M. Turner. Cheyney, PA (Cheyney State College), 1971. 74 pp. Historia Mathematica, 10(1):105–115, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086083900502.

Zaslavsky:1983:BRB

[Zas83b] Claudia Zaslavsky. Book review: Black mathematicians and their works: Edited by Virginia K. Newell, Joella H. Gipson, L. Waldo Rich, and Beauregard Stubblefield. Ardmore, PA (Dorrance & Company), 1980. 327 pp. $18.00 (cloth), $12.50 (paper). Historia Mathematica, 10(1):105–115, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086083900496.

Zhaohua:1988:SNC

[ZD88] Li Zhaohua and Liu Dun. The second national Chinese conference for history of mathematics. Historia Mathematica, 15(3):278–280, August 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086088900705.

Zepeda:2012:BRD

[Zep12] Henry Zepeda. Book review: Defending Hypatia: Ramus, Savile, and the Renaissance Rediscovery of Mathematical History. By Robert Goulding. Archimedes: New Studies in the History and Philosophy of Science and Technology vol. 25. Dordrecht, Heidelberg, London, New York (Springer). 2010. ISBN 978-90-481-3541-7. xx + 201 pp. Historia Mathematica, 39(1):109–111, February 2012. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086011000504.

Zervos:1975:BRE

[Zer75] S. P. Zervos. Book review: English–Greek mathematical dictionary: By C. P. Tzelekis. Athens, 1973. 235 p. Historia Mathematica, 2(3): 355, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086075900877.

Zhmud:1989:PM

[Zhm89] Leonid Zhmud. Pythagoras as a mathematician. Historia Mathematica, 16(3):249–268, August 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/0315086089900207.

Zhou:2021:TPS

[Zho21] Xulin Zhou. Two problems in the [Chinese characters] Suanshu shu (Book of Mathematics): Geometric relations between circles and squares and methods for determining their mutual relations. Historia Mathematica, 57(??):1–23, November 2021. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086021000410.

Zhu:2016:DCC

[Zhu16] Yiwen Zhu. Different cultures of computation in seventh century China from the viewpoint of square root extraction. Historia Mathematica, 43(1):3–25, February 2016. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086015000385.

Zhu:2020:HDW

[Zhu20] Yiwen Zhu. How do we understand mathematical practices in nonmathematical fields? Reflections inspired by cases from 12^{th} and 13^{th} century China. Historia Mathematica, 52(??):1–25, August 2020. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086020300252.

Zitarelli:1989:Aa

[Zit89a] David E. Zitarelli. Abstracts. Historia Mathematica, 16(1):98–103, February 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086089901079.

Zitarelli:1989:Ab

[Zit89b] David E. Zitarelli. Abstracts. Historia Mathematica, 16(4):393–405, November 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086089900918.

Zitaralli:1990:A

[Zit90a] David E. Zitaralli. Abstracts. Historia Mathematica, 17(4):394–414, November 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609090042C.

Zitarelli:1990:Aa

[Zit90b] David E. Zitarelli. Abstracts. Historia Mathematica, 17(1):81–96, February 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609090089V.

Zitarelli:1990:Ab

[Zit90c] David E. Zitarelli. Abstracts. Historia Mathematica, 17(2):187– 197, May 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609090070T.

Zitarelli:1990:Ac

[Zit90d] David E. Zitarelli. Abstracts. Historia Mathematica, 17(3):295–309, August 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086090900226.

Zitarelli:1991:A

[Zit91] David E. Zitarelli. Abstracts. Historia Mathematica, 18(4):393–402, November 1991. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609190385B.

Zitarelli:1992:Aa

[Zit92a] David E. Zitarelli. Abstracts. Historia Mathematica, 19(1):109–121, February 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/031508609290071I.

Zitarelli:1992:Ab

[Zit92b] David E. Zitarelli. Abstracts. Historia Mathematica, 19(2):230, May 1992. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508609290092P.

Zitarelli:1996:Aa

[Zit96a] David E. Zitarelli. Abstracts. Historia Mathematica, 23(1):96–113, February 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086096900117.

Zitarelli:1996:Ab

[Zit96b] David E. Zitarelli. Abstracts. Historia Mathematica, 23(3):333–348, August 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086096900348.

Zitarelli:1996:Ac

[Zit96c] David E. Zitarelli. Abstracts. Historia Mathematica, 23(4):447–464, November 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086096900440.

Zitarelli:1997:Aa

[Zit97a] David E. Zitarelli. Abstracts. Historia Mathematica, 24(1):102–120, February 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086097921418.

Zitarelli:1997:Ab

[Zit97b] David E. Zitarelli. Abstracts. Historia Mathematica, 24(2):214– 237, May 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S031508609792156X.

Zitarelli:1997:Ac

[Zit97c] David E. Zitarelli. Abstracts. Historia Mathematica, 24(3):343–355, August 1997. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086097921650.

Zitarelli:1999:Aa

[Zit99a] David E. Zitarelli. Abstracts. Historia Mathematica, 26(1):83–103, February 1, 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086099922280.

Zitarelli:1999:Ab

[Zit99b] David E. Zitarelli. Abstracts. Historia Mathematica, 26(2):181– 198, May 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086099922401.

Zitarelli:1999:Ac

[Zit99c] David E. Zitarelli. Abstracts. Historia Mathematica, 26(3):295–308, August 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086099922486.

Zitarelli:1999:Ad

[Zit99d] David E. Zitarelli. Abstracts. Historia Mathematica, 26(4):389–409, November 1999. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086099922590.

Zitarelli:2007:BRP

[Zit07a] David E. Zitarelli. Book review: Pursuit of Genius: Flexner, Einstein, and the Early Faculty at the Institute for Advanced Study, Steve Batterson. A. K. Peters, Wellesley, MA (2006), ISBN 978-1-56881-259-5. Historia Mathematica, 34(4):443–446, November 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086007000031.

Zitarelli:2007:SCS

[Zit07b] David E. Zitarelli. Straddling centuries: the struggles of a mathematician and his university to enter the ranks of research mathematics, 1870–1950. Historia Mathematica, 34(3):271–288, August 2007. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0315086006001273.

Zlatopolski:2023:PAV

[Zla23] Dmitry Zlatopolski. “Perfect Arithmetic” by Vaclav Josef Pelikan. Historia Mathematica, 62(??):40–50, February 2023. CODEN HI-MADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/S0315086022000659.

Zormbala:1996:GDP

[Zor96] Konstantina Zormbala. Gauss and the definition of the plane concept in Euclidean elementary geometry. Historia Mathematica, 23 (4):418–436, November 1996. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0315086096900403.

Zund:1983:SCR

[Zun83] J. D. Zund. Some comments on Riemann’s contributions to differential geometry. Historia Mathematica, 10(1):84–89, February 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/031508608390037X.

Zund:2002:GDB

[Zun02] Joseph D. Zund. George David Birkhoff and John von Neumann: a question of priority and the ergodic theorems, 1931–1932. Historia Mathematica, 29(2):138–156, May 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0315086001923389.

Zepeda:2022:MRP

[ZvDF22] Henry Zepeda, Benno van Dalen, and Menso Folkerts. In memoriam: Richard P. Lorch (1942–2021). Historia Mathematica, 58(??):7–16, February 2022. CODEN HIMADS. ISSN 0315-0860 (print), 1090-

249X (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086022000039.

Zygmund:1975:RFS

[Zyg75] Antoni Zygmund. The role of Fourier series in the development of analysis. Historia Mathematica, 2(4):591–594, November 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0315086075901263.